A Complete Bibliography of Publications in Scientometrics: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 October 2025 Version 1.63

Title word cross-reference

3 [BWLX20, LP20]. ^{n} [Tom24]. _{2} [dGIL21]. g [Kos25]. g_{m} [SDD20]. H [CLC^{+}24, TD24c, WAM^{+}25, ALM^{+}23, Bi23a, Bi23b, Bi23c, DLK20, ER21, ELR24, Fas20a, Fas20b, HHKZ20, HBSW24, HWNL20, KDS21, KS22a, LGLM20, MMN20, MNM^{+}21, MHK22, Naz24b, PMHC21, Pra21a, Pra21c, SdL21, SG21, Sin22, TK21, Vin23, dS21c]. h(fa) [SdL21]. h_{u} [LSBC21]. k [VMF22]. p [Pra21c]. π [Vin23]. Ψ [LP23]. Q [For23]. t [Sin22]. xd [NLS24]. x [LNS21]. z [Pra21c].

-an [GCHT20]. -core [CLC^{+}24]. -dimensional [LP20]. -Education [RSF20]. -index [ALM^{+}23, Bi23a, Bi23b, Bi23c, DLK20, ER21, ELR24, Fas20b, HHKZ20, HBSW24, HWNL20, KDS21, KS22a, Kos25, LNS21, LSBC21, MMN20, MNM^{+}21, MHK22, NLS24, Naz24b, PMHC21, Pra21a, Pra21c, SDD20, SdL21, SG21, Sin22, TK21, Vin23, dS21c, Fas20a, LP23, TD24c, WAM^{+}25]. -means [VMF22]. -type [DLK20].

000-index [dS21c].

10 [AHH23]. 10-year [MKWB^{+}23, SJW21]. 100 [MF20b]. 17th [DMG20]. 18th [GDR22]. 19 [AD21, ADD22, ADM22, Ali22, AM22, ARR21, BD21, BdMPH21, BJS22, BTP24, BFG^{+}21, BW22, CFW21, CFA^{+}23, CP21, CCC22, Coc21, CA22, DDVV20, EXT^{+}21, Fas21c, GCGB^{+}24, DR20, HB20, HL21b, HKV20, KHT22, KTB22, KCC21, LCC23, LYZ22, LYZ23, LDAAAB21, MA21, MAS^{+}21, NRGvSTS23, Pal21, RBB22, SSCK21, SSdCRD22, SMG^{+}24, SB21, SP20, SCP22, SFC^{+}23, TZY^{+}23, TTB22, WR21, WL21a, WZM^{+}24, YNO22, YNL25, YWB21, ZCW21, ZLM^{+}23, Zha23b, dSTE21]. 19/SARS [HB21b, PHBG21]. 1948-2020 [Bor21]. 1965 [SdSFB22]. 1985 [HMW24]. 1998-2017 [Gon23b]. 19th [BMW24].

2 [HB21b, PHBG21, Pur24]. 2-stage [ZBKW24]. 2.0 [GC20]. 2015 [CMZ^{+}20a]. 2019 [Ano20a, HB20]. 2019-nCoV [HB20]. 2020 [PSW24]. 2021 [Ano21a, vR21a]. 2022 [Zha23a]. 2023 [Ano23, BMW24, Sug23]. 2SCE [ZBKW24]. 2SCE-4SL [ZBKW24].

3SA [MJRR25].

4.0 [Dha22, GSG22b]. 40-year [CxGdH^{+}24, TKZ^{+}23]. 40years [Nuz20, SYHW20]. 4SL [ZBKW24].

5 [MF20b]. 5-100 [MF20b]. 5-year [EGPGGA^{+}23]. 5.0 [GSG22b]. 5G [RBBP24, vdP24].

9 [SKS^{+}23].

= [SLK^{+}23].

ABCal [Le 24]. ability [BMS21a, PSW24]. ablation [MSV23]. abnormal [Liu23]. above [OSS^{+}23]. above-average [OSS^{+}23]. Abramo [Zha23a]. abstract [Bös21, Cop20a]. abstracting [ASSB20, Wij21]. abstractive [SMA21]. abstracts [AGDE25, BS23, BRG^{+}23, Cop20a, Cop20b, EBW23, GKDB25, GSH24, GCT^{+}20, HBC25, JADG20, JDG21, Li22, WY23, Xie20, dW24]. Abuse [dS21a]. academia [CHT24, FBM21, HFAR23, JT21, MMF23, MOMGNLG24, PL22b, Pur25, SR24, UBROGS23, WZG21, WGA22, YWW22, Zan23, ZO20, dSAKT20]. Academic [BVvEW23, Etz25, FM21, GdM20, KM20, KR22a, LBS23, LR21, MMTOMLC21a, OM21, TYS22, VPR22, WP22, XWGD22, YCL25, AJBBA23, AA25, ALM^{+}23, BBP23, BZZ25, BKT22, Bir21, BM21, CW24, CTH22, CW25, CLP21, CDZZ20, CKO21, Cho21, Cho24, CF23, DS25, DTHM21, DFRS20, DWG23, DH23, EPP23, FH21, FH23, FFJ21, Fur21, GDRPR23, GDM22, GM22, dMGC23, GTB23, Gub24, HBW^{+}21, HYY25, HKV20, HLS23, HY22, JXY^{+}23a, JXY^{+}23b, KFM20, KLJS25, Kos23, KDN22, Kou24, KBDS23, LSA22, LKPS25, Li22, LH24, LZC21, Lin20, LCCH24, LWX21, LH22a, LWW24, LZ25b, LRW25, LC20, LUH21a, LUH21b, MZD22, MMF20, Mal22a, MPS22b, MS21d, MMG23, MR20, MNP21, MN24, MFF^{+}20, NBBL24, NC25, OM20, PWZ25, Plo25b, Poh21, Por25, QD21, RM24, SHL22a, SBF23, SGM22, TKS^{+}23, TDWW24, The20b]. academic [The23, TP20, dGdO^{+}23, WGW21, WZL22c, WJZ24, WLHY24, WZZ25, WPSK25, WKC22, WC20, WL22, WVH21, XZ23, XZD19, XLA23, XDLZ23, YLCY20, YWZ^{+}23, YF24, YLC^{+}20, YXQG21, YLf25, YWW23, YMW^{+}20, mYY22, ZWSC20, ZW21a, ZWC22, ZZZ22, ZC24, ZW24a, ZYZZ25, ZZ25, ZJY22, ZYL25, ZL20b, Tei21, MMTOMLC21b, Mas20]. academic-corporate [Poh21]. academic-social [GDRPR23]. academicians [AYRG23, PLW23]. academics [AJB^{+}21, ADH23, GCGB^{+}24, KSAT24, SMT^{+}23, KN22]. Academy [HB23, PA22]. Accelerated [TLL24]. accept [dSV23]. acceptance [ABKGBR25, Seb23a, TK21, TLL24, WZZ^{+}21]. accepted [MWH20, Ste22]. Access [Agu20, AÁSTMR24, ADS20, AGL22, Asa20a, Asa20b, Asa21, Asa24, BBP21, BPLIdMA^{+}20, BMBA20, BGAMS23, Bös21, BL20b, BGWZ20, Cho20, DMR20, DK21a, DGSVDG20, DMR22, Gra20, HAR24, Hen19, Hen20, HJT21, HHB^{+}20, HNM^{+}24, KL22, Lin21b, Mad20, MLS21, MS23, MBD25a, MBD25b, MMP21, Mor20, NRGvSTS23, Naz24a, NM25, OMFJSC24, PROG21, RPL21, STAÁMR24, SL24, SPS20, SSK22, TPVVPA21, Tay20, Tay23, TD24a, UNK23, VRVH21, WC20, Wei20, WPS22, WLMZ24, YSWC24, ZWW21, ZWHS22, ZM24, Zhu25, ZHH23a, Bou20, MMGC^{+}24]. accessibility [SGC21a, SGC21b, SWL^{+}23, WCL25]. according [ES22]. account [Wal22]. accountability [PROGDG24]. Accounting [DDS22, ABKGBR25, BdCM21, EPP23, HKW24, Lyu21, SCW^{+}23, UBROGS23, UNK23, OFA23]. accounts [BP22, YNO22]. accumulated [LTL24]. accumulation [FC20a, KG20]. Accuracy [Liu20, SdLV21a, CSR23, TP24, Tei24b]. accurate [YCXY20, YMX21, YYC22]. accurately [AJLA21]. achievements [KYKC21]. acknowledged [Kou24]. acknowledgee [XZ23]. Acknowledgement [TXL21, LHZ^{+}23, SM23]. acknowledgements [SM24a]. Acknowledgment [ZX25, BP22, DH23, McC24]. Acknowledgments [Naz25, EO24a, JXZW23]. ACM [FCC22, Tom21]. acquisition [CxGdH^{+}24]. acronyms [BD21, CCC22, KCC21]. across [ADD20, ADM22, AL24, Bir21, CKO21, FCT^{+}20, GDdZ25, GLPN25, GCHT20, HB20, HBO^{+}20, ICB24, LKPS25, LSBC21, LC20, MS22c, MBR^{+}25, PMC21, RNAH21, SLK^{+}23, SHF24, Sch24b, TMW23, WL22, WKC24, WCH^{+}20, YWW23, mYY22, ZSSH22, ZW24a]. act [Cop20d]. Action [DK21a]. active [CPLS20, GC21, ZSS22, ZTY^{+}24]. activities [ANR21, Cho24, QZS^{+}24, YLX25, ZL23c]. activity [BGSRF20, GKK23, GCHT20, HBW^{+}21, HGDRM20, HHB^{+}20, HKNF23, NCD22, RNB22, San21a, San21b, YD21, ZCSK24]. activity-based [BGSRF20]. actors [JDG^{+}20, LOGS21, OVS25]. actually [LYZ^{+}20, SvWC21]. actuarial [YKTM20]. acupuncture [The21]. adaptive [JJY21, LHX^{+}24, Wag25, WWS23a, WLL25]. addiction [SCFSAB^{+}25]. Additional [Kos23]. additions [HKW24]. Additive [Pra21a]. address [SdLV21a, ZZL^{+}20]. addresses [MB22a]. addressing [NWZ23]. Adjectives [WL22]. adjusted [SBP20]. adjusting [XDW22]. administration [MMG21, MMG23, NBBL24]. adopting [LLC22]. Adoption [Mas20, FH23, HJT21, KLJS25, LHCH20, LCCH24, Por25, YLCY24]. adoptions [OVS25]. ADS [MS21e]. advance [Cop20d]. advanced [NHAB21, YKTM20, ZM23b]. advancement [HDY^{+}23]. advances [BMBA20]. advantage [AAD24a, BBP21, BPLIdMA^{+}20, CTH25, CTH22, Cop20a, LZY22, NM25, RPAVFV21, Tay20, WWL24b, ZWW21]. advantages [GDM22, WLLW25a]. adverbs [WL22]. adversarial [CNT25]. advisor [GSOW20]. advisor-match [GSOW20]. advisors [WGW21]. aerospace [KY20]. affect [Ali22, DM25, MTD^{+}23, NBBL24, PXY24]. affecting [jDXjS21, HTX23, LLP21, OŠV^{+}24, SMBD^{+}24, TTDK22]. affects [ACT23, LCL23, TD24c, ZW21b]. affiliated [RSV^{+}20, ZAZS22]. Affiliation [YS24, DS21d, QD21, SK22]. affiliations [GFVK21, Mou20, MJRR25, UA20, dMGDD24]. affinity [CRBRGS21, TKS^{+}23]. Africa [BP20, CMV23, Dha22, EOL23, EO24a, EO24b, HBW^{+}21, NB20, NB22, Ony20, Vie22, Vie24]. African [VC22]. after [Ali25, BZZ25, FG20, HZC21, MSJ23, MF20b, Reh20, SYHW20]. against [BT20, DZ23, MRM20]. age [DGGD22, Dri24, GCHT20, KP23a, KR22a, Tay23, YXDL24, ZC21]. age- [ZC21]. age-related [GCHT20]. agenda [ADDCMG^{+}25, BS20, JVY22, NS23]. agendas [SHL22c, SHL22a, SHL22b]. agents [WZ21]. Ages [Sta21]. aggregation [ZYWA25]. aggregators [KBS21]. aging [GMB22, HMZ21]. agree [Sma24]. agreement [LGY23]. agreements [Sch24b]. agriculture [HGM21, NB20, RBB^{+}24]. Agronomy [dGdO^{+}23]. AGSTA [WWS23a]. AGSTA-NET [WWS23a]. AHP [Can20]. AI [Ano22a, FT24, GWLY24, GLPN25, JCZ22, Kim24, KLJS25, KSS21, LHC25, LLL25, RKCA23, TLM22, WXZ24, YSW^{+}20, ZZMS22]. AI-enabled [LLL25]. aims [SM22a, ZSG21]. Air [Plo25a]. airline [MH24a]. AJG [DMA^{+}20]. al [Bjø20a, Bjø20c, HP21, MHD21b, Pra22, Rus23, SdL23, WFW24, Yur23, Zha23a, KUS22]. Al-Farabi [KUS22]. al. [Cop20a]. alert [FFJ21]. algorithm [Kal20, MJRR25, SWT20, VMF22]. Algorithmic [Don22]. algorithms [Kua20, LL24b, LGXH24, ZWC22]. alignment [CXW^{+}24]. all-author [YLJ21]. All-pervading [Tut23]. alleged [Cop20a]. alliance [ZWD20]. allied [VPC^{+}23]. allocating [AA22]. allocation [AJLA21, BPL25, DLC21, GJ24, GAB21, GM22, Han20, JC22, Sun23, WMK21, ZL23c, AKD23]. alone [KR22b]. along [Fan20, GKK23, LTL23]. Alphabetic [FC20b]. Alphabetical [DK25, Öz24]. Alphabetized [WB22]. also [DV22, PJ24]. alt [HIA^{+}20]. alt-index [HIA^{+}20]. alter [GSG22a]. Alternative [The21, Ans23a, BORROPGV20, Cop20d, GUG^{+}22, KR21]. alters [Tut23]. Altmetric [GSB24, AMZ24, Cop20c, DMA^{+}20, FCT^{+}20, FC20a, HP20, HHB^{+}20, KBS21, LH22a, LC20, dMMdCM22, PSKS22, SdSCdS25, Tay20, Tay23, TSAMT21, VAMBTS25, YXQG21, AMTS24, BS23]. Altmetric.com [FNM^{+}22, FC20a, KBS21]. Altmetrics [Agu20, Cho21, HLZ23, Ara20, AMTSRG21, BFG^{+}21, BST22, CB25, GX22, KTK21, LBP22, MS22b, Mal22a, Mal22b, MH24b, MJYS23, NLS20, Ort20, SS24, TSDAMRG24, YCXY20, YMX21, ZWZ^{+}19, ZWX22, Lec21, Tei24a, YWHS23]. Altmetrics- [Lec21]. Altmetrics-based [HLZ23]. Alzheimer [MPS22a]. ambiguity [SdLV21a]. ambiguous [CBA24]. ambitions [JI22]. America [AdAFR24, Ano21c, BH24, Cor22, GB21, LMLMPA22, MKWB^{+}23, SdSCdS25, SM20, VCD^{+}23, VdBC20, ZYL20]. American [BGP20, GMCRV21, LdGRV^{+}22, PRGTOP^{+}24, RFG^{+}24, RP20, RPAVFV21, SSC21, SCU21, YGO22, dCPT22]. Among [SMS23, AHH21, Ali21a, Ali21b, AGL22, BP23, Ben20, Ben24, BL20b, EPP23, FH23, GB24, GDdZ22, JVY22, KN22, LWL24, LM20, MBD25a, MBD25b, OdSO20, Por25, QZS^{+}24, SSZ22, Sha24b, SDC22, SMZY23, WCY22, ZSS22, ZMK22, dW24]. amplifying [Tei21]. anaesthesia [GWvZ21]. analogy [ZY20b]. analyse [RPL21]. analyses [dOSAL21, Ans23a, Cha23a, CR24, Cot22, EO24a, Le 24, MMG21, SVC20, SS24, The21, WCH^{+}20]. Analysing [LFV21, ZWC22, BVS25, DCdA25, PROGDG24]. Analysis [BKS20, CH20, Fan21, FZL^{+}20a, GS24, Jab25, KLM24, LGD21, Lec21, LL25, NH24, PRS^{+}21, PSS23, Pra24, RSF20, SMZY23, dBGP21, AM21a, AJBBA23, AHH21, Ali25, AKD23, AYS24, AAHY20, ACD^{+}24, ACDS22, AMS21, ARR21, ARR23, BR25, BP20, Bar25, BVvEW25, BBP21, Bat20, Ben24, BSS24, BJ21, Bha21, BdPFdO25, Bir21, BZG22, BDPP23, BM21, Bor21, Bor22, BWH20a, BWH20b, BL20a, Bor24, BI24, CTH25, Can20, Cao20, CCH^{+}24, CGPR24, CMV23, CW25, CMT24, CC21, CxGdH^{+}24, CZQC24, CCH25, CWL^{+}20, CZZN25, Cho21, CCMM24, CKMY20, CDG20, CFC22, Cor22, CBA24, CHJ^{+}25, CSR23, CD22, Day22, DMR20, DTHM21, Dej25, DK21b, DX20, DKWC20, DH23, Don21a, DWM21, DG21, DMGBG20, EGR22, Ela25, EE24, ELR24, FMF20, FCT^{+}20, FCW22, FVLA20, FZhCC21, FFH^{+}24, FRRBDD23, Fil21]. analysis [FRS23, Fog23, FWH^{+}23, GRTPAJ21, GHW22, GVS20, GSP21, Gon23a, GdM20, GTH24, GWvZ21, GSB24, GMKG20, GCHT20, Han20, HC22, HHKZ20, HP21, HVGP21, HBO^{+}20, HG20, HFAR23, HBC25, HKV20, HBN^{+}21, HY20, HYZ23, HTZ24, HKN23, HKNF23, HZ21, HCZ^{+}22, HL24, IQ21, IPP22, IHB21, JLL21, JKK24, JHVC21, KSS20a, KH21, KDIR23, KBSS23, KK23, KP23b, KSAT24, KPS21, Kua20, Kua23, KCS25, KBDS23, LCB21, LCC^{+}20, LMC22, LR25, LS20a, LHXS20, LHC22, LHC25, LSFLZLLV20, LGY23, LTL24, LP20, LLH20, LL21, LWX21, LH22a, LH22b, Liu23, LZ25b, LXL25, LdS25, LMLR24, MPS22a, MTD^{+}23, ML24, MB20b, MB21, MB22b, MCSOAA21, dMMdCM22, MML^{+}20, MPS22b, MRF21, MSL22, MBD25a, MBD25b, MMG23, MGCR24, MMB23, MQS^{+}20, MMP21, MOMGNLG24, MDCBdGVMD24, Mou21b, Mou21c, MNK21, Mu21, NC20]. analysis [Nis23a, Nis23b, NK24, OJKY23, Ony20, OMAAORCL23, Ort20, OA24, PL22a, PMB^{+}23, PRGTOP^{+}24, PLYZ24, PYS22, PA21, PVB^{+}24, PLH20, PL22b, PT20, PT24, PL23, PKCOG21, PM20, PMGÁC23, Pup21, QZF22, QQL21, RG20, ROB21, RNAH21, RKMG20, RTT23, RSA20, RP24b, RLZ^{+}24, RBC21, RBB22, RPL21, RAG25, SSCK21, SJBBR^{+}23, SBP20, SSS^{+}24, SMJ21, SK21, SXLC21, SDW^{+}20, SSM21, Sin22, SM23, SSC21, SAK22, SJW21, SLL23b, SSK22, SLTS25, SCU21, SC20b, SM24b, SSBC22, TZ24, TLL24, TFWC24, Tom23, TSAMT21, Urb25, UA20, UBROGS23, VJCV22, VCD^{+}23, VPDAG22, VEGSMMC22, VPC^{+}23, Vie22, VA22, Wad20, WGC20, WYT21, WWH21, WGW21, WXZ22, WJZ24, WL24, WLHY24, WQ22, WKC22, WLBC20, WG20b, WZZZ20, WGZ22, WLMZ24, Xie20, XHA^{+}20, YRP21, YXQG21, YWX23, YH24, YWN25, YLJ21, YL24]. analysis [YYZ25, YJJ25, YP20, YS20, YMX21, YY21, YY22b, YYC22, YMLL23, YS23, YLX25, Yur22, ZIS21, ZWZ^{+}19, ZZS^{+}20, ZW21a, ZWW21, Zha21b, ZWX22, ZY22, ZGM^{+}23, ZZ23, ZX25, ZS20b, ZPH21, ZO20, ZCL20, ZZ20b, ZQS22, ZMK22, ZYL20, ZXL20, ZHH23a, dSN23, dFBGCGASG24, FZL^{+}20b]. analytical [SR23]. analytics [ADS20, ADU22, HCN21, HCL22, RSPC20]. analyze [WPBH24]. analyzed [PS20]. Analyzing [BSA23, CB25, FT24, Kim24, SD20, SA21, TQA21, YWB21, YWZ^{+}24, ZS24, HSE23, HAT23, LYS22, MGCR24, NC25, SdSFB22, TTB20, WH23, ZAZS22]. anchor [NKY^{+}24]. Angola [AGL22]. ANN [QJI^{+}22]. anniversary [Kos22]. annotated [SF20]. Annotation [NA24, ZC24]. annual [ZYL20]. Annus [YG20]. anomalies [Naz24b]. anomalous [WG22]. anonymous [SSSd22, LZ24a]. ANP [GB24]. antecedents [BMM22]. anthropogenic [GGSA20]. anti [GSG22a, KSP24, vSDC20]. anti-doping [KSP24]. anti-inflammatory [GSG22a]. anti-vaccination [vSDC20]. Anticipating [LHK21]. Anticipation [AMS21]. antiquity [dDM25]. any [CKT21, DAGG24, VRVH21]. anyone [dSV23]. apatride [SK22]. API [SVY22]. APIs [GdMDD24]. applicability [GDG^{+}20, NK24]. applicant [QPH^{+}25]. Application [Mar20, VMF22, ACV23, BBP21, BO20, Cha23a, CLC^{+}24, CFI22, CBA24, EGR22, FESRMB21, FSEJF^{+}20, GS22a, HS22a, QLW25, RLZ^{+}24, VA22, WSZ21, WMCL22, WZL22c, XLL20, YN23, YKS22]. applications [HJLZ23, Iva25, OPS25, Tom23, WFW24, ZWSC20]. Applied [dRRJ20, BJS22, CRBRGS21, Coc20, FYY21, HGM21, HKN23, HKNF23, Li22, LL24b, PJ24, SHL22a, SWT20]. Applying [NHAB21, Ove23, BERD21, LSY21, SH20, Fil21]. appointment [Ger24]. appraisals [FLC22]. Approach [BORROPGV20, GVK24, AdAFR24, AlYT21, Agu20, APR19, ABAAN^{+}23, AED20, ADS20, ADU22, ADDCMG^{+}25, AMTSRG21, Avd21, ARR23, AMIK23, BGB25, Bat20, Bir21, BMM22, BO20, CTH25, CCH25, CPL21, CLR25, CMZ20c, DDS22, DDL23, DRF21, GP21, GRBAM22, GB24, GDDG24, GGSA20, GDG^{+}20, Hal20b, HAT23, HWC22, HZCM23, Hüc23, IOWN20, JPS21, Jia21, KSS20a, KR21, KYL20, Kös20a, KG20, LGD21, LHK21, LLC22, LZF^{+}24, MBL21, MMFF25, MSI21, MGCR24, MBR^{+}25, Mut22, OMC21, PD20a, PD20b, PLH20, PM20, Reh20, RP25, RAS21, RPL21, SV20, SGA22, SB20, Sha24a, SC20a, SMZY23, TA22, TD24a, TMW23, VBF24, WYT21, WYL21, XlYqZ22, XLY21, YN23, YWY^{+}20, YJS21, YKS22, ZLL22]. approaches [FGB25, GSRG24, Jia25, LHC22, RSPC20, SA24, VA22]. appropriate [BB22]. approval [TM25]. Approximate [BO20]. aquatic [SdSFB22]. AR5 [HDY^{+}23]. AR6 [HDY^{+}23]. Arab [BL20b]. Arabia [AJBBA23, AAAA23]. Arabic [Ibr21]. Arbitrariness [BB20]. arc [Kua20]. archetypoid [WG20b]. architecture [DXL21, NS23]. archive [Aba21]. archived [Bou21]. area [AA25, LZH20, LBW^{+}20, LC20, MF20a, VC20]. areas [Azm22, BBP21, LHF23, MFM20, SSK21, WLBC20, ZW24a, vRW25]. arena [LDMVS21]. Argentine [DCdA25]. arguments [CCZL23]. Armenia [ADGS25, SMM^{+}20]. art [JT21]. artefact [HS22a]. Article [HB21a, MS22b, ZC21, ADD23, ASIN24, AÁSTMR24, AGDE25, AGB23, Asa20b, Asa23b, Asa24, BGWZ20, BI24, CPN21, CWZ22, CG22, CCC22, CLS^{+}20, Dia21, ES23, HGMÁW23, JJ23, KDZ21, KCC21, KCC22, KCC23, LCB21, LAT22, Li22, LCC23, LWG21, LMSNZG22, MLA^{+}23, Naz24a, Pea20, Pea21, SLR^{+}23, SJW22, TBA23, TWH^{+}24, Tom24, WSZ21, Wan24a, XL20, ZY20c, ZHH23b]. article-based [CG22]. articles [AMZ24, BBB^{+}24, uHBKK20, uHBKK21, BBP21, BKT20, Bir21, BS20, Bre21, CTH25, CWZ21, Cha21, Che18, Che20, CDZZ20, CNA22, CLC23, DD23, Ela21, Ela22, Emm19, Fas21b, GW23, GGS22, DR20, Hao24, HP21, HZ22, HHB^{+}20, HJCT23, JADG20, JDG21, LS20b, LBP22, LTL23, LHCH20, LLH22, LLSH23, Lin20, Lin21b, LHZ^{+}23, LGR^{+}21, MZD22, MTD^{+}23, MH24b, Mu21, NZL21, PMB^{+}23, PY24, PY25, PD20b, Pra23c, Pra24, Pra25b, Pup21, RBR22, SKS^{+}23, Sch24c, SMBD^{+}24, SP20, SLTS25, Ste23, SCP22, TZY^{+}23, The23, TKM^{+}23, TDM24, UA21, VCJ22, VP21a, WWZW20, WH24, WH25, WAT23, WZNL23, YWX23, YP24, YWW23, mYY22, Zbo25, ZWSC20, ZZZ22, ZM23a, ZC24, ZXL^{+}24, ZYZZ25, ZJ25, ZL20a, Zhu25]. articles-in-press [Lin21b]. Artificial [BCT22, AHH23, AJVACC22, CTLH24, CPTÁ22, GLPN25, HJLZ23, HSE23, LHC22, LSY21, LGXH24, MBL21, RSM24, WD21, ZWZ^{+}19, ZWX22]. artist [OSS^{+}23]. arts [CY25, LNH24]. ARWU [SD20, LMLR24]. arXiv [BST22, LYZ^{+}20]. ascendancy [BD21, CCC22, KCC21]. ASEAN [SSS^{+}24]. Asia [MKWB^{+}23, PDH21, Pur21]. Asian [Cho24, Ove23]. ASJC [ÁLGBdMA24]. aspects [BKS20, GD22]. assess [GDG^{+}20, RSV^{+}20]. assessed [BKT20]. Assessing [ADH23, BH24, JMP20, KL24, MDGR21, PD20a, RBB^{+}24, WSL23, XLL20, BAE24, Bi23a, Bi23b, Bi23c, BMA21, GG22, HBW^{+}21, LCCH24, MMFF25, Mal22a, NLS20, RBBP24, SMT^{+}23, TMNH^{+}24, WPS22, YWHS23]. Assessment [DD21a, KT21, Men25, SAA21, ADD24, ADU22, BTD21, CMMO20, Dag24, DV20, DRS20, EGR22, FFN25, FNK23, FNM^{+}22, GLPC20, GMKG20, HAR24, HBO^{+}20, KG22, LP23, LLZ21, Mal22b, MdMAGB^{+}21, OMBP24, OPMRP20, Pin23, Pra20d, SM22a, Sch20a, Sch22, Tei24a, UMA21, VFL^{+}21, VKY23, VPR22, WD21, XYY25, YS24]. assessments [BR25, BDPP23, BST22, JI22, San21b]. Assigning [Ela25]. assignment [GSRG24, Kal20, Kua20, MRBSFSJ24, ZS24]. assignments [MBH25]. associated [BS23, BRG^{+}23, GAPR20, YWX23, ZY20a]. associated-sleeping-beautie [ZY20a]. Association [Par24, AJB^{+}21, Kim24, SLR^{+}23, SD22, SLTS25, VWR25, WZY21, BP23]. associations [HZJ22, SS24]. assortativity [LLLL24]. Astro [Don21b]. astrobiology [ML24]. attacks [HNO^{+}21]. attempt [Yam20]. Attention [AMTS24, Ara20, BS23, BRG^{+}23, DK21a, DG24, GMPRSFV25, HWW^{+}24, JDG21, JVY22, KBS20, dMMdCM22, MN21b, RAF25b, Tay20, Tay23, WGC20, WWS23a, WLL25, WC20, ZA20, ZHH23b]. attention-based [ZA20]. Attitudes [Ali21a, RS22a, TK21, Ali21b]. attorney [TM25]. attorneys [KS22b]. attract [PSW24, QLW25]. attribute [LH24]. attribute-based [LH24]. attribution [AA22]. atypical [WCM20]. audience [DSL21, KZV20]. auditing [ABK20]. augmentated [SBKK21]. augmentation [ZDL^{+}20]. Australia [CLZ21, GLK^{+}23]. Australian [HB23, JA23, Jam25, OPS25]. Austria [RS22b]. Author [FCC22, GX22, SS23, Sun23, Wij21, ARDA20, ABKGBR25, BZX^{+}23, CKO21, CLC23, DvE20, DJM22, DLC21, DK25, Emm19, GRSFV20a, GRSFV22, GKDB25, GLPN25, HM23, HGDRM20, HGMÁW23, Jia21, JP22, KS22a, KOS21a, KCS25, Le 24, MS21e, MKWB^{+}23, MFM20, PA21, PMC21, Pra25b, Pur24, PFS^{+}20, SAA21, San20, SGA22, SdLV21a, SC20a, UMA21, VA22, WQ21, WQ22, WG22, YLJ21, Yen25, ZMA23, ZY22, ZS20b, Tei24b, YS24]. author-centric [WG22]. Author-level [GX22, SC20a]. author-specified [CLC23]. author-topic [PFS^{+}20]. authored [FMB22]. authoring [SSCK21]. authority [AM21b]. authors [AKW23, AM21b, Bi23a, Bi23b, Bi23c, BSA23, CGPR24, Cha21, CY25, Che24, DLK20, DGRÁTS23, DXL21, DD21b, Ela21, FC20b, GFVK21, GDG^{+}23, GBP21, Ioa23, JKK24, JXZW23, Jok20, KS22a, KK20a, KYKC21, LAT22, MB22a, MB22b, MSJ23, MMGL23, Naz24a, NF22, RBR22, San25, SMJ21, SSZ22, The23, UA21, dFVFA^{+}24, VP21a, ZF22, dS21a]. authors-addresses [MB22a]. Authorship [DBR21, GAB21, HGMÁW23, SR24, ALR24, BBP23, BMS21a, BP23, BOV21, BH24, CLZ21, DS25, Dri24, EGP22, GAPR20, GDdZ25, GdOCRM20, HC22, HX25, JKK24, JA23, Jam25, JYT21, Jok20, KNT20, Kos21b, KS24, LBS23, LR25, LH22b, LZ24b, Nuz20, Nuz21, SAA21, SHH23a, SHH23b, SFJO20, SDRS23, Sun23, TMZ24, TD24c, UA21, UA20, dGdO^{+}23, WKC24, WB22, ZS21, dOCP21]. authorships [CTH25, HFS22, dOPGdAPU22]. automate [WD21]. Automated [CXW^{+}24, ZC24, CNA22, DTHM21, KCC23, MMFF25, PW20, YNO22]. Automatic [ASSB20, CPLS20, CCSX23, GdMDD24, JADG20, HHKZ20, Kal20, LCW^{+}20, SM24a, YWY^{+}20, ZC24]. automatically [BT22]. automation [CD22]. automotive [LHXS20]. autumn [CFW21]. Availability [Bou21, Mal22a, WLHY24]. available [LML21, Wal22]. average [OSS^{+}23, Sin22]. aviation [Mey25b]. Awakening [TTB22]. Award [Sug23, vR21a, CJT22, CZZN25, JYT21]. awarding [vdBM24]. awards [HZC21, TdS20, Wan24b, WZZ24]. aware [DDS22, JHVC23, JJPC20, JMZ^{+}23, WLL25]. awareness [KKK^{+}25, ZSX24].

B [HCZEP20]. back [IHLP21, San20]. back-of-the-envelope [San20]. background [TTB20, VC20]. badges [ZHH23b]. balance [Mut22, RZS23a, RZS23b]. Balkan [ZP20]. Baltic [San21a]. banks [KBDS23]. banquet [ACT23]. Bar [BL20a, Hal20a, HP20, LB20, OM20]. Bar-Ilan [BL20a, Hal20a, HP20, LB20, OM20]. barely [FN23]. barrier [KY20]. barriers [AYRG23, CL21a, CdSK21]. base [HGB20, RSF20]. Based [ZC23, AAD24b, AHH23, AED20, ÁLGBdMA24, ADS20, AGDE25, AAB^{+}24, AU20, AA22, AMIK23, BBB^{+}24, uHBKK20, uHBKK21, BGB25, BB22, BVvEW23, BZX^{+}23, Bir21, BCF21, Bor21, Bor20b, BMS21b, BVS25, BO20, BCG22, BRJ21, BGSRF20, CTLH24, CXZ^{+}20, CZQC24, CG22, CLS21, CLZ21, CLC23, CLS^{+}20, CdSK21, DDL23, DDD^{+}21, DLC21, DCL23, DD23, DDF20, jDXjS21, Dri24, DG21, EGR22, EE24, FZC20, Fan23, FT24, FVLA20, FZL^{+}20a, FZL^{+}20b, FCC22, For23, GG25, HNO^{+}21, Han20, HIQ^{+}20, HHKZ20, HZ22, HHB^{+}20, HFS22, HLZ23, HTZ24, HhC20, HCS25, HLSX20, HDY^{+}23, HZCM23, HWW^{+}24, IQ21, IHB21, CLY^{+}22, JL23, JSLJ22, Kaj22, KS23, KKOL20, KA24, KH21, KR21, KS22c, LCC^{+}20, LNS21, LHXS20, LS21a, LS21b, Li22, LTL23, LHX^{+}24, LTL24, LH24, LWL24, LLSH23, LGLM20, LWW25, LZL20a, LWG21, LH22a, LL24a, LZ25a]. based [LXZ^{+}20, LLZ21, LCT^{+}24, MLD21, MPS22a, MTD^{+}23, ML24, MS22d, MRBSFSJ24, MSI21, MFM20, MNK21, NKY^{+}24, NZ23, NJCL20, OCKY20, OJKY23, OdSO20, OOCD^{+}21, OMC21, OMAAORCL23, ÖT24, dOPGdAPU22, PD20b, PLYZ24, PROGDG24, Poh24, PVA24, QZF22, RBC21, RAL22, SD20, SAA21, SGA22, SHB25, Sch21a, SdLV21a, SdL21, ŞKKT25, SGC21a, SGC21b, SLH24, Sin22, SNK22, SAK22, SLL23b, SCU21, SMZY23, TZZZ23, TDWW24, THJ^{+}23, Thi20, TLL24, VJCV22, VPR22, VP21a, WL21a, WGW21, WSZ21, WZZ^{+}21, WCY22, WCCC24, WQ21, WC20, WLS22, WGZ22, WLMZ24, XW25, XXL25, Yam20, YZH23, YZWH24, YS24, YCYZ24, YLJ21, YL24, YWY^{+}20, YYZ25, YKS22, YWZ^{+}24, YLX25, YXYQ22, ZIS21, ZA20, ZS21, ZAF20, ZY20b, ZY20a, ZLY21, ZW21a, ZSS22, ZLZ^{+}23, ZGM^{+}23, ZMP24, ZZW24, ZWL^{+}25b, ZJY22, ZGBH20, ZW24b, ZZ20b, ZYPC20, ZQS22, ZYL20, Zie22]. baseline [Pra25a]. Basic [XCT20, CBP^{+}24, DLL21, DS20, DS22, FYY21, KPS21, SHL22a, ZL23a, ZL23b]. Basins [Wra20a]. basis [ZLL22]. Batagelj [PM20]. Bayes [Sma24, Sma25]. Bayesian [GDDG24, JJY21, MN24, RKCA23, SMZY23]. BBDT [Bor20b]. be [Asa23b, BAE24, BKT20, Cop20a, Cop20c, DV22, GRSFV22, HB21b, HMW24, HYSY23, LS20b, MRM20, MHK22, Mut22, The20a, YMLL23, dW24]. Beall [TdS20]. Beall-listed [TdS20]. bears [McC24]. beautie [ZY20a]. beauties [HV21, HLZ23, KVZ22, TTB22, vR21b]. Beauty [HYSY23]. become [Hla25]. becoming [mYY22]. bedside [LWW24]. been [GSOCS21, LYZ^{+}20]. bees [AGWG^{+}23]. before [BZZ25, HZC21, KKPJ23, KTB22, MF20b]. behavior [CP21, DX20, HBO^{+}20, ICB24, RdNB21, SB21, VPC^{+}23, YRP21, YLCY24, ZW21c]. behaviors [RS22a, WC20]. behaviour [ADD22, ADH23, KPS21]. Behind [SBF23, Emm19, KK23, Li25, LHZ^{+}23, ŞKKT25, YSW^{+}20]. Belgium [DDVV20, JVY22]. belonging [Ben20]. bench [LWW24]. benchmark [WJ22]. Benchmarking [ADGS25, BL24, MRM20, TBPN^{+}20]. benchmarks [ZWC22]. benefit [WB21]. Benefits [GEFJ24, XYY25, YWN25]. Benford [GSB24]. Bergmark [DKWC20]. Bernini [WG20a]. BERT [JJPC20, MTD^{+}23, SLL23b, YZH23]. BERT-based [MTD^{+}23, SLL23b, YZH23]. BERTopic [BPL25, WCCC24]. Best [ATC20, BVvEW25, CLS^{+}20, Fas21a, Wan24b, ZY20c, Wan24b]. best-fit [CLS^{+}20]. better [BW20b, CXD21, Fag20, LZY22, Mut22, Pol20, WXZ22, YWHS23, ZWL23]. Between [Ali22, AJB^{+}21, Asa21, Bre21, CW24, Cao20, CGPR24, CMT24, Che18, Che20, CZFC24, CLZ21, CLC23, CXW^{+}24, DSL21, DDH21, DD23, FRRBDD23, GRSFV21b, GCS25, GC22, Gon23a, GDRPR23, Han20, HGMÁW23, HZJ22, Hüc22, KK20a, KHK21, KKYR24, LZH20, LML21, LdGRV^{+}22, LH22a, LGXH24, LHZ^{+}23, LDA22, LM20, MMF23, Mal22b, MB22b, MHD21a, MLP^{+}20, MKWB^{+}23, MOMGNLG24, MJYS23, Mu21, Nis23a, Nis23b, OdS20, ON20, Ony20, PDH21, Pin23, Pra21c, Pra23c, RBV25, RBR22, RNB22, RBC21, SHL22a, SdL21, SXLC21, SLR^{+}23, SSB22, SD22, SLTS25, SS24, SZK20, SYL^{+}24, SPA20, TKS^{+}23, Thi20, TWH^{+}24, TQA21, TP20, VP21b, Wan24b, WFW24, WKC22, WC20, WG20a, XZ23, XLY21, ZXS22, ZO20, dMGDD24, dGIL21, dSSA22]. between-sectors [SYL^{+}24]. Beyond [THG23, BS20, GDG^{+}20, PVB^{+}24]. bi [MGA22]. bi-lingual [MGA22]. bias [Cop20a, ES22, FWH^{+}23, FBM21, GRSFV20b, GLPC20, HKW24, KSZ22, Kup22, Le 24, LR21, SD20, TYS22, Tut23, WMK21, vdBM24]. bias-free [GLPC20]. biased [CCMM24, DRS20]. Biases [FCY23]. bibliographic [AL24, AM21b, Bat24b, BOPGOP25, CCSX23, Gus22, HC22, HhC20, LCC^{+}20, LML21, OMCC23, PB24, PD20b, Pur21, YWN25, YLJ21, YL24, ZY22]. bibliographic-patent-coupling [YL24]. bibliography [MFM20]. Bibliometric [BP20, GA21, MdMAGB^{+}21, OPMRP20, OA24, TSAMRG23, VCD^{+}23, AD21, APR19, AJBB22, ASK22, Ans23a, ADDCMG^{+}25, ÁRMS21, Bar25, BH25, BW22, Cha23a, CLP21, CCH25, CCMM24, DSH^{+}20, DTHM21, Dej25, DK21b, DL23, DG21, ES23, EGPGGA^{+}23, FMF20, FNK23, FLC22, FC20b, Fog23, FWH^{+}23, FDM24, FN23, GP21, GGSA20, GHGMSM21, Gor21, GWvZ21, Gub24, HB23, HC22, HMW24, HP20, Hir22, HG20, HdSV23, HKN23, HWC22, IOWN20, KDIR23, KLNW20, LKPS25, LB24a, LSFLZLLV20, LHF20, LZYR24, LL21, Liu23, MMFF25, MPR21, MPMR21, MRF21, MSL22, MJ22, MJ23, MMP21, NS23, NB20, NB22, NF22, PRGTOP^{+}24, PYS22, Pet22, PDH21, PL22b, PL23, RG20, RNAH21, RSA20, RBB22, RPL21, RAG25, Rus23, San21a, SSS^{+}24, SdSCdS25, SB20, Sch20a, Sch22, SdLV21b, SK21]. bibliometric [SAR19, SDW^{+}20, SMQL20, SJW21, SK20, SGG20, SA21, SCU21, SC20b, TG20, TMW23, Tom23, UA21, Urb25, VJCV22, VEGSMMC22, VPC^{+}23, VBF24, Vie22, WJZ24, WZM^{+}24, YYZ25, ZAZS22, ZYL20, LHC21]. Bibliometrics [Bor20b, Pet22, PVB^{+}24, AJBB22, BST22, DFRS20, DS20, DS22, Fas21c, Fas24, Kaj22, KKS21, LZYR24, LC21, MMG23, PYS22, SKBSC20, SSSd22, VKY23]. Bibliometrics-based [Bor20b]. biblioverlap [AL24]. BIC [RZS23a, RZS23b]. Biclustering [FZC20]. bidirectional [SZ20, ZA20]. Big [MHK22, SBKK21, CLP21, GVS20, GC20, JCX22, KLJS25, MYK22, RG20, ZZMS22, Ano22a, Ano22b, HWP22, LC20, LC21]. BiGRU [TBA23]. bilateral [Poh24]. binary [QJI^{+}22]. binomial [FD21]. bio [HC22]. bio-eco-techno [HC22]. bioactive [GVS20]. biodiversity [KTK21]. bioeconomy [BP20]. biographical [CNP21]. bioinformatics [JKY22]. biological [Gra20, KR22a, ZM24]. biology [FFH^{+}24, HLS23, Mu21, Sun23, TKZ^{+}23, WXZ22]. Biomedical [FFRRPR^{+}24, VAMBTS25, dOSAL21, BO20, BPW^{+}21, jDXjS21, DMGBG20, Ela21, Ela22, FMF20, JKK24, LCW^{+}20, LTL23, LTL24, Man24, RvLR^{+}20, WYT21, YK23, YX21]. biomedicine [Ke23]. biopharmaceutical [FMF20, Gu21, ZX25]. bioRxiv [SCP22]. biosecurity [WLLW25b]. bipartite [CB20, NKY^{+}24, PFS^{+}20]. birds [AGWG^{+}23, LCS25]. black [SPS20]. blind [GKDB25]. blockchain [Ant20, CB25, SAK22, YD21, YS20]. blockmodeling [CMZ20c, TMZ24]. Blogs [Ort20]. board [GdM20, Tut24, XAA21, dMGDD24]. boards [AS23, RFMMPJ22, WLLL20]. Bologna [RBL20]. Book [NLS20, Bir21, LZ25b, Mal22b, WZ21, YXQG21, Zho24, ZYPC20]. book-based [ZYPC20]. books [BKT20, BWH20a, BWH20b, Dag24, Mal22a, NLS20, Sch21b, Tay20, WC20, ZZ20a, Don21a]. boost [LCY21, LCY22, ZM21]. booster [HMW24]. boosts [BdPFdO25]. bootstrapped [SV20]. bootstrapped-truncated [SV20]. bootstrapping [KYKC21]. borders [HKV20]. Bornmann [Ano20a]. Borromini [WG20a]. borrowed [HP24]. both [Kav25, LHF20, MS22d, PD20a, WPSK25, JTB23]. botnet [HNO^{+}21]. bottles [JL20]. bottom [JTB23, MS22d]. bound [WRT20]. boundaries [Lie20]. boundary [QZS^{+}24, VP21b]. boundary-crossing [QZS^{+}24]. Boyack [Ano23, Sug23]. Brain [SA21, CPTÁ22, TP20, VPSF23]. brain-drain [VPSF23]. branches [GBP21]. Braun [GS22b]. Brazil [MN21a, VC20]. Brazilian [BdPFdO25, CRC20, CdSS25, GdOCRM20, HdSV23, JDSL20, JMP20, dMMdCM22, MNM^{+}20, MN21b, MFP21, RPCF20, RSF20, SG21, TTDH25, dGdO^{+}23]. Breach [RM24]. breakthrough [LN24, LLLL24, WYT21, WZ25, YT25]. breakthroughs [RWL23, WZ25]. Breimer [RJ20b]. BRICS [AM21a, SR23, SSK21]. bridge [RP23]. bridges [vSDC20]. bridging [CXW^{+}24]. Brief [The20a, LS23]. bring [WZNL23]. Bringing [RJ20a]. broader [GC20]. broker [Min25]. brokerages [RYS24]. Brooks [SdSFB22]. browsed [CDZZ20]. budgets [AÁSTMR24]. build [RB25]. Building [GP21, DV20, Gor21]. bullwhip [YY21]. burden [KKYR24, ZZL^{+}20]. burst [ASK22, CR24]. bursts [CWZ21]. bursty [TNS20]. business [Ant20, BT25, Cor22, FNK23, GP21, GAACBA24, HWX22, JS21, JP22, LMSNZG22, MMG21, NS23, OPMRP20, QZ20, RA20, RP20, SBKK21, SBF23, Sta21, TdS20, WWL24a, YWX23, BH24].

calculation [HB22b, LS20b, Thi20, YWZ^{+}24]. calculations [HLR21]. calculus [BF25]. call [GSRG24, HKV20, PHBG21, Pur25]. Cambridge [BBB^{+}24]. campaigns [FFN25]. Can [BKT20, CTH22, DMA^{+}20, GRSFV22, HB21b, LCY21, PXY24, San24, The20a, VAMBTS25, Wal22, ZHH23a, dW24, HMW24, MM25, Ove23, YMLL23, LCY22]. Canada [ALM^{+}23, RBBP24]. Canadian [GBH21, RA20, TdS20]. cancer [PL22a]. candidate [AED20, NJCL20]. canonical [KBSS23]. capabilities [FYY21]. capability [HY22]. capacity [CWZ22, FD21, For23, FBDM25, WG24]. capillary [CdAdFC21]. capital [CZX24, CWLS23, DDF20, HLS21, YZZT24]. capture [Mey25a]. captured [KBS21]. Capturing [BKS20]. Cardiovascular [PL22a]. care [Lam20, Seb23a, Seb23b, SCU21]. Career [Yur24, DS25, GSOW20, GCMJ25, HLS23, Lee24, LZC21, LCD20, LCS25, Par25, The20b, VWR25, Wad24, YWN25, ZY20c]. Careers [BM21, BdPFdO25, Emm19, FCC22, GZOC24, HF22, KR22a, OSS^{+}23, YG20, ZF22]. Caribbean [Cor22, RPAVFV21, VdBC20, dCPT22]. carrier [CP21]. carrot [HPH21, MHD21a, MHD21b]. Carrying [PG24]. carryovers [XlYqZ22]. Cas-9 [SKS^{+}23]. cascade [HYZ23]. case [AD20, ADGS25, ATW20, AHH23, AM21b, Asa23a, ÁRMS21, BP22, BBB^{+}24, BR25, BMA21, BFG^{+}21, BWH20a, BWH20b, CTLH24, CRC20, CLL22, CGPR24, CMMO20, CSC23, CKO21, CF23, DFRS20, DDH21, DDF20, EEÁ21, EGV22, FNK23, FSEJF^{+}20, GLK^{+}23, GJ24, GEFJ24, GAPR20, GdOCRM20, HAT23, HB21b, HMW24, HP21, HL21a, JS21, JCX22, JXY^{+}23a, JXY^{+}23b, JDSL20, KN21, KNT20, KVZ20, KSAT24, KUS22, LYT^{+}20, LMH22, LL25, LYZ^{+}20, Lin20, Lin21b, LLL25, Liu20, LGMK21, LWZ22, LPY22, LMLMPA22, LC20, LC21, MPS22a, ML24, Mas20, MQS^{+}20, MOMGNLG24, MJYS23, MFF^{+}20, NSM24, OMC21, OMFJ22, PS20, PSKS22, PROG21, PA22, PSH22, Pur21, Pur24, QQL21, QZ20, RSPC20, RS22b, RWL23, RBB^{+}24, RBL20, RM24, SSCK21, SD20, SBP20, SBKK21, STAÁMR24, SdS25, SKS^{+}23, SdSFB22, SPS20, SM23]. case [SGG20, SLX21, SZZ23, SSBC22, TMZ24, TP20, TG24, VPSF23, Vie24, VC20, WH23, Wan24b, WY23, WGA22, Wra20a, WLLL20, WYL21, YD21, YK23, YH24, YWW23, YNO22, YJJ25, YTJ23, mYY22, ZLL21, ZGM^{+}23, ZCSK24, ZWL^{+}25b, ZM23b, vD21, vSDC20, vdPF25]. case-controlled [SSBC22]. case-study [ML24]. cases [CL21a]. cat [dS21b]. catalog [TSAMT21, GMCRV21]. catch [IPP22]. catching [JLL21]. catching-up [JLL21]. catchphrase [TDM24]. Categories [dSSA22, JS21]. categorization [AHH23, SS23]. category [ZS24]. causal [Dej25, WLMZ24, ZHH23a]. causality [LDA22, OJKY23, TWH^{+}24, ZBKW24]. causes [AGB23, GFVK21, KS24, LHW21, OPS25]. causing [GB24]. CCA [BWH20a, BWH20b]. CCRO [IQ21]. CEE [MSJ23, SMJ21]. cell [WXZ22]. censoring [GRSFV21c]. censorship [PHBG21]. central [Bös21, HSE23, WCH^{+}20, CF23, Jok20, MMB23, Ove23, PMB^{+}23]. centrality [Cho20, WGLN23, YX21]. centralization [Zha21a]. centralized [LLL22]. centres [Ali24]. centric [WG22]. centuries [DK21b]. century [BM20b, MMG23, RP24b, dDM25, GMCRV21]. Cernat [Naz24a]. certain [IHLP21]. certification [CCdC20]. CFMf [LLM24]. chain [FAY25, SBP20, YY21]. chain-of-thought [FAY25]. challenge [GC20]. Challenges [HLG21, HG20, KMM22, LDMVS21, PG24, XYY25, ZWSC20]. chambers [SS22]. change [DDD^{+}21, HZC21, HDY^{+}23, JDG^{+}20, KMGS21, KUS22, Liu23, LZ25b, Mey25b, PY25, dGIL21]. change-related [dGIL21]. changed [SMBD^{+}24]. Changes [SZP^{+}24, ARR21, DCdA25, KG22, PY24, PKA24, PFS^{+}20, WVH21, YP24, ZJ22]. Changing [KL22, BM21, MMGP25, ZM23a]. channels [HGM21, TSDAMRG24]. chapters [YYZ25]. character [LZH^{+}22]. characteristic [LRH22]. Characteristics [Cha21, Ela22, DR20, AL21, ACDS22, CDZZ20, FVLA20, GKDB25, GSH24, GS24, HYSY23, HHY22, JC22, KCC22, LL25, LC21, PMB^{+}23, PLW23, PXY24, RWL23, SJW21, SJW22, SLL23b, VCJ22, WL21b, YWHS23]. Characterization [LDDL24, dMMdCM22, FVGBPA23, FFRRPR^{+}24, MMGL23]. characterizations [ER24]. characterize [LWW25]. Characterizing [HWZ21, AMTSRG21, DSL23, DLS23, OSS^{+}23]. charges [AÁSTMR24, Asa20b, Asa23b, Asa24, BGWZ20, MS22b]. charging [QZ20]. Charting [Bou20]. ChatGPT [BZZ25, BL24, BWE24, GdMDD24, Kou24, OMCC23, San24, YNL25, dW24]. checklist [RSV^{+}20]. chemical [GLK^{+}23, Kim23, Sun23]. Chemistry [KPM22, HF22, Pra20d, Reh20]. Chief [Lin20]. China [CH20, CXD21, CTH22, CLL22, CGJM24, DLL21, jDXjS21, GEFJ24, GC22, Gu21, JCK22, JXY^{+}23a, JXY^{+}23b, KK20a, LCC^{+}20, LGD21, LYT^{+}20, LMH22, MYK22, MH24a, NC20, QQL21, QSYL23, RP24a, SG22, SQC^{+}20, SWL^{+}23, SC20b, SZK20, SZP^{+}24, TKZ^{+}23, WCM20, WL21b, WZZ24, WJZ24, WLLW25a, WZZZ20, YH24, YLf25, YX21, YT25, YS23, lYnLyO^{+}20, ZZL^{+}20, Zha23b, ZBG^{+}24, ZXL^{+}24, ZPH21, ZL23c, ZL20a, ZYL20]. Chinese [AlYT21, Cao20, CtRbQ22, CxGdH^{+}24, CFC22, Gon23a, Gon23b, GX22, LL25, LZC21, LWW25, LL24b, NLS20, TZ24, WZL22a, XYY25, XlYqZ22, YLM24a, YMD20, ZM21, ZM23a, ZWL23, ZO20, ZJ25]. chiropractic [The21]. Choice [JC22, AKW23, DD21b, GMPRSFV25].

Choosing [LCS25, QZF22, LLH20]. Chou [LYZ23]. Chronic [GGSA20]. Chronological [MZZL21]. Cinvestav [LMLMPA22]. Circling [vdB25]. citability [JADG20]. Citation [ASK22, BWH20a, Don21a, FFH^{+}24, GVK24, KBZT21, KZH25, KHK21, LNH24, Mou21b, NLS20, NZ23, Par21, SSBC22, WWL24b, AD20, AUM21a, AUM21b, AFH21, AA22, BZL21, BGB25, BVvEW23, BVvEW25, BBP21, BT22, BPLIdMA^{+}20, BZX^{+}23, BA22a, BA24, BZG22, Bor22, BW19, BWH20b, BW20c, BW20b, BMS21b, Bre21, BY23, BRJ21, BI24, CTH25, CGPR24, CPN21, CNT25, CSL22, CDQ25, CWL^{+}20, CZFC24, CKT21, Cop20a, DKWC20, DDH21, DMS20, DBR21, ES22, FWDQ20, Fan20, Fan23, GZSC22a, GZSC22b, GKM^{+}24, GMB22, GBP21, HH21, HM23, HS22b, HB21a, HMPS24, HZ22, HGMÁW23, HFAR23, HM20b, HMZ21, HHY22, HZW22, HCZ^{+}22, IQ21, IHB21, Jab25, CLY^{+}22, JA23, JHVC21, JHVC23, JJPC20, JL25, JL23, JC23, Jia25, KS23, KA24, KCS25, KS22c, LCB21, LCC^{+}20, LS20a, LZH^{+}22, LW24, LZYR24, Lin21b]. citation [LGMK21, LWG21, LL24b, LMSNZG22, LHZ^{+}23, LXZ^{+}20, LYLC23, LTXZ24, MZL20, MZZL21, MLD21, MS22b, MCDFSP24, MPS22b, MS22d, MNM^{+}21, Mou21c, MFP21, NKY^{+}24, Naz24b, NII24, Nis23a, Nis23b, NK24, NM25, NJCL20, OŠV^{+}24, Öz24, PB24, PSKS22, PD20a, Poh24, PLT22, Pra21b, Pra22, QWS^{+}23, QYL21, RBR22, RKMG20, RSM24, RGVQTS^{+}24, RLZ^{+}24, RBC21, SBP20, SdS25, SH20, SHB25, SKS^{+}23, SYHW20, SS23, SVY22, SWT20, SXLC21, SLH24, SFJO20, SVC20, Sin22, SD22, SSC21, SLTS25, SS22, SLY22, SMZY23, SPA20, TZ24, TXL21, Tom23, TA20, TDM24, TTB22, VPC^{+}23, Vie23, VP21a, VRVH21, VA22, Wad24, WZJ^{+}20, WWY^{+}20, WSZ21, WCY22, WWS23a, WZZ23, Wan24b, WL24, Wan24a, WLL25, WC20, WPS22, XW25, XDLZ23, XXL25, YTZ20, YZH23, YWB22, YTJ23, YXYQ22, Yur23, ZMA23, ZA20, ZS21, ZW21a, ZZW^{+}22a]. citation [ZY22, ZZW^{+}22b, ZM23a, ZW24a, ZSX24, ZYWA25, ZS20a, ZS20b, ZZ20a, Zho21, ZW24b, ZYL25, ZXL20, dS21c, dSV21, dSN23, dSVN24, dMALIGBM20, dW24, BWH20b, CG24, Mou21c, The21, SVY22]. citation-based [Fan23, KS22c, MS22d, Poh24, RBC21, XW25]. citation-forecasting [WSZ21]. Citations [FM21, SSC21, Tei24a, ASIN24, ARDA20, ASX22, Ans23b, AGB23, AMZ24, BD22, BS23, BSS24, BDPP23, BRG^{+}23, Bor20a, Bor22, BY23, CCA22, Cop20a, CLRMR22, CLR25, DDH21, DMGBG20, FRRBDD23, GSP21, Gon23a, HP21, HNM^{+}24, cHhC24, JCZ22, KFM20, KTK21, LBP22, LL20, LTL24, LML21, LLSH23, LYS22, LLP21, LH22a, LWW24, LM20, LS23, MS22c, Mal22a, Mal22b, MFMC21, MPC22, MMTOMLC21b, MMG21, MRC22, NZL21, Nis23a, Nis23b, NM25, Pea21, PD20b, PCG^{+}20, PA21, PVB^{+}24, QJI^{+}22, QPH^{+}25, QLW25, RAS21, SF20, San20, SdSFB22, Sha21, SAR19, SSL22, SA24, SRL24, SS24, SDRS23, TFWC24, Tom21, THG23, VEGSMMC22, Wad20, WWZW20, WZJ^{+}20, WGC20, WSZ21, WXZ22, WKC22, WLBC20, WB21, WXD24, XJZW24, XDL22, YXQG21, YWX23, Zbo25, ZW24a, ZJ25, ZYPC20, Tei21, dSV21, dSV23, MMTOMLC21a]. cite [Wad20, CNT25, XXL25]. Cited [GVK24, KK23, ZNNA20, BJS22, BKT20, BL20a, CDE23, CT20, Che23, Cho21, CLC23, DGRÁTS23, DGGD22, ECD24, FN23, FGB25, GSOCS21, GWLY24, GDG^{+}23, KKS21, LS20b, LZYR24, LM20, NLS20, Pra24, RL20, Sch24c, Sta21, Ste23, TZY^{+}23, TDM24, WZL22a, Wra20a, XDL22, dSVN24]. cited-references [BL20a]. citedness [ADD23]. CiteScore [Fan21, Kra20, OdS20, dS21b]. CiteSpace [WRT20]. citie [YX21]. cities [WSL23, ZBG^{+}24, vRW25]. Citing [GVK24, LAT22, HKW24, LRXC21, MFM20, ZW25]. citizen [HCS21, PRS^{+}21, SOSC24]. city [Cot22, jDXjS21, TPVVPA21]. civic [MBR^{+}25]. civilian [LRSB21]. CLARA [BZX^{+}23]. clarity [Kra20]. class [LRP21, XS25]. classes [dSN23]. classic [BWH20a, BWH20b, SdSFB22]. classical [AGDE25, GS22a]. classics [LZYR24]. Classification [GVK24, AFH21, ASX22, AG23, AGDE25, ARR23, BBB^{+}24, BA24, CNT25, CM20, DABSC20, HZCM23, HWW^{+}24, JC23, Jia25, KCC23, LXZ^{+}20, LCT^{+}24, LXS22, LRXC21, MMFF25, OCKY20, PLH20, QJI^{+}22, QWS^{+}23, RGVQTS^{+}24, RAL22, SM24a, WMCL22, XXL25, YL24, ZS21, ZZW^{+}22a, ZZW^{+}22b, ZS24, dS21c, dSN23]. classification-based [OCKY20]. Classifications [SMQL20, ÁLGBdMA24, BY23, TP24, ZSSH22]. classifier [MK23]. classifiers [PJ24]. classify [LCB20, PJ24, WG20b]. classifying [GCT^{+}20]. Climate [SMS23, HDY^{+}23, JDG^{+}20, Liu23, dGIL21, DK21a]. Clinical [THvB21, dOSAL21, CDZ^{+}20, CBP^{+}24, DTHM21, EBB^{+}20, Hua24, LTL24, LWW24, NHAB21, SYHW20, ZL23a, ZL23b, ZF22]. closed [HKV20, ZWHS22]. closely [Pur25, Yur20]. Closer [CZFC24]. Closing [ŞKKT25]. cloud [CR24]. club [NSM23]. clues [ES22]. cluster [CC21, LMC22, LLL22, SVC20]. clustered [KCC23]. Clustering [ZC23, ABAAN^{+}23, AED20, BVvEW25, CGJ22, CZZN25, Don21b, HAS^{+}20, JPS21, JS21, KYPC20, KP23b, LCC^{+}20, LHC22, LHC25, SA24, VMF22, WQ21, XW25, XHQZ20]. clustering-based [AED20]. clusterings [LBW21]. clusters [BVvEW23, Hüc22, LZD^{+}24, Ste21]. CNC [ZDLR21]. CNCI [PSA22, PKA24]. Co [BP23, CGPR24, GDRPR23, PLYZ24, Yur20, ZZ20b, ALR24, BBP23, BMS21a, Ben24, BA22b, BOV21, CLMM21, CLZ21, Cor22, DX20, DLC21, DK25, EGP22, FZC20, FMB22, GAACBA24, GAPR20, HC22, HFS22, HBN^{+}21, HX25, IOWN20, JA23, KNT20, Kos21b, KN22, LBS23, LH22a, LZ24b, MK23, PB24, dOPGdAPU22, Poh21, PKCOG21, RKMG20, SSCK21, SBP20, SHH23a, SHH23b, SH20, SFJO20, SVC20, SDRS23, TMZ24, The23, TD24c, VBF24, WKC24, WB22, XZZ22a, XZZ22b, YWY^{+}20, Zha21a, ZS20b, ZYZW23, dOCP21, dGIL21, SVY22]. co-author [DLC21, DK25]. co-authored [FMB22]. co-authoring [SSCK21]. co-authors [The23]. Co-authorship [BP23, ALR24, BBP23, BMS21a, BOV21, CLZ21, EGP22, GAPR20, HC22, JA23, Kos21b, LBS23, LZ24b, SHH23a, SHH23b, SFJO20, SDRS23, TMZ24, TD24c, WKC24, WB22, dOCP21]. co-authorships [HFS22, dOPGdAPU22]. Co-citation [CGPR24, RKMG20, SBP20, SH20, SVC20, ZS20b, SVY22]. co-creation [GAACBA24]. co-ethnicity [PB24]. Co-evolution [PLYZ24]. co-evolved [IOWN20]. co-first [HX25]. Co-follower [GDRPR23]. co-invention [Zha21a]. co-nomination [KNT20]. co-occurrence [Ben24, DX20, FZC20, LH22a, MK23, VBF24, YWY^{+}20, ZYZW23]. co-publications [BA22b, KN22, Poh21]. Co-word [ZZ20b, Cor22, HBN^{+}21, PKCOG21, XZZ22a, XZZ22b]. co-words [CLMM21]. Co-worker [Yur20]. coaches [ZW25]. coalitional [AdAFR24]. coatings [KY20]. coauthor [LHZ^{+}23]. coauthors [Vin23, WW25]. coauthorship [HPL21, JDSL20, McC24, Xie21]. Cochrane [HFAR23, RAF25b]. COCI [MMTOMLC21a, MMTOMLC21b]. cocoon [HYY25]. coded [Bös21]. codified [ZL23c]. coefficient [BB23]. coexistence [UA20]. coffee [CCdC20]. cogency [OŠV^{+}24]. cognitive [CZX24, QQL21, ZGM^{+}23, vSDC20]. cognitive-based [ZGM^{+}23]. cohesion [Fog23, GYWZ21, Wan24a]. cohort [Lam20, ZAZS22]. coin [CtRbQ22, CG24]. Collaboration [Asa21, Lin21a, LL24a, MB22b, McC24, Mor19, Poh20, dGdO^{+}23, ZAF20, ADD22, AS24, APR25, APR19, AM21b, AG23, Asa20a, Avd21, BZL21, BJ21, CCdC20, CFW21, CTLH24, CFA^{+}23, CMV23, CWK21, CSJW22, CXZM25, CdSS25, CRBRGS21, CLZ21, CMZ20c, CMK24, Dem22, DGRÁTS23, DSL23, DLS23, EEÁ21, FWDQ20, FFL21, FMB22, FSR21, GLK^{+}23, GCS25, GEFJ24, GC22, GdOCRM20, dMGC23, HSE23, HWZ21, Hir22, HX25, Hüc22, HL24, JYT21, Jok20, JI22, KSS21, Kwi20, LHW20, LL25, LdGRV^{+}22, LY20, LLLL24, LZY22, LZ24b, LGR^{+}21, MH24a, MF20b, MSL22, MBF23, MNM^{+}20, MS21f, MOMGNLG24, Mor20, NCD22, PWZ25, PB24, PS20, PSKS22, PLYZ24, PRS^{+}21, Plo25a, PSA22, PKA24, Pur24, PFS^{+}20, QZF22, RP23, RP24b, RP24a, RP25, Sha24b, SXLC21, Söd23, SYK22, SZK20, TXL21, UA20, Vie22, Vie23]. collaboration [Vie24, Wag25, WJZ24, WFW24, YCL25, YSW^{+}20, ZF22, ZQML23, ZCSK24, ZPH21, ZCL20, dFBGCGASG24, dBGP21]. Collaboration-based [LL24a]. collaboration-citation [BZL21]. collaboration-knowledge [CSJW22]. Collaborations [GAB21, AAD22, AAD24b, BSA23, jDXjS21, EGP22, GFVK21, HTGW20, HZJ22, Hüc23, JDSL20, Poh24, RZS23a, RZS23b, SHF24, TG24, VEGSMMC22, XWGD22, dSMNdV^{+}21]. collaborative [BH24, KK20a, KFCK21, LWL24, NSM23, XDW22, YT25, ZGM^{+}23, DG24]. collaborator [LH24]. collaborators [HCN21, YLC^{+}20]. colleagues [SO21]. Collecting [DvE20]. Collection [Bou21, HWL23, LL20, Liu21, LHW21, LNH24]. collective [dFBGCGASG24]. collectively [BMBA20]. College [PSH22, HCS21]. colleges [GdM20, ATC20]. Collel [Yur23]. colloid [Zbo25]. colored [YTJ23]. columns [Fan20]. combat [AD21]. combination [dLBdAPUdLB^{+}25, GHW22, WZZ^{+}21]. combinations [LW24, YTZ20]. combinatorial [MSI21]. combined [Mut22, WSZ21]. Combining [CZZN25, CD22, EGR22, Hal20b, KS22c, YY22b, ZJ22, GJ24, GB24, HZW22, KZH25, LGD21, MBH25]. Comeback [GCMJ25]. Commemorating [BGS20, HP20]. Comment [LCC23, LYZ23, MHD21b, Mou21a, SD23a, Cop20a]. commentary [Naz24a]. Comments [Min22, PM20, Pra23a, BAE24, Bjø20a, Bjø20c, MMGC^{+}24, Pra22, SdL23]. Commerce [LHC21, SBKK21]. commercial [SHB25, SM20, Wal22]. commercialisation [HL21a]. commercialization [SZZ23]. common [Cot22, LZD^{+}24, TSDAMRG24]. commonality [XLY21]. Communication [BP23, YMW^{+}20, FH23, FT21, GdM20, GDG^{+}20, GDdZ22, GDG^{+}23, GTH24, GDdZ25, GAB21, HGM21, KKK^{+}25, MDGR21, Par24, PA22, RS22a, TSDAMRG24, THG23, TDM24, YNO22, YNL25, ZYPC20]. communications [AA22, Iva25, JT21, SK20]. communicators [RS22a, ZW25]. Communities [Ara20, vSDC20, AMTSRG21, CGJ22, DD21a, IVCE20, LC21, ML24, dMMdCM22, SCK21, WKC22, ZLY21]. Community [Min25, WL21b, ATCS20, AUA^{+}23, DLL21, JDSL20, LGD21, MS21d, NF22, XHQZ20, YCL25, YKTM20]. Community-level [WL21b]. companies [BdOdS21, FT24, LWW25]. company [CGJM24, ZX25]. company-funded [ZX25]. comparable [SdS25]. Comparative [AS25, ACD^{+}24, Lec21, MML^{+}20, ZPH21, Bor20a, CW25, GKDB25, GDM22, JHVC23, KP23b, Liu20, LZL22, MMG21, MBR^{+}25, NC20, Ort20, RBC21, SdL21, Seb23a, SMJ21, SSM21, SMA21, TLL24, VA22, WWH21, XL20, YP24, YLJ21, YL24, YMX21, YS23]. compare [KYKC21, MN24, Poh24]. compared [CHJ^{+}25, Mal22a, Seb23a]. Comparing [CSA20, Gus22, LHC22, Mit20, PSA22, QPH^{+}25, WY23, YJJ25, ZSSH22, ZL20a, AÁSTMR24, BPL25, LZF^{+}24, WSL23]. Comparison [BRJ21, Cha23a, CN24, GKM^{+}24, Wan24b, YXQG21, AGDE25, Che18, Che20, FD21, KBS21, LLM24, LML21, MLS21, MS22c, MMTOMLC21a, MMTOMLC21b, MKWB^{+}23, MFM20, ODQ24, PD20b, PA21, SHB25, Sch21a, XW25, ZS20b, dSSA22]. comparisons [GC22, HB20]. competencies [SNK22]. competency [LNS21]. competitive [AAD24a, GMPRSFV25, MMF20, PA22]. competitiveness [ZP20]. complement [Fas20a, Fas20b]. complementarity [PKA24]. completeness [DDS22, MB22a, RBV25]. completeness-aware [DDS22]. Completing [WQ22]. Complex [TBPN^{+}20, LZL20a, OOCD^{+}21, dOPGdAPU22, RNB22, SRL24, TSA24, Wag25]. complexities [MS21f]. Complexity [DDF20, Wou25, JJ23, PY24, PY25, WWS^{+}23b, YP24, Yen25]. Complexity-based [DDF20]. Compliance [STAÁMR24, SCFSAB^{+}25]. component [KT21]. components [YTZ20]. composers [GS22a]. composite [EGR22, FDM24]. composition [Jam25]. compositions [BR25, ZWL25a]. compounds [GSG22a, GVS20]. Comprehensive [SDW^{+}20, ECD24, Sca25, SM23, VFL^{+}21, YYZ25]. comprehensiveness [Men25]. compromise [Pra21c]. computation [Le 24]. computational [DBR21, SM22a]. computationally [GdMDD24]. Computer [FCC22, AHH23, BMA21, DFRS20, DFS22, DKWC20, FRS23, GEFJ24, HAT23, HBW^{+}21, HCZEP20, KSS20b, LFV21, LYZ^{+}20, LLL25, MSV23, MMGL23, San21a, San25, TA22, TNS20, VPSF23, dFVFA^{+}24, Wan24b, Kup22, MAS^{+}21]. computing [AHH23, CR24, LPD21]. Concentration [KHC^{+}25, Fer21]. concentrations [Ste21]. concept [BWH20a, BWH20b, DLC21, Don22, TKS^{+}23]. concepts [BWH20a, BWH20b, GS24, SSC^{+}23, SCW^{+}23]. conceptual [ACD^{+}24, GMB22, MPR21, MPMR21, NS23, ZGM^{+}23, ZNW22]. conceptualizing [CLMM21]. conditions [Vie24]. conducted [KK20a]. Conference [BMW24, DMG20, GDR22, Pur21, KG24, LGE^{+}23, OPZP25, San25, TFWC24, WXZ24, vR21a]. conferences [FH21, FH23, KN22, LPKG20, LRW25, MAS^{+}21, OPZP25, Wan24b, ZYL20]. confidence [dOSAL21, WXZ24]. confident [WVH21]. Configuration [WZZ24]. configurations [CW25, LWX21]. Confirmatory [GRSFV20b]. confirms [Pra24]. conflating [MS22d]. conflict [LRSB21, RBBP24]. conflicts [SR24]. confronting [RPCF20]. confusion [WH24]. Congress [SDC22]. congruent [WH25]. conjoint [MBD25a, MBD25b]. Connecting [CPTÁ22, She20]. connection [Tut24]. Connections [CFI22, dMGDD24]. connects [Kav25]. consensus [MRBSFSJ24]. Consequences [PSA22, EBW23, KG22, LHW21]. consider [TTB20]. considered [VWR25]. considering [DABSC20, cHhC24, KG20, WZZZ20, YF24]. Consistency [YTM21, Gon23a, ZNW22]. consistent [ÁRMS21, WXZ24, XJZW24]. consolidation [LC22]. CONSORT [OMAAORCL23]. consortia [FRRBDD23]. construct [CLMM21]. Constructing [CNA22, ON20, PW20, LSY21]. construction [Bor22, JL23]. consumption [Par21]. contagion [Mou21b, Mou21c]. contain [Pra24]. Contemporary [LZYR24]. content [AM21a, KSS20a, KP23b, LDDL24, LWL24, LB25, MZL20, Nuz21, SDW^{+}20, Ste22, VRVH21, WMCL22, WZHW25, WXZ24, YJJ25, YMX21, YYC22, ZW21a]. content-based [LWL24, ZW21a]. contents [Ela25]. Contested [Lec21]. Context [JHVC23, CDQ25, DHSS20, GRTPAJ21, Gu21, JHVC21, JJPC20, JVY22, LPD21, MZL20, NB22, Nis23a, Nis23b, NK24, WYT21, WLL25, ZZ24, Zho21]. Context-aware [JHVC23, JJPC20, WLL25]. context-enhanced [ZZ24]. contexts [Par21, PT24, RPCF20]. Contextual [LP23, GLPC20, HIQ^{+}20, MZD22, WZJ^{+}20]. Contextualised [JC23]. continue [Sch24c]. Continued [SYHW20, ZQML23]. continuum [LTL23]. Contradiction [Cao20]. contrast [LCC^{+}20]. Contrasting [SS24, CL21a, LS23]. contrastive [Bir21, FAY25, Mu21, TZZZ23]. contribute [ATC20, LWX21, VCV22]. contributed [AHRH24]. Contribution [MSJ23, dSSA22, CCZL23, CCMM24, FMF20, GUG^{+}22, KM20, LWG21, LH22b, SMJ21, VCJ22]. contributions [CCH20, CNA22, CDQ25, DK25, HCS21, Old22, QZS^{+}24, Sun23, THL23, WWY^{+}20, XZ23]. contributor [DLC21]. contributors [OPZP25]. control [dLBdAPUdLB^{+}25, BdPFdO25, KYL20, RB25]. controlled [SSBC22]. controlling [DGSVDG20]. Controversial [Naz20]. converge [Pin23]. Convergence [CMMO20, EPP23, AMS21, BCG22, CLS21, DG21, KYL20, LHK21, LS21a, LS21b, LHC22, LMG20, WH23, YWL25, ZMK22, ZM23b]. conversational [Ara20]. conversion [Kua20]. converts [MMP21]. convolutional [JJPC20]. Conway [FD21]. Cooperation [MF20a, SSK21, KLM24, MYK22]. cooperative [CLL22, jDXjS21]. Coopetition [Hüc22, YJS21]. COPE [Lin20]. copy [dS21b]. copy-cat [dS21b]. Core [Bou21, HWL23, LL20, Liu21, LHW21, LNH24, MJ22, MJ23, CLC^{+}24, KFCK21, LN24, LNS21, Pra23a, SD23a, SD23b]. Core-periphery [MJ22, MJ23]. Coronavirus [BMBA20, HV21, PKCOG21, SB21]. coronavirus-related [SB21]. coronavirus-seeking [SB21]. corporate [CLMM21, CL20, IHLP21, MR20, Poh21, SVC20]. corporation [YLCY20]. corporations [YWX23]. corpus [Bir21, VPDAG22, Li22]. corpus-based [Bir21, Li22]. correct [GMPRSFV25]. correcting [CCA22]. Correction [Ali21b, AUM21a, uHBKK21, Bi23a, Bi23b, Bjø20a, BWH20b, BW20c, CMZ^{+}20a, Che20, DS22, Fas20a, FZL^{+}20b, Fuk22a, Hen20, JXY^{+}23a, LCY22, LUH21a, MPR21, MS21b, MMTOMLC21a, MBD25a, MJ23, MSP23a, Mou21c, Nis23a, RZS23a, SH21, SHL22c, SHH23a, Sch22, SGC21a, XZZ22a, ZL23a, ZWX22, ZZW^{+}22a, Bor20a, MA21]. correlated [Moo21]. correlates [RSV^{+}20]. Correlating [ADD23, TdS20]. Correlation [OdS20, RBR22, CZFC24, Cho20, DD23, HLSX20, LCD20, MHD21a, SXLC21, SS24, XZ23, ZJ25]. correlations [MJYS23]. corresponding [CTH25, GdOCRM20, dGdO^{+}23]. cost [AÁSTMR24, ZC24]. costs [AL20, Pol20, dSAKT20]. could [Cop20d]. Council [BMM22, RPL21, WJZ24, XYY25]. count [DMGBG20, FD21, GAB21, HM23, MLD21, RBR22, SB25, SXLC21, TXL21, WWS23a, Wra20b, ZYWA25, ZXL20]. counterfactual [CMT24, CLR25]. counterparts [ZWL23]. counting [Pra21b, Sun23, WLLW25b, ZS20a]. countrie [RNB22, TM20]. countries [ADM22, Ali22, AAHY20, Azm22, BGP20, CRC20, CLP21, CF23, DLS23, ES23, HTGW20, JME20, KDIR23, MML^{+}20, MSL22, MBF23, MAN22, Naz24a, Ony20, Poh20, Pur21, RKCA23, San21a, SB21, SAR19, SR23, SSK21, UA20, VC22, Vie23, VdBC20, XYY25, YMW^{+}20, ZL23c, ZP20]. Country [Jam24, AAA^{+}23, CPFVPA25, CMV23, CT20, FZhCC21, Fer21, GSP21, GMRVU24, KG24, MS21a, MS22a, MS21b, MS21c, MGCR24, MDGR21, MN23, Old22, Ort20, SSCK21, Sha24b, SAR19, SSB22, UNK23, VPSF23]. country-case [SSCK21]. country-level [KG24]. country-undefined [MN23]. counts [Cop20a, ES22, LLH20, LH22a, MNM^{+}21, SWT20, She20, SLTS25, TWH^{+}24, TFWC24, Tom24, WXZ22, dW24]. coupled [CdAdFC21]. coupler [CGPR24]. coupling [HC22, HhC20, LLL22, PB24, YLJ21, YL24, ZY22, ZQS22, HAS^{+}20]. CoV [HB21b, PHBG21]. covariation [WH25]. Cover [WLS22, WXZ22]. coverage [AOM24, CHJ^{+}25, DMGBG20, GKM^{+}24, Gus22, KBS21, KL24, LNH24, Lyu21, MMTOMLC21a, MMTOMLC21b, Ort20, ODQ24, SSM21]. covering [SdSCdS25]. COVID [AD21, Ali22, AM22, ARR21, BD21, BdMPH21, BJS22, BTP24, BFG^{+}21, CFW21, CFA^{+}23, CP21, CCC22, Coc21, CA22, DDVV20, EXT^{+}21, GCGB^{+}24, DR20, HB21b, HL21b, HKV20, KHT22, KTB22, KCC21, LCC23, LYZ22, LYZ23, LDAAAB21, MA21, MAS^{+}21, NRGvSTS23, Pal21, PHBG21, SSCK21, SSdCRD22, SMG^{+}24, SB21, SP20, SCP22, SFC^{+}23, TMW23, TTB22, WR21, WL21a, WZM^{+}24, YNO22, YNL25, YWB21, ZCW21, ZLM^{+}23, Zha23b, ADD22, ADM22, BW22, Fas21c, HB20, RBB22, TZY^{+}23, dSTE21]. COVID-19 [AD21, Ali22, AM22, ARR21, BD21, BdMPH21, BJS22, BTP24, BFG^{+}21, CFW21, CFA^{+}23, CP21, CCC22, Coc21, CA22, DDVV20, EXT^{+}21, GCGB^{+}24, DR20, HB21b, HL21b, HKV20, KHT22, KTB22, KCC21, LCC23, LYZ22, LYZ23, LDAAAB21, MA21, MAS^{+}21, NRGvSTS23, Pal21, PHBG21, SSCK21, SSdCRD22, SMG^{+}24, SB21, SP20, SCP22, SFC^{+}23, TTB22, WR21, WL21a, WZM^{+}24, YNO22, YNL25, YWB21, ZCW21, ZLM^{+}23, Zha23b, ADD22, ADM22, BW22, Fas21c, HB20, RBB22, TZY^{+}23, dSTE21]. COVID-19/SARS-CoV-2 [HB21b, PHBG21]. create [KSAT24]. creating [VCV22]. creation [CLL22, CNA22, GAACBA24]. creativity [FLDD20]. credibility [AA25]. credit [AA22, DS25, DLC21, DGGBDG21, GAB21, HCZEP20, Sun23, dSDE21]. credit-like [dSDE21]. CRExplorer [BL20a]. CRF [HTZ24]. criminology [Kav25]. crises [Coc21]. crisis [ADD22, CFW21, Dej25, HB20, SCU21]. CRISPR [SKS^{+}23]. CRISPR/Cas [SKS^{+}23]. CRISPR/Cas-9 [SKS^{+}23]. criteria [Cop20c, Hug24, MSP23a, MSP23b]. criterion [JI22, OŠV^{+}24]. Critical [Bor22, dSVN24, KK20a, RBC21]. criticism [TK21]. critique [Ste22]. critiques [BAE24]. Cross [Fer21, GSP21, LXY23, UA20, AAD22, AAA^{+}23, Ali21a, Ali21b, BRG^{+}23, CMV23, Che23, CG22, jDXjS21, ECD24, FCW22, FZhCC21, GC22, Hao24, KKK^{+}25, Liu24, LCT^{+}24, MS21a, MS22a, OMAAORCL23, PY24, SR24, SdLV21a, SS24, UNK23, WH24, Wan24a, WH25, ZGM^{+}23]. cross-city [jDXjS21]. cross-correlation [SS24]. Cross-country [Fer21, GSP21, AAA^{+}23, CMV23, FZhCC21, MS21a, MS22a, UNK23]. Cross-cultural [LXY23]. cross-database [GC22]. Cross-disciplinary [UA20, CG22, FCW22, Hao24, KKK^{+}25, PY24, WH24, Wan24a, WH25]. cross-discipline [SR24]. cross-domain [Liu24, ZGM^{+}23]. cross-field [Che23, ECD24]. cross-generic [Hao24]. cross-lingual [LCT^{+}24]. cross-sectional [Ali21a, Ali21b, BRG^{+}23, OMAAORCL23, SdLV21a]. cross-sector [AAD22]. Crossing [OM20, QZS^{+}24]. Crossref [CCA22, GKM^{+}24, MNK21]. crowd [Gor21]. crowds [SZY24]. Crowdsourcing [Gor21]. cryptocurrency [Ant20]. CSC [LWX21]. CSC-funded [LWX21]. CSSCI [CFC22, GC22, YS23]. CSSCI-indexed [CFC22]. Cuba [RP23, RP24a]. Cuban [GRTPAJ21, RP24b]. cube [Wou25]. cultural [DBR21, ELR24, LXY23, MRBSFSJ24]. culture [DDU24, MS23, WPBH24, dSV21]. cultures [YGO22]. cum [vdBM24]. Cumulative [RPAVFV21]. Current [dOSAL21, CD22, HG20, SvWC21]. curricula [Cha23a]. curriculum [PSH22]. cursed [FC21]. curtains [SBF23]. curve [RZS23a, RZS23b, ZSL24]. curves [Naz24b]. cut [GAPR20]. cut-points [GAPR20]. cutting [Pra20a]. CV [WLHY24]. cyberinfrastructure [SCW^{+}23]. cycle [GAACBA24, GMB22, HTZ24, RTT23]. cycles [TNS21]. Czech [TMNH^{+}24].

D [BWLX20]. daily [Pet22]. Danish [DDD^{+}21]. Data [DSB22, The20a, WD21, ZMP24, AMMN21, AHH23, AM21b, ACV23, ANR21, ADS20, BP22, BD22, BH25, CMV23, CNT25, CC21, CCH25, CL20, CPL21, CCA22, CBP^{+}24, DvE20, DDS22, DD23, Don21a, Don21b, DGGBDG21, EGR22, FCT^{+}20, FC20a, FNM^{+}22, FD21, GVS20, GSB24, HMPS24, HM20b, JS21, JL23, JCX22, Kav25, KTK21, KOS21a, KLJS25, KK23, LCC^{+}20, LS21a, LML21, LGMK21, LWW22, Liu24, LDAAAB21, LS23, MYK22, MB22a, MCDFSP24, MLP^{+}20, Mey25b, MDCBdGVMD24, MNK21, MFP21, MBH25, NBBL24, NII24, NH24, OMAAORCL23, Ort20, Par21, PHBG21, PD20a, PROGDG24, PLH20, PT20, Pra20a, PSH22, RG20, RA20, RTT23, RL20, SD20, SF20, SBKK21, SYHW20, SCFSAB^{+}25, SCP22, SA21, TKS^{+}23, THvB21, VDY21, WH23, WLHY24, Wan25, YZH23, YKTM20, YMX21, ZM21, ZWC22, ZZMS22, ZM23a, ZC24, ZAZS22]. data [ZDL^{+}20, Ano22a, Ano22b, LC20, LC21, HWP22]. data-based [PROGDG24]. Data-driven [WD21, ZMP24, AHH23, JS21, NH24]. data-drivenness [Kav25]. data-sharing [DGGBDG21]. database [APR25, Asa23b, Bös21, CPN21, CM20, EBB^{+}20, GMRVU24, GC22, HMPS24, YYZ25, YCXY20]. databases [BdOdS21, BRJ21, CKO21, Gus22, KCS25, LKPS25, MMFF25, NII24, OdSO20, ODQ24, PD20a, PD20b, Pur21, SK22, SHB25, Tom21, Wal22, Wij21, ZL20b, dSTE21]. DataCite [GKM^{+}24]. datanomics [ODF20]. dataset [ADD23, BT22, CNA22, Don21b, PDH21, SMA21]. datasets [AL24, CCSX23, GKM^{+}24, KTK21, KS22c, ZY20a]. date [ZWZ22]. dates [Liu21]. Day [COB21, BFG^{+}21, RdNB21]. Day-to-day [COB21]. days [HL21b]. DBLP [CKO21]. De-Westernization [AM21a]. DEA [FZhCC21, Jab25, XlYqZ22]. deaths [FG20]. debate [BdMPH21, OMBP24]. decadal [TD24a]. decade [IHB21, LFV21, MQS^{+}20, RdNB21, Tay23]. decades [GGS22, Li22, RP24a, SdSCdS25, Sha24b, SMBD^{+}24]. decades-long [RP24a]. decay [HTZR24]. decide [GRSFV22]. deciding [Bor20b]. Deciphering [KPM22, BA24]. decision [Bor20b, FFN25, LGXH24]. decisions [LR23, SdVvdR24]. declaration [OMBP24]. decline [BF25]. declining [CtRbQ22, Hua24]. decolonisation [PSH22]. decomposing [KG24, NM25]. decomposition [CLC^{+}24]. deconstruction [WZJW24, WZM^{+}24]. decoupling [HM21]. decrease [LLSH23]. decreasing [WLZ22]. deduplicating [THvB21]. Deep [GVK24, GKD23, KMGS21, ZS20a, AUA^{+}23, CXZ^{+}20, DXL21, GK22, HAL25, CLY^{+}22, JJY21, KYPC20, KYL20, KDZ21, LZD^{+}24, LXZ^{+}20, MLD21, PLH20, RAL22, TA22, WYT21, YZWH24, YZZT24, ZDL^{+}20, ZDLR21]. deep-learning [MLD21]. default [CGA21]. defined [FSEJF^{+}20]. Defining [FN23]. definition [Bor22, FGB25]. definitions [ZWSC20]. Degree [LLLL24, YZWH24]. degrees [Bat24b]. delay [PS22]. delays [ABKGBR25, CMZ^{+}20a, CMZ^{+}20b, DD23, XAA21]. delineation [Lie20]. demand [CL21b, YLX25]. demanding [BPH21]. DEMATEL [GB24, OJKY23]. demise [KSP24]. Demonstrating [BD21, CCC22, KCC21]. demonstration [KP23b, PLH20]. denialism [Rus23, TSAMRG23]. density [Ste21, XDW22, DBR21]. departmental [MHK22]. departments [AHH23, BdOdS21, CC21, DT20, MSV23, RNAH21, GG25]. departure [ZAZS22]. dependence [RNB22, RP25]. dependency [HCS25]. dependent [HB22b]. Depicting [IVCE20]. depreciation [LGMK21]. depth [AA22, BVvEW23, ZCL20]. Derek [Ano20a, Ano21a, Ano23, GDRPR23, Sug23, vR21a]. derived [Bat24b, YJJ25]. dermatology [ALM^{+}23]. described [PR23]. description [HMPS24, MB20b, MRF21]. descriptions [ZW25, dS21c]. descriptive [JKK24]. desensitization [ZZ24]. deserve [WLLW25a]. deserved [JXZW23]. Design [VHM24, Can24, CD22, Mey25b, SB25, ZY20b, ZGM^{+}23]. despite [Nuz20]. destabilization [LC22]. Destructive [SB20]. detect [HB21b, JHVC21, TSA24]. Detecting [Aba21, CCH^{+}24, LYS22, TNS20, WG22, BFG^{+}21, Cop20d, WYL21, XLY21]. detection [GdMDD24, HS22a, LCW^{+}20, LGD21, LCT^{+}20, Naz24b, VCD^{+}23, WLS22, ZY20b]. Determinants [Asa24, Gu21, RA20, DHSS20, Lee24, Plo25b, SHL22c, SHL22b, YN23, dFBGCGASG24]. determine [dSDE21]. Determining [Sha24a, YX21, ADG24, Hir22, LWW24]. developed [CRC20, LRSB21, WSL23]. Developing [JJY21, VDY21, ES23, JME20, KW24, KR22a, MQS^{+}20, Naz24a, SSK21]. Development [AJLA21, Bar25, SSZ22, VHM24, ZP20, ADS20, CPTÁ22, CxGdH^{+}24, CdSK21, DLL21, HS22a, HL21b, HdSV23, HGZ21, JL25, KT21, LCY21, LCY22, LSFLZLLV20, LT20, MB20a, MBH25, OCKY20, OJKY23, Ony20, Poh20, Pra20c, SHL22a, Vie24, WG24]. developmental [YA21]. developments [CFW21, FMB22, KY20, LSY21]. deviance [AHH21]. deviations [KG24]. devising [dSN23]. devoted [HBO^{+}20]. Diachronic [PY24, PY25, YP24, DFG21, LMC22, LZ25b, SKS^{+}23, WWS^{+}23b, WH24, Wan24a, XL20, Xie20, YWW23, mYY22]. Diagnosing [Hua24]. diagnosis [HdSV23]. Diaspora [DDF20]. Did [KZV20, Ali22]. differ [Bor20b, GDdZ22, SSL22]. difference [Ali24, Nuz20, SPA20, WSC^{+}24]. difference-in-difference [Ali24]. differences [AAA^{+}23, CT20, CLC23, FBDM25, GDG^{+}23, GTH24, KSP24, LXY23, LRW25, MTD^{+}23, MS21a, SSZ22, SM20, THJ^{+}23, THJ^{+}24, THG23, VP21b, Wal22, WKC24, YWZ^{+}23, ZSG21, ZSS22, MS22a]. Different [NB22, SL22, ZM24, AKW23, CMV23, CG24, CKO21, CRBRGS21, DDH21, Ela25, FWDQ20, FGB25, HL21a, HTGW20, LL20, LCCH24, LHW21, LH22a, LS23, MBF23, Mor19, MBH25, PD20b, See20, SSL22, Sun23, WPS22, Wij21, XLL25, ZWZ^{+}19, ZWX22, ZSSH22, ZXL^{+}24]. differential [YXYQ22]. differentials [JP22]. differentiated [DGRÁTS23]. differs [MKWB^{+}23]. Diffusion [LHCH20, CLL22, HYZ23, JCZ22, LHC25, LCW21, MRC22, PHBG21, SHZ22, SZY24, XDW22, YH24, YWY^{+}20, YS20, YY21, YXYQ22, ZLL21]. Digital [Cop20b, CF23, FBM21, Tom21, GJ24, HWC22, RTT23, WG24]. dilemma [DDU24, FFN25, SMM^{+}20]. dilute [Zha21b]. dimension [CxGdH^{+}24]. dimensional [FH22, GZSC22a, GYWZ21, HWC22, LHF20, LP20, MY25, TFWC24, YWZ^{+}24, ZS21]. dimensionality [Pra20b]. Dimensionless [Pra21b]. Dimensions [GS24, MMTOMLC21a, FYY20, KFLS20, KSS21, ODQ24, PMC21, TSVV23, MMTOMLC21b, SVY22, SSM21, SS22, Tei24a]. diplomacy [RAG25]. dipping [Asa23a]. direct [SA24]. directed [CLC^{+}24, DLL21]. direction [ZLL21]. directions [CD22, MPMR21, OCKY20, MPR21]. Dirichlet [AKD23, BPL25, GJ24, Han20]. dis [TTDH25]. dis- [TTDH25]. disadvantage [Öz24]. disagreement [IHLP21]. disambiguate [ACV23]. Disambiguation [MS21e, DvE20, KOS21a, PJ24, PMC21, SGA22, WQ21, WQ22, YMD20, MJRR25]. Disciplinary [Ano22a, AG23, KG24, BSS24, CG22, DGGD22, FCW22, Gus22, Hao24, KKK^{+}25, LZ23, Mil20, PY24, UA20, VC20, WH24, Wan24a, WH25, YXQG21, YWL25, YXYQ22, ZZMS22]. discipline [CWL^{+}20, DGSVDG20, Gon23a, MB20a, PS20, SR24, SA21, YWZ^{+}23, YJJ25, ZAZS22, Zho21]. disciplines [AAD24a, AA25, Bir21, CN24, Coc20, Gon23a, GLPN25, GCHT20, LKPS25, LSBC21, LZ23, LS23, LUH21b, MS22c, MQS^{+}20, MKWB^{+}23, Mor19, Nis23a, Nis23b, Öz24, Pin23, QZS^{+}24, RJ20a, RBC21, SLK^{+}23, Sch24b, SDC22, Sta21, TDM24, VCJ22, WWH21, WCH^{+}20, LUH21a]. disclosing [CZFC24]. disconnected [ATCS20]. discontinued [GCMJ25]. discourse [HGMÁW23, LCB20, WKC22]. discoveries [FMF20, PS22, WYT21, Zha23b]. Discovering [CCH20, HM20a, KLJS25, SSdCRD22, ZLY21, HMW24, WZL22b]. Discovery [SSC^{+}23, AKD23, COB21, GS22a, HAL25, KH21, Sma25, WH23]. discrepancies [ARR23]. discrimination [Wal22]. discriminative [CWZ22]. discriminatory [MNM^{+}21]. discuss [AGWG^{+}23]. discussion [AJBB22, ANR21, KHT22, WZL22a, ZXL^{+}24]. disease [GVS20, KKYR24, MPS22a, WZY21]. diseases [CDZ^{+}20, ZZL^{+}20]. Disentangling [CLMM21, Par24, Sch25b, CF22, OMBP24]. Disparities [OdSO20, AOM24, Che24, GDdZ25, LSY20, PL22b, PVA24, The20b, UBROGS23]. disparity [HYY25, Mut22, RNAH21, SDC22]. Dispersion [SG21, FD21]. displacement [SMG^{+}24]. disrupt [WXD24]. disruption [BDTC20, DZ23, JL23, LB24a, LB24b, LGR^{+}21, SLR^{+}23, WZZ23, YGW^{+}24, YLM^{+}24b, ZCW21, ZZW24, ZXL22]. Disruptive [BDTC20, BWE24, CGDS24, CXZM25, Fas21c, LLH22, LWW22, OB21, WLZ^{+}23, YK23, ZLM^{+}23, ZZW24, ZWL25a]. disruptiveness [BF25]. dissemination [AAB^{+}24, MBL21, YSWC24, YNO22]. dissension [Cor22]. dissertation [AAB^{+}24, Che24, DFG21, RBV25]. dissertations [Hao24]. dissimilarity [ZJ22]. distance [MJRR25, Sha24a, TLM22, ZLY21, dMGDD24]. distinct [FZC20]. distinguished [lYnLyO^{+}20]. distributed [Xie21]. Distribution [GMRVU24, CWZ22, ECD24, HHY22, LHX^{+}24, Sta21, ZSL24]. Distributional [GS24]. distributions [Bor24, Che23, CL20, Cot22, GHGMSM21, Sch20a, Sch22, YGW^{+}24]. disturbances [GGSA20]. diverge [Pin23]. divergence [BCG22]. Divergent [Liu24]. diverse [HJT21, HNM^{+}24, XLA23]. diversification [JME20, ZSX24]. diversified [XCT20]. diversify [KJL23]. diversifying [ADD24]. Diversity [Mut22, WZG21, ZCLW20, BGB25, CdSS25, DX20, DSL21, GDG^{+}20, GDdZ22, HYY25, JJ23, KN21, KSS20a, KZH25, LGXH24, MFM20, NLS24, RYS24, TTDH25, WGLN23, WLLL20, YS24, YWZ^{+}24, YLM^{+}24b, ZLL21, Zie22]. diversity-based [YS24]. divide [CFA^{+}23, DD21a]. divided [DG24, WHÅ24]. divides [vD21]. division [CH20, CDQ25]. Do [ADD24, AAD24a, AKW23, BBP21, BPLIdMA^{+}20, BT25, BDPP23, CTH25, CY25, CLR25, Fil21, GMPRSFV25, GZOC24, HHB^{+}20, IHLP21, JXZW23, JDG21, KDS21, KVZ20, Par25, Pin23, SM22a, SO21, SdVvdR24, SvWC21, SZP^{+}24, WXZ22, WLLW25a, XAA21, XDL22, XLA23, ZZL^{+}20, ZHH23b, AYS24, Ans23b, BdCM21, Cot22, GWLY24, GCH^{+}22, Gub24, HZC21, Hug24, JI22, KJL23, KSAT24, LB24b, LC20, NBBL24, Sch24c, See20, SSL22, TKM^{+}23, TDM24, VEGSMMC22, VCV22, Wad20, WL24, WZ25]. Doc2vec [YKS22, GJ24, ZJY22]. Doc2vec-based [YKS22]. Docampo [Pra23a]. doctoral [AAB^{+}24, BKO25, CMK24, DFG21, FGCGFC20, GB24, Hao24, HGDRM20, LCD20, MRBSFSJ24, PWZ25, RJ20a, SOdLGB21, SJBBR^{+}23, WGW21, XS25, YA21]. Document [ZLZ^{+}23, ASSB20, AL24, BO20, CPLS20, Ela25, HB22b, HZCM23, cHhC24, KYPC20, LZL20a, MMFF25, OdSO20, PXY24, RGVQTS^{+}24, SBP20, Sin22, TZZZ23, Thi20, YCXY20, YMLL23]. document-structure-based [LZL20a]. document-type [MMFF25]. Documentary [SLMCSV21]. Documents [Ano24, Bös21, CXZ^{+}20, LAT22, MFM20, QZ20, RAL22, ZMLZ24]. Dodson [SdSFB22]. Does [ADD20, Asa23a, Che23, CKT21, DM25, DMGBG20, Fas21a, GSOW20, GdOCRM20, HWX22, KBS20, KZZ^{+}20, Kua23, LGR^{+}21, MTD^{+}23, NM25, PSKS22, SMM^{+}20, SFC^{+}23, TWH^{+}24, TA20, THL23, VC20, WLLL20, WXD24, XDLZ23, YLM24a, ZM21, ZXL20, BSFS^{+}22, Dip21, FDM24, KSS20a, KBZT21, LCW^{+}20, Lax23b, Lin21a, MM24, Mal22b, Sma24, SHZ22, TSVV23, WZNL23, WSRM23, YLf25]. DOI [CCA22]. DOIs [MFP21]. Domain [ZZC22, AKYH21, BO20, CGPR24, CCSX23, CCH25, GRBAM22, Liu24, OM21, WZJW24, WRT20, YS20, ZGM^{+}23]. domain-specific [CCSX23]. domains [CXW^{+}24, HGZ21, LCCH24, OdSO20, SM23, UNK23, Wij21]. domestic [AD20, AAD22, TZ24]. dominant [LBW^{+}20]. domino [LAT22]. Don’t [Ell23]. door [ŞKKT25]. doping [KSP24]. Doppelgängers [RB25]. DORA [OMBP24]. dormant [HYSY23]. dose [SCU21]. Dot [OM21]. Dot-science [OM21]. double [Asa23a, CTH25, GKDB25, XS25, ZM23a]. double-blind [GKDB25]. double-edged [ZM23a]. doubt [JDG^{+}20]. download [GCH^{+}22]. downloaded [CDZZ20]. downloads [DDH21, FRRBDD23, OA24]. Dr. [AM22, LZ24a, LYZ23, Naz24a, RJ20b]. drain [SA21, TP20, VPSF23]. Draw [CLD22]. Drawing [Lie20]. drinking [TPVVPA21]. drive [OMFJSC24]. Driven [Ano22b, AS24, AHH23, Coc21, DDVV20, HWP22, JS21, NH24, ON20, dFVFA^{+}24, WD21, ZMP24]. drivenness [Kav25]. Drivers [BGWZ20, KBSS23]. drives [LWW24]. driving [FFL21, YNO22, YLM^{+}24b, ZYZW23, ZM23b]. dropout [GCMJ25, XZD19]. drug [SvWC21, WZY21]. drugs [AKH22]. dual [CSJW22, CxGdH^{+}24]. dual-layer [CSJW22]. due [ARR21, CCA22, FFRRPR^{+}24, GMPRSFV25, LDDL24]. dumpsites [OM21]. duplicate [KKOL20]. duplication [Day22]. duration [BMA21, HLS23]. during [BPH21, BTP24, CFW21, DR20, KTB22, LRSB21, LGE^{+}23, LCD20, Naz25, SOdLGB21, TTB22, WL21a, YLW20, ZAF20]. Dynamic [HCN21, LHC25, CCH^{+}24, FZhCC21, GHW22, HBN^{+}21, JMZ^{+}23, LCT^{+}20, PT20, RYS24, TMZ24, WGZ22, YLC^{+}20]. Dynamics [CKMY20, RWL23, CGJ22, CSJW22, CLZ21, Coc20, GKK23, HhC20, KL22, LHW20, LGD21, LR25, LZ23, MMGP25, MMB23, OTH22, PFS^{+}20, RKMG20, RP24b, SK21, Söd23, SC20b, YWZ^{+}23, YXDL24, YXYQ22, Zha21a].

E- [Lec21]. E-learning [BLL21]. e-mails [LBP22]. EAEU [SMM^{+}20]. Early [KYL20, LCD20, MYX22, WH23, ZY20b, Cop20d, GSOW20, GCMJ25, HB21b, HWX22, Lee24, MCDFSP24, MMG23, Mey25b, Par25, ZY20c, ZCW21, ZW24a, ZDLR21, Zhu25]. early-career [Lee24, Par25]. earthers [ZW25]. easily [LZ24b]. easing [RJ20a]. East [Cho24, EOL23, EO24a, EO24b, LL21, SB21]. Eastern [CF23, MMB23, Jok20]. easy [SRL24]. Eco [LZL20b, HC22]. Eco-system [LZL20b]. ecologists [AGWG^{+}23]. ecology [GGSA20, HLS23, RKMG20, SdSFB22, dSMNdV^{+}21]. econometric [Yur22]. economic [Azm22, FFL21, GXC25, GAACBA24, JLL21, JS21, KT21, LWW25, MS22b, MNK21, Ony20, PT20, PT24, PSS23, RNB22]. economics [Ant20, BP22, BW20a, BCT22, BW19, BW22, Can20, DM25, DGRÁTS23, DMA^{+}20, FNK23, Fas21a, Fas21b, Kos23, LSY20, LR21, LA20, ODF20, PY25, Reh20, RP20, WG20b, WB21, WB22, WLLL20, YGO22, Yur20, ZY20a, ZM21, ZM23a, BH24, BW20c, BBB^{+}24]. economies [AdAFR24, AJBB22, CTH22, ES23, GG22, MDCBdGVMD24, Sha24a]. economists [FG20, GSOW20, SBF23, vD21]. economy [BSFS^{+}22, WG24]. Econophysics [SK21]. ecosystem [CPL21, GVS20, HSE23, KA20, PROG21, XHQZ20, vdP24]. edge [Pra20a]. edged [ZM23a]. edit [MJRR25]. editing [Cho24, LZH20, Tei24a]. editor [ABKGBR25, GRSFV21a, Lax23a, Nuz21, Pra20a, Pra20b, Pra20c, Pra21a, Pra21b, Pra21c, Pra22, Pra23b, WSRM23, FC21, PM20, Pra23a]. Editorial [CL21c, WW25, Zha23a, AS23, BMW24, Ela22, GdM20, Lin20, RFMMPJ22, Tut24, WLLL20, XAA21, ZZMS22, ZMLZ24, dMGDD24]. editormetrics [SM22b]. editors [MFF^{+}20, Lin20]. Editors-in-Chief [Lin20]. educated [SKBSC20]. Education [TMNH^{+}24, UNK23, YS24, AlYT21, AGL22, Avd21, BJ21, BL20b, BCM20, BKO25, FMB22, HC22, HFS22, LCD20, LCP22, Mar24, MRM20, PFPCMS20, RPCF20, SJBBR^{+}23, VDY21, VC20, WCH^{+}20, WZZZ20, ZP20, RSF20]. Educational [KW24, GG22]. Effect [FYY21, FM21, PROG21, ZW21c, ADD20, AAD22, AAD24b, dOSAL21, AGB23, Asa20a, CXD21, Cho24, CKT21, CLRMR22, Dej25, DSL21, DGRÁTS23, DGSVDG20, FGB25, GZOC24, GAACBA24, GDdZ25, HYZ23, HMZ21, JS21, JXY^{+}23a, JXY^{+}23b, LS20b, LAT22, LLSH23, MMGC^{+}24, MWH20, MJYS23, OFA23, SdSFB22, SP20, Ste22, WLZ^{+}23, WGLN23, WG24, XZD19, YY21, Tei21, YS23]. Effecting [ZZZZ20]. effective [HPH21, MHD21b, WP22]. effectiveness [BVvEW25, VMF22, ZWZ^{+}19, ZWX22]. Effects [Fuk22a, Fuk22b, Lin21b, ADH23, BSPR20, CL21b, CSL22, EGP22, FFJ21, GXC25, Hüc22, JC22, KMM22, MMF20, MBF23, Min25, Pea21, PSS23, SHF24, STAÁMR24, SMQL20, SAK22, TG24, VRVH21, Wad24, Wal22, ZWD20, ZBG^{+}24, ZHH23b]. Efficiency [BJ21, Can20, CC21, FZhCC21, JSLJ22, KZZ^{+}20, RA20, SV20, WZZZ20]. efficient [CLS^{+}20, PMGÁC23]. effort [JMP20]. efforts [HB20]. ego [MMG21]. ego-network [MMG21]. egocentric [YCYZ24]. Egypt [Ali21b, Ali21a]. Egyptian [Ali24, Ali25, OA24]. eight [GLPN25, San21a]. eighteenth [dDM25]. Ekhad [HCZEP20]. elasticity [Asa24]. electric [QZ20]. Electronic [LHC21, CD22, FFJ21, Mal22a, Mal22b]. electronics [Bre21]. electrophoresis [CdAdFC21]. elements [TDM24, WWC22, WFSZ24]. Elevating [WJZ24]. eleven [ZYPC20]. elite [AlYT21, CDE23, DT20]. elites [JYT21, LZY22]. EM’-index [BTD21]. emails [AM22, dSAKT20]. embargo [LBP22]. embeddedness [LMG20, LZ21]. Embedding [SM24a, CCZ^{+}20, GPYR^{+}20, HAT23, HCL22, cHhC24, JCX22, LT20, PLT22, TSA24]. embedding-enhanced [JCX22]. embeddings [CPLS20, KYPC20, LXS22, SGA22, Thi20]. emergence [CBA24, KSS21, LP20, MGA22, RSPC20]. emergencies [ZZS^{+}20]. Emerging [PJKK24, AJBB22, AED20, BSFS^{+}22, CTH22, CPL21, CR24, GG22, JPS21, LPKG20, OTH22, PFPCMS20, RBB^{+}24, RFG^{+}24, SAK22, SLL23b, SC20b, WH23, WLS22, XHA^{+}20, YRP21, YZWH24, ZDL^{+}20, ZDLR21, ZL23c]. eminence [HB23]. eminent [Ben24]. Emirates [BL20b]. emissions [dGIL21]. emotion [Ben25]. emotionality [ACT23]. emotions [WL22]. Empirical [GS24, MMF23, Pra23c, ADD23, Azm22, Bor24, CC21, CRBRGS21, Coc20, FESRMB21, FLDD20, FGB25, GTH24, IHB21, JLL21, JL23, MYK22, MTD^{+}23, MPS22b, SHB25, SSB22, SZK20, Wad24, WGW21, WSC^{+}24, WQ22, YLM24a, ZXL^{+}24, ZDLR21, dFBGCGASG24, XCML25]. employ [AKW23]. Employee [Zha21a]. employees [LS21a, LZ21]. employing [AQK^{+}22, ZZW^{+}22a, ZZW^{+}22b]. Enabled [SMS23, LLL25, SSCK21]. enablers [AYRG23]. enables [PJ24]. encoders [ZNNA20]. Encoding [GMB22]. encompassing [ADS20]. ends [HM21, Hla25]. Energy [MML^{+}20, CLL22, CZX24, DG21, Lin21b, QLW25, RLZ^{+}24, SH20]. engage [KKK^{+}25]. engagement [FCW22, MYX22, MBR^{+}25, RBV25, TDM24]. Engels [KKS21]. engineering [BOPGOP25, DFRS20, GMRVU24, PA21, YP24, Tom21]. engines [KY20, OM20]. England [Pin23]. English [DWG23, KMM22, Mu21, Pea20, YMW^{+}20]. English-language [YMW^{+}20]. enhance [MNM^{+}21, ON20, Par25]. Enhanced [ZY22, GHW22, JCX22, KR21, LTXZ24, SSS^{+}24, WLL25, ZC24, ZZ24]. enhances [YWN25]. Enhancing [DZ23, FAY25, MZD22, MBR^{+}25, PLT22, ZZZ22, ZYZZ25]. enriched [MZL20]. ensemble [ABAAN^{+}23]. Ensembling [Jia25]. entirely [SMT^{+}23]. entitie [ZW21a]. Entities [Ano24, KKOL20, LYS22, MS21e, NKY^{+}24, SM24a, WZL22c, YWB21, YWZ^{+}24, ZMP24, ZMLZ24]. Entity [MLK22, CR24, Mar20, MJRR25, SLY22, YWB22]. entity-linking [MJRR25]. entitymetrics [SLY22, YWB21, YWB22]. entrepreneur [EE24]. entrepreneurship [CLMM21]. entropy [BGB25, EE24, LGLM20, Sin22, ZMK22]. entropy-based [BGB25, EE24, LGLM20]. envelope [San20]. envelopment [CC21, EGR22]. environment [AJB^{+}21, DD23, OVS25, PG24]. environment-based [DD23]. environmental [Coc21, GMRVU24, ROB21, WZZZ20, ZBG^{+}24]. environments [dMGC23]. envision [CPL21]. epidemic [CP21]. epidemiology [HBSW24]. Epistemic [NF22, Par25, Pol20]. Epistemological [CPFVPA25]. EPO [Wad20]. Eponyms [MM25, SGS22]. equal [Nuz20, WLBC20, WAM^{+}25]. equality [BSPR20, SSZ22]. equalizer [KSP24]. equally [Ste22]. equation [BERD21, MBL21]. equations [Moo21, TNS21]. equilibrium [dLBdAPUdLB^{+}25, Zha21a]. equipment [LCY21, LCY22]. Equivalent [BB23]. equivalents [SDRS23]. Era [Ano22a, Dha22, SH20, ZZMS22]. ERC [PROG21]. Érdi [Sch20b]. erratum [MA21]. error [LCW^{+}20]. Errors [RGVQTS^{+}24, SdSCdS25, CCA22, GMPRSFV25, MMFF25]. Essential [SYK22]. esteem [SGG20]. estimate [GCS25, KDS21, LP23, MN24]. estimates [FD21, FBDM25]. estimating [CYK22, ZHH23b]. estimation [For23, Mut22, Ste21, ZMK22]. ETD [KBDS23]. Ethics [SYK22, DV20, LRP21, Pol21, Tei20, dSVN24]. ethnic [DSL21]. ethnicity [PB24]. ETTs [SLL23b]. EU [ACV23, CL21a, Fog23, MJ22, MJ23, Mor19, RKCA23, SMM^{+}20, ZP20]. EU-EAEU [SMM^{+}20]. Euclidean [Hal20b]. Europe [FFRRPR^{+}24, MLS21, MKWB^{+}23, VCV22]. Europe-wide [VCV22]. European [AlYT21, AAHY20, BMM22, BCM20, CF23, DDL23, DMR20, GEFJ24, GDM22, HPL21, Jok20, KHC^{+}25, LDMVS21, MCSOAA21, MMB23, MBR^{+}25, PL22a, PSW24, SB25, UA20, Vie23, Tei24a]. Eurozone [MDCBdGVMD24]. evaluate [AMIK23, XDLZ23, ZC21]. Evaluating [AlYT21, BSPR20, GCT^{+}20, HZW22, KSS20b, LPKG20, OMAAORCL23, RSPC20, SZY24, WZZ23, WC20, ZZ20a, Zho24, CFI22, FBDM25, KOS21a, LP20, LWG21, Liu23, PW20, PKA24, YF24, ZXL22, BR25]. Evaluation [Ano24, CMZ^{+}20b, Lyu21, MBD25b, RPCF20, SdL21, Vin21, ZLL22, ASIN24, AJBB22, AUM21a, AUM21b, Ali24, Ali25, AED20, BVvEW23, BERD21, BDPP23, BW20b, BM20a, CSC23, CLWlY24, DV20, DV22, DS25, DLC21, EBB^{+}20, FFN25, FLC22, FZL^{+}20a, FZL^{+}20b, GG25, GX22, HHKZ20, Hug24, KZH25, LMH22, LL24b, LdS25, MK23, MBH25, ÖT24, Par21, PJKK24, PSA22, QJI^{+}22, ŞKKT25, SA24, VP21b, WR21, WJ22, WZL22c, Wei20, WFSZ24, YY22a, ZMLZ24, ZNW22, CMZ^{+}20a, MBD25a]. evaluations [CKT21, SMQL20]. evaluative [DWG23]. evaluative- [DWG23]. Evans [BDTC20]. event [FVLA20]. events [LFV21, VFL^{+}21]. Everyone [San21b]. Evidence [AAAA23, uHBKK20, uHBKK21, HZ22, AAA^{+}23, ABKGBR25, BF22, BCG22, BW22, CGJM24, CL20, CM20, DDL23, DLL21, EGP22, FWH^{+}23, Fuk22a, Fuk22b, HKW24, HL23, HDY^{+}23, Kaj22, Ke23, Kim23, KPY21, Kos23, KKYR24, Li22, LB25, LLH22, LL24a, LGXH24, MYK22, MS21a, MS22a, MMF23, Min25, RA20, SPS20, Sma24, TWH^{+}24, VAMBTS25, VY23, WL21b, WGW21, WZZ24, WZZ25, XYY25, XLL25, XCML25, YWW22, YWN25, YT25, ZM21, ZM24, ZBG^{+}24, ZL23c, BW20a, BT25, CXD21, CTH22, DM25, DGRÁTS23, FFL21, Fog23, Hug24, JDG21, LWW25, LGR^{+}21, MH24a, SO21, Sch24b, SZP^{+}24, WLLW25a, WZNL23, WB21, WXZ24, YLf25, ZXL20]. Evidence-based [uHBKK20, uHBKK21, HZ22, Kaj22]. evidenced [AD20]. evidential [ZLL22]. Evolution [AGB23, Coc21, CR24, Han20, KDIR23, LCW21, SFJO20, ZGM^{+}23, ZWL^{+}25b, ZQS22, Ali24, BLL21, BOV21, Cho20, CCMM24, Coc20, CBA24, DFS22, Dri24, EXT^{+}21, GHW22, GKK23, HV21, HAT23, HBC25, HCL22, JHVC21, KS23, KMGS21, LFV21, LYT^{+}20, LC22, LGE^{+}23, LGLM20, LCT^{+}20, LT20, LZF^{+}24, MRF21, MGA22, PLYZ24, PRS^{+}21, QPH^{+}25, RG20, RKMG20, SZN^{+}21, SSK22, TLM22, TKZ^{+}23, TNSF21, WCCC24, WZJW24, WWL24b, Xie21, YCYZ24, YTJ23, YSW^{+}20, ZCW21, ZMP24, ZYWA25, dDM25]. Evolutionary [AJVACC22, WWH21, HLS23, LXL25, SLX21, ZXS22]. evolutions [WYL21]. evolve [WL24]. evolved [IOWN20]. Evolving [ICB24, FMB22, HMZ21, Tay23, ZSL24]. examination [CJT22, FLDD20, KK20a, LMH22, Nuz21, Wad20, YLCY20, dSVN24]. examine [MMG21]. examiner [QPH^{+}25, TM25]. examiners [LTXZ24, Wad20, Wad24]. Examining [BZZ25, HX25, LS23, NKY^{+}24, SYL^{+}24, WWL24a, YNL25, dGIL21, BI24, GDdZ25, MMG23, PVA24, RP24a, ZWD20, ZJ25, JL20, TDM24, WXD24]. example [BH25]. examples [ER24, NB22]. Exaptation [LLH25]. excellence [BMS21b, BKO25, GUG^{+}22, HLS21, KHC^{+}25, LdS25, MF20b, Pin23]. excellencemapping.net [BMS21b]. excellent [WZZ25]. except [RAF25b]. excess [SPA20]. excluded [LS20b]. executives [CGJM24]. exemplary [BWH20a, BWH20b]. exercise [Lax23b, Nuz20, Nuz21, WSRM23]. exercises [CMMO20, DRS20]. exist [TWH^{+}24]. existing [WRT20]. expansion [GMCRV21, LNH24, MGA22, NJCL20, TMNH^{+}24, YWY^{+}20, dS21c]. expectations [VdOS^{+}20]. expected [FLDD20]. expenditure [AÁSTMR24]. expenditures [Gu21]. expensive [LCY21, LCY22]. Experience [Wad24, BBP21, BTP24, CxGdH^{+}24, MLP^{+}20, Rod22]. experiences [YA21]. experiential [Par25]. experiment [SdVvdR24, TSAMT21]. experimental [AS25, CCH25, Coc20, FZL^{+}20a, FZL^{+}20b, Plo25a, SHL22a]. experiments [ASSB20]. Expert [WL21a, BGSRF20, CZQC24, JPS21, SSC21, ZLL22]. expertise [GBP21, NLS24, ZZC22]. explain [HLS21, MPS22b]. Explainable [BJS22]. explained [HMW24]. explaining [SV20]. explanation [HZCM23]. explanations [ER24, Liu23]. explanatory [FBDM25, LHCH20, RAS21]. explication [Sta21]. exploitation [Lin21a]. exploitative [ZWD20]. Exploiting [CCZ^{+}20, LCB20, PMC21, BZL21, WZJ^{+}20]. exploration [Bir21, BWE24, HWNL20, LCB21, Lin21a, WWH21, YT25, YKTM20, lYnLyO^{+}20, dW24]. Exploratory [Bar25, Day22, ATW20, BM21, DDU24, Ela25, LZH20, MGCR24, RYS24, SKS^{+}23, WWZW20, ZWD20, Zha21a]. explore [CGJ22]. Exploring [ANR21, BGB25, Ben24, BMM22, CZX24, CDZZ20, CDQ25, DFG21, DLK20, DCL23, GJ24, GB24, HCS21, HHY22, JCZ22, JCX22, Lee24, LH22a, MR20, NK24, SSB22, TSAMT21, WSC^{+}24, Wan25, XZ23, YTJ23, ZL23b, ZLL21, GC20, HDY^{+}23, KN22, NCD22, SW24, SCU21, WRT20, NC25, WLHY24, ZL23a]. exportation [BPMRSL21]. expression [ZZ24]. Expressions [WH24, WH25]. extensive [FCT^{+}20]. extensively [Ste23]. extent [DGGBDG21, LCW^{+}20, OŠV^{+}24, VCV22]. external [CdSS25, Don21b, JC22, VCJ22, WZZZ20]. extra [MNM^{+}21]. extract [Bös21, WAT23]. Extracting [ZZ24, CXZ^{+}20, LZL20a]. Extraction [Ano24, ABAAN^{+}23, GPYR^{+}20, KP23b, PVB^{+}24, SM24a, WZL22c, WFSZ24, ZZZ22, ZLZ^{+}23, ZMLZ24, ZBKW24, ZYZZ25, ZZC22]. extractive [SMA21]. extreme [ICB24]. extremities [San25]. eye [CNP21].

F [DGN22]. F1000Prime [WWZW20]. Facebook [YMX21, YWHS23]. faceted [MJYS23]. facets [LCB20]. facilitate [XDW22]. Facilitating [QZS^{+}24]. facilities [LL25]. facility [EBB^{+}20]. facing [LDAAAB21]. factions [WPS22]. Factor [KBZT21, AGB23, Bar25, BB22, BT20, GUG^{+}22, KHK21, LS20b, THL23, Urb25, VJCV22, VP21b, MDGR21, OdS20]. Factors [LLP21, TTDK22, ADG24, BF22, CBP^{+}24, jDXjS21, ELR24, FFL21, GB24, HS22b, HTX23, HL24, MFP21, Öz24, PL22a, SOSC24, SMBD^{+}24, SMZY23, TLM22, YX21, ZS21, Zha21b, ZYZW23, Jam25]. factory [Lin24]. Facts [TTB20]. faculties [DT20, WG20b]. faculty [Ali21a, Ali21b, CZQC24, FT21, GG22, JC22, Lam20, LKPS25, RNAH21, SO21, SL22, VC20, YLM24a]. Fair [DS25, BR25, SCFSAB^{+}25]. fairest [Pra23b]. Fairness [GZSC22a]. fake [dS21b]. falsifying [SYHW20]. famed [SGS22]. familiarity [Par25]. families [Pra20a]. family [GP21, HYSY23, Tol24, Wad20]. family-to-family [Wad20]. famous [FG20]. Farabi [KUS22]. Fassin [Mou21a]. fast [Naz24b]. Fauci [AM22]. favor [SOSC24]. favoritism [TYS22]. Feasibility [BGSRF20]. feather [LCS25]. feature [LCC^{+}20, XXL25]. featured [LGE^{+}23, SSC21]. Features [ASIN24, DCL23, KA24, LWL24, LXL25, MLD21, MFMC21, MPC22, MNK21, Mu21, OŠV^{+}24, QJI^{+}22, SSK22, TQA21, WZZ^{+}21, WQ22, XLL23, YZZT24]. featuring [LGD21]. federal [LS21c]. federally [YK23]. Federation [GKK23]. feedback [Pea20]. feel [AYS24]. fees [OFA23]. feet [CLWlY24]. fellow [PB24]. Fellows [BP23, CMT24, Yur22, FCC22, Par24]. fellowship [HB23]. Female [MKWB^{+}23, AKW23, GDG^{+}23, KR22b, LXY23, PMGÁC23, SW24, SM20, XS25]. females [Lax23b, WSRM23]. fertiliser [HPH21, MHD21b]. few [OMFJSC24]. fewer [CLR25, QLW25, SO21]. Field [HAL25, Lec21, ADD23, AS23, AM21b, BLL21, BMS21a, BCT22, CNP21, CTLH24, CT20, Cha21, Che23, Cho21, Coo25, ECD24, EEÁ21, FFH^{+}24, GDDG24, GA21, GDG^{+}23, HB22b, HBN^{+}21, HWX22, KSS20a, LZYR24, Lie20, LLL25, LSY20, LMLR24, MB20b, MB22b, MRF21, MDGR21, Par24, PFPCMS20, PKCOG21, Pra23a, RFG^{+}24, RdNB21, RAG25, SD23a, SD23b, San25, SdS25, SHB25, Sch24a, See20, SDW^{+}20, SG21, SdVvdR24, SLL23b, SC20b, Sun23, SYL^{+}24, TG20, The20b, TP24, TSAMT21, TSDAMRG24, WMCL22, WLLW25b, WUH21, WLLL20, WYL21, WXD24, YRP21, YWHS23, ZMA23, ZC21, ZWL^{+}25b, ZZZZ20]. field- [GDDG24]. field-normalized [HB22b, SHB25, ZC21]. field-specific [Che23]. field-weighted [SdS25]. fields [ADD20, Che24, CG24, Coc21, CR24, ES23, FCT^{+}20, FVLA20, FC20b, IOWN20, KNT20, Kös20a, KPY21, LYT^{+}20, LH22a, LZL20b, MS22c, Mor20, NRGvSTS23, PD20a, PA21, PT20, RFG^{+}24, SHF24, SCDP21, TKS^{+}23, Tas21, TKM^{+}23, YY22a, YWW22, ZJ22, ZMK22, dMGDD24, Rod22]. fifth [Kos22]. filtered [ZS20a]. final [Ste22]. Finally [ACT23, Wra20b]. finance [dOPGdAPU22, OFA23]. financial [AAAA23, Fas21a]. financing [MNM^{+}20]. find [BL24, Gus22]. Finding [DSH^{+}20, LML21, RB25, YRP21, IOWN20, ZZC22]. findings [Cop20d, KR21, RAG25]. finds [DTHM21]. Fine [BBB^{+}24, QD21, ZMP24, ZYWA25, ZZ20a]. Fine-grained [BBB^{+}24, QD21, ZMP24, ZYWA25, ZZ20a]. finer [dS21c]. finer-scale [dS21c]. Finland [PA22]. Finnish [HHB^{+}20, SCDP21]. fire [dSMNdV^{+}21]. Firm [SB25, KPS21, LY20, MYX22, WL21b, Zha21a]. firm-level [KPS21]. firms [HZ21, KJL23, KPS21, MR20]. First [JP22, WRT20, Ali22, BT25, BWE24, DS25, GDdZ25, HL21b, HX25, Lin21b, MMG23, SSZ22, SLTS25, The23, YLJ21, YCXY20, ZL20a]. First-author [JP22]. first-citation [Lin21b]. first-generation [BT25]. fish [JMP20, MHK22]. fisheries [TdSFB20]. fit [BB22, CLS^{+}20]. fitness [ZW25]. fitting [LHX^{+}24]. five [BP22, BW22, GWvZ21, LML21, LR21, MKWB^{+}23, Tay23, Yur20, dSN23]. fixed [XlYqZ22]. fixed-sum [XlYqZ22]. Flagging [LCW^{+}20]. Flanders [JVY22]. flat [ZW25]. flat-earthers [ZW25]. flavour [KNT20]. flexibility [JJ23]. Flexible [HLT25, HM23, SC20a]. flipping [BPLIdMA^{+}20]. flock [LCS25]. flourish [QSYL23]. Flow [HGM21, WLLW25b, WGZ22, lYnLyO^{+}20]. flows [AD20, GSP21, SHH23a, SHH23b]. focal [WGLN23]. focus [Pur24, ZWL^{+}25b]. Focused [DG24, SSZ22]. focusing [LHC25, Nis23a, Nis23b]. fold [CSC23]. follow [AS25, SSBC22]. follow-up [AS25, SSBC22]. followee [YLCY20]. follower [GDRPR23, YLCY20]. food [HGM21, RdNB21]. footprint [CWK21]. force [ZM23b]. forces [GMPRSFV25, Zan23]. Forecasting [Tas21, ZDL^{+}20, DDVV20, HJLZ23, LZF^{+}24, WSZ21]. foreign [Mor19]. forest [CLC23, Hüc23]. forestry [ZZZZ20]. forgotten [LNH24]. form [BWH20a, BWH20b]. formal [DGGBDG21, GSG22a, WVH21, YZH23, YLH24, KN22]. formalism [ER21]. format [LCW^{+}20, Mal22b]. formation [CZX24, CKMY20, MYK22, NF22, Sch24a]. formats [WL24]. formed [HM20b, YLCY20]. formula [SSC^{+}23]. formulaic [ZZ24]. Formulation [AUA^{+}23]. fortification [dS21c]. forums [JDG^{+}20]. FoS [ZMA23]. foster [NM25]. Foundation [WLLW25a, NH24]. foundations [Ant20, CZZN25]. founding [AL21]. Four [Bi23c, BW20a, FVLA20, LRW25, Öz24, SM23, UA20, WWZW20, ZS20b, dSN23, Bi23a, Bi23b]. four-tier [dSN23]. fractional [ÁLGBdMA24, Bat20, GAB21, PSA22, PKA24, PM20, Pra21b, WLLW25b]. Fractionalization [Dem22]. Fragmentation [vdPF25]. framed [SGS22]. Framework [ACV23, CFI22, PSW24, AHH23, BMA21, BZG22, CL21a, CCZL23, FYY20, GLPC20, KA20, KHC^{+}25, KW24, LNS21, LS21a, LLZ25, MY25, PR23, QZF22, VFL^{+}21, VPR22, WQ21, WLS22, WFSZ24, XWGD22, XHQZ20, ZBKW24, ZDLR21]. frameworks [Pin23]. franca [DWG23]. France [CDE23]. Frangopol [Ben22]. Fraud [MCSOAA21, BT20, Day22]. fraudulent [SYHW20]. free [GLPC20, VdOS^{+}20, SMT^{+}23]. Freeloading [RvLR^{+}20]. freely [LML21]. frenemies [Hüc22]. frequency [FZC20, GGS22, HZ22, LM20, Pup21]. frequent [HHB^{+}20]. friends [WW25]. front [YY22b]. frontier [AlYT21, FZhCC21]. fronts [Mar20, WZL22b]. fsQCA [CW25, LWX21]. FT [HKW24]. FT50 [Fas21a, Mou21a, Zha21b]. fuel [LHXS20, ÖT24]. Fuels [MML^{+}20, Lin21b]. full [BM21, LCB21, SF20, SDRS23, WLLW25b]. full-text [BM21, LCB21, SF20]. full-time [SDRS23]. fuller [CM22]. fully [Asa24]. function [BA24, BY23, CNT25, JC23, Jia25, MZD22, Sha24a, ZZW^{+}22a, ZZW^{+}22b]. Functional [DBR21, HM23, RTT23, YLCY24, NBBL24]. functions [BT22]. fundamental [CNA22]. funded [LB25, LWX21, PROGDG24, VKY23, WJZ24, XLL25, YK23, ZX25]. Funding [DM25, LTH20, MN21a, AJLA21, BGAMS23, BCG22, DLL21, EO24a, FYY21, HLS21, HPH21, Hir22, JCX22, KM20, Liu20, LR23, MMF20, MHD21b, MYX22, Mor19, Mor20, MRC22, NCD22, OPS25, PROGDG24, PSW24, PA22, RBC21, She20, SdVvdR24, TSVV23, WJZ24, WLLW25a, XLL25, YK23, ZZL^{+}20, ZCL20, vdP24]. funds [Cha23b, LHX^{+}24, MHD21a]. funnel [FLDD20]. further [HWL23, OG24]. fusing [LLZ21]. Fusion [ZC23, BBB^{+}24, GJ24, Liu24, LYLC23, XXL25, ZWSC20]. future [CLS21, CPL21, CGA21, CDG20, JPS21, Kup22, MPR21, MPMR21, NS23, SFC^{+}23, Tas21, VdOS^{+}20, ZWZ22, dRRJ20]. fuzzy [Can20, CW25, GB24, JPS21, MK23, dOCP21]. fuzzy-set [CW25]. FWCI [SdS25].

Gagolewski [Pra22]. gain [CTH22, SA21, vR21b]. gained [VP21a]. galleries [KSAT24]. game [AdAFR24, GRSFV20a, GRSFV21a, LBS23, Sch20b]. gamification [MBD25a, MBD25b]. gamma [CYK22, Sta21]. gamma/Gompertz/NBD [CYK22]. GAN [CNT25]. GAN-CITE [CNT25]. gap [CXW^{+}24, KR22b, Lax23b, SdLV21b, TD24b, WSRM23, YKTM20, ZO20]. gaps [APR19, FJO20, JA23, JHVC23]. Garfield [GUG^{+}22, MDGR21]. gate [AHRH24]. gatecrashers [Tut24]. Gatekeepers [LZH20, Tut24]. gatekeeping [ŞKKT25]. GDP [Pra25a]. Gender [APR19, CT20, Che24, DMS20, DL23, FWH^{+}23, FJO20, GTH24, GDdZ25, HBSW24, JA23, LSY20, MRBSFSJ24, PS20, PL22b, PVA24, RNAH21, SJBBR^{+}23, SdLV21b, SSZ22, SM20, UBROGS23, WMK21, ZSG21, ZSS22, ADM22, ABKGBR25, BSPR20, CCMM24, DD21a, DGRÁTS23, FBM21, GBH21, GSRG24, GdMDD24, HLS21, HSE23, HdSV23, JP22, KDIR23, KSP24, Kos23, KZZ^{+}20, Kup22, KR22b, LCD20, MTD^{+}23, OPZP25, PMGÁC23, RPL21, SS23, The20b, TD24b, WGA22, WKC24, XS25, YKTM20, ZAZS22, vdBM24, LRW25, SLK^{+}23]. gender-based [PVA24]. gender-biased [CCMM24]. gender-related [KSP24]. gender-responsible [GBH21]. gene [BPW^{+}21, LZH20, WZY21]. genealogy [GP21, HWC22, KBDS23]. general [KMGS21, LHW20]. Generalized [Kos25, For23]. Generalizing [KR21]. generated [LDDL24]. Generating [KCC23, LLC22]. generation [ASSB20, BT25, BZG22, CTLH24, OTH22, SOSC24]. generations [ÁLGBdMA24, OPMRP20]. generative [CPN21, CNT25]. generic [BZG22, Hao24]. Geodiversity [Pur24]. geographic [ADD20, AAD22, AAD24b, PLW23]. Geographical [AS23, FVGBPA23, BI24, KSZ22, MGCR24, Pur24]. geographically [HWZ21, HBO^{+}20]. geography [AAHY20, GXC25, KMGS21, SFC^{+}23, ZYL20]. geohistoriometric [CPFVPA25]. geopolitical [RBBP24, THG23]. geopolitics [ZCSK24]. geoscience [Mit20]. geosciences [MHD21a]. Geothermal [QLW25]. GERGM [CTLH24]. German [DK21b, HLS21, KSS21, KN22, Reh20, SHB25, SGM22, THJ^{+}23, THJ^{+}24, THL23, ZAZS22, vRW25]. German-affiliated [ZAZS22]. German-language [DK21b]. Germany [Ano21d, BG21, BKO25, HJT21, HKN23, HKNF23, KPS21, Sch24b]. get [JXZW23, MM25, Sch24c, VEGSMMC22, WXZ22]. gets [PSKS22]. Getting [MWH20]. ghost [OMCC23]. Gini [BB23]. GIS [HBN^{+}21]. GIS-T [HBN^{+}21]. give [BDPP23]. given [Sma24]. Global [ER23, LDDL24, Li25, LPY22, Old22, RP25, SWL^{+}23, AS24, AHH21, BWLX20, Dej25, Dem22, DT20, DCdA25, Jam25, LXY23, MGA22, MAN22, Pal21, PL22a, RAG25, SD20, SSS^{+}24, SL22, SYK22, TZ24, TBPN^{+}20, ZZS^{+}20, ZT23, ZSX24, Cha23b]. globality [ZYL20]. globally [Cha23b]. gloria [FG20]. go [FH21, SKBSC20]. Goal [SMS23, SSZ22]. goals [HS22a, MBH25, Bar25]. Going [SRL24]. gold [BPLIdMA^{+}20, MBH25, Naz24a, SPS20]. gold-green [SPS20]. Golden [Gra20, GCHT20]. Gompertz [CYK22, TNS21]. gone [BW20a]. gonorrhea [SLMCSV21]. good [DV20, DV22, HY22]. Google [OMFJSC24, ALR24, BGP20, Don21a, Fur20, GKM^{+}24, GTH24, MMTOMLC21a, MMTOMLC21b, MRM20, OMFJ22, San20, SSL22, Tei24a, TTBA20, dS21c, dCPT22]. governance [TKZ^{+}23]. Government [NCD22, ZM23b, BKT20, DLL21, MYX22, PSKS22, Pur25, XCT20, ZCL20]. governments [PHBG21]. governors [AL20]. Graduate [KPM22, Yur23]. graduate-level [Yur23]. graduates [RBL20]. grained [BBB^{+}24, QD21, ZMP24, ZYWA25, ZZ20a]. Granger [LDA22, TWH^{+}24]. grant [Hug24, LB25, NSM23, SdVvdR24, YN23]. grantee [PROG21]. granting [FT21, SO21]. grants [DM25, TQA21, WMK21]. graph [HCS25, HWW^{+}24, HTZR24, JJPC20, KLJS25, KBDS23, LFV21, LZ25a, LTXZ24, LXS22, MS21e, SGA22, VPDAG22, WWS23a, ZLZ^{+}23]. graphene [SZZ23, WYL21]. Graphical [BS23, GSH24, Hal20b]. graphs [GS22a, LCC^{+}20]. gravity [Avd21]. Graz [RS22b]. greater [BS23, ZNW22]. greatest [Ell23, dOPGdAPU22]. Greek [SKBSC20]. Green [NS23, BPMRSL21, LL24b, SPS20, ZBG^{+}24]. grey [BKT20, BKT22, EE24, KTB22, SSSd22]. GRIs [Kim23]. grounds [EGV22]. group [LR21, YY22a, ZY20a]. groupings [NBBL24]. groups [ATCS20, BMS21b, RB25]. Growing [CWK21, MSJ23, OMCC23, Poh20]. Growth [SK21, Azm22, Liu23, NS23, NRGvSTS23, Naz24a, ÖT24, PT20, PT24, PSS23, SD22, Sta21, UA21, VV21, VdOS^{+}20, WZZ25]. guide [BB22, EO24b]. guideline [Sch21a, THL23]. guideline-based [Sch21a]. guidelines [EBB^{+}20, GLPN25, OMAAORCL23, See20].

H [Fur20]. Habermas [dCFdCA^{+}23]. Half [BM20b]. hand [Pra20a]. handful [Sch21b]. Handling [DDU24, HB22b]. happens [MMP21]. hard [LZ23, MZE20]. harnessing [Tom24]. Harvard [Lam20]. hashtag [AKYH21]. hate [TNSF21]. HCR [CDE23, Pra20d]. head [HWW^{+}24, MWH20]. headings [CLC23, SDC22]. Health [PMB^{+}23, BD22, FJO20, GC20, GTH24, HB20, HBC25, KVZ22, KKYR24, Lam20, MF20a, NB20, Seb23a, VPC^{+}23, WH23, ZZS^{+}20, PSH22]. heard [BdMPH21]. heart [CPTÁ22]. heat [Ste21, ZSL24]. heavy [KNT20]. hedging [YWW23]. heeding [Pur25]. Helix [JSLJ22, AdAFR24, Coo25, EL25, HLR21, Mey25a]. help [KSAT24, XDLZ23]. helps [FBM21]. Henk [DGN22]. here [TDM24]. hesitant [Can20]. Hess [Wra20a]. Heterogeneity [LY20, BCM20]. Heterogeneous [SLH24, ZBG^{+}24, CCZ^{+}20, HTZR24, LH24, LXS22, YLC^{+}20, ZW21a]. heuristic [DDL23]. heuristics [WQ21, WAT23]. Heyard [MHD21b]. HF [Fas20a, Fas20b]. HF-rating [Fas20a, Fas20b]. Hidden [Ans23a, LB24a, Li25, ML24]. hiding [ZW21c]. Hierarchical [MN24, KCC23, YLM^{+}24b, ZLZ^{+}23]. hierarchy [Ben20, HB23, JL20]. high [CH20, CW25, Cha21, CNA22, DWM21, FZC20, GWvZ21, Hen19, Hen20, HM20b, Ioa23, KKK^{+}25, KTB22, LWX21, MH24a, MR20, SS23, Seb23a, TDWW24, WZZ24, Wan24a]. high-frequency [FZC20]. high-impact [GWvZ21, HM20b, SS23, Seb23a]. high-profile [Hen19, Hen20]. high-quality [CH20, CNA22, DWM21, TDWW24]. high-speed [MH24a]. high-value [WZZ24]. Higher [NSM23, SLK^{+}23, TMNH^{+}24, AlYT21, AGL22, Avd21, BJ21, BL20b, BCM20, CZFC24, FC20c, FMB22, GWLY24, GG22, HFS22, JDG21, KBS20, LBP22, LZH^{+}22, LCP22, Mar24, PSKS22, PFPCMS20, TKM^{+}23, VDY21, WZZZ20, XAA21, XCML25, YWX23]. Higher-order [NSM23, FC20c, LZH^{+}22]. highest [ZZL^{+}20]. highly [CWZ21, CDE23, CDZZ20, Che23, Cho21, ECD24, FGB25, KKS21, LS20b, LZYR24, LM20, NLS20, RL20, SKBSC20, TDM24, WZL22a, dSVN24, KK23]. highly-cited [LM20, NLS20, RL20, TDM24, dSVN24]. hijacked [Aba21]. Hildrun [GKML25]. hinder [ADD24]. hiring [FT21]. Historical [BL20a, LYT^{+}20, NS23, KKS21]. histories [Mit20]. history [CSR23, LHW20, LCC^{+}20, RTT23, YGO22, dSMNdV^{+}21, Wra20a]. HNRWalker [YLC^{+}20]. hoax [Kos22]. hold [AAD24a]. holders [YWW22]. holding [Mal22b]. holdings [Mal22a, Mal22b]. holes [WLZ^{+}23]. home [XYY25]. Homeland [DDF20]. homeopathy [The21]. homogeneous [YK23]. Homophily [HFS22, SHF24]. Horizon [PSW24]. hospital [SdL21]. hospital-based [SdL21]. Hospitality [BB22]. hospitals [SdLV21b]. host [XYY25]. Hot [GWvZ21, OSS^{+}23, CKMY20, HH21, Kos20b, SP20]. hotness [ZW24b]. hotspots [HG20, KCC22, SJW21, SJW22]. Huawei [RBBP24]. Hub [CLRMR22, MS23, OA24, RM24]. human [BPW^{+}21, CXD21, DDF20, GCT^{+}20, Gra20, HLS21, San21a, San25]. Humanities [LNH24, RJ20a, AG23, BZG22, CY25, CM20, DABSC20, EGV22, MN21b, Nis23a, Nis23b, SCDP21, Tay20, YS23, dMALIGBM20, CxGdH^{+}24]. humanization [GSG22b]. hundred [DT20]. Hungarian [TMNH^{+}24]. Hungary [CF23]. hunger [Pur24]. hurdles [MFF^{+}20]. hybrid [Asa23a, AMIK23, DWM21, GK22, KFLS20, LGD21, QJI^{+}22, WAT23]. hyper [KNT20]. hyper-authorship [KNT20]. hypergraph [SLH24, Xie21]. hyperprolific [MMGL23, dFVFA^{+}24]. hypothesis [FLDD20]. HypTrails [KSS21]. i10 [dS21c]. i100 [dS21c]. i100-index [dS21c]. i1000 [dS21c]. i1000-index [dS21c]. Iberian [SOdLGB21]. Ibero [BGP20]. ICA [Par24, PB24]. iceberg [PJKK24]. ICT [QQL21, SMS23]. ICT-Enabled [SMS23]. ideas [CBA24, DH23, HGM21, RTT23]. idem [GSG22a]. Identification [HCS25, HCL22, HWL23, KG20, SLL23b, WLMZ24, ASX22, BORROPGV20, Don22, HAL25, HYSY23, HY22, JKY22, KYL20, MZD22, SBKK21, WZJ^{+}20, WMCL22, XHQZ20, YZH23, ZNNA20]. identified [LL20, Pup21, SGG20, WKC22]. identifiers [SdLV21a]. identify [CZZN25, CLC23, DDL23, Mar20, SKBSC20, San24, TSDAMRG24, WPSK25, WZ25, Yam20, YY22b, ZDLR21]. Identifying [AMTSRG21, CPL21, CCA22, Cor22, FNM^{+}22, JPS21, Kös20a, KPY21, LNS21, LGMK21, LWW22, PVP^{+}23, PFS^{+}20, Ste21, WSY21, WJ22, WCCC24, ZY20a, ZYZW23, BdOdS21, BKT22, BWE24, CCZL23, OSS^{+}23, TD24a, WYT21, YJS21]. identities [TSAMT21]. identity [CGPR24, HGMÁW23]. IEEE [TWH^{+}24, Tom21]. if [JTB23, dSV23]. II [LHX^{+}24]. Ilan [BL20a, Hal20a, HP20, LB20, OM20]. illogicality [MS22d]. illustrated [BWH20a, BWH20b]. image [PJ24]. Imaginary [HPH21, MHD21b, MHD21a]. imbalance [SJBBR^{+}23, WGA22]. iMetrics [LYT^{+}20, MB20a]. IMF [EGP22]. immersive [Wan25]. immigrant [BT25]. immunology [WPBH24]. IMO [Yur24]. Impact [ADM22, BCF21, CXZM25, Egg22, FM21, GKDB25, KS22a, KBZT21, KSP24, MFMC21, MDGR21, OdS20, SCDP21, SAR19, STM^{+}24, Vin23, WPS22, ZMA23, Zie22, ADD23, ADG24, AS24, AAAA23, Ali24, ARDA20, AM21b, AGB23, BWLX20, BZL21, BD22, dLBdAPUdLB^{+}25, BB22, BT20, Bi23a, Bi23b, Bi23c, BKT20, BSFS^{+}22, BW19, BW20c, BW20b, BMS21b, BVS25, BH24, CSA20, CMMO20, CSL22, CG24, CCMM24, Cop20a, Cop20b, Cop20d, DABSC20, DS25, Dej25, Dem22, DSL21, DHSS20, DRF21, DLL21, ER23, Emm19, FWDQ20, Fan20, FCY23, FT24, Fas21b, FZhCC21, GZSC22a, GZSC22b, GWLY24, GC20, GAACBA24, GDDG24, GC22, GUG^{+}22, GTH24, GWvZ21, GdOCRM20, Gu21, HAR24, HM20b, HZC21, HZJ22, HHY22, HZW22, Ioa23, Jab25, JYT21, JVY22, Ke23, KTK21, KSAT24, KSZ22, KHK21, Kup22, LS20a, LKPS25, LS20b, LCL23]. impact [LHF23, LW24, LL25, LdGRV^{+}22, LWG21, LZ23, LMSNZG22, MS22b, Mal22b, MR20, MNM^{+}20, MNP21, Mey25a, Mil20, Mor19, Mor20, NLS20, NSM24, NCD22, OdSO20, OPZP25, OMAAORCL23, OŠV^{+}24, Öz24, PWZ25, PS20, PSKS22, PD20a, PT20, PSS23, Pra22, RSF20, RPAVFV21, RAS21, SD20, SSS^{+}24, SdS25, SHB25, SS23, Seb23a, SAS23, SW24, SWL^{+}23, SD22, SRL24, SDRS23, TLM22, TM20, TZ24, TKM^{+}23, Tom23, TA20, THG23, TMNH^{+}24, THL23, TTB22, UKP21, VPSF23, Vie23, VP21a, VP21b, WZL22a, WGC20, Wan24b, WSC^{+}24, WWL24a, WZHW25, WC20, Wei20, WZZZ20, XLL23, XDL22, XCML25, YTZ20, YF24, YXQG21, YXDL24, ZC21, ZM21, ZW21a, Zha21b, ZLM^{+}23, ZW24a, ZWL25a, ZS20a, ZPZ21, ZWZ22, ZYL25, ZJ25, ZHH23a, Tei24b, dMALIGBM20, dCFdCA^{+}23, vR21b, vdP24]. impact-factor [KHK21]. impacted [Cer24b]. Impactful [Zha23b, BFG^{+}21, Ste23]. Impacts [CWLS23, ZL23c, OPS25, RBL20, SL22, SS22, ZSG21, ZM24, ZZ20a, Zho24, ZHH23b, Tei21]. impede [LCW^{+}20]. imperfect [GMPRSFV25]. Imperial [PSH22]. implantology [EGPGGA^{+}23]. Implementation [AMSS^{+}23, DV22, DMR20, DMR22, SVY22, VWR25]. implications [BA22b, DJM22, KR22a, MH24b, MMGL23, SRL20, WL22, ZCS^{+}24, GBH21, SFC^{+}23]. implicit [DS20, DS22, VWR25]. importance [KDS21, KS22b, SdL23, Wal22]. Important [ASX22, WZJ^{+}20, AKH22, HCS25, Jia25, LZ24b, The23, ZWD20]. importation [BPMRSL21]. impossible [Lie20]. impression [OŠV^{+}24]. improve [NHAB21, Ove23, VMF22, XZZ22a, XZZ22b, YLM24a, ZXL20]. Improved [LBW21, BDTC20, HWW^{+}24, ZZW24]. improvement [CG22, KS22a, ZMK22]. improves [YSWC24]. Improving [BGAMS23, GSG22b, SdS25, WCL25]. in-depth [BVvEW23, ZCL20]. in-group [LR21]. in-text [AFH21, IHB21, SF20]. inaccurate [Tei24b]. inactive [HFAR23]. inbound [SZ20]. Incentive [CL21b, MOMGNLG24]. incentives [AAAA23, BL20b, GRSFV21b, SRL20]. inception [MSJ23]. incidence [NT24]. inclined [KR22b]. including [BWH20a, BWH20b]. incomplete [BC21, LZ25a]. Inconsistent [VY23, Cer24b]. incorporates [LMH22]. Incorporating [LS20a, SS24]. incorrect [BPW^{+}21, LCW^{+}20]. increase [BRG^{+}23, TA20, WLLL20, Zha21b, ZHH23a]. Increasing [KK20a, CF23, HZ22]. incremental [WZJW24]. Independent [DAGG24, CPFVPA25, HB22b, VP21b]. index [ALM^{+}23, BORROPGV20, Bi23a, Bi23b, Bi23c, BTD21, BDTC20, CRBRGS21, Cho20, DLK20, ER21, EE24, ELR24, Fas20b, Fur20, GZSC22b, GG22, HIA^{+}20, HHKZ20, HBSW24, HWNL20, HY22, JL23, JSLJ22, KDS21, KS22a, Kos25, LNS21, LP23, LB24a, LB24b, LHF20, LSBC21, LLH22, LGLM20, LWG21, MMN20, MNM^{+}21, MHK22, NLS24, Naz24b, PMHC21, Pra21a, Pra21c, Pra22, RFG^{+}24, SDD20, SMT^{+}23, SdL21, SG21, Sin22, TK21, TD24c, VJCV22, Vin23, WAM^{+}25, YZWH24, YGW^{+}24, YTM21, ZZW24, ZL23c, Zhu25, Zie22, dS21c, DCdA25, HMPS24, LNH24, MJYS23, NLS20, PHP22, YS24, Fas20a]. indexation [OdSO20, ODQ24, SS22]. indexed [CFC22, EO24a, LHW21, NT24, PD20b, SMJ21, SSSd22, SSB22, ZL20a, dCPT22]. indexes [DLK20, Fan23, WSL23]. Indexing [AKH22, BO20, Wij21]. India [GG22, KL22, KKYR24, PS20, PSKS22, Sha24b]. Indian [DSL23, Ela21, KBDS23, RS22a, SPS20, SNK22]. indicator [AMSS^{+}23, CG22, CM22, FDM24, GG25, HB22b, HGMÁW23, Mad20, MSI21, RZS23b, VCV22, WP22, YF24, RZS23a]. indicators [AJBB22, AMS21, dLBdAPUdLB^{+}25, Bar25, CKT21, DD23, DDF20, EGR22, FLC22, FZL^{+}20a, FZL^{+}20b, FDM24, Gub24, HKN23, KY20, KCS25, LB24a, LWW25, LP20, LdS25, LMLR24, MAN22, ON20, Ove23, Pet22, Poh24, Pra20b, Pra20c, Pra21b, SAR19, SGG20, TM20, Urb25, WGC20, YXQG21, ZS21]. indices [DMA^{+}20, FSEJF^{+}20, Hal20b, MFM20, NLS24, SAA21, ZWZ^{+}19, ZWX22]. indirect [ZWD20]. Individual [Plo25b, AHH21, ANR21, Bi23a, Bi23b, Bi23c, DvE20, FBDM25, LR25, LLL25, LC21, PLW23, WZZ^{+}21, ZWZ22, ZW24b, dFBGCGASG24]. Individual-level [Plo25b, AHH21, ANR21, LC21]. individuals [LSBC21]. industrial [KM20, LHC22, LLL22, LZD^{+}24, SAK22]. industrialized [MML^{+}20]. industry [AMS21, BWLX20, BJ21, CLL22, CZX24, CSJW22, Fil21, GLK^{+}23, Hua24, LHXS20, LY20, LGMK21, MOMGNLG24, NCD22, OSS^{+}23, PSKS22, Pur25, QQL21, WL21b, WGZ22, YX21, ZO20, Dha22, GSG22b]. industry-university [LY20]. industry-university-research [CLL22, CZX24, WGZ22]. Inequalities [VV21]. inequality [Gra20, KBZT21]. inequities [AD21]. inequity [SLK^{+}23]. inertia [SZ20]. infer [GdMDD24]. inference [RKCA23, ZHH23a]. inflammatory [GSG22a]. inflate [TM20]. inflation [Nuz21]. Influence [DST23, EEÁ21, ELR24, LW24, SGC21b, SMS23, SZP^{+}24, TG24, AG23, Asa20a, AA22, BSA23, BW22, Che23, DMGBG20, FC20c, GB24, Gon23a, GdM20, GdOCRM20, GMKG20, KZZ^{+}20, LGD21, LB25, LGR^{+}21, MPC22, MMN20, NZL21, PROGDG24, SVY22, SW24, SPA20, TSVV23, TDM24, Vie23, WGW21, WWL24a, XAA21, ZAF20, ZCSK24, ZCLW20, ZXL22, SGC21a]. influences [PFS^{+}20, TTB22, Zbo25]. influencing [BF22, Jam25, Öz24, SMZY23]. influential [HV21, KPY21, LR25, QD21]. inform [DSB22, PSH22]. informal [KN22, YLH24]. informatics [KVZ22]. Information [Ano20c, LHC21, PFPCMS20, Pra23a, RS22b, SD23b, SSS^{+}24, Sma25, VEGSMMC22, WGC20, AS23, ADU22, AMTS24, BBB^{+}24, BMS21a, BC21, BdOdS21, BVvEW23, Bre21, CFM20, Cha21, CXZ^{+}20, Cho21, CP21, Cop20c, DX20, FH23, FGCGFC20, FSEJF^{+}20, GS22a, GA21, HKV20, HhC20, HLSX20, Ibr21, IOWN20, Jok20, LHK21, LCW21, LLC22, Liu20, LTH20, LXL25, LLZ21, LM20, MZD22, MK23, MIY^{+}20, MKI24, Mut22, OdS20, SD23a, SLMCSV21, Tas21, TTDK22, TSAMT21, WZJ^{+}20, WWH21, WLBC20, XZZ22a, XZZ22b, YH24, YCL25, YJJ25, ZZW^{+}22a, ZY22, ZZZ22, ZZW^{+}22b, ZT23, ZMP24, ZYZZ25, ZYWA25, ZS20b, ZPH21, HCZ^{+}22]. information-sharing [CP21]. information-theoretical [Mut22]. informative [CKT21]. Informetrics [BMW24, SK20, CD22, HGMÁW23, ZZMS22, Ano22a, DMG20, GDR22, CN24]. infospheres [BGP20]. infoveillance [AYS24]. infrastructure [AL20, AAD24b]. infrastructures [CRC20, ZWL25a]. inhibit [SHZ22]. initial [Fuk22a, Fuk22b, DR20, Naz25, Öz24, YWW22]. initiative [LdS25, MF20b]. initiatives [DMR20, OMAAORCL23]. Innovation [EL25, FZhCC21, MKG23, RvdSvMK20, Vie23, VdBC20, AdAFR24, BWLX20, dLBdAPUdLB^{+}25, BORROPGV20, BGP20, BVS25, BSPR20, CW24, CXD21, CLL22, CZX24, CL21b, CXZM25, CPL21, Coo25, CM22, FYY21, FESRMB21, GBH21, GSP21, HJLZ23, HS22a, HVGP21, HL23, JSLJ22, KM20, KN21, Kim23, KA20, LLH25, LLL22, LHF23, LLLL24, LLL25, LMG20, LT20, LLZ25, LZ21, MOMGNLG24, MKI24, PR22, Poh21, QZF22, RYS24, SM22a, SM22b, Sch25a, SZ20, SW24, SvWC21, SC20b, VDY21, VdOS^{+}20, WL21b, WWC22, WGLN23, WFW24, WGZ22, XHQZ20, XDW22, YD21, YXDL24, YX21, YP20, YWL25, ZWD20, Zha21a, ZCS^{+}23, ZBG^{+}24, ZZW24, ZWL^{+}25b, vdP24]. innovations [OVS25, WZ25]. Innovative [SMS23, CL20, CBA24, EE24, LY20, LGXH24, ZW21c, ZC24]. Innovativeness [MPMR21, MPR21]. input [Cao20, HL24, Sha24a, ZS21]. input-process-output [HL24]. inputs [LN24]. Insert [TDM24]. insider [TYS22, Tut23]. insight [Gor21, Plo25a, ZX25]. Insightful [Dha22]. Insights [BBP23, Bös21, AJBBA23, BDPP23, MBD25a, MBD25b, MDCBdGVMD24, Plo25b, QPH^{+}25, WWL24a, WCH^{+}20, YWZ^{+}23, YSWC24, YNL25, ZTY^{+}24, DDVV20]. instability [YLW20]. Institute [CF23, dBGP21, RS22b, WGZ22, RS22b]. institutes [Asa21, BKO25, LCP22]. Institution [HLSX20, PS20, ZQML23, MJRR25]. Institutional [Agu20, CL21a, KBSS23, ATC20, BL24, CWLS23, FWDQ20, FZhCC21, HLSX20, JL20, KG24, OMFJSC24, PLW23, PSA22, RBV25, Sha24b, SNK22, WZZ^{+}21, WLLL20, dMGDD24]. institutionalization [LMLMPA22]. institutions [AGL22, Asa20a, BJ21, BL24, BL20b, BCM20, GdOCRM20, GG22, LNS21, Mar24, NLS24, SO21, SDRS23, dGdO^{+}23, WFW24, WG20b, WZZZ20, ZZ23, ZCS^{+}24]. instruct [See20]. instructional [HGB20]. instrument [Söd23]. instruments [DSB22, LDA22]. InsurTech [OVS25]. integrate [Hug24]. Integrated [WMCL22, LZF^{+}24, PYS22, WYL21, YF24, YK23]. integrating [LWL24, LWW22, QYL21, WLL25, Zho24]. integration [DDS22, VC22]. Integrity [SYK22]. Intellectual [AL21, Cho20, GXC25, HBN^{+}21, KT21, Kös20a, KPY21, MPR21, MPMR21, PKCOG21, SB25, SVC20, VBF24, YJJ25]. intelligence [AHH23, AJVACC22, BCT22, CTLH24, CPTÁ22, GLPN25, HJLZ23, HSE23, LHC22, LSY21, LGXH24, RSM24, WD21, ZWZ^{+}19, ZWX22]. intelligence-based [AHH23]. Intelligent [TDWW24, WZM^{+}24]. intensity [FSR21, MY25, WXD24, XCT20, ZL23a, ZL23b]. intensive [Men25]. Intent [GVK24, QWS^{+}23, XXL25]. intents [SLH24]. inter [ARR23, CWLS23, Iva25, Sha24b, SAK22, WSC^{+}24, ZCS^{+}23]. inter- [ZCS^{+}23]. inter-country [Sha24b]. inter-industrial [SAK22]. inter-institutional [CWLS23]. inter-organizational [WSC^{+}24]. inter-social [Iva25]. inter-system [ARR23]. interacting [dMMdCM22]. interaction [LRW25, ON20, San21a, San25, TG24, TM25, VRVH21, WZHW25, YWZ^{+}23, dW24]. interactional [DFG21]. Interactions [Ano22a, HCZ^{+}22, Par24, ZZMS22]. interactive [DFG21, GBP21]. interconnectedness [HKN23]. intercontinental [KLM24]. interdependencies [LUH21a, LUH21b]. interdependent [SFJO20]. Interdisciplinarity [CSL22, CZQC24, LCL23, ÁRMS21, BPMRSL21, BMM22, CB20, CGDS24, CXZM25, DX20, GD22, HCZ^{+}22, Kav25, KFLS20, LRH22, Mut22, NWZ23, QZS^{+}24, SMG^{+}24, UIM21, YLM24a, ZNW22, vdB25, LB20].

Interdisciplinary [Ke23, YWN25, YLM^{+}24b, Can24, DDU24, EGV22, HZCM23, JDSL20, LXL25, MS21f, NM25, Par25, PVP^{+}23, RWL23, RLZ^{+}24, VC20, WMCL22, WCCC24, WLLW25b, XCML25, YRP21, YTJ23, YWL25, ZLL21, ZTY^{+}24]. interest [ML24, OMBP24, PFS^{+}20, TNSF21]. interests [XWGD22, XLA23, dDM25, vSDC20]. interface [Zbo25]. interlocking [GdM20, dMGDD24]. Internal [HL24, BA22b, Seb23a]. International [Avd21, BdPFdO25, BMW24, CFW21, CMV23, DMG20, GDR22, MSL22, MNM^{+}20, Vie22, Wag25, WGA22, APR19, Ali24, CCdC20, CTLH24, CTH22, CFA^{+}23, Cho24, DSL23, EEÁ21, EBB^{+}20, FH21, FRS23, FMB22, GEFJ24, HL23, JCZ22, KK20a, KLM24, Kwi20, Lin20, LZ24b, LDMVS21, LMSNZG22, MS21d, Mor19, Mu21, NB20, PWZ25, PS20, PLYZ24, PDH21, Plo25b, Poh20, RPCF20, RSF20, RP24b, RP25, SKBSC20, SK22, SR24, SL22, SA21, VEGSMMC22, VC22, Wad20, WJZ24, XYY25, ZZS^{+}20, ZCSK24, ZCL20, dFBGCGASG24, BP23, LGE^{+}23, Par24, Yur24]. Internationalisation [Poh21]. Internationalists [Kwi20]. Internationality [SS20a, Asa20a, BGAMS23]. Internationalization [HBO^{+}20, GP21, Gon23b, NSM24]. Internationalizing [TLM22]. internet [UNK23, FFJ21]. interplay [ABKGBR25, GRSFV21b, Hir22]. interpretable [WLMZ24, YN23, ZA20]. interpretation [Zho21]. interpreted [YMLL23]. interpreting [YN23]. interregional [YX21]. interrelations [Hüc22]. intersection [MY25]. intersectional [RPL21]. Intertwining [ZYL20]. interval [dOSAL21, CZFC24, NF21]. Intra [SMS23, AAD22, ARR23, Sha24b, ZCS^{+}23]. intra-country [Sha24b]. intra-organisational [ZCS^{+}23]. intra-sector [AAD22]. intra-system [ARR23]. Intra-Team [SMS23]. Introducing [HIA^{+}20, MSI21, VJCV22, Ara20, PVB^{+}24, ZS20a]. Introduction [BG21, BH25, HWP22, Kös20a, MH24a]. introductions [uHBKK20, uHBKK21]. invalid [CCA22]. invention [WL21b, YD21, Zha21a]. inventions [Fer21, KVZ20, KG20, SLL23b]. Inventor [RYS24, YL24, CF22, WSC^{+}24, WPSK25]. inventor-patent-classification-coupling [YL24]. inventors [SZZ23, XLA23, YMD20, ZL23c]. investigate [LYT^{+}20]. Investigating [AJB^{+}21, Ali25, AM21b, Fan20, FLDD20, HWC22, KFCK21, RLZ^{+}24, TP20, ZIS21, AA22, QSYL23]. Investigation [Dha22, ADD23, Azm22, BKT20, HHB^{+}20, HhC20, HWL23, PY24, PLW23, SSB22, UNK23, WH24, WSC^{+}24, Wan24a, WH25, ZWC22, ZSSH22, ZSS22, dSSA22]. investigators [HLS23]. investment [SCW^{+}23]. invisible [GdM20, HCS21]. involved [YWHS23]. involvement [ACDS22]. involving [Par24, YWX23]. invulnerability [LLL22]. IoT [BKS20, HNO^{+}21]. IoT-based [HNO^{+}21]. IPCC [HDY^{+}23]. Iran [RSV^{+}20]. Iranian [GFVK21]. Iraqi [AMMN21]. irregularities [ARR23]. isolation [CC22]. ISPD [GG25]. Israeli [WZG21]. ISSI [BMW24, vR21a]. Issue [Ano21e, BG21, CL21c, Cop20b, HH21, LM20]. Issues [SLY22, FWH^{+}23, HHY22, KHT22, LGY23, TdS20, YWB22, dSAKT20]. Italian [ATCS20, BBP23, BDPP23, CMMO20, CM20, DDL23, DRS20, MPS22b, MN24, Pet22, RPL21, TP20, Zan23, dBGP21]. Italy [AD20, BR25, DFRS20, VPC^{+}23]. items [LS20b, ZYPC20]. Iterative [BTD21]. ITGInsight [WZL22b]. Ivory [WCL25].

Japan [Fuk22a, Fuk22b, KN21, Mas20, NC25, NSM24]. Japanese [KKK^{+}25, Min25]. jargon [WY23]. JATS [Bös21]. JATSdecoder [Bös21]. JCR [BB22, DMA^{+}20, RPAVFV21]. JCRs [LBW21]. jerk [FLC22]. job [BM21, TP20]. John [Ano20a, Ano21a, Ano23, GTB23]. joint [CCZL23, HL23]. jointly [HM23]. Journal [AGDE25, BSS24, DS21d, GMRVU24, KBZT21, Pra20b, RL20, ŚT21, SZP^{+}24, ADD23, AKW23, AOM24, ARR23, BBB^{+}24, BB22, BBP21, BPW^{+}21, CB20, CLWlY24, Cho24, CM20, DD23, DMA^{+}20, DD21b, ES22, EGPGGA^{+}23, Fan20, Fan23, Fas21a, Fas21b, FZL^{+}20a, FZL^{+}20b, GRSFV21b, GMPRSFV25, GSH24, GUG^{+}22, GKD23, HM23, HKW24, Jam24, JL23, JP22, JTB23, KBS20, Kim24, KCC23, LS20b, Lin20, LHW21, MS22c, MS22d, MMG21, MMG23, MLA^{+}23, MMP21, MSP23a, MSP23b, MFF^{+}20, Mu21, NZL21, OdSO20, OFA23, Pra24, Sam23, SM22a, Sha24b, SSM21, TKS^{+}23, TZY^{+}23, THJ^{+}23, The20a, The23, TFWC24, THL23, UKP21, VMF22, VP21a, VP21b, WB21, WLLL20, YS24, Yen25, YP20, YTM21, ZS21, ZM21, Zha21b, ZS24, ZWL^{+}25b, ZZZZ20, ZJY22, Zho21, OdS20, CN24]. journal-based [OdSO20, THJ^{+}23]. journal-discipline [Zho21]. journalism [AM21a, GTH24]. Journals [CxGdH^{+}24, MB21, Aba21, AM21a, APR25, ABKGBR25, AGB23, Asa20a, Asa20b, Asa21, Asa23a, Asa24, BP22, BPLIdMA^{+}20, BPMRSL21, BSS24, BMA21, Bjø20a, Bjø20b, Bjø20c, BdCM21, Bou20, CSA20, CLWlY24, CLC23, CFC22, CSR23, DABSC20, DK21b, DGGD22, EPP23, Fas21a, FRRBDD23, GFVK21, GC21, GMRVU24, Gon23a, GdM20, GLPN25, GWvZ21, HBO^{+}20, HM21, HHB^{+}20, HWX22, HY22, Ibr21, IHLP21, Ioa23, Jab25, Jam24, Jam25, JP22, Jok20, KHT22, KHK21, LAT22, LBW21, LZH^{+}22, Lin21b, LCCH24, LLP21, LOGS21, LM20, LR21, LA20, LZL20b, Lyu21, MSJ23, MMG23, MR20, MWH20, MMB23, MdMAGB^{+}21, MKWB^{+}23, Mou21b, Mou21a, Naz20, Naz24a, NF22, Nuz20, Nuz21, OdS20, ODF20, OPMRP20, Pra23a, Pra20d, RFMMPJ22, RPAVFV21, RAF25b, RL20, SD23a, SD23b, SCDP21, Sch24a, SS23, Seb23a]. journals [Seb23b, See20, SMJ21, SSB22, STM^{+}24, Ste23, TD24a, TKM^{+}23, TP24, TYS22, Tut23, Tut24, TD24b, VY23, Wal22, WWZW20, WXZ22, WWL24a, Wei20, WLS22, WZNL23, WWL24b, WLMZ24, WAM^{+}25, XL20, YWX23, YMW^{+}20, Yur20, Zha21b, ZS24, ZM24, dSDE21, dSSA22, dCPT22, LHC21, Mou21c]. journey [SDW^{+}20]. judging [KSZ22]. judgment [Hug24]. Judit [BL20a, BGS20, Hal20a, HP20, LB20, OM20]. junior [DGRÁTS23]. Jürgen [dCFdCA^{+}23]. just [DV22, HLZ23, OG24]. justification [AAA^{+}23]. justify [FDM24]. JYUcite [SVY22].

KAKEN [CKO21]. Kardashian [PHP22]. Karl [BWH20b, BWH20a]. Kazakh [KUS22]. Keeping [BKO25]. Kendall [CR24]. kernel [Ste21]. Kevin [Ano23, Sug23]. Key [YT25, CXD21, CTH22, CZZN25, GAPR20, HSE23, KY20, KL24, LN24, OVS25, XHA^{+}20, YN23, YWL25, TSVV23]. key-node [KL24]. key-route [XHA^{+}20]. keyphrase [ZZZ22, ZYZZ25]. Keyword [BVS25, CWL^{+}20, Sam23, Cho20, HZCM23, MK23, OJKY23, YLCY24, YWY^{+}20, ZLZ^{+}23, ZSL24]. keyword-based [HZCM23]. Keyword-citation-keyword [CWL^{+}20]. keywords [CLC23, LZL20a, MB20b, TSA24, VBF24, ZXS22]. kind [CW25]. Kingdom [LSA22]. Klavans [Ano23, Sug23]. knee [FLC22]. knee-jerk [FLC22]. knockdown [BPW^{+}21]. know [KDS21, LB24b]. Knowledge [AKD23, Ano24, HL23, JCK22, LMG20, LTXZ24, MIY^{+}20, SMS23, YK23, YS20, YY21, ZWSC20, ADD20, AD20, BMBA20, Bor22, CLL22, CSJW22, CWL^{+}20, CZZN25, CLZ21, CXW^{+}24, Dip21, DG21, FESRMB21, GAACBA24, Gon23b, HGB20, HL21b, HYZ23, HGZ21, JLL21, KH21, KG20, LFV21, LN24, LZH20, LPD21, LC22, LTL23, LW24, LLLL24, LT20, LGMK21, Liu24, LOGS21, MY25, MSI21, MS21f, MYX22, MRC22, NKY^{+}24, NC20, Pal21, PR22, RYS24, SOSC24, SGA22, SSdCRD22, SLR^{+}23, SCK21, SHZ22, SZY24, Sta21, SYL^{+}24, SM24b, TDWW24, Vie24, WMCL22, WWC22, WSC^{+}24, WZJW24, WLLW25b, WFSZ24, WG24, WRT20, WCH^{+}20, WGZ22, WZM^{+}24, WZ25, XDW22, XCML25, YLH24, YCYZ24, YP20, YWB21, YWZ^{+}24, YXYQ22, ZLL21, Zha21a, ZW21c, ZXS22, ZLM^{+}23, ZMLZ24, ZL23c, ZZ20b, dMGDD24]. knowledge-a [SM24b]. knowledge-based [MSI21, TDWW24]. knowledge-employee [Zha21a]. Knowledge-integrated [YK23]. KNOWMAK [MLP^{+}20]. knowmetrics [LPD21]. Korea [CLZ21, JC22, KK20a, Kim23, LKPS25, YP20]. Korean [KK20a, Sha24a]. Kretschmer [GKML25]. Kuhn [BWH20b, BWH20a].

Lab [CWK21]. label [LCT^{+}24, RAL22, SMA21]. labeled [KOS21a]. Labor [SDRS23, CH20, CGJM24, CDQ25]. laboratories [LS21c, YT25]. lack [EOL23, SMG^{+}24, VP21b]. lacks [AAA^{+}23]. lag [Lin21b]. laggard [PT24]. lags [GFVK21]. lakes [HG20]. Lancet [DDH21]. landscape [Bou20, DDL23, HJT21, MGA22, NC25, RvdSvMK20, RBB22, THJ^{+}23, THJ^{+}24, Wan25, ZZ23]. Language [Ano20c, CFM20, CXW^{+}24, DK21b, EXT^{+}21, IHB21, JDG^{+}20, KP23b, Kos23, LZ25b, NK24, NJCL20, Ort20, Pea20, RFMMPJ22, RAL22, Sch25b, SLL^{+}23a, SLX21, WLZ22, WL24, YMW^{+}20]. languages [NSM24, Tei20]. Lankan [RBR22]. Laplacian [FZL^{+}20b, FZL^{+}20a]. lapse [ACT23]. Large [Nuz20, YMD20, ADD23, Asa20a, Asa21, CPN21, CXW^{+}24, DvE20, DGSVDG20, FCW22, GCS25, JHVC21, KBS21, KA20, KR22a, LTL24, LL25, LML21, MMFF25, NF21, NK24, SF20, SHB25, Sch25b, WWS^{+}23b, Xie21, YWN25, YWW23, mYY22, ZWL25a, ZDLR21]. Large-scale [YMD20, ADD23, CPN21, DvE20, DGSVDG20, FCW22, JHVC21, KBS21, KA20, KR22a, LTL24, LL25, LML21, MMFF25, NF21, WWS^{+}23b, YWN25, YWW23, mYY22, ZWL25a, ZDLR21]. larger [CXD21]. Largest [VCJ22]. last [FFRRPR^{+}24, GGS22, MQS^{+}20]. latency [Man24, SKS^{+}23]. latent [BPL25, CF22, DK21a, GJ24, Han20, TD24a, AKD23]. Latin [AdAFR24, Ano21c, BH24, Cor22, GB21, GMCRV21, LMLMPA22, PRGTOP^{+}24, RFG^{+}24, RP20, RPAVFV21, SdSCdS25, SM20, VdBC20, dCPT22]. Laudan [Sma24]. Laudation [Sug23, vR21a]. laude [vdBM24]. launch [BZZ25]. launching [Ali25]. laureate [HF22, JL20]. laureates [CJT22, DCL23, Kos20b, LL24a]. Law [Li25, CL20, GSB24, GS24]. laws [MRC22, She20]. Laxdal [WSRM23]. lay [WY23]. layer [CSJW22, LHXS20]. layer-based [LHXS20]. layered [XHQZ20]. layers [BBB^{+}24]. laziness [CHT24]. LDA [LLM24, YJJ25]. lead [CW25, ZWHS22]. leader [Ger24]. leaders [DH23, HGB20]. Leadership [CFA^{+}23, BGAMS23, CTLH24, CMT24, HGB20, HWZ21, SB20]. leading [BB22, CSR23, Fas24, GdM20, GLPN25, LCW^{+}20, LBCO21, MDCBdGVMD24, Mou21a, RAF25b, Sch24a, YRP21, lYnLyO^{+}20]. leads [CdSS25]. leaky [WZZ25]. Lean [BdOdS21]. learned [Fur20, KHC^{+}25]. Learning [SMS23, AUA^{+}23, AGDE25, BLL21, BJS22, CTH25, CPLS20, CXZ^{+}20, CCSX23, CLS21, CPL21, EXT^{+}21, FAY25, GJ24, GSP21, GCT^{+}20, GK22, GKD23, HJLZ23, HS22a, HAL25, HZCM23, IHB21, CLY^{+}22, KMGS21, KYL20, KDZ21, KG20, LHK21, LS21b, LLL25, LZD^{+}24, LLZ25, LZ25a, LXZ^{+}20, LUH21a, LUH21b, MLD21, MS21e, MKG23, PLH20, PR22, QWS^{+}23, RAL22, SSC^{+}23, SLH24, TZZZ23, VA22, WYT21, WCL25, WLS22, WLMZ24, YZWH24, YN23, YZZT24, YY22b, ZDL^{+}20, ZDLR21]. learning-based [CLY^{+}22, YZWH24]. leaving [OŠV^{+}24]. legacy [Hal20a, LCP22]. legal [LZ24a]. legitimacy [dSDE21]. Leiden [Bor20b, DDL23, Lec21, MJYS23]. length [GW23, HB21a, Wad24]. Lenin [KKS21]. lens [GC21, HSE23, UBROGS23, ODQ24]. lenses [PS20]. less [CRC20, KR22b, Li22, Ste23, WVH21]. lessons [KHC^{+}25, SMG^{+}24, Pra20a]. Letter [FC21, PM20, Pra20a, Pra20b, Pra20c, Pra21a, Pra21b, Pra21c, Pra22, Pra23a, Pra23b, Rus23, WSRM23, Lax23a, RJ20b, Zha23a]. Letters [Nuz21, FFJ21, Nuz20]. Leuven [vR21a]. level [AHH21, AAHY20, ANR21, AMS21, CMZ20c, DvE20, FWDQ20, GUG^{+}22, GX22, Hir22, KPY21, KG24, KLM24, KPS21, LZL20b, LC21, MS21c, McC24, MMG23, MGCR24, MAN22, MDGR21, OM21, Plo25b, PSA22, RA20, SS20a, SVY22, SXLC21, SC20a, SA24, SM24b, TWH^{+}24, Wad20, WL21b, WZZ^{+}21, WCY22, YJS21, Yur23, ZSSH22, ZS24, ZBG^{+}24, MS21b]. levels [SNK22, ZWZ^{+}19, ZWX22]. Leveraging [GVK24, LCB21, CNT25, DWM21, LFV21, MKG23]. lexical [Dia21, PY25]. lexicon [BCF21]. lexicon-based [BCF21]. Leydesdorff [BH25, Coo25, Etz25, Mey25a, Raf25a, Sma25, WW25, Wag25]. Li [SdL23]. libraries [Mal22a]. Library [HCZ^{+}22, Pra23a, SD23b, SSS^{+}24, AS23, AMTS24, Cha21, Cho21, Dag24, FRRBDD23, GA21, HhC20, Ibr21, Jok20, LXL25, Mal22a, Mal22b, MIY^{+}20, OdS20, SD23a, Tas21, TSAMT21, YCL25, ZT23, ZMP24, ZS20b, ZPH21, Dia21, SDC22, Tom21, VEGSMMC22, WGC20]. licensing [MLYK23]. Life [BTP24, VAMBTS25, Bha21, CH20, GMB22, HTZ24, MRC22, OTH22, RTT23, TNS21, WL22, YK23]. life-cycle [GMB22]. Lifespan [CMT24, CJT22]. lifetime [CYK22]. light [SFC^{+}23]. like [ZL20a, dSDE21]. likely [CGDS24, Cop20c]. limitations [Ans23a, DLK20]. limited [CNT25, Ste22]. line [BPW^{+}21]. linear [For23, HCL22]. lines [Sma24]. lingua [DWG23]. lingual [LCT^{+}24, MGA22]. Linguistic [BA24, LZ23, PMB^{+}23, WWS^{+}23b, BZZ25, CDZZ20, LZ25b, Mu21, NA24, WKC22]. linguistics [Li22, MMB23, Pea21, RFMMPJ22, WLZ22, XL20]. lining [GCGB^{+}24]. Link [OMC21, CLS21, DSL21, DG21, GCS25, NZ23, OCKY20, QZF22, RBV25, RNB22, VA22, WL21a, XZZ22a, XZZ22b, YKS22]. Link-based [OMC21]. linkage [MKI24, TZ24, WZJW24, YY22b]. linkages [LZL20b]. linked [Bat20, KOS21a, OMFJ22, PROGDG24, PM20]. LinkedIn [HKN23]. Linking [KFLS20, CR24, Mar20, MLK22, MJRR25, THvB21]. links [COB21, MB22a, SF20]. LIS [Pra23a, SD23a, SD23b, SSS^{+}24, CCMM24, Han20, HZJ22, Kim24, TD24a, UA21, UA20, VCJ22, YWHS23, ZHH23a]. List [dSSA22, Fas21a, GW23, HKW24, KK23, MFMC21, WG22]. listed [CGJM24, TdS20]. lists [SdLV21a]. literacy [BMS21a, LCW21, PFPCMS20]. literals [SGA22]. literature [AKH22, AKD23, Ano20c, BGB25, BJS22, BKT20, BKT22, CFM20, CPLS20, CG22, CGA21, CBA24, DGGD22, DMA^{+}20, Ela21, Ela22, FAY25, FESRMB21, GMB22, HB20, HV21, HNO^{+}21, HWC22, KH21, KZH25, KTB22, KS22c, Lee24, LB24b, MF20a, MLK22, OŠV^{+}24, ODQ24, dOPGdAPU22, PXY24, Pol20, PW20, QZ20, Sch25b, SMD22, SLH24, VdBC20, WZL22b, WWC22, WZL22c, XL20, ZSL24, ZBKW24]. literature-aging [GMB22]. literature-based [KH21]. literatures [ZZ20a]. litigated [ACDS22]. litigation [ACDS22]. Liu [LGD21, Ano21e, CL21c, ZLL21]. Local [Jam25, MGA22, DCdA25, KK20b, Mu21, VA22, WLL25, ZWL23]. local-oriented [KK20b]. locality [ZYL20]. localization [AD20]. locals [Kwi20]. location [Pur24, QLW25, ZS20a]. lock [LMC22]. lock-in [LMC22]. lockdown [GCGB^{+}24]. Loet [BH25, Coo25, Etz25, Mey25a, Raf25a, WW25, Wag25]. logic [dOCP21]. logical [ARR23]. Logistic [TNS21, UMA21]. London [PSH22]. Lone [LR25]. lonely [LR25]. long [Cha23a, HWX22, MM25, RS22b, RP24a, YWN25, ZA20]. long-term [Cha23a, HWX22, YWN25]. longer [Kua23, Seb23a]. longevity [CJT22]. Longitudinal [CMK24, FGCGFC20, ARR21, FJO20, HAR24, MB22a, PCG^{+}20, PT24, SSK22, WL24]. look [CB20, ECD24, Lin24, YCXY20]. Lorenz [RZS23a, RZS23b]. Lost [MN23, SK22, JT21]. Low [KKK^{+}25, LRH22, OVS25, Wal22]. low-tech [OVS25]. lower [KS22a]. lower-ranked [KS22a]. LSTM [HTZ24]. Ludo [vR21a, Ano21a]. lunches [SMT^{+}23]. LUSTRUM [HGDRM20]. Lutz [Ano20a].

MacDonald [Lax23a]. Machine [HJLZ23, LUH21b, AGDE25, BJS22, CTH25, CLS21, CPL21, EXT^{+}21, GJ24, GCT^{+}20, HS22a, IHB21, KYL20, KG20, LHK21, MKG23, SSC^{+}23, VA22, WAT23, WLS22, WLMZ24, YN23, YY22b, ZDLR21, LUH21a]. machines [GS22a]. made [Nis23a, Nis23b]. Magnitude [LLSH23]. mails [LBP22]. Main [AYRG23, LDMVS21, OJKY23, AdAFR24, Fil21, Kua20, Kua23, KL24, LLH20, MB20b, SSK22, XHA^{+}20, YS20, YY21, YY22b]. mainstream [LOGS21, dSSA22]. maintain [BPH21]. Major [jDXjS21, BW20a, KLLG20, Öz24, dSTE21]. majority [Old22]. MAK [Kös20a]. make [GZOC24]. Making [Rod22, SdS25, BW20b, Kaj22]. male [AKW23, DT20, LXY23]. Malmquist [DCdA25]. malpractice [SSBC22]. malware [MRF21]. managed [AGWG^{+}23]. Management [Ano22b, BVS25, CW25, Cor22, DDU24, Fas21a, Fas21b, HL23, HWP22, KP23a, KPY21, MMG21, MWH20, MS21f, NLS24, OPMRP20, SdS25, See20, Sta21, UNK23, WJ22, ZWL^{+}25b, WHÅ24]. Managing [GM22, KUS22]. mandates [DMR22, STAÁMR24]. mangrove [GVS20]. manifestations [NB22]. Manifesto [Lec21]. manipulation [WG22]. Mann [CR24]. manpower [SKBSC20]. manufacturing [Fil21]. manuscript [Asa24, CMZ^{+}20a, CMZ^{+}20b, GRSFV21a, Moo21, TD24b]. manuscripts [JXZW23, Ste22]. many [JKK24, LYZ^{+}20, LZL22, MB20a, San25, TSDAMRG24]. map [HGB20, MLP^{+}20, Ste21]. Mapping [BPL25, BMS21b, CCdC20, CDZ^{+}20, Dag24, DX20, ECD24, HS22a, HBW^{+}21, HBN^{+}21, HTGW20, JL25, KNT20, LCT^{+}20, LZF^{+}24, MDCBdGVMD24, PKCOG21, RKMG20, RdNB21, RBB22, SVC20, WUH21, XHQZ20, YJS21, YWL25, ZZ23, dMGDD24, dSMNdV^{+}21, AA25, ACD^{+}24, BMBA20, FH22, HJLZ23, HNO^{+}21, HGZ21, KH21, KA20, KFLS20, LZL20b, NHAB21, PYS22, QD21, Raf25a, RFG^{+}24, RvdSvMK20, SNK22, VCD^{+}23, XW25, YY21, ZIS21, BCF21, KMGS21]. maps [BVvEW25, BVS25, KFLS20, KCC23]. Marginal [FM21]. marginalizing [OFA23]. Marine [TdSFB20, Mit20]. mark [LHCH20]. marker [HM20a]. Market [Asa20b, Zan23, BM21, CGJM24, CF23]. marketing [Mou21b, Mou21c]. markets [AJBB22]. married [XS25]. Marshall [GHGMSM21]. Marx [KKS21]. Mas [Yur23]. Mason [JTB23]. mass [CdAdFC21, YWY^{+}20]. master [Wal22, XLL20, Xie20]. Masters [PSH22]. match [GSOW20, SM22a]. matching [AL24, BO20, CKO21, FJO20, Pup21, ZXL20]. matching-based [BO20]. Materials [AGB23, YTJ23]. Math [GPYR^{+}20]. Math-word [GPYR^{+}20]. Mathematical [HLR21, SSC21, GDG^{+}20, ODF20]. mathematicians [HCZEP20]. Mathematics [Yur24, DK25, HC22, SG21]. MathSciNet [Tom21]. Matilda [GDdZ25]. Matrix [ZC23, Fan20]. matter [AGWG^{+}23, CFI22, Cop20a, GSOW20, KBZT21, Lin21a, Liu21, MM24, Mal22b, SdVvdR24, SFC^{+}23, WGLN23]. matters [CdSS25, MH24a, SCK21]. Matthew [LLSH23, YS23, Tei21]. maturing [dMGC23]. Maturity [LN24, KPY21, Tay23]. Max [BKO25]. maximization [LCC^{+}20]. maximum [DJM22]. Maxwell [FD21]. may [Cop20a, DDD^{+}21]. MCDM [AMIK23]. MDPI [Cer24a]. me [CLD22]. Mean [Sch20a, KBZT21, Sch22]. meaning [Iva25]. meaningful [BDTC20, HBP20, QJI^{+}22, VP21b]. meanings [BY23]. means [HM21, Hla25, VMF22]. means-ends [HM21]. Mearsheimer [GTB23]. Measure [Pra23b, Can24, CFI22, DX20, GD22, HBP20, KLNW20, LTL23, LGLM20, Mey25b, SvWC21, UKP21, VDY21, WFW24, dOCP21]. measured [VCJ22]. Measurement [ZZW24, Can24, DCL23, DD23, NWZ23, RGVQTS^{+}24, ZBG^{+}24]. measurements [ZZ25]. Measures [ZC23, ÁRMS21, BBP21, BW20b, DDD^{+}21, Egg22, ER23, SL22, SG21, SvWC21, ZJ22]. Measuring [AMMN21, ALM^{+}23, BD22, CC21, CGJM24, CC22, DSL23, DLS23, FSR21, GUG^{+}22, GDdZ22, HYY25, HCZ^{+}22, KTK21, KLLG20, LC22, Mad20, MPC22, MKI24, QQL21, SV20, Sca25, Sch25a, SW20, SDC22, TZ24, WWC22, WLLW25b, ZW21a, ZSL24, ZW24b, CWZ22, DLK20, GDDG24, HIA^{+}20, cHhC24, LLZ25, SC20a, THG23, WZL22a, WZG21, YTM21, ZNW22, GZOC24, VdOS^{+}20]. mechanical [YP24]. mechanism [CTLH24, CZX24, HWW^{+}24, MYK22]. mechanisms [AL21, DGGBDG21, Lin21b]. Medal [Ano20a, Ano21a, Ano23, GDRPR23, Rod22]. medalist [HZC21]. medalists [Yur24]. media [AMTSRG21, CWZ21, CB25, DK21a, DMA^{+}20, GTH24, GDRPR23, HAS^{+}20, HM20b, HYZ23, JVY22, KBS20, MBL21, SZN^{+}21, TDM24, WC20, YLf25, ZIS21, Zha23b, ZJ25, ZHH23b, dW24]. median [KBZT21]. mediating [DSL21]. Medical [Lam20, RSV^{+}20, SJW21, Ste23, AKH22, CPLS20, CLC23, DAGG24, FGCGFC20, HBC25, HJCT23, LPKG20, LRSB21, LPD21, MKWB^{+}23, NF22, RAF25b, Sch21a, SdLV21b, SS23, STM^{+}24, SJW22, SLTS25, THL23, dSSA22, KCC22, PMB^{+}23]. medicine [CDZ^{+}20, GJ24, HF22, Lax23b, OTH22, PY25, RNAH21, Seb23a, WZY21, WSRM23, YP24, ZL23a, ZL23b, ZF22]. medicines [The21]. Mediterranean [Ali22]. MEDLINE [Bou21]. medRxiv [SCP22]. meets [CNP21]. mega [The20a]. mega-journal [The20a]. megajournals [Hen19, Hen20]. Member [GDM22, AHRH24, WL21b]. member-contributed [AHRH24]. Members [SMS23, Ali21a, Ali21b, SO21, XAA21, YLM24a, ZW21c]. membership [CFI22]. Memorial [Ano21e, GDRPR23, CL21c]. memoriam [Coo25]. memory [AQK^{+}22, Etz25, ZA20]. MENA [Azm22]. menagerie [Lin24]. Mendeley [RBR22, TFWC24]. mention [FNM^{+}22, PVB^{+}24, WWY^{+}20, YLf25, ZLY21]. mentioning [YMLL23]. mentions [KBS21, VAMBTS25, YCXY20, YYC22, YMLL23]. Mentoring [YA21]. mentors [Che24]. mentorship [SDD20]. merging [NII24]. MERS [HB20]. MeSH [CLC23, FZC20, IPP22]. Meta [GVS20, SS24, AlYT21, FZhCC21, Ioa23, Le 24, LRXC21, Ony20, Pup21, SXLC21, ZZ20b]. Meta-analyses [SS24, Le 24]. Meta-analysis [GVS20, Ony20, Pup21, SXLC21]. meta-frontier [AlYT21, FZhCC21]. meta-path [ZZ20b]. meta-research [Ioa23]. meta-synthesis [LRXC21]. metadata [ACD^{+}24, Bös21, Dag24, HIQ^{+}20, MLD21, RBV25, SF20, VPDAG22, WAT23]. metadiscourse [Bir21, DFG21]. Metapath [LH24]. metaphor [RLZ^{+}24]. metasemantic [TDWW24]. metaverse [ZXL^{+}24, Wan25]. meteorology [SH20]. method [AA22, dLBdAPUdLB^{+}25, BKT22, BERD21, CGJ22, CXZ^{+}20, CWZ22, CLWlY24, DLC21, GP21, HM23, HAL25, Hir22, HTZ24, HTZR24, KYKC21, LCC^{+}20, LLC22, LWZ22, LZD^{+}24, LXS22, MRBSFSJ24, PYS22, Pra21a, Ste22, Vin21, Vin23, WWC22, WZL22c, WLLW25b, WPSK25, YRP21, YZWH24, ZS21, ZMP24, ZZ24, ZJY22, ZZ20b]. Methodological [LGY23, ANR21, BBP21, GdMDD24, VCV22]. methodology [AD21, ACV23, DS21d, HAS^{+}20, JME20, NC25, PROGDG24, YS24]. methods [AGDE25, Can20, CN24, CR24, HCS21, HMW24, Ibr21, JS21, LMSNZG22, MS21e, OOCD^{+}21, ODF20, Sun23, Ven24, XLL23, YZH23, ZSSH22, ZT23, ZMP24]. metric [DZ23, EEÁ21, GDRPR23, KZH25, LPKG20, ÖT24, San21b, ŞKKT25, SVY22]. metric-based [ÖT24, ŞKKT25]. Metrics [BH25, Lec21, ATC20, Cop20d, Fas21b, FVLA20, Kim24, LKPS25, MS22d, Naz24b, OdSO20, Ove23, SSL22, VP21a, VA22, WP22, dS21b, Tei21, dSVN24, dCPT22]. metrics-based [FVLA20]. metrological [SDW^{+}20]. Mexican [MOMGNLG24]. Mexico [BOV21, CPFVPA25, EEÁ21, FVGBPA23, HGM21, LSFLZLLV20, LOGS21, RP23, SRL20, TPVVPA21, dRRJ20]. Mf [XXL25]. Mf-cite [XXL25]. MFCOPIs [SLL23b]. MHCF [WQ21]. micro [MY25, WHÅ24]. micro- [WHÅ24]. microeconomics [Yur23]. Microsoft [MMTOMLC21a, MMTOMLC21b]. Microwave [SdSFB22]. Mid [The20b]. Mid-career [The20b]. Middle [EOL23, EO24a, EO24b, LL21, SB21]. migration [SA21, WGA22, ZAZS22]. Mikhailidis [RJ20b]. million [HBC25, ZYPC20]. mills [Day22, OG24, dS21a]. mind [Pra20a]. mineralogy [CPFVPA25]. Ming [SdL23]. Mining [Ano22b, XZZ22a, XZZ22b, YD21, ZXS22, AS25, Ano20c, BCF21, BGSRF20, CFM20, CLC23, Fer21, GRBAM22, HWP22, Jia21, KA20, LZD^{+}24, LZ25b, MKI24, WPSK25, ZPZ21, ZZ20a, dRRJ20]. Mirabilis} [YG20]. MIS [KSS20a, LHC21]. misalignment [AA25, KKYR24]. misclassification [LUH21a, LUH21b]. misconduct [FFRRPR^{+}24, LdS25, RM24, SdSCdS25, Sha24b]. misinformation [AYS24]. misleading [dS21b]. mismatch [TdSFB20]. miss [KZV20]. Missing [ZCS^{+}24, LYS22]. missions [KFCK21]. mitigation [FCY23]. mixed [For23, GP21, YRP21]. mixed-method [GP21, YRP21]. MLP [QJI^{+}22]. Mobile [KDN22, PFPCMS20, SBKK21, YRP21]. mobilisation [HB20]. mobilities [PDH21]. mobility [BdPFdO25, Cha23a, CWLS23, FRS23, GMKG20, HLS21, JL20, JXY^{+}23a, JXY^{+}23b, KFCK21, LWX21, LDMVS21, MDCBdGVMD24, PWZ25, PLYZ24, PLW23, Plo25b, PL23, SKBSC20, WSC^{+}24, XYY25, YSW^{+}20]. mobilization [WCH^{+}20]. mock [Pra21c]. mode [Bat24a, vR21b]. model [AFH21, ASX22, ADS20, Avd21, BAE24, Can24, CF22, CPN21, CCZ^{+}20, CYK22, CLS^{+}20, CdSK21, DJM22, DG21, DD21b, ER24, FZhCC21, For23, GZSC22a, GHW22, GRSFV22, GW23, GMB22, Han20, JJPC20, KS23, KKOL20, LLM24, LHX^{+}24, LHCH20, LZL20a, LCT^{+}20, Liu24, LLZ21, LTXZ24, LCT^{+}24, MLD21, MCDFSP24, MN24, NZ23, PSS23, QJI^{+}22, QWS^{+}23, QYL21, Reh20, SB25, Sch25b, SLL^{+}23a, SLL23b, WSZ21, WLL25, Xie21, YZZT24, ZLZ^{+}23, ZYWA25, Zie22]. modeled [TSA24]. Modeling [CLL22, DGGD22, YXYQ22, ZA20, BMM22, DDVV20, JPS21, JJY21, JMZ^{+}23, MBL21, MIY^{+}20, PLH20, TD24a, VPR22, XW25]. modeling-artificial [MBL21]. Modelling [AdAFR24, CBP^{+}24, Mal22b, TNS21, BERD21]. modelmania [ODF20]. models [AUM21b, ADU22, CXW^{+}24, FD21, cHhC24, Jab25, KDZ21, NS23, NK24, NJCL20, SWT20, SAS23, SM24a, VPDAG22, Ven24, XLL25, ZA20, ZW24a, AUM21a]. moderating [AAD24b, CXZM25, FYY21, GYWZ21, LLLL24, LMG20, MJYS23, RYS24, SS24, TXL21, XCT20]. moderation [Zha21a]. modern [GMCRV21]. modes [HTX23]. modification [Fan23]. modified [CYK22, PR23, SWT20]. Moed [DGN22]. molecular [FFH^{+}24]. momentum [ZC21]. monetary [GCS25, TA20, VP21a]. monoculture [DT20]. monographs [KK20b]. monopolising [Fog23]. month [MWH20]. months [DR20]. most [GDdZ22, GDG^{+}23, Gus22, HV21, HPL21, KP23a, Rod22, SAR19, WZL22a]. most-highly [WZL22a]. moth [dSVN24]. motifs [YTJ23]. motivated [DGGBDG21]. motivates [VdOS^{+}20]. motivation [Ali22, GMKG20]. motivations [LRXC21, PXY24]. move [ZQML23]. movement [vSDC20]. movie [NT24]. moving [Sch25a]. much [BSFS^{+}22, CCH25, Pur25, SSL22, SPA20]. Multi [Ano22a, Cop20c, GVK24, GYWZ21, LZ21, QWS^{+}23, BAE24, Bha21, CCH25, HWW^{+}24, KS24, LHK21, LWW22, LLZ21, LCT^{+}24, LXS22, MJYS23, NKY^{+}24, NZ23, RAL22, SAA21, TFWC24, WH23, XHQZ20, XXL25, YJS21, YWZ^{+}24, ZZMS22, Zho24]. multi-anchor [NKY^{+}24]. multi-authorship [KS24, SAA21]. Multi-criteria [Cop20c]. Multi-dimensional [GYWZ21, TFWC24, YWZ^{+}24]. Multi-Disciplinary [Ano22a, ZZMS22]. multi-faceted [MJYS23]. multi-feature [XXL25]. multi-head [HWW^{+}24]. multi-label [LCT^{+}24, RAL22]. multi-layered [XHQZ20]. multi-level [YJS21]. Multi-network [LZ21]. multi-perspective [Bha21]. multi-relational [NZ23]. multi-scenario [CCH25]. multi-source [LWW22, Zho24]. Multi-task [QWS^{+}23, BAE24]. Multi-Tasking [GVK24]. multi-technology [LHK21, WH23]. multi-view [LXS22]. multi-viewpoint [LLZ21]. multicluster [KFLS20]. Multidimensional [GRTPAJ21, Xie20, Cao20, GS22a, Mu21]. multidisciplinarity [MFM20]. Multidisciplinary [CF23, AJVACC22, MMTOMLC21a, MMTOMLC21b, SCU21, ZS24]. multifaceted [CF22, SMBD^{+}24]. Multilateral [LZ24b]. Multilayer [WQ21]. Multilingual [CKO21, MMB23]. multimedia [Cop20b]. multinational [LH22b, Mey25b]. Multiple [CWZ21, BBB^{+}24, BBP21, CTH25, CR24, FYY20, KA24, LWW25, LYLC23, NII24, PMC21, Raf25a, Sma24, SLL23b, THvB21, WCY22, ZXS22, ZW24a]. multiple-field [SLL23b]. multirank [Dip21]. multivariate [DRF21, KBSS23, PSS23]. multiverse [LB24a]. mundi [FG20]. museum [DDU24]. museums [KSAT24]. Music [GS22a, OSS^{+}23]. Muslim [Old22]. Muslim-majority [Old22]. Mutual [LDA22, SM24b]. my [Fur20].

Name [MJRR25, DvE20, KOS21a, PJ24, PMC21, SGA22, SdLV21a, WQ21, WQ22, YMD20, Tei24b]. named [KKOL20, SM24a]. names [ACV23, GdMDD24, Jam24, KMM22, MB20a]. nanoparticles [LSFLZLLV20]. nanotechnology [GBH21, HGM21]. narrow [ZS20a]. nation [KL22, MQS^{+}20]. National [AS24, BDPP23, HF22, HM21, KN21, KUS22, RPL21, STAÁMR24, BRJ21, GKK23, HWL23, KLLG20, MLS21, MPS22b, MGCR24, MGA22, MdMAGB^{+}21, Naz20, NB20, RPCF20, SRL20, SB25, She20, SWL^{+}23, TYS22, Tut24, TD24b, ZZL^{+}20, vdPF25, WLLW25a]. nationally [LPY22]. Nations [MBH25]. native [BT25, YMW^{+}20, ZZW^{+}22a, ZZW^{+}22b]. Natural [Ano20c, WLLW25a, CFM20, EXT^{+}21, IHB21, KP23b, Nis23a, Nis23b]. nature [AJVACC22, DBR21, MJYS23, ZSSH22, Emm19, Mil20, TTBA20]. Navigating [RBBP24, BPL25]. Navigation [NJCL20]. Navigation-based [NJCL20]. Nazarovets [Cer24a]. NBD [CYK22]. nCoV [HB20]. nearly [HB22b]. necessarily [HY20]. necessity [HLZ23]. need [Ans23b, LPKG20, Pol20, OŠV^{+}24]. needs [LZD^{+}24, TdSFB20]. negative [Bor20a, FD21, XDL22]. neglected [Zha23b]. neonatal [Vie24]. Nepal [PG24]. nepotism [SH08, SH21]. nested [VBF24]. NET [WWS23a]. Netherlands [vdPF25]. Network [BKS20, CSJW22, EGP22, GLK^{+}23, Lie20, SZ20, Yur22, Aba21, AM21a, AQK^{+}22, ASK22, AMTSRG21, BBP23, BBB^{+}24, BR25, BWLX20, BZL21, BMS21a, Bat24a, BZG22, BVS25, CLL22, CCH^{+}24, CZX24, CW25, CWL^{+}20, CFI22, CDG20, CBA24, DMR20, DT20, DD23, FT21, FZhCC21, FFH^{+}24, FMB22, FSR21, GS22a, GdM20, GYWZ21, HWZ21, HZ21, HCN21, HCL22, HWW^{+}24, JJY21, JL25, KS23, KA24, Kim23, LNS21, LHC22, LZL20a, LT20, LZY22, LL24b, LXL25, LZ21, MBL21, MB21, MB22b, MK23, MBF23, MMG21, MMG23, NKY^{+}24, OOCD^{+}21, PLT22, QYL21, ROB21, RBBP24, RSM24, SVY22, SK21, SCK21, SLY22, SMZY23, TA22, Thi20, TXL21, dFVFA^{+}24, WWY^{+}20, WGW21, WZZ^{+}21, WWS23a, WKC22, WCH^{+}20, WZ25, XWGD22, XZZ22a, XZZ22b, XHQZ20, XDW22, YWZ^{+}23, YCYZ24, YP20, YKS22, YWB22, YTJ23, YXYQ22, Yur20, Zha21a, ZW21a, ZGBH20, ZZ20b]. network [ZQS22, dBGP21]. network-based [BVS25, SMZY23]. network-driven [dFVFA^{+}24]. network-standardized [SVY22]. networkedness [KN22]. networking [HBW^{+}21, PB24, RFG^{+}24, YLCY20, YWZ^{+}23]. networks [ABK20, Ara20, BVvEW25, Bat20, Bat24b, CW24, CB20, CPN21, CNT25, CLC^{+}24, DSH^{+}20, GAPR20, dMGC23, HSE23, HPL21, HMZ21, JJPC20, JL20, JCX22, JDSL20, KS23, KFCK21, LHW20, LZH^{+}22, LLL22, LH24, LLLL24, LZ25a, MPS22a, MYK22, ML24, MB20b, MF20a, MPS22b, Min25, MLYK23, MOMGNLG24, NSM23, dOPGdAPU22, PRS^{+}21, PFS^{+}20, RKMG20, RFG^{+}24, SSCK21, SHH23a, SHH23b, SFJO20, SA24, Söd23, SS22, TDWW24, TPVVPA21, TSA24, TBPN^{+}20, VC22, VA22, WWH21, WKC24, WP22, Xie21, YLCY20, YLC^{+}20, YWW22, YCL25, YX21, YA21, ZA20, ZWD20, ZCS^{+}23, dOCP21, Mas20, PM20]. neural [HWW^{+}24, MBL21, NZ23, TA22, Thi20, WWY^{+}20]. neural-network [Thi20]. neurodegenerative [GVS20]. neuroscience [RL20]. neurosciences [LGR^{+}21]. never [GSOCS21]. News [ATC20, DK21a, KPM22, FNM^{+}22, HJCT23, MKG23, Ort20, YYC22]. newspapers [DMGBG20, PL22a, Pet22]. next [Ant20, FFN25]. nexus [MJ22, MJ23]. NIH [LB25, XLL25]. NIH-funded [LB25, XLL25]. Nine [ZYPC20]. NISO [Bös21]. NISO-JATS [Bös21]. NLP [IQ21, LNS21]. NLP-based [IQ21]. no [MHD21a]. Nobel [Bjø20a, Bjø20c, HMW24, AKD23, ACT23, Bjø20b, CJT22, CSC23, DCL23, HF22, KJY24, Kos20b, LRH22, LLH22, LLSH23, LL24a, LHZ^{+}23, PS22, RWL23, Rod22, Tol24, TTB20, ZZ23]. Nobel-winning [RWL23]. node [KKOL20, Kua20, KL24]. noise [CCSX23, DZ23, MKG23]. nomination [KNT20]. nominations [Fur23]. Non [ZT23, ADG24, AGWG^{+}23, BPW^{+}21, Gra20, Ioa23, KDN22, LUH21a, LUH21b, MMF20, MPC22, MSJ23, MNP21, SV20, SMJ21, Ste23, dGdO^{+}23, VdOS^{+}20, YMW^{+}20, UIM21]. non-academic [MNP21, dGdO^{+}23]. non-CEE [MSJ23, SMJ21]. non-competitive [MMF20]. non-human [Gra20]. non-managed [AGWG^{+}23]. non-mobile [KDN22]. non-native [YMW^{+}20]. non-parametric [SV20]. non-pecuniary [VdOS^{+}20]. non-questionable [Ste23]. non-research [Ioa23]. non-scientific [ADG24, MPC22]. non-STEM [UIM21]. Non-synchronism [ZT23]. non-targeting [BPW^{+}21]. non-trivial [LUH21a, LUH21b]. nonimpact [HY20]. nonlinear [WSZ21, XZD19]. nonprofits [Pur25]. Nordic [San21a]. norm [MAS^{+}21]. Normalisation [BW19, BW20c]. Normalising [WAM^{+}25]. Normalization [SA24]. normalizations [MBF23]. normalized [BT20, GDDG24, HB22b, LdGRV^{+}22, Mad20, Men25, SHB25, SS22, ZC21]. Normative [BP22, DV20]. Norms [San25]. North [CFA^{+}23, EOL23, EO24a, EO24b, KK20a, SCU21, YP20]. Norway [WGA22]. Norwegian [SCDP21]. Note [LHC21, MS22a, BC21, EPP23, LLH20, PS22, Pra25b, ZMLZ24]. notices [AYS24]. nouns [WH25]. novel [ACV23, ADS20, ACD^{+}24, AMIK23, CGDS24, GK22, HTZ24, KZH25, LCT^{+}20, LLZ25, Mad20, PJ24, PSH22, THvB21, VWR25, WLZ^{+}23, ZWL25a]. Novelty [Kav25, DSL21, MSI21, WWL24a, WFW24, WFSZ24, WXD24, XLL25, YF24, ZZ25, ZJ25]. NPE [YL24]. NSFC [LWZ22, WZZ25]. NSGA [LHX^{+}24]. nucleotide [BPW^{+}21, LCW^{+}20]. nucleus [CNP21]. Number [NF21, AGB23, MFMC21, Moo21, PD20b, Pra23c, Vin23, Zbo25]. Numbers [MS22c, BM24, LHW21]. numerical [KYL20]. Numismatic [APR25]. nursing [KVZ20, KZV20]. nutrition [RdNB21]. NUTS2 [AAHY20]. NUTS2-level [AAHY20].

OA [BPLIdMA^{+}20, THJ^{+}23, THJ^{+}24]. obesity [KFLS20]. Obituary [Ano20b, Ano21b, Che21, DGN22, GS22b]. objective [SD20]. objectivity [BAE24]. Obscure [ZWD20]. observation [FYY20]. observational [CBP^{+}24, RSV^{+}20]. Observations [ZZL^{+}20]. observed [Cho20, Tay23]. obsolescence [CG24, DGGD22]. obtained [SHB25]. occasion [Sug23, vR21a]. occur [Asa23a]. occurrence [Ben24, DX20, FZC20, LH22a, MK23, VBF24, YWY^{+}20, ZYZW23]. occurrences [Sam23]. Ocean [Wra20a, OM20]. oceanography [Mit20]. Ockham [GZSC22b, Pra22]. OCLC [Mal22a]. OCRA [Can20]. Octopus [Mou20]. oeuvre [GTB23, OM20, LB20]. off [LRSB21, WH25]. offering [Fuk22a, Fuk22b]. official [WL24]. often [Sma24]. old [GDM22, HP24, JL20]. older [LZYR24]. oligopoly [SL24]. Olkin [GHGMSM21]. Olympiad [Yur24]. onboard [San21b]. oncology [GGS22]. One [LZL22, DT20, KL22, Cop20d, DLC21]. one-hundred [DT20]. One-to-many [LZL22]. Online [HC22, DD23, JDG21, OMAAORCL23, Par24, RAF25b, SCK21, STM^{+}24, ZQS22]. online-only [STM^{+}24]. only [KS22b, STM^{+}24]. ontologies [MLP^{+}20, Raf25a]. ontology [ADS20, CLS^{+}20]. ontology-based [ADS20, CLS^{+}20]. Open [AGL22, DMR20, DGSVDG20, HJT21, HNM^{+}24, LT20, MLS21, OMFJSC24, SCP22, WWP20, YSWC24, AÁSTMR24, Asa20a, Asa20b, Asa21, Asa24, BBP21, BPLIdMA^{+}20, BMBA20, BDPP23, BGAMS23, Bös21, BGWZ20, Cho20, DST23, DK21a, Dej25, DDU24, DMR22, Gra20, HAR24, HTX23, HP21, HMPS24, Hen19, Hen20, HHB^{+}20, HKV20, JL23, JVY22, KL22, KW24, Lin21b, Mad20, MS23, MBD25a, MBD25b, MMP21, Mor20, NRGvSTS23, Naz24a, NII24, NM25, NH24, PROG21, PROGDG24, RPL21, STAÁMR24, SGC21a, SGC21b, SZ20, SL24, SPS20, SSK22, Tay20, Tay23, TD24a, VRVH21, WC20, Wei20, WZNL23, WPS22, WLMZ24, ZM21, ZWW21, ZWHS22, ZM23a, ZM24, ZXL20, ZHH23a, ZHH23b, Agu20, Bou20, BL20b, KW24, MMGC^{+}24, PMB^{+}23]. open-access [Gra20, HAR24]. open-source [NII24]. OpenAlex [CHJ^{+}25, ODQ24, SHB25, YYZ25, ZCS^{+}24]. OpenCitation [MMTOMLC21a]. OpenCitations [HMPS24, MMTOMLC21b, ZYPC20]. opening [LBW^{+}20, NZL21]. Openness [MFP21, VdOS^{+}20]. Operationalization [Iva25, GMB22, GDG^{+}20]. operations [RL20]. ophthalmology [KYKC21, SMBD^{+}24]. opinion [AS25, DH23]. opinions [CZQC24, JPS21, PHP22, SSC21, SSK22]. opioid [SCU21]. Opium [BM24]. opportunities [CdSK21, LS21b, MPS22a, WH23, Wan25]. opportunity [HAL25, Nuz20, ZY20b]. optimal [GM22]. Optimization [ABAAN^{+}23]. optimized [LLC22]. oral [EGPGGA^{+}23]. ORCID [ANR21, KOS21a, MDCBdGVMD24, Por25, WLHY24, dS21a]. ORCID-linked [KOS21a]. ORCIDs [Bou21]. order [DK25, Dri24, FC20b, FC20c, HGMÁW23, LZH^{+}22, NSM23]. organisation [NB20]. organisational [ZCS^{+}23]. organisers [FH23]. Organization [MPS22a, ACV23, JCX22, LMG20, WGLN23]. Organization-oriented [MPS22a]. organization-topic [JCX22]. Organizational [WHÅ24, FFL21, HF22, LL25, SB20, WSC^{+}24]. organizations [CMZ20c, ZZL^{+}20]. Orientation [SMS23, HM21]. orientations [DG24]. oriented [CdSK21, HS22a, KK20b, MPS22a, RBB^{+}24, YH24]. origin [KN21, TZY^{+}23]. original [HZ22]. originality [SW20]. originally [BDTC20]. originate [WZ25]. origins [SCU21]. ORMS [Jab25]. orthodontic [DRF21]. orthopedic [CMZ^{+}20a, CMZ^{+}20b]. osteopathy [The21]. Other [Cop20d, Ben25, BWH20a, BWH20b, BM24, CtRbQ22, LZY22, LM20, Mor19, PA21, RAF25b, TZY^{+}23, dSN23]. our [BGAMS23]. outbound [VdOS^{+}20]. outbreak [WL21a]. outburst [Pal21]. outcome [FGB25, SZK20]. outcomes [AAB^{+}24, CGJM24, GKDB25, HB21a, WJZ24]. outlets [UA21]. outlier [ZDLR21]. outliers [DDL23]. outline [VCV22]. outperform [BT25]. outperformance [AHH21]. Output [SZP^{+}24, ADM22, Azm22, Cao20, Cer24b, CB25, CL20, CF23, GRTPAJ21, GCGB^{+}24, HHB^{+}20, HWL23, HL24, LCD20, LPY22, MDGR21, PL22a, PWZ25, PT20, PT24, PSS23, RPCF20, SVY22, SPS20, SSB22, TZ24, YYC22, ZS21, dGIL21]. output-level [SVY22]. outputs [Hir22, HNM^{+}24, KKYR24, SGM22, WGLN23]. overall [Hug24, NLS24, SNK22]. Overcoming [Öz24, LCP22]. overdrive [vR21b]. overlooked [LZ24b]. overseas [Ayk21, HLT25, ZL23c]. overtime [GCH^{+}22]. overview [AUM21a, AUM21b, IHB21, LCC^{+}20, LB24b, LTH20, Old22].

package [AL24, Bös21, Le 24]. packages [DBR21, WL24]. paid [GRSFV22]. pairs [NC25, WPSK25, WKC24]. Pakistan [MQS^{+}20]. Pakistani [UA21]. Pandemic [MA21, ARR21, BTP24, CP21, DDVV20, DR20, HB20, LDAAAB21, SB21, TTB22]. panel [BR25, CMV23, PT20]. panels [ALR24]. panic [Ali22]. panoramic [ZS20b]. Paper [GVK24, WZZ^{+}21, Wan24b, AQK^{+}22, BW20a, BW22, CLRMR22, Day22, DS21d, GK22, HIQ^{+}20, HMW24, HWW^{+}24, HTZR24, JMZ^{+}23, KA24, LXS22, MLD21, NC25, PM20, Pra22, Pra23a, SdSFB22, SXLC21, SLY22, SMZY23, TZ24, WPSK25, WB21, Wra20a, WXD24, YG20, YTZ20, YWB22, YWHS23, ZWC22, ZSSH22, ZS24, dS21a, dSVN24, EGP22]. paper-entity [SLY22, YWB22]. paper-level [ZS24]. Papers [GMCRV21, ÁLGBdMA24, AYS24, AS25, BZZ25, BW20a, BdCM21, BDTC20, BMS21b, BWE24, BM20b, Cho21, Cop20d, DLC21, DDH21, FCW22, GWLY24, GCH^{+}22, GFVK21, Gon23a, GKD23, HH21, HM20a, HB21b, HMW24, HM20b, HY20, HYZ23, JHVC23, Kal20, KKPJ23, KVZ20, KG24, LCB20, LCW^{+}20, LLC22, LTL24, LLP21, LH22a, LH22b, LZF^{+}24, LLZ21, LYLC23, LM20, MBL21, Man24, dMMdCM22, MWH20, MFM20, MKI24, MN23, MBH25, NT24, Naz25, OPMRP20, QLW25, RJ20a, San25, SM22a, SMJ21, TDWW24, Tei24a, TFWC24, TA20, VEGSMMC22, WZL22a, WSY21, WXZ22, Wan24b, WLS22, WG22, XDL22, XDLZ23, XXL25, Yam20, YF24, YLf25, YP20, YMLL23, ZL23a, ZL23b, ZZ24, ZW24a, ZX25, ZZ25, ZXL22, ZWZ22, ZW24b, ZYL25, ZL20b, dSN23, KZV20]. paradigm [MSP23a, MSP23b]. Paradigms [Wra20b, ZWD20]. paradox [Bor22, MS23, SWL^{+}23, WR21]. Paradoxical [Naz24a]. paragraph [Thi20]. parameter [BB23]. parameters [UMA21]. parametric [SV20]. Part [THJ^{+}24, GMPRSFV25, KW24, Pra21a, THJ^{+}23, Vin21, Vin23]. part-impact [Vin23]. part-set [Pra21a, Vin21]. Partial [dSN23]. participants [YWZ^{+}23, YNL25]. Participation [NB20, CL21a, LOGS21, SOSC24, SW24, SM20, San21b]. particularism [KJY24]. partition [Liu24]. Partner [WGLN23, LCS25, QZF22]. partners [MR20, WGLN23]. Partnership [BMS21a, RP24a]. partnerships [SDW^{+}20, YT25]. party [HLS21]. passivized [Li22]. past [CD22, LFV21, Li22, MPR21, MPMR21, RWL23, RAG25, dRRJ20]. Pasteur [GWLY24]. Patent [KYPC20, LS21a, SHZ22, WZZ24, ACDS22, AMS21, CXZ^{+}20, CCZ^{+}20, CL20, CPL21, DWM21, GWLY24, HYSY23, HZ21, CLY^{+}22, JCZ22, KS22b, LHK21, LHC25, LWW25, LGMK21, Liu24, LXZ^{+}20, LTXZ24, LCT^{+}24, MPS22a, Mey25b, MKI24, NC25, NCD22, OCKY20, Pra20a, QZ20, RSM24, RAL22, Sch25b, SMA21, TZ24, TM25, VCV22, Wad24, WH23, Wan25, WPSK25, WYL21, WXD24, WZ25, YL24, YN23, YMD20, YKS22, ZL23c, ZYL25, ZMK22]. patent-cited [GWLY24]. patent-level [AMS21]. patent-paper [NC25, WXD24]. patent-to-paper [TZ24]. patentable [dMALIGBM20]. patented [SHZ22, SLL23b]. patenting [KS22b, SM20, Ste21, WSL23]. patenting-not [KS22b]. PatentNet [RAL22]. Patents [MAN22, ACDS22, BGSRF20, jDXjS21, GSOCS21, GSP21, LZL22, LZD^{+}24, LZF^{+}24, LL24b, LLZ25, OMFJ22, RKCA23, SG22, SAK22, Wad24, WZZ24, WFSZ24, XLY21, XCML25, ZBG^{+}24, ZX25, ZDLR21, vRW25]. path [Fil21, HL21b, Kua20, Kua23, KL24, LCT^{+}20, LLH20, OJKY23, XHA^{+}20, YS20, YY21, YY22b, YWL25, ZZ20b]. pathology [HWC22]. paths [AA22, DDVV20, ECD24, Fil21, GBP21, Kua23, KL24, LZF^{+}24, WZZ24, YS20, Yur24]. Pattern [YP20, ARR21, Che18, Che20, CXZM25, LH22b, LZ24b, SLX21, VC20, ZM23a]. Patterns [AAB^{+}24, BOV21, GC22, Ali25, BRJ21, CLS21, CLZ21, CMK24, DDD^{+}21, DDH21, DSL23, FWDQ20, FGCGFC20, GLK^{+}23, GDG^{+}23, HF22, HKV20, ICB24, JCK22, JCZ22, JYT21, JCX22, KLLG20, KLJS25, KHK21, LHC25, McC24, NA24, PB24, Pea21, Sha21, SYK22, UA21, Wad24, WCM20, WZY21, WZJW24, WFW24, WZHW25, WG22, YLCY24, YCYZ24, ZZS^{+}20, ZF22, ZPH21]. pause [PHBG21]. Pay [FM21, GRSFV22, SLK^{+}23]. pay-for-productivity [SLK^{+}23]. paying [ZWHS22]. PCT [ZL23c]. peak [YG20]. peaks [DDVV20]. pecuniary [VdOS^{+}20]. pediatric [SMBD^{+}24]. Peer [BPH21, Fur23, Ste22, BAE24, BMA21, BDPP23, Bor20a, BB20, Day22, GRSFV20b, GRSFV21c, GRSFV22, GKDB25, HTX23, Hug24, Jia21, KKPJ23, KHT22, LR23, MM24, MS22c, NZL21, PMB^{+}23, PJKK24, SH08, See20, Wan24b, WZNL23, WWP20, YLX25, ZXL20, SH21]. peer-review [BDPP23, Day22, PJKK24, SH08, SH21]. peer-reviewed [KHT22, MS22c]. peer-selected [Wan24b]. PeerJ [ZXL20]. peers [AS25, BAE24]. penalty [XS25]. Peninsula [SOdLGB21]. pennant [ATW20]. Perceived [SMS23]. Percentile [PD20b, SVY22, BW20b, YY22a]. perception [CBP^{+}24, Jam24, Wij21]. perceptions [ÁRMS21, ES22, PG24, RS22a]. perform [ZWHS22]. Performance [ADS20, AMMN21, AJB^{+}21, ADH23, ADGS25, AS24, ADU22, BWLX20, BSFS^{+}22, Bor20b, BCG22, Can20, CT20, CCH20, Cha21, CWLS23, Cho24, DDD^{+}21, DLK20, FVGBPA23, FBDM25, FJO20, GCT^{+}20, GM22, GG22, GCHT20, HHKZ20, HL21a, HLS23, JXY^{+}23a, JXY^{+}23b, JC22, KN21, KN22, LSA22, Lee24, LY20, LWW25, LCD20, LWX21, LGXH24, LZ21, MTD^{+}23, Mar24, MMGP25, MML^{+}20, Ove23, ÖT24, PR22, PA22, PSA22, PKA24, Pra25a, RS22b, RBL20, RP20, RBC21, SD20, Sca25, SG22, SC20a, STM^{+}24, TMNH^{+}24, UNK23, VBF24, WZG21, WP22, XCT20, XZ23, XZZ22a, XZZ22b, YLM24a, Yur23, ZWL23, ZYL25]. performance-based [BCG22, DDD^{+}21]. performances [LCY21, LCY22, Pin23, YTM21]. period [CKT21, RS22b, SOdLGB21, SMJ21, ZQML23]. periods [OSS^{+}23]. peripheral [HM21]. peripheries [CNP21]. periphery [Gub24, MJ22, MJ23]. perish [Fur21, vD21]. permafrost [Bor21]. Perovskite [YTJ23]. Persian [KMM22]. Persistent [SH08, SH21]. personality [DHSS20]. Personalized [ZSX24, DWM21]. perspective [AMIK23, Bha21, CPFVPA25, CSJW22, CWL^{+}20, DDH21, For23, GXC25, Gon23b, HFS22, HJCT23, JCZ22, KL22, KZZ^{+}20, LXY23, LT20, LMSNZG22, LOGS21, MY25, MJ22, MJ23, MFF^{+}20, Pra20b, RBB22, SM22b, SKS^{+}23, SLL23b, SZY24, WZJW24, WG20a, WZZZ20, Xie20, YXQG21, YLCY24, YCYZ24, YY22b, lYnLyO^{+}20, ZMP24, ZYZW23, ZGBH20, ZLL22]. perspectives [Ara20, BA24, CWLS23, CGA21, Hao24, HP20, HYY25, Lin20, ZS21]. pertinent [MS22c]. Peruvian [ADDCMG^{+}25]. pervading [Tut23]. Péter [Sch20b]. Petre [Ben22]. Ph.D. [BSFS^{+}22, Don21a, Don22, RBL20, SO21]. pharmaceutical [RvdSvMK20, WXD24, ZO20]. Pharmacology [HVGP21, GSH24, SYL^{+}24]. Pharmacy [HVGP21]. PhD [AL20, BM20b, FT21, GSOW20, QSYL23, YWW22, vdBM24]. PhD-granting [FT21]. phenomenon [ER24, Fas21c, NSM23]. pheromone [dSVN24]. Philanthropic [LHF23]. Philosopher [FLC22]. Philosophical [SLX21]. philosophy [AM21b, CC22, DK21b, Ven24]. phone [YRP21]. photovoltaic [LGMK21]. photovoltaics [HAL25]. phrasal [PY24]. phrase [ZY20b]. phylum [McC24]. physical [CKMY20, HGDRM20, Nuz20, Nuz21]. Physics [HMW24, Bjø20b, Coc20, HF22, KNT20, PRGTOP^{+}24, RL20, WGW21, WLZ^{+}23, WZHW25, YGO22, ZY20a, dRRJ20, DCL23]. physiology [HF22]. physiology/medicine [HF22]. picture [CLP21, CM22]. PIDs [HLSX20]. piecewise [HCL22]. pilot [KA20]. pioneer [LBW^{+}20]. pipeline [WZZ25]. PIs [LB25]. pivot [LCB20]. pivoting [WSL23]. place [Ant20]. placement [FCC22, YWW22]. plagiarism [AL20, Ali21a, Ali21b, Lin20, OG24, PJKK24, Pup21]. plagiarized [AL20]. Plan [STAÁMR24]. Plancks [BKO25]. planning [OCKY20]. platforms [CB25, DSB22]. play [JI22, Sch20b, VC20]. players [GAPR20]. Please [BAE24]. PLOS [BA22a, DLC21, DD23, Cop20d, WZNL23]. plots [CLC23]. plotting [Le 24]. plugins [KBS20]. PlumX [KBS21]. pluralistic [Pol20]. plus [SSK21]. PM [WRT20]. PM-bound [WRT20]. PNAS [AHRH24, Mil20]. pneumonia [GAPR20]. point [EGR22]. points [GAPR20]. Poisson [FD21]. polarisation [CMMO20]. polarity [HZW22, XDLZ23]. policies [dLBdAPUdLB^{+}25, BL20b, DMR22, GKK23, KG22, PROG21, RL20, SS22]. Policy [AMZ24, HAR24, AU20, BA22b, DST23, DSB22, Fog23, HZJ22, JI22, Kaj22, Kim24, KFLS20, KL22, Lam20, LMH22, NBBL24, RNB22, YH24, YCXY20, YMLL23, ZHH23a]. policy-related [KFLS20]. policymaker [MLP^{+}20]. policymaking [NBBL24]. Polish [Plo25b, TMNH^{+}24]. polite [BAE24]. political [DMS20, DHSS20, GDG^{+}20, GDdZ22, GTH24, GDdZ25, GTB23, HLS21, LW24, MMG21, NSM24, YLW20]. politics [SMM^{+}20]. ponds [MHK22]. poor [DTHM21, Kwi20]. Popper [BWH20a, BWH20b, BF25]. popular [SHB25]. popularity [HKN23, HKNF23, ZW21b, ZCLW20]. population [CTH22]. portal [SNK22]. portfolio [NLS24]. portfolios [TTDH25]. portraits [SMD22]. Portraying [RTT23]. position [MAN22, SCK21, TMNH^{+}24, WCH^{+}20]. positions [HSE23]. Positive [WH25, AHH21, CCSX23, EBW23, WWL24b, mYY22]. positive-unlabeled [CCSX23]. positivity [LZ23, LZ25b]. possibilities [RS22b]. possible [AGB23, AYRG23, Liu23, ZCS^{+}24]. post [ADGS25, Bor20a, CLP21, KDIR23, LPY22, MSL22, MBF23, Mit20, SYHW20, Wan24b, YWHS23]. post-publication [Bor20a, Wan24b]. post-retraction [SYHW20]. post-Soviet [ADGS25, CLP21, KDIR23, LPY22, MSL22, MBF23]. post-WWII [Mit20]. Posthumous [Kos21b]. posting [BTP24]. potential [ADD24, DG21, LLH25, OMCC23, QJI^{+}22, WSZ21, WZY21, WSC^{+}24, XYY25, YZZT24, ZTY^{+}24]. potentially [MS21b, MS21c]. Power [CL20, Asa20b, MNM^{+}21, OB21, RAS21, Tom24, UKP21]. Power-law [CL20]. power-weakness [UKP21]. PPPs [SDW^{+}20]. practical [DvE20, DMR22]. practice [BPL25, BA22a, EBB^{+}20, HB22a, HTX23, LRSB21, Naz20, PROG21, PCG^{+}20, RJ20a, Sch25a]. Practices [SMS23, DV20, DMS20, HKW24, SdSCdS25, VWR25, YA21]. Prathap [SD23a]. pre [CA22, LCT^{+}24, SLL^{+}23a, Wan24b, ZNNA20]. pre-and [Wan24b]. pre-prints [CA22]. pre-trained [LCT^{+}24, SLL^{+}23a, ZNNA20]. precision [FNM^{+}22, RBB^{+}24, WZY21]. Predatory [MS21a, MS21b, MS21c, Mou21b, Mou21c, WLMZ24, MS22a, AAA^{+}23]. predications [LPD21]. predict [AJLA21, DMA^{+}20, HWX22, dW24, dRRJ20]. predictability [WWL24b]. Predicting [CLS21, CDG20, DWG23, DXL21, DG21, GSOW20, Kup22, LWL24, OCKY20, SLTS25, YZZT24, ZW24a, ZWZ22, ASIN24, MBL21, MPS22a, XLL20, YZWH24]. Prediction [WSZ21, YN23, ASK22, CLS21, CGA21, DG21, HB23, JADG20, MLD21, MCDFSP24, NZ23, OCKY20, QZF22, TA22, VA22, WL21a, WZZ^{+}21, WWS23a, WLS22, XLL23, XZZ22a, XZZ22b, YLH24, YKS22, ZYWA25, ZCLW20]. predictive [AFH21]. Predictors [DRF21, AHH21]. Preface [CFM20, CL21c]. prefecture [ZBG^{+}24]. prefecture-level [ZBG^{+}24]. prefer [Kua23]. preference [DS20, DS22, MZZL21, ZWL^{+}25b]. preferences [CY25, Cop20b, JMZ^{+}23, LBCO21, SHL22a]. preferentially [Tei21]. Preliminary [HKV20]. Preprint [BA22a, BZZ25, COB21, SCP22, TLL24]. preprint-publication [COB21]. preprinting [KHT22]. preprints [BTP24, KHT22, LYZ^{+}20, WGC20]. Presence [EBW23, RFMMPJ22, FCT^{+}20, GMCRV21, KBS20, SZN^{+}21, Mor20]. present [BH25, MSJ23, RdNB21]. presentation [Sug23, vR21a]. presenting [RS22b]. presidents [CMT24]. Press [Pet22, JVY22, Lin21b]. pressure [Ell23]. prestige [ES22]. prestigious [WLLW25a, Zha21b]. presumptions [MLP^{+}20]. pretrained [NJCL20]. prevails [IHLP21]. Prevalence [GSH24, MS21c, FCY23, MS21b]. Price [Ano20a, Ano21a, Ano23, GDRPR23, Sug23, vR21a, Wal22, CZZN25, HZC21]. prices [Wal22]. primary [Seb23a, Seb23b]. principal [Cer24b, HLS23, ODQ24]. principals [HGB20]. principle [vD21]. principles [SCFSAB^{+}25]. print [Mal22a, Mal22b, STM^{+}24]. printed [LYZ^{+}20]. printing [BWLX20]. prints [CA22]. prior [SLR^{+}23]. Priorities [KKYR24, BL20b]. private [FMF20, FYY21, Kim23, Sha24a, SDW^{+}20]. privatization [MYX22]. Prize [Bjø20a, Bjø20c, AKD23, ACT23, Bjø20b, CSC23, HMW24, KJY24, LRH22, LLH22, LLSH23, PS22, Rod22, TTB20, ZZ23]. Prize-producing [ZZ23]. Prize-winning [LLH22]. proactive [Lin21a, Poh20]. probabilistic [CRBRGS21]. probabilities [YLC^{+}20]. probability [BF25, KKOL20]. problem [GRBAM22, LLC22, MJRR25, ZZ24]. problem-solving [GRBAM22]. problems [Bi23a, Bi23b, Bi23c, HB21b, Hüc22, HL24]. procedure [GdMDD24, GAB21, LR23]. proceedings [Pur21, San25, WXZ24]. process [BB20, CMZ^{+}20a, CMZ^{+}20b, HC22, HL24, LMLMPA22, WGZ22, YMLL23, ZXS22]. processes [BF22]. Processing [Ano20c, AÁSTMR24, Asa20b, Asa23b, Asa24, BGWZ20, CFM20, EXT^{+}21, IHB21, KP23b, MS22b]. produce [CGDS24, SO21]. produced [UA21]. producers [KLLG20]. producing [WLZ^{+}23, ZZ23]. product [JL25, Mut22, OCKY20, ZGM^{+}23]. Production [MN21b, AAAA23, BERD21, BRJ21, Cha23b, CLZ21, ELR24, EGPGGA^{+}23, JLL21, JCK22, LOGS21, NC20, PRGTOP^{+}24, SV20, SSdCRD22, YP20, dCFdCA^{+}23]. productive [GDdZ22, HPL21, KP23a, KK20b, SAR19]. Productivity [FWDQ20, Jok20, ZF22, AlYT21, ALM^{+}23, Bi23a, Bi23b, Bi23c, Bor24, BSA23, CF22, CW25, DJM22, DM25, DCdA25, DHSS20, DXL21, FRS23, FBDM25, GSOW20, GLPC20, GCS25, GDG^{+}23, HLS21, HPH21, JYT21, KBSS23, Lam20, LP23, LW24, LL24a, MHD21a, MHD21b, MN24, Par25, PLW23, PROGDG24, RNAH21, RPL21, SLK^{+}23, SHL22c, SHL22b, SdL21, SDRS23, TQA21, THG23, XS25, XLL25, YG20, YGO22, YY22a, ZPH21]. prof [LRW25]. Prof. [LGD21, vR21a]. profession [SGM22, VPC^{+}23]. professional [DRF21]. professionalism [KCC22, SJW21, SJW22]. professor [LZC21, ZLL21, SD23a]. professoriate [MRM20]. professors [WZG21]. professorship [ADH23]. profile [Hen19, Hen20, LRP21, PA22]. Profiler [VPR22]. profiles [ALR24, FH22, HSE23, MN21b, Mil20, San20, WLHY24]. profiling [BGSRF20]. profit [VdOS^{+}20]. profits [GBH21]. program [BSFS^{+}22, Bor24, DLL21, MOMGNLG24, PSW24, ZM23b]. Programme [CFI22]. programmes [XYY25, ACV23]. programming [WL24]. Programs [KPM22, CL21a, FT21, KHC^{+}25]. progress [Pin23, RSM24]. project [AÁSTMR24, LHX^{+}24, MLP^{+}20, MS21f, XDW22, MF20b]. projections [Bat24a]. projects [OPS25, PROGDG24, PVP^{+}23, XDW22, XLL25, ZLL22]. Prolific [Ioa23, BSA23, CY25]. Prolificacy [MZE20, HH21]. PROMETHEE [EE24]. prominent [San24]. promising [KG20, SSK21]. promote [SHZ22]. promoted [LBP22]. promoter [Ben22]. Promoting [YWW23, HL24, PROG21, WWP20]. promotion [Ali25]. promotions [DFRS20, MPS22b]. Prompt [BOPGOP25]. prompted [HB20]. prompting [FAY25]. proof [DLC21, Don22]. proof-of-concept [Don22]. Propaganda [TG20]. propensity [ZXL20]. properly [ZZL^{+}20]. properties [Coc20]. property [AL21, GXC25, SB25]. proportions [KG24]. proposal [ADDCMG^{+}25, BW20b, CdSK21, DvE20, KP23b, LR23, MBH25, SMT^{+}23, SdVvdR24]. proposals [JI22, LWZ22]. proposed [BDTC20, KL22, XlYqZ22]. protection [AL21, GXC25, SB25, SHZ22]. protocols [OMAAORCL23]. provided [THJ^{+}24]. providers [MBH25, Ort20]. providing [HMPS24]. proximities [TG24]. Proximity [KSS21, MH24a, ADD20, AAD22, AAD24b, FFL21, GYWZ21, QQL21]. proxy [CFI22]. psychological [Jia21, MSJ23, SMJ21]. psychologists [Ben24, LSA22]. psychology [GDdZ22, GTH24, RTT23, WVH21, WUH21]. Psychosocial [SOSC24]. Public [FMF20, PROGDG24, PSH22, AÁSTMR24, AU20, BCF21, BKO25, DST23, DCdA25, DMR22, Fuk22a, Fuk22b, GC20, HIQ^{+}20, KBSS23, KHT22, LdGRV^{+}22, MMG21, MMG23, MYX22, NBBL24, RS22a, SDW^{+}20, XlYqZ22, ZZS^{+}20, ZXL^{+}24]. Public-private [FMF20]. Publication [AHH21, ARR21, FM21, LZ25a, SZP^{+}24, WPBH24, AJB^{+}21, AD20, ADH23, Ali25, AAB^{+}24, BC21, BPLIdMA^{+}20, Bor20a, BH25, BOV21, COB21, CMZ^{+}20a, CMZ^{+}20b, CZZN25, CG24, CF23, DvE20, DDD^{+}21, DD23, DXL21, ES22, Fan20, FN23, GKK23, GDDG24, GDdZ22, GTB23, GCHT20, HS22b, HB22a, KL22, KPS21, KUS22, LCW^{+}20, LBCO21, LCD20, Liu21, MSV23, Mit20, MA21, MDGR21, Mor20, MNK21, PWZ25, PD20a, PLW23, PROG21, QSYL23, RGVQTS^{+}24, RBL20, SF20, SRL20, SH20, SGM22, Seb23a, She20, SMBD^{+}24, SA24, SYK22, SP20, ŚT21, TTDK22, TWH^{+}24, VC20, VRVH21, Wan24b, XAA21, YGO22, YY22a, ZW21b, Zha21b, ZPH21, ZWZ22]. publication-based [AAB^{+}24]. publication-citation [Fan20]. publication-level [AHH21, SA24]. publications [ADG24, ABAAN^{+}23, Ali24, BD22, BdOdS21, BMBA20, BKT20, BF22, BA22b, Bou21, BGSRF20, BW22, BPW^{+}21, Cer24a, CLP21, CDG20, DMR20, Don22, DMGBG20, EO24a, FCT^{+}20, FT24, FYY20, FC20b, FC20c, FN23, GKDB25, GCS25, GZOC24, GMKG20, HAR24, HAS^{+}20, HMW24, HGM21, HGDRM20, HZW22, IVCE20, JKK24, JKY22, KBS20, Kou24, KSS20b, KN22, LMC22, LKPS25, LXY23, LZYR24, LS21c, Liu23, LDA22, LUH21a, LUH21b, Mad20, MMGC^{+}24, MS22c, MCSOAA21, MR20, MNP21, MMN20, MNM^{+}21, MA21, Naz20, Naz24a, OMFJSC24, ÖT24, PRS^{+}21, Poh21, PVA24, Pra20a, Pur21, QWS^{+}23, RdNB21, SO21, SK22, SdS25, SdLV21a, SB21, SGC21a, SGC21b, SSSd22, SQC^{+}20, SSL22, SR23, SD22, SJW21, SDRS23, TA22, TdS20, Tut24, UBROGS23, Vin21, WZL22a, WJZ24, XLY21, XAA21, YNO22, YNL25, ZW25]. publications [ZLL21, ZWZ^{+}19, ZAF20, ZSSH22, ZF22, ZCLW20, ZWX22]. publicity [WZ21]. publicize [KKPJ23]. publics [TSDAMRG24]. Publish [Fur21, QSYL23, Bjø20a, Bjø20b, Bjø20c, Fag20, FC21, KR22b, TKM^{+}23, vD21]. publish-or-perish [vD21]. published [BdCM21, BDTC20, BM20b, DLC21, FNK23, DR20, Gon23a, LYS22, LHW21, LPY22, PMB^{+}23, RJ20a, RAF25b, SMJ21, SCP22, Tei24a, Yur20, ZL23a, ZL23b]. publisher [CF23, GMRVU24, Kim24, Wal22]. publishers [Asa20a, Asa20b, Asa21, WZ21, Zho21, dSDE21]. Publishing [CF23, KHT22, dSTE21, ADH23, AAA^{+}23, Ali22, Asa21, AYRG23, CY25, GDM22, HWX22, ICB24, JA23, JP22, KFM20, KL22, KK20b, LZ24a, Lin21b, MS21a, MS22a, MB21, MS21b, MS21c, MSJ23, MBD25a, MBD25b, MMB23, MFF^{+}20, NRGvSTS23, NSM24, PS20, SBF23, SL24, SC20a, Tei20, The20b, The23, UA21, WB21, ZY20c, Tei21, Tei24b]. Publons [ZSS22, ZTY^{+}24]. PubMed} [SdLV21a, Bös21, Bou21, BST22, CKO21, LTL24, LML21, ODQ24, TLL24, WG22]. PubMed}-based [SdLV21a]. PubPeer [MMGC^{+}24]. punctuated [dLBdAPUdLB^{+}25]. purpose [HZW22, KMGS21]. pursuits [ADD24]. put [Ell23]. Putting [DHSS20]. Python [Le 24].

Q&A [SCK21, YWZ^{+}23]. QCA [FESRMB21]. QS [AMSS^{+}23, AA25, Hla25, MJYS23]. quadrant [GWLY24]. qualification [MPS22b, BDPP23]. qualified [ZY20a]. Qualis [RSF20]. Qualitative [Ano21c, Ara20, BDPP23, CW25, GB21, HP21, LMSNZG22]. Quality [GRSFV21c, ADG24, ADU22, Azm22, BCG22, CH20, CF22, CNA22, DDS22, DDD^{+}21, DDF20, DWM21, FNK23, FZhCC21, FLDD20, Fuk22a, Fuk22b, GG25, GRSFV21b, GSB24, Jam24, KHT22, KG20, LPKG20, LHF20, LWW25, LLZ21, MB22a, MSV23, MS22d, MSP23a, MSP23b, RSV^{+}20, SLMCSV21, Sch25b, Sha24b, SZK20, TDWW24, TKM^{+}23, UNK23, VFL^{+}21, VAMBTS25, VY23, WZL22a, WLLL20, WWL24b, XLL20, XDLZ23, ZPH21]. quantifications [BM24]. quantified [Ove23]. quantify [GD22, GBP21, LSBC21]. Quantifying [BZL21, FC20c, HY22, LBCO21, NWZ23, RSM24, Wan24a, MY25, WYL21, ZJ22]. Quantitative [Ano21c, Ano21d, BERD21, DDD^{+}21, GB21, HHKZ20, BG21, BST22, CKMY20, DX20, GRTPAJ21, HP21, HY20, HX25, ODF20, RJ20a, XYY25]. Quantity [Azm22, CF22, FLDD20, LHF20, WAM^{+}25]. quantity-quality [FLDD20]. quantum [JCK22, WZHW25]. Quarterly [JJ23]. quartiles [VP21b]. quasi [BFG^{+}21, Plo25a]. quasi-experimental [Plo25a]. quasi-zero-day [BFG^{+}21]. quest [BKO25, WSL23]. question [LHCH20, SdL23, SCK21]. questionable [KHK21, ÖT24, SdSCdS25, Ste23, VWR25]. Questions [KCC21, KCC23, HCS21, SCK21, SRL24]. quick [Tay23]. quotation [CSR23]. Quoted [Pea21].

R [BWH20a, BWH20b, AL24, WL24]. R&D [jDXjS21, Gu21, KZZ^{+}20, LS21a, LY20, LZ21, OVS25, RvdSvMK20, STAÁMR24, VdBC20, ZM23b]. R&D&I [KHC^{+}25]. race [KDN22, SP20]. raction [BPW^{+}21]. radical [SvWC21]. radioactive [CHT24]. railways [MH24a]. raising [MHD21a]. ramifications [KS24]. random [Hüc23, PSS23, Sin22]. randomised [OMAAORCL23]. randomly [LDDL24]. Rank [SZP^{+}24, MSV23, MHK22, Naz24b, See20, Vin23, YY22a, GMRVU24, SVY22]. rank-citation [Naz24b]. ranked [KS22a, LFV21, OdS20, OŠV^{+}24, QD21, ZPH21]. rankers [KCS25]. Ranking [CLWlY24, GG25, GC21, KCS25, KPM22, SD20, Sch20b, TMNH^{+}24, UMA21, AMSS^{+}23, AA25, AU20, BZX^{+}23, Bor20b, LNS21, Li25, LWG21, LL24b, LA20, MJYS23, Pra25a, SAA21, SAR19, VPSF23, ZWC22, ZZZZ20]. ranking-related [LWG21]. Ranking-the [Sch20b]. rankings [ATC20, DDL23, DJM22, Fas21b, FSEJF^{+}20, KS22a, LKPS25, Lin24, Pra24, QD21, RGVQTS^{+}24, SD20, SL22, SS20b, ŚT21, TBPN^{+}20, YS24, Hla25, TMNH^{+}24]. ranks [DJM22, ZL20a]. Rao [Zie22]. rapidly [BMBA20, Poh20]. RAR [TBA23]. RAR-SB [TBA23]. rate [LGMK21, PR23]. rates [AG23, LCCH24, Öz24]. rating [CZQC24, Fas20a, Fas20b, dSDE21]. ratings [DS21d]. ratio [UKP21]. rationale [MS22b, OŠV^{+}24]. rationale-cogency-extent [OŠV^{+}24]. ratios [NF21, San21b]. RCE [OŠV^{+}24]. re [Liu23]. re-evaluating [Liu23]. Reach [CF22]. reactions [Jia21]. reacts [ZZS^{+}20]. Readability [WLZ22, WL22, WY23]. readable [JDG21]. reader [BMS21b]. readership [RBR22, TFWC24, dW24]. reading [Pra23a, SD23a, SD23b]. reagents [BPW^{+}21, LCW^{+}20]. real [AL20]. reallocation [XlYqZ22]. really [HWX22, Mal22b, Mou21a, SM22a]. reason [IQ21]. reasoning [WCH^{+}20, ZLL22]. Reasons [KG22, GGS22, Hüc23, Sha24b, ZCS^{+}24]. recall [FNM^{+}22, KMM22]. receive [CLR25, HNM^{+}24, JDG21, LBP22]. received [RAF25b]. recency [WXD24]. reception [ACT23]. recipients [CZZN25, DM25, Wan24b]. Reciprocal [VHM24]. reciprocality [ALR24]. Recognition [CJT22, CTH22, DGGBDG21, FLC22, SSC^{+}23, TDWW24, ZC24]. recognize [HY22]. recognized [SMD22]. recombination [CCH^{+}24, WG24]. Recommendation [LS21b, VHM24, ZJY22, AUM21a, AUM21b, AQK^{+}22, AKYH21, CCZ^{+}20, CLS^{+}20, DWM21, GK22, GKD23, HTZR24, CLY^{+}22, JHVC23, JJPC20, JMZ^{+}23, KA24, LH24, LWZ22, LZ25a, LYLC23, LTXZ24, MZL20, MZZL21, NJCL20, PLT22, QYL21, TZZZ23, TBA23, WLL25, ZSX24]. recommendations [GdMDD24, HZ22, KJY24, WL21a]. recommended [VMF22, WWZW20]. recommender [FCY23, HIQ^{+}20, MLA^{+}23, Pol21, VMF22, WCY22, XWGD22]. recommending [HCN21, LZD^{+}24, YLC^{+}20]. Reconfiguring [SZZ23]. reconsidered [WB22]. reconstruction [HLG21, ZC24]. recorded [TSVV23]. records [ATW20, LHW21]. recruitment [HLT25]. recurring [NF22]. reduce [The20b, XDL22]. reducing [CFA^{+}23, YWW23]. reduction [CCSX23, SCU21, ZC24]. redundant [Cop20c]. Redux [DKWC20]. REF [AJLA21]. REF2021 [KSAT24]. RefCit2vec [cHhC24]. referees [Hug24]. Reference [CHJ^{+}25, HB22a, BC21, CG22, EGR22, GW23, GTB23, MFMC21, SH20, WG22, ZZZ22, Tei24b, WPBH24]. referenced [CZZN25]. References [QZ20, ÁLGBdMA24, BKT22, BL20a, DGGD22, ES22, EBB^{+}20, HGM21, cHhC24, LRH22, NF21, OMCC23, Pra24, Sta21, TTB20, TTB22, Yam20, Yur23, ZLM^{+}23]. referencing [KLLG20, WCM20, WG22]. refinement [JKY22]. reflect [Fil21, THL23, VAMBTS25]. reflected [SG22, ZXL^{+}24]. reflecting [LHF20]. Reflections [LWR20, LMLMPA22, BBP21, HLR21, KFCK21]. reform [Cer24b, RBL20]. Reframing [Kaj22]. refuse [dSV21, dSV23]. Regarding [Kua20]. region [GMRVU24, LPY22, RPAVFV21, WKC24]. region-subject [WKC24]. Regional [AOM24, PSW24, AAD24b, Che23, Coo25, GRTPAJ21, KZZ^{+}20, Li25, MLS21, MMB23, Ove23, PVA24, VY23, WZZZ20]. regional-based [AAD24b]. regions [AL20, HTGW20, MBR^{+}25, TG24]. Registered [LCCH24]. registries [THvB21]. regression [SV20, ZHH23a, UMA21]. regulations [ZBG^{+}24]. reinforcement [SM24b, WCL25]. reinventions [SdSFB22]. reject [Ans23b]. rejection [FLC22, Fur21, Wad20]. rejoinder [Cop20b, HPH21, MHD21b]. Related [ATW20, ASSB20, CLC23, Don22, GCHT20, IVCE20, KFLS20, KSP24, LLC22, LWG21, MFP21, SB21, TDM24, Yur20, ZS21, dSTE21, dGIL21]. relatedness [KJL23]. relates [BGAMS23]. Relating [HKN23, UA21]. relation [Ger24, vRW25]. relational [CZX24, EE24, KDZ21, LMG20, NZ23]. Relations [CW24, RKCA23, ZYZW23]. Relationship [GVK24, LdGRV^{+}22, Bre21, CZFC24, DDH21, FRRBDD23, FLDD20, Gor21, LWL24, LZ23, LGXH24, LHZ^{+}23, LM20, MMF23, Mal22b, Ony20, PLH20, Pra23c, RBC21, SHL22a, SZ20, SSB22, SZK20, TQA21, TP20, WKC22, WC20, ZW24b, dGIL21]. relationships [Ben24, LH22a, NKY^{+}24, SMZY23, YLCY20, ZXS22]. relative [CCMM24, FSR21, MNM^{+}21]. relatively [Wal22]. relativity [LHW20]. release [BMS21b]. Relevance [BST22, HB22b, ANR21, PB24]. relevant [BF22, Yam20]. Reliability [FD21, ZLL22, dOCP21]. Reliable [FBDM25, CCH25, WLHY24, WLS22]. RelPath [GBP21]. remains [FG20]. Remembrance [Ben22]. remote [VCD^{+}23, dSMNdV^{+}21]. remove [FBM21]. renewal [LB25]. reorientation [Reh20]. repeat [PR23]. Repeated [TM25, SHF24]. RePEc [Lin24]. replications [CLR25]. Reply [MHD21b, YWB22, JTB23]. Report [ATC20, LCCH24, SYHW20, WY23, KPM22]. reporting [OMAAORCL23, RSV^{+}20]. reports [BRG^{+}23, GSG22a, HDY^{+}23, MM24, OMAAORCL23, OG24, PMB^{+}23, PJKK24, SNK22]. repositories [OMFJSC24, WPS22, Agu20]. repository [KBDS23, RBV25, THJ^{+}24]. repository-provided [THJ^{+}24]. represent [RZS23a, RZS23b]. representation [AS23, DL23, GDG^{+}23, GTH24, HCL22, LS21b, LTL23, LLZ21, LYLC23, MKWB^{+}23, PJ24, ZY20b, Tei24b]. representations [KR21, QYL21, YJJ25, Pet22]. representative [HBO^{+}20, WZZ23]. representativeness [KSS20a]. represented [BVvEW25]. representing [LPD21]. Reproducibility [CA22, TdS20, KR21]. reputation [CW24, MR20, MZE20]. reputational [SD20]. required [Kra20]. requirements [ADH23, QSYL23]. requisite [ZCS^{+}23]. Rescaling [YGW^{+}24]. Research [AKYH21, ADDCMG^{+}25, BWLX20, BMM22, CRC20, CFI22, EO24a, Fas21c, Fas24, Ger24, GDG^{+}23, dMGC23, HIQ^{+}20, HG20, HL21a, KA24, KP23b, LSA22, Lam20, LBP22, LRP21, LHC21, LH22b, LXZ^{+}20, LYLC23, LCT^{+}24, MYK22, Mar24, OMBP24, PMB^{+}23, Pea20, PT24, PL23, PA22, RPL21, SHH23b, SMG^{+}24, SYK22, SGG20, TA22, TPVVPA21, UIM21, Vie24, WWY^{+}20, Wei20, XLL25, YWW22, YLW20, YT25, AMMN21, AMSS^{+}23, AJB^{+}21, ADD20, AAD22, ADD22, ADM22, ADH23, ADD24, AAD24a, AAD24b, ADGS25, AS24, AM21a, APR19, AJBB22, AJBBA23, AAAA23, Ali24, Ali25, dOSAL21, ABK20, Ant20, AJVACC22, Asa20a, Asa21, AU20, Avd21, ÁRMS21, Ayk21, Azm22, BR25, BLL21, BD22, uHBKK20, uHBKK21, BP20, BD21, BB22, BGP20, BJS22, BKS20, Bi23a, Bi23b, Bi23c, BKT22, BSFS^{+}22, Bir21, Bjø20a, Bjø20b, Bjø20c]. research [BERD21, BS20, BFG^{+}21, BDPP23, BCF21, BRG^{+}23, Bor21, BGAMS23, BL20a, BM20a, Bor24, BVS25, BCG22, BKO25, BSPR20, CCdC20, CNP21, CH20, CTLH24, CGJ22, Can20, Can24, Cao20, CLL22, CZX24, CFA^{+}23, CMV23, Cer24b, CW25, CB25, CCH20, Cha21, Cha23b, CMMO20, Che24, CDZ^{+}20, CNA22, CWLS23, CxGdH^{+}24, CGDS24, CC22, Cho20, CLZ21, CCC22, CKT21, Coc21, CR24, Cop20b, Cor22, CBP^{+}24, Dag24, DV20, DV22, DST23, Dej25, Dem22, DJM22, DRS20, DFS22, DX20, Dia21, DLK20, DCL23, Dip21, DHSS20, DDF20, DRF21, DLL21, DS20, DGGD22, DSL23, DLS23, EXT^{+}21, EBB^{+}20, EGP22, FMF20, FWDQ20, FYY21, FCT^{+}20, FFN25, FT24, FRS23, Fog23, FMB22, FH22, GG25, GCGB^{+}24, GC20, GD22, GCT^{+}20, GAPR20, GDG^{+}20, GDdZ22, GLPN25, GKD23, GG22, Han20, Hao24, HAT23]. research [HBW^{+}21, HIA^{+}20, HWZ21, HBP20, HVGP21, Hir22, HL21b, HJT21, HHB^{+}20, HFS22, HKNF23, HhC20, HJCT23, HTGW20, HZJ22, HY22, HDY^{+}23, HWL23, Hua24, HNM^{+}24, Hüc22, Hüc23, HL24, Ioa23, JA23, JHVC21, JL23, JJ23, JDG21, JXY^{+}23a, JXY^{+}23b, JVY22, JI22, JC22, Kaj22, KSS20a, KNT20, KKPJ23, KBS20, KDIR23, Ke23, KR21, KBSS23, KK20a, KFLS20, KSP24, KZV20, KVZ22, KSS21, Kos23, KG22, KPS21, KKYR24, KCC21, KYKC21, KCC22, Kwi20, KR22a, KR22b, LGD21, LBCO21, LMC22, LNS21, Lax23b, LS20a, LHXS20, LCY21, LCY22, Lee24, LKPS25, LSFLZLLV20, LRSB21, LYT^{+}20, Li22, LMH22, LRH22, LCL23, LL25, LB25, LGLM20, LCC23, LCCH24, LS21c, LBW^{+}20, LSY21, LWZ22, LZY22, LYZ22, LZ23, LYZ23, Liu23, LZ24b, LGXH24, LMSNZG22, LPY22, LCP22, LdS25, LDAAAB21, LMLMPA22, LMLR24, LC20, LGR^{+}21, LC21, MTD^{+}23]. research [MH24a, MMF23, Man24, MPR21, MPMR21, MMGP25, MHD21a, Mar20, MK23, MSI21, MLP^{+}20, MN21a, Men25, MN24, MGCR24, MGA22, Min25, MS21f, MYX22, MDGR21, Mor19, Mor20, MRC22, MKI24, MNK21, Mu21, MBH25, NS23, NSM23, NLS24, NC25, NB20, NB22, NH24, OPS25, Ony20, OB21, Ove23, Pal21, PL22a, PY24, PY25, PWZ25, Par24, PS20, PSKS22, PHBG21, PG24, Pea21, PD20a, PDH21, PVP^{+}23, Pin23, PT20, Plo25a, Poh20, Poh24, PS22, PKA24, PSS23, Pra23c, Pur24, Pur25, RG20, RNAH21, RBBP24, RKMG20, RPCF20, RS22b, RWL23, RA20, RFG^{+}24, RNB22, RvdSvMK20, RP20, RBC21, RL20, RBB22, RvLR^{+}20, RAG25, SD20, SLK^{+}23, San21a, San21b, SSS^{+}24, SHL22c, SHL22a, SHL22b, STAÁMR24, SO21, Sca25, SB20, Sch20a, Sch22, SdLV21b, SdL23, ŞKKT25, SAS23]. research [SSZ22, She20, SDW^{+}20, SW24, SPS20, SSB22, SNK22, SCFSAB^{+}25, SD22, SRL24, SJW21, SJW22, SCW^{+}23, SDRS23, SFC^{+}23, SZK20, SSBC22, SPA20, TKS^{+}23, TG20, TKZ^{+}23, TNS20, TNS21, Tay23, TdSFB20, TD24a, TSVV23, TBA23, TMW23, TNSF21, THG23, TDM24, TdS20, TTDH25, UA21, UNK23, VCJ22, dGdO^{+}23, VWR25, VKY23, VDY21, VC22, Vie22, Vie23, VC20, VP21b, WR21, WWH21, WJ22, WZL22b, WLZ^{+}23, WH24, WJZ24, Wan24a, WLLW25a, WH25, WZHW25, WLBC20, WUH21, WSRM23, WCH^{+}20, WPBH24, WGZ22, XWGD22, XCT20, XL20, XLL20, XHA^{+}20, XS25, YD21, YK23, YWX23, YP24, YLM24a, YZWH24, YWN25, YWW23, YKTM20, YY22b, YS23, YWZ^{+}24, YLX25, mYY22, lYnLyO^{+}20, ZIS21, ZZS^{+}20, ZSG21, ZCW21, ZWHS22, ZGM^{+}23, ZT23, ZMP24, ZXL^{+}24, ZCSK24, ZSL24, ZW24a, ZTY^{+}24, ZWL25a, ZS20b, ZPH21, ZWL23, ZPZ21, ZJ22, ZCL20, ZL20a]. research [ZLL22, ZHH23a, ZHH23b, dFBGCGASG24, dGIL21, dSMNdV^{+}21, vdPF25, HYSY23, SHH23a, SYK22, DS22]. research-based [AU20]. research-intensive [Men25]. research-teaching [Sca25]. Researcher [For23, OTH22, PLW23, TK21, Cho24, DV20, DGGBDG21, Dri24, FD21, FBDM25, HKNF23, Jam24, JMZ^{+}23, MMN20, PHP22, VPR22, WD21, XYY25, dS21c]. Researchers [KK20b, AHH21, BT25, BdPFdO25, BWH20a, BWH20b, CDE23, Che23, CMZ20c, DvE20, DDD^{+}21, DGRÁTS23, ECD24, EOL23, FJO20, FGB25, GMPRSFV25, GCMJ25, HPL21, HKN23, KDS21, KKPJ23, KDN22, KKS21, Kup22, Lee24, MS22c, MMF20, MBD25a, MBD25b, MMB23, Naz25, NBBL24, OPS25, OA24, Par25, Por25, PMGÁC23, QD21, SRL20, SHL22c, SHL22b, SDD20, SdL21, ŞKKT25, SG21, SA21, WLLW25a, WGA22, YY22a, YLCY24, Yur20, ZAZS22, KK23]. ResearchGate [GDRPR23, MS21d, SSL22]. ResearchGate-way [MS21d]. researching [The21]. reservoirs [HG20]. ResGAT [HWW^{+}24]. residual [AA22, FLDD20, HWW^{+}24]. resilience [CBP^{+}24, TMNH^{+}24, YWN25, ZCW21]. resistance [KKOL20]. resistance-based [KKOL20]. resolving [KKOL20]. resonance [SS22]. resource [GM22, Kwi20]. resource-poor [Kwi20]. resources [XlYqZ22, KW24]. respect [Cop20c]. respond [MAS^{+}21]. Response [Bjø20c, Lax23a, LYZ23, RJ20b, Rus23, SD23a, SJW22, Cer24a, SdL23, SMG^{+}24, YWZ^{+}23, ZZS^{+}20, Zha23a, Bjø20a]. responses [BPW^{+}21, SBF23]. responsible [GBH21]. rest [JP22]. restrictions [GCGB^{+}24]. restructuring [BVS25]. result [KBS20]. Results [ZC23, BGP20, BDTC20, DTHM21, KMM22, OMFJSC24, SZK20, VMF22, ZGBH20, BBP21]. ret [BPW^{+}21]. Rethinking [CGA21, dSV21, GCS25]. Retracted [Ela21, GGS22, GFVK21, SR23, SP20, TZY^{+}23, AYS24, AS25, Cop20d, Ela22, FYY20, HB21b, HP21, LAT22, LXY23, ODQ24, Sch24c, SYHW20, Sha21, SSBC22, YYZ25, ZAF20, ZF22, dSN23]. Retraction [AAA^{+}23, LHC21, MS22a, Sha24b, AYS24, SYHW20, ZAF20]. Retractions [LL21, Seb23b, Bha21, Cop20d, ES23, FFH^{+}24, FFRRPR^{+}24, LDDL24, LRP21, MCSOAA21, SdSCdS25, SGC21b, Sha21, SGC21a]. Retrieval [Ano20c, Liu23, ZC23, ATW20, BVvEW23, CFM20, SLH24, WZM^{+}24]. retrieved [KMM22]. retrospective [JKK24, LTL24]. Return [ZAZS22, SCW^{+}23]. returnees [ZWL23]. retweeted [CWZ21]. retweeting [MH24b]. Reuben [SSBC22]. reunification [Reh20]. reuse [KTK21]. reveal [HS22b, KSS20a, Sha21]. Revealed [GDM22, Cop20b]. Revealing [KY20, LZH^{+}22, TMZ24, WZY21, BWH20a, BWH20b, VdOS^{+}20]. reveals [Day22, Dem22, LUH21a, LUH21b, YGW^{+}24]. Reverse [FH23]. Review [GMKG20, HJLZ23, XHA^{+}20, AJB^{+}21, ASIN24, AS23, AJBB22, ADU22, BPH21, BAE24, BMA21, Bir21, BS20, BDPP23, Bor20a, Bor21, Bös21, BB20, CMZ^{+}20a, CMZ^{+}20b, CRBRGS21, Cho24, CGA21, Day22, Fas21b, FESRMB21, GRSFV20b, GRSFV21c, GRSFV22, GKDB25, GLPN25, HTX23, Hug24, KKPJ23, Lee24, LS20b, LWR20, LR23, MZL20, MM24, MF20a, MSP23b, NS23, NZL21, OPS25, OOCD^{+}21, OG24, PMB^{+}23, PDH21, PJKK24, PW20, Pra23c, Pup21, SH08, SdSCdS25, Sch25b, SMD22, SS20b, SK20, SCU21, Tut23, TD24b, VdBC20, WZL22c, WZNL23, WRT20, WWP20, WXZ24, XLL23, YLW20, YLX25, ZIS21, ZZ25, ZXL20, dCFdCA^{+}23, MSP23a, SH21]. reviewed [KHT22, MS22c]. Reviewer [LWZ22, GRSFV20a, GRSFV21b, XJZW24, See20]. reviewers [CLS^{+}20, JXZW23, Kal20, See20, Ste22, WXZ24, YTM21, ZSS22, ZTY^{+}24]. reviewing [LZ25b]. reviews [BVvEW23, CBA24, HFAR23, HP24, Jia21, Le 24, MCDFSP24, MLK22, OŠV^{+}24, Pol20, RAF25b, Sch21a, SSC21, Zho24]. revisit [HWL23, LZ24a]. revisited [JYT21, Kos21b]. Revisiting [DDH21, LLH22, MRBSFSJ24]. revivals [HMW24]. revolutionary [Wra20a]. reward [GBH21]. rewarding [Naz20]. rewards [VP21a]. rhythm [LZC21]. rich [NSM23]. rich-club [NSM23]. Richard [Ano23, Sug23]. right [Ans23b, GZOC24, QZF22, dSV21, dSV23]. rights [GXC25, SB25]. rigor [LGY23]. rigorous [BPH21]. RIRE [SYK22]. rise [MMGL23, ODF20, HP24]. rising [AS24, DSH^{+}20, LHCH20, ZZC22]. risk [JJY21, PL22a, YLW20, dS21b]. risk-adaptive [JJY21]. risky [LBS23]. Ritz [Lax23a]. rivalry [WG20a]. roadmap [JJY21]. Robust [PMHC21, BT20]. robustness [DZ23, RBBP24]. Role [BPMRSL21, AAD24b, ADG24, ARDA20, BY23, BKO25, CZX24, CXZM25, FYY21, GYWZ21, HX25, Kaj22, KN22, LLL22, LLLL24, LMG20, MPS22b, RYS24, SQC^{+}20, SZK20, TXL21, VC20, XCT20, YLH24, YCXY20, ZCS^{+}23]. roles [DLC21, MR20, PLW23, XDW22, YWL25]. Roma [IVCE20, SOdLGB21]. Romani [IVCE20]. Romania [Ben22, Cer24a, Cer24b]. Romanian [KBSS23]. RootCite [LYT^{+}20]. roots [BL20a, GTB23, LYT^{+}20, NS23]. route [XHA^{+}20]. routes [MH24a, ZM24]. rows [Fan20]. Royal [SLX21, CMT24, ZM24]. RPYS [WPBH24, HB22a]. RT [Sca25]. rules [AUA^{+}23, EO24b, Pra21a, RJ20a, Sch20b]. Russia [Che24, SA21, ZCSK24]. Russian [AL20, DAGG24, GKK23, KG22, LdS25, MML^{+}20, MF20b, PL22b]. Russo [Naz25].

S [BWH20a, BWH20b, SSBC22]. S&T [SSK21, YMW^{+}20]. SAARC [DLS23]. Safón [Pra23a]. Saharan [Ony20]. salient [BKS20, WCY22]. Salinas [Rus23]. Salton [DKWC20]. Same [LL20, LHW21, Che18, Che20, CG24, San25]. Sample [RSA20]. sampling [JMP20]. SAO [OJKY23, YKS22, ZC24]. SAP [VPR22]. SARS [HB20]. SARS-CoV-2 [HB21b, PHBG21]. satellite [SH20]. satisfaction [TP20]. Satoshi [Ant20]. Saudi [AJBBA23, AAAA23]. says [TTBA20]. SB [TBA23]. scale [ADD23, CXD21, CPN21, DvE20, DGSVDG20, FCW22, JHVC21, KBS21, KOS21a, KA20, KR22a, LTL24, LL25, LML21, MMFF25, NF21, Sha24a, WWS^{+}23b, YWN25, YWW23, YMD20, mYY22, ZWL25a, ZDLR21, dS21c]. scaling [GS22a, vRW25]. scams [BT20]. Scenario [DDVV20, CCH25]. Scenario-driven [DDVV20]. scenes [LHZ^{+}23]. scheme [ÁLGBdMA24]. schemes [RGVQTS^{+}24, ZS20b]. Scholar [MMTOMLC21a, dCPT22, CB20, LR25, LWL24, RA20, SZY24, YZZT24, YLX25, YTM21, ALR24, Fur20, GKM^{+}24, GTH24, MMTOMLC21b, MRM20, San20, SSL22, Tei24a, dS21c]. scholar-level [RA20]. scholar/journal [CB20]. Scholarly [Ano20c, CFM20, FVLA20, WZ21, ADG24, AFH21, ALM^{+}23, BGB25, Bou20, CB25, Che18, Che20, CFC22, Dag24, DABSC20, FCW22, FCY23, FC20b, GW23, GCHT20, HB20, HZC21, KL22, KTB22, LFV21, LMC22, LZ24a, LZL20b, NLS24, OPZP25, ODQ24, Par21, Par24, PA22, PVA24, SF20, SSdCRD22, Sca25, SGC21a, SGC21b, STM^{+}24, TA20, WAT23, WZG21, YNO22, YNL25, YYC22, YMLL23, ZXL^{+}24, ZYPC20, dDM25, VHM24]. scholars [BT25, BTD21, CW25, GDdZ22, GX22, KP23a, LWL24, LWX21, LRW25, Par24, PG24, Plo25b, San24, SBF23, YRP21, YWZ^{+}24, lYnLyO^{+}20]. Scholarship [WJZ24, XYY25, BSFS^{+}22, HBO^{+}20, KW24, RAG25, WCM20]. school [TdS20, Lam20]. schools [AU20, BT25, RA20, UNK23, ZPH21]. SCI [ZL20a, CLRMR22, MS23, OA24, RM24]. Sci-Hub [CLRMR22, MS23, OA24, RM24]. SCI-indexed [ZL20a]. SciBERT [TBA23, WLL25]. Science [Ano21c, Ano21d, BGP20, BM21, Bou21, CLD22, Che18, Che20, Dia21, FCC22, HWL23, LL20, LGE^{+}23, Liu21, LHW21, LNH24, MM25, MMTOMLC21a, MN21b, Mil25, Pra23a, RS22b, RAG25, SD23b, SSS^{+}24, SNK22, Tei24a, WLLW25a, WCL25, YSW^{+}20, ZW25, ADGS25, AAHY20, ACD^{+}24, AMTS24, BVvEW25, BPMRSL21, BMBA20, BS23, BMA21, BG21, Bor20a, BM24, BWE24, BVS25, BA22b, Cha21, CtRbQ22, CSC23, Cho21, CKMY20, DST23, Dej25, DFRS20, DFS22, DKWC20, DSB22, DMS20, DHSS20, DDF20, DAGG24, EO24b, EEÁ21, FVLA20, FESRMB21, FRS23, FNM^{+}22, Fuk22a, Fuk22b, GZOC24, GKK23, GEFJ24, Gon23a, GA21, GB21, GDG^{+}20, GDdZ22, GDdZ25, GLPN25, GMCRV21, HLS21, HJLZ23, HCS21, HS22a, HGB20, HAT23, HBW^{+}21, HBC25, HKV20, HhC20, Ibr21, ICB24, JS21, Jok20, JVY22]. science [KLLG20, KLNW20, KJY24, KKK^{+}25, KDZ21, KSS20b, KCS25, KS24, LFV21, LZH20, LMH22, LYZ^{+}20, LCC23, LLL25, LWR20, LYZ22, LYZ23, LXL25, LM20, LZL20b, MB22a, MSV23, MSJ23, MMG21, MNM^{+}20, MQS^{+}20, MS21f, MIY^{+}20, MBR^{+}25, MMGL23, MRC22, MKI24, NSM24, Nuz20, Nuz21, OdS20, ON20, Old22, OM21, Ove23, PRS^{+}21, PYS22, PT20, Pol20, Pra20c, Pra25b, RG20, Raf25a, Rod22, RPAVFV21, RP25, RAS21, SD23a, SMM^{+}20, Sch24a, SGS22, SMJ21, SAS23, SLL^{+}23a, SW24, SW20, SMQL20, SWL^{+}23, SM20, SK20, SC20b, TA22, Tas21, TTDK22, TNS20, TZ24, TSDAMRG24, VPSF23, Ven24, dFVFA^{+}24, VC20, Wag25, Wan24b, WWP20, WXD24, XW25, YGO22, YWW22, YP24, YSWC24, YGW^{+}24, YCL25, YJJ25, YA21, YKTM20, YY21, YY22b, Zbo25, ZT23, ZM23a, ZMP24, ZCSK24, ZS20b, ZPH21, ZP20, ZHH23b]. science [Tei21, dSV21, dMGDD24, vD21, AGB23, AOM24, ARR23, BdOdS21, CHJ^{+}25, EO24a, FCT^{+}20, GKM^{+}24, HB23, HWNL20, HCZ^{+}22, KMM22, KS22c, Kup22, LCC^{+}20, LGE^{+}23, Lie20, LTH20, MMTOMLC21b, MMB23, MN23, MAS^{+}21, OdSO20, ODQ24, PLH20, SSSd22, SQC^{+}20, SSM21, SM23, SS22, VV21, VEGSMMC22, WGC20, YWW23, ZL20b, Emm19, JDG21, Mil20, TTBA20, mYY22]. science-technology [TZ24, YY22b]. science-wide [KCS25]. Sciences [BOV21, RJ20a, VAMBTS25, AS23, AG23, AAB^{+}24, Ben20, Ben25, BSS24, Bha21, CH20, CM20, Coc20, DABSC20, DM25, EGV22, FJO20, GLK^{+}23, GC22, Gon23b, GdM20, GDG^{+}20, Gra20, HGDRM20, JS21, KDIR23, MN21b, MJ22, MJ23, MZE20, Nis23a, Nis23b, OTH22, Öz24, ROB21, RBB22, SRL20, SHL22c, SHL22b, Tay20, TMZ24, TMW23, WSY21, WL22, YK23, YS23, ZS24, ZGBH20, dMALIGBM20, CxGdH^{+}24, PMB^{+}23, RSV^{+}20]. Scientific [ARDA20, Ano24, BDPP23, CLZ21, CMZ20c, EGPGGA^{+}23, GMCRV21, HB23, JLL21, LS21c, LZY22, MF20b, MBF23, MNP21, MRC22, NC20, PRGTOP^{+}24, AD21, ADG24, ASSB20, ABAAN^{+}23, Ali22, AQK^{+}22, AUA^{+}23, AYS24, AA22, BZL21, BD22, BMBA20, BdPFdO25, BTD21, BGAMS23, BF25, BRJ21, BGSRF20, Cao20, CXD21, CGPR24, Cer24b, CSA20, CDE23, CMT24, Cha23a, CCZL23, CWLS23, CxGdH^{+}24, CDQ25, CdSS25, CRBRGS21, CCMM24, Coc20, CDG20, CMK24, DD21a, DSL21, DM25, DH23, DMA^{+}20, DSL23, DMGBG20, EBW23, EO24a, ELR24, FCW22, FAY25, FC20b, FGCGFC20, FRRBDD23, FFL21, FVGBPA23, FLDD20, FC20c, FJO20, GRTPAJ21, GWLY24, GCH^{+}22, GKDB25, GCS25, DR20, GC20, GAACBA24, GdOCRM20, GMKG20, GCHT20, HSE23, Hal20a, HIA^{+}20, HAS^{+}20, HB21b, HMW24, HGDRM20, HBC25, HWX22, HYZ23, HMZ21]. scientific [HCN21, HZJ22, HZW22, HDY^{+}23, HTZR24, HX25, Ioa23, IVCE20, JME20, JHVC23, JKY22, JADG20, JDSL20, KSZ22, KLM24, KN22, LCB20, LLC22, LMH22, LC22, LCL23, LHX^{+}24, LW24, LL25, LdGRV^{+}22, LYS22, LLL25, LWZ22, LL24a, LZ24b, LZF^{+}24, LLZ21, LYLC23, LDA22, LMLMPA22, LMLR24, LXS22, LC21, LUH21a, LUH21b, MBL21, MH24b, MPC22, MCSOAA21, dMMdCM22, MPS22b, MSI21, MSL22, MR20, MdMAGB^{+}21, MDCBdGVMD24, MBH25, NRGvSTS23, Naz25, NA24, OMC21, PLYZ24, PRS^{+}21, PROG21, PROGDG24, Pol21, Pra25a, Pra20d, Pup21, QWS^{+}23, QYL21, Reh20, RWL23, RdNB21, RNB22, RBL20, RP23, RP24b, RP24a, SOSC24, SHF24, SMG^{+}24, Sha24b, SXLC21, SDC22, Sma25, SM23, SM24a, Sta21, SLX21, SM24b, TZZZ23, TM20, TNS21, TZ24, TSDAMRG24, THG23, TMNH^{+}24, TG24, TTB20, VFL^{+}21, VPDAG22, Ven24, VPC^{+}23]. scientific [WZL22b, WWC22, WWS^{+}23b, WWL24a, WZNL23, WY23, Wra20a, XLL23, XlYqZ22, XLY21, XAA21, XXL25, XCML25, YG20, YF24, YWX23, YZH23, YLH24, YXDL24, YCYZ24, YT25, YP20, ZW25, ZC21, Zbo25, ZLL21, ZNNA20, ZZS^{+}20, ZF22, ZZ23, ZMLZ24, ZZ24, ZSL24, ZBKW24, ZCL20, ZCLW20, ZXL22, ZJ25, ZLL22, dSDE21, dOCP21, dW24, dBGP21, dCFdCA^{+}23, dSSA22]. scientifically [KR22a]. scientist [ÁRMS21, BTP24, Bor24, GTB23, HSE23, YWN25, ZWZ22]. Scientists [GCH^{+}22, ADD24, AHH23, CT20, CCH20, DLL21, FFJ21, JVY22, KR22b, LZY22, RS22a, RB25, SMD22, WZZ23, WWL24a, WZZ25, ZL23a, ZL23b, ZY20c, ZQML23]. Scientometric [EO24b, HB20, KH21, LdS25, MQS^{+}20, SMD22, She20, VKY23, APR25, Ali24, Ali25, AMZ24, ARR21, dLBdAPUdLB^{+}25, Bor21, DRF21, Kaj22, KW24, Le 24, RSV^{+}20, RP24b, RP25, SOdLGB21, SSdCRD22, SNK22, SSBC22, THvB21, WG20a, YLW20, dCFdCA^{+}23]. Scientometrics [BMW24, DMG20, GDR22, MS21f, ALR24, Ben22, BS20, CGPR24, CZZN25, Fur21, Fur23, LFV21, LB24b, dOPGdAPU22, RFG^{+}24, RGVQTS^{+}24, VKY23, WGW21, XLL20, Dia21, GC20, TSAMT21, Zha23a, BDTC20, CN24, YLJ21]. SciFinder [Tom24]. Scilit [ODQ24]. SCImago [GMRVU24]. Scite [LS23]. scite.ai [LS23]. SciVal [DJM22]. scope [MGCR24, SM22a, Sha24a]. Scopes [BdCM21]. scoping [MSP23a, MSP23b, OPS25, SK20]. Scopus [AAA^{+}23, MS22a, MS21b, MMTOMLC21a, AMMN21, APR25, ÁLGBdMA24, AOM24, ARR23, BdOdS21, BRJ21, CHJ^{+}25, Don21a, Ela25, Fan21, GKM^{+}24, HWL23, KG24, KS22c, LL20, Liu20, LHW21, MS21a, MS21c, MMTOMLC21b, MFM20, MN23, NT24, OdSO20, ODQ24, PD20a, PLYZ24, PDH21, RB25, SMJ21, SSM21, SS22, SA21, TKS^{+}23, Tei24a, TP24, ZAZS22, ZL20b]. Scopus-based [PLYZ24]. Scopus-indexed [NT24]. Scopus/Web [HWL23]. score [FZL^{+}20a, HB22b, Sca25, YY22a, ZXL20, AMTS24, FZL^{+}20b]. Scoreboard [Vie23]. scores [BS23, BMS21b, Cop20c, HB22b, SHB25, WXZ24, XJZW24]. Scott [SSBC22]. scraping [BOPGOP25]. screening [CPLS20, Jia25]. scrutinized [MMGC^{+}24]. Scrutinizing [JI22]. SDCF [BT22]. SDG [HS22a, Pur24]. SDG-oriented [HS22a]. SDGs [HS22a, MBH25]. Search [Gus22, AFH21, CDE23, FNK23, GPYR^{+}20, KL24, LS21a, Lin21a, LSY21, Liu24, OM20, PR22, Pol20, OMFJSC24]. second [Pea20]. secret [AHRH24, Par24]. section [ASSB20, LLC22, ZYZZ25]. sectional [Ali21a, Ali21b, BRG^{+}23, OMAAORCL23, SdLV21a]. sectioned [Bös21]. sector [AAD22, Kim23, OVS25, QZ20, SZZ23, VPC^{+}23]. sector-specific [OVS25]. sectoral [PSW24]. sectors [LHC22, MNP21, SYL^{+}24]. security [RdNB21, WLBC20]. seed [WSY21]. seeking [SB21]. segment [JC23]. segment-wise [JC23]. segments [WCY22]. select [Fas21a]. selected [CKT21, Mou21b, Mou21c, Ony20, Wan24b]. Selecting [BB22, SSK21]. selection [Coo25, Cop20a, DD21b, GMPRSFV25]. Self [Bor20a, FSEJF^{+}20, Lin20, TZZZ23, VPC^{+}23, Zho21, ARDA20, BSS24, BRJ21, Cop20a, DMS20, HH21, KFM20, KLLG20, LS21b, LL20, OG24, PCG^{+}20, San20, SAR19, SPA20, WLL25]. self- [OG24]. self-adaptive [WLL25]. Self-citation [VPC^{+}23, Zho21, BRJ21, DMS20, HH21, SPA20]. self-citations [ARDA20, BSS24, Cop20a, KFM20, LL20, PCG^{+}20, San20, SAR19]. Self-correction [Bor20a]. Self-defined [FSEJF^{+}20]. Self-plagiarism [Lin20]. self-referencing [KLLG20]. self-selection [Cop20a]. Self-supervised [TZZZ23, LS21b]. Semantic [AA22, GHW22, KDZ21, AMTSRG21, CCH^{+}24, CXZ^{+}20, CDG20, DCL23, FZC20, GPYR^{+}20, HCZ^{+}22, LPD21, LYLC23, MLD21, PMB^{+}23, QQL21, SBP20, SLL23b, VPR22, XZZ22a, XZZ22b, YP20, ZY20b, ZY22, ZLZ^{+}23, ZYZW23, VPR22]. Semantic-enhanced [GHW22]. semantics [KZH25, LYLC23]. semantometrics [KSS20b]. semi [AS25, ASX22, BT22, CNT25, HM21, KP23b, MMFF25, ZC24]. semi-automated [MMFF25]. semi-automatic [ZC24]. semi-automatically [BT22]. semi-experimental [AS25]. semi-peripheral [HM21]. semi-supervised [ASX22, CNT25, KP23b]. semiconductor [Fil21]. seminal [HM20a]. senior [LRW25, SO21]. seniority [DMS20]. sensing [VCD^{+}23, dSMNdV^{+}21]. sensors [PHBG21]. SentCite [WCY22]. sentence [WCY22, ZZ24]. sentence-level [WCY22]. sentences [NKY^{+}24, ZC24]. sentiment [AKD23, AYS24, BY23, MTD^{+}23, SSK22]. separate [MKG23]. sequence [BPW^{+}21, LCW^{+}20]. Sequencing [ZM23b]. Serbian [DDF20]. serials [UA20]. series [BW20a, BW22, LHC25, LGE^{+}23, SCU21, TWH^{+}24]. series-based [SCU21]. serious [SMT^{+}23]. servers [SCP22]. service [SCW^{+}23]. services [Ans23a, MS22c]. Sesame [BL20b]. set [ARR23, CW25, NLS24, Pra21a, SF20, Vin21]. setting [Asa20b, CPLS20]. settings [Zie22]. seven [Hüc22, ODQ24, WSL23]. Seventh [CFI22, PSW24]. sex [Lax23b, Nuz20, WSRM23]. Shallow [DXL21]. Shalosh [HCZEP20]. Shanghai [SD20]. ShanghaiRanking [Li25]. Shannon [Sma25]. shape [BB23, DSB22, DH23]. shaped [ADD22]. shapes [PS22]. shaping [HAR24, Jam24]. share [CF23]. shared [DS25, SCFSAB^{+}25, vSDC20]. Sharing [SMS23, ANR21, Ben20, CP21, CBP^{+}24, DTHM21, DSB22, DGGBDG21, LDAAAB21, RL20, SCP22, ZM23a]. shell [WH25]. shift [LC20]. shifts [Ali25, BZZ25, KY20, KG24]. Shiraz [RSV^{+}20]. shock [JC22]. Shodhganga [KBDS23]. Short [MMF20, ARR21, JT21, LWR20, ZA20]. Short-term [MMF20, ARR21, ZA20]. shortens [ZQML23]. shorter [XAA21]. Should [LS20b, MRM20, Mut22, SKBSC20, ZWHS22, Asa23b, DV22, HBP20, KKPJ23, MHK22]. show [BPLIdMA^{+}20, Tay23]. shows [WCM20]. shrimp [TdSFB20]. Sic [FG20]. side [CtRbQ22, SP20]. side-effect [SP20]. sides [CG24, WPSK25]. sight [GSOCS21]. signalize [BdCM21]. signals [MKG23, SYK22, VY23]. significance [ZXL22]. significant [CCH20, SPS20]. significantly [TZY^{+}23]. signs [Cop20d]. Silver [LSFLZLLV20, GCGB^{+}24]. similar [BL24, Ela25, SHB25]. similarities [CLC23, GSG22a, GDRPR23, PMC21, Thi20]. Similarity [ZC23, AA22, BBB^{+}24, BVvEW25, BZX^{+}23, CZFC24, DCL23, HGMÁW23, cHhC24, LHX^{+}24, LXZ^{+}20, Par25, SBP20, TTDH25, VA22, WCY22]. similarity-based [AA22, BZX^{+}23]. simple [San20]. simplex [HS22b]. Simpson [WR21]. simulating [Xie21]. simulation [CLL22, HS22b, SS20b]. simultaneous [CL21b, KVZ20, MBL21]. simultaneously [PJ24]. since [KYKC21, RdNB21, ZL20a]. Singh [JTB23]. single [GKDB25, MMF20, MJYS23, Pra25b]. single- [GKDB25]. single-author [Pra25b]. sites [HBW^{+}21, YLCY20, YWZ^{+}23]. SIUR [JVY22]. six [MS22c, OdSO20, SA24]. Size [BA22b, dOSAL21, CTH22, CXZM25, Cot22, DRS20, Hir22, LKPS25, MM24, RSA20, Sha21, Wal22]. size-biased [DRS20]. skepticism [JDG^{+}20]. Skewed [Bor24, Sch20a, Sch22]. sleep [KVZ20]. Sleeping [KVZ22, vR21b, HV21, HLZ23, KVZ20, TTB22, ZY20a, HYSY23, KZV20]. Slovak [TMNH^{+}24]. Slovenia [BRJ21, RBL20]. Slovenian [TMZ24]. Slow [Tay23]. small [ABK20, Asa21, ER24, MHK22, OB21, RJ20a, SM22b, TdS20, ZY20a]. small-teams [OB21]. small-world [ER24]. Smart [PYS22, Fog23, WH23]. SME [CGA21, VCV22]. SMEs [VCV22, WSL23]. SMN [AQK^{+}22]. SNIP [DMA^{+}20]. soccer [VBF24]. Social [BKS20, BOV21, CxGdH^{+}24, Mas20, RJ20a, YLCY20, ABK20, AG23, AMTSRG21, AAB^{+}24, BRG^{+}23, CWZ21, CZX24, CB25, CtRbQ22, CWLS23, CM20, CDG20, CBA24, DABSC20, DD23, DM25, DDF20, DMA^{+}20, EGV22, GC22, Gon23a, Gon23b, GdM20, GLPN25, GDRPR23, HIA^{+}20, HAS^{+}20, HM20b, HYZ23, Iva25, JVY22, KBS20, KDIR23, LCC23, LYZ22, LYZ23, LOGS21, MBL21, MS22c, ML24, MB20b, MB21, MB22b, MMG23, MN21b, MJ22, MJ23, MQS^{+}20, Nis23a, Nis23b, Öz24, PHBG21, RTT23, RBB22, SRL20, SHL22c, SHL22b, SZN^{+}21, Sch20b, SLL^{+}23a, Tay20, TMZ24, TMW23, WGC20, WKC22, WC20, WP22, XlYqZ22, YWZ^{+}23, YWW22, YLf25, YZZT24, YS23, ZIS21, ZGBH20, ZJ25, ZHH23b, dMALIGBM20, dW24, Lie20, PMB^{+}23]. social-spatial [PHBG21]. Societal [JVY22, ZPZ21, DRF21, KSAT24, Mey25a, Pur25, ZM24]. society [BMBA20, BM24, ON20, SK20, Yur22, CMT24, SSC21, ZM24, SLX21]. socio [AMTSRG21]. socio-semantic [AMTSRG21]. socioeconomic [ELR24, vRW25]. sociologists [ATCS20, Gub24]. sociology [DT20, GDdZ25]. soft [LZ23]. Software [Bös21, ADU22, HB22a, HCS25, JKY22, OMC21, PA21, Pup21, SS20b, Tom23, WL24, WRT20]. Sokal [Kos22]. solar [DG21, LGMK21, SH20]. solid [DBR21]. Solla [Ano20a, Ano21a, Ano23, GDRPR23, Sug23, vR21a]. solo [Emm19, KR22b, LR25]. solo-authorship [LR25]. solution [PSH22, THvB21, WZM^{+}24]. solutions [Don21b, LHW21, ZCS^{+}24]. solving [GRBAM22]. Some [KK23, Sch24c, TM20]. Something [HP24]. sonification [BH25]. Sorry [Gra20]. sosia [RB25]. soulmate [Etz25]. source [ANR21, DGSVDG20, LWW22, NII24, Par21, Tay23, TSAMT21, WLHY24, WLS22, Zho24]. Sources [BCG22, JDG^{+}20, Kra20, Gon23b, LML21, LC20, Mor19, Ort20, OVS25, Tay23, YK23]. South [Kim23, LKPS25, AHH21, BP20, CFA^{+}23, Cha23b, CLZ21, Dha22, LdGRV^{+}22, VCD^{+}23]. Southeast [Pur21]. Southern [NB22]. Soviet [ADGS25, CLP21, GKK23, KDIR23, LPY22, LCP22, MSL22, MBF23]. space [SSH21, Yen25]. spaces [BPL25, KDZ21]. Spain [RFMMPJ22]. Spam [dSAKT20]. span [WKC24, ZNNA20]. Spanish [DABSC20, FGCGFC20, FRRBDD23, FFL21, DMR22, HGDRM20, LMLR24, PMGÁC23, SV20, SJBBR^{+}23, SCDP21, STAÁMR24]. spanning [XWGD22]. spatial [FMB22, HZ21, PHBG21, RKCA23, Ste21, TdSFB20, WCH^{+}20, ZWW21]. spatial-temporal [HZ21]. spatiotemporal [Gu21, WWS23a]. special [BG21, FWH^{+}23, HHY22, MN21b]. specialisation [Fog23]. specialist [GSH24]. specialization [BPMRSL21, Fer21, Sam23]. specialize [AAD24a]. specialties [EGV22]. specialty [XLY21]. specific [CCSX23, Che23, CLS^{+}20, GRBAM22, HB22b, MNK21, OVS25, WKC22]. Specification [LB24a, HLSX20]. specificity [BI24]. specified [CLC23]. SPECTER [AQK^{+}22]. spectrometry [CdAdFC21]. Spectroscopy [WPBH24, BC21, CZZN25, GTB23, HB22a, SH20]. speech [Pea21, TNSF21]. speeches [ACT23]. speed [Lin21b, MH24a]. spending [ZM23b]. sphere [DST23]. spillover [SAK22]. spillovers [ADD20, Wad20]. spin [LRSB21]. spin-off [LRSB21]. spine [CLC23]. spine-related [CLC23]. spinoffs [Fuk22a, Fuk22b]. SPIRIT [OMAAORCL23]. sponsor [DTHM21]. sponsors [IHLP21, MK23]. sponsorship [Hua24]. sport [BS23, BSS24, SdS25]. sports [HGDRM20, Lax23b, WSRM23]. Spotting [ZTY^{+}24]. SPR [AQK^{+}22]. SPR-SMN [AQK^{+}22]. Springer [Che18, Che20]. Sri [RBR22]. SRRS [VHM24]. SSCI [GC22, YS23]. SsciBERT [SLL^{+}23a]. SSH [WZ21]. ST&I [Ano22b, HWP22]. stability [GYWZ21]. stacking [HH21]. stage [FZhCC21, GCMJ25, CLY^{+}22, ZSL24, ZBKW24]. stages [AJVACC22]. stakeholders [Pur25]. Stalin [KKS21]. stamp [KS23]. stance [DWG23, Pol20]. stance-taking [DWG23]. Standard [BF22, GRSFV21b, MBH25, PSA22, PKA24, Yur23]. Standard-relevant [BF22]. standardised [KCS25]. standardization [BVS25]. standardized [Sca25, SVY22]. standards [BPH21]. star [LHCH20, SZZ23]. Stars [ABK20, CF22, DSH^{+}20, ZZC22]. Starstruck [ES22]. start [MWH20]. state [DBR21, HM21, LCC23, LYZ22, LYZ23, YT25]. stated [Cop20b]. States [MKWB^{+}23, GEFJ24, GDM22, PL23, SO21, ZPH21]. Statistical [Ibr21, ELR24, GG25, LGY23, MMB23, Mut22]. statisticians [BBP23, MN24]. statistics [APR25, Bre21]. status [BMS21b, Tay23, WKC24, Tei21]. stay [SKBSC20]. steady [GAACBA24]. STEM [Tom21, UIM21]. step [MBL21, OG24, PLH20, WPSK25]. STI [SC20b]. still [LZ24a, LZYR24, MH24a, MSJ23, Ste23]. stochastic [PD20b]. stochastic-based [PD20b]. stochastically [BT20]. stock [VCV22]. Stone [Lax23a]. storms [RBBP24]. story [Cer24a, SH20]. strabismus [SMBD^{+}24]. Stranger [AMTS24]. strategic [BP22, KPY21, SHL22c, SHL22a, SHL22b, SBF23, WJZ24, LMLR24]. strategies [AKW23, HS22b, LLL25, SSS^{+}24, dGdO^{+}23]. strategy [FNK23, JME20, JKY22, Lin21a, LSY21, Liu23, ŚT21, VWR25, WAT23, ZY20c]. stratification [ATC20]. streaks [OSS^{+}23]. stream [LBW^{+}20]. streams [JMP20]. Streetlight [MMGC^{+}24]. strength [SW24, vRW25]. strengths [LNS21, SGG20]. stress [GB24]. stretching [Liu24]. STROBE [RSV^{+}20]. strong [XZD19]. strongly [HB22b]. structural [BERD21, CZX24, KZH25, MGA22, Reh20, WLZ^{+}23]. Structure [MLYK23, WKC24, Wra20b, ACD^{+}24, CWL^{+}20, Cho20, Coc21, HBN^{+}21, KZH25, Kös20a, KPY21, KZZ^{+}20, LLL22, LWL24, LZL20a, MZD22, MIY^{+}20, OJKY23, OPZP25, PKCOG21, SVC20, Söd23, TMZ24, VBF24, WZJW24, YXDL24, YLCY24, ZYZZ25, dDM25]. structure-based [LWL24]. structured [BT22, SMD22]. structures [MPR21, MPMR21, MS21f, YKS22]. structuring [OOCD^{+}21]. student [AMIK23, ES22, LDMVS21, PWZ25, PDH21, WMK21]. students [GB24, QSYL23, WGW21, XS25]. Studies [Ano21c, Ano21d, BG21, BWH20a, BWH20b, CN24, CMK24, DMGBG20, EEÁ21, FNK23, FESRMB21, GKDB25, GA21, GB21, GBP21, HV21, HdSV23, IVCE20, IHB21, JMP20, KPY21, KSAT24, KSZ22, LZ25b, OM20, PVA24, RSV^{+}20, RWL23, SM22a, SM22b, Sch25a, Sch21a, SC20b, WHÅ24]. study [ATW20, Ali21a, AHH23, AM21b, AHRH24, ANR21, AMZ24, BZZ25, dLBdAPUdLB^{+}25, BMA21, BRG^{+}23, Bor20a, BST22, CLL22, CGPR24, CLP21, CDZ^{+}20, CC21, CG22, CKO21, Cop20d, CF23, CBP^{+}24, DDU24, DFRS20, DX20, DFG21, DDD^{+}21, Dia21, DD23, DL23, Don22, DGSVDG20, DMA^{+}20, ES23, EE24, EGPGGA^{+}23, FZL^{+}20a, FZL^{+}20b, FC20b, FCC22, FGCGFC20, FJO20, FN23, GLK^{+}23, GCMJ25, GEFJ24, GAPR20, GHGMSM21, Gor21, GdOCRM20, HNO^{+}21, HAT23, HB21b, HMW24, HTX23, HGM21, HM20b, HX25, Ioa23, JHVC23, JCX22, KA20, KKK^{+}25, KVZ20, KSS21, KUS22, KKS21, KS22c, KR22a, LFV21, Lam20, LZH20, LYT^{+}20, LL25, LLSH23, LYZ^{+}20, LHC21, Liu20, LGMK21, LS23, LC20, MTD^{+}23, MB22a, MSV23, ML24, Mas20, MMG21, MQS^{+}20, MFF^{+}20, NHAB21, NB20, NB22, NF21, NF22, Ony20]. study [OMC21, OMFJ22, PS20, PSKS22, PROG21, PCG^{+}20, PHP22, Pur21, Pur24, QQL21, QZ20, RSV^{+}20, RBR22, RJ20a, RL20, SOdLGB21, San21a, SHF24, SKS^{+}23, SGM22, SdLV21a, SdL21, SdLV21b, SS23, Seb23a, SdSFB22, SG21, SR23, SM23, SK20, SGG20, SSK22, SMA21, SCW^{+}23, SLX21, TG20, TD24a, Tom21, UA21, VAMBTS25, dFVFA^{+}24, Wad24, WWZW20, WWS^{+}23b, WYL21, XL20, YZH23, YP24, YLM24a, YH24, YLJ21, YWW23, YCL25, YNO22, YWL25, mYY22, ZMA23, ZLL21, ZGM^{+}23, ZXL^{+}24, ZWL^{+}25b, ZPZ21, ZDLR21, ZYPC20, ZM23b, Zhu25, Zbo25, Ali21b]. Studying [BS20, BCM20, FC20a, LC21, WZJW24]. style [MTD^{+}23]. stylistic [WVH21]. stylometry [JADG20]. sub [Ony20, Tas21]. sub-fields [Tas21]. sub-Saharan [Ony20]. subdomains [HAT23, RSM24]. subfields [CSA20, HhC20]. subgroups [SMA21]. Subject [EGV22, SDC22, AA25, ARR23, BBP21, CPN21, CG24, CLC23, FCT^{+}20, JS21, Li25, LMLR24, Men25, MFM20, OdSO20, Ort20, PLH20, WKC24, Wij21, ZZ20b]. subject-normalized [Men25]. subjects [Ben20, Ben25]. submission [MWH20, MLA^{+}23, OFA23, RJ20a, ZJY22]. submissions [AHRH24, Asa24, TLL24]. submitted [Ali25]. subscription [HHB^{+}20, KL22]. subscription-based [HHB^{+}20]. subset [Cop20c, GSG22a]. substances [Tom24]. substantially [Bor20b]. substitution [Wal22]. succeed [Hüc23]. success [Cer24a, CDG20, Fur21, HWX22, OSS^{+}23, WZG21]. Successful [KS22b, The23, VWR25]. such [GG22, Pur25]. sufficient [CCH25]. Suggestions [CCC22, KCC22, SJW22]. suggestive [WG22]. sum [Mut22, XlYqZ22]. summaries [CNA22, WY23]. summarisation [ZNNA20]. summarization [FAY25, LZL22, SMA21]. summarize [LCB20]. summing [FDM24]. super [DGRÁTS23]. super-cited [DGRÁTS23]. Superior [HY22]. supervised [ASX22, CNT25, KP23b, LS21b, MS21e, SM24a, TZZZ23]. supplement [HLZ23]. supply [SBP20, YY21, YLX25]. Support [LHZ^{+}23, CBA24, FFN25, GRBAM22, GS22a, Gub24, MRC22, SMM^{+}20, TA20, WAT23, XCT20, ZLM^{+}23]. supported [LHF23, YK23]. supporting [Poh20]. supportive [LS23]. surface [VCD^{+}23]. surge [Cer24a]. surgery [CMZ^{+}20a, CMZ^{+}20b]. surname [Öz24]. survey [AGL22, KSS20a, KKK^{+}25, NBBL24, Plo25b, RS22a, SR24, ZWSC20]. surveys [FFJ21]. survival [SLTS25]. sustainability [GSG22b, RBB^{+}24, Sch24a, YLX25]. sustainability-oriented [RBB^{+}24]. sustainable [BPMRSL21, CPTÁ22, CdSK21, GRSFV22, HS22a, MBH25, NS23, Bar25, SSZ22]. sustained [ATC20]. swan [ZY20a]. swan-group [ZY20a]. swift [HL21b]. Swiss [CL21a, SdLV21b]. switch [Ger24]. switches [The20b]. sword [ZM23a]. synchronism [ZT23]. synergetic [WGZ22]. Synergy [LB20]. syntactic [Dia21, WZJ^{+}20, YP24, ZY22]. synthesis [HJLZ23, LRXC21]. Synthetic [TKZ^{+}23, dLBdAPUdLB^{+}25, BdPFdO25]. syphilis [SLMCSV21]. System [WGZ22, AMSS^{+}23, ADGS25, AdAFR24, AFH21, ARR23, BCG22, DWM21, FFN25, GBH21, GK22, GKD23, HIQ^{+}20, Hir22, CLY^{+}22, KA24, KCC23, Kwi20, LS21b, LMH22, LZL20b, NRGvSTS23, ROB21, RPAVFV21, SJBBR^{+}23, SRL20, SQC^{+}20, VMF22, Wag25, YLX25, ZS24, dSDE21, dSN23, VHM24]. Systematic [HKW24, AJB^{+}21, BVvEW23, BKT22, CGA21, CBA24, GKDB25, HNO^{+}21, HP24, MB20b, MCDFSP24, MF20a, MRF21, NHAB21, OOCD^{+}21, OŠV^{+}24, Pol20, Pup21, RAF25b, Sch25b, Sch21a, SCU21, VdBC20, WRT20, ZIS21, ZNW22]. systemic [Can24, DT20]. Systems [LHC21, ADD24, BSPR20, Coo25, DDS22, EO24b, FCY23, HHKZ20, JSLJ22, KR22a, MLK22, MJYS23, ÖT24, Pol21, PW20, ŞKKT25, ZSSH22, vdPF25].

T [Ben22, HBN^{+}21]. tackle [QJI^{+}22]. Taiwan [HL21a, HLS23, KP23a]. take [CTH22]. taking [DWG23, LZY22]. tale [ZL20b]. talent [HLT25, Rod22]. talents [lYnLyO^{+}20]. talks [LRW25]. Tamquam [GSG22a]. TAPRec [JMZ^{+}23]. Tardigrada [McC24]. Target [YH24, KZV20, Sch25a]. Target-oriented [YH24]. targeted [NCD22]. targeting [BPW^{+}21]. task [BAE24, QWS^{+}23]. Tasking [GVK24]. tasks [XLL23]. taxonomy [BPMRSL21, CXW^{+}24, DLC21]. TBC [KY20]. Teaching [KW24, MMF23, Sca25]. Team [LGE^{+}23, Sha21, SMS23, Ano23, BA22b, LWR20, LGXH24, YXDL24, ZW21c, ZWL25a]. teams [AL21, CdSS25, LGXH24, OB21]. Tech [Ano22b, GRBAM22, HWP22, OVS25]. technical [DCdA25, GRBAM22, ZMK22]. techniques [ASIN24, IHB21, Öz24, SMA21, ZWSC20]. techno [HC22, LZL20b]. techno-science [LZL20b]. Technological [Kim23, KJL23, OVS25, BORROPGV20, CXD21, CCH^{+}24, CLS21, DG24, Fil21, Ke23, KY20, KYL20, LLH25, LLL22, LCL23, LGMK21, LZD^{+}24, LLZ25, MAN22, OJKY23, PT24, PMGÁC23, RSPC20, RNB22, SG22, SHZ22, SM24b, TZ24, Wan25, WSL23, WZ25, XCML25, YP20, YJS21, ZZ23]. technologically [Yam20]. technologies [AED20, CPL21, JPS21, LWW22, NHAB21, QZ20, RSPC20, RBB^{+}24, Wan25, ZDL^{+}20, ZDLR21]. Technology [Ano21c, CdSK21, LLL22, ZMK22, BdOdS21, BPMRSL21, Bre21, CB25, CYK22, DG24, ES23, EEÁ21, FH23, FESRMB21, FBM21, GJ24, GB21, HJLZ23, HS22a, HAL25, HTZ24, IOWN20, JCK22, JJY21, JCZ22, Kim23, KJL23, KMGS21, LN24, LHK21, LS21a, LS21b, LHC22, LHC25, LZH20, LMH22, LP20, LCT^{+}20, Liu24, LZF^{+}24, MPS22a, MYK22, QPH^{+}25, SAK22, SLL23b, Ste21, SC20b, SZK20, SM24b, TZ24, VCV22, WH23, WFSZ24, WYL21, WXD24, YD21, YH24, YY22b, ZY20b, ZBG^{+}24, ZWL^{+}25b, ZM23b, dMALIGBM20, dBGP21]. technology-level [SM24b]. Technostress [SMS23]. TeknoAssistant [GRBAM22]. Telescopic [ZS20b]. tell [Dip21]. Telling [SH20]. tells [LB24a]. Temporal [HV21, HLS23, DFS22, EXT^{+}21, HZ21, LZ23, RKCA23, SSdCRD22, WWL24b, ZWW21, ZSL24, ZYWA25, ZCLW20]. temporary [Cha23a]. Ten [Hüc23, MCDFSP24, SSBC22]. ten-year [SSBC22]. tend [GMPRSFV25]. tensions [LW24]. tentative [Cop20d]. tenure [JC22, SQC^{+}20]. tenured [WZG21]. term [ARR21, Cha23a, HWX22, MMF20, YWN25, ZA20]. terms [BD22, CLC23, FZC20, IPP22, PB24, TNS20, WB21]. territorial [AD21, AAD24b]. territory [LBW^{+}20]. tertiary [FGCGFC20]. Teseo [MRBSFSJ24]. TESOL [JJ23]. test [CR24, San20, VWR25, WFW24, ZWC22]. testing [GDG^{+}23, Sun23]. Text [LLZ21, OOCD^{+}21, ZC23, AFH21, BVvEW25, BCF21, BM21, Bös21, BGSRF20, CPLS20, CLC23, Day22, IHB21, Jia21, LCB21, LZL20a, LZ25b, MKI24, PJ24, Pup21, QYL21, RSPC20, SF20, TQA21, WPSK25, ZNNA20, ZPZ21, Zie22]. text-based [Zie22]. text-matching [Pup21]. text-mining [LZ25b]. textbook [Yur23]. textbooks [DL23]. texts [AKD23, SLL^{+}23a, SdVvdR24, SM23, TSA24, WL22]. textual [Bor21, MZL20, MTD^{+}23, RTT23]. their [ADD24, ADG24, AS25, Bor20b, CTH22, Cha21, Che24, DMR20, DABSC20, DJM22, DGGD22, FLC22, FG20, HZC21, Hug24, JXZW23, JADG20, KKPJ23, KJL23, KZV20, KDN22, KL24, LGE^{+}23, LTL24, MCSOAA21, MNP21, MFM20, Öz24, RS22a, SHL22a, SO21, SdSFB22, SL22, SMQL20, STM^{+}24, TZY^{+}23, TNSF21, TG24, TTB22, VP21b, WWZW20, WZZ23, WCCC24, WH25, XZ23, Yam20, ZZZ22, ZWL23, Zho21, vRW25]. them [BW20b, Pra23b]. thematic [FMB22, LNS21, TdSFB20, Vie22]. theme [BLL21]. themes [CCMM24, GGS22, SVC20]. theorem [BM20a]. theoretical [CdSK21, Mut22]. theories [SdSFB22]. theory [AdAFR24, ARR23, dLBdAPUdLB^{+}25, Iva25, MRBSFSJ24, MS21e, dOPGdAPU22, Ste22, DBR21]. therapy [Nuz20, Nuz21]. there [Ali24, DGRÁTS23, DAGG24, GCHT20, Kup22, LZ24a, MS22b, San25, SZ20, vdBM24]. thermal [KY20]. thermodynamic [PR23]. theses [AL20, AAB^{+}24, BM20b, Don21a, FGCGFC20, HGDRM20, MRBSFSJ24, SOdLGB21, XLL20]. thesis [Don22, RJ20a, Xie20, vdBM24]. thesis-related [Don22]. thin [BPW^{+}21]. things [AMTS24]. third [HLS21]. Thirty [TNSF21]. Thomas [BWH20b, BWH20a, BM20a]. those [Cha21, GD22, Ste23]. thought [FAY25]. thoughts [KK23]. thousands [ZGBH20]. threats [Coc21]. Three [DK21b, Bir21, GZSC22a, GGS22, Gon23a, HB21b, Li22, LM20, MY25, MFM20, ODF20, RSPC20, SHB25, YGO22, dSSA22]. three-dimensional [GZSC22a, MY25]. throughout [GAACBA24]. Tibor [GS22b]. tier [dSN23]. ties [RP25]. tilted [FLDD20]. Time [KS23, SCU21, Tom24, ACT23, BFG^{+}21, CSL22, CZFC24, GFVK21, GDDG24, HTZR24, JMZ^{+}23, JC22, LHC25, Liu21, MZZL21, Pra20c, RS22b, SLTS25, SDRS23, TTDK22, TLL24, TWH^{+}24, Tut23, TD24b, WL24, YWZ^{+}23, Zhu25]. time-aware [JMZ^{+}23]. time-normalized [GDDG24]. time-series [TWH^{+}24]. Time-stamp [KS23]. Time-tracking [Tom24]. time-windows [BFG^{+}21]. Times [Fas21a, BPH21, HKV20, Seb23a, SMBD^{+}24, TM20, WWY^{+}20, TMNH^{+}24]. timing [CJT22]. tip [PJKK24]. title [TDM24]. titled [MHD21b, YP20]. Titles [Hao24, BI24, Dia21, Jam24, JJ23, JADG20, NT24, Pea20, Pea21, QLW25, WZ21, Wan24a, XL20]. today [Zan23]. together [LCS25]. toll [SGC21a, SGC21b]. toll-based [SGC21a, SGC21b]. tone [BAE24]. too [JKK24, San25, SPA20]. tool [AJLA21, BMS21b, GBP21, HS22b, HLSX20, KYL20, MBD25a, MBD25b, NII24, ZDLR21]. toolkit [RTT23]. tools [Cop20b, Tom23, YKTM20, ZC21]. Top [RS22a, BP22, CT20, CLC23, DT20, DLL21, HWX22, JP22, JTB23, LFV21, LM20, LR21, MS22d, OdS20, ODF20, OM21, OŠV^{+}24, Par24, PA21, QD21, WZL22a, WWL24a, WLS22, WXZ24, YWX23, Yur20, ZL23a, ZL23b, ZPH21, Zho21, ZYL20, AHH23]. top-1 [WZL22a]. top-cited [CLC23]. top-five [BP22]. top-ranked [LFV21, OdS20, OŠV^{+}24]. Top2Vec [LLM24]. Topic [SSH21, ZCW21, ABAAN^{+}23, BB22, BPL25, BMM22, CCZ^{+}20, CZZN25, CLS^{+}20, CKMY20, GHW22, Ger24, Han20, HLG21, HCL22, JPS21, JJY21, JCX22, LGD21, LLM24, LMC22, LCT^{+}20, ML24, MIY^{+}20, NZ23, Pra25b, Pur24, PFS^{+}20, QZF22, ROB21, Reh20, SHH23a, SHH23b, SMG^{+}24, SD22, TD24a, TWH^{+}24, VPDAG22, VPR22, WWH21, WYL21, XW25, ZGM^{+}23, ZSL24, ZWL^{+}25b, ZQS22]. topic-based [ML24]. topic-journal [BB22]. topic-level [TWH^{+}24]. topic-model [LLM24]. topic-specific [CLS^{+}20]. topicality [HHY22]. topics [AKYH21, BVvEW25, CN24, CC22, CR24, DK21a, FCT^{+}20, GWvZ21, Han20, HAT23, IHLP21, JHVC21, KY20, KFLS20, LCW21, LGE^{+}23, LC20, Lyu21, MB20b, MK23, OTH22, SLL23b, TD24a, TNSF21, WCCC24, WLS22, WUH21, XHA^{+}20, YZWH24, ZQS22]. topology [XWGD22]. TOPSIS [DD21b]. Torres [Rus23]. Torres-Salinas [Rus23]. Tourism [BB22, SBP20, SFC^{+}23]. Tower [WCL25]. Trace [WPSK25]. tracer [CHT24]. Tracing [HGZ21, Cha23a]. tracked [FC20a]. tracker [DTHM21]. Tracking [CHT24, KFM20, LTL23, LGE^{+}23, LSY21, PKA24, LGD21, LCT^{+}20, Tom24]. trade [WH25]. trade-off [WH25]. Trademarks [CM22]. trading [DWM21, EGV22]. Traditional [TM20]. traditions [CNP21]. trained [LCT^{+}24, SLL^{+}23a, ZNNA20]. training [YA21]. traits [WKC22]. trajectories [Fil21, GSP21, GDdZ22, JL25, LMC22, SSH21]. trajectory [LHXS20, NS23]. Transactions [SLX21, SAK22]. Transdisciplinary [LOGS21, Pur25, Kaj22]. transfer [CdSK21, Dip21, DG24, FESRMB21, HAL25, HZCM23, LZH20, SZK20, SYL^{+}24, ZW24b]. transfers [HZ21]. transform [EO24b]. transformations [CPFVPA25]. transformative [Sch24b]. transformer [AGDE25, ZZ24]. transformer-based [AGDE25]. transformers [MLA^{+}23]. transit [FG20]. transition [CPFVPA25, GXC25, MIY^{+}20, STM^{+}24, YLC^{+}20, YLH24, ZQML23]. translation [ZL23a, ZL23b]. translational [KLNW20, LTL23, LBW^{+}20]. transliteration [KMM22]. transparency [GSRG24, KK23, WWP20]. transportation [ZIS21]. travel [Plo25a]. traversal [LLH20]. treat [GVS20]. tree [Bor20b, FZC20, LGXH24]. trend [AJBBA23, CD22, LHXS20, LT20, MSJ23, SBKK21, TA22, TNS21, YP20, YSW^{+}20, ZMA23, Zha21b]. Trends [WUH21, APR25, AKYH21, ADDCMG^{+}25, BLL21, BSS24, BERD21, CKMY20, CR24, Dem22, FWDQ20, FFRRPR^{+}24, GMKG20, HLS23, IPP22, JHVC23, KK20a, KHT22, LS20a, LDDL24, LHF23, LZF^{+}24, Mar20, MFP21, PFPCMS20, SYK22, Tay23, TD24a, YD21, ZWHS22, dSMNdV^{+}21, TTBA20]. trendy [Pra23a, SD23a, SD23b]. triadic [Wad24]. trial [DTHM21, OMAAORCL23, SYHW20, THvB21]. trials [LWW24, NHAB21, THvB21]. triangular [GB24]. tribes [WHÅ24]. tribology [LLP21]. tribute [Mey25a, Raf25a, WW25]. trip [Ayk21]. Triple [JSLJ22, AdAFR24, Coo25, EL25, HLR21, Mey25a]. trivial [LUH21a, LUH21b]. troubles [Hen19, Hen20]. true [Man24]. Trump [AM22]. truncated [Bat24b, SV20]. trustworthiness [ES22]. truth [Li25]. turbulence [LLL22]. Turing [JYT21]. Turki [Bjø20a, Bjø20c]. Turkish [TG24, TYS22, Tut23, Tut24, TD24b]. Türkiye [Mar24]. turn [MWH20, ROB21]. turn-of-the-month [MWH20]. tutorial [HB22a]. tweet [WZHW25, HAS^{+}20]. tweeted [HJCT23]. tweeters [AYS24]. Tweeting [MH24b, YNL25]. tweets [AM22, AYS24, AS25, FCW22, HB21b, WZHW25]. Twenty [Gon23b, Kos22, MMG23, RAG25]. Twenty-fifth [Kos22]. twenty-first [MMG23]. twins [KDN22]. Twitter [YNL25, CP21, HKN23, HKNF23, dMMdCM22, OMFJ22, OMBP24, PHBG21, PHP22, SSK22, TK21, WZ21, WZHW25, YNO22, YMX21, ZW21b, vSDC20]. Twitterverse [CP21]. Two [CG24, FH22, LHXS20, ÁLGBdMA24, Bat24a, Bor20b, CSC23, CLC23, CMZ20c, FZhCC21, FRRBDD23, GZOC24, HWC22, CLY^{+}22, KBS21, LHC22, LHF20, MBL21, NSM24, OdSO20, PLH20, RSPC20, RJ20a, SdSCdS25, Sha24b, SMBD^{+}24, Tei20, WPSK25, ZS21, ZL20b]. Two-dimensional [FH22, HWC22, LHF20, ZS21]. two-fold [CSC23]. two-level [CMZ20c]. two-mode [Bat24a]. two-stage [FZhCC21, CLY^{+}22]. two-step [MBL21, PLH20, WPSK25]. two-way [LHC22]. type [AG23, CCZL23, DLK20, DGSVDG20, HB22b, KM20, MMFF25, PLH20, PT24]. types [CCH20, CG24, Ela25, FZC20, HB22b, HL21a, PS20, RGVQTS^{+}24, VdOS^{+}20, WZHW25, WPS22, ZPZ21].

U [Dip21]. U-multirank [Dip21]. U.S. [ATC20, CL20, HZ21, LS21c, RP23]. UAE [Bou20]. UK [AJLA21, BJ21, BKT20, CL21a, FMF20, KSAT24, LRSB21, RJ20a, ZZL^{+}20]. Ukrainian [ADH23, HM21, Hla25, MNK21, Naz25]. unarXive [SF20]. Unbiased [LL24b]. Uncertainty [LA20, CL21b, JJY21, LB24a, LPD21, NWZ23, NA24, YWW23, YLW20]. uncited [FN23, HY20, MMN20, MNM^{+}21]. uncitedness [DGSVDG20]. uncontainable [GA21]. uncontrollable [GA21]. uncover [Wan25]. uncovering [KL24]. undefined [MN23]. underexplored [GXC25]. underlying [ALR24, Cop20c, YMLL23]. undermines [VP21b]. understand [CD22]. Understanding [BY23, CF23, EXT^{+}21, HKNF23, Jia21, LGXH24, LXL25, MBL21, Por25, THJ^{+}23, THJ^{+}24, YLCY24, YCYZ24, Yen25, YLX25, uHBKK20, uHBKK21, BGAMS23, DD21b, MGA22, PR22, PKA24, RAL22, SM20, YWHS23]. Unequal [LKPS25]. unexpected [LRH22]. unified [LSBC21]. Union [GDM22, GKK23, GEFJ24, MCSOAA21, PSW24]. unit [FWDQ20, LPD21]. United [BL20b, GEFJ24, LSA22, MKWB^{+}23, MBH25, PL23, SO21, ZPH21]. units [Sch20a, WJ22, Sch22]. Universal [MNK21, Fas20a, Fas20b]. Universalism [KJY24]. universality [WWL24b, YGW^{+}24]. universitie [FYY21, TTDH25]. Universities [ACDS22, LCP22, AMMN21, AAD24a, AlYT21, Ali24, Ali25, ADDCMG^{+}25, AMIK23, Bor20b, CXD21, CTH22, CGJM24, DDL23, DMR20, DJM22, DCdA25, EGR22, Fas24, FRRBDD23, FFL21, DMR22, GAACBA24, HLS21, Hla25, HL21a, HHB^{+}20, HZ21, KBSS23, Kim23, KFCK21, LBCO21, LRP21, LdGRV^{+}22, MF20b, Men25, MHK22, Pin23, SV20, San21b, SZN^{+}21, SHB25, Sha24a, TMNH^{+}24, TBPN^{+}20, Wal22, WLLW25a, XlYqZ22, YLM24a]. University [Hla25, RSV^{+}20, RS22b, TMNH^{+}24, AJB^{+}21, BJ21, BA22b, BCG22, BKO25, CLL22, CZX24, CC21, CSJW22, Dem22, DJM22, EE24, FFN25, FRRBDD23, FSEJF^{+}20, FH22, Fuk22a, Fuk22b, GLK^{+}23, Ger24, JVY22, KKK^{+}25, LKPS25, Li25, LY20, LZD^{+}24, LdS25, dMMdCM22, MLYK23, MJYS23, Naz24b, NCD22, PSKS22, PMGÁC23, SD20, SJBBR^{+}23, SS20a, SdLV21b, SVC20, SS20b, SZK20, THJ^{+}23, Tut23, WL21b, WP22, WGZ22, KUS22, SGG20, THJ^{+}24]. university-industry [BJ21, CSJW22, GLK^{+}23, NCD22, WL21b]. university-industry-government [PSKS22]. unlabeled [CCSX23]. Unpacking [LMC22, RAS21]. unprecedented [HB20]. unprincipled [Cer24b]. unravels [OŠV^{+}24]. unscholarly [FN23]. unseen [BPL25]. Unsuccessful [OPS25]. unsupervised [BO20, FZL^{+}20a, FZL^{+}20b, LLZ25]. until [RdNB21]. unveil [AD21]. Unveiling [VBF24, YXDL24, LLH25]. Unverified [CSR23]. unwanted [Ans23b, dSV21, dSV23]. unweighted [Hal20b]. unwritten [Sch20b]. updated [LTH20]. upon [dMALIGBM20]. uptake [CFC22, HJT21, HKNF23, THJ^{+}23, THJ^{+}24]. Urban [vRW25, WG24]. USA [LRSB21]. Usage [Che18, VRVH21, Bir21, Bre21, DMA^{+}20, Mas20, OMC21, PXY24, TWH^{+}24, TFWC24, WGC20, WXZ22, ZWW21, ZT23, Che20]. Use [AJBB22, FRRBDD23, HFAR23, MNM^{+}21, SCW^{+}23, dOSAL21, Ben25, BL24, BWE24, CM20, DWG23, FLC22, FESRMB21, GLPN25, JHVC21, Kim24, KYKC21, LCC^{+}20, LS23, MSV23, Moo21, MFP21, MBH25, Öz24, Pea21, SdLV21a, Urb25, Ven24, VP21b, WY23, ZL20b, dS21a]. used [Asa23b, HB21b, Ibr21, MRM20, MHK22, dW24]. useful [Mor20, The20a]. usefulness [BWH20a, BWH20b]. User [FCW22, RBV25, WZHW25, YWHS23]. users [BCF21, YLCY20]. uses [Pet22]. Using [BFG^{+}21, CLC23, DV20, MLP^{+}20, MMG21, Mey25b, NLS20, Poh24, RSV^{+}20, RFG^{+}24, RZS23b, Thi20, TSA24, WG20b, WPBH24, AMMN21, AlYT21, AUA^{+}23, AGDE25, AMS21, BD21, BVvEW23, Ben24, BDTC20, BL20a, BH25, BY23, BGSRF20, Can20, CCH^{+}24, CB20, CPLS20, CB25, CC21, CCSX23, CLS21, CCC22, DX20, DLC21, FZhCC21, GK22, GSB24, GS24, HBW^{+}21, HC22, HMW24, HM20b, HCZ^{+}22, HWL23, HTZR24, IQ21, IPP22, JPS21, JJY21, KNT20, KP23b, KCC21, LNS21, LHK21, LYT^{+}20, LPD21, LMH22, LGMK21, Liu24, LGXH24, LS23, LC21, MZD22, MB20b, MCDFSP24, MK23, MRF21, MLA^{+}23, MS21e, MDGR21, MKI24, MJRR25, NKY^{+}24, NA24, OCKY20, OJKY23, PHBG21, PD20a, PLT22, PSS23, QJI^{+}22, QQL21, RSPC20, RKMG20, RSM24, RB25, RAL22, SV20, San20, San24, SGA22, Sca25, SSC^{+}23, Sch25b, SVY22]. using [SDC22, Sin22, SLL23b, SLTS25, Ste21, SA21, TKS^{+}23, TNS21, Tei24a, TD24a, TBA23, TSDAMRG24, UMA21, VCD^{+}23, VPDAG22, WWY^{+}20, WYT21, Wan25, WKC22, WRT20, XHQZ20, YZWH24, YWY^{+}20, YKS22, YKTM20, YWB21, ZA20, ZNNA20, ZWC22, ZY22, ZS24, ZZ24, ZBG^{+}24, ZW24a, ZYZZ25, ZAZS22, ZGBH20, ZDL^{+}20, ZYPC20, dOCP21, RZS23a, SKBSC20, ZHH23a]. USPTO [Wad20, XCML25]. utilizing [WAT23, WZ25]. Uzzi [WFW24].

vaccination [vSDC20]. valid [AS25]. validated [PD20a]. validating [Liu24, MSI21]. Validation [Don21b, VPDAG22, HTX23, WFW24]. validity [HLG21, Sun23, ZZW24, ZNW22]. valuable [SCK21]. Valuative [Lec21]. value [Ayk21, CKT21, GCS25, HJCT23, KTB22, Kra20, WZZ24, YMW^{+}20]. values [AMTS24, RM24, Sch20a, Sch22]. valuing [HM23]. vanishing [AMTS24]. variable [CF22, FBDM25]. variables [SS24, ZZZZ20, ZGBH20]. variance [FLDD20]. variant [BDTC20]. variants [CRBRGS21, NLS24]. variation [DWG23, LZ23, TSVV23]. variations [PSW24, SPS20]. varied [WWL24b]. variety [Mut22, ZCS^{+}23]. Various [GD22, BMS21b, YY22a, ZMK22]. vary [ADD20]. varying [SS22, ZM24]. vector [GS22a, WAT23]. vectorisation [KDZ21]. vehicle [CZX24]. vehicles [CLL22, QZ20]. velocity [FC20a]. ventilator [GAPR20]. ventilator-associated [GAPR20]. venture [HL23]. ventures [WSL23]. venue [MSV23]. veracity [SLMCSV21]. verbal [KR21]. versus [BP22, DDF20, GKDB25, GCT^{+}20, JP22, KHC^{+}25, LZY22, MZE20, SLMCSV21, SB25, SB21, Sma24, SSC21, UIM21, UA20]. via [Ans23a, Cor22, CXW^{+}24, FAY25, GAB21, JMZ^{+}23, LXS22, MMTOMLC21a, MMTOMLC21b, ZZ20a, Zho24]. Video [BRG^{+}23, Cop20a, Cop20b, ZW25]. Vienna [SGG20]. view [AAD24b, LXS22]. Viewing [DKWC20]. viewpoint [LLZ21]. views [BRG^{+}23, IHLP21, THG23, YWX23, ZS20b]. Village [Tom21]. virtual [FH21]. virtualization [OPZP25]. virtues [DV20]. vis [AAD22, ADG24]. vis-à-vis [AAD22, ADG24]. Visibility [Tom23, BSA23, CB25, DMR20, DABSC20, DGSVDG20, GC21, HM23, HIA^{+}20, HKN23, Jok20, LPY22, MZE20, SCDP21, STAÁMR24]. visible [Cha23b]. vision [MPR21, MPMR21, ON20]. vision-driven [ON20]. visiting [LWX21]. visits [OMFJSC24]. visual [ADS20, ADU22]. Visualization [Naz24b, CDZ^{+}20, GS22a, MIY^{+}20]. visualizations [LBW21]. visualize [GBP21]. Visualizing [BLL21, ML24, Pal21, PFPCMS20, WZL22b]. vitae [Cha23a]. Vocabulary [Ben20]. voices [BdMPH21]. volume [Kim24]. volumes [Zha21b, dSTE21]. VOSviewer [OMC21]. voting [CLWlY24]. vs [KR22a, LCP22, MJYS23, PY25, YJJ25].

W [Ano23, Sug23]. wake [RBBP24]. Wakefield [HP21, LAT22]. wall [Pra23b]. Waltman [Ano21a, vR21a]. Wang [BDTC20]. war [Naz25]. warning [Cop20d]. was [BJS22, SYHW20]. water [McC24, TPVVPA21, VCD^{+}23, WRT20]. wave [Ali22]. way [LHC22, MS21d]. ways [BH25, Hir22, Mey25a]. weakest [Ell23]. weakness [UKP21]. weaknesses [dS21a]. wealth [AD21, RNB22]. Web [Che20, MMTOMLC21a, BOPGOP25, Bre21, JS21, MB22a, OMAAORCL23, Par24, SNK22, AOM24, ARR23, BdOdS21, Bou21, Che18, CHJ^{+}25, EO24a, FCT^{+}20, GKM^{+}24, HWNL20, HWL23, KMM22, KA20, KS22c, LL20, LTH20, Liu21, LHW21, LNH24, MMTOMLC21b, MMB23, MN23, OdSO20, ODQ24, PLH20, SSSd22, SQC^{+}20, SSM21, SM23, SS22, VV21, ZL20b]. web-based [OMAAORCL23, SNK22]. web-scraping [BOPGOP25]. webometrics [Agu20, SZN^{+}21]. website [KBS20]. websites [OM21, SLMCSV21]. WeChat [CFC22, YLf25, ZXL^{+}24]. weight [Kua20]. Weighted [Bat24b, LWG21, Bat24a, BTD21, GAB21, Hal20b, HWZ21, RZS23a, RZS23b, SdS25]. weighting [ZS20b]. weights [Kua20]. welcome [Min22]. we’re [Gra20]. Western [DT20, ZP20]. Westernization [AM21a]. WGI [HDY^{+}23]. Where [BW20a, WZ25, AAD24a, Gus22, IHLP21, IVCE20]. whether [Bor20b, GRSFV22]. Which [ATC20, Asa23b, BVvEW25, HWNL20, LZYR24, SCK21, PJ24, TKM^{+}23, VP21a]. while [AYRG23]. Who [Cha23b, CdSS25, GBH21, LRW25, WZHW25, dSV23, IVCE20, KDN22, Lax23b, LZY22, Pra23b, WSRM23, Yur20, ZW25, ZQML23, dS21a]. whom [BCF21, KS22b]. wide [KCS25, VCV22]. wider [ZGBH20, ZZ20a]. Wikipedia [SKS^{+}23, YSWC24]. Wiley [Asa23a]. Will [HYSY23, XCML25, Zha21b, Gus22]. Williams [Lax23a]. willingly [Gub24]. willingness [KKK^{+}25]. window [CSL22]. windows [BFG^{+}21]. wine [JL20]. winners [AKD23, GDRPR23]. winning [DM25, KDN22, LRH22, LLH22, RWL23, WZZ24]. wins [Ano20a, Ano21a, Ano23]. wisdom [Gor21, SZY24]. wise [JC23, CWK21]. withdrawing [KPS21]. within [FSR21, GDG^{+}23, HWNL20, LM20, PR23, San21b, San25, SdS25, TZY^{+}23, THJ^{+}23, THJ^{+}24, Thi20, VPC^{+}23]. within-issue [LM20]. without [Kra20]. witty [TDM24]. Wolf [LR25]. Women [HSE23, SW24, EOL23, HdSV23, RFMMPJ22, WWL24a]. Word [TMNH^{+}24, CPLS20, CCZ^{+}20, Cor22, GPYR^{+}20, HBN^{+}21, HhC20, HCL22, JCX22, PKCOG21, TSA24, XZZ22a, XZZ22b, ZYZW23, ZZ20b, ZQS22]. word-topic [ZQS22]. word2vec [GHW22]. Words [AGWG^{+}23, Ben25, CLMM21, EBW23, LCB20, WWL24b]. work [AJB^{+}21, ASSB20, LGD21, LLC22, Pur25, RLZ^{+}24, TYS22, Wag25, ZHH23b]. work-energy [RLZ^{+}24]. worker [Yur20]. Working [EGP22, BW20a, BW22, GCH^{+}22, WB21]. works [HM20a, WZZ23]. World [ATC20, Hla25, KPM22, AHH23, ABK20, DT20, ER24, LRP21, MMGP25, McC24, Old22, Rod22, SM22b, ZL23a, ZL23b, Fas24]. world-systemic [DT20]. WorldCat [TSAMT21]. worth [The21]. worthiness [ZA20]. WoS [Che18, Che20, BBP21, BRJ21, MML^{+}20, PD20a]. Write [Fag20]. writing [BZZ25, DFG21, DWG23, HBC25, Li22, MTD^{+}23, Pea20, Ste22, WWS^{+}23b, WVH21, YWW23, mYY22]. writings [SLX21]. written [JVY22, Pea20]. wrongs [GZOC24]. Wu [BDTC20]. WURS [SS20b]. WWII [Mit20].

X [YNL25]. XGBoost [ZJY22]. XIX [GMCRV21]. XML [Bös21]. Xplore [TWH^{+}24, Tom21]. XSEDE [SCW^{+}23].

Year [WPBH24, BC21, CxGdH^{+}24, CZZN25, EGPGGA^{+}23, GTB23, HB22a, MCDFSP24, MKWB^{+}23, SH20, SJW21, SSBC22, TKZ^{+}23]. years [ASIN24, FFRRPR^{+}24, GSG22a, Gon23b, HBO^{+}20, HBC25, MCDFSP24, MSJ23, Naz25, RWL23, RAG25, SGM22, TNSF21, WL22, YG20, YWW23, mYY22]. YERUN [DMR20]. yesterday [Zan23]. young [CW25, KDN22, San21b, ŞKKT25, WZZ25, lYnLyO^{+}20]. younger [SO21]. YouTube [SAS23, ZW25]. Yule [WR21].

Zealand [BCG22]. zero [BFG^{+}21, Pur24, SWT20]. zero-modified [SWT20]. Zeyuan [Ano21e, CL21c, LGD21, ZLL21]. Zimbabwe [NB20, NB22]. Zipf [GS24]. zone [SSSd22]. Zong [Cop20a].

References

Asubiaro:2022:SSB

[AA22] Toluwase Victor Asubiaro and Isola Ajiferuke. Semantic similarity-based credit attribution on citation paths: a method for allocating residual citation to and investigating depth of influence of scientific communications. Scientometrics, 127(11):6257–6277, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04522-3.

Alshraideh:2025:CQR

[AA25] Hussam Alshraideh and Mohamed Abdelgawad. On the credibility of QS and THE ranking by subject area: misalignment of subject mapping to academic disciplines. Scientometrics, 130(4):2237–2272, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05266-6.

Abramo:2023:REP

[AAA^{+}23] Giovanni Abramo, Isidro F. Aguillo, Dag W. Aksnes, Kevin Boyack, Quentin L. Burrell, Juan Miguel Campanario, Zaida Chinchilla-Rodríguez, Rodrigo Costas, Ciriaco Andrea D’Angelo, Anne-Wil Harzing, Hamid R. Jamali, Vincent Larivière, Loet Leydesdorff, Marc Luwel, Ben Martin, Philipp Mayr, Katherine W. McCain, Isabella Peters, Ismael Rafols, Nicolas Robinson-Garcia, Torben Schubert, Henry Small, Cassidy R. Sugimoto, Mike Thelwall, Peter van den Besselaar, Thed van Leeuwen, and Ludo Waltman. Retraction of Predatory publishing in Scopus: evidence on cross-country differences lacks justification. Scientometrics, 128(2):1459–1461, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04565-6.

AlShareef:2023:IFI

[AAAA23] Mohammed R. AlShareef, Ibrahim A. Alrammah, Nasser A. Alshoukani, and Abdulaziz M. Almalik. The impact of financial incentives on research production: Evidence from Saudi Arabia. Scientometrics, 128(5):3067–3089, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04692-8.

Asanov:2024:PDD

[AAB^{+}24] Anastasiya-Mariya Asanov, Igor Asanov, Guido Buenstorf, Valon Kadriu, and Pia Schoch. Patterns of dissertation dissemination: publication-based outcomes of doctoral theses in the social sciences. Scientometrics, 129(4):2389–2405, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04952-1.

Abramo:2022:GPE

[AAD22] Giovanni Abramo, Francesca Apponi, and Ciriaco Andrea D’Angelo. The geographic proximity effect on domestic cross-sector vis-à-vis intra-sector research collaborations. Scientometrics, 127(6):3505–3521, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04404-8.

Abramo:2024:DRU

[AAD24a] Giovanni Abramo, Francesca Apponi, and Ciriaco Andrea D’Angelo. Do research universities specialize in disciplines where they hold a competitive advantage? Scientometrics, 129(9):5453–5468, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05136-7.

Abramo:2024:MRT

[AAD24b] Giovanni Abramo, Francesca Apponi, and Ciriaco Andrea D’Angelo. The moderating role of the territorial research infrastructure on the geographic proximity effect in research collaborations: a regional-based view. Scientometrics, 129(6):3149–3168, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05025-z.

Andersson:2020:GSE

[AAHY20] David Emanuel Andersson, Åke E. Andersson, Björn Hårsman, and Xiyi Yang. The geography of science in 12 European countries: a NUTS2-level analysis. Scientometrics, 124(2):1099–1125, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03510-9.

Alonso-Alvarez:2024:COA

[AÁSTMR24] Patricia Alonso-Álvarez, Pablo Sastrón-Toledo, and Jorge Mañana-Rodriguez. The cost of open access: comparing public projects’ budgets and article processing charges expenditure. Scientometrics, 129(10):6149–6166, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04988-3.

Abalkina:2021:DNH

[Aba21] Anna Abalkina. Detecting a network of hijacked journals by its archive. Scientometrics, 126(8):7123–7148, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04056-0.

Al-Betar:2023:OSP

[ABAAN^{+}23] Mohammed Azmi Al-Betar, Ammar Kamal Abasi, Ghazi Al-Naymat, Kamran Arshad, and Sharif Naser Makhadmeh. Optimization of scientific publications clustering with ensemble approach for topic extraction. Scientometrics, 128(5):2819–2877, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04674-w.

Andrikopoulos:2020:SSW

[ABK20] Andreas Andrikopoulos, Michael Bekiaris, and Konstantinos Kostaris. Stars in a small world: social networks in auditing research. Scientometrics, 122(1):625–643, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03272-z.

Argiles-Bosch:2025:IAE

[ABKGBR25] Josep Maria Argilés-Bosch, Yuliya Kasperskaya, Josep Garcia-Blandon, and Diego Ravenda. The interplay of author and editor gender in acceptance delays: evidence from accounting journals. Scientometrics, 130(3):1939–1965, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05259-5.

Aria:2024:CSM

[ACD^{+}24] Massimo Aria, Corrado Cuccurullo, Luca D’Aniello, Michelangelo Misuraca, and Maria Spano. Comparative science mapping: a novel conceptual structure analysis with metadata. Scientometrics, 129(11):7055–7081, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05161-6.

Ascione:2022:UIP

[ACDS22] Grazia Sveva Ascione, Laura Ciucci, Claudio Detotto, and Valerio Sterzi. Universities involvement in patent litigation: an analysis of the characteristics of US litigated patents. Scientometrics, 127(12):6855–6879, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04500-9.

Aranzales:2023:FHT

[ACT23] Iván Aranzales, Ho Fai Chan, and Benno Torgler. Finally! How time lapse in Nobel Prize reception affects emotionality in the Nobel Prize banquet speeches. Scientometrics, 128(7):4089–4115, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04739-w.

Ancona:2023:NMD

[ACV23] Andrea Ancona, Roy Cerqueti, and Gianluca Vagnani. A novel methodology to disambiguate organization names: an application to EU Framework Programmes data. Scientometrics, 128(8):4447–4474, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04746-x.

Abramo:2020:DLK

[AD20] Giovanni Abramo and Ciriaco Andrea D’Angelo. The domestic localization of knowledge flows as evidenced by publication citation: the case of Italy. Scientometrics, 125(2):1305–1329, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03487-5.

Abramo:2021:BMU

[AD21] Giovanni Abramo and Ciriaco Andrea D’Angelo. A bibliometric methodology to unveil territorial inequities in the scientific wealth to combat COVID-19. Scientometrics, 126(8):6601–6624, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04017-7.

Acuna:2024:MTH

[AdAFR24] Miguel A. Ortiz Acuña, Adiel T. de Almeida Filho, and Francisco S. Ramos. Modelling the triple helix system innovation of the main economies from Latin America: a coalitional game theory approach. Scientometrics, 129(6): 3245–3270, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05020-4.

Abramo:2020:DGP

[ADD20] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Does the geographic proximity effect on knowledge spillovers vary across research fields? Scientometrics, 123(2):1021–1036, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03411-x.

Abramo:2022:HCC

[ADD22] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. How the Covid-19 crisis shaped research collaboration behaviour. Scientometrics, 127(8):5053–5071, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04450-2.

Abramo:2023:CAC

[ADD23] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Correlating article citedness and journal impact: an empirical investigation by field on a largescale dataset. Scientometrics, 128(3):1877–1894, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04622-0.

Abramo:2024:DRA

[ADD24] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Do research assessment systems have the potential to hinder scientists from diversifying their research pursuits? Scientometrics, 129(10):5915–5935, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04959-8.

Arcila-Diaz:2025:RTP

[ADDCMG^{+}25] Juan Arcila-Diaz, Jorge Delgado-Caramutti, Pablo A. Millones-Gómez, Joel Figueroa-Quiñones, and Alejandro Valencia-Arias. Research trends in Peruvian universities: proposal for a research agenda with a bibliometric approach. Scientometrics, 130(1):489–514, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05211-z.

Abramo:2024:RNS

[ADG24] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Leonardo Grilli. The role of non-scientific factors vis-à-vis the quality of publications in determining their scholarly impact. Scientometrics, 129(8):5003–5019, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05106-z.

Abramo:2025:BRP

[ADGS25] Giovanni Abramo, Ciriaco Andrea D’Angelo, Edita Gzoyan, and Shushanik Sargsyan. Benchmarking research performance in a post-Soviet science system: the case of Armenia. Scientometrics, 130(4):2213–2235, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05312-3.

Abramo:2023:AEP

[ADH23] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Myroslava Hladchenko. Assessing the effects of publication requirements for professorship on research performance and publishing behaviour of Ukrainian academics. Scientometrics, 128(8):4589–4609, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04753-y.

Abramo:2022:ICR

[ADM22] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Ida Mele. Impact of Covid-19 on research output by gender across countries. Scientometrics, 127(12):6811–6826, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04245-x.

Angelini:2020:PMD

[ADS20] Marco Angelini, Cinzia Daraio, and Giuseppe Santucci. Performance model’s development: a novel approach encompassing ontology-based data access and visual analytics. Scientometrics, 125(2):865–892, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03689-x.

Angelini:2022:VAA

[ADU22] Marco Angelini, Cinzia Daraio, and Luca Urban. A visual analytics approach for the assessment of information quality of performance models — a software review. Scientometrics, 127(12):6827–6853, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04399-2.

Altuntas:2020:CBA

[AED20] Serkan Altuntas, Zulfiye Erdogan, and Turkay Dereli. A clustering-based approach for the evaluation of candidate emerging technologies. Scientometrics, 124(2):1157–1177, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03535-0.

Aljohani:2021:TCC

[AFH21] Naif Radi Aljohani, Ayman Fayoumi, and Saeed-Ul Hassan. An in-text citation classification predictive model for a scholarly search system. Scientometrics, 126(7): 5509–5529, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03986-z.

Arhiliuc:2023:DCR

[AG23] Cristina Arhiliuc and Raf Guns. Disciplinary collaboration rates in the social sciences and humanities: what is the influence of classification type? Scientometrics, 128(6): 3419–3436, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04719-0.

Ariza-Guerrero:2023:ENC

[AGB23] Ana M. Ariza-Guerrero and J. Sebastián Blázquez. Evolution of number of citations per article in Materials Science: possible causes and effect on the impact factor of journals. Scientometrics, 128(12):6589–6609, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04863-7.

Arhiliuc:2025:JAC

[AGDE25] Cristina Arhiliuc, Raf Guns, Walter Daelemans, and Tim C. E. Engels. Journal article classification using abstracts: a comparison of classical and transformer-based machine learning methods. Scientometrics, 130(1):313–342, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05217-7.

Artigas:2022:OAA

[AGL22] Wileidys Artigas, Eurico Wongo Gungula, and Mikael Laakso. Open access in Angola: a survey among higher education institutions. Scientometrics, 127(7):3977–3993, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04410-w.

Aguillo:2020:AOA

[Agu20] Isidro F. Aguillo. Altmetrics of the Open Access Institutional Repositories: a webometrics approach. Scientometrics, 123(3):1181–1192, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03424-6.

Argueta-Guzman:2023:WMH

[AGWG^{+}23] Magda Argueta-Guzmán, Mari West, Marilia P. Gaiarsa, Christopher W. Allen, Jacob M. Cecala, Lauren Gedlinske, Quinn S. McFrederick, Amy C. Murillo, Madison Sankovitz, and Erin E. Wilson Rankin. Words matter: how ecologists discuss managed and non-managed bees and birds. Scientometrics, 128(3):1745–1764, March 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04620-2.

Albanna:2021:POA

[AHH21] Basma Albanna, Julia Handl, and Richard Heeks. Publication outperformance among global South researchers: an analysis of individual-level and publication-level predictors of positive deviance. Scientometrics, 126(10):8375–8431, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04128-1.

Ali:2023:AIB

[AHH23] Nisar Ali, Zahid Halim, and Syed Fawad Hussain. An artificial intelligence-based framework for data-driven categorization of computer scientists: a case study of world’s Top 10 computing departments. Scientometrics, 128(3): 1513–1545, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04627-9.

Andersen:2024:TSG

[AHRH24] Jens Peter Andersen, Serge P. J. M. Horbach, and Tony Ross-Hellauer. Through the secret gate: a study of member-contributed submissions in PNAS. Scientometrics, 129(9):5673–5687, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05115-y.

Aboagye:2021:IAB

[AJB^{+}21] Emmanuel Aboagye, Irene Jensen, Gunnar Bergström, Elisabeth Björk Brämberg, Oscar Javier Pico-Espinosa, and Christina Björklund. Investigating the association between publication performance and the work environment of university research academics: a systematic review. Scientometrics, 126(4):3283–3301, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03820-y;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03820-y.pdf.

Al-Jamimi:2022:UBR

[AJBB22] Hamdi A. Al-Jamimi, Galal M. BinMakhashen, and Lutz Bornmann. Use of bibliometrics for research evaluation in emerging markets economies: a review and discussion of bibliometric indicators. Scientometrics, 127(10):5879– 5930, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04490-8.

Al-Jamimi:2023:SAR

[AJBBA23] Hamdi A. Al-Jamimi, Galal M. BinMakhashen, Lutz Bornmann, and Yousif Ahmed Al Wajih. Saudi Arabia research: academic insights and trend analysis. Scientometrics, 128(10):5595–5627, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04797-0.

Al-Janabi:2021:DTA

[AJLA21] Shahd Al-Janabi, Lee Wei Lim, and Luca Aquili. Development of a tool to accurately predict UK REF funding allocation. Scientometrics, 126(9):8049–8062, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04030-w.

Arencibia-Jorge:2022:ESM

[AJVACC22] Ricardo Arencibia-Jorge, Rosa Lidia Vega-Almeida, and Humberto Carrillo-Calvet. Evolutionary stages and multidisciplinary nature of artificial intelligence research. Scientometrics, 127(9):5139–5158, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04477-5.

Alkan:2023:KDT

[AKD23] Bilal Baris Alkan, Leyla Karakus, and Bekir Direkci. Knowledge discovery from the texts of Nobel Prize winners in literature: sentiment analysis and Latent Dirichlet Allocation. Scientometrics, 128(9):5311–5334, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04783-6.

Alharbey:2022:IID

[AKH22] Riad Alharbey, Jong In Kim, and Malik Khizar Hayat. Indexing important drugs from medical literature. Scientometrics, 127(5):2661–2681, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04340-7.

Amirkhanyan:2023:DMF

[AKW23] Hayk Amirkhanyan, Michał Krawczyk, and Maciej Wilamowski. Do male and female authors employ different journal choice strategies? Scientometrics, 128(11):5905– 5928, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04829-9.

Amiri:2021:RTT

[AKYH21] Babak Amiri, Ramin Karimianghadim, Navid Yazdanjue, and Liaquat Hossain. Research topics and trends of the hashtag recommendation domain. Scientometrics, 126(4): 2689–2735, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-021-03874-6.

Abalkina:2020:RCP

[AL20] Anna Abalkina and Alexander Libman. The real costs of plagiarism: Russian governors, plagiarized PhD theses, and infrastructure in Russian regions. Scientometrics, 125(3):2793–2820, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03716-x.

Amoroso:2021:IPP

[AL21] Sara Amoroso and Albert N. Link. Intellectual property protection mechanisms and the characteristics of founding teams. Scientometrics, 126(9):7329–7350, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04098-4.

AlvesVieira:2024:PBR

[AL24] Gabriel Alves Vieira and Jacqueline Leta. biblioverlap: an R package for document matching across bibliographic datasets. Scientometrics, 129(7):4513–4527, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05065-5.

Alvarez-Llorente:2024:NFC

[ÁLGBdMA24] Jesús M. Álvarez-Llorente, Vicente P. Guerrero-Bote, and Félix de Moya-Anegón. New fractional classifications of papers based on two generations of references and on the ASJC Scopus scheme. Scientometrics, 129(6):3493–3515, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05030-2.

Ali:2021:ATP

[Ali21a] Mona Farouk Ali. Attitudes towards plagiarism among faculty members in Egypt: a cross-sectional study. Scientometrics, 126(4):3535–3547, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03872-8. See correction [Ali21b].

Ali:2021:CAT

[Ali21b] Mona Farouk Ali. Correction to: Attitudes towards plagiarism among faculty members in Egypt: a crosssectional study. Scientometrics, 126(4):3549, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03935-w;   http:
//link.springer.com/content/pdf/10.1007/s11192-
021-03935-w.pdf. See [Ali21a].

Ali:2022:BPM

[Ali22] Mona Farouk Ali. Between panic and motivation: did the first wave of COVID-19 affect scientific publishing in Mediterranean countries? Scientometrics, 127(6): 3083–3115, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04391-w.

Ali:2024:TDD

[Ali24] Mona Farouk Ali. Is there a “difference-in-difference”? The impact of scientometric evaluation on the evolution of international publications in Egyptian universities and research centres. Scientometrics, 129(2):1119–1154, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04911-2.

Ali:2025:ISP

[Ali25] Mona Farouk Ali. Investigating shifts in publication patterns after launching scientometric evaluation at Egyptian universities: an analysis of submitted research for promotion. Scientometrics, 130(3):1751–1787, March 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05257-7.

Azar:2023:MIS

[ALM^{+}23] Marleine Azar, Francois Lagacé, Anastasiya Muntyanu, Elena Netchiporouk, Youwen Zhou, Charles Lynde, Linda Moreau, Steve Mathieu, Denis Sasseville, Rachel Asiniwasis, Neil H. Shear, Robert Gniadecki, Elham Rahme, and Ivan V. Litvinov. Measuring h-index and scholarly productivity in academic dermatology in Canada.

Scientometrics, 128(2):1071–1090, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04589-y.

Alexi:2024:SRU

[ALR24] Ariel Alexi, Teddy Lazebnik, and Ariel Rosenfeld. The scientometrics and reciprocality underlying co-authorship panels in Google Scholar profiles. Scientometrics, 129(6): 3303–3313, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05026-y.

Agasisti:2021:EHE

[AlYT21] Tommaso Agasisti, Guo liang Yang, and Carolyn-Thi Thanh Dung Tran. Evaluating the higher education productivity of Chinese and European “elite” universities using a meta-frontier approach. Scientometrics, 126(7): 5819–5853, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03978-z.

Ai:2021:WJR

[AM21a] Minwei Ai and Muhammad Masood. De-westernization in journalism research: a content and network analysis of the BRICS journals. Scientometrics, 126(12):9477–9498, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04194-5.

Amjad:2021:IIC

[AM21b] Tehmina Amjad and Javeria Munir. Investigating the impact of collaboration with authority authors: a case study of bibliographic data in field of philosophy. Scientometrics, 126(5):4333–4353, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03930-1.

Allen:2022:TCT

[AM22] David E. Allen and Michael McAleer. Trump’s COVID-19 tweets and Dr. Fauci’s emails. Scientometrics, 127(3):

1643–1655, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04243-z.

Ayyildiz:2023:NHM

[AMIK23] Ertugrul Ayyildiz, Mirac Murat, Gul Imamoglu, and Yildiz Kose. A novel hybrid MCDM approach to evaluate universities based on student perspective. Scientometrics, 128(1):55–86, January 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04534-z.

Abdul-Majeed:2021:MRP

[AMMN21] Ghassan Abdul-Majeed, Wissam Mahmood, and Nasri S. M. Namer. Measuring research performance of Iraqi universities using Scopus data. Scientometrics, 126(3): 2349–2363, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03866-6.

Ashouri:2021:AAI

[AMS21] Sajad Ashouri, Anne-Laure Mention, and Kosmas X. Smyrnios. Anticipation and analysis of industry convergence using patent-level indicators. Scientometrics, 126 (7):5727–5758, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04025-7.

Abdul-Majeed:2023:INR

[AMSS^{+}23] Ghassan Abdul-Majeed, Elameer Amer Saleem, Drai A. Smait, Sadiq H. Abdulhussain, Sadiq M. Sait, Hasan S. Majdi, Haydar Abdulameer Marhoon, and Waleed Khalid Al-Azzawi. Implementation of a new research indicator to QS ranking system. Scientometrics, 128(2):1351– 1365, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04611-3.

Arroyo-Machado:2024:STV

[AMTS24] Wenceslao Arroyo-Machado and Daniel Torres-Salinas. Stranger things: the vanishing of the Altmetric Attention Score values in information and library science.

Scientometrics, 129(10):6287–6300, October 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05011-5.

Arroyo-Machado:2021:ICS

[AMTSRG21] Wenceslao Arroyo-Machado, Daniel Torres-Salinas, and Nicolas Robinson-Garcia. Identifying and characterizing social media communities: a socio-semantic network approach to altmetrics. Scientometrics, 126(11):9267– 9289, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04167-8.

Atapour:2024:PCS

[AMZ24] Hashem Atapour, Robabeh Maddahi, and Rasoul Zavaraqi. Policy citations of scientometric articles: an altmetric study. Scientometrics, 129(7):4423–4436, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05091-3.

Anonymous:2020:LBW

[Ano20a] Anonymous. Lutz Bornmann wins the 2019 Derek John de Solla Price Medal. Scientometrics, 122(1):1–2, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03290-x.

Anonymous:2020:O

[Ano20b] Anonymous. Obituary. Scientometrics, 125(1):829–830, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03653-9.

Anonymous:2020:SLM

[Ano20c] Anonymous. Scholarly literature mining with information retrieval and natural language processing. Scientometrics, 125(3):2833, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03781-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03781-2.pdf.

Anonymous:2021:LWW

[Ano21a] Anonymous. Ludo Waltman wins the 2021 Derek John de Solla Price Medal. Scientometrics, 126(10):8233–8234, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04116-5.

Anonymous:2021:O

[Ano21b] Anonymous. Obituary. Scientometrics, 126(4):3693–3695, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03903-4.

Anonymous:2021:QQS

[Ano21c] Anonymous. Quantitative and qualitative studies of science and technology in Latin America. Scientometrics, 126(3):2409, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03921-2.

Anonymous:2021:QSS

[Ano21d] Anonymous. Quantitative studies of science in Germany. Scientometrics, 126(12):9639, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04202-8.

Anonymous:2021:ZLM

[Ano21e] Anonymous. Zeyuan Liu memorial issue. Scientometrics, 126(7):6131, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04074-y.

Anonymous:2022:AIM

[Ano22a] Anonymous. AI + Informetrics: Multi-disciplinary interactions in the era of Big Data. Scientometrics, 127(11): 6501, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04560-x.

Anonymous:2022:BDD

[Ano22b] Anonymous. “Big Data” driven tech mining and ST&I management. Scientometrics, 127(9):5225, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04506-3.

Anonymous:2023:TKW

[Ano23] Anonymous. The team Kevin W. Boyack and Richard Klavans wins the 2023 Derek John de Solla Price Medal. Scientometrics, 128(12):6307–6308, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04847-7.

Anonymous:2024:EEK

[Ano24] Anonymous. Extraction and evaluation of knowledge entities from scientific documents. Scientometrics, 129(11): 7167, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05203-z.

Andrea:2021:ERO

[ANR21] Sixto-Costoya Andrea, Robinson-Garcia Nicolas, and Costas Rodrigo. Exploring the relevance of ORCID as a source of study of data sharing activities at the individuallevel: a methodological discussion. Scientometrics, 126(8): 7149–7165, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04043-5.

Ansorge:2023:HLA

[Ans23a] Libor Ansorge. Hidden limitations of analyses via alternative bibliometric services. Scientometrics, 128(3): 2031–2033, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04626-w.

Ansorge:2023:RRU

[Ans23b] Libor Ansorge. The right to reject an unwanted citations: do we need it? Scientometrics, 128(7):4147–4150, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04702-9.

Ante:2020:PNS

[Ant20] Lennart Ante. A place next to Satoshi: foundations of blockchain and cryptocurrency research in business and economics. Scientometrics, 124(2):1305–1333, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03492-8;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03492-8.pdf.

Asubiaro:2024:RDW

[AOM24] Toluwase Asubiaro, Sodiq Onaolapo, and David Mills. Regional disparities in Web of Science and Scopus journal coverage. Scientometrics, 129(3):1469–1491, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04948-x.

Aksnes:2019:GGI

[APR19] Dag W. Aksnes, Fredrik Niclas Piro, and Kristoffer Rørstad. Gender gaps in international research collaboration: a bibliometric approach. Scientometrics, 120(2): 747–774, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-019-03155-3.

Akbash:2025:NJS

[APR25] Kateryna Akbash, Natalya Pasichnyk, and Renat Rizhniak. Numismatic journals from the Scopus scientometric database: statistics, trends, collaboration. Scientometrics, 130(3):1967–1987, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05273-7.

Ali:2022:SSS

[AQK^{+}22] Zafar Ali, Guilin Qi, Pavlos Kefalas, Shah Khusro, Inayat Khan, and Khan Muhammad. SPR-SMN: scientific paper recommendation employing SPECTER with memory network. Scientometrics, 127(11):6763–6785, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04425-3.

Araujo:2020:CAN

[Ara20] Ronaldo Ferreira Araujo. Communities of attention networks: introducing qualitative and conversational perspectives for altmetrics. Scientometrics, 124(3):1793– 1809, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03566-7.

Amjad:2020:SIA

[ARDA20] Tehmina Amjad, Yusra Rehmat, Ali Daud, and Rabeeh Ayaz Abbasi. Scientific impact of an author and role of selfcitations. Scientometrics, 122(2):915–932, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03334-2.

Avila-Robinson:2021:BMC

[ÁRMS21] Alfonso Ávila-Robinson, Cristian Mejia, and Shintaro Sengoku. Are bibliometric measures consistent with scientists’ perceptions? The case of interdisciplinarity in research. Scientometrics, 126(9):7477–7502, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04048-0.

Aviv-Reuven:2021:PPC

[ARR21] Shir Aviv-Reuven and Ariel Rosenfeld. Publication patterns’ changes due to the COVID-19 pandemic: a longitudinal and short-term scientometric analysis. Scientometrics, 126(8):6761–6784, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04059-x.

Aviv-Reuven:2023:LST

[ARR23] Shir Aviv-Reuven and Ariel Rosenfeld. A logical set theory approach to journal subject classification analysis: intrasystem irregularities and inter-system discrepancies in Web of Science and Scopus. Scientometrics, 128(1):157–175, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04576-3.

Akca:2023:GRE

[AS23] Sümeyye Akça and Özlem Senyurt. Geographical representation of editorial boards: a review in the field of library and information sciences. Scientometrics, 128(2): 1409–1427, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04614-0.

Adams:2024:NRI

[AS24] Jonathan Adams and Martin Szomszor. National research impact is driven by global collaboration, not rising performance. Scientometrics, 129(5):2883–2896, May 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05010-6.

Amiri:2025:COM

[AS25] Mahsa Amiri and Hajar Sotudeh. Comparative opinion mining of tweets on retracted papers and their valid peers: a semi-experimental follow-up. Scientometrics, 130(2): 1159–1179, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05231-3.

Asai:2020:ECL

[Asa20a] Sumiko Asai. The effect of collaboration with large publishers on the internationality and influence of open access journals for research institutions. Scientometrics, 124 (1):663–677, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03426-4.

Asai:2020:MPP

[Asa20b] Sumiko Asai. Market power of publishers in setting article processing charges for open access journals. Scientometrics, 123(2):1037–1049, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03402-y.

Asai:2021:CBR

[Asa21] Sumiko Asai. Collaboration between research institutes and large and small publishers for publishing open access journals. Scientometrics, 126(6):5245–5262, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03949-4.

Asai:2023:DDD

[Asa23a] Sumiko Asai. Does double dipping occur? The case of Wiley’s hybrid journals. Scientometrics, 128(9):5159– 5168, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04800-8.

Asai:2023:WDA

[Asa23b] Sumiko Asai. Which database with article processing charges should be used? Scientometrics, 128(11):6293– 6298, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04841-z.

Asai:2024:DMS

[Asa24] Sumiko Asai. Determinants of manuscript submissions to fully open access journals: elasticity to article processing charges. Scientometrics, 129(3):1687–1696, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04934-3.

Aiza:2024:FTE

[ASIN24] Wan Siti Nur Aiza, Liyana Shuib, Norisma Idris, and Nur Baiti Afini Normadhi. Features, techniques and evaluation in predicting articles’ citations: a review from years 2010– 2023. Scientometrics, 129(1):1–29, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04845-9.

Amjad:2022:CBP

[ASK22] Tehmina Amjad, Nafeesa Shahid, and Asma Khatoon. Citation burst prediction in a bibliometric network. Scientometrics, 127(5):2773–2790, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04344-3.

AbuRaed:2020:ARW

[ASSB20] Ahmed AbuRa’ed, Horacio Saggion, Alexander Shvets, and Àlex Bravo. Automatic related work section generation: experiments in scientific document abstracting. Scientometrics, 125(3):3159–3185, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03630-2.

An:2022:ICI

[ASX22] Xin An, Xin Sun, and Shuo Xu. Important citations identification with semi-supervised classification model. Scientometrics, 127(11):6533–6555, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04212-6.

Alsmadi:2020:UNW

[ATC20] Izzat Alsmadi, Z. W. Taylor, and Joshua Childs. U.S. News & World Report Best Colleges rankings: Which institutional metrics contribute to sustained stratification? Scientometrics, 124(3):1851–1869, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03560-z.

Akbaritabar:2020:ISC

[ATCS20] Aliakbar Akbaritabar, Vincent Antonio Traag, Alberto Caimo, and Flaminio Squazzoni. Italian sociologists: a community of disconnected groups. Scientometrics, 124(3):2361–2382, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03555-w.

Akbulut:2020:RRR

[ATW20] Müge Akbulut, Yasar Tonta, and Howard D. White. Related records retrieval and pennant retrieval: an exploratory case study. Scientometrics, 122(2):957–987, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03303-9.

Ash:2020:RBR

[AU20] Elliott Ash and Miguel Urquiola. A research-based ranking of public policy schools. Scientometrics, 125 (1):499–531, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03625-z;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03625-z.pdf.

Alshdadi:2023:FRS

[AUA^{+}23] Abdulrahman A. Alshdadi, Muhammad Usman, Madini O. Alassafi, Muhammad Tanvir Afzal, and Rayed Al-Ghamdi. Formulation of rules for the scientific community using deep learning. Scientometrics, 128(3):1825–1852, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04633-5.

Ali:2021:COE

[AUM21a] Zafar Ali, Irfan Ullah, and Khan Muhammad. Correction to: An overview and evaluation of citation recommendation models. Scientometrics, 126(10):8771, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04065-z. See [AUM21b].

Ali:2021:OEC

[AUM21b] Zafar Ali, Irfan Ullah, and Khan Muhammad. An overview and evaluation of citation recommendation models. Scientometrics, 126(5):4083–4119, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03909-y. See correction [AUM21a].

Avdeev:2021:ICH

[Avd21] Stanislav Avdeev. International collaboration in higher education research: a gravity model approach. Scientometrics, 126(7):5569–5588, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04008-8.

Aykac:2021:VOR

[Ayk21] Gokhan Aykac. The value of an overseas research trip. Scientometrics, 126(8):7097–7122, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04052-4.

Aydin:2023:MBP

[AYRG23] Abdulkerim Aydin, Süleyman Eren Yürük, İlknur Reisoğlu, and Yuksel Goktas. Main barriers and possible enablers of academicians while publishing. Scientometrics, 128(1): 623–650, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04528-x.

Amiri:2024:HDT

[AYS24] Mahsa Amiri, Maryam Yaghtin, and Hajar Sotudeh. How do tweeters feel about scientific misinformation: an infoveillance sentiment analysis of tweets on retraction notices and retracted papers. Scientometrics, 129(1):261– 287, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04871-7.

Azmeh:2022:QQR

[Azm22] Chadi Azmeh. Quantity and quality of research output and economic growth: empirical investigation for all research areas in the MENA countries. Scientometrics, 127(11):6147–6163, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04512-5.

Bertin:2022:PCP

[BA22a] Marc Bertin and Iana Atanassova. Preprint citation practice in PLOS. Scientometrics, 127(12):6895–6912, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04388-5.

Bouabid:2022:SST

[BA22b] Hamid Bouabid and Hind Achachi. Size of science team at university and internal co-publications: science policy implications. Scientometrics, 127(12):6993–7013, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04285-x.

Bertin:2024:LPD

[BA24] Marc Bertin and Iana Atanassova. Linguistic perspectives in deciphering citation function classification. Scientometrics, 129(10):6301–6313, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05082-4.

Bharti:2024:PPY

[BAE24] Prabhat Kumar Bharti, Mayank Agarwal, and Asif Ekbal. Please be polite to your peers: a multi-task model for assessing the tone and objectivity of critiques of peer review comments. Scientometrics, 129(3):1377–1413, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04938-z.

Barrantes:2025:EFA

[Bar25] Barbara S. Lancho Barrantes. Exploratory factor analysis of bibliometric indicators for the Sustainable Development Goals. Scientometrics, 130(3):1701–1729, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05256-8.

Batagelj:2020:FAA

[Bat20] Vladimir Batagelj. On fractional approach to analysis of linked networks. Scientometrics, 123(2):621–633, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03383-y. See com-

ments [PM20].

Batagelj:2024:WTM

[Bat24a] Vladimir Batagelj. On weighted two-mode network projections. Scientometrics, 129(6):3565–3571, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05041-z.

Batagelj:2024:WDT

[Bat24b] Vladimir Batagelj. Weighted degrees and truncated derived bibliographic networks. Scientometrics, 129(8):4863– 4883, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05092-2.

Brezis:2020:APR

[BB20] Elise S. Brezis and Aliaksandr Birukou. Arbitrariness in the peer review process. Scientometrics, 123(1):393–411, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03348-1;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03348-1.pdf.

Barrera-Barrera:2022:SAL

[BB22] Ramón Barrera-Barrera. Selecting the appropriate leading journal in hospitality and tourism research: a guide based on the topic-journal fit and the JCR impact factor. Scientometrics, 127(4):1801–1823, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04288-8.

Bertoli-Barsotti:2023:EGC

[BB23] Lucio Bertoli-Barsotti. Equivalent Gini coefficient, not shape parameter! Scientometrics, 128(1):867–870, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04571-8.

Baccini:2024:FGC

[BBB^{+}24] Alberto Baccini, Federica Baccini, Lucio Barabesi, Martina Cioni, Eugenio Petrovich, and Daria Pignalosa. Finegrained classification of journal articles based on multiple layers of information through similarity network fusion: The case of the Cambridge Journal of Economics. Scientometrics, 129(1):373–400, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04884-2.

Basson:2021:DOA

[BBP21] Isabel Basson, Jaco P. Blanckenberg, and Heidi Prozesky. Do open access journal articles experience a citation advantage? Results and methodological reflections of an application of multiple measures to an analysis by WoS subject areas. Scientometrics, 126(1):459–484, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03734-9;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03734-9.pdf.

Bacci:2023:ICA

[BBP23] Silvia Bacci, Bruno Bertaccini, and Alessandra Petrucci. Insights from the co-authorship network of the Italian academic statisticians. Scientometrics, 128(8):4269–4303, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04761-y.

Ballandonne:2021:NRP

[BC21] Matthieu Ballandonne and Igor Cersosimo. A note on reference publication year spectroscopy with incomplete information. Scientometrics, 126(6):4927–4939, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03976-1.

Bonaccorsi:2021:IWM

[BCF21] Andrea Bonaccorsi, Filippo Chiarello, and Gualtiero Fantoni. Impact for whom? Mapping the users of public research with lexicon-based text mining. Scientometrics, 126(2):1745–1774, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03803-z;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03803-z.pdf.

Buckle:2022:SCD

[BCG22] Robert A. Buckle, John Creedy, and Norman Gemmell. Sources of convergence and divergence in university research quality: evidence from the performance-based research funding system in New Zealand. Scientometrics, 127 (6):3021–3047, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04376-9.

Bruni:2020:SHE

[BCM20] Renato Bruni, Giuseppe Catalano, and Henk F. Moed. Studying the heterogeneity of European higher education institutions. Scientometrics, 125(2):1117–1144, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03717-w.

Bickley:2022:AIF

[BCT22] Steve J. Bickley, Ho Fai Chan, and Benno Torgler. Artificial intelligence in the field of economics. Scientometrics, 127(4):2055–2084, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04294-w.

Barnett:2021:DAC

[BD21] Adrian Barnett and Zoë Doubleday. Demonstrating the ascendancy of COVID-19 research using acronyms. Scientometrics, 126(7):6127–6130, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04016-8.

Bai:2022:MIH

[BD22] Yongmei Bai and Jian Du. Measuring the impact of health research data in terms of data citations by scientific publications. Scientometrics, 127(12):6881–6893, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04559-4.

Borba:2021:SAJ

[BdCM21] José Alonso Borba, Alessanderson Jacó de Carvalho, and Fábio Minatto. Scopes of accounting journals and published papers: what do they signalize? Scientometrics, 126 (7):5665–5685, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03992-1.

Benjamens:2021:AVH

[BdMPH21] Stan Benjamens, Vincent E. de Meijer, Robert A. Pol, and Martijn P. D. Haring. Are all voices heard in the COVID-19 debate? Scientometrics, 126(1):859–862, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03730-z;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03730-z.pdf.

Barud:2021:LIT

[BdOdS21] Naia Augusto Barud, Renata Araujo de Oliveira, and Marcos dos Santos. Lean in information technology departments or companies: identifying publications on the Scopus and Web of Science databases. Scientometrics, 126(3): 2437–2457, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03662-8.

Biazoli:2025:IMB

[BdPFdO25] Leonardo Biazoli, Bruna de Paula Fonseca, and Izabela Regina Cardoso de Oliveira. International mobility boosts scientific careers: a synthetic control analysis of Brazilian researchers. Scientometrics, 130(5):3013–3028, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05327-w.

Bologna:2023:DOC

[BDPP23] Federica Bologna, Angelo Di Iorio, Silvio Peroni, and Francesco Poggi. Do open citations give insights on the qualitative peer-review evaluation in research assessments? An analysis of the Italian National Scientific Qualification. Scientometrics, 128(1):19–53, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04581-6.

Bornmann:2020:DPP

[BDTC20] Lutz Bornmann, Sitaram Devarakonda, Alexander Tekles, and George Chacko. Disruptive papers published in Scientometrics: meaningful results by using an improved variant of the disruption index originally proposed by Wu, Wang, and Evans (2019). Scientometrics, 123(2):1149–1155, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03406-8;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03406-8.pdf.

Benjafield:2020:VSA

[Ben20] John G. Benjafield. Vocabulary sharing among subjects belonging to the hierarchy of sciences. Scientometrics, 125(3):1965–1982, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03671-7;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03671-7.pdf.

Benga:2022:RPF

[Ben22] Gheorghe Benga. Remembrance of Petre T. Frangopol (1933–2020) the promoter of scientometrics in Romania. Scientometrics, 127(2):687–701, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04213-5.

Benjafield:2024:ERA

[Ben24] John G. Benjafield. Exploring relationships among eminent psychologists using co-occurrence analysis. Scientometrics, 129(3):1787–1799, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04930-7.

Benjafield:2025:UEW

[Ben25] John G. Benjafield. The use of emotion words by the sciences and other subjects. Scientometrics, 130(1):21–42, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05215-9.

Blanco-Encomienda:2021:QEP

[BERD21] Francisco Javier Blanco-Encomienda and Elena Rosillo-Díaz. Quantitative evaluation of the production and trends in research applying the structural equation modelling method. Scientometrics, 126(2):1599–1617, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03794-x.

Blind:2022:SRP

[BF22] Knut Blind and Alex Fenton. Standard-relevant publications: evidence, processes and influencing factors. Scientometrics, 127(1):577–602, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04210-8.

Bornmann:2025:PPC

[BF25] Lutz Bornmann and Russell J. Funk. Popper’s probability calculus and the decline of scientific disruptiveness. Scientometrics, 130(8):4773–4799, August 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05404-0.

Boetto:2021:UAD

[BFG^{+}21] Erik Boetto, Maria Pia Fantini, Aldo Gangemi, Davide Golinelli, Manfredi Greco, Andrea Giovanni Nuzzolese, Valentina Presutti, and Flavia Rallo. Using altmetrics for detecting impactful research in quasi-zero-day time-windows: the case of COVID-19. Scientometrics, 126(2):1189–1215, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03809-7;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03809-7.pdf.

Blumel:2021:ISI

[BG21] Clemens Blümel and Stephan Gauch. Introduction to special issue: quantitative studies of science in Germany. Scientometrics, 126(12):9641–9647, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04203-7.

Bordons:2023:IOU

[BGAMS23] María Bordons, Borja González-Albo, and Luz Moreno-Solano. Improving our understanding of open access: how it relates to funding, internationality of research and scientific leadership. Scientometrics, 128(8):4651–4676, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04726-1.

Banerjee:2025:ECD

[BGB25] Suchismita Banerjee, Abhik Ghosh, and Banasri Basu. Exploring citation diversity in scholarly literature: an entropy-based approach. Scientometrics, 130(5):2639– 2671, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05313-2.

Belli:2020:SRI

[BGP20] Simone Belli and Carlos Gonzalo-Penela. Science, research, and innovation infospheres in Google results of the Ibero–American countries. Scientometrics, 123(2): 635–653, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03399-4.

Braun:2020:CJ

[BGS20] Tibor Braun, Wolfgang Glänzel, and András Schubert. Commemorating Judit. Scientometrics, 123(3):1175–1179, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03442-4;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03442-4.pdf.

Bukowski:2020:FAB

[BGSRF20] Mark Bukowski, Sandra Geisler, Thomas Schmitz-Rode, and Robert Farkas. Feasibility of activity-based expert profiling using text mining of scientific publications and patents. Scientometrics, 123(2):579–620, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03414-8.

Budzinski:2020:DAP

[BGWZ20] Oliver Budzinski, Thomas Grebel, Jens Wolling, and Xijie Zhang. Drivers of article processing charges in open access. Scientometrics, 124(3):2185–2206, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03578-3;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03578-3.pdf.

Brambila:2024:AIC

[BH24] Claudia N. Gonzalez Brambila and Renata Herrerias. Assessing the impact of collaborative authorship in Business Economics in Latin America. Scientometrics, 129(7): 4623–4660, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05072-6.

Bornmann:2025:MSI

[BH25] Lutz Bornmann and Rouven Lazlo Haegner. Metrics sonification: The introduction of new ways to present bibliometric data using publication data of Loet Leydesdorff as an example. Scientometrics, 130(6):??, June 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05320-3.

Bhatt:2021:MPA

[Bha21] Bhumika Bhatt. A multi-perspective analysis of retractions in life sciences. Scientometrics, 126(5):4039–4054, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03907-0.

Bi:2023:CFPa

[Bi23a] Henry H. Bi. Correction to: Four problems of the h-index for assessing the research productivity and impact of individual authors. Scientometrics, 128(5):2693–2699, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04455-x. See [Bi23c].

Bi:2023:CFPb

[Bi23b] Henry H. Bi. Correction to: Four problems of the hindex for assessing the research productivity and impact of individual authors. Scientometrics, 128(5):2701, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04665-x. See [Bi23c].

Bi:2023:FPE

[Bi23c] Henry H. Bi. Four problems of the h-index for assessing the research productivity and impact of individual authors. Scientometrics, 128(5):2677–2691, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04323-8. See corrections

[Bi23a, ?, Bi23b].

Burns:2024:CAE

[BI24] C. Sean Burns and Md. Anwarul Islam. A citation analysis examining geographical specificity in article titles. Scientometrics, 129(7):4317–4328, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05075-3.

Birhan:2021:EMU

[Bir21] Amare Tesfie Birhan. An exploration of metadiscourse usage in book review articles across three academic disciplines: a contrastive analysis of corpus-based research approach. Scientometrics, 126(4):2885–2902, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03822-w.

Bertoletti:2021:EUI

[BJ21] Alice Bertoletti and Geraint Johnes. Efficiency in university-industry collaboration: an analysis of UK higher education institutions. Scientometrics, 126(9): 7679–7714, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04076-w.

Bjork:2020:CRC

[Bjø20a] R. Bjørk. Correction to: Response to the comments of Turki et al. on “The journals that publish Nobel Prize research”. Scientometrics, 124(1):795, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03565-8;   http:
//link.springer.com/content/pdf/10.1007/s11192-

020-03565-8.pdf. See [Bjø20a, Bjø20c, TTB20].

Bjork:2020:JPP

[Bjø20b] R. Bjørk. The journals in physics that publish Nobel Prize research. Scientometrics, 122(2):817–823, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03312-8. See comments [TTB20]

and response [Bjø20c, Bjø20a].

Bjork:2020:RCT

[Bjø20c] R. Bjørk. Response to the comments of Turki et al. on “The journals that publish Nobel Prize research”. Scientometrics, 124(1):791–793, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

020-03459-9. See [Bjø20b, TTB20, Bjø20a].

Beranova:2022:WWC

[BJS22] Lucie Beranová, Marcin P. Joachimiak, and Vilém Sklenák. Why was this cited? Explainable machine learning applied to COVID-19 research literature. Scientometrics, 127(5):2313–2349, May 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04314-9.

Buenstorf:2025:KMP

[BKO25] Guido Buenstorf, Johannes Koenig, and Anne Otto. Keeping up with the Max Plancks? Germany’s quest for university excellence and the role of public research institutes in doctoral education. Scientometrics, 130(1):67–108, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05195-w.

Bhattacharya:2020:CSA

[BKS20] Sujit Bhattacharya, Ravinder Kumar, and Shubham Singh. Capturing the salient aspects of IoT research: A social network analysis. Scientometrics, 125(1):361– 384, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03620-4.

Bickley:2020:CIG

[BKT20] Matthew S. Bickley, Kayvan Kousha, and Michael Thelwall. Can the impact of grey literature be assessed? An investigation of UK government publications cited by articles and books. Scientometrics, 125(2):1425–1444, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03628-w.

Bickley:2022:SMI

[BKT22] Matthew S. Bickley, Kayvan Kousha, and Michael Thelwall. A systematic method for identifying references to academic research in grey literature. Scientometrics, 127(12):6913–6933, December 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04408-4.

Bornmann:2020:HRJ

[BL20a] Lutz Bornmann and Loet Leydesdorff. Historical roots of Judit Bar-Ilan’s research: a cited-references analysis using CRExplorer. Scientometrics, 123(3):1193–1200, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03438-0;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03438-0.pdf.

Boufarss:2020:OSO

[BL20b] Mohamed Boufarss and Mikael Laakso. Open Sesame? Open access priorities, incentives, and policies among higher education institutions in the United Arab Emirates. Scientometrics, 124(2):1553–1577, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03529-y;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03529-y.pdf.

Bornmann:2024:UCF

[BL24] Lutz Bornmann and Benedetto Lepori. The use of ChatGPT to find similar institutions for institutional benchmarking. Scientometrics, 129(6):3593–3598, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05039-7.

Bai:2021:VRT

[BLL21] Yang Bai, Hongxiu Li, and Yong Liu. Visualizing research trends and research theme evolution in e-learning field: 1999–2018. Scientometrics, 126(2):1389–1414, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03760-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03760-7.pdf.

Bornmann:2020:TTR

[BM20a] Lutz Bornmann and Werner Marx. Thomas theorem in research evaluation. Scientometrics, 123(1):553–555, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03389-6;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03389-6.pdf.

Breimer:2020:HCM

[BM20b] Lars H. Breimer and Dimitri P. Mikhailidis. Half a century and more of PhD theses by published papers. Scientometrics, 125(1):813–816, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03622-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03622-2.pdf.

Boothby:2021:EFT

[BM21] Clara Boothby and Stasa Milojević. An exploratory fulltext analysis of science careers in a changing academic job market. Scientometrics, 126(5):4055–4071, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03905-2.

Bornmann:2024:OSS

[BM24] Lutz Bornmann and Julian N. Marewski. Opium in science and society: numbers and other quantifications. Scientometrics, 129(9):5313–5346, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05104-1.

Bilalli:2021:FAP

[BMA21] Besim Bilalli, Rana Faisal Munir, and Alberto Abelló. A framework for assessing the peer review duration of journals: case study in computer science. Scientometrics, 126(1):545–563, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03742-9.

Belli:2020:CMS

[BMBA20] Simone Belli, Rogério Mugnaini, Joan Baltà, and Ernest Abadal. Coronavirus mapping in scientific publications: When science advances rapidly and collectively, is access to this knowledge open to society? Scientometrics, 124(3):2661–2685, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03590-7;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03590-7.pdf.

Bonaccorsi:2022:EAI

[BMM22] Andrea Bonaccorsi, Nicola Melluso, and Francesco Alessandro Massucci. Exploring the antecedents of interdisciplinarity at the European Research Council: a topic modeling approach. Scientometrics, 127(12):6961–6991, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04368-9.

Baji:2021:PAC

[BMS21a] Fatima Baji, Ismail Mostafavi, and Zivar Sabaghinejad. Partnership ability and co-authorship network of information literacy field. Scientometrics, 126(9):8205–8216, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04062-2.

Bornmann:2021:MIP

[BMS21b] Lutz Bornmann, Rüdiger Mutz, and Moritz Stefaner. Mapping the impact of papers on various status groups in excellencemapping.net: a new release of the excellence mapping tool based on citation and reader scores. Scientometrics, 126(11):9305–9331, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04141-4.

Borner:2024:ICS

[BMW24] Katy Börner, Stasa Milojević, and Caroline Wagner. 19th International Conference on Scientometrics & Informetrics (ISSI) 2023 editorial. Scientometrics, 129(10):5815–5819, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05185-y.

Boukhari:2020:AMB

[BO20] Kabil Boukhari and Mohamed Nazih Omri. Approximate matching-based unsupervised document indexing approach: application to biomedical domain. Scientometrics, 124(2):903–924, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03474-w.

Blazquez-Ochando:2025:PEB

[BOPGOP25] Manuel Blázquez-Ochando, Juan José Prieto-Gutiérrez, and María Antonia Ovalle-Perandones. Prompt engineering for bibliographic web-scraping. Scientometrics, 130 (7):3405–3432, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05372-5.

Bordignon:2020:SCS

[Bor20a] Frederique Bordignon. Self-correction of science: a comparative study of negative citations and post-publication peer review. Scientometrics, 124(2):1225–1239, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03536-z.

Bornmann:2020:BBD

[Bor20b]  Lutz Bornmann.  Bibliometrics-based decision tree

(BBDT) for deciding whether two universities in the Leiden ranking differ substantially in their performance. Scientometrics, 122(2):1255–1258, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03319-1.

Bordignon:2021:SRP

[Bor21] Frederique Bordignon. A scientometric review of permafrost research based on textual analysis (1948-2020). Scientometrics, 126(1):417–436, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03747-4.

Bordignon:2022:CCK

[Bor22] Frederique Bordignon. Critical citations in knowledge construction and citation analysis: from paradox to definition. Scientometrics, 127(2):959–972, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04226-0.

Bornmann:2024:SDS

[Bor24] Lutz Bornmann. Skewed distributions of scientists’ productivity: a research program for the empirical analysis. Scientometrics, 129(4):2455–2468, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04962-z.

Barragan-Ocana:2020:AIA

[BORROPGV20] Alejandro Barragán-Ocaña, Gerardo Reyes-Ruiz, Samuel Olmos-Peña, and Hortensia Gómez-Viquez. Approach to the identification of an alternative technological innovation index. Scientometrics, 122(1):23–45, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03292-9.

Boschen:2021:SRJ

[Bös21] Ingmar Böschen. Software review: The JATSdecoder package — extract metadata, abstract and sectioned text from NISO-JATS coded XML documents; insights to PubMed central’s open access database. Scientometrics, 126(12):9585–9601, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04162-z.

Boufarss:2020:COA

[Bou20] Mohamed Boufarss. Charting the Open Access scholarly journals landscape in the UAE. Scientometrics, 122 (3):1707–1725, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03349-0;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03349-0.pdf.

Boudry:2021:AOP

[Bou21] Christophe Boudry. Availability of ORCIDs in publications archived in PubMed, MEDLINE, and Web of Science Core Collection. Scientometrics, 126(4):3355–3371, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03825-7.

Brambila:2021:PEP

[BOV21] Claudia N. González Brambila and José Luis Olivares-Vázquez. Patterns and evolution of publication and coauthorship in social sciences in Mexico. Scientometrics, 126(3):2595–2626, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-020-
03644-w.

Bambo:2020:BAB

[BP20] Thabang Lazarus Bambo and Anastassios Pouris. Bibliometric analysis of bioeconomy research in South Africa. Scientometrics, 125(1):29–51, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03626-y.

Baccini:2022:NVS

[BP22] Alberto Baccini and Eugenio Petrovich. Normative versus strategic accounts of acknowledgment data: The case of the top-five journals of economics. Scientometrics, 127(1): 603–635, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04185-6.

Barnett:2023:CAA

[BP23] George A. Barnett and Han Woo Park. Co-authorship among the fellows of the International Communication Association. Scientometrics, 128(6):3401–3418, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04705-6.

Benjamens:2021:PRD

[BPH21] Stan Benjamens, Robert A. Pol, and Martijn P. D. Haring. Peer review during demanding times: maintain rigorous standards. Scientometrics, 126(7):6115–6117, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04007-9.

Benz:2025:MUP

[BPL25] Pierre Benz, Carolina Pradier, and Vincent Larivière. Mapping the unseen in practice: comparing latent Dirichlet allocation and BERTopic for navigating topic spaces. Scientometrics, 130(7):3813–3838, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05339-6.

Bautista-Puig:2020:DJF

[BPLIdMA^{+}20] Nuria Bautista-Puig, Carmen Lopez-Illescas, Felix de Moya-Anegon, Vicente Guerrero-Bote, and Henk F. Moed. Do journals flipping to gold open access show an OA citation or publication advantage? Scientometrics, 124(3):2551– 2575, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03546-x.

Bautista-Puig:2021:RTG

[BPMRSL21] Núria Bautista-Puig, Jorge Mañana-Rodríguez, and Antonio Eleazar Serrano-López. Role taxonomy of green and sustainable science and technology journals: exportation, importation, specialization and interdisciplinarity. Scientometrics, 126(5):3871–3892, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03939-6.

Byrne:2021:TRR

[BPW^{+}21] Jennifer A. Byrne, Yasunori Park, Rachael A. West, Amanda Capes-Davis, Bertrand Favier, Guillaume Cabanac, and Cyril Labbé. The thin ret(raction) line: biomedical journal responses to incorrect non-targeting nucleotide sequence reagents in human gene knockdown publications. Scientometrics, 126(4):3513–3534, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03871-9;   http:
//link.springer.com/content/pdf/10.1007/s11192-
021-03871-9.pdf.

Baccini:2025:PFE

[BR25] Alberto Baccini and Cristina Re. Is the panel fair? Evaluating panel compositions through network analysis. The case of research assessments in Italy. Scientometrics, 130 (4):2093–2135, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05255-9.

Breitzman:2021:RBW

[Bre21] Anthony Breitzman. The relationship between web usage and citation statistics for electronics and information technology articles. Scientometrics, 126(3):2085–2105, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03851-5.

Bonnevie:2023:VAA

[BRG^{+}23] Tristan Bonnevie, Aurore Repel, Francis-Edouard Gravier, Joel Ladner, Louis Sibert, Jean-François Muir, Antoine Cuvelier, and Marc-Olivier Fischer. Video abstracts are associated with an increase in research reports citations, views and social attention: a cross-sectional study. Scientometrics, 128(5):3001–3015, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04675-9.

Budimir:2021:CSC

[BRJ21] Gordana Budimir, Sophia Rahimeh, and Primoz Juznic. Comparison of self-citation patterns in WoS and Scopus databases based on national scientific production in Slovenia (1996–2020). Scientometrics, 126(3):2249–2267, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03862-w.

Blumel:2020:SRA

[BS20] Clemens Blümel and Alexander Schniedermann. Studying review articles in scientometrics and beyond: a research agenda. Scientometrics, 124(1):711–728, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03431-7;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03431-7.pdf.

Bennett:2023:GAA

[BS23] Hunter Bennett and Flynn Slattery. Graphical abstracts are associated with greater Altmetric attention scores, but not citations, in sport science. Scientometrics, 128(6): 3793–3804, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04709-2.

Brito:2023:AIP

[BSA23] Ana C. M. Brito, Filipi N. Silva, and Diego R. Amancio. Analyzing the influence of prolific collaborations on authors productivity and visibility. Scientometrics, 128(4): 2471–2487, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04669-7.

Bin:2022:HMD

[BSFS^{+}22] Adriana Bin, Sergio Salles-Filho, Ana Carolina Spatti, Jesús Pascual Mena-Chalco, and Fernando Antonio Basile Colugnati. How much does a Ph.D. scholarship program impact an emerging economy research performance? Scientometrics, 127(12):6935–6960, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04487-3.

Buhrer:2020:EGE

[BSPR20] Susanne Bührer, Evanthia Kalpazidou Schmidt, Rachel Palmén, and Sybille Reidl. Evaluating gender equality effects in research and innovation systems. Scientometrics, 125(2):1459–1475, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03596-1;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03596-1.pdf.

Bennett:2024:JSC

[BSS24] Hunter Bennett, Ben Singh, and Flynn Slattery. Journal self-citations trends in sport sciences: an analysis of disciplinary journals from 2013 to 2022. Scientometrics, 129(8): 4741–4750, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05096-y.

Breuer:2022:RAB

[BST22] Timo Breuer, Philipp Schaer, and Dirk Tunger. Relevance assessments, bibliometrics, and altmetrics: a quantitative study on PubMed and arXiv. Scientometrics, 127 (5):2455–2478, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04319-4.

Belkadhi:2020:TSR

[BT20] Khaled Belkadhi and Adel Trabelsi. Toward a stochastically robust normalized impact factor against fraud and scams. Scientometrics, 124(3):1871–1884, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03577-4.

Basuki:2022:SSA

[BT22] Setio Basuki and Masatoshi Tsuchiya. SDCF: semiautomatically structured dataset of citation functions. Scientometrics, 127(8):4569–4608, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04471-x.

Bi:2025:DFG

[BT25] Weilong Bi and Benno Torgler. Do first-generation immigrant scholars outperform native researchers? Evidence from US business schools. Scientometrics, 130(7): 3839–3870, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05374-3.

Bihari:2021:IWE

[BTD21] Anand Bihari, Sudhakar Tripathi, and Akshay Deepak. Iterative weighted EM and iterative weighted EM’-index for scientific assessment of scholars. Scientometrics, 126 (7):5551–5568, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03937-8.

Biesenbender:2024:LSE

[BTP24] Kristin Biesenbender, Ralf Toepfer, and Isabella Peters. Life scientists’ experience with posting preprints during the COVID-19 pandemic. Scientometrics, 129(10):6407– 6434, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04982-9.

Borsi:2025:KSR

[BVS25] Balázs Borsi, Zsófia Vida, and Sándor Soós. Keyword standardization and restructuring: the impact on analysing network-based science maps in innovation management research. Scientometrics, 130(2):593–617, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05232-2.

Bascur:2023:AIR

[BVvEW23] Juan Pablo Bascur, Suzan Verberne, Nees Jan van Eck, and Ludo Waltman. Academic information retrieval using citation clusters: in-depth evaluation based on systematic reviews. Scientometrics, 128(5):2895–2921, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04681-x.

Bascur:2025:WTB

[BVvEW25] Juan Pablo Bascur, Suzan Verberne, Nees Jan van Eck, and Ludo Waltman. Which topics are best represented by science maps? An analysis of clustering effectiveness for citation and text similarity networks. Scientometrics, 130(2):1181–1199, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05218-6.

Bornmann:2019:NCI

[BW19] Lutz Bornmann and Klaus Wohlrabe. Normalisation of citation impact in economics. Scientometrics, 120(2):841–884, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03140-w;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03140-w.pdf. See correction

[BW20c].

Baumann:2020:WAW

[BW20a] Alexandra Baumann and Klaus Wohlrabe. Where have all the working papers gone? Evidence from four major economics working paper series. Scientometrics, 124 (3):2433–2441, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03570-x;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03570-x.pdf.

Bornmann:2020:EPM

[BW20b] Lutz Bornmann and Richard Williams. An evaluation of percentile measures of citation impact, and a proposal for making them better. Scientometrics, 124(2): 1457–1478, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03512-7;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03512-7.pdf.

Bornmann:2020:CNC

[BW20c] Lutz Bornmann and Klaus Wohlrabe. Correction to: Normalisation of citation impact in economics. Scientometrics, 123(2):1167, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03255-0;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03255-0.pdf. See [BW19].

Burgi:2022:ICP

[BW22] Constantin Bürgi and Klaus Wohlrabe. The influence of Covid-19 on publications in economics: bibliometric evidence from five working paper series. Scientometrics, 127(9):5175–5189, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04473-9.

Bornmann:2024:UCI

[BWE24] Lutz Bornmann, Lingfei Wu, and Christoph Ettl. The use of ChatGPT for identifying disruptive papers in science: a first exploration. Scientometrics, 129(11):7161– 7165, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05176-z.

Bornmann:2020:CCA

[BWH20a] Lutz Bornmann, K. Brad Wray, and Robin Haunschild. Citation concept analysis (CCA): a new form of citation analysis revealing the usefulness of concepts for other researchers illustrated by exemplary case studies including classic books by Thomas S. Kuhn and Karl R. Popper. Scientometrics, 122(2):1051–1074, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03326-2;   http:
//link.springer.com/content/pdf/10.1007/s11192-

019-03326-2.pdf. See correction [BWH20b].

Bornmann:2020:CCC

[BWH20b] Lutz Bornmann, K. Brad Wray, and Robin Haunschild. Correction to: Citation concept analysis (CCA): a new form of citation analysis revealing the usefulness of concepts for other researchers illustrated by exemplary case studies including classic books by Thomas S. Kuhn and Karl R. Popper. Scientometrics, 124(3):2737, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03495-5;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03495-5.pdf. See [BWH20a].

Bai:2020:RIG

[BWLX20] Xu Bai, Jinxi Wu, Yun Liu, and Yihan Xu. Research on the impact of global innovation network on 3D printing industry performance. Scientometrics, 124(2):1015– 1051, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03534-1.

Budi:2023:UMC

[BY23] Indra Budi and Yaniasih Yaniasih. Understanding the meanings of citations using sentiment, role, and citation function classifications. Scientometrics, 128(1):735–759, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04567-4.

Blidstein:2022:TNG

[BZG22] Moshe Blidstein and Maayan Zhitomirsky-Geffet. Towards a new generic framework for citation network generation and analysis in the humanities. Scientometrics, 127(7): 4275–4297, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04438-y.

Bai:2021:QSC

[BZL21] Xiaomei Bai, Fuli Zhang, and Ivan Lee. Quantifying scientific collaboration impact by exploiting collaborationcitation network. Scientometrics, 126(9):7993–8008, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04078-8.

Bedru:2023:CCS

[BZX^{+}23] Hayat D. Bedru, Chen Zhang, Feng Xie, Shuo Yu, and Iftikhar Hussain. CLARA: citation and similarity-based author ranking. Scientometrics, 128(2):1091–1117, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04590-5.

Bao:2025:ELS

[BZZ25] Tong Bao, Yi Zhao, and Chengzhi Zhang. Examining linguistic shifts in academic writing before and after the launch of ChatGPT: a study on preprint papers. Scientometrics, 130(7):3571–3595, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05341-y.

Collins:2022:RCP

[CA22] Annie Collins and Rohan Alexander. Reproducibility of COVID-19 pre-prints. Scientometrics, 127(8):4655–4673, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04418-2.

Candan:2020:EPA

[Can20] Gökçe Candan. Efficiency and performance analysis of economics research using hesitant fuzzy AHP and OCRA methods. Scientometrics, 124(3):2645–2659, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03584-5.

Cantone:2024:HMI

[Can24] Giulio Giacomo Cantone. How to measure interdisciplinary research? A systemic design for the model of measurement. Scientometrics, 129(8):4937–4982, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05085-1.

Cao:2020:CBI

[Cao20] Qinwei Cao. Contradiction between input and output of Chinese scientific research: a multidimensional analysis. Scientometrics, 123(1):451–485, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03377-w.

Carusi:2020:LIU

[CB20] Chiara Carusi and Giuseppe Bianchi. A look at interdisciplinarity using bipartite scholar/journal networks. Scientometrics, 122(2):867–894, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03309-3.

Chandrakala:2025:ASV

[CB25] B. R. Chandrakala and H. Rajendra Babu. Analyzing the scholarly visibility of blockchain technology research output on social media platforms using altmetrics. Scientometrics, 130(7):3679–3706, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05354-7.

Cowhitt:2024:EEA

[CBA24] Thomas Cowhitt, Joshua Travis Brown, and Anthony Lising Antonio. The emergence and evolution of ambiguous ideas: an innovative application of social network analysis to support systematic literature reviews. Scientometrics, 129(11):7005–7033, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05144-7.

Cugnata:2024:MPR

[CBP^{+}24] Federica Cugnata, Chiara Brombin, Chiara Maria Poli, Roberto Buccione, and Clelia Di Serio. Modelling perception and resilience factors to data sharing in clinical and basic research: an observational study. Scientometrics, 129 (6):3169–3192, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05015-1.

Chen:2021:MEU

[CC21] Shih-Pin Chen and Chung-Wei Chang. Measuring the efficiency of university departments: an empirical study using data envelopment analysis and cluster analysis. Scientometrics, 126(6):5263–5284, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03982-3.

Chi:2022:MIR

[CC22] Pei-Shan Chi and Stijn Conix. Measuring the isolation of research topics in philosophy. Scientometrics, 127(4): 1669–1696, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04276-y.

Cioffi:2022:ICI

[CCA22] Alessia Cioffi, Sara Coppini, and Nooshin Shahidzadeh Asadi. Identifying and correcting invalid citations due to DOI errors in Crossref data. Scientometrics, 127(6): 3593–3612, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04367-w.

Chow:2022:SAD

[CCC22] Julie Chi Chow, Tsair-Wei Chien, and Willy Chou. Suggestions to the article: demonstrating the ascendancy of COVID-19 research using acronyms. Scientometrics, 127 (5):2897–2899, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04302-z.

Cabrera:2020:MCI

[CCdC20] Lilian Cervo Cabrera, Carlos Eduardo Caldarelli, and Marcia Regina Gabardo da Camara. Mapping collaboration in international coffee certification research. Scientometrics, 124(3):2597–2618, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03549-8.

Chang:2020:DTR

[CCH20] Yu-Wei Chang, Dar-Zen Chen, and Mu-Hsuan Huang. Discovering types of research performance of scientists with significant contributions. Scientometrics, 124(2): 1529–1552, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03533-2.

Cao:2024:DTR

[CCH^{+}24] Xiaoli Cao, Xiang Chen, Lu Huang, Lijie Deng, Yijie Cai, and Hang Ren. Detecting technological recombination using semantic analysis and dynamic network analysis. Scientometrics, 129(11):7385–7416, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04812-4.

Chen:2025:HMD

[CCH25] Guo Chen, Shuya Chen, and Jiming Hu. How much data is sufficient for reliable bibliometric domain analysis? A multi-scenario experimental approach. Scientometrics, 130 (5):2901–2921, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05335-w.

Choji:2024:SIL

[CCMM24] Thamyres T. Choji, Manuel J. Cobo, and Jose A. Moral-Munoz. Is the scientific impact of the LIS themes genderbiased? A bibliometric analysis of the evolution, scientific impact, and relative contribution by gender from 2007 to 2022. Scientometrics, 129(10):6023–6047, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05005-3.

Chen:2023:ANR

[CCSX23] Guo Chen, Jing Chen, Yu Shao, and Lu Xiao. Automatic noise reduction of domain-specific bibliographic datasets using positive-unlabeled learning. Scientometrics, 128(2):

1187–1204, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04598-x.

Chen:2020:EWE

[CCZ^{+}20] Jie Chen, Jialin Chen, Shu Zhao, Yanping Zhang, and Jie Tang. Exploiting word embedding for heterogeneous topic model towards patent recommendation. Scientometrics, 125(3):2091–2108, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03666-4.

Chao:2023:JFI

[CCZL23] Wenhan Chao, Mengyuan Chen, Xian Zhou, and Zhunchen Luo. A joint framework for identifying the type and arguments of scientific contribution. Scientometrics, 128(6): 3347–3376, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04694-6.

Curiac:2022:CIT

[CD22] Christian-Daniel Curiac and Alex Doboli. Combining informetrics and trend analysis to understand past and current directions in electronic design automation. Scientometrics, 127(10):5661–5689, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04481-9.

Ceribeli:2021:HCC

[CdAdFC21] Caroline Ceribeli, Henrique Ferraz de Arruda, and Luciano da Fontoura Costa. How coupled are capillary electrophoresis and mass spectrometry? Scientometrics, 126 (5):3841–3851, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03923-0.

Chaignon:2023:SSE

[CDE23] Lauranne Chaignon, Domingo Docampo, and Daniel Egret. In search of a scientific elite: highly cited researchers (HCR) in France. Scientometrics, 128(10):5801–5827, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04805-3.

Colladon:2020:PFS

[CDG20] Andrea Fronzetti Colladon, Ciriaco Andrea D’Angelo, and Peter A. Gloor. Predicting the future success of scientific publications through social network and semantic analysis. Scientometrics, 124(1):357–377, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03479-5.

Chen:2025:ESC

[CDQ25] Liyue Chen, Jielan Ding, and Zihao Qu. Exploring scientific contributions through citation context and division of labor. Scientometrics, 130(5):2875–2899, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05318-x.

Corsi:2021:TTO

[CdSK21] Alana Corsi, Fabiane Florencio de Souza, and João Luiz Kovaleski. Technology transfer oriented to sustainable development: proposal of a theoretical model based on barriers and opportunities. Scientometrics, 126(6):5081–5112, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03969-0.

Chiarini:2025:WLM

[CdSS25] Tulio Chiarini, Emerson Gomes dos Santos, and Leandro Alves Silva. Who leads matters: diversity and external collaboration in Brazilian scientific teams. Scientometrics, 130(8):4723–4748, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05387-y.

Chen:2020:MDC

[CDZ^{+}20] Yilong Chen, Yiting Dong, Yu Zeng, Xiaoyan Yang, Jiantong Shen, Lang Zheng, Jingwen Jiang, Liming Pu, and Qilin Bao. Mapping of diseases from clinical medicine research — a visualization study. Scientometrics, 125(1):171–185, October 2020. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03646-8;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03646-8.pdf.

Chen:2020:ELC

[CDZZ20] Bikun Chen, Dannan Deng, Zhouyan Zhong, and Chengzhi Zhang. Exploring linguistic characteristics of highly browsed and downloaded academic articles. Scientometrics, 122(3):1769–1790, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03361-4.

Cernat:2024:RSM

[Cer24a] Vasile Cernat. Is Romania’s surge in MDPI publications a success story? A response to Nazarovets (2024). Scientometrics, 129(12):7985–7988, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05187-w. See [Cer24b, Naz24a].

Cernat:2024:UPH

[Cer24b] Vasile Cernat. The unprincipled principal: how Romania’s inconsistent research reform impacted scientific output. Scientometrics, 129(9):5557–5575, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05118-9. See [Naz24a, Cer24a].

Caviggioli:2022:RSD

[CF22] Federico Caviggioli and Boris Forthmann. Reach for the stars: disentangling quantity and quality of inventors’ productivity in a multifaceted latent variable model. Scientometrics, 127(12):7015–7040, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04328-3.

Csomos:2023:UIM

[CF23] György Csomós and Jenő Zsolt Farkas. Understanding the increasing market share of the academic publisher “Multidisciplinary Digital Publishing Institute” in the publication output of Central and Eastern European countries: a case study of Hungary. Scientometrics, 128(1):803–824, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04586-1.

Carvalho:2023:LIC

[CFA^{+}23] Danilo Silva Carvalho, Lucas Lopes Felipe, Priscila Costa Albuquerque, Fabio Zicker, and Bruna de Paula Fonseca. Leadership and international collaboration on COVID-19 research: reducing the North–South divide? Scientometrics, 128(8):4689–4705, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04754-x.

Cong:2022:WUC

[CFC22] Ting Cong, Zhichao Fang, and Rodrigo Costas. WeChat uptake of Chinese scholarly journals: an analysis of CSSCIindexed journals. Scientometrics, 127(12):7091–7110, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04347-0.

Cinelli:2022:CMP

[CFI22] Matteo Cinelli, Giovanna Ferraro, and Antonio Iovanella. Connections matter: a proxy measure for evaluating network membership with an application to the Seventh Research Framework Programme. Scientometrics, 127(7): 3959–3976, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04414-6.

Cabanac:2020:SLM

[CFM20] Guillaume Cabanac, Ingo Frommholz, and Philipp Mayr. Scholarly literature mining with information retrieval and natural language processing: Preface. Scientometrics, 125(3):2835–2840, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03763-4;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03763-4.pdf.

Cai:2021:ICD

[CFW21] X. Cai, C. V. Fry, and C. S. Wagner. International collaboration during the COVID-19 crisis: autumn 2020 developments. Scientometrics, 126(4):3683–3692, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03873-7;   http://link.springer.
com/content/pdf/10.1007/s11192-021-03873-7.pdf.

Chi:2022:ABC

[CG22] Pei-Shan Chi and Wolfgang Glänzel. An article-based cross-disciplinary study of reference literature for indicator improvement. Scientometrics, 127(12):7077–7089, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04262-w.

Chi:2024:TSS

[CG24] Pei-Shan Chi and Wolfgang Glänzel. Two sides of the same coin? Citation obsolescence and impact of different publication types and subject fields. Scientometrics, 129(10): 6373–6386, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04926-9.

Ciampi:2021:RSD

[CGA21] Francesco Ciampi, Alessandro Giannozzi, and Edward I. Altman. Rethinking SME default prediction: a systematic literature review and future perspectives. Scientometrics, 126(3):2141–2188, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03856-0.

Chen:2024:IML

[CGDS24] Shiji Chen, Yanan Guo, Alvin Shijie Ding, and Yanhui Song. Is interdisciplinarity more likely to produce novel or disruptive research? Scientometrics, 129(5):2615–2632, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04981-w.

Cambe:2022:NCM

[CGJ22] Jordan Cambe, Sebastian Grauwin, and Pablo Jensen. A new clustering method to explore the dynamics of research communities. Scientometrics, 127(8):4459–4482, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04463-x.

Chen:2024:MLM

[CGJM24] Pei Chen, Shan Gao, Fan Jiang, and Yifang Ma. Measuring the labor market outcomes of universities: evidence from China’s listed company executives. Scientometrics, 129(9):5715–5730, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05138-5.

Castanha:2024:CCA

[CGPR24] Rafael Gutierres Castanha, Maria Claudia Cabrini Grácio, and Antonio Perianes-Rodríguez. Co-citation analysis between coupler authors of a scientific domain’s citation identity: a case study in scientometrics. Scientometrics, 129(3):1545–1566, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04927-8.

Cai:2020:ADL

[CH20] Xiaoyu Cai and Tao Han. Analysis of the division of labor in China’s high-quality life sciences research. Scientometrics, 125(2):1077–1094, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03582-7.

Chang:2021:CHR

[Cha21] Yu-Wei Chang. Characteristics of high research performance authors in the field of library and information science and those of their articles. Scientometrics, 126(4): 3373–3391, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-021-03898-y.

Chang:2023:CAC

[Cha23a] Yu-Wei Chang. Comparison of the application of curricula vitae and bibliometric analyses for tracing longterm and temporary scientific mobility. Scientometrics, 128(12):6509–6526, December 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-023-04853-9.

Chankseliani:2023:WFP

[Cha23b] Maia Chankseliani. Who funds the production of globally visible research in the Global South? Scientometrics, 128(1):783–801, January 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04583-4.

Chen:2018:UPC

[Che18] Bikun Chen. Usage pattern comparison of the same scholarly articles between Web of Science (WoS) and Springer. Scientometrics, 115(1):519–537, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2616-3. See correction [Che20].

Chen:2020:CUP

[Che20] Bikun Chen. Correction to: Usage pattern comparison of the same scholarly articles between Web of Science (WoS) and Springer. Scientometrics, 123(2):1173, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03283-w;   http:
//link.springer.com/content/pdf/10.1007/s11192-
019-03283-w.pdf. See [Che18].

Chen:2021:O

[Che21] Yue Chen. Obituary. Scientometrics, 126(7):6135–6136, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03983-2.

Chen:2023:DCF

[Che23] Xinyi Chen. Does cross-field influence regional and fieldspecific distributions of highly cited researchers? Scientometrics, 128(1):825–840, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04584-3.

Chechik:2024:GDR

[Che24] Elena Chechik. Gender disparities in research fields in Russia: dissertation authors and their mentors. Scientometrics, 129(6):3341–3358, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05018-y.

Huang:2024:REM

[cHhC24] Chien chih Huang and Kuang hua Chen. RefCit2vec: embedding models considering references and citations for measuring document similarity. Scientometrics, 129(8): 4669–4693, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05067-3.

Culbert:2025:RCA

[CHJ^{+}25] Jack H. Culbert, Anne Hobert, Najko Jahn, Nick Haupka, Marion Schmidt, Paul Donner, and Philipp Mayr. Reference coverage analysis of OpenAlex compared to Web of Science and Scopus. Scientometrics, 130(4):2475–2492, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05293-3.

Cho:2020:ISE

[Cho20] Jane Cho. Intellectual structure evolution of open access research observed through correlation index of keyword centrality. Scientometrics, 125(3):2617–2635, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03682-4.

Cho:2021:AAH

[Cho21] Jane Cho. Altmetrics analysis of highly cited academic papers in the field of library and information science. Scientometrics, 126(9):7623–7635, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04084-w.

Cho:2024:EEA

[Cho24] Jane Cho. The effect East Asian researcher’s academic performance on international journal review and editing activities. Scientometrics, 129(3):1825–1839, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04918-9.

Chan:2024:TRT

[CHT24] Ho Fai Chan, Ella Hugo, and Benno Torgler. Tracking a “radioactive tracer”: laziness in academia. Scientometrics, 129(1):431–443, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04908-x.

Chan:2022:RLE

[CJT22] Ho Fai Chan, Franklin G. Mixon Jr, and Benno Torgler. Recognition and longevity: an examination of award timing and lifespan in Nobel laureates. Scientometrics, 127 (6):3629–3659, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04379-6.

Chumachenko:2020:DTF

[CKMY20] A. V. Chumachenko, B. G. Kreminskyi, Iu. L. Mosenkis, and A. I. Yakimenko. Dynamics of topic formation and quantitative analysis of hot trends in physical science. Scientometrics, 125(1):739–753, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03610-6.

Chikazawa:2021:MAM

[CKO21] Yuto Chikazawa, Marie Katsurai, and Ikki Ohmukai. Multilingual author matching across different academic databases: a case study on KAKEN, DBLP, and PubMed. Scientometrics, 126(3):2311–2327, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03861-3.

Clermont:2021:DCP

[CKT21] Marcel Clermont, Johanna Krolak, and Dirk Tunger. Does the citation period have any effect on the informative value of selected citation indicators in research evaluations? Scientometrics, 126(2):1019–1047, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03782-1;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03782-1.pdf.

Choi:2020:PLD

[CL20] Mincheol Choi and Chang-Yang Lee. Power-law distributions of corporate innovative output: evidence from U.S. patent data. Scientometrics, 122(1):519–554, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03304-8.

Cavallaro:2021:IBP

[CL21a] Marco Cavallaro and Benedetto Lepori. Institutional barriers to participation in EU framework programs: contrasting the Swiss and UK cases. Scientometrics, 126(2):1311–1328, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03810-0;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03810-0.pdf.

Chen:2021:IUS

[CL21b] Jun Chen and Jia Liu. Incentive and uncertainty: the simultaneous effects of demand on innovation. Scientometrics, 126(9):7743–7757, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04093-9.

Chen:2021:EPZ

[CL21c] Yue Chen and Jean-Charles Lamirel. Editorial preface to the Zeyuan Liu memorial issue. Scientometrics, 126(7): 6133–6134, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03997-w.

Chou:2023:UTM

[CLC23] Po-Hsin Chou, Jui-Chung John Lin, and Tsair-Wei Chien. Using text mining and forest plots to identify similarities and differences between two spine-related journals based on medical subject headings (MeSH terms) and author-specified keywords in 100 top-cited articles. Scientometrics, 128(1):1–17, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04549-6.

Chen:2024:CDD

[CLC^{+}24] Xiaoyu Chen, Yang Liu, Zhenxin Cao, Xiaopeng Li, and Jinde Cao. H-core decomposition for directed networks and its application. Scientometrics, 129(11):6571– 6596, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05170-5.

Chavalarias:2022:DMS

[CLD22] David Chavalarias, Quentin Lobbé, and Alexandre Delanoë. Draw me science. Scientometrics, 127(1):545–575, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04186-5.

Cao:2022:MSK

[CLL22] Xia Cao, Chuanyun Li, and Yunchang Li. Modeling and simulation of knowledge creation and diffusion in an industry-university-research cooperative innovation network: a case study of China’s new energy vehicles. Scientometrics, 127(7):3935–3957, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04416-4.

Castriotta:2021:DCE

[CLMM21] Manuel Castriotta, Michela Loi, Elona Marku, and Ludovica Moi. Disentangling the corporate entrepreneurship construct: conceptualizing through co-words. Scientometrics, 126(4):2821–2863, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03846-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03846-2.pdf.

Chankseliani:2021:BPB

[CLP21] Maia Chankseliani, Andrey Lovakov, and Vladimir Pislyakov. A big picture: bibliometric study of academic publications from post-Soviet countries. Scientometrics, 126(10):8701–8730, October 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04124-5.

Coupe:2025:DRR

[CLR25] Tom Coupé, Thomas Logchies, and W. Robert Reed. Do replications receive fewer citations? A counterfactual approach. Scientometrics, 130(4):2403–2423, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05288-0.

Correa:2022:SHE

[CLRMR22] Juan C. Correa, Henry Laverde-Rojas, and Fernando Marmolejo-Ramos. The Sci-Hub effect on papers’ citations. Scientometrics, 127(1):99–126, January 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03806-w.

Chughtai:2020:EOB

[CLS^{+}20] Gohar Rehman Chughtai, Jia Lee, Mahnoor Shahzadi, Asif Kabir, and Muhammad Arshad Shehzad Hassan. An efficient ontology-based topic-specific article recommendation model for best-fit reviewers. Scientometrics, 122(1):

249–265, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03261-2.

Cho:2021:PFT

[CLS21] Joon Hyung Cho, Jungpyo Lee, and So Young Sohn. Predicting future technological convergence patterns based on machine learning using link prediction. Scientometrics, 126 (7):5413–5429, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03999-8.

Chen:2024:RJV

[CLWlY24] Kun Chen, Xu Liu, Abduhalik Wupur, and Guo liang Yang. Ranking journals by voting with feet: a new method for journal evaluation. Scientometrics, 129(3):1567–1588, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04888-y.

Jaewoong:2022:TSD

[CLY^{+}22] Jaewoong Choi, Jiho Lee, Janghyeok Yoon, Sion Jang, Jaeyoung Kim, and Sungchul Choi. A two-stage deep learning-based system for patent citation recommendation. Scientometrics, 127(11):6615–6636, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04301-0.

Choi:2021:SKP

[CLZ21] Minsoo Choi, Heejin Lee, and Hanah Zoo. Scientific knowledge production and research collaboration between Australia and South Korea: patterns and dynamics based on co-authorship. Scientometrics, 126(1):683–706, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03765-2.

Cicero:2020:UJC

[CM20] Tindaro Cicero and Marco Malgarini. On the use of journal classification in social sciences and humanities: evidence from an Italian database. Scientometrics, 125(2):1689– 1708, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03581-8.

CostaRibeiro:2022:TII

[CM22] Leonardo Costa Ribeiro and Ulisses Pereira dos Santos and Valbona Muzaka. Trademarks as an indicator of innovation: towards a fuller picture. Scientometrics, 127(1): 481–508, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04197-2.

Cugmas:2024:LPS

[CMK24] Marjan Cugmas, Franc Mali, and Luka Kronegger. Longitudinal patterns of scientific collaboration in doctoral studies. Scientometrics, 129(2):1055–1077, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04900-5.

Checchi:2020:CPI

[CMMO20] Daniele Checchi, Irene Mazzotta, Sandro Momigliano, and Francesco Olivanti. Convergence or polarisation? The impact of research assessment exercises in the Italian case. Scientometrics, 124(2):1439–1455, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03517-2.

Chan:2024:LSL

[CMT24] Ho Fai Chan, Franklin G. Mixon, Jr., and Benno Torgler. Lifespan and scientific leadership: a counterfactual analysis between presidents and fellows of the Royal Society. Scientometrics, 129(3):1615–1635, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04922-z.

Cerdeira:2023:IRC

[CMV23] Jorge Cerdeira, João Mesquita, and Elizabeth S. Vieira. International research collaboration: is Africa different? A cross-country panel data analysis. Scientometrics, 128 (4):2145–2174, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04659-9.

Charen:2020:CEP

[CMZ^{+}20a] Daniel A. Charen, Nolan A. Maher, Nicole Zubizarreta, Jashvant Poeran, Calin S. Moucha, and Shai Shemesh. Correction to: Evaluation of publication delays in the orthopedic surgery manuscript review process from 2010 to 2015. Scientometrics, 124(2):1137, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03603-5;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03603-5.pdf. See [CMZ^{+}20b].

Charen:2020:EPD

[CMZ^{+}20b] Daniel A. Charen, Nolan A. Maher, Nicole Zubizarreta, Jashvant Poeran, Calin S. Moucha, and Shai Shemesh. Evaluation of publication delays in the orthopedic surgery manuscript review process from 2010 to 2015. Scientometrics, 124(2):1127–1135, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03493-7. See correction [CMZ^{+}20a].

Cugmas:2020:SCR

[CMZ20c] Marjan Cugmas, Franc Mali, and Ales Ziberna. Scientific collaboration of researchers and organizations: a two-level blockmodeling approach. Scientometrics, 125(3):2471– 2489, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03708-x.

Chang:2024:CDT

[CN24] Yu-Wei Chang and Majid Nabavi. Comparison of disciplines, topics, and methods in studies in Journal of Informetrics and Scientometrics from 2016 to 2020. Scientometrics, 129(3):1415–1439, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04947-y.

Chen:2022:CHQ

[CNA22] Haihua Chen, Huyen Nguyen, and Asmaa Alghamdi. Constructing a high-quality dataset for automated creation of summaries of fundamental contributions of research articles. Scientometrics, 127(12):7061–7075, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04380-z.

Caetano:2021:MTM

[CNP21] Ana Caetano, Magda Nico, and Anabela Pereira. More than meets the eye: traditions, nucleus and peripheries of the biographical research field. Scientometrics, 126(7): 5707–5726, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04020-y.

Chatrinan:2025:GCL

[CNT25] Krittin Chatrinan, Thanapon Noraset, and Suppawong Tuarob. GAN-CITE: leveraging semi-supervised generative adversarial networks for citation function classification with limited data. Scientometrics, 130(2):679–703, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05233-1.

Cabanac:2021:DDD

[COB21] Guillaume Cabanac, Theodora Oikonomidi, and Isabelle Boutron. Day-to-day discovery of preprint-publication links. Scientometrics, 126(6):5285–5304, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03900-7.

Coccia:2020:ESD

[Coc20] Mario Coccia. The evolution of scientific disciplines in applied sciences: dynamics and empirical properties of experimental physics. Scientometrics, 124(1):451–487, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03464-y.

Coccia:2021:ESR

[Coc21] Mario Coccia. Evolution and structure of research fields driven by crises and environmental threats: the COVID-19 research. Scientometrics, 126(12):9405–9429, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04172-x.

Cooke:2025:THS

[Coo25] Philip Cooke. Triple Helix selection in the regional innovation systems field. In memoriam: Loet Leydesdorff. Scientometrics, 130(6):3127–3138, June 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05321-2.

Copiello:2020:ACA

[Cop20a] Sergio Copiello. The alleged citation advantage of video abstracts may be a matter of self-citations and self-selection bias. Comment on ”The impact of video abstract on citation counts” by Zong et al. Scientometrics, 122(1): 751–757, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03173-1.

Copiello:2020:DMT

[Cop20b] Sergio Copiello. Digital multimedia tools, research impact, stated and revealed preferences: a rejoinder on the issue of video abstracts. Scientometrics, 123(1):543–551, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03370-3.

Copiello:2020:MCA

[Cop20c] Sergio Copiello. Multi-criteria altmetric scores are likely to be redundant with respect to a subset of the underlying information. Scientometrics, 124(1):819–824, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03491-9.

Copiello:2020:OTD

[Cop20d] Sergio Copiello. Other than detecting impact in advance, alternative metrics could act as early warning signs of retractions: tentative findings of a study into the papers retracted by PLoS ONE. Scientometrics, 125(3):2449– 2469, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03698-w.

Cortes:2022:IDM

[Cor22] Julián D. Cortés. Identifying the dissension in management and business research in Latin America and the Caribbean via co-word analysis. Scientometrics, 127(12):7111–7125, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04259-5.

Cottineau:2022:WDA

[Cot22] Clémentine Cottineau. What do analyses of city size distributions have in common? Scientometrics, 127(3): 1439–1463, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04256-8.

Chong:2021:CTE

[CP21] Miyoung Chong and Han Woo Park. COVID-19 in the Twitterverse, from epidemic to pandemic: informationsharing behavior and Twitter as an information carrier. Scientometrics, 126(8):6479–6503, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04054-2.

Campos-Perez:2025:ETM

[CPFVPA25] Lara Campos-Pérez, Xochitl Flores-Vargas, and Miguel Ángel Pérez-Angón. Epistemological transformations on mineralogy in Mexico’s transition as an independent country: a geohistoriometric perspective. Scientometrics, 130(7): 3987–4009, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05371-6.

Choi:2021:IET

[CPL21] Youngjae Choi, Sanghyun Park, and Sungjoo Lee. Identifying emerging technologies to envision a future innovation ecosystem: a machine learning approach to patent data. Scientometrics, 126(7):5431–5476, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04001-1.

Carvallo:2020:ADS

[CPLS20] Andres Carvallo, Denis Parra, Hans Lobel, and Alvaro Soto. Automatic document screening of medical literature using word and text embeddings in an active learning setting. Scientometrics, 125(3):3047–3084, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03648-6.

Chang:2021:GMA

[CPN21] Livia Lin-Hsuan Chang, Frederick Kin Hing Phoa, and Junji Nakano. A generative model of article citation networks of a subject from a large-scale citation database. Scientometrics, 126(9):7373–7395, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04037-3.

Chavarro:2022:CBH

[CPTÁ22] Diego Chavarro, Jaime Andrés Perez-Taborda, and Alba Ávila. Connecting brain and heart: artificial intelligence for sustainable development. Scientometrics, 127(12): 7041–7060, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04299-5.

Coccia:2024:ETT

[CR24] Mario Coccia and Saeed Roshani. Evolution of topics and trends in emerging research fields: multiple analyses with entity linking, Mann–Kendall test and burst methods in cloud computing. Scientometrics, 129(9):5347– 5371, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05139-4.

Chinchilla-Rodriguez:2021:ERD

[CRBRGS21] Zaida Chinchilla-Rodríguez, Yi Bu, Nicolás Robinson-García, and Cassidy R. Sugimoto. An empirical review of the different variants of the probabilistic affinity index as applied to scientific collaboration. Scientometrics, 126(2): 1775–1795, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03815-9.

Caliari:2020:RIL

[CRC20] Thiago Caliari, Márcia Siqueira Rapini, and Tulio Chiarini. Research infrastructures in less developed countries: the Brazilian case. Scientometrics, 122(1):451–475, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03245-2.

Chacon:2020:CIS

[CSA20] Xiomara S. Q. Chacon, Thiago C. Silva, and Diego R. Amancio. Comparing the impact of subfields in scientific journals. Scientometrics, 125(1):625–639, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03651-x.

Chen:2023:TFE

[CSC23] Lingzhi Chen, Yutao Sun, and Cong Cao. A two-fold evaluation in science: the case of Nobel Prize. Scientometrics, 128(11):6267–6291, November 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04830-2.

Chen:2022:NDU

[CSJW22] Hongshu Chen, Xinna Song, Qianqian Jin, and Ximeng Wang. Network dynamics in university-industry collaboration: a collaboration-knowledge dual-layer network perspective. Scientometrics, 127(11):6637–6660, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04330-9.

Chen:2022:IIE

[CSL22] Shiji Chen, Yanhui Song, and Vincent Larivière. Interdisciplinarity and impact: the effects of the citation time window. Scientometrics, 127(5):2621–2642, May 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04338-1.

Cumberledge:2023:UHA

[CSR23] Aaron Cumberledge, Neal Smith, Jr., and Benjamin W. Riley. Unverified history: an analysis of quotation accuracy in leading history journals. Scientometrics, 128(8): 4677–4687, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04755-w.

Chan:2020:GDP

[CT20] Ho Fai Chan and Benno Torgler. Gender differences in performance of top cited scientists by field and country. Scientometrics, 125(3):2421–2447, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03733-w.

Cao:2022:CEE

[CTH22] Qinwei Cao, Manqing Tan, and Jian Huang. Can emerging economies take advantage of their population size to gain international academic recognition? Evidence from key universities in China. Scientometrics, 127(2):927– 957, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04218-0.

Cai:2025:DAM

[CTH25] Ruonan Cai, Wencan Tian, and Zhigang Hu. Do articles with multiple corresponding authorships have a citation advantage? A double machine learning analysis approach. Scientometrics, 130(5):2497–2522, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05242-0.

Cai:2024:GMR

[CTLH24] Ruonan Cai, Wencan Tian, Rundong Luo, and Zhigang Hu. The generation mechanism of research leadership in international collaboration based on GERGM: a case from the field of artificial intelligence. Scientometrics, 129(10): 5821–5839, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04974-9.

Chen:2022:OSC

[CtRbQ22] Kun Chen, Xian tong Ren, and Hai bo Qin. The other side of the coin: The declining of Chinese social science. Scientometrics, 127(1):127–143, January 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04208-2.

Calignano:2024:RBA

[CW24] Giuseppe Calignano and Elisabeth Winsents. Relations between academic reputation and innovation networks. Scientometrics, 129(2):889–908, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04869-1.

Chai:2025:WKR

[CW25] Xinhua Chai and Qiang Wu. What kind of research network configurations lead to high academic productivity for young management scholars? — A fuzzy-set qualitative comparative analysis (fsQCA). Scientometrics, 130 (5):2673–2704, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05329-8.

Chen:2021:GCF

[CWK21] Yue Chen, Zhiqi Wang, and Hildrun Kretschmer. Growing with collaboration: footprint of WISE Lab. Scientometrics, 126(7):6147–6167, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03687-z.

Cheng:2020:KCK

[CWL^{+}20] Qikai Cheng, Jiamin Wang, Wei Lu, Yong Huang, and Yi Bu. Keyword-citation-keyword network: a new perspective of discipline knowledge structure analysis. Scientometrics, 124(3):1923–1943, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03576-5.

Chen:2023:III

[CWLS23] Yitong Chen, Keye Wu, Yue Li, and Jianjun Sun. Impacts of inter-institutional mobility on scientific performance from research capital and social capital perspectives. Scientometrics, 128(6):3473–3506, June 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04690-w.

Cao:2021:MBH

[CWZ21] Renmeng Cao, Xianwen Wang, and Jianlin Zhou. Multiple bursts of highly retweeted articles on social media. Scientometrics, 126(6):5165–5179, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03970-7.

Chen:2022:MMA

[CWZ22] Yuetong Chen, Hao Wang, and Wei Zhang. A method of measuring the article discriminative capacity and its distribution. Scientometrics, 127(6):3317–3341, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04371-0.

Cao:2021:LST

[CXD21] Qinwei Cao, Peng Xie, and Wanchun Duan. The larger scientific and technological human scale, the better innovation effect? Evidence from key universities in China. Scientometrics, 126(7):5623–5649, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04019-5.

Chen:2024:DDS

[CxGdH^{+}24] Kun Chen, Xia xia Gao, Yi di Huang, Wen tao Xu, and Guo liang Yang. The dual dimension of scientific research experience acquisition and its development: a 40-year analysis of Chinese Humanities and Social Sciences journals. Scientometrics, 129(5):2827–2853, May 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05002-6.

Cui:2024:ATA

[CXW^{+}24] Wentao Cui, Meng Xiao, Ludi Wang, Xuezhi Wang, Yi Du, and Yuanchun Zhou. Automated taxonomy alignment via large language models: bridging the gap between knowledge domains. Scientometrics, 129(9):5287–5312, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05111-2.

Chen:2020:DLB

[CXZ^{+}20] Liang Chen, Shuo Xu, Lijun Zhu, Jing Zhang, Xiaoping Lei, and Guancan Yang. A deep learning based method for extracting semantic information from patent documents. Scientometrics, 125(1):289–312, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03634-y.

Cheng:2025:IID

[CXZM25] Deng Cheng, Zhang Xue, Yang Zhibo, and Zhang Mingze. Impact of interdisciplinarity on disruptive innovation: the moderating role of collaboration pattern and collaboration size. Scientometrics, 130(4):2379–2401, April 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05285-3.

Chang:2025:DPA

[CY25] Yu-Wei Chang and Hsuan-Tung Yeh. Do prolific arts and humanities authors have publishing preferences? Scientometrics, 130(2):1281–1302, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05247-9.

Choi:2022:MGG

[CYK22] Myoungjae Choi, Sun-Hi Yoo, and Byunghoon Kim. A modified gamma/Gompertz/NBD model for estimating technology lifetime. Scientometrics, 127(10):5731–5751, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04489-1.

Cheng:2024:CTH

[CZFC24] Wei Cheng, Dejun Zheng, Shaoxiong Fu, and Jingfeng Cui. Closer in time and higher correlation: disclosing the relationship between citation similarity and citation interval. Scientometrics, 129(7):4495–4512, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05080-6.

Chen:2024:IER

[CZQC24] Shiji Chen, Kaiqi Zhang, Junping Qiu, and Jiaqi Chai. Interdisciplinarity and expert rating: an analysis based on faculty opinions. Scientometrics, 129(11):6597–6628, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05145-6.

Cao:2024:EFM

[CZX24] Xia Cao, Xin Zhang, and ZeYu Xing. Exploring the formation mechanism of new energy vehicle industry-university-research innovation network: the role of structural, cognitive and relational social capital. Scientometrics, 129(6): 3045–3065, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05050-y.

Cheng:2025:CRP

[CZZN25] Wei Cheng, Dejun Zheng, Xiaomin Zheng, and Huanhuan Ni. Combining referenced publication year spectroscopy and topic clustering to identify key knowledge foundations in scientometrics: an analysis of recipients of the Price Award. Scientometrics, 130(2):1077–1099, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05230-4.

DeFilippo:2020:TCS

[DABSC20] Daniela De Filippo, Rafael Aleixandre-Benavent, and Elías Sanz-Casado. Toward a classification of Spanish scholarly journals in social sciences and humanities considering their impact and visibility. Scientometrics, 125(2):1709– 1732, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03665-5.

Dagiene:2024:MSB

[Dag24] Eleonora Dagiene. Mapping scholarly books: library metadata and research assessment. Scientometrics, 129(9): 5689–5714, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-024-
05120-1.

Dyachenko:2024:IRM

[DAGG24] Ekaterina Dyachenko, Iurii Agafonov, Katerina Guba, and Alexander Gelvikh. Independent Russian medical science: is there any? Scientometrics, 129(9):5577–5597, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05135-8.

Day:2022:EAT

[Day22] Adam Day. Exploratory analysis of text duplication in peer-review reveals peer-review fraud and paper mills. Scientometrics, 127(10):5965–5987, October 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04504-5.

diBella:2021:ASC

[dBGP21] Enrico di Bella, Luca Gandullia, and Sara Preti. Analysis of scientific collaboration network of Italian Institute of Technology. Scientometrics, 126(10):8517–8539, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04120-9.

Dumaz:2021:ACC

[DBR21] Marie Dumaz, Reese Boucher, and Aldo H. Romero. Authorship and citation cultural nature in Density Functional Theory from solid state computational packages. Scientometrics, 126(8):6681–6695, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04057-z.

Dip:2025:PAP

[DCdA25] Juan Antonio Dip, David Dominguez Casoratti, and Facundo Costa de Arguibel. The productivity of Argentine public universities from 2013 to 2022: analysing the Malmquist Index through global and local technical changes. Scientometrics, 130(7):3961–3985, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05333-y.

doCarmo:2023:IJH

[dCFdCA^{+}23] Gisleine do Carmo, Luiz Flávio Felizardo, Valderí de Castro Alcântara, Cristiane Aparecida da Silva, and José Willer do Prado. The impact of Jürgen Habermas’s scientific production: a scientometric review. Scientometrics, 128(3):1853–1875, March 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04625-x.

Ding:2023:ERF

[DCL23] Jingda Ding, Yifan Chen, and Chao Liu. Exploring the research features of Nobel laureates in Physics based on the semantic similarity measurement. Scientometrics, 128(9):5247–5275, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04786-3.

doCanto:2022:LAC

[dCPT22] Fabio Lorensi do Canto, Adilson Luiz Pinto, and Marcos Talau. Latin American and Caribbean journals indexed in Google Scholar Metrics. Scientometrics, 127(2): 763–783, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04237-x.

DeNicola:2021:AGD

[DD21a] Antonio De Nicola and Gregorio D’Agostino. Assessment of gender divide in scientific communities. Scientometrics, 126(5):3807–3840, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03885-3.

Durmusoglu:2021:TMU

[DD21b] Zeynep Didem Unutmaz Durmusoglu and Alptekin Durmusoglu. A TOPSIS model for understanding the authors choice of journal selection. Scientometrics, 126(1): 521–543, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03770-5.

Ding:2023:SCB

[DD23] Jingda Ding and Dehui Du. A study of the correlation between publication delays and measurement indicators of journal articles in the social network environment-based on online data in PLOS. Scientometrics, 128(3):1711–1743, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04640-6.

Deutz:2021:QQS

[DDD^{+}21] Daniella Bayle Deutz, Thea Marie Drachen, Dorte Drongstrup, Niels Opstrup, and Charlotte Wien. Quantitative quality: a study on how performance-based measures may change the publication patterns of Danish researchers. Scientometrics, 126(4):3303–3320, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03881-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
021-03881-7.pdf.

Djuric:2020:CBQ

[DDF20] Mladen Djuric, Marina Dobrota, and Jovan Filipovic. Complexity-based quality indicators for human and social capital in science and research: the case of Serbian Homeland versus Diaspora. Scientometrics, 124(1): 303–328, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03428-2.

Ding:2021:RRB

[DDH21] Yang Ding, Xianlei Dong, and Beibei Hu. Revisiting the relationship between downloads and citations: a perspective from papers with different citation patterns in the case of the Lancet. Scientometrics, 126(9):7609–7621, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04099-3.

Daraio:2023:HAB

[DDL23] Cinzia Daraio, Simone Di Leo, and Loet Leydesdorff. A heuristic approach based on Leiden rankings to identify outliers: evidence from Italian universities in the European landscape. Scientometrics, 128(1):483–510, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04551-y.

deDampierre:2025:SES

[dDM25] Charles de Dampierre and Hugo Mercier. The structure and evolution of scholarly interests from antiquity to the eighteenth century. Scientometrics, 130(7):3367–3382, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05340-z.

Daraio:2022:AQD

[DDS22] Cinzia Daraio, Simone Di Leo, and Monica Scannapieco. Accounting for quality in data integration systems: a completeness-aware integration approach. Scientometrics, 127(3):1465–1490, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04266-0.

DellaLucia:2024:HOC

[DDU24] Maria Della Lucia, Giulia Dore, and Rana Muhammad Umar. Handling the open culture dilemma in museum management: an exploratory interdisciplinary study. Scientometrics, 129(12):7699–7733, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05164-3.

Decock:2020:SDF

[DDVV20] Kristof Decock, Koenraad Debackere, Anne-Mieke Vandamme, and Bart Van Looy. Scenario-driven forecasting: modeling peaks and paths. Insights from the COVID-19 pandemic in Belgium. Scientometrics, 124 (3):2703–2715, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03591-6;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03591-6.pdf.

Deja:2025:CEG

[Dej25] Marek Deja. The causal effect of the global crisis on open science research impact: a bibliometric causal analysis. Scientometrics, 130(2):1303–1325, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05241-1.

Demaine:2022:FRI

[Dem22] Jeffrey Demaine. Fractionalization of research impact reveals global trends in university collaboration. Scientometrics, 127(5):2235–2247, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04246-w.

deFrutos-Belizon:2024:EAI

[dFBGCGASG24] Jesús de Frutos-Belizón, Natalia García-Carbonell, Félix Guerrero-Alba, and Gonzalo Sánchez-Gardey. An empirical analysis of individual and collective determinants of international research collaboration. Scientometrics, 129 (5):2749–2770, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04999-0.

Deng:2021:EII

[DFG21] Liming Deng, Bagheri Fatemeh, and Xiaoping Gao. Exploring the interactive and interactional metadiscourse in doctoral dissertation writing: a diachronic study. Scientometrics, 126(8):7223–7250, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04064-0.

Demetrescu:2020:BAP

[DFRS20] Camil Demetrescu, Irene Finocchi, Andrea Ribichini, and Marco Schaerf. On bibliometrics in academic promotions: a case study in computer science and engineering in Italy. Scientometrics, 124(3):2207–2228, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03548-9.

Demetrescu:2022:CSR

[DFS22] Camil Demetrescu, Irene Finocchi, and Marco Schaerf. On computer science research and its temporal evolution. Scientometrics, 127(8):4913–4938, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04445-z.

Vieira:2024:NDS

[dFVFA^{+}24] Vinícius da F. Vieira, Carlos H. G. Ferreira, Jussara M. Almeida, Edré Moreira, Alberto H. F. Laender, Wagner Meira, and Marcos André Gonçalves. A networkdriven study of hyperprolific authors in computer science. Scientometrics, 129(4):2255–2283, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04940-5.

Duan:2021:PPK

[DG21] Yueran Duan and Qing Guan. Predicting potential knowledge convergence of solar energy: bibliometric analysis based on link prediction model. Scientometrics, 126(5): 3749–3773, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03901-6.

Duan:2024:FDC

[DG24] Weiyu Duan and Ying Guo. Focused or divided? Collaborative attention and technological orientations in technology transfer. Scientometrics, 129(3):1441–1467, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04897-x.

Vanz:2023:CSC

[dGdO^{+}23] Samile Andréa de Souza Vanz, Maria Claudia Cabrini Gracio, Sandra Cristina de Oliveira, Zaida Chinchilla-Rodríguez, and Domingo Docampo. Collaboration strategies and corresponding authorship in agronomy research of Brazilian academic and non-academic institutions. Scientometrics, 128(12):6403–6426, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04857-5.

Dorta-Gonzalez:2021:WER

[DGGBDG21] Pablo Dorta-González, Sara M. González-Betancor, and María Isabel Dorta-González. To what extent is researchers’ data-sharing motivated by formal mechanisms of recognition and credit? Scientometrics, 126(3):2209–2225, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03869-3.

Dorta-Gonzalez:2022:MOR

[DGGD22] Pablo Dorta-González and Emilio Gómez-Déniz. Modeling the obsolescence of research literature in disciplinary journals through the age of their cited references. Scientometrics, 127(6):2901–2931, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04359-w.

deGouveia:2021:ERB

[dGIL21] M. de Gouveia and R. Inglesi-Lotz. Examining the relationship between climate change-related research output and CO_{2} emissions. Scientometrics, 126(11):9069– 9111, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04148-x.

Daraio:2022:OHF

[DGN22] Cinzia Daraio, Wolfgang Glänzel, and Ed Noyons. Obituary: [Henk F. Moed (1951–2021)]. Scientometrics, 127(2):

683–685, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04257-7.

Dorantes-Gilardi:2023:TDG

[DGRÁTS23] Rodrigo Dorantes-Gilardi, Aurora A. Ramírez-Álvarez, and Diana Terrazas-Santamaría. Is there a differentiated gender effect of collaboration with super-cited authors? Evidence from junior researchers in economics. Scientometrics, 128(4):2317–2336, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04656-y.

Dorta-Gonzalez:2020:OAE

[DGSVDG20] Pablo Dorta-González, Rafael Suárez-Vega, and María Isabel Dorta-González. Open access effect on uncitedness: a large-scale study controlling by discipline, source type and visibility. Scientometrics, 124(3):2619–2644, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03557-8.

Doehne:2023:HAO

[DH23] Malte Doehne and Catherine Herfeld. How academic opinion leaders shape scientific ideas: an acknowledgment analysis. Scientometrics, 128(4):2507–2533, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04623-z.

Dhamija:2022:SAE

[Dha22] Pavitra Dhamija. South Africa in the era of Industry 4.0: An insightful investigation. Scientometrics, 127(9):5083– 5110, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04461-z.

Djupe:2020:PPC

[DHSS20] Paul A. Djupe, Kim Quaile Hill, Amy Erica Smith, and Anand E. Sokhey. Putting personality in context: determinants of research productivity and impact in political science. Scientometrics, 124(3):2279–2300, September 2020.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03592-5.

Diao:2021:LSS

[Dia21] Junli Diao. A lexical and syntactic study of research article titles in Library Science and Scientometrics. Scientometrics, 126(7):6041–6058, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04018-6.

Dip:2021:WDU

[Dip21] Juan Antonio Dip. What does U-multirank tell us about knowledge transfer and research? Scientometrics, 126(4): 3011–3039, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03838-2.

Demeter:2022:MMP

[DJM22] Marton Demeter, Agnes Jele, and Zsolt Balázs Major. The model of maximum productivity for research universities SciVal author ranks, productivity, university rankings, and their implications. Scientometrics, 127(8):4335–4361, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04432-4.

Dehdarirad:2021:NMA

[DK21a] Tahereh Dehdarirad and Kalle Karlsson. News media attention in Climate Action: latent topics and open access. Scientometrics, 126(9):8109–8128, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04095-7.

Demin:2021:TCG

[DK21b] Maxim Demin and Alexei Kouprianov. Three centuries of German-language philosophy journals (1765–1953): a bibliometric analysis. Scientometrics, 126(7):5651–5664, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04009-7.

Donner:2025:AAO

[DK25] Paul Donner and Przemysław Korytkowski. Alphabetical author order and co-author contributions in mathematics. Scientometrics, 130(7):4117–4140, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05369-0.

Devarakonda:2020:VCS

[DKWC20] Sitaram Devarakonda, Dmitriy Korobskiy, Tandy Warnow, and George Chacko. Viewing computer science through citation analysis: Salton and Bergmark redux. Scientometrics, 125(1):271–287, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03624-0.

Dong:2023:GRT

[DL23] Yijie Dong and Danyang Li. Gender representation in textbooks: a bibliometric study. Scientometrics, 128(11):5969– 6001, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04834-y.

Barcellos:2025:IIP

[dLBdAPUdLB^{+}25] Rodrigo da Luz Barcellos, Daniel de Abreu Pereira Uhr, Rodrigo da Luz Barcellos, Daniel de Abreu Pereira Uhr, Eliana de Lemos Crestani, Lígia Mori Madeira, and Marcelo Nogueira Cortimiglia. The impact of innovation policies on scientometric indicators: a study through the combination of punctuated equilibrium theory and synthetic control method. Scientometrics, 130(4):2185–2212, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05272-8.

Ding:2021:NMC

[DLC21] Jingda Ding, Chao Liu, and Wei Cai. A new method of co-author credit allocation based on contributor roles taxonomy: proof of concept and evaluation using papers published in PLOS ONE. Scientometrics, 126(9):7561– 7581, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04075-x.

Ding:2020:ELH

[DLK20] Jingda Ding, Chao Liu, and Goodluck Asobenie Kandonga. Exploring the limitations of the h-index and htype indexes in measuring the research performance of authors. Scientometrics, 122(3):1303–1322, March 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03364-1.

Dong:2021:ITS

[DLL21] Jinyang Dong, Jiamou Liu, and Tiezhong Liu. The impact of top scientists on the community development of basic research directed by government funding: evidence from program 973 in China. Scientometrics, 126(10):8561–8579, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04092-w.

Dua:2023:MCR

[DLS23] Jyoti Dua, Hiran H. Lathabai, and Vivek Kumar Singh. Measuring and characterizing research collaboration in SAARC countries. Scientometrics, 128(2):1265–1294, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04606-0.

Ding:2025:FPD

[DM25] Yang Ding and Fernando Moreira. Funding and productivity: Does winning grants affect the scientific productivity of recipients? Evidence from the social sciences and economics. Scientometrics, 130(3):1831–1870, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05277-3.

Drongstrup:2020:CSM

[DMA^{+}20] Dorte Drongstrup, Shafaq Malik, Naif Radi Aljohani, Salem Alelyani, Iqra Safder, and Saeed-Ul Hassan. Can social media usage of scientific literature predict journal indices of AJG, SNIP and JCR? An altmetric study of economics. Scientometrics, 125(2):1541–1558, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03613-3.

deMoya-Anegon:2020:CIS

[dMALIGBM20] Felix de Moya-Anegon, Carmen Lopez-Illescas, Vicente Guerrero-Bote, and Henk F. Moed. The citation impact of social sciences and humanities upon patentable technology. Scientometrics, 125(2):1665–1687, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03530-5.

Daraio:2020:ICS

[DMG20] Cinzia Daraio, Henk F. Moed, and Wolfgang Glänzel. The 17th International Conference on Scientometrics and Informetrics. Scientometrics, 125(2):831–834, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03768-z.

Dumas-Mallet:2020:DNC

[DMGBG20] Estelle Dumas-Mallet, André Garenne, Thomas Boraud, and François Gonon. Does newspapers coverage influence the citations count of scientific publications? An analysis of biomedical studies. Scientometrics, 123(1): 413–427, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03380-1.

Grochocki:2023:RCN

[dMGC23] Luís Filipe de Miranda Grochocki and Andrea Felippe Cabello. Research collaboration networks in maturing academic environments. Scientometrics, 128(4):2535–2556, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04671-z.

de-Marcos:2024:MST

[dMGDD24] Luis de Marcos, Manuel Goyanes, and Adrián Domínguez-Díaz. Mapping science through editorial board interlocking: connections and distance between fields of knowledge and institutional affiliations. Scientometrics, 129(6):

3385–3406, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05027-x.

Maricato:2022:CCA

[dMMdCM22] João de Melo Maricato and Bruno Lara de Castro Manso. Characterization of the communities of attention interacting with scientific papers on Twitter: altmetric analysis of a Brazilian university. Scientometrics, 127(7):3815–3835, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04442-2.

DeFilippo:2020:OAI

[DMR20] Daniela De Filippo and Jorge Mañana-Rodríguez. Open access initiatives in European universities: analysis of their implementation and the visibility of publications in the YERUN network. Scientometrics, 125(3):2667–2694, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03705-0.

Filippo:2022:PIO

[DMR22] Daniela De Filippo and Jorge Mañana-Rodriguez. The practical implementation of open access policies and mandates in Spanish public universities. Scientometrics, 127 (12):7147–7167, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04261-x.

Dion:2020:GSS

[DMS20] Michelle L. Dion, Sara McLaughlin Mitchell, and Jane L. Sumner. Gender, seniority, and self-citation practices in political science. Scientometrics, 125(1):1–28, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03615-1.

deOliveira:2021:MRS

[dOCP21] Sandra Cristina de Oliveira, Juliana Cobre, and Danilo Florentino Pereira. A measure of reliability for scientific coauthorship networks using fuzzy logic. Scientometrics, 126 (6):4551–4563, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03915-0.

Donner:2021:CAP

[Don21a] Paul Donner. Citation analysis of Ph.D. theses with data from Scopus and Google Books. Scientometrics, 126(12): 9431–9456, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04173-w.

Donner:2021:VAD

[Don21b] Paul Donner. Validation of the Astro dataset clustering solutions with external data. Scientometrics, 126(2):1619– 1645, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03780-3.

Donner:2022:AIP

[Don22] Paul Donner. Algorithmic identification of Ph.D. thesisrelated publications: a proof-of-concept study. Scientometrics, 127(10):5863–5877, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04480-w.

Passos:2022:GCA

[dOPGdAPU22] Marcelo de Oliveira Passos, Priscila Lujan Gonzalez, and Daniel de Abreu Pereira Uhr. The greatest co-authorships of finance theory literature (1896–2006): scientometrics based on complex networks. Scientometrics, 127(10):5841– 5862, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04482-8.

Amaral:2021:CUE

[dOSAL21] Emilyane de Oliveira Santana Amaral and Sergio Roberto Peres Line. Current use of effect size or confidence interval analyses in clinical and biomedical research. Scientometrics, 126(11):9133–9145, November 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04150-3.

Girolamo:2020:CSA

[DR20] Nicola Di Girolamo and Reint Meursinge Reynders. Characteristics of scientific articles on COVID-19 published during the initial 3 months of the pandemic. Scientometrics, 125(1):795–812, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03632-0;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03632-0.pdf.

Dogramaci:2021:PSP

[DRF21] Esma J. Dogramaci and Giampiero Rossi-Fedele. Predictors of societal and professional impact of orthodontic research. A multivariate, scientometric approach. Scientometrics, 126(11):9223–9248, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04163-y.

Drivas:2024:EOA

[Dri24] Kyriakos Drivas. The evolution of order of authorship based on researchers’ age. Scientometrics, 129(9):5615– 5633, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05124-x.

delRio:2020:APM

[dRRJ20] J. Antonio del Río, J. M. Russell, and Daniela Juárez. Applied physics in Mexico: mining the past to predict the future. Scientometrics, 125(1):187–212, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03639-7.

Demetrescu:2020:IRA

[DRS20] Camil Demetrescu, Andrea Ribichini, and Marco Schaerf. Are Italian research assessment exercises size-biased? Scientometrics, 125(1):533–549, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03643-x;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03643-x.pdf.

Donner:2020:IPB

[DS20] Paul Donner and Ulrich Schmoch. The implicit preference of bibliometrics for basic research. Scientometrics, 124(2): 1411–1419, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03516-3.

See correction [DS22].

daSilva:2021:AOW

[dS21a] Jaime A. Teixeira da Silva. Abuse of ORCID’s weaknesses by authors who use paper mills. Scientometrics, 126(7): 6119–6125, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03996-x.

daSilva:2021:CRC

[dS21b] Jaime A. Teixeira da Silva. CiteScore: risk of copy-cat, fake and misleading metrics. Scientometrics, 126(2):1859– 1862, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03791-0.

daSilva:2021:III

[dS21c] Jaime A. Teixeira da Silva. The i100-index, i1000-index and i10,000-index: expansion and fortification of the Google Scholar h-index for finer-scale citation descriptions and researcher classification. Scientometrics, 126(4): 3667–3672, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03831-9.

Docampo:2021:JRP

[DS21d] Domingo Docampo and Vicente Safón. Journal ratings: a paper affiliation methodology. Scientometrics, 126(9): 8063–8090, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04045-3.

Donner:2022:CIP

[DS22] Paul Donner and Ulrich Schmoch. Correction to: The implicit preference of bibliometrics for basic research.

Scientometrics, 127(3):1661, March 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04181-w. See [DS20].

Decius:2025:FCI

[DS25] Julian Decius and Miriam Schilbach. Fair credit? The impact of shared first authorship on academic career evaluation. Scientometrics, 130(3):1731–1750, March 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05262-w.

daSilva:2020:SEA

[dSAKT20] Jaime A. Teixeira da Silva, Aceil Al-Khatib, and Panagiotis Tsigaris. Spam emails in academia: issues and costs. Scientometrics, 122(2):1171–1188, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03315-5.

Devriendt:2022:DSP

[DSB22] Thijs Devriendt, Mahsa Shabani, and Pascal Borry. Data sharing platforms: instruments to inform and shape science policy on data sharing? Scientometrics, 127(6): 3007–3019, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04361-2.

daSilva:2021:CLR

[dSDE21] Jaime A. Teixeira da Silva, Daniel J. Dunleavy, and Joshua Eykens. A credit-like rating system to determine the legitimacy of scientific journals and publishers. Scientometrics, 126(10):8589–8616, October 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04118-3.

Daud:2020:FRS

[DSH^{+}20] Ali Daud, Min Song, Malik Khizar Hayat, Tehmina Amjad, Rabeeh Ayaz Abbasi, Hassan Dawood, and Anwar Ghani. Finding rising stars in bibliometric networks. Scientometrics, 124(1):633–661, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03466-w.

Ding:2021:LBE

[DSL21] Jielan Ding, Zhesi Shen, and Johan Lyhagen. The link between ethnic diversity and scientific impact: the mediating effect of novelty and audience diversity. Scientometrics, 126(9):7759–7810, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04071-1.

Dua:2023:MCI

[DSL23] Jyoti Dua, Vivek Kumar Singh, and Hiran H. Lathabai. Measuring and characterizing international collaboration patterns in Indian scientific research. Scientometrics, 128(9):5081–5116, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04794-3.

deSantana:2021:MRH

[dSMNdV^{+}21] Mariana M. M. de Santana, Eduardo Mariano-Neto, Rodrigo N. de Vasconcelos, Pavel Dodonov, and José M. M. Medeiros. Mapping the research history, collaborations and trends of remote sensing in fire ecology. Scientometrics, 126(2):1359–1388, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03805-x.

daSilva:2023:PCA

[dSN23] Jaime A. Teixeira da Silva and Serhii Nazarovets. Partial citation analysis of five classes of retracted papers, and devising a new four-tier citation classification system for retracted (and other) papers. Scientometrics, 128(8): 4887–4894, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04769-4.

deSouza:2022:SCL

[dSSA22] Edson Melo de Souza, Jose Eduardo Storopoli, and Wonder Alexandre Luz Alves. Scientific Contribution List Categories Investigation: a comparison between three mainstream medical journals. Scientometrics, 127(5): 2249–2276, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04315-8.

DeFilippo:2023:IRO

[DST23] Daniela De Filippo and Pablo Sastrón-Toledo. Influence of research on open science in the public policy sphere. Scientometrics, 128(3):1995–2017, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04645-1.

daSilva:2021:PVM

[dSTE21] Jaime A. Teixeira da Silva, Panagiotis Tsigaris, and Mohammadamin Erfanmanesh. Publishing volumes in major databases related to Covid-19. Scientometrics, 126 (1):831–842, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03675-3;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03675-3.pdf.

daSilva:2021:RRU

[dSV21] Jaime A. Teixeira da Silva and Quan-Hoang Vuong. The right to refuse unwanted citations: rethinking the culture of science around the citation. Scientometrics, 126(6): 5355–5360, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03960-9.

daSilva:2023:WIA

[dSV23] Jaime A. Teixeira da Silva and Quan-Hoang Vuong. Who, if anyone, has the right to accept or refuse unwanted citations? Scientometrics, 128(7):4151–4154, July 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04732-3.

daSilva:2024:CMC

[dSVN24] Jaime A. Teixeira da Silva, Neil J. Vickers, and Serhii Nazarovets. From citation metrics to citation ethics: Critical examination of a highly-cited 2017 moth pheromone paper. Scientometrics, 129(1):693–703, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04855-7.

Demeter:2020:WSN

[DT20] Marton Demeter and Tamas Toth. The world-systemic network of global elite sociology: the Western male monoculture at faculties of the top one-hundred sociology departments of the world. Scientometrics, 124 (3):2469–2495, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03563-w;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03563-w.pdf.

Decullier:2021:WAT

[DTHM21] E. Decullier, P. V. Tang, L. Huot, and H. Maisonneuve. Why an automated tracker finds poor sharing of clinical trial results for an academic sponsor: a bibliometric analysis. Scientometrics, 126(2):1239–1248, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03775-0.

Daraio:2020:UNE

[DV20] Cinzia Daraio and Alessio Vaccari. Using normative ethics for building a good evaluation of research practices: towards the assessment of researcher’s virtues. Scientometrics, 125(2):1053–1075, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03658-4.

Daraio:2022:HSE

[DV22] Cinzia Daraio and Alessio Vaccari. How should evaluation be? Is a good evaluation of research also just? Towards the implementation of good evaluation. Scientometrics, 127(12):7127–7146, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04329-2.

DAngelo:2020:CLS

[DvE20] Ciriaco Andrea D’Angelo and Nees Jan van Eck. Collecting large-scale publication data at the level of individual researchers: a practical proposal for author name disambiguation. Scientometrics, 123(2):883–907, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03410-y.

deWinter:2024:CCU

[dW24] Joost de Winter. Can ChatGPT be used to predict citation counts, readership, and social media interaction? An exploration among 2222 scientific abstracts. Scientometrics, 129(4):2469–2487, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04939-y.

Deng:2023:PVS

[DWG23] Liming Deng, Meiling Wang, and Xiaoping Gao. Predicting the variation in stance-taking: the use of evaluativethat in English as a lingua franca academic writing. Scientometrics, 128(6):3283–3311, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04700-x.

Du:2021:PRS

[DWM21] Wei Du, Yibo Wang, and Jian Ma. A personalized recommendation system for high-quality patent trading by leveraging hybrid patent analysis. Scientometrics, 126(12):9369–9391, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04180-x.

Deng:2020:MII

[DX20] Shengli Deng and Sudi Xia. Mapping the interdisciplinarity in information behavior research: a quantitative study using diversity measure and co-occurrence analysis. Scientometrics, 124(1):489–513, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03465-x.

Du:2021:PPP

[DXL21] Wumei Du, Zheng Xie, and Yiqin Lv. Predicting publication productivity for authors: Shallow or deep architecture? Scientometrics, 126(7):5855–5879, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04027-5.

Deng:2023:ERD

[DZ23] Nan Deng and An Zeng. Enhancing the robustness of the disruption metric against noise. Scientometrics, 128(4): 2419–2428, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04644-2.

Eriksson:2020:NDR

[EBB^{+}20] Magnus Eriksson, Annika Billhult, Tommy Billhult, Elena Pallari, and Grant Lewison. A new database of the references on international clinical practice guidelines: a facility for the evaluation of clinical research. Scientometrics, 122(2):1221–1235, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03318-2;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03318-2.pdf.

Edlinger:2023:PCP

[EBW23] Moritz Edlinger, Finn Buchrieser, and Guilherme Wood. Presence and consequences of positive words in scientific abstracts. Scientometrics, 128(12):6633–6657, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04864-6.

Egret:2024:MPH

[ECD24] Daniel Egret, Lauranne Chaignon, and Domingo Docampo. Mapping the paths of highly cited researchers: a comprehensive look at the 2023 cross-field distribution. Scientometrics, 129(11):7107–7129, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05151-8.

Elevli:2024:SEI

[EE24] Sermin Elevli and Birol Elevli. A study of entrepreneur and innovative university index by entropy-based grey relational analysis and PROMETHEE. Scientometrics, 129 (6):3193–3223, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05033-z.

Elena:2021:IIC

[EEÁ21] Luna-Morales Maria Elena, Luna-Morales Evelia, and Pérez-Angón Miguel Ángel. Influence of the international collaboration in the field of metric studies of science and technology: the case of Mexico (1971–2018). Scientometrics, 126(3):2485–2511, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03522-5.

Egghe:2022:IMW

[Egg22] Leo Egghe. Impact measures: What are they? Scientometrics, 127(1):385–406, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04053-3.

Essers:2022:NER

[EGP22] Dennis Essers, Francesco Grigoli, and Evgenia Pugacheva. Network effects and research collaborations: evidence from IMF Working Paper co-authorship. Scientometrics, 127(12):7169–7192, December 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04335-4.

Espinosa-Gimenez:2023:SPO

[EGPGGA^{+}23] Julián Espinosa-Giménez, Vanessa Paredes-Gallardo, María Dolores Gómez-Adrián, Carlos Bellot-Arcís, and Verónica García-Sanz. Scientific production of an oral implantology journal: a 5-year bibliometric study. Scientometrics, 128(6):3535–3554, June 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-023-04696-4.

ElGibari:2022:CRP

[EGR22] Samira El Gibari, Trinidad Gómez, and Francisco Ruiz. Combining reference point based composite indicators with data envelopment analysis: application to the assessment of universities. Scientometrics, 127(8):4363–4395, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04436-0.

Eykens:2022:SSI

[EGV22] Joshua Eykens, Raf Guns, and Raf Vanderstraeten. Subject specialties as interdisciplinary trading grounds: the case of the social sciences and humanities. Scientometrics, 127(12):7193–7213, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04254-w.

Etzkowitz:2025:ITH

[EL25] Henry Etzkowitz and Loet Leydesdorff. Innovation and the Triple Helix. Scientometrics, 130(6):3275–3277, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05337-8.

Elango:2021:RAB

[Ela21] Bakthavachalam Elango. Retracted articles in the biomedical literature from Indian authors. Scientometrics, 126 (5):3965–3981, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03895-1.

Elango:2022:CRE

[Ela22] Bakthavachalam Elango. Characteristics of retracted editorial articles in the biomedical literature. Scientometrics, 127(3):1431–1438, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04263-9.

Elango:2025:ADD

[Ela25] Bakthavachalam Elango. Assigning different document types by Scopus for similar contents: an exploratory analysis. Scientometrics, 130(3):1989–2003, March 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05270-w.

Ellinger:2023:DPG

[Ell23] Jan Ellinger. Don’t put the greatest pressure on the weakest. Scientometrics, 128(10):5853–5857, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04816-0.

Esparza:2024:ICS

[ELR24] Luz Judith R. Esparza, Ángel Lee, and Carmen Rubio. Influence of cultural and socioeconomic factors on scientific production: a statistical analysis of the hindex. Scientometrics, 129(4):2083–2099, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04965-w.

Emmer:2019:CBI

[Emm19] Adam Emmer. The careers behind and the impact of solo author articles in Nature and Science. Scientometrics, 120 (2):825–840, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-019-03145-5.

El-Ouahi:2024:RFM

[EO24a] Jamal El-Ouahi. Research funding in the Middle East and North Africa: analyses of acknowledgments in scientific publications indexed in the Web of Science (2008– 2021). Scientometrics, 129(6):2933–2968, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04983-8.

El-Ouahi:2024:SRG

[EO24b] Jamal El-Ouahi. Scientometric rules as a guide to transform science systems in the Middle East and North Africa.

Scientometrics, 129(2):869–888, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04916-x.

El-Ouahi:2023:LWR

[EOL23] Jamal El-Ouahi and Vincent Larivière. On the lack of women researchers in the Middle East and North Africa. Scientometrics, 128(8):4321–4348, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04768-5.

Eleftheriou:2023:CAA

[EPP23] Konstantinos Eleftheriou, Patroklos Patsoulis, and Michael Polemis. Convergence among academic journals in accounting: a note. Scientometrics, 128(2):1055–1069, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04588-z.

Egghe:2021:IF

[ER21] Leo Egghe and Ronald Rousseau. The h-index formalism. Scientometrics, 126(7):6137–6145, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03699-9.

Egghe:2023:GIM

[ER23] Leo Egghe and Ronald Rousseau. Global impact measures. Scientometrics, 128(1):699–707, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04553-w.

Egghe:2024:SWP

[ER24] Leo Egghe and Ronald Rousseau. The small-world phenomenon: a model, explanations, characterizations, and examples. Scientometrics, 129(9):5599–5614, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05119-8.

Eika:2022:SJP

[ES22] Evelyn Eika and Frode Eika Sandnes. Starstruck by journal prestige and citation counts? On students’ bias and perceptions of trustworthiness according to clues in publication references. Scientometrics, 127(11):6363– 6390, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04521-4.

Eldakar:2023:BSA

[ES23] Metwaly Ali Mohamed Eldakar and Ahmed Maher Khafaga Shehata. A bibliometric study of article retractions in technology fields in developing economies countries. Scientometrics, 128(11):6047–6083, November 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04823-1.

Etzkowitz:2025:ASM

[Etz25] Henry Etzkowitz. Academic soulmate: a memory of Loet Leydesdorff. Scientometrics, 130(6):3267–3274, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05336-9.

Ebadi:2021:UTE

[EXT^{+}21] Ashkan Ebadi, Pengcheng Xi, Stéphane Tremblay, Bruce Spencer, Raman Pall, and Alexander Wong. Understanding the temporal evolution of COVID-19 research through machine learning and natural language processing. Scientometrics, 126(1):725–739, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03744-7;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03744-7.pdf.

Fages:2020:WBP

[Fag20] Diego Marino Fages. Write better, publish better. Scientometrics, 122(3):1671–1681, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03332-4;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03332-4.pdf.

Fang:2020:IJI

[Fan20] Hui Fang. Investigating the journal impact along the columns and rows of the publication-citation matrix. Scientometrics, 125(3):2265–2282, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03715-y.

Fang:2021:ANS

[Fan21] Hui Fang. Analysis of the new Scopus CiteScore. Scientometrics, 126(6):5321–5331, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03964-5.

Fang:2023:MCB

[Fan23] Hui Fang. A modification of citation-based journal indexes. Scientometrics, 128(2):1119–1132, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04594-1.

Fassin:2020:CHR

[Fas20a] Yves Fassin. Correction to: The HF-rating as a universal complement to the h-index. Scientometrics, 125(2): 991, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03771-4.

See [Fas20b].

Fassin:2020:HRU

[Fas20b] Yves Fassin. The HF-rating as a universal complement to the h-index. Scientometrics, 125(2):965–990, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03611-5. See correc-

tion [Fas20a].

Fassin:2021:DFT

[Fas21a] Yves Fassin. Does the Financial Times FT50 journal list select the best management and economics journals? Scientometrics, 126(7):5911–5943, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03988-x. See comment [Mou21a].

Fassin:2021:IRA

[Fas21b] Yves Fassin. The impact of review articles in management and economics journal rankings and metrics. Scientometrics, 126(12):9623–9632, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04160-1.

Fassin:2021:RCD

[Fas21c] Yves Fassin. Research on Covid-19: a disruptive phenomenon for bibliometrics. Scientometrics, 126(6):5305– 5319, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03989-w.

Fassin:2024:RBW

[Fas24] Yves Fassin. Research on bibliometrics and the World’s leading universities. Scientometrics, 129(11):6693–6718, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05141-w.

Feng:2025:ESL

[FAY25] Yu Feng, Wenkang An, and Zhen Yin. Enhancing scientific literature summarization via contrastive learning and chain-of-thought prompting. Scientometrics, 130(8): 4749–4772, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05397-w.

Forthmann:2025:RID

[FBDM25] Boris Forthmann, Marie Beisemann, Philipp Doebler, and Rüdiger Mutz. Reliable individual differences in researcher performance capacity estimates: evaluating productivity as explanatory variable. Scientometrics, 130(1):43–66, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05210-0.

Fortin:2021:DTH

[FBM21] Julie Fortin, Bjarne Bartlett, and Zia Mehrabi. Digital technology helps remove gender bias in academia. Scientometrics, 126(5):4073–4081, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03911-4.

Fang:2020:SAV

[FC20a] Zhichao Fang and Rodrigo Costas. Studying the accumulation velocity of altmetric data tracked by Altmetric.com. Scientometrics, 123(2):1077–1101, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03405-9;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03405-9.pdf.

Fernandes:2020:AOA

[FC20b] João M. Fernandes and Paulo Cortez. Alphabetic order of authors in scholarly publications: a bibliometric study for 27 scientific fields. Scientometrics, 125(3):2773– 2792, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03686-0.

Franceschet:2020:QHO

[FC20c] Massimo Franceschet and Giovanni Colavizza. Quantifying the higher-order influence of scientific publications. Scientometrics, 125(2):951–963, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03580-9.

Fernandez-Cano:2021:LEP

[FC21] Antonio Fernandez-Cano. Letter to the Editor: publish, publish … cursed! Scientometrics, 126(4):3673–3682, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03833-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03833-7.pdf.

Fernandes:2022:APC

[FCC22] João M. Fernandes, António Costa, and Paulo Cortez. Author placement in computer science: a study based on the careers of ACM Fellows. Scientometrics, 127(1):351–368, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04035-5.

Fang:2020:EAP

[FCT^{+}20] Zhichao Fang, Rodrigo Costas, Wencan Tian, Xianwen Wang, and Paul Wouters. An extensive analysis of the presence of altmetric data for Web of Science publications across subject fields and research topics. Scientometrics, 124(3):2519–2549, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03564-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03564-9.pdf.

Fang:2022:UES

[FCW22] Zhichao Fang, Rodrigo Costas, and Paul Wouters. User engagement with scholarly tweets of scientific papers: a large-scale and cross-disciplinary analysis. Scientometrics, 127(8):4523–4546, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04468-6.

Farber:2023:BSR

[FCY23] Michael Färber, Melissa Coutinho, and Shuzhou Yuan. Biases in scholarly recommender systems: impact, prevalence, and mitigation. Scientometrics, 128(5):2703–2736, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04636-2.

Forthmann:2021:RRC

[FD21] Boris Forthmann and Philipp Doebler. Reliability of researcher capacity estimates and count data dispersion: a comparison of Poisson, negative binomial, and Conway–Maxwell–Poisson models. Scientometrics, 126 (4):3337–3354, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-021-
03864-8;   http://link.springer.com/content/pdf/
10.1007/s11192-021-03864-8.pdf.

Forthmann:2024:WSB

[FDM24] Boris Forthmann, Philipp Doebler, and Rüdiger Mutz. Why summing up bibliometric indicators does not justify a composite indicator. Scientometrics, 129(12):7475– 7499, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05194-x.

Fernandez:2021:CCC

[Fer21] Viviana Fernandez. Cross-country concentration and specialization of mining inventions. Scientometrics, 126(8): 6715–6759, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04044-4.

Fernandez-Esquinas:2021:UQS

[FESRMB21] Manuel Fernández-Esquinas, María Isabel Sánchez-Rodríguez, and Rocío Muñoz-Benito. The use of QCA in science, technology and innovation studies: a review of the literature and an empirical application to knowledge transfer. Scientometrics, 126(8):6349–6382, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04012-y.

Feng:2024:CNA

[FFH^{+}24] Sida Feng, Lingzi Feng, Fang Han, Ye Zhang, Yanqing Ren, Lixue Wang, and Junpeng Yuan. Citation network analysis of retractions in molecular biology field. Scientometrics, 129(8):4795–4817, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05101-4.

Frandell:2021:EEA

[FFJ21] Ashlee Frandell, Mary K. Feeney, and Heyjie Jung. The effects of electronic alert letters for Internet surveys of academic scientists. Scientometrics, 126(8):7167–7181, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04029-3.

Fernandez:2021:OEP

[FFL21] Ana Fernández, Esther Ferrándiz, and M. Dolores León. Are organizational and economic proximity driving factors of scientific collaboration? Evidence from Spanish universities, 2001–2010. Scientometrics, 126(1):579–602, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03748-3.

Fantozzi:2025:URA

[FFN25] Paolo Fantozzi, Valerio Ficcadenti, and Maurizio Naldi. The university research assessment dilemma: a decision support system for the next evaluation campaigns. Scientometrics, 130(3):1617–1658, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05263-9.

Freijedo-Farinas:2024:BRD

[FFRRPR^{+}24] Fabián Freijedo-Farinas, Alberto Ruano-Ravina, Mónica Pérez-Ríos, Joseph Ross, and Cristina Candal-Pedreira. Biomedical retractions due to misconduct in Europe: characterization and trends in the last 20 years. Scientometrics, 129(5):2867–2882, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04992-7.

Frey:2020:STG

[FG20] Bruno S. Frey and Anthony Gullo. Sic transit gloria mundi: What remains of famous economists after their deaths? Scientometrics, 123(1):283–298, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03393-w.

Frietsch:2025:DHC

[FGB25] Rainer Frietsch, Sonia Gruber, and Lutz Bornmann. The definition of highly cited researchers: the effect of different approaches on the empirical outcome. Scientometrics, 130(2):881–907, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05158-1.

Fernandez-Guerrero:2020:LPS

[FGCGFC20] Inés M. Fernández-Guerrero, Zoraida Callejas, David Griol, and Antonio Fernández-Cano. Longitudinal patterns in Spanish doctoral theses on scientific medical information: a tertiary study. Scientometrics, 124(2):1241– 1260, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03494-6.

Falk:2021:WIA

[FH21] Martin Thomas Falk and Eva Hagsten. When international academic conferences go virtual. Scientometrics, 126(1):707–724, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03754-5;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03754-5.pdf.

Fuchs:2022:TDM

[FH22] Joel Emanuel Fuchs and Thomas Heinze. Two-dimensional mapping of university profiles in research. Scientometrics, 127(12):7215–7228, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04356-z.

Falk:2023:RAI

[FH23] Martin Thomas Falk and Eva Hagsten. Reverse adoption of information and communication technology among organisers of academic conferences. Scientometrics, 128(3): 1963–1985, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04616-y.

Filippin:2021:DMP

[Fil21] Flavia Filippin. Do main paths reflect technological trajectories? Applying main path analysis to the semiconductor manufacturing industry. Scientometrics, 126(8):

6443–6477, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04023-9.

Frandsen:2020:GGS

[FJO20] Tove Faber Frandsen, Rasmus Højbjerg Jacobsen, and Jakob Ousager. Gender gaps in scientific performance: a longitudinal matching study of health sciences researchers. Scientometrics, 124(2):1511–1527, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03528-z.

Feenstra:2022:PAB

[FLC22] Ramón A. Feenstra and Emilio Delgado López-Cózar. Philosophers’ appraisals of bibliometric indicators and their use in evaluation: from recognition to knee-jerk rejection. Scientometrics, 127(4):2085–2103, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04265-1.

Forthmann:2020:IQQ

[FLDD20] Boris Forthmann, Mark Leveling, Yixiao Dong, and Denis Dumas. Investigating the quantity-quality relationship in scientific creativity: an empirical examination of expected residual variance and the tilted funnel hypothesis. Scientometrics, 124(3):2497–2518, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03571-w;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03571-w.pdf.

Faria:2021:MIP

[FM21] João Ricardo Faria and Franklin G. Mixon, Jr. The marginal impact of a publication on citations, and its effect on academic pay. Scientometrics, 126(9):8217– 8226, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04073-z.

Fu:2022:EIR

[FMB22] Yuan Chih Fu, Marcelo Marques, and David P. Baker. An evolving international research collaboration network: spatial and thematic developments in co-authored higher education research, 1998–2018. Scientometrics, 127(3): 1403–1429, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04200-w.

Fabiano:2020:PPC

[FMF20] Gianluca Fabiano, Andrea Marcellusi, and Giampiero Favato. Public-private contribution to biopharmaceutical discoveries: a bibliometric analysis of biomedical research in UK. Scientometrics, 124(1):153–168, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03429-1;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03429-1.pdf.

Frandsen:2023:DUP

[FN23] Tove Faber Frandsen and Jeppe Nicolaisen. Defining the unscholarly publication: a bibliometric study of uncited and barely cited publications. Scientometrics, 128(2): 1337–1350, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04610-4.

Farooq:2023:AQS

[FNK23] Umar Farooq, Adeel Nasir, and Kanwal Iqbal Khan. An assessment of the quality of the search strategy: a case of bibliometric studies published in business and economics. Scientometrics, 128(8):4855–4874, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04765-8.

Fleerackers:2022:ISN

[FNM^{+}22] Alice Fleerackers, Lise Nehring, Lauren A. Maggio, Asura Enkhbayar, Laura Moorhead, and Juan Pablo Alperin. Identifying science in the news: an assessment of the precision and recall of Altmetric.com news mention data. Scientometrics, 127(11):6109–6123, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04510-7.

Foglia:2023:SSM

[Fog23] Francesco Foglia. Is smart specialisation monopolising the research on the EU cohesion policy? Evidence from a bibliometric analysis. Scientometrics, 128(2):1001– 1021, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04585-2.

Forthmann:2023:RCE

[For23] Boris Forthmann. Researcher capacity estimation based on the Q model: a generalized linear mixed model perspective. Scientometrics, 128(8):4753–4764, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04756-9.

Fernandez-Ramos:2023:USJ

[FRRBDD23] Andrés Fernández-Ramos, Blanca Rodríguez-Bravo, and Ángela Diez-Diez. Use of scientific journals in Spanish universities: analysis of the relationship between citations and downloads in two university library consortia. Scientometrics, 128(4):2489–2505, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04670-0.

Finocchi:2023:AIM

[FRS23] Irene Finocchi, Andrea Ribichini, and Marco Schaerf. An analysis of international mobility and research productivity in computer science. Scientometrics, 128(11):6147– 6175, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04837-9.

Ferrer-Sapena:2020:SDI

[FSEJF^{+}20] A. Ferrer-Sapena, E. Erdogan, E. Jiménez-Fernández, E. A. Sánchez-Pérez, and F. Peset. Self-defined information indices: application to the case of university rankings. Scientometrics, 124(3):2443–2456, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03575-6.

Fuchs:2021:MRI

[FSR21] Joel Emanuel Fuchs, Gunnar Sivertsen, and Ronald Rousseau. Measuring the relative intensity of collaboration within a network. Scientometrics, 126(10):8673–8682, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04110-x.

Feeley:2021:FHN

[FT21] Thomas Hugh Feeley and Frank Tutzauer. The faculty hiring network for PhD-granting communication programs. Scientometrics, 126(5):3983–4003, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03917-y.

Farber:2024:AIC

[FT24] Michael Färber and Lazaros Tampakis. Analyzing the impact of companies on AI research based on publications. Scientometrics, 129(1):31–63, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04867-3.

Fukugawa:2022:CEQ

[Fuk22a] Nobuya Fukugawa. Correction to: Effects of the quality of science on the initial public offering of university spinoffs: evidence from Japan. Scientometrics, 127(8):4457, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04474-8.

Fukugawa:2022:EQS

[Fuk22b] Nobuya Fukugawa. Effects of the quality of science on the initial public offering of university spinoffs: evidence from Japan. Scientometrics, 127(8):4439–4455, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04433-3.

Furnham:2020:WLM

[Fur20] Adrian Furnham. What I have learned from my Google Scholar and H index. Scientometrics, 122(2):1249– 1254, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03316-4.

Furnham:2021:PPR

[Fur21] Adrian Furnham. Publish or perish: rejection, scientometrics and academic success. Scientometrics, 126(1): 843–847, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03694-0.

Furnham:2023:PNS

[Fur23] Adrian Furnham. Peer nominations as scientometrics. Scientometrics, 128(2):1451–1458, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04597-y.

Flores-Vargas:2023:GCS

[FVGBPA23] Xochitl Flores-Vargas, Claudia Noemí González-Brambila, and Miguel Ángel Pérez-Angón. Geographical characterization of the scientific performance in Mexico (1995–2015). Scientometrics, 128(3):1785–1799, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04632-6.

Fathalla:2020:SEC

[FVLA20] Said Fathalla, Sahar Vahdati, Christoph Lange, and Sören Auer. Scholarly event characteristics in four fields of science: a metrics-based analysis. Scientometrics, 123(2):677–705, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03391-y;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03391-y.pdf.

Fan:2020:PTC

[FWDQ20] Lipeng Fan, Yuefen Wang, Shengchun Ding, and Binbin Qi. Productivity trends and citation impact of different institutional collaboration patterns at the research units’ level. Scientometrics, 125(2):1179–1196, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03609-z.

Formanowicz:2023:GBS

[FWH^{+}23] Magdalena Formanowicz, Marta Witkowska, Weronika Hryniszak, Zuzanna Jakubik, and Aleksandra Cisłak. Gender bias in special issues: evidence from a bibliometric analysis. Scientometrics, 128(4):2283–2299, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04639-z.

Feng:2020:OFR

[FYY20] Lingzi Feng, Junpeng Yuan, and Liying Yang. An observation framework for retracted publications in multiple dimensions. Scientometrics, 125(2):1445–1457, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03702-3.

Fan:2021:EBR

[FYY21] Xia Fan, Xiaowan Yang, and Zhou Yu. Effect of basic research and applied research on the universities’ innovation capabilities: the moderating role of private research funding. Scientometrics, 126(7):5387–5411, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03998-9.

Fang:2020:BHF

[FZC20] Li Fang, Xiaobei Zhou, and Lei Cui. Biclustering high-frequency MeSH terms based on the co-occurrence of distinct semantic types in a MeSH tree. Scientometrics, 124(2):1179–1190, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03496-4.

Feng:2021:IEI

[FZhCC21] Yongqi Feng, Haolin Zhang, Yung ho Chiu, and Tzu-Han Chang. Innovation efficiency and the impact of the institutional quality: a cross-country analysis using the two-stage meta-frontier dynamic network DEA model. Scientometrics, 126(4):3091–3129, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03829-3.

Feng:2020:AJE

[FZL^{+}20a] Lin Feng, Jian Zhou, Sheng-Lan Liu, Ning Cai, and Jie Yang. Analysis of journal evaluation indicators: an experimental study based on unsupervised Laplacian score. Scientometrics, 124(1):233–254, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03422-8. See correction [FZL^{+}20b].

Feng:2020:CAJ

[FZL^{+}20b] Lin Feng, Jian Zhou, Sheng-Lan Liu, Ning Cai, and Jie Yang. Correction to: Analysis of journal evaluation indicators: an experimental study based on unsupervised Laplacian score. Scientometrics, 124(3):2739–2740, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03504-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03504-7.pdf. See [FZL^{+}20a].

Gonzalez-Alcaide:2021:BSO

[GA21] Gregorio González-Alcaide. Bibliometric studies outside the information science and library science field: uncontainable or uncontrollable? Scientometrics, 126(8):6837– 6870, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04061-3.

Gomez-Aguayo:2024:SEK

[GAACBA24] Ana María Gómez-Aguayo, Joaquín M. Azagra-Caro, and Carlos Benito-Amat. The steady effect of knowledge cocreation with universities on business scientific impact throughout the economic cycle. Scientometrics, 129(5): 2771–2799, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04986-5.

Griffin:2021:CCA

[GAB21] Darrin J. Griffin, Zachary W. Arth, and San Bolkan. Collaborations in communication: Authorship credit allocation via a weighted fractional count procedure. Scientometrics, 126(5):4355–4372, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03927-w.

Gonzalez-Alcaide:2020:CPK

[GAPR20] Gregorio González-Alcaide, Héctor Pinargote, and José M. Ramos. From cut-points to key players in co-authorship networks: a case study in ventilator-associated pneumonia research. Scientometrics, 123(2):707–733, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03404-w.

Gonzalez-Brambila:2021:QQS

[GB21] Claudia N. Gonzalez-Brambila. Quantitative and qualitative studies of science and technology in Latin America. Scientometrics, 126(3):2411–2412, March 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03902-5.

Garg:2024:EIF

[GB24] Shanky Garg and Rashmi Bhardwaj. Exploring the influence of factors causing stress among doctoral students by combining fuzzy DEMATEL–ANP with a triangular approach. Scientometrics, 129(8):4695–4719, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05108-x.

Ghiasi:2021:WPC

[GBH21] Gita Ghiasi, Catherine Beaudry, and Matthew Harsh. Who profits from the Canadian nanotechnology reward system? Implications for gender-responsible innovation. Scientometrics, 126(9):7937–7991, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04022-w.

Guilarte:2021:RIT

[GBP21] Orlando Fonseca Guilarte, Simone Diniz Junqueira Barbosa, and Sinesio Pesco. RelPath: an interactive tool to visualize branches of studies and quantify the expertise of authors by citation paths. Scientometrics, 126(6): 4871–4897, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03959-2.

Glanzel:2020:BCS

[GC20] Wolfgang Glänzel and Pei-Shan Chi. The big challenge of Scientometrics 2.0: exploring the broader impact of scientific research in public health. Scientometrics, 125(2):1011–1031, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03473-x.

Giri:2021:RJT

[GC21] Rabishankar Giri and Sabuj Kumar Chaudhuri. Ranking journals through the lens of active visibility. Scientometrics, 126(3):2189–2208, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03850-6.

Gong:2022:PIC

[GC22] Kaile Gong and Ying Cheng. Patterns and impact of collaboration in China’s social sciences: cross-database comparisons between CSSCI and SSCI. Scientometrics, 127(10):5947–5964, October 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04483-7.

Garcia-Costa:2024:SLC

[GCGB^{+}24] Daniel García-Costa, Francisco Grimaldo, Giangiacomo Bravo, Bahar Mehmani, and Flaminio Squazzoni. The silver lining of COVID-19 restrictions: research output of academics under lockdown. Scientometrics, 129(3): 1771–1786, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04929-0.

Geng:2022:SWO

[GCH^{+}22] Yu Geng, Renmeng Cao, Xiaopu Han, Wencan Tian, Guangyao Zhang, and Xianwen Wang. Scientists are working overtime: when do scientists download scientific papers? Scientometrics, 127(11):6413–6429, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04524-1.

Gyorffy:2020:TGA

[GCHT20] Balázs Győrffy, Gyöngyi Csuka, Péter Herman, and Ádám Török. Is there a golden age in publication activity?-an analysis of age-related scholarly performance across all scientific disciplines. Scientometrics, 124(2):1081–1097, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03501-w;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03501-w.pdf.

Gayen:2025:CDS

[GCMJ25] Avijit Gayen, Somyajit Chakraborty, Saikat Mitra, and Angshuman Jana. Comeback or dropout: study of discontinued researchers at early career stage. Scientometrics, 130(2):1201–1236, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05243-z.

Giffoni:2025:LBL

[GCS25] Francesco Giffoni, Louis Colnot, and Emanuela Sirtori. The link between large scientific collaboration and productivity. Rethinking how to estimate the monetary value of publications. Scientometrics, 130(7):3743–3771, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05348-5.

Goh:2020:EHV

[GCT^{+}20] Yeow Chong Goh, Xin Qing Cai, Walter Theseira, Giovanni Ko, and Khiam Aik Khor. Evaluating human versus machine learning performance in classifying research abstracts. Scientometrics, 125(2):1197–1212, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03614-2;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03614-2.pdf.

Glanzel:2022:VAI

[GD22] Wolfgang Glänzel and Koenraad Debackere. Various aspects of interdisciplinarity in research and how to quantify and measure those. Scientometrics, 127(9):5551– 5569, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04133-4.

Gomez-Deniz:2024:FTN

[GDDG24] Emilio Gómez-Déniz and Pablo Dorta-González. A fieldand time-normalized Bayesian approach to measuring the impact of a publication. Scientometrics, 129(5):2659–2676, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04997-2.

Goyanes:2022:MPD

[GDdZ22] Manuel Goyanes, Márton Demeter, and Homero Gil de Zúñiga. Measuring publication diversity among the most productive scholars: how research trajectories differ in communication, psychology, and political science. Scientometrics, 127(6):3661–3682, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04386-7.

Goyanes:2025:GDF

[GDdZ25] Manuel Goyanes, Marton Demeter, and Homero Gil de Zúñiga. Gender disparities in first authorship: examining the Matilda effect across communication, political science, and sociology. Scientometrics, 130(5):2923–2946, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05303-4.

Goyanes:2020:MAA

[GDG^{+}20] Manuel Goyanes, Márton Demeter, Aurea Grané, Irene Albarrán-Lozano, and Homero Gil de Zúñiga. A mathematical approach to assess research diversity: operationalization and applicability in communication sciences, political science, and beyond. Scientometrics, 125(3):2299– 2322, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03680-6.

Goyanes:2023:RPC

[GDG^{+}23] Manuel Goyanes, Márton Demeter, Aurea Grané, Tamás Tóth, and Homero Gil de Zúñiga. Research patterns in communication (2009–2019): testing female representation and productivity differences, within the most cited authors and the field. Scientometrics, 128(1):137–156, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04575-4.

Goyanes:2020:AII

[GdM20] Manuel Goyanes and Luis de Marcos. Academic influence and invisible colleges through editorial board interlocking in communication sciences: a social network analysis of leading journals. Scientometrics, 123(2):791–811, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03401-z.

Grancay:2022:RCA

[GDM22] Martin Grancay, Tomás Dudás, and Ladislav Mura. Revealed comparative advantages in academic publishing of “old” and “new” European Union Member States 1998– 2018. Scientometrics, 127(3):1247–1271, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04251-z.

Goyanes:2024:AGD

[GdMDD24] Manuel Goyanes, Luis de Marcos, and Adrián Domínguez-Díaz. Automatic gender detection: a methodological procedure and recommendations to computationally infer the gender from names with ChatGPT and gender APIs. Scientometrics, 129(11):6867–6888, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05149-2.

Gracio:2020:DCA

[GdOCRM20] Maria Cláudia Cabrini Grácio, Ely Francina Tannuri de Oliveira, Zaida Chinchilla-Rodríguez, and Henk F. Moed. Does corresponding authorship influence scientific impact in collaboration: Brazilian institutions as a case of study. Scientometrics, 125(2):1349–1369, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03655-7.

Glanzel:2022:ICS

[GDR22] Wolfgang Glänzel, Koenraad Debackere, and Ronald Rousseau. The 18th International Conference on Scientometrics & Informetrics. Scientometrics, 127(12):6809– 6810, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04558-5.

Gracio:2023:CFM

[GDRPR23] Maria Cláudia Cabrini Grácio, Natalia Rodrigues Delbianco, Fábio Sampaio Rosas, and Antonio Perianes-Rodríguez. Co-follower metric on academic-social media ResearchGate: similarities between Derek de Solla Price Memorial Medal winners. Scientometrics, 128(10):5545– 5569, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04793-4.

Gomez-Espes:2024:BIC

[GEFJ24] Alberto Gómez-Espés, Michael Färber, and Adam Jatowt. Benefits of international collaboration in computer science: a case study of China, the European Union, and the United States. Scientometrics, 129(2):1155–1171, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04902-3.

Gerashchenko:2024:RTS

[Ger24] Daria Gerashchenko. Research topic switch and its relation to appointment as university leader. Scientometrics, 129(3):1841–1862, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04958-9.

Ghorbi:2021:RPI

[GFVK21] Ali Ghorbi, Mohsen Fazeli-Varzaneh, and Marcin Kozak. Retracted papers by Iranian authors: causes, journals, time lags, affiliations, collaborations. Scientometrics, 126(9):7351–7371, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04104-9.

Gunthe:2022:NIA

[GG22] Sachin S. Gunthe and Ravindra Gettu. A new index for assessing faculty research performance in higher educational institutions of emerging economies such as India. Scientometrics, 127(8):4959–4976, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04460-0.

Galli:2025:RDB

[GG25] Federica Galli and Fedele Greco. Ranking Departments based on research quality: a statistical evaluation of the ISPD indicator. Scientometrics, 130(2):705–725, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05240-2.

Gholampour:2022:RAO

[GGS22] Behzad Gholampour, Sajad Gholampour, and Ali Akbar Saboury. Retracted articles in oncology in the last three decades: frequency, reasons, and themes. Scientometrics, 127(4):1841–1865, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04305-w.

Goncalves:2020:CAD

[GGSA20] Paulo Henrique Santos Gonçalves, Thiago Gonçalves-Souza, and Ulysses Paulino Albuquerque. Chronic anthropogenic disturbances in ecology: a bibliometric approach. Scientometrics, 123(2):1103–1117, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03403-x.

Gonzalez-Hernandez:2021:MOD

[GHGMSM21] Isidro Jesús González-Hernández, Rafael Granillo-Macías, and Isaías Simón-Marmolejo. Marshall–Olkin distributions: a bibliometric study. Scientometrics, 126(11):9005– 9029, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04156-x.

Gao:2022:SET

[GHW22] Qiang Gao, Xiao Huang, and Jiang Wu. Semanticenhanced topic evolution analysis: a combination of the dynamic topic model and word2vec. Scientometrics, 127 (3):1543–1563, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04275-z.

Gao:2024:ETF

[GJ24] Qiang Gao and Man Jiang. Exploring technology fusion by combining latent Dirichlet allocation with Doc2vec: a case of digital medicine and machine learning. Scientometrics, 129(7):4043–4070, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05069-1.

Gundogan:2022:NHP

[GK22] Esra Gündogan and Mehmet Kaya. A novel hybrid paper recommendation system using deep learning. Scientometrics, 127(7):3837–3855, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04420-8.

Gundogan:2023:DLJ

[GKD23] Esra Gündogan, Mehmet Kaya, and Ali Daud. Deep learning for journal recommendation system of research papers. Scientometrics, 128(1):461–481, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04535-y.

Giannakakos:2025:IAC

[GKDB25] Vasiliki P. Giannakakos, Troy S. Karanfilian, Antonios D. Dimopoulos, and Anne Barmettler. Impact of author characteristics on outcomes of single- versus double-blind peer review: a systematic review of comparative studies in scientific abstracts and publications. Scientometrics, 130(1): 399–421, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05213-x.

Gokhberg:2023:SUR

[GKK23] Leonid Gokhberg, Tatiana Kuznetsova, and Maxim Kotsemir. From the Soviet Union to the Russian Federation: publication activity dynamics along the evolution of national science policies. Scientometrics, 128(11):6195– 6246, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04838-8.

Gerasimov:2024:CDC

[GKM^{+}24] Irina Gerasimov, Binita KC, Armin Mehrabian, James Acker, and Michael P. McGuire. Comparison of datasets citation coverage in Google Scholar, Web of Science, Scopus, Crossref, and DataCite. Scientometrics, 129(7):3681–3704, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05073-5.

Glanzel:2025:HK

[GKML25] Wolfgang Glänzel, Theo Kretschmer, Bernd Markscheffel, and Jean-Charles Lamirel. Hildrun Kretschmer (1947–2024). Scientometrics, 130(3):1327–1328, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05267-5.

Gallagher:2023:NPU

[GLK^{+}23] Colin Gallagher, Dean Lusher, Johan Koskinen, Bopha Roden, Peng Wang, Aaron Gosling, Anastasios Polyzos, Martina Stenzel, Sarah Hegarty, Thomas Spurling, and Gregory Simpson. Network patterns of universityindustry collaboration: a case study of the chemical sciences in Australia. Scientometrics, 128(8):4559–4588, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04749-8.

George:2020:FTB

[GLPC20] Susan George, Hiran H. Lathabai, Thara Prabhakaran, and Manoj Changat. A framework towards bias-free contextual productivity assessment. Scientometrics, 122(1): 127–157, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03286-7.

Goyanes:2025:UAI

[GLPN25] Manuel Goyanes, Carlos Lopezosa, and Valeriano Piñeiro-Naval. The use of artificial intelligence (AI) in research: a review of author guidelines in leading journals across eight social science disciplines. Scientometrics, 130(7): 3707–3723, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05377-0.

Grigoriev:2022:MAP

[GM22] Alexander Grigoriev and Olga Mondrus. Managing academic performance by optimal resource allocation. Scientometrics, 127(5):2433–2453, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04342-5.

Gou:2022:ECL

[GMB22] Zhenyu Gou, Fan Meng, and Yi Bu. Encoding the citation life-cycle: the operationalization of a literature-aging conceptual model. Scientometrics, 127(8):5027–5052, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04437-z.

Gutierrez-Maya:2021:EMS

[GMCRV21] Jazmín I. Gutiérrez-Maya, Francisco Collazo-Reyes, and Rodrigo A. Vega y Ortega Baez. The expansion of modern science through the Catalog of Scientific Papers, XIX century: the Latin American presence. Scientometrics, 126(3):2575–2593, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03606-2.

Gureyev:2020:RAP

[GMKG20] Vadim N. Gureyev, Nikolay A. Mazov, Denis V. Kosyakov, and Andrey E. Guskov. Review and analysis of publications on scientific mobility: assessment of influence, motivation, and trends. Scientometrics, 124(2):1599–1630, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03515-4.

Garcia:2025:DCF

[GMPRSFV25] J. A. García, J. J. Montero-Parodi, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Do competitive forces tend to correct choice errors in journal selection due to imperfect attention on the part of researchers? Scientometrics, 130(3):1421–1441, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05260-y.

Gomez:2024:DCR

[GMRVU24] Luis Fernando Gómez, Andrés Felipe Montoya-Rendón, and Juan Pablo Vélez-Uribe. Distribution by country, region, and publisher in environmental engineering journals in SCImago Journal and Country Rank database (1999–2022). Scientometrics, 129(9):5769–5782, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05121-0.

Gong:2023:IDC

[Gon23a] Kaile Gong. The influence of discipline consistency between papers and published journals on citations: an analysis of Chinese papers in three social science disciplines. Scientometrics, 128(5):3129–3146, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04686-6.

Gong:2023:TYC

[Gon23b] Kaile Gong. Twenty years of Chinese social sciences towards internationalization (1998-2017): a knowledge sources perspective. Scientometrics, 128(12):6373–6402, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04859-3.

Gordon:2021:CRW

[Gor21] Avishag Gordon. Crowdsourcing and its relationship to wisdom of the crowd and insight building: a bibliometric study. Scientometrics, 126(5):4373–4382, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03932-z.

Galvagno:2021:BGF

[GP21] Marco Galvagno and Vincenzo Pisano. Building the genealogy of family business internationalization: a bibliometric mixed-method approach. Scientometrics, 126(1): 757–783, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03755-4.

Greiner-Petter:2020:MWE

[GPYR^{+}20] André Greiner-Petter, Abdou Youssef, Terry Ruas, Bruce R. Miller, Moritz Schubotz, Akiko Aizawa, and Bela Gipp. Math-word embedding in math search and semantic extraction. Scientometrics, 125(3):3017–3046, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03502-9;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03502-9.pdf.

Gray:2020:SWO

[Gra20] Russell J. Gray. Sorry, we’re open: Golden open-access and inequality in non-human biological sciences. Scientometrics, 124(2):1663–1675, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03540-3.

Garechana:2022:TDS

[GRBAM22] Gaizka Garechana, Rosa Río-Belver, and Izaskun Alvarez-Meaza. TeknoAssistant: a domain specific tech mining approach for technical problem-solving support. Scientometrics, 127(9):5459–5473, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04280-2.

Garcia:2020:ARG

[GRSFV20a] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The author–reviewer game. Scientometrics, 124(3):2409–2431, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03559-6.

Garcia:2020:CBP

[GRSFV20b] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Confirmatory bias in peer review. Scientometrics, 123(1):517–533, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03357-0.

Garcia:2021:EMG

[GRSFV21a] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The editor–manuscript game. Scientometrics, 126(5):4277–4295, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
03918-x.

Garcia:2021:IBR

[GRSFV21b] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The interplay between the reviewer’s incentives and the journal’s quality standard. Scientometrics, 126(4): 3041–3061, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03839-1.

Garcia:2021:QCP

[GRSFV21c] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Quality censoring in peer review. Scientometrics, 126(1):825–830, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03693-1.

Garcia:2022:CPM

[GRSFV22] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Can a paid model for peer review be sustainable when the author can decide whether to pay or not? Scientometrics, 127(3):1491–1514, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04248-8.

Galban-Rodriguez:2021:MQA

[GRTPAJ21] Ernesto Galbán-Rodríguez, Deborah Torres-Ponjuán, and Ricardo Arencibia-Jorge. Multidimensional quantitative analysis of the Cuban scientific output and its regional context. Scientometrics, 126(3):2643–2665, March 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03904-3.

Georges:2022:MIV

[GS22a] Patrick Georges and Aylin Seckin. Music information visualization and classical composers discovery: an application of network graphs, multidimensional scaling, and support vector machines. Scientometrics, 127(5):2277–2311, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04331-8.

Glanzel:2022:O

[GS22b] Wolfgang Glänzel and András Schubert. Obituary: [Tibor Braun (1932–2022)]. Scientometrics, 127(11):6061– 6062, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04556-7.

Gupta:2024:DCD

[GS24] Solanki Gupta and Vivek Kumar Singh. Distributional characteristics of dimensions concepts: An empirical analysis using Zipf’s law. Scientometrics, 129(2):1037–1053, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04899-9.

Gupta:2024:ADQ

[GSB24] Solanki Gupta, Vivek Kumar Singh, and Sumit Kumar Banshal. Altmetric data quality analysis using Benford’s law. Scientometrics, 129(7):4597–4621, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05061-9.

Galli:2022:TAI

[GSG22a] Carlo Galli, Roberto Sala, and Stefano Guizzardi. Tamquam alter idem: formal similarities in a subset of reports on anti-inflammatory compounds in the years 2008– 2019. Scientometrics, 127(7):3879–3910, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04434-2.

Grabowska:2022:IIH

[GSG22b] Sandra Grabowska, Sebastian Saniuk, and Bozena Gajdzik. Industry 5.0: improving humanization and sustainability of Industry 4.0. Scientometrics, 127(6): 3117–3144, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04370-1.

Gilling:2024:PCG

[GSH24] Russyl Gilling, Marissa Scandlyn, and Blair Hesp. Prevalence and characteristics of graphical abstracts in a specialist pharmacology journal. Scientometrics, 129(12):7881– 7888, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05146-5.

Gandal:2021:SPN

[GSOCS21] Neil Gandal, Michal Shur-Ofry, Michael Crystal, and Royee Shilony. Out of sight: patents that have never been cited. Scientometrics, 126(4):2903–2929, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03849-z.

Garcia-Suaza:2020:PEC

[GSOW20] Andrés García-Suaza, Jesús Otero, and Rainer Winkelmann. Predicting early career productivity of PhD economists: Does advisor-match matter? Scientometrics, 122(1):429–449, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03277-8;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03277-8.pdf.

Giglio:2021:CCL

[GSP21] Carlo Giglio, Roberto Sbragia, and Roberto Palmieri. Cross-country learning from patents: an analysis of citations flows in innovation trajectories. Scientometrics, 126(9):7917–7936, September 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04094-8.

Gonzalez-Salmon:2024:CTG

[GSRG24] Elvira González-Salmón and Nicolas Robinson-Garcia. A call for transparency in gender assignment approaches. Scientometrics, 129(4):2451–2454, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04995-4.

Gruber:2023:JMA

[GTB23] Anton Gruber, Alexander Tekles, and Lutz Bornmann. John Mearsheimer’s academic roots: a reference publication year spectroscopy of a political scientist’s oeuvre. Scientometrics, 128(7):3867–3877, July 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04721-6.

Goyanes:2024:GDG

[GTH24] Manuel Goyanes, Tamás Tóth, and Gergő Háló. Gender differences in Google Scholar representation and impact: an empirical analysis of political communication, journalism, health communication, and media psychology. Scientometrics, 129(3):1719–1737, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04945-0.

Gu:2021:DBR

[Gu21] Jiafeng Gu. Determinants of biopharmaceutical R&D expenditures in China: the impact of spatiotemporal context. Scientometrics, 126(8):6659–6680, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04058-y.

Guba:2024:WDS

[Gub24] Katerina Guba. Why do sociologists on academic periphery willingly support bibliometric indicators? Scientometrics, 129(1):497–518, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04890-4.

Gorraiz:2022:MEC

[GUG^{+}22] Juan Gorraiz, Ursula Ulrych, Wolfgang Glänzel, Wenceslao Arroyo-Machado, and Daniel Torres-Salinas. Measuring the excellence contribution at the journal level: an alternative to Garfield’s impact factor. Scientometrics, 127(12):7229–7251, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04295-9.

Gusenbauer:2022:SWY

[Gus22] Michael Gusenbauer. Search where you will find most: Comparing the disciplinary coverage of 56 bibliographic databases. Scientometrics, 127(5):2683–2745, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04289-7.

Ghosal:2024:DMT

[GVK24] Tirthankar Ghosal, Kamal Kaushik Varanasi, and Valia Kordoni. A deep multi-tasking approach leveraging on cited–citing paper relationship for citation intent classification. Scientometrics, 129(2):767–783, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04811-5.

George:2020:MAB

[GVS20] Benu George, Pradeep Varathan, and T. V. Suchithra. Meta-analysis on big data of bioactive compounds from mangrove ecosystem to treat neurodegenerative disease. Scientometrics, 122(3):1539–1561, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03355-2.

Ghaffari:2023:MRL

[GW23] Fatemeh Ghaffari and Mark C. Wilson. A model for reference list length of scholarly articles. Scientometrics, 128(9):5335–5350, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04780-9.

Gao:2024:PQA

[GWLY24] Xingyu Gao, Qiang Wu, Yuanyuan Liu, and Ruilu Yang. Pasteur’s quadrant in AI: do patent-cited papers have higher scientific impact? Scientometrics, 129(2):909– 932, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04925-w.

Grace:2021:HTA

[GWvZ21] S. G. Grace, F. S. S. Wiepking, and A. A. J. van Zundert. Hot topics in anaesthesia: a bibliometric analysis of five high-impact journals from 2010–2019. Scientometrics, 126(10):8749–8759, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04129-0.

Guo:2022:ALA

[GX22] Ying Guo and Xiantao Xiao. Author-level altmetrics for the evaluation of Chinese scholars. Scientometrics, 127(2): 973–990, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04228-y.

Gao:2025:UEE

[GXC25] Xing Gao, Senmao Xia, and Mengqiu Cao. The underexplored effects of economic transition on intellectual property rights protection: An economic geography perspective. Scientometrics, 130(7):??, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05352-9.

Guo:2021:MDP

[GYWZ21] Min Guo, Naiding Yang, Jingbei Wang, and Yanlu Zhang. Multi-dimensional proximity and network stability: the moderating role of network cohesion. Scientometrics, 126 (4):3471–3499, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-021-03882-6.

Ginther:2024:DTW

[GZOC24] Donna K. Ginther, Carlos Zambrana, Patricia Oslund, and Wan-Ying Chang. Do two wrongs make a right? Measuring the effect of publications on science careers. Scientometrics, 129(1):289–320, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04849-5.

Gagolewski:2022:FTD

[GZSC22a] Marek Gagolewski, Barbara Zogała-Siudem, and Anna Cena. Fairness in the three-dimensional model for citation impact. Scientometrics, 127(10):6055–6059, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04497-1.

Gagolewski:2022:OIC

[GZSC22b] Marek Gagolewski, Barbara Zogała-Siudem, and Anna Cena. Ockham’s index of citation impact. Scientometrics, 127(5):2829–2845, May 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04345-2. See [Pra22].

Halevi:2020:SLJ

[Hal20a] Gali Halevi. The scientific legacy of Judit Bar-Ilan. Scientometrics, 123(3):1201–1209, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03439-z.

Haley:2020:CWU

[Hal20b] M. Ryan Haley. Combining the weighted and unweighted Euclidean indices: a graphical approach. Scientometrics, 123(1):103–111, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03368-x.

Han:2025:FIO

[HAL25] Ruilian Han, Lu An, and Gang Li. Field identification and opportunity discovery of photovoltaics technology: deep transfer learning method. Scientometrics, 130(8): 4249–4281, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05394-z.

Han:2020:ERT

[Han20] Xiaoyao Han. Evolution of research topics in LIS between 1996 and 2019: an analysis based on latent Dirichlet allocation topic model. Scientometrics, 125(3):2561–2595, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03721-0;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03721-0.pdf.

Hao:2024:TRA

[Hao24] Jialiang Hao. Titles in research articles and doctoral dissertations: cross-disciplinary and cross-generic perspectives. Scientometrics, 129(4):2285–2307, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04941-4.

Hadad:2024:PSI

[HAR24] Shlomit Hadad, Noa Aharony, and Daphne R. Raban. Policy shaping the impact of open-access publications: a longitudinal assessment. Scientometrics, 129(1):237–260, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04875-3.

Hassan:2020:TCS

[HAS^{+}20] Saeed-Ul Hassan, Naif R. Aljohani, Mudassir Shabbir, Umair Ali, Sehrish Iqbal, Raheem Sarwar, Eugenio Martínez-Cámara, Sebastián Ventura, and Francisco Herrera. Tweet Coupling: a social media methodology for clustering scientific publications. Scientometrics, 124(2): 973–991, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03499-1.

Harikandeh:2023:EAA

[HAT23] Seyyed Reza Taher Harikandeh, Sadegh Aliakbary, and Soroush Taheri. An embedding approach for analyzing the evolution of research topics with a case study on computer science subdomains. Scientometrics, 128(3):1567–1582, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04642-4.

Haghani:2020:CPU

[HB20] Milad Haghani and Michiel C. J. Bliemer. Covid-19 pandemic and the unprecedented mobilisation of scholarly efforts prompted by a health crisis: Scientometric comparisons across SARS, MERS and 2019-nCoV literature. Scientometrics, 125(3):2695–2726, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03706-z;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03706-z.pdf.

Hasan:2021:ALC

[HB21a] Syed Hasan and Robert Breunig. Article length and citation outcomes. Scientometrics, 126(9):7583–7608, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04083-x.

Haunschild:2021:CTU

[HB21b] Robin Haunschild and Lutz Bornmann. Can tweets be used to detect problems early with scientific papers? A case study of three retracted COVID-19/SARS-CoV-2 papers. Scientometrics, 126(6):5181–5199, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03962-7.

Haunschild:2022:RPY

[HB22a] Robin Haunschild and Lutz Bornmann. Reference publication year spectroscopy (RPYS) in practice: a software tutorial. Scientometrics, 127(12):7253–7271, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04369-8.

Haunschild:2022:RDT

[HB22b] Robin Haunschild and Lutz Bornmann. Relevance of document types in the scores’ calculation of a specific fieldnormalized indicator: Are the scores strongly dependent on or nearly independent of the document type handling? Scientometrics, 127(8):4419–4438, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04446-y.

Haslam:2023:SES

[HB23] Nick Haslam and Naomi Baes. Scientific eminence and scientific hierarchy: bibliometric prediction of fellowship in the Australian Academy of Science. Scientometrics, 128(12):6659–6674, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04870-8.

Hohmann:2025:ESW

[HBC25] Mollie Hawkes Hohmann, Adrian G. Barnett, and Sean D. Connell. The evolution of scientific writing: an analysis of 20 million abstracts over 70 years in health and medical science. Scientometrics, 130(7):3313–3347, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05353-8.

Hosseini:2021:MIS

[HBN^{+}21] Seyedmohammadreza Hosseini, Hamed Baziyad, Rasoul Norouzi, Sheida Jabbedari Khiabani, Győző Gidófalvi, Amir Albadvi, Abbas Alimohammadi, and Seyedehsan Seyedabrishami. Mapping the intellectual structure of GIS-T field (2008–2019): a dynamic co-word analysis. Scientometrics, 126(4):2667–2688, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03840-8;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03840-8.pdf.

Hida:2020:IGR

[HBO^{+}20] Rahma M. Hida, John C. Begeny, Helen O. Oluokun, Taylor E. Bancroft, Felicia L. Fields-Turner, Brodie D. Ford, Cecily K. Jones, Chynna B. Ratliff, and Andrykah Y. Smith. Internationalization and geographically representative scholarship in journals devoted to behavior analysis: an assessment of 10 journals across 15 years. Scientometrics, 122(1):719–740, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03289-4.

Helmer:2020:WMR

[HBP20] Sven Helmer, David B. Blumenthal, and Kathrin Paschen. What is meaningful research and how should we measure it? Scientometrics, 125(1):153–169, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03649-5;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03649-5.pdf.

Horney:2024:GIE

[HBSW24] Jennifer A. Horney, Adam Bitunguramye, Shazia Shaukat, and Zackery White. Gender and the h-index in epidemiology. Scientometrics, 129(7):3725–3733, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05083-3.

Harsh:2021:MCS

[HBW^{+}21] Matthew Harsh, Ravtosh Bal, Alex Weryha, Justin Whatley, Charles C. Onu, and Lisa M. Negro. Mapping computer science research in Africa: using academic networking sites for assessing research activity. Scientometrics, 126(1):305–334, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03727-8.

Hasumi:2022:OME

[HC22] Toshiyuki Hasumi and Mei-Shiu Chiu. Online mathematics education as bio-eco-techno process: bibliometric analysis using co-authorship and bibliographic coupling. Scientometrics, 127(8):4631–4654, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04441-3.

Huang:2022:ITE

[HCL22] Lu Huang, Xiang Chen, and Jiarun Liu. Identification of topic evolution: network analytics with piecewise linear representation and word embedding. Scientometrics, 127(9):5353–5383, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04273-1.

Huang:2021:DNA

[HCN21] Lu Huang, Xiang Chen, and Xingxing Ni. Dynamic network analytics for recommending scientific collaborators.

Scientometrics, 126(11):8789–8814, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04164-x.

Hajibayova:2021:EIC

[HCS21] Lala Hajibayova, L. P. Coladangelo, and Heather A. Soyka. Exploring the invisible college of citizen science: questions, methods and contributions. Scientometrics, 126(8): 6989–7003, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04050-6.

Hu:2025:IIS

[HCS25] Xiaotong Hu, Ziwei Chen, and Jin Shi. Identification of important software based on software dependency graph. Scientometrics, 130(7):3929–3959, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05366-3.

Huang:2022:MII

[HCZ^{+}22] Lu Huang, Yijie Cai, Erdong Zhao, Shengting Zhang, Yue Shu, and Jiao Fan. Measuring the interdisciplinarity of Information and Library Science interactions using citation analysis and semantic analysis. Scientometrics, 127(11):6733–6761, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04401-x.

Hu:2020:SBE

[HCZEP20] Zhiwen Hu, Yiping Cui, Jian Zhang, and Jacqueline Eviston-Putsch. Shalosh B. Ekhad: a computer credit for mathematicians. Scientometrics, 122(1):71–97, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03305-7.

Hoppen:2023:DBW

[HdSV23] Natascha Helena Franz Hoppen and Samile Andréa de Souza Vanz. The development of Brazilian women’s and gender studies: a bibliometric diagnosis. Scientometrics, 128(1):227–261, January 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04545-w.

Huang:2023:AAE

[HDY^{+}23] Tian-Yuan Huang, Liangping Ding, Yong-Qiang Yu, Lei Huang, and Liying Yang. From AR5 to AR6: exploring research advancement in climate change based on scientific evidence from IPCC WGI reports. Scientometrics, 128(9):5227–5245, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04788-1.

Heneberg:2019:THP

[Hen19] Petr Heneberg. The troubles of high-profile open access megajournals. Scientometrics, 120(2):733–746, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03144-6. See correc-

tion [Hen20].

Heneberg:2020:CTH

[Hen20] Petr Heneberg. Correction to: The troubles of highprofile open access megajournals. Scientometrics, 123 (2):1169–1171, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-019-
03281-y;   http://link.springer.com/content/pdf/
10.1007/s11192-019-03281-y.pdf.

Heinze:2022:NOP

[HF22] Thomas Heinze and Joel Emanuel Fuchs. National and organizational patterns of Nobel laureate careers in physiology/medicine, physics, and chemistry. Scientometrics, 127(12):7273–7288, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04250-0.

Hoffmeyer:2023:UIC

[HFAR23] Bodil Hoffmeyer, Siv Fonnes, Kristoffer Andresen, and Jacob Rosenberg. Use of inactive Cochrane reviews in academia: a citation analysis. Scientometrics, 128(5): 2923–2934, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04691-9.

Horta:2022:HHE

[HFS22] Hugo Horta, Shihui Feng, and João M. Santos. Homophily in higher education research: a perspective based on co-authorships. Scientometrics, 127(1):523–543, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04227-z.

Ho:2020:RHC

[HG20] Long Ho and Peter Goethals. Research hotspots and current challenges of lakes and reservoirs: a bibliometric analysis. Scientometrics, 124(1):603–631, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03453-1.

Hallinger:2020:PIL

[HGB20] Philip Hallinger, Sedat Gümüs, and Mehmet Sükrü Bellibas. ’Are principals instructional leaders yet?’ A science map of the knowledge base on instructional leadership, 1940–2018. Scientometrics, 122(3):1629–1650, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03360-5.

Hernandez-Gonzalez:2020:SDT

[HGDRM20] V. Hernández-González, A. De Pano-Rodríguez, and J. Reverter-Masia. Spanish doctoral theses in physical activity and sports sciences and authors’ scientific publications (LUSTRUM 2013–2017). Scientometrics, 122(1): 661–679, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03295-6.

Hernandez-Garcia:2021:FIS

[HGM21] Yoscelina Iraida Hernandez-García and Mónica Anzaldo Montoya. Flow of ideas in the study of communication channels and references in publications on nanotechnology applied to food and agriculture in Mexico. Scientometrics, 126(2):995–1017, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03793-y.

Hilario:2023:AOI

[HGMÁW23] Carla Mara Hilário, Maria Cláudia Cabrini Grácio, Daniel Martínez-Ávila, and Dietmar Wolfram. Authorship order as an indicator of similarity between article discourse and author citation identity in informetrics. Scientometrics, 128(10):5389–5410, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04791-6.

Huang:2021:TDM

[HGZ21] Ying Huang, Wolfgang Glänzel, and Lin Zhang. Tracing the development of mapping knowledge domains. Scientometrics, 126(7):6201–6224, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03821-x.

Haghighat:2021:HHH

[HH21] Mansour Haghighat and Javad Hayatdavoudi. How hot are hot papers? The issue of prolificacy and self-citation stacking. Scientometrics, 126(1):565–578, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03749-2.

Holmberg:2020:DAO

[HHB^{+}20] Kim Holmberg, Juha Hedman, Timothy D. Bowman, Fereshteh Didegah, and Mikael Laakso. Do articles in open access journals have more frequent altmetric activity than articles in subscription-based journals? An investigation of the research output of Finnish universities. Scientometrics, 122(1):645–659, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03301-x;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03301-x.pdf.

Hsiao:2020:DRS

[HhC20] Tsung-Ming Hsiao and Kuang hua Chen. The dynamics of research subfields for library and information science: an investigation based on word bibliographic coupling. Scientometrics, 125(1):717–737, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03645-9.

Hauer:2020:QAA

[HHKZ20] Marc P. Hauer, Xavier C. R. Hofmann, Tobias D. Krafft, and Katharina A. Zweig. Quantitative analysis of automatic performance evaluation systems based on the hindex. Scientometrics, 123(2):735–751, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03407-7;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03407-7.pdf.

Huang:2022:ECS

[HHY22] Ruhua Huang, Yuting Huang, and Wei Yu. Exploring the characteristics of special issues: distribution, topicality, and citation impact. Scientometrics, 127(9):5233– 5256, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04384-9.

Hassan:2020:IAI

[HIA^{+}20] Saeed-Ul Hassan, Sehrish Iqbal, Naif R. Aljohani, Salem Alelyani, and Alesia Zuccala. Introducing the ‘alt-index’ for measuring the social visibility of scientific research. Scientometrics, 123(3):1407–1419, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03447-z.

Haruna:2020:RPR

[HIQ^{+}20] Khalid Haruna, Maizatul Akmar Ismail, Atika Qazi, Habeebah Adamu Kakudi, Mohammed Hassan, Sanah Abdullahi Muaz, and Haruna Chiroma. Research paper recommender system based on public contextual metadata. Scientometrics, 125(1):101–114, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03642-y.

Hirv:2022:ISR

[Hir22] Tanel Hirv. The interplay of the size of the research system, ways of collaboration, level, and method of funding in determining bibliometric outputs. Scientometrics, 127(3): 1295–1316, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04232-2.

Htoo:2023:WMR

[HJCT23] Tint Hla Hla Htoo, Na Jin-Cheon, and Michael Thelwall. Why are medical research articles tweeted? The news value perspective. Scientometrics, 128(1):207–226, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04578-1.

Hain:2023:MLA

[HJLZ23] Daniel Hain, Roman Jurowetzki, Sungjoo Lee, and Yuan Zhou. Machine learning and artificial intelligence for science, technology, innovation mapping and forecasting: Review, synthesis, and applications. Scientometrics, 128(3): 1465–1472, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04628-8.

Hobert:2021:OAU

[HJT21] Anne Hobert, Najko Jahn, and Niels Taubert. Open access uptake in Germany 2010–2018: adoption in a diverse research landscape. Scientometrics, 126(12):9751–9777, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04002-0.

Howoldt:2023:RPT

[HKN23] David Howoldt, Henning Kroll, and Peter Neuhäusler. Relating popularity on Twitter and LinkedIn to bibliometric indicators of visibility and interconnectedness: an analysis of 8512 applied researchers in Germany. Scientometrics, 128(10):5571–5594, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04799-y.

Howoldt:2023:URT

[HKNF23] David Howoldt, Henning Kroll, Peter Neuhäusler, and Alexander Feidenheimer. Understanding researchers’ Twitter uptake, activity and popularity — an analysis of applied research in Germany. Scientometrics, 128(1): 325–344, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04569-2.

Homolak:2020:PAC

[HKV20] J. Homolak, I. Kodvanj, and D. Virag. Preliminary analysis of COVID-19 academic information patterns: a call for open science in the times of closed borders. Scientometrics, 124(3):2687–2701, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03587-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03587-2.pdf.

Hrazdil:2024:SBC

[HKW24] Karel Hrazdil, Jeong-Bon Kim, and Frederick H. Willeboordse. Systematic bias in citing practices: evidence from accounting journal additions to the FT list. Scientometrics, 129(11):6947–6973, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05168-z.

Ho:2021:RCP

[HL21a] Jonathan C. Ho and Demei Lee. Research commercialisation performance in different types of universities: case from Taiwan. Scientometrics, 126(10):8617–8634, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04109-4.

Ho:2021:SKD

[HL21b] Mei Hsiu-Ching Ho and John S. Liu. The swift knowledge development path of COVID-19 research: the first 150 days. Scientometrics, 126(3):2391–2399, March 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03835-5.

Hsiao:2023:KMI

[HL23] Yung-Chang Hsiao and Jun-You Lin. Knowledge management and innovation: evidence of international joint venture. Scientometrics, 128(1):87–113, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04562-9.

Huckstadt:2024:IFP

[HL24] Malte Hückstädt and Luca M. Leisten. Internal factors promoting research collaboration problems: an inputprocess-output analysis. Scientometrics, 129(4):2007– 2035, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04957-w.

Hladchenko:2025:UUQ

[Hla25] Myroslava Hladchenko. Ukrainian universities in QS World University Rankings: when the means become ends. Scientometrics, 130(2):969–997, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05165-2.

Held:2021:CVT

[HLG21] Matthias Held, Grit Laudel, and Jochen Gläser. Challenges to the validity of topic reconstruction. Scientometrics, 126(5):4511–4536, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03920-3.

Hu:2021:MRT

[HLR21] Xiaojun Hu, Xian Li, and Ronald Rousseau. Mathematical reflections on Triple Helix calculations. Scientometrics, 126(10):8581–8587, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04114-7.

Habicht:2021:HHC

[HLS21] Isabel M. Habicht, Mark Lutter, and Martin Schröder. How human capital, universities of excellence, third party funding, mobility and gender explain productivity in German political science. Scientometrics, 126(12):9649–9675, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04175-8.

Hsu:2023:TTA

[HLS23] Gen-Chang Hsu, Wei-Jiun Lin, and Syuan-Jyun Sun. Temporal trends in academic performance and career duration of principal investigators in ecology and evolutionary biology in Taiwan. Scientometrics, 128(6):3437–3451, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04710-9.

Huang:2020:IIS

[HLSX20] Yongwen Huang, Jiao Li, Tan Sun, and Guojian Xian. Institution information specification and correlation based on institutional PIDs and IND tool. Scientometrics, 122 (1):381–396, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-019-
03268-9.

Hu:2025:FRO

[HLT25] Guangyuan Hu, Hongxu Liu, and Li Tang. Flexible recruitment of overseas talent. Scientometrics, 130(8):4369– 4395, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05379-y.

Hou:2023:ABS

[HLZ23] Jianhua Hou, Hao Li, and Yang Zhang. Altmetricsbased sleeping beauties: necessity or just a supplement? Scientometrics, 128(10):5477–5506, October 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04798-z.

Haunschild:2020:DSW

[HM20a] Robin Haunschild and Werner Marx. Discovering seminal works with marker papers. Scientometrics, 125(3): 2955–2969, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03358-z;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03358-z.pdf.

Hou:2020:HHI

[HM20b] Jianhua Hou and Da Ma. How the high-impact papers formed? A study using data from social media and citation. Scientometrics, 125(3):2597–2615, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03703-2.

Hladchenko:2021:NOU

[HM21] Myroslava Hladchenko and Henk F. Moed. National orientation of Ukrainian journals: means-ends decoupling in a semi-peripheral state. Scientometrics, 126(3):2365–2389, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03844-4.

Haley:2023:FFM

[HM23] M. Ryan Haley and M. Kevin McGee. A flexible functional method for jointly valuing journal visibility and author citation count. Scientometrics, 128(6):3337–3346, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04727-0.

Heibi:2024:OID

[HMPS24] Ivan Heibi, Arianna Moretti, Silvio Peroni, and Marta Soricetti. The OpenCitations Index: description of a database providing open citation data. Scientometrics, 129(12):7923–7942, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05160-7.

Haunschild:2024:HCR

[HMW24] Robin Haunschild, Werner Marx, and Jürgen Weis. How can revivals of scientific publications be explained using bibliometric methods? A case study discovering booster papers for the 1985 Physics Nobel Prize paper. Scientometrics, 129(2):1079–1095, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04906-z.

Hu:2021:AEE

[HMZ21] Feng Hu, Lin Ma, and Zi-Ke Zhang. The aging effect in evolving scientific citation networks. Scientometrics, 126 (5):4297–4309, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03929-8.

Huang:2024:OAR

[HNM^{+}24] Chun-Kai Huang, Cameron Neylon, Lucy Montgomery, Richard Hosking, James P. Diprose, Rebecca N. Handcock, and Katie Wilson. Open access research outputs receive more diverse citations. Scientometrics, 129(2): 825–845, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04894-0.

Hamid:2021:IBB

[HNO^{+}21] Habiba Hamid, Rafidah Md Noor, Syaril Nizam Omar, Ismail Ahmedy, Shaik Shabana Anjum, Syed Adeel Ali Shah, Sheena Kaur, Fazidah Othman, and Emran Mohd Tamil. IoT-based botnet attacks systematic mapping study of literature. Scientometrics, 126(4):2759–2800, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03819-5.

Haustein:2020:CJB

[HP20] Stefanie Haustein and Isabella Peters. Commemorating Judit Bar-Ilan from bibliometric and altmetric perspectives. Scientometrics, 123(3):1211–1224, June 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03448-y.

Heibi:2021:QQA

[HP21] Ivan Heibi and Silvio Peroni. A qualitative and quantitative analysis of open citations to retracted articles: the Wakefield 1998 et al.’s case. Scientometrics, 126(10): 8433–8470, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04097-5.

Hognogi:2024:ESO

[HP24] Gheorghe-Gavrila Hognogi and Ana-Maria Pop. Something old, new, and borrowed. Rise of the systematic reviews. Scientometrics, 129(9):5797–5803, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05133-w.

Heyard:2021:ICE

[HPH21] Rachel Heyard, Tobias Philipp, and Hanna Hottenrott. Imaginary carrot or effective fertiliser? A rejoinder on funding and productivity. Scientometrics, 126(11):9333– 9338, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04130-7.

See reply [MHD21b].

Hancean:2021:CNM

[HPL21] Marian-Gabriel Hâncean, Matjaz Perc, and Jürgen Lerner. The coauthorship networks of the most productive European researchers. Scientometrics, 126(1):201–224, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03746-5.

Hajikhani:2022:MSD

[HS22a] Arash Hajikhani and Arho Suominen. Mapping the sustainable development goals (SDGs) in science, technology and innovation: application of machine learning in SDGoriented artefact detection. Scientometrics, 127(11):6661– 6693, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04358-x.

Hammami:2022:SST

[HS22b] Asma Hammami and Nabil Semmar. The simplex simulation as a tool to reveal publication strategies and citation factors. Scientometrics, 127(1):319–350, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04198-1.

Hajibabaei:2023:WKP

[HSE23] Anahita Hajibabaei, Andrea Schiffauerova, and Ashkan Ebadi. Women and key positions in scientific collaboration networks: analyzing central scientists’ profiles in the artificial intelligence ecosystem through a gender lens. Scientometrics, 128(2):1219–1240, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04601-5.

Hu:2020:MRC

[HTGW20] Zhigang Hu, Wencan Tian, Jiacheng Guo, and Xianwen Wang. Mapping research collaborations in different countries and regions: 1980–2019. Scientometrics, 124(1): 729–745, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03484-8.

He:2023:VSF

[HTX23] Ying He, Kun Tian, and Xiaoran Xu. A validation study on the factors affecting the practice modes of open peer review. Scientometrics, 128(1):587–607, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04552-x.

Hou:2024:NTL

[HTZ24] Jianhua Hou, Shiqi Tang, and Yang Zhang. A novel technology life cycle analysis method based on LSTM and CRF. Scientometrics, 129(3):1173–1196, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04946-z.

Huang:2024:SPR

[HTZR24] Zhenye Huang, Deyou Tang, Rong Zhao, and Wenjing Rao. A scientific paper recommendation method using the time decay heterogeneous graph. Scientometrics, 129(3): 1589–1613, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04933-4.

Huang:2024:DDI

[Hua24] Jianan Huang. Diagnosing the declining industry sponsorship in clinical research. Scientometrics, 129(1):663–679, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04887-z.

Huckstadt:2022:CBF

[Hüc22] Malte Hückstädt. Coopetition between frenemies — interrelations and effects of seven collaboration problems in research clusters. Scientometrics, 127(9):5191–5224, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04472-w.

Huckstadt:2023:TRW

[Hüc23] Malte Hückstädt. Ten reasons why research collaborations succeed — a random forest approach. Scientometrics, 128(3):1923–1950, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04629-7.

Hug:2024:HDR

[Hug24] Sven E. Hug. How do referees integrate evaluation criteria into their overall judgment? Evidence from grant peer review. Scientometrics, 129(3):1231–1253, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04915-y.

Haghani:2021:TEM

[HV21] Milad Haghani and Pegah Varamini. Temporal evolution, most influential studies and sleeping beauties of the coronavirus literature. Scientometrics, 126(8):7005–7050, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04036-4.

Herrera-Vallejera:2021:PPR

[HVGP21] Darlenis Herrera-Vallejera and Salvador Gorbea-Portal. Pharmacology and pharmacy: research and innovation analysis. Scientometrics, 126(3):2513–2522, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03568-5.

Hu:2022:IGL

[HWC22] Dayu Hu, Chengyuan Wang, and Xiaoyu Cui. Investigating the genealogy of the literature on digital pathology: a two-dimensional bibliometric approach. Scientometrics, 127(2):785–801, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04224-2.

Huang:2023:INR

[HWL23] Xiaoling Huang, Lei Wang, and Weishu Liu. Identification of national research output using Scopus/Web of Science Core Collection: a revisit and further investigation. Scientometrics, 128(4):2337–2347, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04649-x.

Hu:2020:WIE

[HWNL20] Guangyuan Hu, Lei Wang, Rong Ni, and Weishu Liu. Which h-index? An exploration within the Web of Science. Scientometrics, 123(3):1225–1233, June 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03425-5.

Huang:2022:BDD

[HWP22] Ying Huang, Xuefeng Wang, and Alan L. Porter. “Big data” driven tech mining and ST&I management: an introduction. Scientometrics, 127(9):5227–5231, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04507-2.

Huang:2024:RIG

[HWW^{+}24] Xuejian Huang, Zhibin Wu, Gensheng Wang, Zhipeng Li, Yuansheng Luo, and Xiaofang Wu. ResGAT: an improved graph neural network based on multi-head attention mechanism and residual network for paper classification. Scientometrics, 129(2):1015–1036, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04898-w.

Hou:2022:DEP

[HWX22] Li Hou, Qiang Wu, and Yundong Xie. Does early publishing in top journals really predict long-term scientific success in the business field? Scientometrics, 127(11):6083– 6107, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04509-0.

He:2021:CRL

[HWZ21] Chaocheng He, Jiang Wu, and Qingpeng Zhang. Characterizing research leadership on geographically weighted collaboration network. Scientometrics, 126(5):4005–4037, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03943-w.

Huang:2025:ERC

[HX25] Tian-Yuan Huang and Jie Xue. Examining the role of cofirst authorship in scientific collaboration: a quantitative study. Scientometrics, 130(4):2357–2377, April 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05296-0.

Hou:2020:UPN

[HY20] Jianhua Hou and Jiantao Ye. Are uncited papers necessarily all nonimpact papers? A quantitative analysis. Scientometrics, 124(2):1631–1662, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03539-w.

Huang:2022:SII

[HY22] Tian-Yuan Huang and Liying Yang. Superior identification index: Quantifying the capability of academic journals to recognize good research. Scientometrics, 127(7): 4023–4043, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04372-z.

Hou:2023:WPF

[HYSY23] Jianhua Hou, Xiucai Yang, Haoyang Song, and Haiyue Yao. Will patent family be dormant? Research on the identification and characteristics of Sleeping Beauty’s patent family. Scientometrics, 128(10):5361–5387, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04784-5.

He:2025:MAC

[HYY25] Guoxiu He, Jia Yuan, and Yunhan Yang. Measuring academic cocoon from disparity and diversity perspectives. Scientometrics, 130(4):2445–2474, April 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05290-6.

Hou:2023:ESM

[HYZ23] Jianhua Hou, Xiucai Yang, and Yang Zhang. The effect of social media knowledge cascade: an analysis of scientific papers diffusion. Scientometrics, 128(9):5169– 5195, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04785-4.

Hu:2021:STN

[HZ21] Tao Hu and Yin Zhang. A spatial-temporal network analysis of patent transfers from U.S. universities to firms. Scientometrics, 126(1):27–54, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03745-6.

Hessler:2022:EBR

[HZ22] Nicole Heßler and Andreas Ziegler. Evidence-based recommendations for increasing the citation frequency of original articles. Scientometrics, 127(6):3367–3381, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04378-7.

Hou:2021:HDP

[HZC21] Jianhua Hou, Bili Zheng, and Chaomei Chen. How do Price medalists’ scholarly impact change before and after their awards? Scientometrics, 126(7):5945–5981, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03979-y.

Huang:2023:TLA

[HZCM23] Xiaoming Huang, Peihu Zhu, Yuwen Chen, and Jian Ma. A transfer learning approach to interdisciplinary document classification with keyword-based explanation. Scientometrics, 128(12):6449–6469, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04825-z.

Huang:2022:ABS

[HZJ22] Zhihong Huang, Qianjin Zong, and Xuerui Ji. The associations between scientific collaborations of LIS research and its policy impact. Scientometrics, 127(11):6453– 6470, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04532-1.

Huang:2022:ESI

[HZW22] Heng Huang, Donghua Zhu, and Xuefeng Wang. Evaluating scientific impact of publications: combining citation polarity and purpose. Scientometrics, 127(9):5257– 5281, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04183-8.

Ibrahim:2021:SMU

[Ibr21] Bahaa Ibrahim. Statistical methods used in Arabic journals of library and information science. Scientometrics, 126 (5):4383–4416, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03913-2.

Ioannidis:2024:EPE

[ICB24] John P. A. Ioannidis, Thomas A. Collins, and Jeroen Baas. Evolving patterns of extreme publishing behavior across science. Scientometrics, 129(9):5783–5796, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05117-w.

Iqbal:2021:DTC

[IHB21] Sehrish Iqbal, Saeed-Ul Hassan, and Lutz Bornmann. A decade of in-text citation analysis based on natural language processing and machine learning techniques: an overview of empirical studies. Scientometrics, 126(8): 6551–6599, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04055-1.

Ingwersen:2021:DJC

[IHLP21] Peter Ingwersen, Soeren Holm, Birger Larsen, and Thomas Ploug. Do journals and corporate sponsors back certain views in topics where disagreement prevails? Scientometrics, 126(1):389–415, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03743-8;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03743-8.pdf.

Ioannidis:2023:PNR

[Ioa23] John P. A. Ioannidis. Prolific non-research authors in high impact scientific journals: meta-research study. Scientometrics, 128(5):3171–3184, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04687-5.

Iwami:2020:BAF

[IOWN20] Shino Iwami, Arto Ojala, Chihiro Watanabe, and Pekka Neittaanmäki. A bibliometric approach to finding fields that co-evolved with information technology. Scientometrics, 122(1):3–21, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03284-9;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03284-9.pdf.

Ilgisonis:2022:HCT

[IPP22] Ekaterina V. Ilgisonis, Mikhail A. Pyatnitskiy, and Elena A. Ponomarenko. How to catch trends using MeSH terms analysis? Scientometrics, 127(4):1953–1967, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04292-y.

Ihsan:2021:NBC

[IQ21] Imran Ihsan and M. Abdul Qadir. An NLP-based citation reason analysis using CCRO. Scientometrics, 126(6): 4769–4791, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03955-6.

Ivanova:2025:OTM

[Iva25] Inga Ivanova. Operationalization of the theory of meaning in inter-social communications and its applications. Scientometrics, 130(6):3089–3108, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05305-2.

Ioannoni:2020:DCR

[IVCE20] Vanessa Ioannoni, Tommaso Vitale, Corrado Costa, and Iris Elliott. Depicting communities of Romani studies: on the who, when and where of Roma related scientific publications. Scientometrics, 122(3):1473–1490, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03352-5.

Jamali:2023:GGA

[JA23] Hamid R. Jamali and Alireza Abbasi. Gender gaps in Australian research publishing, citation and co-authorship. Scientometrics, 128(5):2879–2893, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04685-7.

Jablonsky:2025:ACI

[Jab25] Josef Jablonský. Analysis of citation impact of ORMS journals by DEA models. Scientometrics, 130(1):343–365, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05205-x.

Jimenez:2020:APC

[JADG20] Sergio Jimenez, Youlin Avila, George Dueñas, and Alexander Gelbukh. Automatic prediction of citability of scientific articles by stylometry of their titles and abstracts. Scientometrics, 125(3):3187–3232, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03526-1.

Jamali:2024:CNJ

[Jam24] Hamid R. Jamali. Country names in journal titles: shaping researchers’ perception of journals quality. Scientometrics, 129(2):803–823, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04904-1.

Jamali:2025:LGF

[Jam25] Hamid R. Jamali. Local or global? Factors influencing authorship composition of Australian journals. Scientometrics, 130(7):4073–4091, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05360-9.

Jung:2022:CAC

[JC22] Jung-Kyu Jung and Jae Young Choi. Choice and allocation characteristics of faculty time in Korea: effects of tenure, research performance, and external shock. Scientometrics, 127(5):2847–2869, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04320-x.

Jiang:2023:CSW

[JC23] Xiaorui Jiang and Jingqiang Chen. Contextualised segment-wise citation function classification. Scientometrics, 128(9):5117–5158, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04778-3.

Jang:2022:KPP

[JCK22] Byeongdeuk Jang, Jae-Yong Choung, and Inje Kang. Knowledge production patterns of China and the US: quantum technology. Scientometrics, 127(10):5691–5719, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04478-4.

Jin:2022:EFP

[JCX22] Qianqian Jin, Hongshu Chen, and Fei Xiong. Exploring funding patterns with word embedding-enhanced organization-topic networks: a case study on big data. Scientometrics, 127(9):5415–5440, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04253-x.

Jiang:2022:EPI

[JCZ22] Lidan Jiang, Jingyan Chen, and Fang Zou. Exploring the patterns of international technology diffusion in AI from the perspective of patent citations. Scientometrics, 127(9):5307–5323, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04134-3.

Janko:2020:SDA

[JDG^{+}20] Ferenc Jankó, Áron Drüszler, Borbála Gálos, Norbert Móricz, Judit Papp-Vancsó, Ildikó Pieczka, Rita Pongrácz, Ervin Rasztovits, Zsuzsanna Soósné Dezső, and Orsolya Szabó. Sources of doubt: actors, forums, and language of climate change skepticism. Scientometrics, 124(3):2251–2277, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03552-z;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03552-z.pdf.

Jin:2021:DRA

[JDG21] Tan Jin, Huiqiong Duan, and Kai Guo. Do research articles with more readable abstracts receive higher online attention? Evidence from Science. Scientometrics, 126(10): 8471–8490, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04112-9.

Junior:2020:ICS

[JDSL20] Geraldo J. Pessoa Junior, Thiago M. R. Dias, Thiago H. P. Silva, and Alberto H. F. Laender. On interdisciplinary collaborations in scientific coauthorship networks: the case of the Brazilian community. Scientometrics, 124(3):2341– 2360, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03605-3.

Dong:2021:MFA

[jDXjS21] Zuo jun Dong, Lan Xu, and Guo jun Sun. Major factors affecting biomedical cross-city R&D collaborations based on cooperative patents in China. Scientometrics, 126(3):

1923–1943, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03828-4.

Jebari:2021:UCC

[JHVC21] Chaker Jebari, Enrique Herrera-Viedma, and Manuel Jesus Cobo. The use of citation context to detect the evolution of research topics: a large-scale analysis. Scientometrics, 126(4):2971–2989, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03858-y.

Jebari:2023:CAC

[JHVC23] Chaker Jebari, Enrique Herrera-Viedma, and Manuel Jesus Cobo. Context-aware citation recommendation of scientific papers: comparative study, gaps and trends. Scientometrics, 128(8):4243–4268, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04773-8.

Jonsson:2022:SCC

[JI22] Oskar Jonsson and Susanne Iwarsson. Scrutinizing the collaboration criterion in research: how do policy ambitions play out in proposals and assessments? Scientometrics, 127(8):4675–4696, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04428-0.

Jiang:2021:UAP

[Jia21] Shan Jiang. Understanding authors’ psychological reactions to peer reviews: a text mining approach. Scientometrics, 126(7):6085–6103, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04032-8.

Jiang:2025:EAC

[Jia25] Xiaorui Jiang. Ensembling approaches to citation function classification and important citation screening. Scientometrics, 130(3):1371–1419, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05265-7.

Jiang:2023:MDM

[JJ23] Gianna Kexin Jiang and Yajun Jiang. More diversity, more complexity, but more flexibility: research article titles in TESOL Quarterly, 1967–2022. Scientometrics, 128 (7):3959–3980, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04738-x.

Jeong:2020:CAC

[JJPC20] Chanwoo Jeong, Sion Jang, Eunjeong Park, and Sungchul Choi. A context-aware citation recommendation model with BERT and graph convolutional networks. Scientometrics, 124(3):1907–1922, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03561-y.

Jeong:2021:DRA

[JJY21] Yujin Jeong, Hyejin Jang, and Byungun Yoon. Developing a risk-adaptive technology roadmap using a Bayesian network and topic modeling under deep uncertainty. Scientometrics, 126(5):3697–3722, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03945-8.

Jakab:2024:HMA

[JKK24] Martin Jakab, Eva Kittl, and Tobias Kiesslich. How many authors are (too) many? A retrospective, descriptive analysis of authorship in biomedical publications. Scientometrics, 129(3):1299–1328, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04928-1.

Jiang:2022:RSI

[JKY22] Lu Jiang, Xinyu Kang, and Bo Yang. A refinement strategy for identification of scientific software from bioinformatics publications. Scientometrics, 127(6):3293–3316, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04381-y.

Jiang:2020:NWO

[JL20] Fan Jiang and Nian Cai Liu. New wine in old bottles? Examining institutional hierarchy in laureate mobility networks, 1900–2017. Scientometrics, 125(2):1291– 1304, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03477-7.

Jiang:2023:CER

[JL23] Yuyan Jiang and Xueli Liu. A construction and empirical research of the journal disruption index based on open citation data. Scientometrics, 128(7):3935–3958, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04737-y.

Jeong:2025:MPD

[JL25] Dawoon Jeong and Jeong-Dong Lee. Mapping product development trajectories: product citation network. Scientometrics, 130(8):4621–4665, August 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05395-y.

Jack:2021:SKP

[JLL21] Pablo Jack, Jeremias Lachman, and Andrés López. Scientific knowledge production and economic catching-up: an empirical analysis. Scientometrics, 126(6):4565–4587, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03973-4.

Janavi:2020:MDS

[JME20] Elmira Janavi, Mohammad Javad Mansourzadeh, and Mojgan Samandar Ali Eshtehardi. A methodology for developing scientific diversification strategy of countries. Scientometrics, 125(3):2229–2264, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03685-1.

Junqueira:2020:AFS

[JMP20] Nara Tadini Junqueira, Luiz Fernando Magnago, and Paulo Santos Pompeu. Assessing fish sampling effort in studies of Brazilian streams. Scientometrics, 123(2): 841–860, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03418-4.

Jiang:2023:TTA

[JMZ^{+}23] Chi Jiang, Xiao Ma, Jiangfeng Zeng, Yin Zhang, Tingting Yang, and Qiumiao Deng. TAPRec: time-aware paper recommendation via the modeling of researchers’ dynamic preferences. Scientometrics, 128(6):3453–3471, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04731-4.

Jokic:2020:PVA

[Jok20] Maja Jokić. Productivity, visibility, authorship, and collaboration in library and information science journals: Central and Eastern European authors. Scientometrics, 122(2):1189–1219, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03308-4.

Joanis:2022:FAG

[JP22] Steven T. Joanis and Vivek H. Patil. First-author gender differentials in business journal publishing: top journals versus the rest. Scientometrics, 127(2):733–761, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04235-z.

Jang:2021:IET

[JPS21] Wooseok Jang, Yongtae Park, and Hyeonju Seol. Identifying emerging technologies using expert opinions on the future: a topic modeling and fuzzy clustering approach. Scientometrics, 126(8):6505–6532, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04024-8.

Jesenko:2021:EWS

[JS21] Berndt Jesenko and Christian Schlögl. The effect of web of science subject categories on clustering: the case of data-driven methods in business and economic sciences. Scientometrics, 126(8):6785–6801, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04060-4.

Jovanovic:2022:TTH

[JSLJ22] Milica Jovanović, Gordana Savić, and Maja Levi-Jaksić. Towards a triple helix based efficiency index of innovation systems. Scientometrics, 127(5):2577–2609, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04304-x.

Joaquin:2021:LAS

[JT21] Jeremiah Joven Joaquin and Raymond R. Tan. The lost art of short communications in academia. Scientometrics, 126(12):9633–9637, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04192-7.

Joaquin:2023:WIJ

[JTB23] Jeremiah Joven Joaquin, Raymond R. Tan, and Hazel T. Biana. So, what if a journal is both at the ‘top’ and ‘bottom’: reply to Mason and Singh. Scientometrics, 128(10):5859–5863, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04809-z. See [MS22d].

Jonker:2022:SIU

[JVY22] Hans Jonker, Florian Vanlee, and Walter Ysebaert. Societal impact of university research in the written press: media attention in the context of SIUR and the open science agenda among social scientists in Flanders, Belgium.

Scientometrics, 127(12):7289–7306, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04374-x.

Jing:2023:CEA

[JXY^{+}23a] Song Jing, Pengxin Xie, Qun Yin, Qingzhao Ma, Celestine Chinedu Ogbu, Xia Guo, Daniel M. J. J. Stanley, and Leuta Philatelic Tutaia. Correction: The effect of academic mobility on research performance: the case of China. Scientometrics, 128(10):5851, October 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04832-0.

Jing:2023:EAM

[JXY^{+}23b] Song Jing, Pengxin Xie, Qun Yin, Qingzhao Ma, Celestine Chinedu Ogbu, Xia Guo, Daniel M. J. J. Stanley, and Leuta Philatelic Tutaia. The effect of academic mobility on research performance: the case of China. Scientometrics, 128(10):5829–5850, October 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04814-2.

Jia:2023:DRG

[JXZW23] Pengfei Jia, Weixi Xie, Guangyao Zhang, and Xianwen Wang. Do reviewers get their deserved acknowledgments from the authors of manuscripts? Scientometrics, 128(10): 5687–5703, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04790-7.

Jin:2021:TAE

[JYT21] Yinyu Jin, Sha Yuan, and Jie Tang. Turing Award elites revisited: patterns of productivity, collaboration, authorship and impact. Scientometrics, 126(3):2329–2348, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03860-4.

Kinne:2020:WMI

[KA20] Jan Kinne and Janna Axenbeck. Web mining for innovation ecosystem mapping: a framework and a large-scale pilot study. Scientometrics, 125(3):2011–2041, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03726-9;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03726-9.pdf.

Kanwal:2024:RPR

[KA24] Tayyaba Kanwal and Tehmina Amjad. Research paper recommendation system based on multiple features from citation network. Scientometrics, 129(9):5493–5531, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05109-w.

Kajikawa:2022:REE

[Kaj22] Yuya Kajikawa. Reframing evidence in evidence-based policy making and role of bibliometrics: toward transdisciplinary scientometric research. Scientometrics, 127(9): 5571–5585, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04325-6.

Kalmukov:2020:AAA

[Kal20] Yordan Kalmukov. An algorithm for automatic assignment of reviewers to papers. Scientometrics, 124(3):1811– 1850, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03519-0.

Kavalerchik:2025:NIC

[Kav25] Anne Kavalerchik. Novelty and interdisciplinarity in criminology: how data-drivenness connects both. Scientometrics, 130(2):665–678, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05224-8.

Kumar:2023:BSA

[KBDS23] Dhananjay Kumar, Plaban Kumar Bhowmick, Sumana Dey, and Debarshi Kumar Sanyal. On the banks of Shodhganga: analysis of the academic genealogy graph of an Indian ETD repository. Scientometrics, 128(7):3879–3914, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04728-z.

Karmakar:2020:DPS

[KBS20] Mousumi Karmakar, Sumit Kumar Banshal, and Vivek Kumar Singh. Does presence of social media plugins in a journal website result in higher social media attention of its research publications? Scientometrics, 124(3):2103– 2143, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03574-7.

Karmakar:2021:LSC

[KBS21] Mousumi Karmakar, Sumit Kumar Banshal, and Vivek Kumar Singh. A large-scale comparison of coverage and mentions captured by the two altmetric aggregators: Altmetric.com and PlumX. Scientometrics, 126(5):4465– 4489, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03941-y.

Kifor:2023:IDR

[KBSS23] Claudiu Vasile Kifor, Ana Maria Benedek, Ioan Sîrbu, and Roxana Florenţa Săvescu. Institutional drivers of research productivity: a canonical multivariate analysis of Romanian public universities. Scientometrics, 128(4):2233–2258, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04655-z.

Kiesslich:2021:CIJ

[KBZT21] Tobias Kiesslich, Marlena Beyreis, Georg Zimmermann, and Andreas Traweger. Citation inequality and the journal impact factor: median, mean, (does it) matter? Scientometrics, 126(2):1249–1269, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03812-y;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03812-y.pdf.

Kuo:2021:QAD

[KCC21] Shu-Chun Kuo, Tsair-Wei Chien, and Willy Chou. Questions to the article: demonstrating the ascendancy of COVID-19 research using acronyms. Scientometrics, 126 (10):8761–8764, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04108-5.

Kuo:2022:SAM

[KCC22] Shu-Chun Kuo, Tsair-Wei Chien, and Willy Chou. Suggestions to the article: Medical professionalism research characteristics and hotspots. Scientometrics, 127(2):1191– 1194, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04221-5.

See [SJW21] and response [SJW22].

Kuo:2023:QAG

[KCC23] Shu-Chun Kuo, Tsair-Wei Chien, and Willy Chou. Questions to the article: Generating clustered journal maps: an automated system for hierarchical classification. Scientometrics, 128(4):2615–2617, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04086-8.

Kuc-Czarnecka:2025:RRA

[KCS25] Marta Kuc-Czarnecka and Andrea Saltelli. Ranking the rankers. an analysis of science-wide author databases of standardised citation indicators. Scientometrics, 130(3): 1497–1517, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05253-x.

Kataeva:2023:EGR

[KDIR23] Zumrad Kataeva, Naureen Durrani, Zhanna Izekenova, and Aray Rakhimzhanova. Evolution of gender research in the social sciences in post-Soviet countries: a bibliometric analysis. Scientometrics, 128(3):1639–1666, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04619-9.

Kotsemir:2022:MYR

[KDN22] Maxim Kotsemir, Ekaterina Dyachenko, and Alena Nefedova. Mobile young researchers and their non-mobile ‘twins’: who is winning the academic race? Scientometrics, 127(12):7307–7332, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04488-2.

Kamrani:2021:DRK

[KDS21] Pantea Kamrani, Isabelle Dorsch, and Wolfgang G. Stock. Do researchers know what the h-index is? And how do they estimate its importance? Scientometrics, 126(7): 5489–5508, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03968-1.

Kozlowski:2021:SRS

[KDZ21] Diego Kozlowski, Jennifer Dusdal, and Andreas Zilian. Semantic and relational spaces in science of science: deep learning models for article vectorisation. Scientometrics, 126(7):5881–5910, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03984-1.

Ke:2023:IRT

[Ke23] Qing Ke. Interdisciplinary research and technological impact: evidence from biomedicine. Scientometrics, 128(4): 2035–2077, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04662-0.

Kosztyan:2021:ICM

[KFCK21] Zsolt Tibor Kosztyán, Beáta Fehérvölgyi, Tibor Csizmadia, and Kinga Kerekes. Investigating collaborative and mobility networks: reflections on the core missions of universities. Scientometrics, 126(4):3551–3564, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03865-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
021-03865-7.pdf.

Kiss:2020:LDP

[KFLS20] Anna Kiss, Péter Fritz, Zoltán Lakner, and Sándor Soós. Linking the dimensions of policy-related research on obesity: a hybrid mapping with multicluster topics and interdisciplinarity maps. Scientometrics, 122 (1):159–213, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-019-
03293-8;   http://link.springer.com/content/pdf/
10.1007/s11192-019-03293-8.pdf.

Kacem:2020:TSC

[KFM20] Ameni Kacem, Justin W. Flatt, and Philipp Mayr. Tracking self-citations in academic publishing. Scientometrics, 123(2):1157–1165, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03413-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03413-9.pdf.

Kwon:2020:IPI

[KG20] Uijun Kwon and Youngjung Geum. Identification of promising inventions considering the quality of knowledge accumulation: a machine learning approach. Scientometrics, 125(3):1877–1897, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03710-3.

Kosyakov:2022:RCC

[KG22] Denis Kosyakov and Andrey Guskov. Reasons and consequences of changes in Russian research assessment policies. Scientometrics, 127(8):4609–4630, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04469-5.

Kosyakov:2024:DIS

[KG24] Denis Kosyakov and Andrey Guskov. Disciplinary and institutional shifts: decomposing deviations in the countrylevel proportions of conference papers in Scopus. Scientometrics, 129(3):1697–1717, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04943-2.

Kastrin:2021:SAK

[KH21] Andrej Kastrin and Dimitar Hristovski. Scientometric analysis and knowledge mapping of literature-based discovery (1986–2020). Scientometrics, 126(2):1415–1451, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03811-z.

Kosztyan:2025:CVE

[KHC^{+}25] Zsolt T. Kosztyán, Frigyes Hausz, Tibor Csizmadia, Attila Katona, István Szabó, and Beáta Fehérvölgyi. Concentration versus excellence: lessons learned of European R&D&I framework programs. Scientometrics, 130(2):937– 967, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05220-y.

Kulczycki:2021:CPB

[KHK21] Emanuel Kulczycki, Marek Hołowiecki, and Franciszek Krawczyk. Citation patterns between impact-factor and questionable journals. Scientometrics, 126(10):8541–8560, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04121-8.

Kodvanj:2022:PCP

[KHT22] Ivan Kodvanj, Jan Homolak, and Vladimir Trkulja. Publishing of COVID-19 preprints in peer-reviewed journals, preprinting trends, public discussion and quality issues. Scientometrics, 127(3):1339–1352, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04249-7.

Kim:2023:TIG

[Kim23] Yong Jin Kim. Technological innovation in GRIs, universities, and the private sector: evidence from the chemical technology network in south Korea. Scientometrics, 128(11):5929–5948, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04820-4.

Kim:2024:AAU

[Kim24] Eungi Kim. Analyzing AI use policy in LIS: association with journal metrics and publisher volume. Scientometrics, 129(12):7623–7644, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05189-8.

Kim:2023:TRH

[KJL23] Seung Hwan Kim, Bogang Jun, and Jeong-Dong Lee. Technological relatedness: how do firms diversify their technology? Scientometrics, 128(9):4901–4931, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04775-6.

Ko:2024:UPR

[KJY24] Byoung-Kwon Ko, Yeongkyun Jang, and Jae-Suk Yang. Universalism and particularism in the recommendations of the Nobel Prize for science. Scientometrics, 129(2): 847–868, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04921-0.

Kim:2020:CEI

[KK20a] Eungi Kim and Eun Sil Kim. A critical examination of international research conducted by North Korean authors: Increasing trends of collaborative research between China and North Korea. Scientometrics, 124(1):429–450, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03461-1.

Kulczycki:2020:RPM

[KK20b] Emanuel Kulczycki and Przemysław Korytkowski. Researchers publishing monographs are more productive and more local-oriented. Scientometrics, 125(2):1371– 1387, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03376-x;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03376-x.pdf.

Klein:2023:STT

[KK23] Alexandra-Maria Klein and Nina Kranke. Some thoughts on transparency of the data and analysis behind the Highly Cited Researchers list. Scientometrics, 128(12):6773– 6780, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04852-w.

Kobayashi:2025:LAH

[KKK^{+}25] Yoshihiko Kobayashi, Kuriko Kudo, Toshiya Kobayashi, Hiroko Kinoshita, HyunJung Bang, Hiroshi Ito, Akihiro Kishimura, Yusuke Matsumoto, Masato Miwa, Motoko Unoki, and Tamaki Yoshioka. Low awareness but high willingness to engage in science communication: a crossdisciplinary survey study in a Japanese university. Scientometrics, 130(3):1357–1370, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05251-z.

Kang:2020:NRB

[KKOL20] Namyong Kang, Jeong-Jae Kim, Byung-Won On, and Ingyu Lee. A node resistance-based probability model for resolving duplicate named entities. Scientometrics, 124(3):1721–1743, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03585-4.

Kardos:2023:HSR

[KKPJ23] Peter Kardos, Ádám Kun, Csaba Pléh, and Ferenc Jordán. (How) should researchers publicize their research papers before peer review? Scientometrics, 128(3):2019–2023, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04646-0.

Kulczycki:2021:MEL

[KKS21] Emanuel Kulczycki, Władysław Marek Kolasa, and Krystian Szadkowski. Marx, Engels, Lenin, and Stalin as highly cited researchers? Historical bibliometrics study. Scientometrics, 126(10):8683–8700, October 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04126-3.

Kumar:2024:PHR

[KKYR24] Avinash Kumar, Moumita Koley, Alfredo Yegros, and Ismael Rafols. Priorities of health research in India: evidence of misalignment between research outputs and disease burden. Scientometrics, 129(4):2433–2450, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04980-x.

Koley:2022:CDS

[KL22] Moumita Koley and Kanchan Lala. Changing dynamics of scholarly publication: a perspective towards open access publishing and the proposed one nation, one subscription policy of India. Scientometrics, 127(6):3383–3411, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04375-w.

Kuan:2024:AMP

[KL24] Chung-Huei Kuan and Ssu-Yu Liao. Assessing main paths by uncovering their coverage with key-node path search. Scientometrics, 129(11):6629–6657, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05155-4.

Kim:2025:DAA

[KLJS25] Sang Yoon Kim, Won Kyung Lee, Su Jung Jee, and So Young Sohn. Discovering AI adoption patterns from big academic graph data. Scientometrics, 130(2):809– 831, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05228-4.

Khelfaoui:2020:MNS

[KLLG20] Mahdi Khelfaoui, Julien Larrègue, Vincent Larivière, and Yves Gingras. Measuring national self-referencing patterns of major science producers. Scientometrics, 123 (2):979–996, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03381-0.

Krawczyk:2024:ASC

[KLM24] Małgorzata J. Krawczyk, Mateusz Libirt, and Krzysztof Malarz. Analysis of scientific cooperation at the international and intercontinental level. Scientometrics, 129(8): 4983–5002, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05107-y.

Kim:2020:BMT

[KLNW20] Yeon Hak Kim, Aaron D. Levine, Eric J. Nehl, and John P. Walsh. A bibliometric measure of translational science. Scientometrics, 125(3):2349–2382, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03668-2.

Kang:2020:ACI

[KM20] Byeongwoo Kang and Kazuyuki Motohashi. Academic contribution to industrial innovation by funding type. Scientometrics, 124(1):169–193, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03420-w.

Klinger:2021:DLD

[KMGS21] Joel Klinger, Juan Mateos-Garcia, and Konstantinos Stathoulopoulos. Deep learning, deep change? Mapping the evolution and geography of a general purpose technology. Scientometrics, 126(7):5589–5621, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03936-9.

Kaveh:2022:ECT

[KMM22] Mahsa Kaveh, Mahdieh Mirzabeigi, and Amirsaeid Moloodi. The effects of the challenges in the transliteration of Persian names into English on the recall of retrieved results in the Web of Science. Scientometrics, 127(2):1099– 1128, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04234-0.

Kang:2021:NOD

[KN21] Byeongwoo Kang and Kaoru Nabeshima. National origin diversity and innovation performance: the case of Japan. Scientometrics, 126(6):5333–5351, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03981-4.

Kroll:2022:FIN

[KN22] Henning Kroll and Peter Neuhäusler. “Formal and informal networkedness among German Academics”: exploring the role of conferences and co-publications in scientific performance. Scientometrics, 127(11):6431–6452, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04526-z.

Karaulova:2020:MRF

[KNT20] Maria Karaulova, Maria Nedeva, and Duncan A. Thomas. Mapping research fields using co-nomination: the case of hyper-authorship heavy flavour physics. Scientometrics, 124(3):2229–2249, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03538-x.

Koseoglu:2020:IIS

[Kös20a] Mehmet Ali Köseoglu. Identifying the intellectual structure of fields: introduction of the MAK approach. Scientometrics, 125(3):2169–2197, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03719-8.

Kosmulski:2020:NLH

[Kos20b] Marek Kosmulski. Nobel laureates are not hot. Scientometrics, 123(1):487–495, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03378-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03378-9.pdf.

Kim:2021:OLL

[KOS21a] Jinseok Kim and Jason Owen-Smith. ORCID-linked labeled data for evaluating author name disambiguation at scale. Scientometrics, 126(3):2057–2083, March 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03826-6.

Kosmulski:2021:PCA

[Kos21b] Marek Kosmulski. Posthumous co-authorship revisited. Scientometrics, 126(9):8227–8231, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04107-6.

Kosmulski:2022:TFA

[Kos22] Marek Kosmulski. Twenty-fifth anniversary of Sokal hoax. Scientometrics, 127(2):1187–1190, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04238-w.

Kosnik:2023:AEG

[Kos23] Lea-Rachel Kosnik. Additional evidence on gender and language in academic economics research. Scientometrics, 128(11):5949–5968, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04839-7.

Kosmulski:2025:GI

[Kos25] Marek Kosmulski. Generalized g-index. Scientometrics, 130(1):531–536, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-024-
05221-x.

Kousha:2024:HCA

[Kou24] Kayvan Kousha. How is ChatGPT acknowledged in academic publications? Scientometrics, 129(12):7959– 7969, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05193-y.

Kao:2023:MPA

[KP23a] Chiang Kao and Hui-Lan Pao. The most productive age of the management scholars in Taiwan. Scientometrics, 128(12):6719–6738, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04866-4.

Knisely:2023:RPC

[KP23b] Benjamin M. Knisely and Holly H. Pavliscsak. Research proposal content extraction using natural language processing and semi-supervised clustering: a demonstration and comparative analysis. Scientometrics, 128(5): 3197–3224, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04689-3.

Kuno:2022:DUN

[KPM22] Masaru Kuno, Mary Prorok, and Thurston Miller. Deciphering the US News and World Report ranking of US chemistry graduate programs. Scientometrics, 127(5): 2131–2150, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04317-6.

Krieger:2021:FWB

[KPS21] Bastian Krieger, Maikel Pellens, and Torben Schubert. Are firms withdrawing from basic research? An analysis of firm-level publication behaviour in Germany. Scientometrics, 126(12):9677–9698, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04147-y.

Koseoglu:2021:IIS

[KPY21] Mehmet Ali Köseoglu, John A. Parnell, and Melissa Yan Yee Yick. Identifying influential studies and maturity level in intellectual structure of fields: evidence from strategic management. Scientometrics, 126(2):1271– 1309, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03776-z.

Kenett:2021:GRF

[KR21] Ron S. Kenett and Abraham Rubinstein. Generalizing research findings for enhanced reproducibility: an approach based on verbal alternative representations. Scientometrics, 126(5):4137–4151, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03914-1.

Kwiek:2022:AVB

[KR22a] Marek Kwiek and Wojciech Roszka. Academic vs. biological age in research on academic careers: a large-scale study with implications for scientifically developing systems. Scientometrics, 127(6):3543–3575, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04363-0.

Kwiek:2022:FSL

[KR22b] Marek Kwiek and Wojciech Roszka. Are female scientists less inclined to publish alone? The gender solo research gap. Scientometrics, 127(4):1697–1735, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04308-7.

Krauskopf:2020:SCV

[Kra20] Erwin Krauskopf. Sources without a CiteScore value: more clarity is required. Scientometrics, 122(3):1801–1812, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03350-7.

Khurana:2022:IEH

[KS22a] Parul Khurana and Kiran Sharma. Impact of h-index on author’s rankings: an improvement to the h-index for lower-ranked authors. Scientometrics, 127(8):4483–4498, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04464-w.

Klincewicz:2022:SPO

[KS22b] Krzysztof Klincewicz and Szymon Szumiał. Successful patenting-not only how, but with whom: the importance of patent attorneys. Scientometrics, 127(9):5111– 5137, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04476-6.

Kumpulainen:2022:CWS

[KS22c] Miika Kumpulainen and Marko Seppänen. Combining Web of Science and Scopus datasets in citation-based literature study. Scientometrics, 127(10):5613–5631, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04475-7.

Kammari:2023:TSB

[KS23] Monachary Kammari and Durga Bhavani S. Timestamp based network evolution model for citation networks. Scientometrics, 128(6):3723–3741, June 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04704-7.

Kuperman:2024:CRM

[KS24] Vadim Y. Kuperman and Gerald H. Sokol. On the causes and ramifications of multi-authorship in science. Scientometrics, 129(4):2205–2225, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04963-y.

Kousha:2024:HDM

[KSAT24] Kayvan Kousha, Emma Stuart, Mahshid Abdoli, and Mike Thelwall. How do museums and galleries help academics create societal impact? An analysis of the UK REF2021 impact case studies. Scientometrics, 129(12):7759–7782, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05180-3.

Kiss:2024:IED

[KSP24] Anna Kiss, Sándor Soós, and Andrea Petróczi. Impact as equalizer: the demise of gender-related differences in anti-doping research. Scientometrics, 129(7):4071–4108, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05094-0.

Karanja:2020:WDM

[KSS20a] Erastus Karanja, Aditya Sharma, and Ibrahim Salama. What does MIS survey research reveal about diversity and representativeness in the MIS field? A content analysis approach. Scientometrics, 122(3):1583–1628, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03331-5.

Kreutz:2020:ESC

[KSS20b] Christin Katharina Kreutz, Premtim Sahitaj, and Ralf Schenkel. Evaluating semantometrics from computer science publications. Scientometrics, 125(3):2915–2954, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03409-5;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03409-5.pdf.

Koopmann:2021:PDE

[KSS21] Tobias Koopmann, Maximilian Stubbemann, and Gerd Stumme. Proximity dimensions and the emergence of collaboration: a HypTrails study on German AI research. Scientometrics, 126(12):9847–9868, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03922-1.

Kowal:2022:IGB

[KSZ22] Marta Kowal, Piotr Sorokowski, and Agnieszka Zelaźniewicz. The impact of geographical bias when judging scientific studies. Scientometrics, 127(1):265–273, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04176-7.

Kalenyuk:2021:AIC

[KT21] Iryna Kalenyuk and Liudmyla Tsymbal. Assessment of the intellectual component in economic development. Scientometrics, 126(6):4793–4816, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03958-3.

Kousha:2022:HSV

[KTB22] Kayvan Kousha, Mike Thelwall, and Matthew Bickley. The high scholarly value of grey literature before and during COVID-19. Scientometrics, 127(6):3489–3504, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04398-3.

Khan:2021:MIB

[KTK21] Nushrat Khan, Mike Thelwall, and Kayvan Kousha. Measuring the impact of biodiversity datasets: data reuse, citations and altmetrics. Scientometrics, 126(4):3621–3639, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-021-03890-6.

Kuan:2020:RWA

[Kua20] Chung-Huei Kuan. Regarding weight assignment algorithms of main path analysis and the conversion of arc weights to node weights. Scientometrics, 124(1):775–782, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03468-8.

Kuan:2023:DMP

[Kua23] Chung-Huei Kuan. Does main path analysis prefer longer paths? Scientometrics, 128(1):841–851, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04543-y.

Kuppler:2022:PFI

[Kup22] Matthias Kuppler. Predicting the future impact of Computer Science researchers: Is there a gender bias? Scientometrics, 127(11):6695–6732, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04337-2.

Kudaibergenova:2022:MPC

[KUS22] Renata Kudaibergenova, Sandugash Uzakbay, and Kadyrzhan Smagulov. Managing publication change at Al-Farabi Kazakh National University: a case study. Scientometrics, 127(1):453–479, January 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04139-y.

Kokol:2020:DSI

[KVZ20] Peter Kokol, Helena Blazun Vosner, and Jernej Zavrsnik. Do simultaneous inventions sleep? A case study on nursing sleeping papers. Scientometrics, 125(3):2827–2832, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03695-z.

Kokol:2022:SBH

[KVZ22] Peter Kokol, Helena Blazun Vosner, and Grega Zlahtic. Sleeping beauties in health informatics research. Scientometrics, 127(8):5073–5081, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04454-y.

Kullmann:2024:TPO

[KW24] Sylvia Kullmann and Verena Weimer. Teaching as part of open scholarship: developing a scientometric framework for Open Educational Resources. Scientometrics, 129(10): 6065–6087, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05007-1.

Kwiek:2020:ILI

[Kwi20] Marek Kwiek. Internationalists and locals: international research collaboration in a resource-poor system. Scientometrics, 124(1):57–105, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03460-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03460-2.pdf.

Khor:2020:RKT

[KY20] K. A. Khor and L. G. Yu. Revealing key topics shifts in thermal barrier coatings (TBC) as indicators of technological developments for aerospace engines. Scientometrics, 125(2):1763–1781, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03667-3.

Kuo:2021:UBM

[KYKC21] Shu-Chun Kuo, Yu-Tsen Yeh, Wei-Chih Kan, and Tsair-Wei Chien. The use of bootstrapping method to compare research achievements for ophthalmology authors in the US since 2010. Scientometrics, 126(1):509–520, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03725-w.

Kong:2020:EIT

[KYL20] Dejing Kong, Jianzhong Yang, and Lingfeng Li. Early identification of technological convergence in numerical control machine tool: a deep learning approach. Scientometrics, 125(3):1983–2009, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03696-y.

Kim:2020:PDC

[KYPC20] Jaeyoung Kim, Janghyeok Yoon, Eunjeong Park, and Sungchul Choi. Patent document clustering with deep embeddings. Scientometrics, 123(2):563–577, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03396-7.

Kong:2025:CSD

[KZH25] Mingyue Kong, Yinglong Zhang, and Kaifeng Hong. Citation structural diversity: a novel metric combining structure and semantics for literature evaluation. Scientometrics, 130(7):4011–4026, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05356-5.

Kokol:2020:DSP

[KZV20] Peter Kokol, Jernej Zavrsnik, and Helena Blazun Vosner. Did Sleeping Papers in nursing research miss their target audience? Scientometrics, 122(2):1243–1248, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03323-5.

Kou:2020:DGS

[KZZ^{+}20] Mingting Kou, Yi Zhang, Yu Zhang, Kaihua Chen, Jiancheng Guan, and Senmao Xia. Does gender structure influence R&D efficiency? A regional perspective. Scientometrics, 122(1):477–501, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03282-x.

Lyhagen:2020:URE

[LA20] Johan Lyhagen and Per Ahlgren. Uncertainty and the ranking of economics journals. Scientometrics, 125(3): 2545–2560, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03681-5;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03681-5.pdf.

Lamba:2020:RPH

[Lam20] Manika Lamba. Research productivity of health care policy faculty: a cohort study of Harvard Medical School. Scientometrics, 124(1):107–130, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03433-5.

Leta:2022:CDW

[LAT22] Jacqueline Leta, Kizi Araujo, and Stephanie Treiber. Citing documents of Wakefield’s retracted article: the domino effect of authors and journals. Scientometrics, 127(12): 7333–7349, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04353-2.

Laxdal:2023:RLE

[Lax23a] Aron Laxdal. Response to a letter to the editor by Williams, Stone, Ritz, and MacDonald. Scientometrics, 128(7):4161–4162, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-
04744-z.

Laxdal:2023:SGS

[Lax23b] Aron Laxdal. The sex gap in sports and exercise medicine research: who does research on females? Scientometrics, 128(3):1987–1994, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04641-5. See comments [WSRM23]

and response [Lax23a].

Leydesdorff:2020:ISO

[LB20] Loet Leydesdorff and Lutz Bornmann. “Interdisciplinarity” and “Synergy” in the OEuvre of Judit Bar-Ilan. Scientometrics, 123(3):1247–1260, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03451-3;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03451-3.pdf.

Leibel:2024:SUW

[LB24a] Christian Leibel and Lutz Bornmann. Specification uncertainty: what the disruption index tells us about the (hidden) multiverse of bibliometric indicators. Scientometrics, 129(12):7971–7979, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05201-1.

Leibel:2024:WDW

[LB24b] Christian Leibel and Lutz Bornmann. What do we know about the disruption index in scientometrics? An overview of the literature. Scientometrics, 129(1):601–639, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04873-5.

Li:2025:IGR

[LB25] Baicun Li and Aruhan Bai. The influence of grant renewal on research content: evidence from NIH-funded PIs. Scientometrics, 130(5):2595–2616, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05295-1.

Lancho-Barrantes:2021:QPP

[LBCO21] Barbara S. Lancho-Barrantes and Francisco J. Cantu-Ortiz. Quantifying the publication preferences of leading research universities. Scientometrics, 126(3):2269–2310, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03790-1.

Lemke:2022:RAP

[LBP22] Steffen Lemke, Max Brede, and Isabella Peters. Research articles promoted in embargo e-mails receive higher citations and altmetrics. Scientometrics, 127(1):75–97, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04217-1.

Lazebnik:2023:ACA

[LBS23] Teddy Lazebnik, Stephan Beck, and Labib Shami. Academic co-authorship is a risky game. Scientometrics, 128 (12):6495–6507, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-
04843-x.

Liu:2020:NTP

[LBW^{+}20] Yuxian Liu, Ewelina Biskup, Yueqian Wang, Fengfeng Cai, and Xiaoyan Zhang. A new territory and its pioneer: opening up a dominant research stream for a translational research area. Scientometrics, 125(2):1213– 1228, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03638-8.

Leydesdorff:2021:ICV

[LBW21] Loet Leydesdorff, Lutz Bornmann, and Caroline S. Wagner. Improved clusterings and visualizations of 11,359 journals in the JCRs 2015. Scientometrics, 126(6):5353–5354, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03938-7.

Lyu:2020:HDA

[LC20] Xiaozan Lyu and Rodrigo Costas. How do academic topics shift across altmetric sources? A case study of the research area of Big Data. Scientometrics, 123(2): 909–943, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03415-7.

Lyu:2021:SCS

[LC21] Xiaozan Lyu and Rodrigo Costas. Studying the characteristics of scientific communities using individual-level bibliometrics: the case of Big Data research. Scientometrics, 126(8):6965–6987, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04034-6.

Li:2022:MDC

[LC22] Jiexun Li and Jiyao Chen. Measuring destabilization and consolidation in scientific knowledge evolution. Scientometrics, 127(10):5819–5839, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04479-3.

LaQuatra:2020:EPW

[LCB20] Moreno La Quatra, Luca Cagliero, and Elena Baralis. Exploiting pivot words to classify and summarize discourse facets of scientific papers. Scientometrics, 125(3):3139– 3157, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03532-3.

LaQuatra:2021:LFT

[LCB21] Moreno La Quatra, Luca Cagliero, and Elena Baralis. Leveraging full-text article exploration for citation analysis. Scientometrics, 126(10):8275–8293, October 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04117-4.

Lamirel:2020:OHS

[LCC^{+}20] Jean-Charles Lamirel, Yue Chen, Pascal Cuxac, Shadi Al Shehabi, Nicolas Dugué, and Zeyuan Liu. An overview of the history of Science of Science in China based on the use of bibliographic and citation data: a new method of analysis based on clustering with feature maximization and contrast graphs. Scientometrics, 125(3):2971– 2999, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03503-8.

Lin:2023:CAS

[LCC23] Ju-Kuo Lin, Tsair-Wei Chien, and Willy Chou. Comment on the article: The state of social science research on COVID-19. Scientometrics, 128(2):1429–1436, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04600-6. See [LYZ22,

LYZ23].

Lin:2024:RRA

[LCCH24] Ting-Yu Lin, Hao-Chien Cheng, Li-Fu Cheng, and Tsung-Min Hung. Registered report adoption in academic journals: assessing rates in different research domains. Scientometrics, 129(4):2123–2130, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04896-y.

Lindahl:2020:ECP

[LCD20] Jonas Lindahl, Cristian Colliander, and Rickard Danell. Early career performance and its correlation with gender and publication output during doctoral education. Scientometrics, 122(1):309–330, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03262-1;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03262-1.pdf.

Li:2023:IAT

[LCL23] Bing Li, Shiji Chen, and Vincent Larivière. Interdisciplinarity affects the technological impact of scientific research. Scientometrics, 128(12):6527–6559, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04846-8.

Lovakov:2022:UVR

[LCP22] Andrey Lovakov, Maia Chankseliani, and Anna Panova. Universities vs. research institutes? Overcoming the Soviet legacy of higher education and research. Scientometrics, 127(11):6293–6313, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04527-y.

Liu:2025:CCP

[LCS25] Junwan Liu, Zining Cui, and Yinglu Song. Choosing a career partner: birds of a feather flock together. Scientometrics, 130(5):2705–2748, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05299-x.

Liu:2020:MTE

[LCT^{+}20] Huailan Liu, Zhiwang Chen, Jie Tang, Yuan Zhou, and Sheng Liu. Mapping the technology evolution path: a novel model for dynamic topic detection and tracking. Scientometrics, 125(3):2043–2090, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-020-03700-5;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03700-5.pdf.

Lu:2024:RCL

[LCT^{+}24] Yonghe Lu, Lehua Chen, Xinyu Tong, Yongxin Peng, and Hou Zhu. Research on cross-lingual multi-label patent classification based on pre-trained model. Scientometrics, 129 (6):3067–3087, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05024-0.

Labbe:2020:FIN

[LCW^{+}20] Cyril Labbé, Guillaume Cabanac, Rachael A. West, Thierry Gautier, Bertrand Favier, and Jennifer A. Byrne. Flagging incorrect nucleotide sequence reagents in biomedical papers: To what extent does the leading publication format impede automatic error detection? Scientometrics, 124(2):1139–1156, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03463-z;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03463-z.pdf.

Li:2021:EDI

[LCW21] Yating Li, Ye Chen, and Qiyu Wang. Evolution and diffusion of information literacy topics. Scientometrics, 126 (5):4195–4224, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03925-y.

Lee:2021:CER

[LCY21] Kyoungmi Lee, Sunglok Choi, and Jae-Suk Yang. Can expensive research equipment boost research and development performances? Scientometrics, 126(9):7715– 7742, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04088-6.

See correction [LCY22].

Lee:2022:CCE

[LCY22] Kyoungmi Lee, Sunglok Choi, and Jae-Suk Yang. Correction to: Can expensive research equipment boost research and development performances? Scientometrics, 127(3):

1663–1665, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04204-6.

See [LCY21].

Luan:2022:MGC

[LDA22] Chunjuan Luan, Siming Deng, and John R. Allison. Mutual Granger “causality” between scientific instruments and scientific publications. Scientometrics, 127(11):6209– 6229, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04516-1.

Lucas-Dominguez:2021:SRD

[LDAAAB21] Rut Lucas-Dominguez, Adolfo Alonso-Arroyo, and Rafael Aleixandre-Benavent. The sharing of research data facing the COVID-19 pandemic. Scientometrics, 126(6): 4975–4990, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03971-6.

Lei:2024:GRD

[LDDL24] Fang Lei, Liang Du, Min Dong, and Xuemei Liu. Global retractions due to randomly generated content: Characterization and trends. Scientometrics, 129(12):7943–7958, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05172-3.

Limaymanta:2022:RBC

[LdGRV^{+}22] Cesar H. Limaymanta, Rosalía Quiroz de García, Jesús A. Rivas-Villena, Andrea Rojas-Arroyo, and Orlando Gregorio-Chaviano. Relationship between collaboration and normalized scientific impact in South American public universities. Scientometrics, 127(11):6391–6411, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04523-2.

Lopez-Duarte:2021:MCI

[LDMVS21] Cristina López-Duarte, Jane F. Maley, and Marta M. Vidal-Suárez. Main challenges to international student mobility in the European arena. Scientometrics, 126(11):

8957–8980, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04155-y.

Lovakov:2025:SIR

[LdS25] Andrey Lovakov and Jaime A. Teixeira da Silva. Scientometric indicators in research evaluation and research misconduct: analysis of the Russian university excellence initiative. Scientometrics, 130(3):1813–1829, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05269-3.

LeClercq:2024:APP

[Le 24] Louis-Stéphane Le Clercq. ABCal: a Python package for author bias computation and scientometric plotting for reviews and meta-analyses. Scientometrics, 129(1):581–600, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04880-6.

Leckert:2021:VMC

[Lec21] Max Leckert. (E-) valuative metrics as a contested field: a comparative analysis of the Altmetrics- and the Leiden Manifesto. Scientometrics, 126(12):9869–9903, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04039-1.

Lee:2024:EDR

[Lee24] Danielle Lee. Exploring the determinants of research performance for early-career researchers: a literature review. Scientometrics, 129(1):181–235, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04868-2.

Lackner:2021:AEC

[LFV21] Arthur Lackner, Said Fathalla, and Sahar Vahdati. Analysing the evolution of computer science events leveraging a scholarly knowledge graph: a scientometrics study of top-ranked events in the past decade. Scientometrics, 126(9):8129–8151, September 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04072-0.

Lamirel:2021:ADI

[LGD21] Jean-Charles Lamirel, Younes Gueddari, and Nicolas Dugué. Analysis of the dynamics and influence of the research work of Prof. Liu Zeyuan in China featuring a new hybrid approach combining community detection with topic tracking. Scientometrics, 126(7):6273–6300, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04010-0.

Li:2023:TFT

[LGE^{+}23] Ruinan Li, Raf Guns, Tim C. E. Engels, Lin Zhang, and Ying Huang. Tracking the featured topics of the International Science of Team Science conference series and their evolution during 2010–2019. Scientometrics, 128(4): 2447–2469, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04651-3.

Lin:2020:EBM

[LGLM20] Deming Lin, Tianhui Gong, Wenbin Liu, and Martin Meyer. An entropy-based measure for the evolution of h index research. Scientometrics, 125(3):2283–2298, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03712-1.

Liu:2021:ITK

[LGMK21] Jie Liu, Arnulf Grubler, Tieju Ma, and Dieter F. Kogler. Identifying the technological knowledge depreciation rate using patent citation data: a case study of the solar photovoltaic industry. Scientometrics, 126(1):93–115, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03740-x.

Lyu:2021:DRC

[LGR^{+}21] Dongqing Lyu, Kaile Gong, Xuanmin Ruan, Ying Cheng, and Jiang Li. Does research collaboration influence the ”disruption” of articles? Evidence from neurosciences. Scientometrics, 126(1):287–303, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03757-2.

Liu:2024:URB

[LGXH24] Junwan Liu, Xiaoyun Gong, Shuo Xu, and Chenchen Huang. Understanding the relationship between team diversity and the innovative performance in research teams using decision tree algorithms: evidence from artificial intelligence. Scientometrics, 129(12):7805–7831, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05183-0.

Li:2023:MIS

[LGY23] Ming Li, Qian Gao, and Tianfei Yu. Methodological issues on statistical rigor of agreement analysis. Scientometrics, 128(3):2025–2027, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04591-4.

Liu:2022:ERB

[LH22a] Chieh Liu and Mu-Hsuan Huang. Exploring the relationships between altmetric counts and citations of papers in different academic fields based on co-occurrence analysis. Scientometrics, 127(8):4939–4958, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04456-w.

Liu:2022:RCP

[LH22b] Hsuan-I Liu and Mu-Hsuan Huang. Research contribution pattern analysis of multinational authorship papers. Scientometrics, 127(4):1783–1800, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04277-x.

Li:2024:MAB

[LH24] Hui Li and Yaohua Hu. Metapath and attribute-based academic collaborator recommendation in heterogeneous academic networks. Scientometrics, 129(7):4295–4315, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05043-x.

Lin:2021:RNB

[LHC21] Arthur Jing Lin, Chien-Lung Hsu, and Chun-Hao Chiang. Retraction note to: Bibliometric study of Electronic Commerce Research in Information Systems & MIS Journals. Scientometrics, 126(1):909, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03818-6;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03818-6.pdf.

Lee:2022:CTC

[LHC22] Soyea Lee, Junseok Hwang, and Eunsang Cho. Comparing technology convergence of artificial intelligence on the industrial sectors: two-way approaches on network analysis and clustering analysis. Scientometrics, 127(1):407–452, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04170-z.

Lee:2025:DPA

[LHC25] Soyea Lee, Junseok Hwang, and Eunsang Cho. Dynamic patterns of AI technology diffusion: focusing on time series clustering and patent analysis. Scientometrics, 130(4): 2005–2036, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05283-5.

Liang:2020:DAE

[LHCH20] Guoqiang Liang, Haiyan Hou, Qiao Chen, and Zhigang Hu. Diffusion and adoption: an explanatory model of “question mark” and “rising star” articles. Scientometrics, 124(1):219–232, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03478-6.

Levene:2020:TDB

[LHF20] Mark Levene, Martyn Harris, and Trevor Fenner. A twodimensional bibliometric index reflecting both quality and quantity. Scientometrics, 123(3):1235–1246, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03454-0.

Li:2023:PSI

[LHF23] Guan Cheng Li, Lingyun He, and Lee Fleming. Philanthropic supported innovation: trends, areas, and impact. Scientometrics, 128(10):5507–5520, October 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04796-1.

Lee:2021:AMT

[LHK21] Changyong Lee, Suckwon Hong, and Juram Kim. Anticipating multi-technology convergence: a machine learning approach using patent information. Scientometrics, 126 (3):1867–1896, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-020-
03842-6.

Lalli:2020:DCN

[LHW20] Roberto Lalli, Riaz Howey, and Dirk Wintergrün. The dynamics of collaboration networks and the history of general relativity, 1925–1970. Scientometrics, 122(2): 1129–1170, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-019-
03327-1;   http://link.springer.com/content/pdf/
10.1007/s11192-019-03327-1.pdf.

Liu:2021:SJD

[LHW21] Weishu Liu, Meiting Huang, and Haifeng Wang. Same journal but different numbers of published records indexed in Scopus and Web of Science Core Collection: causes, consequences, and solutions. Scientometrics, 126 (5):4541–4550, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03934-x.

Li:2024:DSP

[LHX^{+}24] Boze Li, Yandong He, Yuxuan Xiu, Bokui Chen, and Wai Kin Victor Chan. The distribution of scientific project funds model based on adaptive similarity fitting and NSGA–II. Scientometrics, 129(12):7585–7622, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05190-1.

Lee:2020:TLB

[LHXS20] Na Kyeong Lee, Yukyeong Han, Wei Xong, and Min Song. Two layer-based trajectory analysis of the research trend in automotive fuel industry. Scientometrics, 124(3):1701– 1719, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03506-5.

Lou:2023:SBS

[LHZ^{+}23] Wen Lou, Jiangen He, Lingxin Zhang, Zhijie Zhu, and Yongjun Zhu. Support behind the scenes: the relationship between acknowledgement, coauthor, and citation in Nobel articles. Scientometrics, 128(10):5767–5790, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04803-5.

Li:2022:AWL

[Li22] Zhijun Li. Is academic writing less passivized? Corpusbased evidence from research article abstracts in applied linguistics over the past three decades (1990–2019). Scientometrics, 127(10):5773–5792, October 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04498-0.

Li:2025:GRH

[Li25] Hui Li. Global or regional: the hidden truth behind ShanghaiRanking’s global university ranking by the subject of law. Scientometrics, 130(1):515–530, January 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05214-w.

Lietz:2020:DIB

[Lie20] Haiko Lietz. Drawing impossible boundaries: field delineation of Social Network Science. Scientometrics, 125(3):2841–2876, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03527-0;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03527-0.pdf.

Lin:2020:SPA

[Lin20] Wen-Yau Cathy Lin. Self-plagiarism in academic journal articles: from the perspectives of international Editors-in-Chief in editorial and COPE case. Scientometrics, 123 (1):299–319, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03373-0.

Lin:2021:CEE

[Lin21a] Jun-You Lin. Collaboration exploitation and exploration: does a proactive search strategy matter? Scientometrics, 126(10):8295–8329, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04136-1.

Lin:2021:EOA

[Lin21b] Wen-Yau Cathy Lin. Effects of open access and articlesin-press mechanisms on publishing lag and first-citation speed: a case on energy and fuels journals. Scientometrics, 126(6):4841–4869, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03956-5.

Linnemer:2024:MRL

[Lin24] Laurent Linnemer. A menagerie of rankings: a look in RePEc’s factory. Scientometrics, 129(1):321–372, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04879-z.

Liu:2020:AFI

[Liu20] Weishu Liu. Accuracy of funding information in Scopus: a comparative case study. Scientometrics, 124(1): 803–811, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03458-w.

Liu:2021:MTP

[Liu21] Weishu Liu. A matter of time: publication dates in Web of Science Core Collection. Scientometrics, 126(1): 849–857, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03697-x.

Liu:2023:RSP

[Liu23] Fang Liu. Retrieval strategy and possible explanations for the abnormal growth of research publications: reevaluating a bibliometric analysis of climate change. Scientometrics, 128(1):853–859, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04540-1.

Liu:2024:DCD

[Liu24] Jie Liu. “Divergent” cross-domain stretching for technology fusion: validating the knowledge partition search model using patent data. Scientometrics, 129(6):3023– 3043, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05049-5.

Lee:2025:UMR

[LKPS25] SooJeung Lee, Su Jin Kim, Sunna Park, and Jung Cheol Shin. Unequal metrics in research publications: the impact of bibliometric databases and faculty size across academic disciplines on university rankings in south Korea. Scientometrics, 130(2):641–663, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05223-9.

Li:2020:SSD

[LL20] Hui Li and Weishu Liu. Same same but different: self-citations identified through Scopus and Web of Science Core Collection. Scientometrics, 124(3):2723–2732, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03573-8.

Liu:2021:RME

[LL21] Wenjun Liu and Lei Lei. Retractions in the Middle East from 1999 to 2018: a bibliometric analysis. Scientometrics, 126(6):4687–4700, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03919-w.

Liu:2024:CBS

[LL24a] Chih-Hsing Liu and Jun-You Lin. Collaboration-based scientific productivity: evidence from Nobel laureates. Scientometrics, 129(7):3735–3768, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05062-8.

Liu:2024:UER

[LL24b] Xipeng Liu and Xinmiao Li. Unbiased evaluation of ranking algorithms applied to the Chinese green patents citation network. Scientometrics, 129(6):2999–3021, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05023-1.

Li:2025:ARC

[LL25] Xinzhe Li and Xiao Lu. Analysis of the research collaboration organizational characteristics and scientific impact of large-scale research facilities: a case study of Chinese largescale research facilities. Scientometrics, 130(5):2617–2638, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05314-1.

Li:2022:GRW

[LLC22] Pengcheng Li, Wei Lu, and Qikai Cheng. Generating a related work section for scientific papers: an optimized approach with adopting problem and method information. Scientometrics, 127(8):4397–4417, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04458-8.

Liu:2020:NCT

[LLH20] John S. Liu, Louis Y. Y. Lu, and Mei Hsiu-Ching Ho. A note on choosing traversal counts in main path analysis. Scientometrics, 124(1):783–785, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03469-7.

Liang:2022:RDI

[LLH22] Guoqiang Liang, Ying Lou, and Haiyan Hou. Revisiting the disruptive index: evidence from the Nobel Prizewinning articles. Scientometrics, 127(10):5721–5730, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04499-z.

Lee:2025:EUP

[LLH25] Seungmin Lee, Jeong-Dong Lee, and Youwei He. Exaptation: unveiling the potential for technological innovation. Scientometrics, 130(7):3455–3482, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05362-7.

Li:2022:TCC

[LLL22] Li Li, Haifen Lin, and Yibo Lyu. Technology cluster coupling and invulnerability of industrial innovation networks: the role of centralized structure and technological turbulence. Scientometrics, 127(3):1209–1231, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04269-x.

Lin:2025:AEI

[LLL25] Runhui Lin, Yalin Li, and Biting Li. AI-enabled individual learning strategies and scientific innovation: a case from the field of computer science. Scientometrics, 130 (7):3629–3650, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05345-8.

Lin:2024:DAC

[LLLL24] Runhui Lin, Biting Li, Yanhong Lu, and Yalin Li. Degree assortativity in collaboration networks and breakthrough innovation: the moderating role of knowledge networks. Scientometrics, 129(7):3809–3839, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05063-7.

Lamirel:2024:CTM

[LLM24] Jean-Charles Lamirel, Francis Lareau, and Christophe Malaterre. CFMf topic-model: comparison with LDA and Top2Vec. Scientometrics, 129(10):6387–6405, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05017-z.

Liskiewicz:2021:FAC

[LLP21] T. Liskiewicz, G. Liskiewicz, and J. Paczesny. Factors affecting the citations of papers in tribology journals. Scientometrics, 126(4):3321–3336, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03870-w;   http://link.springer.
com/content/pdf/10.1007/s11192-021-03870-w.pdf.

Liang:2023:MDM

[LLSH23] Guoqiang Liang, Yaqin Li, Lurui Song, and Chaoguang Huo. Magnitude decrease of the Matthew effect in citations: a study based on Nobel Prize articles. Scientometrics, 128(12):6357–6371, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04874-4.

Lu:2021:TRM

[LLZ21] Yonghe Lu, Jiayi Luo, and Hou Zhu. Text representation model of scientific papers based on fusing multiviewpoint information and its quality assessment. Scientometrics, 126(8):6937–6963, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04028-4.

Liu:2025:NUL

[LLZ25] Xipeng Liu, Xinmiao Li, and Ping Zhang. A novel unsupervised learning framework for measuring the technological innovation of patents. Scientometrics, 130(8):4157–4186, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05380-5.

Lund:2020:RBH

[LM20] Brady D. Lund and Sanjay Kumar Maurya. The relationship between highly-cited papers and the frequency of citations to other papers within-issue among three top information science journals. Scientometrics, 125(3):2491– 2504, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03720-1.

Lascialfari:2022:URL

[LMC22] Matteo Lascialfari, Marie-Benoît Magrini, and Guillaume Cabanac. Unpacking research lock-in through a diachronic analysis of topic cluster trajectories in scholarly publications. Scientometrics, 127(11):6165–6189, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04514-3.

Liu:2020:KCO

[LMG20] Na Liu, Jianqi Mao, and Jiancheng Guan. Knowledge convergence and organization innovation: the moderating role of relational embeddedness. Scientometrics, 125(3):1899– 1921, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03684-2.

Li:2022:HSR

[LMH22] Yashan Li, Jinge Mao, and Ying Huang. How scientific research incorporates policy: an examination using the case of China’s science and technology evaluation system. Scientometrics, 127(9):5283–5306, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04215-3.

Liang:2021:FCP

[LML21] Zhentao Liang, Jin Mao, and Gang Li. Finding citations for PubMed: a large-scale comparison between five freely available bibliographic data sources. Scientometrics, 126(12):9519–9542, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04191-8.

Luna-Morales:2022:RIP

[LMLMPA22] María Elena Luna-Morales, Evelia Luna-Morales, and Miguel Ángel Perez-Angon. Reflections on the institutionalization process of scientific research in Latin America: the case of Cinvestav. Scientometrics, 127(1):661–681, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04211-7.

Luque-Martinez:2024:SSR

[LMLR24] Teodoro Luque-Martínez and Ignacio Luque-Raya. Spanish scientific research by field and subject. Strategic analysis with ARWU indicators. Scientometrics, 129(9):5265– 5285, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05128-7.

Lopez-Morales:2022:IQM

[LMSNZG22] José Satsumi López-Morales, Héctor Francisco Salazar-Núñez, and Claudia Guadalupe Zarrabal-Gutiérrez. The impact of qualitative methods on article citation: an international business research perspective. Scientometrics, 127 (6):3225–3236, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04390-x.

Lai:2024:MKI

[LN24] Xiuping Lai and Libing Nie. Maturity of knowledge inputs and the breakthrough of key core technology. Scientometrics, 129(11):6551–6570, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05051-x.

Liu:2024:WSC

[LNH24] Weishu Liu, Rong Ni, and Guangyuan Hu. Web of Science Core Collection’s coverage expansion: the forgotten Arts & Humanities Citation Index? Scientometrics, 129(2): 933–955, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04917-w.

Lathabai:2021:IIC

[LNS21] Hiran H. Lathabai, Abhirup Nandy, and Vivek Kumar Singh. x-index: Identifying core competency and thematic research strengths of institutions using an NLP and network based ranking framework. Scientometrics, 126(12):9557–9583, December 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04188-3.

Lopez-Olmedo:2021:TKP

[LOGS21] Roberto Lopez-Olmedo and Norma Georgina Gutierrez-Serrano. Transdisciplinary knowledge production in mainstream journals: from the perspective of the participation of social actors in Mexico. Scientometrics, 126(3): 2627–2641, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03656-6.

Liu:2020:DAE

[LP20] Xiaoyu Liu and Alan L. Porter. A 3-dimensional analysis for evaluating technology emergence indicators. Scientometrics, 124(1):27–55, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03432-6.

Lathabai:2023:CEI

[LP23] Hiran H. Lathabai and Thara Prabhakaran. Contextual Ψ-index and its estimate for contextual productivity assessment. Scientometrics, 128(8):4875–4886, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04757-8.

Li:2021:TMK

[LPD21] Xiaoying Li, Suyuan Peng, and Jian Du. Towards medical knowmetrics: representing and computing medical knowledge using semantic predications as the knowledge unit and the uncertainty as the knowledge context. Scientometrics, 126(7):6225–6251, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03880-8.

Lang:2020:EMC

[LPKG20] Raynell Lang, Kholoud Porter, Hartmut B. Krentz, and M. John Gill. Evaluating medical conferences: the emerging need for a quality metric. Scientometrics, 122(1): 759–764, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03291-w.

Lovakov:2022:GVN

[LPY22] Andrey Lovakov, Anna Panova, and Maria Yudkevich. Global visibility of nationally published research output: the case of the post-Soviet region. Scientometrics, 127(5): 2643–2659, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04326-5.

Lutmar:2021:AGB

[LR21] Carmela Lutmar and Yaniv Reingewertz. Academic ingroup bias in the top five economics journals. Scientometrics, 126(12):9543–9556, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04174-9.

Liu:2023:PPR

[LR23] Yuxian Liu and Ronald Rousseau. A proposal for the peer review procedure for funding decisions. Scientometrics, 128(1):861–865, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04538-9.

Lazebnik:2025:HLI

[LR25] Teddy Lazebnik and Ariel Rosenfeld. How lonely or influential is the Lone Wolf? An analysis of individual scholars’ solo-authorship dynamics. Scientometrics, 130 (5):3029–3051, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05332-z.

Li:2022:LIR

[LRH22] Xian Li, Ronald Rousseau, and Xiaojun Hu. Is low interdisciplinarity of references an unexpected characteristic of Nobel Prize winning research? Scientometrics, 127(4): 2105–2122, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04290-0.

Lievore:2021:REP

[LRP21] Caroline Lievore, Priscila Rubbo, and Luiz Alberto Pilatti. Research ethics: a profile of retractions from world class universities. Scientometrics, 126(8):6871–6889, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03987-y.

Lewison:2021:SCM

[LRSB21] Grant Lewison, Philip Roe, Richard Sullivan, and Martin Bricknell. The spin-off to civilian medical practice in the UK and USA from medical research developed during conflict. Scientometrics, 126(2):1829–1839, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03738-5.

Lutter:2025:WTP

[LRW25] Mark Lutter, Jan Riebling, and Linus Weidner. Who talks to the prof? Gender differences in interaction with senior scholars at four academic conferences. Scientometrics, 130(8):4697–4722, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05392-1.

Lyu:2021:CCM

[LRXC21] Dongqing Lyu, Xuanmin Ruan, Juan Xie, and Ying Cheng. The classification of citing motivations: a metasynthesis. Scientometrics, 126(4):3243–3264, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-021-03908-z.

Lee:2020:ICI

[LS20a] Minchul Lee and Min Song. Incorporating citation impact into analysis of research trends. Scientometrics, 124(2): 1191–1224, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03508-3.

Lei:2020:SHC

[LS20b] Lei Lei and Yunmei Sun. Should highly cited items be excluded in impact factor calculation? The effect of review articles on journal impact factor. Scientometrics, 122(3): 1697–1706, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03338-y.

Lee:2021:PDB

[LS21a] Jong Wook Lee and So Young Sohn. Patent data based search framework for IT R&D employees for convergence technology. Scientometrics, 126(7):5687–5705, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04011-z.

Lee:2021:RST

[LS21b] Jungpyo Lee and So Young Sohn. Recommendation system for technology convergence opportunities based on selfsupervised representation learning. Scientometrics, 126 (1):1–25, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03731-y.

Link:2021:SPU

[LS21c] Albert N. Link and John T. Scott. Scientific publications at U.S. federal research laboratories. Scientometrics, 126(3):2227–2248, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03854-2.

Lund:2023:EUS

[LS23] Brady Lund and Amrollah Shamsi. Examining the use of supportive and contrasting citations in different disciplines: a brief study using Scite (scite.ai) data. Scientometrics, 128(8):4895–4900, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04781-8.

Lai:2022:RPA

[LSA22] Katerina Anne Lai, Gaurav Saxena, and Peter J. Allen. Research performance of academic psychologists in the United Kingdom. Scientometrics, 127(7):4139–4166, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04424-4.

Li:2021:IUI

[LSBC21] Siying Li, Huawei Shen, Peng Bao, and Xueqi Cheng. h_{u}- index: a unified index to quantify individuals across disciplines. Scientometrics, 126(4):3209–3226, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03879-1.

Leon-Silva:2020:SNR

[LSFLZLLV20] Sein León-Silva, Fabián Fernández-Luqueño, Edgar Záyago-Lau, and Fernando López-Valdez. Silver nanoparticles, research and development in Mexico: a bibliometric analysis. Scientometrics, 123(1):31–49, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03367-y.

Liu:2020:GDF

[LSY20] Junwan Liu, Yinglu Song, and Sai Yang. Gender disparities in the field of economics. Scientometrics, 125(2):1477– 1498, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03627-x.

Liu:2021:TDA

[LSY21] Na Liu, Philip Shapira, and Xiaoxu Yue. Tracking developments in artificial intelligence research: constructing and applying a new search strategy. Scientometrics, 126(4):3153–3192, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
021-03868-4;   http://link.springer.com/content/
pdf/10.1007/s11192-021-03868-4.pdf.

Liu:2020:OIP

[LT20] Ting Liu and Liu Tang. Open innovation from the perspective of network embedding: knowledge evolution and development trend. Scientometrics, 124(2):1053–1080, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03520-7.

Liu:2020:FIW

[LTH20] Weishu Liu, Li Tang, and Guangyuan Hu. Funding information in Web of Science: an updated overview. Scientometrics, 122(3):1509–1524, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03362-3.

Li:2023:TBA

[LTL23] Xin Li, Xuli Tang, and Wei Lu. Tracking biomedical articles along the translational continuum: a measure based on biomedical knowledge representation. Scientometrics, 128(2):1295–1319, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04607-z.

Li:2024:HBP

[LTL24] Xin Li, Xuli Tang, and Wei Lu. How biomedical papers accumulated their clinical citations: a large-scale retrospective analysis based on PubMed. Scientometrics, 129 (6):3315–3339, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05016-0.

Lu:2024:KGE

[LTXZ24] Yonghe Lu, Xinyu Tong, Xin Xiong, and Hou Zhu. Knowledge graph enhanced citation recommendation model for patent examiners. Scientometrics, 129(4):2181–2203, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04966-9.

Lyutov:2021:CML

[LUH21a] Alexey Lyutov, Yilmaz Uygun, and Marc-Thorsten Hütt. Correction to: Machine learning misclassification of academic publications reveals non-trivial interdependencies of scientific disciplines. Scientometrics, 126(2):1187, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03847-1;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03847-1.pdf. See [LUH21b].

Lyutov:2021:MLM

[LUH21b] Alexey Lyutov, Yilmaz Uygun, and Marc-Thorsten Hütt. Machine learning misclassification of academic publications reveals non-trivial interdependencies of scientific disciplines. Scientometrics, 126(2):1173–1186, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03789-8;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03789-8.pdf. See [LUH21a].

Li:2024:IPT

[LW24] Moxin Li and Yang Wang. Influence of political tensions on scientific productivity, citation impact, and knowledge combinations. Scientometrics, 129(4):2337–2370, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04973-w.

Liu:2021:WCB

[LWG21] Yuanyuan Liu, Qiang Wu, and Yong Gao. Weighted citation based on ranking-related contribution: a new index for evaluating article impact. Scientometrics, 126(10): 8653–8672, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04115-6.

Li:2024:PCR

[LWL24] Xiuxiu Li, Mingyang Wang, and Xu Liu. Predicting collaborative relationship among scholars by integrating scholars’ content-based and structure-based features. Scientometrics, 129(6):3225–3244, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05012-4.

Liu:2020:RSR

[LWR20] Yuxian Liu, Yishan Wu, and Ronald Rousseau. Reflections on and a short review of the science of team science. Scientometrics, 125(2):937–950, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03513-6.

Liu:2022:IDT

[LWW22] Xiwen Liu, Xuezhao Wang, and Yanpeng Wang. Identifying disruptive technologies by integrating multi-source data. Scientometrics, 127(9):5325–5351, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04283-z.

Liu:2024:BBD

[LWW24] Zhifeng Liu, Chenlin Wang, and Ruojia Wang. From bench to bedside: determining what drives academic citations in clinical trials. Scientometrics, 129(11):6813–6837, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05173-2.

Lin:2025:HCP

[LWW25] Jia Lin, Howei Wu, and Ho-Mou Wu. How to characterize patent quality with multiple indicators? Evidence based on economic performance of Chinese companies. Scientometrics, 130(8):4221–4248, August 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05401-3.

Liu:2021:WAM

[LWX21] Junwan Liu, Rui Wang, and Shuo Xu. What academic mobility configurations contribute to high performance: an fsQCA analysis of CSC-funded visiting scholars. Scientometrics, 126(2):1079–1100, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03783-0.

Liu:2022:RRM

[LWZ22] Xiaoyu Liu, Xuefeng Wang, and Donghua Zhu. Reviewer recommendation method for scientific research proposals: a case for NSFC. Scientometrics, 127(6):3343–3366, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04389-4.

Liu:2025:UEF

[LXL25] Yunhan Liu, Xia Xu, and Shuqing Li. Understanding of evolutionary features in the library and information science with interdisciplinary network analysis. Scientometrics, 130(2):781–808, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05236-y.

Lv:2022:MVM

[LXS22] Yiqin Lv, Zheng Xie, and Yiping Song. A multi-view method of scientific paper classification via heterogeneous graph embeddings. Scientometrics, 127(8):4847–4872, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04419-1.

Li:2023:CCD

[LXY23] Shenghui Li, Wenyan Xu, and Jingqi Yin. Cross-cultural differences in retracted publications of male and female from a global perspective. Scientometrics, 128(7):3805– 3826, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04717-2.

Lu:2020:RCS

[LXZ^{+}20] Yonghe Lu, Xin Xiong, Weiting Zhang, Jiaxin Liu, and Ruijie Zhao. Research on classification and similarity of patent citation based on deep learning. Scientometrics, 123 (2):813–839, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03385-w.

Lin:2020:HIU

[LY20] Jun-You Lin and Chih-Hai Yang. Heterogeneity in industry-university R&D collaboration and firm innovative performance. Scientometrics, 124(1):1–25, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03436-2.

Lu:2023:RSR

[LYLC23] Yonghe Lu, Meilu Yuan, Jiaxin Liu, and Minghong Chen. Research on semantic representation and citation recommendation of scientific papers with multiple semantics fusion. Scientometrics, 128(2):1367–1393, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04566-5.

Yue:2020:EFL

[lYnLyO^{+}20] Ming liang Yue, Rui nan Li, Gui yan Ou, Xia Wu, and Ting can Ma. An exploration on the flow of leading research talents in China: from the perspective of distinguished young scholars. Scientometrics, 125(2):1559– 1574, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03562-x.

Lin:2022:DAM

[LYS22] Jialiang Lin, Yao Yu, and Xiaodong Shi. Detecting and analyzing missing citations to published scientific entities. Scientometrics, 127(5):2395–2412, May 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04334-5.

Li:2020:HIH

[LYT^{+}20] Xin Li, Qiang Yao, Xuli Tang, Qian Li, and Mengjia Wu. How to investigate the historical roots and evolution of research fields in China? A case study on iMetrics using RootCite. Scientometrics, 125(2):1253–1274, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03659-3.

Lyudmyla:2021:EAJ

[Lyu21] Shkulipa Lyudmyla. Evaluation of accounting journals by coverage of accounting topics in 2018–2019. Scientometrics, 126(9):7251–7327, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03875-5.

Lin:2020:HMP

[LYZ^{+}20] Jialiang Lin, Yao Yu, Yu Zhou, Zhiyang Zhou, and Xiaodong Shi. How many preprints have actually been printed and why: a case study of computer science preprints on arXiv. Scientometrics, 124(1):555–574, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03430-8.

Liu:2022:SSS

[LYZ22] Yan-Li Liu, Wen-Juan Yuan, and Shao-Hong Zhu. The state of social science research on COVID-19. Scientometrics, 127(1):369–383, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/
10.1007/s11192-021-04206-4. See comments [LYZ23,

LCC23] and response [LYZ23].

Liu:2023:RDC

[LYZ23] Yan-Li Liu, Wen-Juan Yuan, and Shao-Hong Zhu. Response to Dr. Chou’s comment on “The state of social science research on COVID-19”. Scientometrics, 128(2): 1437–1439, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04603-3.

See [LYZ22, LCC23].

Luo:2021:MNE

[LZ21] Taiye Luo and Zhengang Zhang. Multi-network embeddedness and innovation performance of R&D employees. Scientometrics, 126(9):8091–8107, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04106-7.

Liu:2023:LPS

[LZ23] Xueying Liu and Haoran Zhu. Linguistic positivity in soft and hard disciplines: temporal dynamics, disciplinary variation, and the relationship with research impact. Scientometrics, 128(5):3107–3127, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04679-5.

Li:2024:DAS

[LZ24a] Hui Li and Xingmei Zhang. Dr. Anonymous is still there: a revisit of legal scholarly publishing. Scientometrics, 129(1): 681–692, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04912-1.

Liu:2024:MCA

[LZ24b] Weishu Liu and Ruifeng Zhang. Multilateral coauthorship: an important but easily overlooked pattern in international scientific collaboration research. Scientometrics, 129(7):4661–4668, July 2024. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-024-05095-z.

Liu:2025:PRI

[LZ25a] Jiaying Liu and Jun Zhang. Publication recommendation in incomplete networks based on graph learning. Scientometrics, 130(2):565–591, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05219-5.

Liu:2025:DCL

[LZ25b] Xueying Liu and Haoran Zhu. The diachronic change in linguistic positivity in the academic book reviewing of language studies: a text-mining analysis. Scientometrics, 130(1):133–157, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05208-8.

Liang:2021:CPA

[LZC21] Liming Liang, Zhen. Zhong, and Yue Chen. A Chinese professor’s academic career rhythm. Scientometrics, 126 (7):6169–6186, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03767-0.

Liu:2024:DLM

[LZD^{+}24] Zhaobin Liu, Yongxiang Zhang, Weiwei Deng, Jian Ma, and Xia Fan. A deep learning method for recommending university patents to industrial clusters by common technological needs mining. Scientometrics, 129(6):3089–3113, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05052-w.

Liu:2024:MCT

[LZF^{+}24] Peng Liu, Wei Zhou, Lijie Feng, Jinfeng Wang, Kuo-Yi Lin, Xuan Wu, and Dingtang Zhang. Mapping and comparing the technology evolution paths of scientific papers and patents: an integrated approach for forecasting technology trends. Scientometrics, 129(4):1975–2005, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04961-0.

Li:2020:GKT

[LZH20] Xian Li, Dangzhi Zhao, and Xiaojun Hu. Gatekeepers in knowledge transfer between science and technology: an exploratory study in the area of gene editing. Scientometrics, 124(2):1261–1277, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03537-y.

Li:2022:RCJ

[LZH^{+}22] Xiang Li, Chengli Zhao, Zhaolong Hu, Caixia Yu, and Xiaojun Duan. Revealing the character of journals in higherorder citation networks. Scientometrics, 127(11):6315– 6338, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04518-z.

Liu:2020:DSB

[LZL20a] YiJun Liu, Li Zhang, and Xiaoli Lian. A documentstructure-based complex network model for extracting text keywords. Scientometrics, 124(3):1765–1791, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03542-1.

Lyu:2020:ESM

[LZL20b] Xiaozan Lyu, Ping Zhou, and Loet Leydesdorff. Ecosystem mapping of techno-science linkages at the level of scholarly journals and fields. Scientometrics, 124(3):2037– 2055, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03435-3.

Liu:2022:OMC

[LZL22] Zheng Liu, Jialing Zhang, and Yun Li. One-to-many comparative summarization for patents. Scientometrics, 127 (4):1969–1993, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04307-8.

Liu:2022:SEV

[LZY22] Yun Liu, Mengya Zhang, and Xiongxiong You. Scientific elites versus other scientists: who are better at taking advantage of the research collaboration network? Scientometrics, 127(6):3145–3166, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04362-1.

Li:2024:WOP

[LZYR24] Li Li, Jiandong Zhang, Liying Yang, and Ronald Rousseau. Which older publications are still highly cited in the field of bibliometrics? Contemporary bibliometric citation classics. Scientometrics, 129(8):5037–5046, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05102-3.

Moradi:2021:PPC

[MA21] Shima Moradi and Sajedeh Abdi. Pandemic publication: correction and erratum in COVID-19 publications. Scientometrics, 126(2):1849–1857, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03787-w;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03787-w.pdf.

Maddi:2020:MOA

[Mad20] Abdelghani Maddi. Measuring open access publications: a novel normalized open access indicator. Scientometrics, 124(1):379–398, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03470-0.

Maleki:2022:OLH

[Mal22a] Ashraf Maleki. OCLC library holdings: assessing availability of academic books in libraries in print and electronic compared to citations and altmetrics. Scientometrics, 127(2):991–1020, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04220-6.

Maleki:2022:WDL

[Mal22b] Ashraf Maleki. Why does library holding format really matter for book impact assessment?: Modelling the relationship between citations and altmetrics with print and electronic holdings. Scientometrics, 127(2):1129– 1160, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04239-9.

Mora-Apablaza:2022:PIT

[MAN22] Loreto Mora-Apablaza and Carlos Navarrete. Patents as indicators of the technological position of countries on a global level? Scientometrics, 127(3):1233–1246, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04268-y.

Manganaro:2024:TLB

[Man24] Lorenzo Manganaro. The true latency of biomedical research papers. Scientometrics, 129(5):2897–2910, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05008-0.

Marrone:2020:AEL

[Mar20] Mauricio Marrone. Application of entity linking to identify research fronts and trends. Scientometrics, 122 (1):357–379, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-019-
03274-x;   http://link.springer.com/content/pdf/
10.1007/s11192-019-03274-x.pdf.

Maral:2024:RPH

[Mar24] Muammer Maral. Research performance of higher education institutions in Türkiye: 1980–2022. Scientometrics, 129(8):4771–4793, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05097-x.

Mason:2020:AUA

[Mas20] Shannon Mason. Adoption and usage of Academic Social Networks: a Japan case study. Scientometrics, 122(3): 1751–1767, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03345-4.

Mubin:2021:NNC

[MAS^{+}21] Omar Mubin, Fady Alnajjar, Abdullah Shamail, Suleman Shahid, and Simeon Simoff. The new norm: Computer Science conferences respond to COVID-19. Scientometrics, 126(2):1813–1827, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03788-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03788-9.pdf.

Maltseva:2020:IDD

[MB20a] Daria Maltseva and Vladimir Batagelj. iMetrics: the development of the discipline with many names. Scientometrics, 125(1):313–359, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03604-4.

Maltseva:2020:TSD

[MB20b] Daria Maltseva and Vladimir Batagelj. Towards a systematic description of the field using keywords analysis: main topics in social networks. Scientometrics, 123(1): 357–382, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03365-0.

Maltseva:2021:JPS

[MB21] Daria Maltseva and Vladimir Batagelj. Journals publishing social network analysis. Scientometrics, 126(4): 3593–3620, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-021-03889-z.

Maddi:2022:QWS

[MB22a] Abdelghani Maddi and Lesya Baudoin. The quality of the web of science data: a longitudinal study on the completeness of authors-addresses links. Scientometrics, 127(11):6279–6292, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04525-0.

Maltseva:2022:CBA

[MB22b] Daria Maltseva and Vladimir Batagelj. Collaboration between authors in the field of social network analysis. Scientometrics, 127(6):3437–3470, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04364-z.

Mazarakis:2025:CEG

[MBD25a] Athanasios Mazarakis, Paula Bräuer, and Isabelle Dorsch. Correction to: Evaluation of gamification as a tool for open access publishing among researchers: insights from a conjoint analysis. Scientometrics, 130(4):2493–2495, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05284-4. See [MBD25b].

Mazarakis:2025:EGT

[MBD25b] Athanasios Mazarakis, Paula Bräuer, and Isabelle Dorsch. Evaluation of gamification as a tool for open access publishing among researchers: insights from a conjoint analysis. Scientometrics, 130(2):1253–1280, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05226-6. See correction [MBD25a].

Matveeva:2023:SCP

[MBF23] Nataliya Matveeva, Vladimir Batagelj, and Anuska Ferligoj. Scientific collaboration of post-Soviet countries: the effects of different network normalizations. Scientometrics, 128(8):4219–4242, August 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-023-04752-z.

Mutz:2025:HUA

[MBH25] Rüdiger Mutz, Lutz Bornmann, and Robin Haunschild. How to use assignments of United Nations sustainable development goals (SDGs) to scientific papers in research evaluation? The proposal of a gold standard combining assignments from different data providers. Scientometrics, 130(3):1519–1546, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05254-w.

Ma:2021:UPD

[MBL21] Yaxue Ma, Zhichao Ba, and Gang Li. Understanding and predicting the dissemination of scientific papers on social media: a two-step simultaneous equation modelingartificial neural network approach. Scientometrics, 126(8): 7051–7085, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04051-5.

Mocanu:2025:ECE

[MBR^{+}25] Mihaela Mocanu, Anca-Diana Bibiri, Valentina Diana Rusu, Alina Moroşanu, and Iustinian Gabriel Bejan. Enhancing civic engagement with science: a comparative approach across European regions. Scientometrics, 130(1): 447–468, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05198-7.

McCain:2024:CPL

[McC24] Katherine W. McCain. Collaboration at the phylum level: coauthorship and acknowledgment patterns in the world of the water bears (phylum Tardigrada). Scientometrics, 129(10):6089–6125, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05036-w.

Marques-Cruz:2024:TYC

[MCDFSP24] Manuel Marques-Cruz, Daniel Martinho Dias, João A. Fonseca, and Bernardo Sousa-Pinto. Ten year citation prediction model for systematic reviews using early years citation data. Scientometrics, 129(8):4847–4862, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05105-0.

Marco-Cuenca:2021:FSP

[MCSOAA21] Gonzalo Marco-Cuenca, José Antonio Salvador-Oliván, and Rosario Arquero-Avilés. Fraud in scientific publications in the European Union. An analysis through their retractions. Scientometrics, 126(6):5143–5164, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03977-0.

Moreno-Delgado:2024:MSM

[MDCBdGVMD24] Alicia Moreno-Delgado, Marlon Cárdenas-Bonett, Óscar de Gregorio-Vicente, and Julio Montero-Díaz. Mapping scientific mobility in leading Eurozone economies: insights from ORCID data analysis. Scientometrics, 129(11):6889– 6907, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05153-6.

Moreno-Delgado:2021:APO

[MDGR21] Alicia Moreno-Delgado, Juan Gorraiz, and Rafael Repiso. Assessing the publication output on country level in the research field communication using Garfield’s Impact Factor. Scientometrics, 126(7):5983–6000, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04006-w.

Moed:2021:BAN

[MdMAGB^{+}21] Henk F. Moed, Felix de Moya-Anegon, Vicente Guerrero-Bote, Carmen Lopez-Illescas, and Myroslava Hladchenko.

Bibliometric assessment of national scientific journals. Scientometrics, 126(4):3641–3666, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03883-5.

Mendes:2025:ASN

[Men25] Saulo Mendes. Assessment of subject-normalized comprehensiveness of research-intensive universities. Scientometrics, 130(1):109–131, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05178-x.

Meyer:2025:LLT

[Mey25a] Martin Meyer. Loet Leydesdorff, the Triple Helix and ways to capture societal impact — a tribute. Scientometrics, 130 (6):3139–3148, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05307-0.

Meyer:2025:UMP

[Mey25b] Peter B. Meyer. Using multinational patent data to measure a design change in early aviation. Scientometrics, 130(1):187–204, January 2025. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-024-05148-3.

Marques:2020:CNA

[MF20a] Isabel Cristina Panziera Marques and Mário Franco. Cooperation networks in the area of health: systematic literature review. Scientometrics, 122(3):1727–1750, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03341-3.

Matveeva:2020:SCR

[MF20b] Nataliya Matveeva and Anuska Ferligoj. Scientific collaboration in Russian universities before and after the excellence initiative Project 5-100. Scientometrics, 124 (3):2383–2407, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03602-6.

Mrowinski:2020:HAP

[MFF^{+}20] Maciej J. Mrowinski, Agata Fronczak, Piotr Fronczak, Olgica Nedic, and Aleksandar Dekanski. The hurdles of academic publishing from the perspective of journal editors: a case study. Scientometrics, 125(1):115–133, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03619-x;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03619-x.pdf.

Moschini:2020:CTM

[MFM20] Ugo Moschini, Elena Fenialdi, and Elisa Molinari. A comparison of three multidisciplinarity indices based on the diversity of Scopus subject areas of authors’ documents, their bibliography and their citing papers. Scientometrics, 125(2):1145–1158, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03481-x.

Mammola:2021:IRL

[MFMC21] Stefano Mammola, Diego Fontaneto, Alejandro Martínez, and Filipe Chichorro. Impact of the reference list features on the number of citations. Scientometrics, 126 (1):785–799, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03759-0;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03759-0.pdf.

Mugnaini:2021:OTB

[MFP21] Rogério Mugnaini, Grischa Fraumann, and Abel L. Packer. Openness trends in Brazilian citation data: factors related to the use of DOIs. Scientometrics, 126(3):2523–2556, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03663-7.

Milia:2022:LEG

[MGA22] Matias Federico Milia, Ariadna Nebot Giralt, and Rigas Arvanitis. Local emergence, global expansion: understanding the structural evolution of a bi-lingual national research landscape. Scientometrics, 127(12):7369–7395, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04403-9.

Miguel:2024:TNA

[MGCR24] Sandra Miguel, Claudia M. González, and Zaida Chinchilla-Rodríguez. Towards a new approach to analyzing the geographical scope of national research. an exploratory analysis at the country level. Scientometrics, 129(7):3659–3679, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05045-9.

Ma:2024:PSM

[MH24a] Xin Ma and Tao Huang. Proximity still matters in research collaboration! Evidence from the introduction of new airline routes and high-speed railways in China. Scientometrics, 129(4):2227–2253, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04910-3.

Maleki:2024:TRS

[MH24b] Ashraf Maleki and Kim Holmberg. Tweeting and retweeting scientific articles: implications for altmetrics. Scientometrics, 129(10):6197–6220, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05127-8.

Mariethoz:2021:ICN

[MHD21a] Gregoire Mariethoz, Frédéric Herman, and Amelie Dreiss. The imaginary carrot: no correlation between raising funds and research productivity in geosciences. Scientometrics, 126(3):2401–2407, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03855-1.

Mariethoz:2021:RCH

[MHD21b] Gregoire Mariethoz, Frédéric Herman, and Amelie Dreiss. Reply to the comment by Heyard et al. titled “Imaginary carrot or effective fertiliser? A rejoinder on funding and productivity”. Scientometrics, 126(11):9339– 9342, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04131-6.

See [HPH21].

Mryglod:2022:BFS

[MHK22] O. Mryglod, Yu. Holovatch, and R. Kenna. Big fish and small ponds: why the departmental h-index should not be used to rank universities. Scientometrics, 127(6): 3279–3292, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04373-y.

Milojevic:2020:NSP

[Mil20] Stasa Milojević. Nature, Science, and PNAS: disciplinary profiles and impact. Scientometrics, 123(3):1301–1315, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03441-5.

Milojevic:2025:SS

[Mil25] Staša Milojević. Science of science. Scientometrics, 130 (6):3183–3193, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05322-1.

Minondo:2022:CW

[Min22] Asier Minondo. Comments are welcome. Scientometrics, 127(3):1565–1582, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04267-z.

Min:2025:CBE

[Min25] Jungwon Min. Community broker effects: evidence from Japanese research networks. Scientometrics, 130(3):1469– 1496, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05261-x.

Mitchell:2020:CPW

[Mit20] Neil C. Mitchell. Comparing the post-WWII publication histories of oceanography and marine geoscience. Scientometrics, 124(2):843–866, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03498-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03498-2.pdf.

Miyata:2020:KST

[MIY^{+}20] Yosuke Miyata, Emi Ishita, Fang Yang, Michimasa Yamamoto, Azusa Iwase, and Keiko Kurata. Knowledge structure transition in library and information science: topic modeling and visualization. Scientometrics, 125 (1):665–687, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03657-5;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03657-5.pdf.

Mervar:2022:CPN

[MJ22] Andrea Mervar and Maja Jokić. Core-periphery nexus in the EU social sciences: bibliometric perspective. Scientometrics, 127(10):5793–5817, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04493-5. See correction [MJ23].

Mervar:2023:CCP

[MJ23] Andrea Mervar and Maja Jokić. Correction: Coreperiphery nexus in the EU social sciences: bibliometric perspective. Scientometrics, 128(7):4165, July 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04531-2. See [MJ22].

Munoz-Jordan:2025:ELA

[MJRR25] David Muñoz-Jordán, Gonzalo Ruiz, and Alejandro Rivero. 3SA: an entity-linking algorithm for the Institution Name Disambiguation problem in affiliations using edit distance. Scientometrics, 130(7):4061–4071, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05368-1.

Moshtagh:2023:MEA

[MJYS23] Maryam Moshtagh, Tahereh Jowkar, Maryam Yaghtin, and Hajar Sotudeh. The moderating effect of altmetrics on the correlations between single and multi-faceted university ranking systems: the case of THE and QS vs. Nature Index and Leiden. Scientometrics, 128(1):761–781, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04548-7.

Masoumi:2023:FCE

[MK23] Najmeh Masoumi and Reza Khajavi. A fuzzy classifier for evaluation of research topics by using keyword co-occurrence network and sponsors information. Scientometrics, 128(3):1485–1512, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04618-w.

Muhlroth:2023:ISL

[MKG23] Christian Mühlroth, Laura Kölbl, and Michael Grottke. Innovation signals: leveraging machine learning to separate noise from news. Scientometrics, 128(5):2649–2676, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04672-y.

Motohashi:2024:MSI

[MKI24] Kazuyuki Motohashi, Hitoshi Koshiba, and Kenta Ikeuchi. Measuring science and innovation linkage using text mining of research papers and patent information. Scientometrics, 129(4):2159–2179, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04949-w.

Molwitz:2023:FAR

[MKWB^{+}23] Isabel Molwitz, Sarah Keller, Liesa Wolf-Baldauf, Ann-Kathrin Ozga, Thai-An Nguyen, Ilka Wedekind, Jing Zhao, Elif Can, Minobu Kamo, and Jin Yamamura. Female author representation differs between journals from the United States of America, Europe, and Asia: a 10-year comparison of five medical disciplines. Scientometrics, 128(3):1583–1600, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04612-2.

Malaterre:2024:VHC

[ML24] Christophe Malaterre and Francis Lareau. Visualizing hidden communities of interest: a case-study analysis of topic-based social networks in astrobiology. Scientometrics, 129(10):6167–6181, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05047-7.

Michail:2023:JRA

[MLA^{+}23] Seth Michail, Joseph William Ledet, Taha Yiğit Alkan, Muhammed Numan İnce, and Melih Günay. A journal recommender for article submission using transformers. Scientometrics, 128(2):1321–1336, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04609-x.

Ma:2021:DLB

[MLD21] Anqi Ma, Yu Liu, and Tao Dong. A deep-learning based citation count prediction model with paper metadata semantic features. Scientometrics, 126(8):6803–6823, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04033-7.

Marrone:2022:ELS

[MLK22] Mauricio Marrone, Sascha Lemke, and Lutz M. Kolbe. Entity linking systems for literature reviews. Scientometrics, 127(7):3857–3878, July 2022. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04423-5.

Maynard:2020:UOM

[MLP^{+}20] Diana Maynard, Benedetto Lepori, Johann Petrak, Xingyi Song, and Philippe Laredo. Using ontologies to map between research data and policymakers’ presumptions: the experience of the KNOWMAK project. Scientometrics, 125(2):1275–1290, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03664-6;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03664-6.pdf.

Maddi:2021:OAE

[MLS21] Abdelghani Maddi, Esther Lardreau, and David Sapinho. Open access in Europe: a national and regional comparison. Scientometrics, 126(4):3131–3152, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03887-1.

Modic:2023:SUL

[MLYK23] Dolores Modic, Borut Luzar, and Tohru Yoshioka-Kobayashi. Structure of university licensing networks. Scientometrics, 128(2):901–932, February 2023. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-022-04564-7.

Maddi:2024:PRR

[MM24] Abdelghani Maddi and Luis Miotti. On the peer review reports: does size matter? Scientometrics, 129(10):5893– 5913, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04977-6.

Manz:2025:ESH

[MM25] Niklas Manz and Ian McCullough. Eponyms in science: how long can they get? Scientometrics, 130(7):3433–3453, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05328-9.

Mironescu:2023:RDM

[MMB23] Andreea Mironescu, Alina Morosanu, and Anca-Diana Bibiri. The regional dynamics of multilingual publishing in Web of Science: a statistical analysis of Central and Eastern European journals and researchers in linguistics. Scientometrics, 128(2):1133–1162, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04595-0.

Maisano:2020:STE

[MMF20] Domenico A. Maisano, Luca Mastrogiacomo, and Fiorenzo Franceschini. Short-term effects of non-competitive funding to single academic researchers. Scientometrics, 123(3): 1261–1280, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03449-x.

Maisano:2023:EER

[MMF23] Domenico A. Maisano, Luca Mastrogiacomo, and Fiorenzo Franceschini. Empirical evidence on the relationship between research and teaching in academia. Scientometrics, 128(8):4475–4507, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04770-x.

Maisano:2025:LSS

[MMFF25] D. A. Maisano, L. Mastrogiacomo, L. Ferrara, and F. Franceschini. A large-scale semi-automated approach for assessing document-type classification errors in bibliometric databases. Scientometrics, 130(3):1901–1938, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05244-y.

McGuigan:2021:UEN

[MMG21] Glenn S. McGuigan, Göktug Morçöl, and Travis Grosser. Using ego-network analyses to examine journal citations: a comparative study of public administration, political science, and business management. Scientometrics, 126(12):9345–9368, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04189-2.

McGuigan:2023:SNA

[MMG23] Glenn S. McGuigan, Göktug Morçöl, and Travis Grosser. A social network analysis of academic journals in public administration in the early twenty-first century: examining journal level bibliometrics with network analysis. Scientometrics, 128(12):6561–6588, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04861-9.

Maddi:2024:SEP

[MMGC^{+}24] Abdelghani Maddi, Emmanuel Monneau, Catherine Guaspare-Cartron, Floriana Gargiulo, and Michel Dubois. Streetlight effect in PubPeer comments: are Open Access publications more scrutinized? Scientometrics, 129(7): 4231–4247, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05053-9.

Moreira:2023:RHA

[MMGL23] Edré Moreira, Wagner Meira, Marcos André Gonçalves, and Alberto H. F. Laender. The rise of hyperprolific authors in computer science: characterization and implications. Scientometrics, 128(5):2945–2974, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04676-8.

Marhl:2025:CWD

[MMGP25] Marko Marhl, Rene Markovič, Vladimir Grubelnik, and Matjaž Perc. The changing world dynamics of research performance. Scientometrics, 130(1):469–488, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05199-6.

Markusova:2020:CAR

[MML^{+}20] Valentina Markusova, Levan Mindeli, Alexander Libkind, Anna Zolotova, and Mark Akoev. Comparative analysis of Russian and industrialized countries performance on energy and fuels, WoS, 2008–2017. Scientometrics, 123(3): 1281–1300, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03440-6.

Mohammed:2020:IUP

[MMN20] Shaibu Mohammed, Anthony Morgan, and Emmanuel Nyantakyi. On the influence of uncited publications on a researcher’s h-index. Scientometrics, 122(3):1791–1799, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03356-1.

Momeni:2021:WHW

[MMP21] Fakhri Momeni, Philipp Mayr, and Isabella Peters. What happens when a journal converts to open access? A bibliometric analysis. Scientometrics, 126(12):9811–9827, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03972-5.

Martin-Martin:2021:CGS

[MMTOMLC21a] Alberto Martín-Martín, Mike Thelwall, Enrique Orduna-Malea, and Emilio Delgado López-Cózar. Correction to: Google Scholar, Microsoft Academic, Scopus, Dimensions, Web of Science, and OpenCitations’ COCI: a multidisciplinary comparison of coverage via citations. Scientometrics, 126(1):907–908, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03792-z;   http:
//link.springer.com/content/pdf/10.1007/s11192-

020-03792-z.pdf. See [MMTOMLC21b].

Martin-Martin:2021:GSM

[MMTOMLC21b] Alberto Martín-Martín, Mike Thelwall, Enrique Orduna-Malea, and Emilio Delgado López-Cózar. Google Scholar, Microsoft Academic, Scopus, Dimensions, Web of Science, and OpenCitations’ COCI: a multidisciplinary comparison of coverage via citations. Scientometrics, 126(1):871–906, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03690-4;   http:
//link.springer.com/content/pdf/10.1007/s11192-

020-03690-4.pdf. See correction [MMTOMLC21a].

McManus:2021:FRB

[MN21a] Concepta McManus and Abilio Afonso Baeta Neves. Funding research in Brazil. Scientometrics, 126(1):801–823, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03762-5.

McManus:2021:PPB

[MN21b] Concepta McManus and Abilio Afonso Baeta Neves. Production profiles in Brazilian science, with special attention to social sciences and humanities. Scientometrics, 126(3): 2413–2435, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03452-2.

Mryglod:2023:LCC

[MN23] Olesya Mryglod and Serhii Nazarovets. Lost for the country: country-undefined papers in Web of Science and Scopus. Scientometrics, 128(4):2619–2622, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04661-1.

Mezzetti:2024:HBM

[MN24] Maura Mezzetti and Ilia Negri. Hierarchical Bayesian model to estimate and compare research productivity of Italian academic statisticians. Scientometrics, 129(12): 7443–7474, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-024-
05154-5.

Mryglod:2021:USF

[MNK21] O. Mryglod, S. Nazarovets, and S. Kozmenko. Universal and specific features of Ukrainian economic research: publication analysis based on Crossref data. Scientometrics, 126(9):8187–8203, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04079-7.

McManus:2020:ICB

[MNM^{+}20] Concepta McManus, Abilio Afonso Baeta Neves, Andrea Queiroz Maranhão, Antonio Gomes Souza Filho, and Jaime Martins Santana. International collaboration in Brazilian science: financing and impact. Scientometrics, 125(3):2745–2772, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03728-7;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03728-7.pdf.

Mohammed:2021:URE

[MNM^{+}21] Shaibu Mohammed, Emmanuel K. Nyantakyi, Anthony Morgan, Prosper Anumah, and Justice Sarkodiekyeremeh. Use of relative extra citation counts and uncited publications to enhance the discriminatory power of the h-index. Scientometrics, 126(1):181–199, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03777-y.

McManus:2021:SPN

[MNP21] Concepta McManus, Abilio Afonso Baeta Neves, and Alvaro Toubes Prata. Scientific publications from nonacademic sectors and their impact. Scientometrics, 126 (11):8887–8911, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04159-8.

Montes-Orozco:2024:ACN

[MOMGNLG24] Edwin Montes-Orozco, Karen Miranda, Abel García-Nájera, and Juan-Carlos López-García. On the analysis of collaboration networks between industry and academia: the Mexican case of the innovation incentive program. Scientometrics, 129(3):1523–1544, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04903-2.

Moorhead:2021:LUC

[Moo21] Althea V. Moorhead. Is L^{A}TEX use correlated with the number of equations in a manuscript? Scientometrics, 126(10):8259–8273, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04119-2.

Morillo:2019:CIR

[Mor19] Fernanda Morillo. Collaboration and impact of research in different disciplines with international funding (from the EU and other foreign sources). Scientometrics, 120(2): 807–823, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-019-03150-8.

Morillo:2020:OAP

[Mor20] Fernanda Morillo. Is open access publication useful for all research fields? Presence of funding, collaboration and impact. Scientometrics, 125(1):689–716, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03652-w.

Moustafa:2020:OA

[Mou20] Khaled Moustafa. Octopus affiliations. Scientometrics, 124(3):2733–2735, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03600-8.

Moussa:2021:FJR

[Mou21a] Salim Moussa. Are FT50 journals really leading? A comment on Fassin. Scientometrics, 126(12):9613–9622, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04158-9.

See [Fas21a].

Moussa:2021:CCCa

[Mou21b] Salim Moussa. Citation contagion: a citation analysis of selected predatory marketing journals. Scientometrics, 126(1):485–506, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03729-6. See correction [Mou21c].

Moussa:2021:CCCb

[Mou21c] Salim Moussa. Correction to: Citation contagion: a citation analysis of selected predatory marketing journals. Scientometrics, 126(1):507, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03817-7;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03817-7.pdf. See

[Mou21b].

Mammola:2022:MIN

[MPC22] Stefano Mammola, Elena Piano, and Dan Chamberlain. Measuring the influence of non-scientific features on citations. Scientometrics, 127(7):4123–4137, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04421-7.

Marchiori:2021:IBV

[MPMR21] Danilo Magno Marchiori, Silvio Popadiuk, Emerson Wagner Mainardes, and Ricardo Gouveia Rodrigues. Innovativeness: a bibliometric vision of the conceptual and intellectual structures and the past and future research directions. Scientometrics, 126(1):55–92, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03753-6. See correction [MPR21].

Marchiori:2021:CIB

[MPR21] Danilo Magno Marchiori, Silvio Popadiuk, and Ricardo Gouveia Rodrigues. Correction to: Innovativeness: a bibliometric vision of the conceptual and intellectual structures and the past and future research directions. Scientometrics, 126(11):9343, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03994-z. See [MPMR21].

Ma:2022:OOT

[MPS22a] Jing Ma, Yaohui Pan, and Chih-Yi Su. Organizationoriented technology opportunities analysis based on predicting patent networks: a case of Alzheimer’s disease.

Scientometrics, 127(9):5497–5517, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04219-z.

Martini:2022:RCN

[MPS22b] Maria Cristiana Martini, Elvira Pelle, and Andrea Sciandra. The role of citation networks to explain academic promotions: an empirical analysis of the Italian national scientific qualification. Scientometrics, 127(10):5633–5659, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04485-5.

Missen:2020:SAS

[MQS^{+}20] Malik Muhammad Saad Missen, Sajeeha Qureshi, Nadeem Salamat, Nadeem Akhtar, Hina Asmat, Mickaël Coustaty, and V. B. Surya Prasath. Scientometric analysis of social science and science disciplines in a developing nation: a case study of Pakistan in the last decade. Scientometrics, 123(1):113–142, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03379-8.

McKelvey:2020:ESP

[MR20] Maureen McKelvey and Bastian Rake. Exploring scientific publications by firms: what are the roles of academic and corporate partners for publications in high reputation or high impact journals? Scientometrics, 122(3): 1323–1360, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03344-5.

Matias-Rayme:2024:GAD

[MRBSFSJ24] Nataly Matias-Rayme, Iuliana Botezan, Mari Carmen Suárez-Figueroa, and Rodrigo Sánchez-Jiménez. Gender assignment in doctoral theses: revisiting Teseo with a method based on cultural consensus theory. Scientometrics, 129(7):4553–4572, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05079-z.

Mosleh:2022:SLR

[MRC22] Melika Mosleh, Saeed Roshani, and Mario Coccia. Scientific laws of research funding to support citations and diffusion of knowledge in life science. Scientometrics, 127 (4):1931–1951, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04300-1.

Mat:2021:TSD

[MRF21] Sharfah Ratibah Tuan Mat, Mohd Faizal Ab Razak, and Ahmad Firdaus. Towards a systematic description of the field using bibliometric analysis: malware evolution. Scientometrics, 126(3):2013–2055, March 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03834-6.

Merga:2020:SGS

[MRM20] Margaret K. Merga, Sayidi Mat Roni, and Shannon Mason. Should Google Scholar be used for benchmarking against the professoriate in education? Scientometrics, 125(3):2505–2522, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03691-3;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03691-3.pdf.

Machacek:2021:PPS

[MS21a] Vít Machácek and Martin Srholec. Predatory publishing in Scopus: evidence on cross-country differences. Scientometrics, 126(3):1897–1921, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s11192-020-03852-4. See retraction notice [MS22a] and comments [AAA^{+}23, Zha23a].

Marina:2021:CPP

[MS21b] Tatiana Marina and Ivan Sterligov. Correction to: Prevalence of potentially predatory publishing in Scopus on the country level. Scientometrics, 126(6):5079, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04042-6. See [MS21c].

Marina:2021:PPP

[MS21c] Tatiana Marina and Ivan Sterligov. Prevalence of potentially predatory publishing in Scopus on the country level. Scientometrics, 126(6):5019–5077, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03899-x. See correction [MS21b].

Mason:2021:RWI

[MS21d] Shannon Mason and Yusuke Sakurai. A ResearchGateway to an international academic community? Scientometrics, 126(2):1149–1171, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03772-3.

Mihaljevic:2021:DAE

[MS21e] Helena Mihaljević and Lucía Santamaría. Disambiguation of author entities in ADS using supervised learning and graph theory methods. Scientometrics, 126(5):3893–3917, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03951-w.

Miyashita:2021:SMS

[MS21f] Shuto Miyashita and Shintaro Sengoku. Scientometrics for management of science: collaboration and knowledge structures and complexities in an interdisciplinary research project. Scientometrics, 126(9):7419–7444, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04080-0.

Machacek:2022:RNP

[MS22a] Vít Machácek and Martin Srholec. Retraction note to: Predatory publishing in Scopus: evidence on cross-country differences. Scientometrics, 127(3):1667, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04149-w. See [MS21a].

Maddi:2022:APC

[MS22b] Abdelghani Maddi and David Sapinho. Article processing charges, altmetrics and citation impact: Is there an economic rationale? Scientometrics, 127(12):7351–7368, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04284-y.

Madison:2022:NPC

[MS22c] Guy Madison and Knut Sundell. Numbers of publications and citations for researchers in fields pertinent to the social services: a comparison of peer-reviewed journal publications across six disciplines. Scientometrics, 127(10): 6029–6046, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04495-3.

Mason:2022:WJB

[MS22d] Shannon Mason and Lenandlar Singh. When a journal is both at the ‘top’ and the ‘bottom’: the illogicality of conflating citation-based metrics with quality. Scientometrics, 127(6):3683–3694, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04402-w. See reply [JTB23].

Maddi:2023:COA

[MS23] Abdelghani Maddi and David Sapinho. On the culture of open access: the Sci-Hub paradox. Scientometrics, 128(10):5647–5658, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04792-5.

Matsumoto:2021:INI

[MSI21] Kuniko Matsumoto, Sotaro Shibayama, and Masatsura Igami. Introducing a novelty indicator for scientific research: validating the knowledge-based combinatorial approach. Scientometrics, 126(8):6891–6915, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04049-z.

Martincevic:2023:CCA

[MSJ23] Marina Martincević, Darja Maslić Sersić, and Davor Jokić. Contribution of CEE authors to psychological science: is the growing trend of publishing in non-CEE journals still present 10 years after its inception? Scientometrics, 128 (6):3703–3721, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04695-5.

Matveeva:2022:ISC

[MSL22] Nataliya Matveeva, Ivan Sterligov, and Andrey Lovakov. International scientific collaboration of post-Soviet countries: a bibliometric analysis. Scientometrics, 127(3): 1583–1607, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04274-0.

Moradzadeh:2023:CTN

[MSP23a] Mina Moradzadeh, Shahram Sedghi, and Sirous Panahi. Correction to: Towards a new paradigm for ‘journal quality’ criteria: a scoping review. Scientometrics, 128(1): 323, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04630-0.

See [MSP23b].

Moradzadeh:2023:TNP

[MSP23b] Mina Moradzadeh, Shahram Sedghi, and Sirous Panahi. Towards a new paradigm for ‘journal quality’ criteria: a scoping review. Scientometrics, 128(1):279–321, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04520-5. See correction

[MSP23a].

Maiti:2023:ASU

[MSV23] Aniruddha Maiti, Sai Shi, and Slobodan Vucetic. An ablation study on the use of publication venue quality to rank computer science departments. Scientometrics, 128(8): 4197–4218, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04733-2.

Ma:2023:DWS

[MTD^{+}23] Yongchao Ma, Ying Teng, Zhongzhun Deng, Li Liu, and Yi Zhang. Does writing style affect gender differences in the research performance of articles?: An empirical study of BERT-based textual sentiment analysis. Scientometrics, 128(4):2105–2143, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04666-w.

Mu:2021:MCA

[Mu21] Congjun Mu. A multidimensional contrastive analysis of linguistic features between international and local biology journal English research articles. Scientometrics, 126(9):7901–7916, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04102-x.

Mutz:2022:DIS

[Mut22] Rüdiger Mutz. Diversity and interdisciplinarity: Should variety, balance and disparity be combined as a product or better as a sum? An information-theoretical and statistical estimation approach. Scientometrics, 127(12):7397– 7414, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04336-3.

Meng:2020:GHS

[MWH20] Liang Meng, Haifeng Wang, and Pengfei Han. Getting a head start: turn-of-the-month submission effect for accepted papers in management journals. Scientometrics, 124(3):2577–2595, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03556-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03556-9.pdf.

Ma:2025:TDF

[MY25] Jianbing Ma and Kexin Yang. A three-dimensional framework for quantifying knowledge intersection intensity: from a micro perspective. Scientometrics, 130(1):367–398, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05222-w.

Ma:2022:RFM

[MYK22] Yonghong Ma, Xiaomeng Yang, and Lingkai Kong. Research on the formation mechanism of big data technology cooperation networks: empirical evidence from China. Scientometrics, 127(3):1273–1294, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04270-4.

Mo:2022:EFE

[MYX22] Zhou Mo, Zhang Yujie, and Tan Xiaowen. Early firm engagement, government research funding, and the privatization of public knowledge. Scientometrics, 127(8): 4797–4826, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04448-w.

Yuan:2022:AWB

[mYY22] Zhou min Yuan and Mingxin Yao. Is academic writing becoming more positive? A large-scale diachronic case study of Science research articles across 25 years. Scientometrics, 127(11):6191–6207, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04515-2.

Ma:2022:EIS

[MZD22] Bowen Ma, Chengzhi Zhang, and Sanhong Deng. Enhancing identification of structure function of academic articles using contextual information. Scientometrics, 127(2): 885–925, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04225-1.

Montazerian:2020:PVV

[MZE20] Maziar Montazerian, Edgar Dutra Zanotto, and Hellmut Eckert. Prolificacy and visibility versus reputation in the hard sciences. Scientometrics, 123(1):207–221, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03369-w.

Ma:2020:RCR

[MZL20] Shutian Ma, Chengzhi Zhang, and Xiaozhong Liu. A review of citation recommendation: from textual content to enriched context. Scientometrics, 122(3):1445–1472, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03336-0.

Ma:2021:CCR

[MZZL21] Shutian Ma, Heng Zhang, Chengzhi Zhang, and Xiaozhong Liu. Chronological citation recommendation with time preference. Scientometrics, 126(4):2991–3010, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-021-03878-2.

Ningrum:2024:ASU

[NA24] Panggih Kusuma Ningrum and Iana Atanassova. Annotation of scientific uncertainty using linguistic patterns. Scientometrics, 129(10):6261–6285, October 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05009-z.

Nazarovets:2020:CPR

[Naz20] Serhii Nazarovets. Controversial practice of rewarding for publications in national journals. Scientometrics, 124 (1):813–818, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03485-7.

Nazarovets:2024:PGP

[Naz24a] Serhii Nazarovets. Paradoxical growth of publications by authors from developing countries in gold open access journals: a commentary on Dr. Cernat’s, 2024 article. Scientometrics, 129(12):7981–7984, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05186-x. See [Cer24b] and response

[Cer24a].

Nazarovets:2024:VRC

[Naz24b] Serhii Nazarovets. Visualization of rank-citation curves for fast detection of h-index anomalies in university metrics. Scientometrics, 129(1):705–711, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04886-0.

Nazarovets:2025:ASP

[Naz25] Serhii Nazarovets. Acknowledgments in scientific papers by Ukrainian researchers during the initial years of the Russo– Ukrainian war. Scientometrics, 130(2):755–762, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05235-z.

Ngwenya:2020:PIN

[NB20] Similo Ngwenya and Nelius Boshoff. Participation of ‘international national organisations’ in Africa’s research: a bibliometric study of agriculture and health in Zimbabwe. Scientometrics, 124(1):533–553, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03480-y.

Ngwenya:2022:DMC

[NB22] Similo Ngwenya and Nelius Boshoff. Different manifestations of ‘context’: examples from a bibliometric study of research in Zimbabwe in Southern Africa. Scientometrics, 127(7):3911–3933, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04435-1.

Nelson:2024:HDA

[NBBL24] John P. Nelson, Barry Bozeman, Stuart Bretschneider, and Spencer L. Lindsay. How do academic public administration and public policy researchers affect policymaking? Functional groupings from survey data. Scientometrics, 129(1):65–93, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04860-w.

Nguyen:2020:SKP

[NC20] Chi Mai Nguyen and Jae-Yong Choung. Scientific knowledge production in China: a comparative analysis. Scientometrics, 124(2):1279–1303, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03507-4.

Nguyen:2025:LLE

[NC25] Van Thien Nguyen and Rene Carraz. “Exploring academic patent-paper pairs: a new methodology for analyzing Japan’s research landscape”. Scientometrics, 130(3): 1329–1356, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05275-5.

Nugent:2022:GFU

[NCD22] Annita Nugent, Ho Fai Chan, and Uwe Dulleck. Government funding of university-industry collaboration: exploring the impact of targeted funding on university patent activity. Scientometrics, 127(1):29–73, January 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04153-0.

Nicolaisen:2021:NRL

[NF21] Jeppe Nicolaisen and Tove Faber Frandsen. Number of references: a large-scale study of interval ratios. Scientometrics, 126(1):259–285, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03764-3.

Nicolaisen:2022:ECF

[NF22] Jeppe Nicolaisen and Tove Faber Frandsen. Epistemic community formation: a bibliometric study of recurring authors in medical journals. Scientometrics, 127(7): 4167–4189, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04409-3.

Numajiri:2024:AOD

[NH24] Honami Numajiri and Takayuki Hayashi. Analysis on open data as a foundation for data-driven research. Scientometrics, 129(10):6315–6332, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04956-x.

Ngayua:2021:AAT

[NHAB21] Esther Nanzayi Ngayua, Jianjia He, and Kwabena Agyei-Boahene. Applying advanced technologies to improve clinical trials: a systematic mapping study. Scientometrics, 126(2):1217–1238, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03774-1;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03774-1.pdf.

Nikolic:2024:OST

[NII24] Dusan Nikolić, Dragan Ivanović, and Lidija Ivanović. An open-source tool for merging data from multiple citation databases. Scientometrics, 129(7):4573–4595, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05076-2.

Nishikawa:2023:CHW

[Nis23a] Kai Nishikawa. Correction: How and why are citations between disciplines made? A citation context analysis focusing on natural sciences and social sciences and humanities. Scientometrics, 128(5):2999, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04701-w. See [Nis23b].

Nishikawa:2023:HWC

[Nis23b] Kai Nishikawa. How and why are citations between disciplines made? A citation context analysis focusing on natural sciences and social sciences and humanities. Scientometrics, 128(5):2975–2997, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04664-y. See correction [Nis23a].

Nogueira:2020:NBC

[NJCL20] Rodrigo Nogueira, Zhiying Jiang, Kyunghyun Cho, and Jimmy Lin. Navigation-based candidate expansion and pretrained language models for citation recommendation. Scientometrics, 125(3):3001–3016, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03718-9.

Nishikawa:2024:EAL

[NK24] Kai Nishikawa and Hitoshi Koshiba. Exploring the applicability of large language models to citation context analysis. Scientometrics, 129(11):6751–6777, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05142-9.

Nam:2024:EKE

[NKY^{+}24] Dongin Nam, Jiwon Kim, Jeeyoung Yoon, Chaemin Song, Seongdeok Kim, and Min Song. Examining knowledge entities and its relationships based on citation sentences using a multi-anchor bipartite network. Scientometrics, 129(11):7197–7228, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04824-0.

Nan:2020:UAA

[NLS20] Xia Nan, Ming Li, and Jin Shi. Using altmetrics for assessing impact of highly-cited books in Chinese Book Citation Index. Scientometrics, 122(3):1651–1669, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03347-2.

Nandy:2024:IVS

[NLS24] Abhirup Nandy, Hiran H. Lathabai, and Vivek Kumar Singh. xd-index and its variants: a set of overall scholarly expertise diversity indices for the research portfolio management of institutions. Scientometrics, 129(10): 5937–5962, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05131-y.

Nishikawa:2025:DOA

[NM25] Kai Nishikawa and Akiyoshi Murakami. Does open access foster interdisciplinary citations? Decomposing open access citation advantage. Scientometrics, 130(5):2947–2961, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05297-z.

Nane:2023:CSP

[NRGvSTS23] Gabriela F. Nane, Nicolas Robinson-Garcia, François van Schalkwyk, and Daniel Torres-Salinas. COVID-19 and the scientific publishing system: growth, open access and scientific fields. Scientometrics, 128(1):345–362, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04536-x.

Najmaei:2023:GSB

[NS23] Arash Najmaei and Zahra Sadeghinejad. Green and sustainable business models: historical roots, growth trajectory, conceptual architecture and an agenda for future research — a bibliometric review of green and sustainable business models. Scientometrics, 128(2):957–999, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04577-2.

Nakajima:2023:HOR

[NSM23] Kazuki Nakajima, Kazuyuki Shudo, and Naoki Masuda. Higher-order rich-club phenomenon in collaborative research grant networks. Scientometrics, 128(4):2429–2446, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04621-1.

Nishikawa:2024:IIP

[NSM24] Masaru Nishikawa, Daisuke Sakai, and Akira Matsui. The impact of the internationalization of political science on publishing in two languages: the case of Japan, 1971–2023. Scientometrics, 129(11):6975–7003, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05179-w.

Nazarovets:2024:IMT

[NT24] Serhii Nazarovets and Jaime A. Teixeira da Silva. The incidence of 11 movie titles in the titles of Scopus-indexed papers. Scientometrics, 129(5):2911–2922, May 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05004-4.

Nuzzo:2020:LSD

[Nuz20] James L. Nuzzo. Large sex difference despite equal opportunity: authorship of over 3000 letters in exercise science and physical therapy journals over 56 40years. Scientometrics, 124(1):679–695, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03427-3.

Nuzzo:2021:LEE

[Nuz21] James L. Nuzzo. Letters to the editor in exercise science and physical therapy journals: an examination of content and “authorship inflation”. Scientometrics, 126(8): 6917–6936, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04068-w.

Nakhoda:2023:QAU

[NWZ23] Maryam Nakhoda, Peter Whigham, and Sander Zwanenburg. Quantifying and addressing uncertainty in the measurement of interdisciplinarity. Scientometrics, 128(11): 6107–6127, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-
04822-2.

Niu:2023:CLP

[NZ23] Fenggao Niu and Yating Zhao. Citation link prediction based on multi-relational neural topic model. Scientometrics, 128(9):5277–5292, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04766-7.

Ni:2021:IOP

[NZL21] Jue Ni, Zhenyue Zhao, and Jiang Li. The influence of opening up peer review on the citations of journal articles. Scientometrics, 126(12):9393–9404, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04182-9.

Osman:2024:BAS

[OA24] Ismail Ragab Osman and Hendy Abdullah Hendy Ahmed. Bibliometric analysis of Sci-Hub downloads by Egyptian researchers. Scientometrics, 129(4):2371–2388, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04951-2.

Osorio:2021:DPS

[OB21] António Osório and Lutz Bornmann. On the disruptive power of small-teams research. Scientometrics, 126(1): 117–133, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03756-3.

Oh:2020:PPD

[OCKY20] Seunghyun Oh, Jaewoong Choi, Namuk Ko, and Janghyeok Yoon. Predicting product development directions for new product planning using patent classification-based link prediction. Scientometrics, 125(3):1833–1876, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03709-w.

Oliveira:2020:MDR

[ODF20] Thiago Dumont Oliveira and Marwil J. Dávila-Fernández. From modelmania to datanomics? The rise of mathematical and quantitative methods in three top economics journals. Scientometrics, 123(1):51–70, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03375-y.

Ortega:2024:IRL

[ODQ24] José Luis Ortega and Lorena Delgado-Quirós. The indexation of retracted literature in seven principal scholarly databases: a coverage comparison of dimensions, OpenAlex, PubMed, Scilit, Scopus, The Lens and Web of Science. Scientometrics, 129(7):3769–3785, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05034-y.

Okagbue:2020:CBC

[OdS20] Hilary I. Okagbue and Jaime A. Teixeira da Silva. Correlation between the CiteScore and Journal Impact Factor of top-ranked library and information science journals. Scientometrics, 124(1):797–801, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03457-x.

Okagbue:2020:DDI

[OdSO20] Hilary I. Okagbue, Jaime A. Teixeira da Silva, and Abiodun A. Opanuga. Disparities in document indexation in two databases (Scopus and Web of Science) among six subject domains, and the impact on journal-based metrics. Scientometrics, 125(3):2821–2825, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03704-1.

Oldford:2023:MEJ

[OFA23] Erin Oldford, John Fiset, and Anahit Armenakyan. The marginalizing effect of journal submission fees in Accounting and Finance. Scientometrics, 128(8):4611–4650, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04758-7.

Oviedo-Garcia:2024:RMJ

[OG24] M. Ángeles Oviedo-García. The review mills, not just (self-)plagiarism in review reports, but a step further. Scientometrics, 129(9):5805–5813, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05125-w.

Oh:2023:MPA

[OJKY23] Myeongji Oh, Hyejin Jang, Sunhye Kim, and Byungun Yoon. Main path analysis for technological development using SAO structure and DEMATEL based on keyword causality. Scientometrics, 128(4):2079–2104, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04652-2.

Oldac:2022:GSM

[Old22] Yusuf Ikbal Oldac. Global science and the Muslim world: overview of Muslim-majority country contributions to global science. Scientometrics, 127(11):6231–6255, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04517-0.

Orduna-Malea:2020:CAO

[OM20] Enrique Orduña-Malea. Crossing the academic ocean? Judit Bar-Ilan’s oeuvre on search engines studies. Scientometrics, 123(3):1317–1340, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03450-4.

Orduna-Malea:2021:DST

[OM21] Enrique Orduña-Malea. Dot-science top level domain: Academic websites or dumpsites? Scientometrics, 126(4): 3565–3591, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03832-8.

Orduna-Malea:2023:EOI

[OMAAORCL23] Enrique Orduña-Malea, Adolfo Alonso-Arroyo, José-Antonio Ontalba-Ruipérez, and Ferrán Catalá-López. Evaluating the online impact of reporting guidelines for randomised trial reports and protocols: a cross-sectional web-based data analysis of CONSORT and SPIRIT initiatives. Scientometrics, 128(1):407–440, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04542-z.

Orduna-Malea:2024:RAU

[OMBP24] Enrique Orduña-Malea and Núria Bautista-Puig. Research assessment under debate: disentangling the interest around the DORA declaration on Twitter. Scientometrics, 129(1):537–559, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04872-6.

Orduna-Malea:2021:LBA

[OMC21] Enrique Orduña-Malea and Rodrigo Costas. Link-based approach to study scientific software usage: the case of VOSviewer. Scientometrics, 126(9):8153–8186, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04082-y.

Orduna-Malea:2023:CPG

[OMCC23] Enrique Orduña-Malea and Álvaro Cabezas-Clavijo. Chat-GPT and the potential growing of ghost bibliographic references. Scientometrics, 128(9):5351–5355, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04804-4.

Orduna-Malea:2022:PLT

[OMFJ22] Enrique Orduña-Malea and Cristina I. Font-Julián. Are patents linked on Twitter? A case study of Google patents. Scientometrics, 127(11):6339–6362, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04519-y.

Orduna-Malea:2024:OAP

[OMFJSC24] Enrique Orduña-Malea, Cristina I. Font-Julián, and Jorge Serrano-Cobos. Open access publications drive few visits from Google Search results to institutional repositories. Scientometrics, 129(11):7131–7152, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05175-0.

Okamura:2020:CVD

[ON20] Asako Okamura and Keisuke Nishijo. Constructing vision-driven indicators to enhance the interaction between science and society. Scientometrics, 125(2):1575– 1589, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03598-z.

Onyancha:2020:MAS

[Ony20] Omwoyo Bosire Onyancha. A meta-analysis study of the relationship between research and economic development in selected countries in sub-Saharan Africa. Scientometrics, 123(2):655–675, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03390-z.

Oliva:2021:TSM

[OOCD^{+}21] Samuel Zanferdini Oliva, Livia Oliveira-Ciabati, Denise Gazotto Dezembro, Mário Sérgio Adolfi Júnior, Maísa de Carvalho Silva, Hugo Cesar Pessotti, and Juliana Tarossi Pollettini. Text structuring methods based on complex network: a systematic review. Scientometrics, 126(2):1471– 1493, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03785-y.

Ortiz-Pimentel:2020:BAP

[OPMRP20] Maximiano Ortiz-Pimentel, Carlos Molina, and Guillermo Armando Ronda-Pupo. Bibliometric assessment of papers on generations in management and business journals. Scientometrics, 125(1):445–469, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03670-8.

Odeyemi:2025:URF

[OPS25] Olumide A. Odeyemi, Yvonne Parry, and Nina Sivertsen. Unsuccessful research funding applications: a scoping review of causes and impacts on Australian researchers and research projects. Scientometrics, 130(5):2783–2797, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05330-1.

Olechnicka:2025:IVS

[OPZP25] Agnieszka Olechnicka, Adam Ploszaj, and Ewa Zegler-Poleska. The impact of the virtualization of scholarly conferences on the gender structure of conference contributors. Scientometrics, 130(1):423–445, January 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05212-y.

Ortega:2020:BNS

[Ort20] José Luis Ortega. Blogs and news sources coverage in altmetrics data providers: a comparative analysis by country, language, and subject. Scientometrics, 122(1):555– 572, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03299-2.

Oliveira:2023:HSM

[OSS^{+}23] Gabriel P. Oliveira, Mariana O. Silva, Danilo B. Seufitelli, Gabriel R. G. Barbosa, Bruna C. Melo, and Mirella M. Moro. Hot streaks in the music industry: identifying and characterizing above-average success periods in artists’ careers. Scientometrics, 128(11):6029–6046, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04835-x.

Orosnjak:2024:RRC

[OŠV^{+}24] Marko Orošnjak, Branko Štrbac, Sr– dan Vulanović, Biserka Runje, Amalija Horvatić Novak, and Andrej Razumić. RCE (rationale-cogency-extent) criterion unravels features affecting citation impact of top-ranked systematic literature reviews: leaving the impression … is all you need. Scientometrics, 129(3):1891–1947, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04935-2.

Ozturk:2024:HMB

[ÖT24] Onur Öztürk and Zehra Taskin. How metric-based performance evaluation systems fuel the growth of questionable publications? Scientometrics, 129(5):2729–2748, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04991-8.

Ohniwa:2022:RDG

[OTH22] Ryosuke L. Ohniwa, Kunio Takeyasu, and Aiko Hibino. Researcher dynamics in the generation of emerging topics in life sciences and medicine. Scientometrics, 127(2): 871–884, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04233-1.

Ovezmyradov:2023:AQI

[Ove23] Berdymyrat Ovezmyradov. Applying quantified indicators in Central Asian science: can metrics improve the regional research performance? Scientometrics, 128(1): 177–206, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04544-x.

Ozcan:2025:TAS

[OVS25] Sercan Ozcan, Dominik Brian Vogel, and Ozcan Saritas. Technological adoptions and sector-specific innovations in a low-tech environment: key actors and sources of R&D in InsurTech. Scientometrics, 130(1):1–20, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05209-7.

Oz:2024:OAD

[Öz24] Ali Baris Öz. Overcoming alphabetical disadvantage: factors influencing the use of surname initial techniques and their impact on citation rates in the four major disciplines of social sciences. Scientometrics, 129(8):4885–4908, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05100-5.

Petersen:2021:ATA

[PA21] Kai Petersen and Nauman Bin Ali. An analysis of top author citations in software engineering and a comparison with other fields. Scientometrics, 126(11):9147– 9183, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04144-1.

Polonen:2022:RPS

[PA22] Janne Pölönen and Otto Auranen. Research performance and scholarly communication profile of competitive research funding: the case of Academy of Finland. Scientometrics, 127(12):7415–7433, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04385-8.

Pal:2021:VKO

[Pal21] Jiban K. Pal. Visualizing the knowledge outburst in global research on COVID-19. Scientometrics, 126(5):4173–4193, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03912-3.

Parinov:2021:CCD

[Par21] Sergey Parinov. Citation contexts as a data source for evaluation of scholarly consumption. Scientometrics, 126(11):9249–9265, November 2021. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-021-04165-w.

Park:2024:DSW

[Par24] Han Woo Park. Disentangling a secret web of online scholarly interactions involving the research of top scholars in the field of communication: International Communication Association (ICA) Fellows. Scientometrics, 129(12):7565– 7583, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05181-2.

Park:2025:DES

[Par25] Dong Joon Park. Do epistemic similarity and experiential familiarity enhance the productivity of early-career interdisciplinary researchers? Scientometrics, 130(5): 2799–2828, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05301-6.

Park:2024:IFN

[PB24] Han Woo Park and George A. Barnett. ICA fellows’ networking patterns in terms of collaboration, citation, and bibliographic coupling and the relevance of co-ethnicity. Scientometrics, 129(9):5433–5452, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05112-1.

Peroni:2020:PSC

[PCG^{+}20] Silvio Peroni, Paolo Ciancarini, Aldo Gangemi, Andrea Giovanni Nuzzolese, Francesco Poggi, and Valentina Presutti. The practice of self-citations: a longitudinal study. Scientometrics, 123(1):253–282, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03397-6.

Pech:2020:API

[PD20a] Gerson Pech and Catarina Delgado. Assessing the publication impact using citation data from both Scopus and WoS databases: an approach validated in 15 research fields. Scientometrics, 125(2):909–924, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03660-w.

Pech:2020:PSB

[PD20b] Gerson Pech and Catarina Delgado. Percentile and stochastic-based approach to the comparison of the number of citations of articles indexed in different bibliographic databases. Scientometrics, 123(1):223–252, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03386-9.

Pham:2021:BRR

[PDH21] Hiep-Hung Pham, Thi-Kieu-Trang Dong, and Manh-Tung Ho. A bibliometric review of research on international student mobilities in Asia with Scopus dataset between 1984 and 2019. Scientometrics, 126(6):5201–5224, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03965-4.

Pearson:2020:RAT

[Pea20] William S. Pearson. Research article titles in written feedback on English as a second language writing. Scientometrics, 123(2):997–1019, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03388-7.

Pearson:2021:QSL

[Pea21] William S. Pearson. Quoted speech in linguistics research article titles: patterns of use and effects on citations. Scientometrics, 126(4):3421–3442, April 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03827-5;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03827-5.pdf.

Petrovich:2022:BPR

[Pet22] Eugenio Petrovich. Bibliometrics in press. Representations and uses of bibliometric indicators in the Italian daily newspapers. Scientometrics, 127(5):2195–2233, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04341-6.

Pinto:2020:ILT

[PFPCMS20] María Pinto, Rosaura Fernández-Pascual, David Caballero-Mariscal, and Dora Sales. Information literacy trends in higher education (2006–2019): visualizing the emerging field of mobile information literacy. Scientometrics, 124(2): 1479–1510, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03523-4.

Purwitasari:2020:ICD

[PFS^{+}20] Diana Purwitasari, Chastine Fatichah, Surya Sumpeno, Christian Steglich, and Mauridhi Hery Purnomo. Identifying collaboration dynamics of bipartite author-topic networks with the influences of interest changes. Scientometrics, 122(3):1407–1443, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03342-2.

Paudel:2024:CRN

[PG24] Prakash Kumar Paudel and Basant Giri. Carrying out research in Nepal: perceptions of scholars about research environment and challenges. Scientometrics, 129 (5):2499–2519, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04989-2.

Patel:2021:CGP

[PHBG21] Vanash M. Patel, Robin Haunschild, Lutz Bornmann, and George Garas. A call for governments to pause Twitter censorship: using Twitter data as social-spatial sensors of COVID-19/SARS-CoV-2 research diffusion. Scientometrics, 126(4):3193–3207, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03843-5;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03843-5.pdf.

Powell:2022:KIS

[PHP22] Kerrington Powell, Alyson Haslam, and Vinay Prasad. The Kardashian Index: a study of researchers’ opinions on Twitter 2014–2021. Scientometrics, 127(4):1923–1930, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04281-1.

Pinar:2023:DRP

[Pin23] Mehmet Pinar. Do research performances of universities and disciplines in England converge or diverge? An assessment of the progress between research excellence frameworks in 2014 and 2021. Scientometrics, 128(10):5731– 5766, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04802-6.

Petrie:2024:NTR

[PJ24] Stephen M. Petrie and T’Mir D. Julius. A novel text representation which enables image classifiers to also simultaneously classify text, applied to name disambiguation. Scientometrics, 129(2):719–743, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04712-7.

Piniewski:2024:EPP

[PJKK24] Mikołaj Piniewski, Ivan Jarić, Demetris Koutsoyiannis, and Zbigniew W. Kundzewicz. Emerging plagiarism in peer-review evaluation reports: a tip of the iceberg? Scientometrics, 129(4):2489–2498, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04960-1.

Potter:2024:TCC

[PKA24] Ross W. K. Potter, Milovan Kovac, and Jonathan Adams. Tracking changes in CNCI: the complementarity of standard, collaboration and fractional CNCI in understanding and evaluating research performance. Scientometrics, 129(10):6183–6196, October 2024. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-024-05028-w.

Pourhatami:2021:MIS

[PKCOG21] Aliakbar Pourhatami, Mohammad Kaviyani-Charati, and Carlos Olmeda-Gómez. Mapping the intellectual structure of the coronavirus field (2000–2020): a co-word analysis. Scientometrics, 126(8):6625–6657, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04038-2.

Pallari:2022:CCR

[PL22a] Elena Pallari and Grant Lewison. Cardiovascular and cancer risk factors analysis for 2001–2020 from the global research output and European newspapers. Scientometrics, 127(9):5159–5174, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04465-9.

Pilkina:2022:GDR

[PL22b] Marina Pilkina and Andrey Lovakov. Gender disparities in Russian academia: a bibliometric analysis. Scientometrics, 127(6):3577–3591, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04383-w.

Poitras:2023:RMU

[PL23] Constance Poitras and Vincent Larivière. Research mobility to the United States: a bibliometric analysis. Scientometrics, 128(4):2601–2614, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04657-x.

Phoa:2020:TSD

[PLH20] Frederick Kin Hing Phoa, Hsin-Yi Lai, and Keisuke Honda. A two-step deep learning approach to data classification and modeling and a demonstration on subject type relationship analysis in the Web of Science. Scientometrics, 125(2):851–863, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03599-y.

Ploszaj:2025:ATR

[Plo25a] Adam Ploszaj. Air travel and research collaboration: a quasi-experimental insight. Scientometrics, 130(4):2167– 2183, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05291-5.

Ploszaj:2025:ILD

[Plo25b] Adam Ploszaj. Individual-level determinants of international academic mobility: insights from a survey of Polish scholars. Scientometrics, 130(4):2273–2290, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05281-7.

Pornprasit:2022:ECR

[PLT22] Chanathip Pornprasit, Xin Liu, and Suppawong Tuarob. Enhancing citation recommendation using citation network embedding. Scientometrics, 127(1):233–264, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04196-3.

Peng:2023:RGM

[PLW23] Cheng Peng, Zhepeng (Lionel) Li, and Chaojiang Wu. Researcher geographic mobility and publication productivity: an investigation into individual and institutional characteristics and the roles of academicians. Scientometrics, 128(1):379–406, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04546-9.

Pei:2024:CEI

[PLYZ24] Ruimin Pei, Langqiu Li, Yiying Yang, and Quan Zhou. Co-evolution of international scientific mobility and international collaboration: a Scopus-based analysis. Scientometrics, 129(7):4353–4378, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05081-5.

Prathap:2020:LEC

[PM20] Gangan Prathap and Somenath Mukherjee. Letter to the Editor: Comments on the paper of Batagelj — On fractional approach to analysis of linked networks. Scientometrics, 124(3):2717–2722, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03541-2. See [Bat20].

Palos:2023:LSC

[PMB^{+}23] Andrijana Perković Paloš, Antonija Mijatović, Ivan Buljan, Daniel Garcia-Costa, Elena Álvarez-García, Francisco Grimaldo, and Ana Marušić. Linguistic and semantic characteristics of articles and peer review reports in Social Sciences and Medical and Health Sciences: analysis of articles published in Open Research Central. Scientometrics, 128(8):4707–4729, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04771-w.

Pooja:2021:ESA

[PMC21] KM. Pooja, Samrat Mondal, and Joydeep Chandra. Exploiting similarities across multiple dimensions for author name disambiguation. Scientometrics, 126(9):7525– 7560, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04101-y.

Puertas:2023:FRM

[PMGÁC23] Rosa Puertas, Luisa Marti, and Jose M. García-Álvarez-Coque. Are female researchers more efficient? An analysis of gender in a Spanish technological university. Scientometrics, 128(12):6611–6632, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04844-w.

Poirrier:2021:RI

[PMHC21] Maurice Poirrier, Sebastián Moreno, and Gonzalo Huerta-Cánepa. Robust h-index. Scientometrics, 126(3):1969– 1981, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-020-03857-z.

Pohl:2020:CCR

[Poh20] Hans Pohl. Collaboration with countries with rapidly growing research: supporting proactive development of international research collaboration. Scientometrics, 122(1): 287–307, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03287-6.

Pohl:2021:IIA

[Poh21] Hans Pohl. Internationalisation, innovation, and academiccorporate co-publications. Scientometrics, 126(2):1329– 1358, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03799-6.

Pohl:2024:UCB

[Poh24] Hans Pohl. Using citation-based indicators to compare bilateral research collaborations. Scientometrics, 129(8): 4751–4770, August 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05087-z.

Polonioli:2020:SBS

[Pol20] Andrea Polonioli. In search of better science: on the epistemic costs of systematic reviews and the need for a pluralistic stance to literature search. Scientometrics, 122(2): 1267–1274, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03333-3.

Polonioli:2021:ESR

[Pol21] Andrea Polonioli. The ethics of scientific recommender systems. Scientometrics, 126(2):1841–1848, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03766-1.

Porter:2025:UOA

[Por25] Stephen R. Porter. Understanding ORCID adoption among academic researchers. Scientometrics, 130(5):2749– 2781, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05300-7.

Pigola:2022:SUA

[PR22] Angélica Pigola and Priscila Rezende Da Costa. In search of understanding about knowledge and learning on innovation performance. Scientometrics, 127(7):3995–4022, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04417-3.

Prathap:2023:MRR

[PR23] Gangan Prathap and Ronald Rousseau. The modified repeat rate described within a thermodynamic framework. Scientometrics, 128(5):3185–3195, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04698-2.

Prathap:2020:LEH

[Pra20a] Gangan Prathap. Letter to the editor: is the hand the cutting edge of the mind? Lessons from publications and patent families data. Scientometrics, 123(1): 559–561, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03395-8.

Prathap:2020:LEJ

[Pra20b] Gangan Prathap. Letter to the editor: Journal indicators from a dimensionality perspective. Scientometrics, 122(2): 1259–1265, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03320-8.

Prathap:2020:LES

[Pra20c] Gangan Prathap. Letter to the editor: science indicators in development time. Scientometrics, 123(1):557–558, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03394-9.

Praus:2020:HAS

[Pra20d] Petr Praus. HCR for assessment of scientific journals in chemistry. Scientometrics, 122(2):1237–1242, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03322-6.

Prathap:2021:LEA

[Pra21a] Gangan Prathap. Letter to the editor: Additive rules for h-index for the part-set method. Scientometrics, 126(6): 5369–5371, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03950-x.

Prathap:2021:LED

[Pra21b] Gangan Prathap. Letter to the editor: Dimensionless citation indicators for fractional counting. Scientometrics, 126(10):8765–8769, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04113-8.

Prathap:2021:LEH

[Pra21c] Gangan Prathap. Letter to the editor: Is the h-index a mock compromise between the p-index and the zindex? Scientometrics, 126(5):4537–4539, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03928-9.

Prathap:2022:LEC

[Pra22] Gangan Prathap. Letter to the editor: comments on the paper of Gagolewski et al.: Ockham’s index of citation impact. Scientometrics, 127(10):6051–6054, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04496-2. See [GZSC22b].

Prathap:2023:LEC

[Pra23a] Gangan Prathap. Letter to the Editor: Comments on the paper of Safón and Docampo?: what are you reading? From core journals to trendy journals in the Library and Information Science (LIS) field. Scientometrics, 128 (7):4137–4142, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04742-1.

Prathap:2023:LEM

[Pra23b] Gangan Prathap. Letter to the editor: Measure measure on the wall who is the fairest of them all? Scientometrics, 128(1):871–872, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04580-7.

Praus:2023:ERB

[Pra23c] Petr Praus. Empirical relationship between the number of review and research articles. Scientometrics, 128(4): 2201–2209, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04654-0.

Praus:2024:AJR

[Pra24] Petr Praus. Analysis of journal rankings confirms that more cited articles contain more references. Scientometrics, 129(11):7153–7160, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05169-y.

Prathap:2025:GBR

[Pra25a] Gangan Prathap. GDP baseline ranking of scientific performance. Scientometrics, 130(2):1237–1251, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05250-0.

Praus:2025:NTS

[Pra25b] Petr Praus. A note on the topic of single-author articles in science. Scientometrics, 130(5):3053–3069, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05315-0.

Pedraja-Rejas:2024:SPL

[PRGTOP^{+}24] Liliana Pedraja-Rejas, Miguel-A. Garrido-Tamayo, Ignacio Ortega-Piwonka, Emilio Rodríguez-Ponce, and David Laroze. Scientific production in Latin American physics: a bibliometric analysis. Scientometrics, 129(7):4189–4230, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05035-x.

Perianes-Rodriguez:2021:EPP

[PROG21] Antonio Perianes-Rodríguez and Carlos Olmeda-Gómez. Effect of policies promoting open access in the scientific ecosystem: case study of ERC grantee publication practice. Scientometrics, 126(8):6825–6836, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03966-3.

Perianes-Rodriguez:2024:PFA

[PROGDG24] Antonio Perianes-Rodríguez, Carlos Olmeda-Gómez, Natalia R. Delbianco, and Maria Cláudia Cabrini Grácio. Public funding accountability: a linked open data-based methodology for analysing the scientific productivity and influence of funded projects. Scientometrics, 129(10): 5841–5868, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04975-8.

Pelacho:2021:AEC

[PRS^{+}21] M. Pelacho, G. Ruiz, F. Sanz, A. Tarancón, and J. Clemente-Gallardo. Analysis of the evolution and collaboration networks of citizen science scientific publications. Scientometrics, 126(1):225–257, January 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03724-x;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03724-x.pdf.

Paswan:2020:GRP

[PS20] Jyoti Paswan and Vivek Kumar Singh. Gender and research publishing analyzed through the lenses of discipline, institution types, impact and international collaboration: a case study from India. Scientometrics, 123 (1):497–515, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03398-5.

Polemis:2022:WSD

[PS22] Michael L. Polemis and Thanasis Stengos. What shapes the delay in the Nobel Prize discoveries? A research note. Scientometrics, 127(2):803–811, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04241-1.

Potter:2022:CSC

[PSA22] Ross W. K. Potter, Martin Szomszor, and Jonathan Adams. Comparing standard, collaboration and fractional CNCI at the institutional level: Consequences for performance evaluation. Scientometrics, 127(12):7435–7448, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04303-y.

Price:2022:NDS

[PSH22] Robyn Price, Mark Skopec, and Matthew Harris. A novel data solution to inform curriculum decolonisation: the case of the Imperial College London Masters of Public Health. Scientometrics, 127(2):1021–1037, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04231-3.

Paswan:2022:DUI

[PSKS22] Jyoti Paswan, Vivek Kumar Singh, Mousumi Karmakar, and Prashasti Singh. Does university-industry-government collaboration in research gets higher citation and altmetric impact? A case study from India. Scientometrics, 127(11):6063–6082, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04508-1.

Pourghaz:2023:AIR

[PSS23] Arman Pourghaz, Ehsan Bahrami Samani, and Babak Shokri. Analysis of the impact of research output on economic growth with using a multivariate random effects model. Scientometrics, 128(4):2259–2282, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04638-0.

Piro:2024:RSV

[PSW24] Fredrik Niclas Piro, Marco Seeber, and Lili Wang. Regional and sectoral variations in the ability to attract funding from the European Union’s Seventh Framework Program and Horizon 2020. Scientometrics, 129(3):1493–1521, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04942-3.

Pinto:2020:IRO

[PT20] Tânia Pinto and Aurora A. C. Teixeira. The impact of research output on economic growth by fields of science: a dynamic panel data analysis, 1980–2016. Scientometrics, 123(2):945–978, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03419-3.

Pinto:2024:ROE

[PT24] Tânia Pinto and Aurora A. C. Teixeira. Research output and economic growth in technological laggard contexts: a longitudinal analysis (1980–2019) by type of research. Scientometrics, 129(3):1197–1230, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04893-1.

Pupovac:2021:FPI

[Pup21] Vanja Pupovac. The frequency of plagiarism identified by text-matching software in scientific articles: a systematic review and meta-analysis. Scientometrics, 126(11):8981– 9003, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04140-5.

Purnell:2021:CPP

[Pur21] Philip J. Purnell. Conference proceedings publications in bibliographic databases: a case study of countries in Southeast Asia. Scientometrics, 126(1):355–387, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03773-2;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03773-2.pdf.

Purnell:2024:GRG

[Pur24] Philip J. Purnell. Geodiversity of research: geographical topic focus, author location, and collaboration. A case study of SDG 2: zero hunger. Scientometrics, 129(5):

2701–2727, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04994-5.

Purnell:2025:TRH

[Pur25] Philip James Purnell. Transdisciplinary research: how much is academia heeding the call to work more closely with societal stakeholders such as industry, government, and nonprofits? Scientometrics, 130(7):3537–3569, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05367-2.

Prakash:2024:GGS

[PVA24] Arjun Prakash, Jeevan John Varghese, and Shruti Aggarwal. Gender of gender studies: examining regional and gender-based disparities in scholarly publications. Scientometrics, 129(7):4471–4493, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05084-2.

Petrovich:2024:BBC

[PVB^{+}24] Eugenio Petrovich, Sander Verhaegh, Gregor Bös, Claudia Cristalli, Fons Dewulf, Ties van Gemert, and Nina IJdens. Bibliometrics beyond citations: introducing mention extraction and analysis. Scientometrics, 129(9):5731– 5768, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05116-x.

Pham:2023:IIR

[PVP^{+}23] Hoang-Son Pham, Bram Vancraeynest, Hanne Poelmans, Sadia Vancauwenbergh, and Amr Ali-Eldin. Identifying interdisciplinary research in research projects. Scientometrics, 128(10):5521–5544, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04810-6.

Portenoy:2020:CEA

[PW20] Jason Portenoy and Jevin D. West. Constructing and evaluating automated literature review systems. Scientometrics, 125(3):3233–3251, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03490-w.

Pan:2025:IIA

[PWZ25] Xuelian Pan, Rui Wu, and Yuanyuan Zhai. The impact of international academic mobility on doctoral students’ research collaboration and publication output. Scientometrics, 130(7):3773–3811, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05358-3.

Peng:2024:CDC

[PXY24] VXia Peng, Zequan Xiong, and Li Yang. Can document characteristics affect motivations for literature usage? Scientometrics, 129(6):3543–3563, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05044-w.

Pan:2024:DCP

[PY24] Fan Pan and Yiying Yang. Diachronic changes in the phrasal complexity of research articles (1970–2020): a cross-disciplinary investigation. Scientometrics, 129(7): 4395–4421, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05060-w.

Pan:2025:DCL

[PY25] Fan Pan and Yiying Yang. Diachronic change in lexical complexity of research articles (1970–2020): economics vs. medicine. Scientometrics, 130(3):1789–1812, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05258-6.

Pessin:2022:SBI

[PYS22] Vilker Zucolotto Pessin, Luciana Harue Yamane, and Renato Ribeiro Siman. Smart bibliometrics: an integrated method of science mapping and bibliometric analysis. Scientometrics, 127(6):3695–3718, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04406-6.

Qureshi:2021:FGA

[QD21] Muhammad Sajid Qureshi and Ali Daud. Fine-grained academic rankings: mapping affiliation of the influential researchers with the top ranked HEIs. Scientometrics, 126(10):8331–8361, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04138-z.

Qayyum:2022:TPH

[QJI^{+}22] Faiza Qayyum, Harun Jamil, Naeem Iqbal, DoHyeun Kim, and Muhammad Tanvir Afzal. Toward potential hybrid features evaluation using MLP–ANN binary classification model to tackle meaningful citations. Scientometrics, 127(11):6471–6499, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04530-3.

Qin:2025:GEA

[QLW25] Yinghong Qin, Fanghua Li, and Tianyu Wang. Geothermal energy application papers with location titles attract fewer citations. Scientometrics, 130(8):4545–4572, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05381-4.

Qiao:2025:CEC

[QPH^{+}25] Yali Qiao, Alan L. Porter, Ying Huang, Haiyun Xu, and Xuefeng Wang. Comparing examiner citations and applicant citations: insights into technology evolution. Scientometrics, 130(2):537–563, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05245-x.

Qin:2021:MCP

[QQL21] Yawen Qin, Xiaozhen Qin, and Wei Lang. Measuring cognitive proximity using semantic analysis: a case study of China’s ICT industry. Scientometrics, 126(7):6059–6084, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04021-x.

Quan:2023:PFI

[QSYL23] Wei Quan, Fei Shu, Meijia Yang, and Vincent Larivière. Publish and flourish: investigating publication requirements for PhD students in China. Scientometrics, 128(12):6675–6693, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04854-8.

Qi:2023:MTL

[QWS^{+}23] Ruihua Qi, Jia Wei, Zhen Shao, Zhengguang Li, Heng Chen, Yunhao Sun, and Shaohua Li. Multi-task learning model for citation intent classification in scientific publications. Scientometrics, 128(12):6335–6355, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04858-4.

Qiu:2021:SCR

[QYL21] Tianshuang Qiu, Chuanming Yu, and Gang Li. A scientific citation recommendation model integrating network and text representations. Scientometrics, 126(11):9199– 9221, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04161-0.

Qu:2020:RLB

[QZ20] Zhao Qu and Shanshan Zhang. References to literature from the business sector in patent documents: a case study of charging technologies for electric vehicles. Scientometrics, 124(2):867–886, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03518-1.

Qi:2022:CRC

[QZF22] Yan Qi, Xin Zhang, and Shu Fang. Choosing the right collaboration partner for innovation: a framework based on topic analysis and link prediction. Scientometrics, 127(9):5519–5550, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04306-9.

Qi:2024:FIC

[QZS^{+}24] Fan Qi, Hongyu Zhou, Beibei Sun, Ying Huang, and Lin Zhang. Facilitating interdisciplinarity: the contributions of boundary-crossing activities among disciplines. Scientometrics, 129(10):6435–6453, October 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04924-x.

Rhaiem:2020:DRE

[RA20] Mehdi Rhaiem and Nabil Amara. Determinants of research efficiency in Canadian business schools: evidence from scholar-level data. Scientometrics, 125(1):53–99, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03633-z.

Rafols:2025:TMO

[Raf25a] Ismael Rafols. Towards multiple ontologies in science mapping. A tribute to Loet Leydesdorff. Scientometrics, 130 (6):3213–3227, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05323-0.

Rosengaard:2025:CRR

[RAF25b] Louise Olsbro Rosengaard, Mikkel Zola Andersen, and Siv Fonnes. Cochrane reviews received more online attention than other systematic reviews — except when published in leading medical journals. Scientometrics, 130(5): 2829–2859, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05302-5.

Ruland:2025:SDG

[RAG25] Anna-Lena Rüland, Lise H. Andersen, and Bruno Iochins Grisci. Science diplomacy: A global research field? Findings from a bibliometric analysis of the science diplomacy scholarship of the past twenty years. Scientometrics, 130(8):4667–4696, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05396-x.

Roudsari:2022:PML

[RAL22] Arousha Haghighian Roudsari, Jafar Afshar, and Suan Lee. PatentNet: multi-label classification of patent documents using deep learning based language understanding. Scientometrics, 127(1):207–231, January 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04179-4.

Rudiger:2021:EPC

[RAS21] Matthias Sebastian Rüdiger, David Antons, and Torsten-Oliver Salge. The explanatory power of citations: a new approach to unpacking impact in science. Scientometrics, 126(12):9779–9809, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04103-w.

Rose:2025:FDS

[RB25] Michael E. Rose and Stefano H. Baruffaldi. Finding Doppelgängers in Scopus: how to build scientists control groups using sosia. Scientometrics, 130(5):2987–3011, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05298-y.

Roychowdhury:2022:MRL

[RBB22] Koel Roychowdhury, Radhika Bhanja, and Sushmita Biswas. Mapping the research landscape of Covid-19 from social sciences perspective: a bibliometric analysis. Scientometrics, 127(8):4547–4568, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04447-x.

Rennings:2024:AES

[RBB^{+}24] Michael Rennings, Philipp Baaden, Carolin Block, Marcus John, and Stefanie Bröring. Assessing emerging sustainability-oriented technologies: the case of precision agriculture. Scientometrics, 129(6):2969–2998, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05022-2.

Ramdani:2024:NGS

[RBBP24] Anas Ramdani, Catherine Beaudry, Mario Bourgault, and Davide Pulizzotto. Navigating geopolitical storms: assessing the robustness of Canada’s 5G research network in the wake of the Huawei conflict. Scientometrics, 129(10): 6221–6259, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05078-0.

Roshani:2021:WRB

[RBC21] Saeed Roshani, Mohammad-Reza Bagherylooieh, and Mario Coccia. What is the relationship between research funding and citation-based performance? A comparative analysis between critical disciplines. Scientometrics, 126(9):7859–7874, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04077-9.

Rojko:2020:BRI

[RBL20] Katarina Rojko, Brankica Bratić, and Borut Luzar. The Bologna reform’s impacts on the scientific publication performance of Ph.D. graduates — the case of Slovenia. Scientometrics, 124(1):329–356, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03482-w.

Ravikumar:2022:CSB

[RBR22] S. Ravikumar, Bidyut Bikash Boruah, and M. N. Ravikumar. Correlation study between citation count and Mendeley readership of the articles of Sri Lankan authors. Scientometrics, 127(8):4873–4885, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04470-y.

Rasuli:2025:LBD

[RBV25] Behrooz Rasuli, Michael Boock, and Brenda Van Wyk. The link between dissertation metadata completeness and user engagement in an institutional repository. Scientometrics, 130(5):2861–2874, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05331-0.

Rodrigues:2021:MBS

[RdNB21] Bethânia Ávila Rodrigues, Mariana Machado Fidelis do Nascimento, and Juliana Vitória Messias Bittencourt. Mapping of the behavior of scientific publications since the decade of 1990 until the present day in the field of food and nutrition security. Scientometrics, 126(3):2459–2483, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03679-z.

Rehs:2020:STM

[Reh20] Andreas Rehs. A structural topic model approach to scientific reorientation of economics and chemistry after German reunification. Scientometrics, 125(2):1229– 1251, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03640-0;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03640-0.pdf.

Rivera:2024:USM

[RFG^{+}24] Reynaldo Gustavo Rivera, Carlos Orellana Fantoni, Eunice Gálvez, Priscilla Jimenez-Pazmino, Carmen Karina Vaca Ruiz, and Arturo Fitz Herbert. Using scientometrics to mapping Latin American research networks in emerging fields: the field networking index. Scientometrics, 129(4): 2309–2335, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04970-z.

Rodriguez-Faneca:2022:PWE

[RFMMPJ22] Cristina Rodríguez-Faneca, Alexander Maz-Machado, and Cristina Pedrosa-Jesús. Presence of women on the editorial boards of the language and linguistics journals in Spain. Scientometrics, 127(7):4237–4249, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04412-8.

Raban:2020:EDS

[RG20] Daphne R. Raban and Avishag Gordon. The evolution of data science and big data research: a bibliometric analysis. Scientometrics, 122(3):1563–1581, March 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03371-2;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03371-2.pdf.

Robinson-Garcia:2024:EMS

[RGVQTS^{+}24] Nicolas Robinson-Garcia, Benjamín Vargas-Quesada, Daniel Torres-Salinas, Zaida Chinchilla-Rodríguez, and Juan Gorraiz. Errors of measurement in scientometrics: classification schemes and document types in citation and publication rankings. Scientometrics, 129(10):6455–6475, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05159-0.

Rigby:2020:BDT

[RJ20a] John Rigby and Barbara Jones. Bringing the doctoral thesis by published papers to the social sciences and the humanities: A quantitative easing? A small study of doctoral thesis submission rules and practice in two disciplines in the UK. Scientometrics, 124(2): 1387–1409, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03483-9;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03483-9.pdf.

Rigby:2020:RDB

[RJ20b] John Rigby and Barbara Jones. Response to Dr. Breimer’s and Dr. Mikhailidis’ letter. Scientometrics, 125(1):817– 818, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03623-1.

Rusek:2023:BIS

[RKCA23] Krzysztof Rusek, Agnieszka Kleszcz, and Albert Cabellos-Aparicio. Bayesian inference of spatial and temporal relations in AI patents for EU countries. Scientometrics, 128 (6):3313–3335, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04699-1.

Reale:2020:MDR

[RKMG20] Denis Réale, Mahdi Khelfaoui, Pierre-Olivier Montiglio, and Yves Gingras. Mapping the dynamics of research networks in ecology and evolution using co-citation analysis (1975–2014). Scientometrics, 122(3):1361–1385, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03340-4.

Rousi:2020:JRD

[RL20] Antti M. Rousi and Mikael Laakso. Journal research data sharing policies: a study of highly-cited journals in neuroscience, physics, and operations research. Scientometrics, 124(1):131–152, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03467-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03467-9.pdf.

Rong:2024:IAW

[RLZ^{+}24] Guoyang Rong, Changling Li, Zhijian Zhang, Shuaipu Chen, and Yuxing Qian. Investigating the application of work-energy metaphor in interdisciplinary citation analysis. Scientometrics, 129(6):3573–3591, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05019-x.

Rossello:2024:BAV

[RM24] Giulia Rossello and Arianna Martinelli. Breach of academic values and misconduct: the case of Sci-Hub. Scientometrics, 129(9):5227–5263, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05046-8.

Rachid:2021:GDR

[RNAH21] Elza Rachid, Tania Noureddine, and Christiane Al-Haddad. Gender disparity in research productivity across departments in the faculty of medicine: a bibliometric analysis. Scientometrics, 126(6):4715–4731, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03953-8.

Rodriguez-Navarro:2022:LBC

[RNB22] Alonso Rodríguez-Navarro and Ricardo Brito. The link between countries’ economic and scientific wealth has a complex dependence on technological activity and research policy. Scientometrics, 127(5):2871–2896, May 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04313-w.

Rabitz:2021:TNA

[ROB21] Florian Rabitz, Alin Olteanu, and Agne Budzyte. A topic network analysis of the system turn in the environmental sciences. Scientometrics, 126(3):2107–2140, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03824-8.

Rodriguez:2022:MMW

[Rod22] Juan Gabriel Rodríguez. Making the most of world talent for science? The Nobel Prize and Fields Medal experience. Scientometrics, 127(2):813–847, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04236-y.

Ronda-Pupo:2020:PLA

[RP20] Guillermo Armando Ronda-Pupo. The performance of Latin American research on economics & business. Scientometrics, 122(1):573–590, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03300-y.

Ronda-Pupo:2023:MBC

[RP23] Guillermo Armando Ronda-Pupo. Mexico: a bridge in Cuba–U.S. scientific collaboration. Scientometrics, 128(4): 2301–2315, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04668-8.

Ronda-Pupo:2024:DLP

[RP24a] Guillermo Armando Ronda-Pupo. A decades-long partnership: examining the Cuba–China scientific collaboration. Scientometrics, 129(2):785–802, February 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04883-3.

Ronda-Pupo:2024:DCI

[RP24b] Guillermo Armando Ronda-Pupo. The dynamics of Cuban international scientific collaboration: a scientometric analysis over a century. Scientometrics, 129(9):5211–5226, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05137-6.

Ronda-Pupo:2025:GTS

[RP25] Guillermo Armando Ronda-Pupo. Global ties in science: a scientometric approach to international collaboration dependence. Scientometrics, 130(7):3871–3901, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05363-6.

Ronda-Pupo:2021:CAI

[RPAVFV21] Guillermo Armando Ronda-Pupo, Rodrigo Alda-Varas, and Nelson Fenández-Vergara. Cumulative advantage of the impact of the Latin American and Caribbean science system on JCR journals outside the region. Scientometrics, 126(11):9291–9304, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04168-7.

Reategui:2020:EBR

[RPCF20] Eliseo Reategui, Alause Pires, Michel Carniato, and Sergio Roberto Kieling Franco. Evaluation of Brazilian research output in education: confronting international and national contexts. Scientometrics, 125(1):427–444, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03617-z.

Ruggieri:2021:IAA

[RPL21] Roberta Ruggieri, Fabrizio Pecoraro, and Daniela Luzi. An intersectional approach to analyse gender productivity and open access: a bibliometric analysis of the Italian National Research Council. Scientometrics, 126(2): 1647–1673, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03802-0;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03802-0.pdf.

Rajput:2022:TIS

[RS22a] Abhay S. D. Rajput and Sangeeta Sharma. Top Indian scientists as public communicators: a survey of their perceptions, attitudes and communication behaviors. Scientometrics, 127(6):3167–3192, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04405-7.

Reichmann:2022:PPR

[RS22b] Gerhard Reichmann and Christian Schlögl. On the possibilities of presenting the research performance of an institute over a long period of time: the case of the Institute of Information Science at the University of Graz in Austria. Scientometrics, 127(6):3193–3223, June 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04377-8.

Rogers:2020:SSB

[RSA20] Gordon Rogers, Martin Szomszor, and Jonathan Adams. Sample size in bibliometric analysis. Scientometrics, 125 (1):777–794, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03647-7;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03647-7.pdf.

Rocha:2020:AII

[RSF20] Fabio Gomes Rocha, Rosimeri Ferraz Sabino, and Alejandro C. Frery. Analysis of the international impact of the Brazilian base ”Qualis”-Education. Scientometrics, 125(3):1949–1963, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03713-0.

Rezazadegan:2024:QPA

[RSM24] Reza Rezazadegan, Mahdi Sharifzadeh, and Christopher L. Magee. Quantifying the progress of artificial intelligence subdomains using the patent citation network. Scientometrics, 129(5):2559–2581, May 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04996-3.

Ranaei:2020:ETE

[RSPC20] Samira Ranaei, Arho Suominen, Alan Porter, and Stephen Carley. Evaluating technological emergence using text analytics: two case technologies and three approaches. Scientometrics, 122(1):215–247, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03275-w;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03275-w.pdf.

Rahmani:2020:USC

[RSV^{+}20] Negin Rahmani, Alireza Salehi, Hossein Molavi Vardanjani, Maryam Marzban, and Arezoo Behbood. Using STROBE checklist to assess the reporting quality of observational studies affiliated with Shiraz University of Medical Sciences, and its correlates: a scientometric study from Iran. Scientometrics, 122(2):989–1001, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03317-3.

Rizzoli:2023:PLC

[RTT23] Valentina Rizzoli, Matilde Trevisani, and Arjuna Tuzzi. Portraying the life cycle of ideas in social psychology through functional (textual) data analysis: a toolkit for digital history. Scientometrics, 128(9):5197–5226, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04722-5.

Rushforth:2023:LRT

[Rus23] Alex Rushforth. Letter: Response to Torres-Salinas et al. on “bibliometric denialism”. Scientometrics, 128(12): 6781–6784, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-

04842-y. See [TSAMRG23].

Romasanta:2020:IPR

[RvdSvMK20] Angelo Kenneth S. Romasanta, Peter van der Sijde, and Jacqueline van Muijlwijk-Koezen. Innovation in pharmaceutical R&D: mapping the research landscape. Scientometrics, 125(3):1801–1832, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03707-y;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03707-y.pdf.

Rozing:2020:FBR

[RvLR^{+}20] M. P. Rozing, T. N. van Leeuwen, P. H. Reitsma, F. R. Rosendaal, and N. A. Aziz. Freeloading in biomedical research. Scientometrics, 122(1):47–55, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2984-3.

Ren:2023:DCI

[RWL23] Jingjing Ren, Fang Wang, and Minglu Li. Dynamics and characteristics of interdisciplinary research in scientific breakthroughs: case studies of Nobel-winning research in the past 120 years. Scientometrics, 128(8):4383–4419, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04762-x.

Rong:2024:IBD

[RYS24] Xueyun Rong, Zhongkai Yang, and Yutao Sun. Inventors’ brokerages dynamic and exploratory innovation: the moderating role of knowledge diversity. Scientometrics, 129 (5):2533–2558, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04993-6.

Rousseau:2023:CUW

[RZS23a] Ronald Rousseau, Lin Zhang, and Gunnar Sivertsen. Correction: Using the weighted Lorenz curve to represent balance in collaborations: the BIC indicator. Scientometrics, 128(7):4163–4164, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04708-3. See [RZS23b].

Rousseau:2023:UWL

[RZS23b] Ronald Rousseau, Lin Zhang, and Gunnar Sivertsen. Using the weighted Lorenz curve to represent balance in collaborations: the BIC indicator. Scientometrics, 128(1): 609–622, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04533-0.

See correction [RZS23a].

Subbotin:2021:BDB

[SA21] Alexander Subbotin and Samin Aref. Brain drain and brain gain in Russia: Analyzing international migration of researchers by discipline using Scopus bibliometric data 1996–2020. Scientometrics, 126(9):7875–7900, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04091-x.

Sjogaarde:2024:NDC

[SA24] Peter Sjögårde and Per Ahlgren. Normalization of direct citations for clustering in publication-level networks: evaluation of six approaches. Scientometrics, 129(3):1949– 1968, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04932-5.

Salman:2021:AAR

[SAA21] Muhammad Salman, Mohammad Masroor Ahmed, and Muhammad Tanvir Afzal. Assessment of author ranking indices based on multi-authorship. Scientometrics, 126(5): 4153–4172, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03906-1.

Someda:2022:ASE

[SAK22] Hiroshi Someda, Takanori Akagi, and Yuya Kajikawa. An analysis of the spillover effects based on patents and interindustrial transactions for an emerging blockchain technology. Scientometrics, 127(8):4299–4314, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04457-9.

Sampagnaro:2023:KOJ

[Sam23] Gabriele Sampagnaro. Keyword occurrences and journal specialization. Scientometrics, 128(10):5629–5645, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04815-1.

Sandnes:2020:SBE

[San20] Frode Eika Sandnes. A simple back-of-the-envelope test for self-citations using Google Scholar author profiles. Scientometrics, 124(2):1685–1689, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03521-6.

Sandnes:2021:BSH

[San21a] Frode Eika Sandnes. A bibliometric study of human– computer interaction research activity in the Nordic–Baltic Eight countries. Scientometrics, 126(6):4733–4767, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03940-z.

Sandnes:2021:EOP

[San21b] Frode Eika Sandnes. Everyone onboard? Participation ratios as a metric for research activity assessments within young universities. Scientometrics, 126(7):6105–6113, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04004-y.

Sandnes:2024:CWI

[San24] Frode Eika Sandnes. Can we identify prominent scholars using ChatGPT? Scientometrics, 129(1):713–718, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04882-4.

Sandnes:2025:TTM

[San25] Frode Eika Sandnes. Are there too many papers by the same authors within the same conference proceedings? Norms and extremities within the field of human– computer interaction. Scientometrics, 130(3):1659–1699, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05271-9.

Shehatta:2019:ICS

[SAR19] Ibrahim Shehatta and Abdullah M. Al-Rubaish. Impact of country self-citations on bibliometric indicators and ranking of most productive countries. Scientometrics, 120(2): 775–791, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-019-03139-3.

Shaikh:2023:YSM

[SAS23] Abdul Rahman Shaikh, Hamed Alhoori, and Maoyuan Sun. YouTube and science: models for research impact. Scientometrics, 128(2):933–955, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04574-5.

Scheffler:2020:DLO

[SB20] Maximilian Scheffler and Johannes Brunzel. Destructive leadership in organizational research: a bibliometric approach. Scientometrics, 125(1):755–775, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03621-3;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03621-3.pdf.

SeyyedHosseini:2021:CPM

[SB21] Shohreh SeyyedHosseini and Reza BasirianJahromi. COVID-19 pandemic in the Middle East countries: coronavirus-seeking behavior versus coronavirus-related publications. Scientometrics, 126(9):7503–7523, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04066-y.

Schartinger:2025:FIP

[SB25] Doris Schartinger and Michael Barber. Firms’ intellectual property protection with national versus European design rights: a count model. Scientometrics, 130(8):??, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05351-w.

Scheidegger:2023:BCA

[SBF23] Fabian Scheidegger, Andre Briviba, and Bruno S. Frey. Behind the curtains of academic publishing: strategic responses of economists and business scholars. Scientometrics, 128(8):4765–4790, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04772-9.

Saritas:2021:BDA

[SBKK21] Ozcan Saritas, Pavel Bakhtin, Ilya Kuzminov, and Elena Khabirova. Big data augmentated business trend identification: the case of mobile commerce. Scientometrics, 126(2):1553–1579, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03807-9;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03807-9.pdf.

Sanguri:2020:SSA

[SBP20] Kamal Sanguri, Atanu Bhuyan, and Sabyasachi Patra. A semantic similarity adjusted document co-citation analysis: a case of tourism supply chain. Scientometrics, 125(1): 233–269, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03608-0.

Simoes:2020:FAM

[SC20a] Nadia Simoes and Nuno Crespo. A flexible approach for measuring author-level publishing performance. Scientometrics, 122(1):331–355, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03278-7.

Sun:2020:DSC

[SC20b] Yutao Sun and Cong Cao. The dynamics of the studies of China’s science, technology and innovation (STI): a bibliometric analysis of an emerging field. Scientometrics, 124(2):1335–1365, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03500-x.

Scafetta:2025:MSP

[Sca25] Nicola Scafetta. Measuring scholarly performance using comprehensive standardized research-teaching (RT) score. Scientometrics, 130(5):2577–2593, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05317-y.

Sanz-Casado:2021:IVN

[SCDP21] Elías Sanz-Casado, Daniela De Filippo, and Janne Pölönen. Impact and visibility of Norwegian, Finnish and Spanish journals in the fields of humanities. Scientometrics, 126(11):9031–9049, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04169-6.

Sixto-Costoya:2025:CFP

[SCFSAB^{+}25] Andrea Sixto-Costoya, Antonia Ferrer-Sapena, Rafael Aleixandre-Benavent, Fernanda Peset, Juan Carlos Valderrama-Zurián, and Luiza Petrosyan. The compliance to FAIR principles of shared data in addiction research. Scientometrics, 130(2):763–779, February 2025. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-024-05227-5.

Schmoch:2020:MVS

[Sch20a] Ulrich Schmoch. Mean values of skewed distributions in the bibliometric assessment of research units. Scientometrics, 125(2):925–935, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03476-8. See correction [Sch22].

Schubert:2020:PER

[Sch20b] András Schubert. Péter Érdi: Ranking-the unwritten rules of the social game we all play. Scientometrics, 122(3): 1813–1815, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03335-1.

Schniedermann:2021:CSR

[Sch21a] Alexander Schniedermann. A comparison of systematic reviews and guideline-based systematic reviews in medical studies. Scientometrics, 126(12):9829–9846, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04199-0.

Schubert:2021:HB

[Sch21b] András Schubert. A handful of books. Scientometrics, 126 (6):5379–5385, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03975-2.

Schmoch:2022:CMV

[Sch22] Ulrich Schmoch. Correction to: Mean values of skewed distributions in the bibliometric assessment of research units. Scientometrics, 127(3):1659, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04177-6. See [Sch20a].

Schirone:2024:FFS

[Sch24a] Marco Schirone. The formation of a field: sustainability science and its leading journals. Scientometrics, 129(1): 401–429, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04877-1.

Schmal:2024:HTT

[Sch24b] W. Benedikt Schmal. How transformative are transformative agreements? Evidence from Germany across disciplines. Scientometrics, 129(3):1863–1889, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04955-y.

Schmidt:2024:WDS

[Sch24c] Marion Schmidt. Why do some retracted articles continue to get cited? Scientometrics, 129(12):7535–7563, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05147-4.

Scharnhorst:2025:MMT

[Sch25a] Andrea Scharnhorst. Measuring a moving target: innovation studies in practice. Scientometrics, 130(6):3149–3154, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05308-z.

Schmitt:2025:DPQ

[Sch25b] Valentin J. Schmitt. Disentangling patent quality: using a large language model for a systematic literature review. Scientometrics, 130(1):267–311, January 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05206-w.

Shi:2021:WQV

[SCK21] Yanqing Shi, Si Chen, and Lele Kang. Which questions are valuable in online Q&A communities? A question’s position in a knowledge network matters. Scientometrics, 126(10):8239–8258, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04135-2.

Strcic:2022:ODD

[SCP22] Josip Strcic, Antonia Civljak, and Livia Puljak. Open data and data sharing in articles about COVID-19 published in preprint servers medRxiv and bioRxiv. Scientometrics, 127 (5):2791–2802, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04346-1.

Sud:2021:TSB

[SCU21] Abhimanyu Sud, Darren K. Cheng, and Ross Upshur. Time series-based bibliometric analysis of a systematic review of multidisciplinary care for opioid dose reduction: exploring the origins of the North American opioid crisis. Scientometrics, 126(11):8935–8955, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04154-z.

Stewart:2023:UAC

[SCW^{+}23] Craig A. Stewart, Claudia M. Costa, Julie A. Wernert, Winona Snapp-Childs, Marques Bland, Philip Blood, Terry Campbell, Peter Couvares, Jeremy Fischer, David Y. Hancock, David L. Hart, Harmony Jankowski, Richard Knepper, Donald F. McMullen, Susan Mehringer, Marlon Pierce, Gary Rogers, Robert S. Sinkovits, and John Towns. Use of accounting concepts to study research: return on investment in XSEDE, a US cyberinfrastructure service. Scientometrics, 128(6):3225–3255, June 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04539-8.

Safon:2020:AIR

[SD20] Vicente Safón and Domingo Docampo. Analyzing the impact of reputational bias on global university rankings based on objective research performance data: the case of the Shanghai Ranking (ARWU). Scientometrics, 125(3):2199–2227, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03722-z.

Sjogaarde:2022:ABT

[SD22] Peter Sjögårde and Fereshteh Didegah. The association between topic growth and citation impact of research publications. Scientometrics, 127(4):1903–1921, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04293-x.

Safon:2023:RPP

[SD23a] Vicente Safón and Domingo Docampo. Response to Professor Prathap’s comment on “What are you reading? From core journals to trendy journals in the library and information science (LIS) field”. Scientometrics, 128(7): 4143–4146, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04743-0.

Safon:2023:WYR

[SD23b] Vicente Safón and Domingo Docampo. What are you reading? From core journals to trendy journals in the library and information science (lis) field. Scientometrics, 128 (5):2777–2801, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04673-x.

Shu:2022:MDA

[SDC22] Fei Shu, Jesse David Dinneen, and Shiji Chen. Measuring the disparity among scientific disciplines using Library of Congress Subject Headings. Scientometrics, 127(6): 3613–3628, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04387-6.

Sanyal:2020:INM

[SDD20] Debarshi Kumar Sanyal, Sumana Dey, and Partha Pratim Das. g_{m}-index: a new mentorship index for researchers. Scientometrics, 123(1):71–102, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03384-x.

Sebo:2021:EPH

[SdL21] Paul Sebo and Sylvain de Lucia. Evaluation of the productivity of hospital-based researchers: comparative study between the h-index and the h(fa)-index. Scientometrics, 126(8):7087–7096, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04040-8.

Sebo:2023:AIR

[SdL23] Paul Sebo and Sylvain de Lucia. About the importance of the research question: a response to Ming Li et al.’s comments. Scientometrics, 128(3):2029–2030, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04592-3.

Sebo:2021:APB

[SdLV21a] Paul Sebo, Sylvain de Lucia, and Nathalie Vernaz. Accuracy of PubMed-based author lists of publications and use of author identifiers to address author name ambiguity: a cross-sectional study. Scientometrics, 126(5):4121–4135, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03845-3.

Sebo:2021:GGM

[SdLV21b] Paul Sebo, Sylvain de Lucia, and Nathalie Vernaz. Gender gap in medical research: a bibliometric study in Swiss university hospitals. Scientometrics, 126(1):741–755, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03741-w;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03741-w.pdf.

Stock:2023:LPL

[SDRS23] Wolfgang G. Stock, Isabelle Dorsch, Gerhard Reichmann, and Christian Schlögl. Labor productivity, labor impact, and co-authorship of research institutions: publications and citations per full-time equivalents. Scientometrics, 128(1):363–377, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04582-5.

Scelles:2025:MIP

[SdS25] Nicolas Scelles and Jaime A. Teixeira da Silva. Making the impact of publications within a field comparable by improving the field-weighted citation impact (FWCI): the case of sport management. Scientometrics, 130(3): 1571–1586, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05268-4.

Santos-dAmorim:2025:EQP

[SdSCdS25] Karen Santos-d’Amorim, Elías Sanz-Casado, and Raimundo Nonato Macedo dos Santos. Errors, questionable practices, or misconduct? A bibliometric and altmetric review covering two decades of retractions in Latin America. Scientometrics, 130(7):3651–3677, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05343-w.

Setubal:2022:MEA

[SdSFB22] Rayanne Barros Setubal, Daniel da Silva Farias, and Reinaldo Luiz Bozelli. Microwave effect: analyzing citations from classic theories and their reinventions — a case study from a classic paper in aquatic ecology — Brooks & Dodson, 1965. Scientometrics, 127(8):4751–4767, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04459-7.

Simsek:2024:DGP

[SdVvdR24] Müge Simsek, Mathijs de Vaan, and Arnout van de Rijt. Do grant proposal texts matter for funding decisions? A field experiment. Scientometrics, 129(5):2521–2532, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04968-7.

Shi:2020:CMC

[SDW^{+}20] Jiangang Shi, Kaifeng Duan, Guangdong Wu, Rui Zhang, and Xiaowei Feng. Comprehensive metrological and content analysis of the public–private partnerships (PPPs) research field: a new bibliometric journey. Scientometrics, 124(3):2145–2184, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03607-1.

Sebo:2023:APT

[Seb23a] Paul Sebo. Are acceptance and publication times longer in primary health care journals compared to internal medicine journals? A comparative study of 117 highimpact journals. Scientometrics, 128(1):873–876, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04593-2.

Sebo:2023:RPC

[Seb23b] Paul Sebo. Retractions in primary care journals (2000– 2022). Scientometrics, 128(12):6739–6760, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04850-y.

Seeber:2020:HDJ

[See20] Marco Seeber. How do journals of different rank instruct peer reviewers? Reviewer guidelines in the field of management. Scientometrics, 122(3):1387–1405, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03343-1.

Saier:2020:ULS

[SF20] Tarek Saier and Michael Färber. unarXive: a large scholarly data set with publications’ full-text, annotated intext citations, and links to metadata. Scientometrics, 125(3):3085–3108, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03382-z;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03382-z.pdf.

Sulyok:2023:DGM

[SFC^{+}23] Judit Sulyok, Beáta Fehérvölgyi, Tibor Csizmadia, Attila I. Katona, and Zsolt T. Kosztyán. Does geography matter? Implications for future tourism research in light of COVID-19. Scientometrics, 128(3):1601–1637, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04615-z.

Singh:2020:EIC

[SFJO20] Chakresh Kumar Singh, Demival Vasques Filho, Shivakumar Jolad, and Dion R. J. O’Neale. Evolution of interdependent co-authorship and citation networks. Scientometrics, 125(1):385–404, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03616-0.

Silva:2021:DMI

[SG21] Deise Deolindo Silva and Maria Cláudia Cabrini Grácio. Dispersion measures for h-index: a study of the Brazilian researchers in the field of mathematics. Scientometrics, 126(3):1983–2011, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03848-0.

Schmoch:2022:CTP

[SG22] Ulrich Schmoch and Birgit Gehrke. China’s technological performance as reflected in patents. Scientometrics, 127(1):299–317, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-
04193-6.

Santini:2022:KGE

[SGA22] Cristian Santini, Genet Asefa Gesese, and Mehwish Alam. A knowledge graph embeddings based approach for author name disambiguation using literals. Scientometrics, 127(8):4887–4912, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04426-2.

Shah:2021:CIA

[SGC21a] Tariq Ahmad Shah, Sumeer Gul, and Kanu Chakraborty. Correction to: Influence of accessibility (open and tollbased) of scholarly publications on retractions. Scientometrics, 126(6):4607, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-021-

04041-7. See [SGC21b].

Shah:2021:IAO

[SGC21b] Tariq Ahmad Shah, Sumeer Gul, and Kanu Chakraborty. Influence of accessibility (open and toll-based) of scholarly publications on retractions. Scientometrics, 126(6): 4589–4606, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03990-3.

See correction [SGC21a].

Sorz:2020:RSI

[SGG20] Johannes Sorz, Wolfgang Glänzel, and Juan Gorraiz. Research strengths identified by esteem and bibliometric indicators: a case study at the University of Vienna. Scientometrics, 125(2):1095–1116, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03672-6.

Schneijderberg:2022:SYP

[SGM22] Christian Schneijderberg, Nicolai Götze, and Lars Müller. A study of 25 years of publication outputs in the German academic profession. Scientometrics, 127(1):1–28, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04216-2.

Schubert:2022:ESF

[SGS22] András Schubert, Wolfgang Glänzel, and Gábor Schubert. Eponyms in science: famed or framed? Scientometrics, 127(3):1199–1207, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04298-6.

Sandstrom:2008:PNP

[SH08] Ulf Sandström and Martin Hällsten. Persistent nepotism in peer-review. Scientometrics, 74(2):175–189, February 2008. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-008-0211-3. See correc-

tion [SH21].

Scheidsteger:2020:TSS

[SH20] Thomas Scheidsteger and Robin Haunschild. Telling the story of solar energy meteorology into the satellite era by applying (co-citation) reference publication year spectroscopy. Scientometrics, 125(2):1159–1177, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03597-0.

Sandstrom:2021:CPN

[SH21] Ulf Sandström and Martin Hällsten. Correction to: Persistent nepotism in peer-review. Scientometrics, 126(2): 1863–1865, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03739-4;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03739-4.pdf. See [SH08].

Sharma:2021:TSR

[Sha21] Kiran Sharma. Team size and retracted citations reveal the patterns of retractions from 1981 to 2020. Scientometrics, 126(10):8363–8374, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04125-4.

Shamohammadi:2024:DSS

[Sha24a] Mehdi Shamohammadi. Determining scale and scope economies in Korean private universities: an input distance function approach. Scientometrics, 129(5):2583– 2613, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04967-8.

Sharma:2024:TDS

[Sha24b] Kiran Sharma. Over two decades of scientific misconduct in India: Retraction reasons and journal quality among inter-country and intra-country institutional collaboration. Scientometrics, 129(12):7735–7757, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05192-z.

Scheidsteger:2025:HSF

[SHB25] Thomas Scheidsteger, Robin Haunschild, and Lutz Bornmann. How similar are field-normalized citation impact scores obtained from OpenAlex and three popular commercial databases? An empirical comparison based on large German universities. Scientometrics, 130(7): 3505–3536, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05338-7.

Shelton:2020:SLC

[She20] R. D. Shelton. Scientometric laws connecting publication counts to national research funding. Scientometrics, 123 (1):181–206, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03392-x.

Santos:2024:HEC

[SHF24] João M. Santos, Hugo Horta, and Shihui Feng. Homophily and its effects on collaborations and repeated collaborations: a study across scientific fields. Scientometrics, 129(3):1801–1823, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04950-3.

Schafermeier:2023:CRT

[SHH23a] Bastian Schäfermeier, Johannes Hirth, and Tom Hanika. Correction: Research topic flows in co-authorship networks. Scientometrics, 128(9):5079–5080, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04573-6. See [SHH23b].

Schafermeier:2023:RTF

[SHH23b] Bastian Schäfermeier, Johannes Hirth, and Tom Hanika. Research topic flows in co-authorship networks. Scientometrics, 128(9):5051–5078, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04529-w. See correction [SHH23a].

Santos:2022:RBA

[SHL22a] J. M. Santos, H. Horta, and H. Luna. The relationship between academics’ strategic research agendas and their preferences for basic research, applied research, or experimental development. Scientometrics, 127(7):4191–4225, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04431-5.

Santos:2022:SRA

[SHL22b] João M. Santos, Hugo Horta, and Huan Li. Are the strategic research agendas of researchers in the social sciences determinants of research productivity? Scientometrics, 127(7):3719–3747, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04324-7. See correction [SHL22c].

Santos:2022:CSR

[SHL22c] João M. Santos, Hugo Horta, and Huan Li. Correction to: Are the strategic research agendas of researchers in the social sciences determinants of research productivity? Scientometrics, 127(7):3749, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04440-4. See [SHL22b].

Song:2022:PPD

[SHZ22] Haoyang Song, Jianhua Hou, and Yang Zhang. Patent protection: does it promote or inhibit the patented technological knowledge diffusion? Scientometrics, 127(5): 2351–2379, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04348-z.

Singh:2022:IEB

[Sin22] Prem Kumar Singh. t-index: entropy based random document and citation analysis using average h-index. Scientometrics, 127(1):637–660, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04222-4.

Sanchez-Jimenez:2023:GID

[SJBBR^{+}23] Rodrigo Sánchez-Jiménez, Iuliana Botezan, Jesús Barrasa-Rodríguez, Mari Carmen Suárez-Figueroa, and Manuel Blázquez-Ochando. Gender imbalance in doctoral education: an analysis of the Spanish university system (1977– 2021). Scientometrics, 128(4):2577–2599, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04648-y.

Song:2021:MPR

[SJW21] Xinzhi Song, Nan Jiang, and Deliang Wen. Medical professionalism research characteristics and hotspots: a 10-year bibliometric analysis of publications from 2010 to 2019. Scientometrics, 126(9):8009–8027, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03993-0. See comments [KCC22] and

response [SJW22].

Song:2022:RF

[SJW22] Xinzhi Song, Nan Jiang, and Deliang Wen. Response for “Suggestions to the article: medical professionalism research characteristics and hotspots”. Scientometrics, 127(2):1195–1197, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04223-3. See [SJW21, KCC22].

Soos:2020:ISS

[SK20] Sándor Soós and Anna Kiss. Informetrics and the study of science–society communications: a bibliometric scoping review. Scientometrics, 124(2):825–842, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03444-2;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03444-2.pdf.

Sharma:2021:GDE

[SK21] Kiran Sharma and Parul Khurana. Growth and dynamics of econophysics: a bibliometric and network analysis. Scientometrics, 126(5):4417–4436, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03884-4.

Savchenko:2022:LAA

[SK22] Igor Savchenko and Denis Kosyakov. Lost in affiliation: apatride publications in international databases. Scientometrics, 127(6):3471–3487, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04392-9.

Sachini:2020:SSS

[SKBSC20] Evi Sachini, Nikolaos Karampekios, Pierpaolo Brutti, and Konstantinos Sioumalas-Christodoulou. Should I stay or should I go? Using bibliometrics to identify the international mobility of highly educated Greek manpower. Scientometrics, 125(1):641–663, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03618-y.

Sener:2025:CDB

[ŞKKT25] Hakan Soner Şener, İdris Semih Kaya, Mücella Sena Köksal, and Zehra Taşkın. Closing the door behind: metric-based research evaluation systems and gatekeeping towards young researchers. Scientometrics, 130(4): 2291–2310, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05282-6.

Schmidt:2023:DPC

[SKS^{+}23] Marion Schmidt, Wolfgang Kircheis, Arno Simons, Martin Potthast, and Benno Stein. A diachronic perspective on citation latency in Wikipedia articles on CRISPR/Cas-9: an exploratory case study. Scientometrics, 128(6):3649–3673, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04703-8.

Shin:2022:DMI

[SL22] Jung Cheol Shin and Soo Jeung Lee. Different measures of international faculty and their impacts on global rankings. Scientometrics, 127(11):6125–6145, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04511-6.

Shu:2024:OOA

[SL24] Fei Shu and Vincent Larivière. The oligopoly of open access publishing. Scientometrics, 129(1):519–536, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04876-2.

Shi:2024:HHL

[SLH24] Kaiwen Shi, Kan Liu, and Xinyan He. Heterogeneous hypergraph learning for literature retrieval based on citation intents. Scientometrics, 129(7):4167–4188, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05066-4.

Samaniego:2023:HRP

[SLK^{+}23] Charissa Samaniego, Peggy Lindner, Maryam A. Kazmi, Bobbie A. Dirr, Dejun Tony Kong, Evonzia Jeff-Eke, and Christiane Spitzmueller. Higher research productivity = more pay? Gender pay-for-productivity inequity across disciplines. Scientometrics, 128(2):1395–1407, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04513-4.

Shen:2023:SPT

[SLL^{+}23a] Si Shen, Jiangfeng Liu, Litao Lin, Ying Huang, Lin Zhang, Chang Liu, Yutong Feng, and Dongbo Wang. SsciB-ERT: a pre-trained language model for social science texts. Scientometrics, 128(2):1241–1263, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04602-4.

Song:2023:IET

[SLL23b] Bowen Song, Chunjuan Luan, and Danni Liang. Identification of emerging technology topics (ETTs) using BERT-based model and semantic analysis: a perspective of multiple-field characteristics of patented inventions (MFCOPIs). Scientometrics, 128(11):5883–5904, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04819-x.

Sanz-Lorente:2021:DQV

[SLMCSV21] María Sanz-Lorente, Natalia Moles-Caballero, and Javier Sanz-Valero. Documentary quality versus veracity of information of the websites on syphilis and gonorrhea. Scientometrics, 126(11):8775–8788, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04123-6.

Sheng:2023:ABP

[SLR^{+}23] Libo Sheng, Dongqing Lyu, Xuanmin Ruan, Hongquan Shen, and Ying Cheng. The association between prior knowledge and the disruption of an article. Scientometrics, 128(8):4731–4751, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04751-0.

Srisawad:2025:PTF

[SLTS25] Sutthisak Srisawad, Kullacha Lertsittiphan, Sirirut Tunsirirut, and Pennapa Saenkla. Predicting the time to first citation for medical articles using survival analysis and the association between this time and citation counts. Scientometrics, 130(2):861–880, February 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05237-x.

Sun:2021:EPL

[SLX21] Kun Sun, Haitao Liu, and Wenxin Xiong. The evolutionary pattern of language in scientific writings: A case study of Philosophical Transactions of Royal Society (1665–1869). Scientometrics, 126(2):1695–1724, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03816-8;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03816-8.pdf.

Sun:2022:IAE

[SLY22] Fengjun Sun, Yingqiu Li, and Xiaolin Yao. Issues about entitymetrics and paper-entity citation network. Scientometrics, 127(4):2123–2125, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04316-7. See reply [YWB22].

Sifontes:2020:GDP

[SM20] Domingo Sifontes and Rosa Morales. Gender differences and patenting in Latin America: understanding female participation in commercial science. Scientometrics, 124(3):2009–2036, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03567-6.

Santos:2022:DPR

[SM22a] Ana Teresa Santos and Sandro Mendonça. Do papers (really) match journals’ “aims and scope”? A computational assessment of innovation studies. Scientometrics, 127(12):7449–7470, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04327-4.

Santos:2022:SWI

[SM22b] Ana Teresa Santos and Sandro Mendonça. The small world of innovation studies: an “editormetrics” perspective. Scientometrics, 127(12):7471–7486, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04279-9.

Smirnova:2023:CAA

[SM23] Nina Smirnova and Philipp Mayr. A comprehensive analysis of acknowledgement texts in Web of Science: a case study on four scientific domains. Scientometrics, 128(1): 709–734, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04554-9.

Smirnova:2024:EMS

[SM24a] Nina Smirnova and Philipp Mayr. Embedding models for supervised automatic extraction and classification of named entities in scientific acknowledgements. Scientometrics, 129(11):7261–7285, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04806-2.

Szczygielski:2024:MRS

[SM24b] Krzysztof Szczygielski and Jerzy Mycielski. The mutual reinforcement of scientific and technological knowledge-a technology-level analysis. Scientometrics, 129(11):6533– 6549, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05162-5.

Souza:2021:CSA

[SMA21] Cinthia M. Souza, Magali R. G. Meireles, and Paulo E. M. Almeida. A comparative study of abstractive and extractive summarization techniques to label subgroups on patent dataset. Scientometrics, 126(1):135–156, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03732-x.

Small:2024:LVB

[Sma24] Henry Small. Laudan versus Bayes: how often does Bayes agree with Laudan given multiple lines of evidence? Scientometrics, 129(12):7913–7922, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05197-8.

Small:2025:ISD

[Sma25] Henry Small. Information and scientific discovery: Shannon, Bayes and Leydesdorff. Scientometrics, 130(6):

3109–3126, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05325-y.

Shlezinger:2024:MFA

[SMBD^{+}24] Einav Baharav Shlezinger, Rasha Mosleh, Gil Ben-David, Eedy Mezer, and Tamara Wygnanski-Jaffe. The multifaceted factors affecting the publication times of pediatric ophthalmology and strabismus articles: what has changed in two decades. Scientometrics, 129(9):5047– 5073, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05122-z.

Serenko:2022:SPR

[SMD22] Alexander Serenko, Mauricio Marrone, and John Dumay. Scientometric portraits of recognized scientists: a structured literature review. Scientometrics, 127(8):4827–4846, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04466-8.

Seidlmayer:2024:RTD

[SMG^{+}24] Eva Seidlmayer, Tetyana Melnychuk, Lukas Galke, Lisa Kühnel, Klaus Tochtermann, Carsten Schultz, and Konrad U. Förstner. Research topic displacement and the lack of interdisciplinarity: lessons from the scientific response to COVID-19. Scientometrics, 129(9):5141–5179, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05132-x.

Sersic:2021:CCA

[SMJ21] Darja Maslić Sersić, Marina Martincević, and Maja Jokić. The contribution of CEE authors to psychological science: a comparative analysis of papers published in CEE and non-CEE journals indexed by Scopus in the period 1996– 2013. Scientometrics, 126(2):1453–1469, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03784-z.

Sargsyan:2020:EED

[SMM^{+}20] Sh. A. Sargsyan, D. A. Maisano, A. R. Mirzoyan, A. A. Manukyan, and E. G. Gzoyan. EU-EAEU dilemma of Armenia: Does science support politics? Scientometrics, 122(3):1491–1507, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03337-z.

Shu:2020:CST

[SMQL20] Fei Shu, Yue Ma, Junping Qiu, and Vincent Larivière. Classifications of science and their effects on bibliometric evaluations. Scientometrics, 125(3):2727–2744, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03701-4.

Song:2023:ITL

[SMS23] Linpei Song, Zhuang Ma, and Junyi Sun. The influence of technostress, learning goal orientation, and perceived team learning climate on intra-team knowledge sharing and innovative practices among ICT-enabled team members. Scientometrics, 128(1):115–136, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04568-3.

Scanff:2023:FLI

[SMT^{+}23] Alexandre Scanff, Nicolas Mauhe, Marion Taburet, Pierre-Etienne Savourat, Thomas Clément, Benjamin Bastian, Ioana Cristea, Alain Braillon, Nicolas Carayol, and Florian Naudet. The “Free lunches” index for assessing academics: a not entirely serious proposal. Scientometrics, 128(12):6761–6772, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04862-8.

Sun:2023:ARA

[SMZY23] Mingyue Sun, Tingcan Ma, Lewei Zhou, and Mingliang Yue. Analysis of the relationships among paper citation and its influencing factors: a Bayesian network-based approach. Scientometrics, 128(5):3017–3033, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04697-3.

Singh:2022:ISR

[SNK22] Vivek Kumar Singh, Abhirup Nandy, and Anurag Kanaujia. Indian Science Reports: a web-based scientometric portal for mapping Indian research competencies at overall and institutional levels. Scientometrics, 127(7):4227–4236, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04395-6.

Savage:2021:DSF

[SO21] William E. Savage and Anthony J. Olejniczak. Do senior faculty members produce fewer research publications than their younger colleagues? Evidence from Ph.D. granting institutions in the United States. Scientometrics, 126(6): 4659–4686, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03957-4.

Soderstrom:2023:SDI

[Söd23] Kristofer Rolf Söderström. The structure and dynamics of instrument collaboration networks. Scientometrics, 128 (6):3581–3600, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04658-w.

Salgado-Orellana:2021:SSD

[SOdLGB21] Norma Salgado-Orellana, Emilio Berrocal de Luna, and Calixto Gutiérrez-Braojos. A scientometric study of doctoral theses on the Roma in the Iberian Peninsula during the 1977–2018 period. Scientometrics, 126(1):437– 458, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03723-y.

Sanchez:2024:PFF

[SOSC24] Flor Sánchez, Ricardo Olmos, Leyla Angélica Sandoval, and Fernando Casani. Psychosocial factors that favor citizen participation in the generation of scientific knowledge. Scientometrics, 129(8):5021–5036, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05099-9.

Soltani:2020:RCA

[SP20] Parisa Soltani and Romeo Patini. Retracted COVID-19 articles: a side-effect of the hot race to publication. Scientometrics, 125(1):819–822, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03661-9;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03661-9.pdf.

Szomszor:2020:HMT

[SPA20] Martin Szomszor, David A. Pendlebury, and Jonathan Adams. How much is too much? The difference between research influence and self-citation excess. Scientometrics, 123(2):1119–1147, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03417-5;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03417-5.pdf.

Singh:2020:CSV

[SPS20] Vivek Kumar Singh, Rajesh Piryani, and Satya Swarup Srichandan. The case of significant variations in goldgreen and black open access: evidence from Indian research output. Scientometrics, 124(1):515–531, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03472-y.

Shu:2020:RWS

[SQC^{+}20] Fei Shu, Wei Quan, Bikun Chen, Junping Qiu, Cassidy R. Sugimoto, and Vincent Larivière. The role of Web of Science publications in China’s tenure system. Scientometrics, 122(3):1683–1695, March 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03339-x.

Siva:2023:RPB

[SR23] N. Siva and P. Rajendran. Retracted publications in BRICS countries: an analytical study. Scientometrics, 128(12):6313–6333, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04856-6.

Savchenko:2024:ACA

[SR24] Elizaveta Savchenko and Ariel Rosenfeld. Authorship conflicts in academia: an international cross-discipline survey. Scientometrics, 129(4):2101–2121, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04972-x.

Sandoval-Romero:2020:NSR

[SRL20] Vanessa Sandoval-Romero and Vincent Larivière. The national system of researchers in Mexico: implications of publication incentives for researchers in social sciences. Scientometrics, 122(1):99–126, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03285-8.

Solarino:2024:GCG

[SRL24] Angelo M. Solarino, Elizabeth L. Rose, and Cristian Luise. Going complex or going easy? The impact of research questions on citations. Scientometrics, 129(1):127–146, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04907-y.

Schubert:2020:IUL

[SS20a] András Schubert and Gábor Schubert. Internationality at university level. Scientometrics, 123(3):1341–1364, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03443-3;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03443-3.pdf.

Siniksaran:2020:WSS

[SS20b] Enis Siniksaran and M. Hakan Satman. WURS: a simulation software for university rankings — software review. Scientometrics, 122(1):701–717, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03269-8.

Stahlschmidt:2022:IPT

[SS22] Stephan Stahlschmidt and Dimity Stephen. From indexation policies through citation networks to normalized citation impacts: Web of Science, Scopus, and Dimensions as varying resonance chambers. Scientometrics, 127(5): 2413–2431, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04309-6.

Sebo:2023:AGC

[SS23] Paul Sebo and Amrollah Shamsi. Author gender and citation categorization: a study of high-impact medical journals. Scientometrics, 128(11):6299–6306, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04827-x.

Stephen:2024:CCC

[SS24] Dimity Stephen and Stephan Stahlschmidt. Contrasting cross-correlation: Meta-analyses of the associations between citations and 13 altmetrics, incorporating moderating variables. Scientometrics, 129(10):6049–6063, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05006-2.

Singh:2022:ERB

[SSB22] Vivek Kumar Singh, Prashasti Singh, and Sujit Bhattacharya. Exploring the relationship between journals indexed from a country and its research output: an empirical investigation. Scientometrics, 127(6):2933–2966, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04366-x.

Szilagyi:2022:CRR

[SSBC22] Istvan-Szilard Szilagyi, Gregor A. Schittek, and Helmar Bornemann-Cimenti. Citation of retracted research: a case-controlled, ten-year follow-up scientometric analysis of Scott s. Reuben’s malpractice. Scientometrics, 127(5): 2611–2620, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04321-w.

Smolinsky:2021:CVE

[SSC21] Lawrence Smolinsky, Daniel S. Sage, and Aaron Cao. Citations versus expert opinions: citation analysis of featured reviews of the American Mathematical Society. Scientometrics, 126(5):3853–3870, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03894-2.

Scharpf:2023:DRF

[SSC^{+}23] Philipp Scharpf, Moritz Schubotz, Howard S. Cohl, Corinna Breitinger, and Bela Gipp. Discovery and recognition of formula concepts using machine learning. Scientometrics, 128(9):4971–5025, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04667-9.

Sachini:2021:CEC

[SSCK21] E. Sachini, K. Sioumalas-Christodoulou, and N. Karampekios. COVID-19 enabled co-authoring networks: a country-case analysis. Scientometrics, 126(6):5225–5244, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03952-9.

Santos:2022:DTS

[SSdCRD22] Breno Santana Santos, Ivanovitch Silva, and Marcel da Câmara Ribeiro-Dantas. Discovering temporal scientometric knowledge in COVID-19 scholarly production. Scientometrics, 127(3):1609–1642, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04260-y.

Schaefermeier:2021:TST

[SSH21] Bastian Schaefermeier, Gerd Stumme, and Tom Hanika. Topic space trajectories. Scientometrics, 126(7):5759– 5795, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03931-0.

Sokolov:2021:BBP

[SSK21] Alexander Sokolov, Sergey Shashnov, and Maxim Kotsemir. From BRICS to BRICS plus: selecting promising areas of S&T cooperation with developing countries. Scientometrics, 126(11):8815–8859, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04142-3.

Sotudeh:2022:LSE

[SSK22] Hajar Sotudeh, Zeinab Saber, and Farshad Khunjush. A longitudinal study of the evolution of opinions about open access and its main features: a Twitter sentiment analysis. Scientometrics, 127(10):5587–5611, October 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04502-7.

Singh:2022:RGS

[SSL22] Vivek Kumar Singh, Satya Swarup Srichandan, and Hiran H. Lathabai. ResearchGate and Google Scholar: how much do they differ in publications, citations and different metrics and why? Scientometrics, 127(3):1515–1542, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04264-2.

Singh:2021:JCW

[SSM21] Vivek Kumar Singh, Prashasti Singh, and Philipp Mayr. The journal coverage of Web of Science, Scopus and Dimensions: a comparative analysis. Scientometrics, 126(6): 5113–5142, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03948-5.

Sani:2024:ALI

[SSS^{+}24] Mad Khir Johari Abdullah Sani, Sharunizam Shari, Noor Zaidi Sahid, Norshila Shaifuddin, Zuraidah Abdul Manaf, and Alexander van Servellen. ASEAN Library and Information Science (LIS) research (2018–2022): a bibliometric analysis with strategies for enhanced global impact. Scientometrics, 129(1):95–125, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04878-0.

Shamsi:2022:GZB

[SSSd22] Amrollah Shamsi, Rafaela Carolina Silva, and Karen Santos-d’Amorim. A grey zone for bibliometrics: publications indexed in Web of Science as anonymous. Scientometrics, 127(10):5989–6009, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04494-4.

Shang:2022:GDA

[SSZ22] Yuanyuan Shang, Gunnar Sivertsen, and Lin Zhang. Gender differences among first authors in research focused on the Sustainable Development Goal of Gender Equality. Scientometrics, 127(8):4769–4796, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04430-6.

Spiewanowski:2021:JRP

[ŚT21] Piotr Śpiewanowski and Oleksandr Talavera. Journal rankings and publication strategy. Scientometrics, 126 (4):3227–3242, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-021-
03891-5;   http://link.springer.com/content/pdf/
10.1007/s11192-021-03891-5.pdf.

Stacey:2021:ACR

[Sta21] Anthony G. Stacey. Ages of cited references and growth of scientific knowledge: an explication of the gamma distribution in business and management disciplines. Scientometrics, 126(1):619–640, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03761-6.

Sastron-Toledo:2024:COA

[STAÁMR24] Pablo Sastrón-Toledo, Patricia Alonso-Álvarez, and Jorge Mañana-Rodríguez. Compliance with open access mandates and its effects on research visibility: the case of the Spanish National Plan of R&D. Scientometrics, 129(4): 2057–2081, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04978-5.

Stek:2021:IST

[Ste21] Pieter E. Stek. Identifying spatial technology clusters from patenting concentrations using heat map kernel density estimation. Scientometrics, 126(2):911–930, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03751-8;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03751-8.pdf.

Stephen:2022:PRE

[Ste22] Dimity Stephen. Peer reviewers equally critique theory, method, and writing, with limited effect on the final content of accepted manuscripts. Scientometrics, 127(6): 3413–3435, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04357-y.

Stephen:2023:MAQ

[Ste23] Dimity Stephen. Medical articles in questionable journals are less impactful than those in non-questionable journals but still extensively cited. Scientometrics, 128(8): 4509–4522, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04763-w.

Singla:2024:ITM

[STM^{+}24] Priyanka Singla, Siny Tsang, Michael Mazzeffi, Lucy Carr Jones, Andrea H. Denton, and Nabil Elkassabany. Impact of transition of medical journals from print to online-only on their scholarly performance. Scientometrics, 129(5): 2855–2866, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04998-1.

Sugimoto:2023:LOP

[Sug23] Cassidy R. Sugimoto. Laudation on the occasion of the presentation of the Derek de Solla Price Award 2023 to Kevin W. Boyack and Richard Klavans. Scientometrics, 128(12):6309–6312, December 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04848-6.

Sundling:2023:ACA

[Sun23] Pär Sundling. Author contributions and allocation of authorship credit: testing the validity of different counting methods in the field of chemical biology. Scientometrics, 128(5):2737–2762, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04680-y.

Salas-Velasco:2020:MEP

[SV20] Manuel Salas-Velasco. Measuring and explaining the production efficiency of Spanish universities using a nonparametric approach and a bootstrapped-truncated regression. Scientometrics, 122(2):825–846, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03324-4.

Singh:2020:MTI

[SVC20] Vibhav Singh, Surabhi Verma, and Sushil S. Chaurasia. Mapping the themes and intellectual structure of corporate university: co-citation and cluster analyses. Scientometrics, 122(3):1275–1302, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03328-0.

Stiller:2021:DCR

[SvWC21] Ingo Stiller, Arjen van Witteloostuijn, and Bart Cambré. Do current radical innovation measures actually measure radical drug innovation? Scientometrics, 126(2):1049– 1078, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03778-x.

Seppanen:2022:CCP

[SVY22] Janne-Tuomas Seppänen, Hanna Värri, and Irene Ylönen. Co-citation Percentile Rank and JYUcite: a new networkstandardized output-level citation influence metric and its implementation using Dimensions API. Scientometrics, 127(6):3523–3541, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-022-
04393-8.

Shibayama:2020:MOS

[SW20] Sotaro Shibayama and Jian Wang. Measuring originality in science. Scientometrics, 122(1):409–427, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03263-0;   http:
//link.springer.com/content/pdf/10.1007/s11192-
019-03263-0.pdf.

Shi:2024:WSS

[SW24] Wenxuan Shi and Renli Wu. Women’s strength in science: exploring the influence of female participation on research impact and innovation. Scientometrics, 129(7): 4529–4551, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05089-x.

Shu:2023:GIN

[SWL^{+}23] Fei Shu, Xiaojian Wang, Sichen Liu, Junping Qiu, and Vincent Larivière. Global impact or national accessibility? A paradox in China’s science. Scientometrics, 128(1): 263–277, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04537-w.

Shahmandi:2020:NAZ

[SWT20] Marzieh Shahmandi, Paul Wilson, and Mike Thelwall. A new algorithm for zero-modified models applied to citation counts. Scientometrics, 125(2):993–1010, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03654-8.

Shen:2021:CBS

[SXLC21] Hongquan Shen, Juan Xie, Jiang Li, and Ying Cheng. The correlation between scientific collaboration and citation count at the paper level: a meta-analysis. Scientometrics, 126(4):3443–3470, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03888-0.

Schneider:2020:CPR

[SYHW20] Jodi Schneider, Di Ye, Alison M. Hill, and Ashley S. Whitehorn. Continued post-retraction citation of a fraudulent clinical trial report, 11 40years after it was retracted for falsifying data. Scientometrics, 125(3):2877–2913, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03631-1;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03631-1.pdf.

Soehartono:2022:ESP

[SYK22] A. M. Soehartono, L. G. Yu, and K. A. Khor. Essential signals in publication trends and collaboration patterns in global Research Integrity and Research Ethics (RIRE). Scientometrics, 127(12):7487–7497, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04400-y.

Syafiandini:2024:EBS

[SYL^{+}24] Arida Ferti Syafiandini, Jeeyoung Yoon, Soobin Lee, Chaemin Song, Erjia Yan, and Min Song. Examining between-sectors knowledge transfer in the pharmacology field. Scientometrics, 129(6):3115–3147, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05040-0.

Shi:2020:NII

[SZ20] Xiaoxiao Shi and Qingpu Zhang. Network inertia and inbound open innovation: is there a bidirectional relationship? Scientometrics, 122(2):791–815, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03321-7.

Sun:2020:RRO

[SZK20] Yutao Sun, Chen Zhang, and Robert A. W. Kok. The role of research outcome quality in the relationship between university research collaboration and technology transfer: empirical results from China. Scientometrics, 122(2):1003– 1026, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03330-6.

Sarwar:2021:WES

[SZN^{+}21] Raheem Sarwar, Afifa Zia, Raheel Nawaz, Ayman Fayoumi, Naif Radi Aljohani, and Saeed-Ul Hassan. Webometrics: evolution of social media presence of universities. Scientometrics, 126(2):951–967, February 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03804-y.

Sun:2024:DCJ

[SZP^{+}24] Zhuanlan Sun, Chenwei Zhang, Ka Lok Pang, Ying Tang, and Yiwei Li. Do changes in journal rank influence publication output? Evidence from China. Scientometrics, 129(11):7035–7054, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05167-0.

Song:2024:EWS

[SZY24] Le Song, Guilong Zhu, and Xiao Yin. Evaluating the wisdom of scholar crowds from the perspective of knowledge diffusion. Scientometrics, 129(9):5103–5139, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05090-4.

Sun:2023:RSI

[SZZ23] Yutao Sun, Ying Zhang, and Xiaofei Zhang. Reconfiguring star inventors with commercialization: a case of the graphene sector. Scientometrics, 128(10):5411–5440, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04795-2.

Tonta:2020:DMS

[TA20] Yasar Tonta and Müge Akbulut. Does monetary support increase citation impact of scholarly papers? Scientometrics, 125(2):1617–1641, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03688-y.

Taheri:2022:RTP

[TA22] Soroush Taheri and Sadegh Aliakbary. Research trend prediction in computer science publications: a deep neural network approach. Scientometrics, 127(2):849–869, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04240-2.

Taskin:2021:FFL

[Tas21] Zehra Taskin. Forecasting the future of library and information science and its sub-fields. Scientometrics, 126(2):1527–1551, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03800-2;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03800-2.pdf.

Taylor:2020:AAA

[Tay20] Michael Taylor. An altmetric attention advantage for open access books in the humanities and social sciences. Scientometrics, 125(3):2523–2543, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03735-8.

Taylor:2023:SSQ

[Tay23] Michael Taylor. Slow, slow, quick, quick, slow: five altmetric sources observed over a decade show evolving trends, by research age, attention source maturity and open access status. Scientometrics, 128(4):2175–2200, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04653-1.

Thierry:2023:RSR

[TBA23] Nimbeshaho Thierry, Bing-Kun Bao, and Zafar Ali. RAR-SB: research article recommendation using SciBERT with BiGRU. Scientometrics, 128(12):6427–6448, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04840-0.

Tuesta:2020:CNB

[TBPN^{+}20] Esteban Fernández Tuesta, Máxima Bolaños-Pizarro, Daniel Pimentel Neves, Geziel Fernández, and Justin Axel-Berg. Complex networks for benchmarking in global universities rankings. Scientometrics, 125(1):405–425, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03637-9.

Thakuria:2024:DSI

[TD24a] Abhijit Thakuria and Dipen Deka. A decadal study on identifying latent topics and research trends in open access LIS journals using topic modeling approach. Scientometrics, 129(7):3841–3869, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05058-4.

Tutuncu:2024:GGM

[TD24b] Lokman Tutuncu and Rahman Dag. A gender gap in the manuscript review time of Turkish national journals? Scientometrics, 129(12):7783–7803, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05177-y.

Tzitzikas:2024:HCA

[TD24c] Yannis Tzitzikas and Giorgos Dovas. How co-authorship affects the H-index? Scientometrics, 129(7):4437–4469, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05088-y.

Tran:2024:IWC

[TDM24] Huu Dat Tran, Pham Phuong Uyen Diep, and Nabila Mushtarin. “Insert witty catchphrase here!”: do title elements influence engagement and citation? — Examining highly-cited research articles in media, communication, and related disciplines. Scientometrics, 129(11):6675– 6691, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05152-7.

Tsigaris:2020:RIC

[TdS20] Panagiotis Tsigaris and Jaime A. Teixeira da Silva. Reproducibility issues with correlating Beall-listed publications and research awards at a small Canadian business school. Scientometrics, 123(1):143–157, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03353-4.

Teixeira:2020:MSF

[TdSFB20] Elizabeth C. Teixeira, Victor E. L. da Silva, Nidia N. Fabré, and Vandick S. Batista. Marine shrimp fisheries research — a mismatch on spatial and thematic needs. Scientometrics, 122(1):591–606, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03276-9.

Tang:2024:IRH

[TDWW24] Xiaobo Tang, Xin Du, Qiongfu Wang, and Jialin Wu. Intelligent recognition of high-quality academic papers: based on knowledge-based metasemantic networks. Scientometrics, 129(11):6779–6812, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05157-2.

TeixeiradaSilva:2020:EPT

[Tei20] Jaime A. Teixeira da Silva. The ethics of publishing in two languages. Scientometrics, 123(1):535–541, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03363-2.

daSilva:2021:MEI

[Tei21] Jaime A. Teixeira da Silva. The Matthew effect impacts science and academic publishing by preferentially amplifying citations, metrics and status. Scientometrics, 126(6): 5373–5377, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03967-2.

TeixeiradaSilva:2024:CPP

[Tei24a] Jaime A. Teixeira da Silva. Citations to papers published in European Science Editing from 2020 to 2022: assessment using Scopus, Dimensions, Google Scholar, and Altmetrics. Scientometrics, 129(3):1969–1974, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04953-0.

daSilva:2024:IRA

[Tei24b] Jaime A. Teixeira da Silva. The inaccurate representation of an author’s publishing name, and impact on reference accuracy. Scientometrics, 129(5):2923–2932, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05029-9.

Tian:2024:MDA

[TFWC24] Wencan Tian, Zhichao Fang, Xianwen Wang, and Rodrigo Costas. A multi-dimensional analysis of usage counts, Mendeley readership, and citations for journal and conference papers. Scientometrics, 129(2):985–1013, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04909-w.

Tal:2020:PRF

[TG20] Diana Tal and Avishag Gordon. Propaganda as a research field: a bibliometric study. Scientometrics, 122(1): 741–750, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03298-3.

Tuncer:2024:IPT

[TG24] A. Tuncer and F. Gezici. Influence of proximities and their interaction effects on scientific collaborations: the case of Turkish regions. Scientometrics, 129(3):1277–1298, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04919-8.

Thelwall:2020:DBC

[The20a] Mike Thelwall. Data in brief: Can a mega-journal for data be useful? Scientometrics, 124(1):697–709, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03437-1.

Thelwall:2020:MCF

[The20b] Mike Thelwall. Mid-career field switches reduce gender disparities in academic publishing. Scientometrics, 123(3): 1365–1383, June 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03445-1.

Thelwall:2021:AMW

[The21] Mike Thelwall. Alternative medicines worth researching? Citation analyses of acupuncture, chiropractic, homeopathy, and osteopathy 1996–2017. Scientometrics, 126(10): 8731–8747, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04145-0.

Thelwall:2023:SCA

[The23] Mike Thelwall. Are successful co-authors more important than first authors for publishing academic journal articles? Scientometrics, 128(4):2211–2232, April 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04663-z.

Toth:2023:BVP

[THG23] János József Tóth, Gergő Háló, and Manuel Goyanes. Beyond views, productivity, and citations: measuring geopolitical differences of scientific impact in communication research. Scientometrics, 128(10):5705–5729, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04801-7.

Thijs:2020:UNN

[Thi20] Bart Thijs. Using neural-network based paragraph embeddings for the calculation of within and between document similarities. Scientometrics, 125(2):835–849, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03583-6.

Taubert:2023:UDO

[THJ^{+}23] Niels Taubert, Anne Hobert, Najko Jahn, Andre Bruns, and Elham Iravani. Understanding differences of the OA uptake within the German university landscape (2010– 2020): part 1 — journal-based OA. Scientometrics, 128 (6):3601–3625, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04716-3.

Taubert:2024:UDO

[THJ^{+}24] Niels Taubert, Anne Hobert, Najko Jahn, Andre Bruns, and Elham Iravani. Understanding differences of the OA uptake within the German University landscape (2010–2020): Part 2 — repository-provided OA. Scientometrics, 129(5):2801–2825, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05003-5.

Traylor:2023:DJI

[THL23] Christopher Traylor and Christoph Herrmann-Lingen. Does the journal impact factor reflect the impact of German medical guideline contributions? Scientometrics, 128 (3):1951–1962, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-
04647-z.

Thiele:2021:CTR

[THvB21] Christian Thiele, Gerrit Hirschfeld, and Ruth von Brachel. Clinical trial registries as scientometric data: A novel solution for linking and deduplicating clinical trials from multiple registries. Scientometrics, 126(12):9733–9750, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04111-w.

Thelwall:2021:RAT

[TK21] Mike Thelwall and Kayvan Kousha. Researchers’ attitudes towards the h-index on Twitter 2007–2020: criticism and acceptance. Scientometrics, 126(6):5361–5368, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03961-8.

Thelwall:2023:WFD

[TKM^{+}23] Mike Thelwall, Kayvan Kousha, Meiko Makita, Mahshid Abdoli, Emma Stuart, Paul Wilson, and Jonathan Levitt. In which fields do higher impact journals publish higher quality articles? Scientometrics, 128(7):3915–3933, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04735-0.

Takahashi:2023:NCA

[TKS^{+}23] Ryo Takahashi, Kenji Kaibe, Kazuyuki Suzuki, Sayaka Takahashi, Kotaro Takeda, Marc Hansen, and Michiaki Yumoto. New concept of the affinity between research fields using academic journal data in Scopus. Scientometrics, 128(6):3507–3534, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04711-8.

Tang:2023:SBG

[TKZ^{+}23] Li Tang, Jennifer Kuzma, Xi Zhang, Xinyu Song, Yin Li, Hongxu Liu, and Guangyuan Hu. Synthetic biology and governance research in China: a 40-year evolution. Scientometrics, 128(9):5293–5310, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04789-0.

Tian:2024:AAT

[TLL24] Dan Tian, Xin Liu, and Jiang Li. Accelerated acceptance time for preprint submissions: a comparative analysis based on PubMed. Scientometrics, 129(7):3787–3807, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05056-6.

Tang:2022:IAE

[TLM22] Xuli Tang, Xin Li, and Feicheng Ma. Internationalizing AI: evolution and impact of distance factors. Scientometrics, 127(1):181–205, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04207-3.

Tarkhan-Mouravi:2020:TII

[TM20] Sandro Tarkhan-Mouravi. Traditional indicators inflate some countries’ scientific impact over 10 times. Scientometrics, 123(1):337–356, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03372-1.

Tur:2025:REA

[TM25] Elena M. Tur and Arjan Markus. Repeated examiner– attorney interaction and patent approval. Scientometrics, 130(2):619–639, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-025-
05239-9.

Toth:2024:RRA

[TMNH^{+}24] Barbara Tóth, Hossein Motahari-Nezhad, Nicki Horseman, László Berek, Levente Kovács, Áron Hölgyesi, Márta Péntek, Seyedali Mirjalili, László Gulácsi, and Zsombor Zrubka. Ranking resilience: assessing the impact of scientific performance and the expansion of the Times Higher Education Word University Rankings on the position of Czech, Hungarian, Polish, and Slovak universities. Scientometrics, 129(3):1739–1770, March 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04920-1.

Thomson:2023:CRA

[TMW23] Ryan Thomson, Rebecca Mosier, and Michelle Worosz. COVID research across the social sciences in 2020: a bibliometric approach. Scientometrics, 128(6):3377–3399, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04714-5.

Telarico:2024:RDC

[TMZ24] Fabio Ashtar Telarico, Franc Mali, and Ales Ziberna. Revealing dynamic co-authorship structure in the social sciences through blockmodeling: the Slovenian case (1991–2020). Scientometrics, 129(9):5635–5672, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05130-z.

Tattershall:2020:DBT

[TNS20] E. Tattershall, G. Nenadic, and R. D. Stevens. Detecting bursty terms in computer science research. Scientometrics, 122(1):681–699, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03307-5;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03307-5.pdf.

Tattershall:2021:MTL

[TNS21] E. Tattershall, G. Nenadic, and R. D. Stevens. Modelling trend life cycles in scientific research using the Logistic and Gompertz equations. Scientometrics, 126(11):9113– 9132, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04137-0.

Tontodimamma:2021:TYR

[TNSF21] Alice Tontodimamma, Eugenia Nissi, Annalina Sarra, and Lara Fontanella. Thirty years of research into hate speech: topics of interest and their evolution. Scientometrics, 126(1):157–179, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03737-6;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03737-6.pdf.

Tol:2024:NF

[Tol24] Richard S. J. Tol. The Nobel family. Scientometrics, 129(3):1329–1346, March 2024. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://arxiv.org/abs/2303.06106;   https:
//link.springer.com/article/10.1007/s11192-024-
04936-1.

Tomaszewski:2021:SCS

[Tom21] Robert Tomaszewski. A study of citations to STEM databases: ACM Digital Library, Engineering Village, IEEE Xplore, and MathSciNet. Scientometrics, 126(2): 1797–1811, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03795-w.

Tomaszewski:2023:VIA

[Tom23] Robert Tomaszewski. Visibility, impact, and applications of bibliometric software tools through citation analysis. Scientometrics, 128(7):4007–4028, July 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04725-2.

Tomaszewski:2024:TTS

[Tom24] Robert Tomaszewski. Time-tracking substances: harnessing the power of article counts over time with SciFinder^{n}. Scientometrics, 129(8):4819–4845, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05098-w.

Torrisi:2020:IRB

[TP20] Benedetto Torrisi and Giuseppe Pernagallo. Investigating the relationship between job satisfaction and academic brain drain: the Italian case. Scientometrics, 124(2): 925–952, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03509-2.

Thelwall:2024:AFC

[TP24] Mike Thelwall and Stephen Pinfield. The accuracy of field classifications for journals in Scopus. Scientometrics, 129(2):1097–1117, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04901-4.

Tapia-Pacheco:2021:RNA

[TPVVPA21] Diana Tapia-Pacheco, Laura Liliana Villa-Vázquez, and Miguel Ángel Pérez-Angón. Research networks on the access of drinking water in Mexico City (2004–2018). Scientometrics, 126(3):2557–2573, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03569-4.

Tohalino:2021:ARB

[TQA21] Jorge A. V. Tohalino, Laura V. C. Quispe, and Diego R. Amancio. Analyzing the relationship between text features and grants productivity. Scientometrics, 126(5): 4255–4275, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03926-x.

Tohalino:2024:UWE

[TSA24] Jorge A. V. Tohalino, Thiago C. Silva, and Diego R. Amancio. Using word embedding to detect keywords in texts modeled as complex networks. Scientometrics, 129(7): 3599–3623, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05055-7.

Torres-Salinas:2023:BD

[TSAMRG23] Daniel Torres-Salinas, Wenceslao Arroyo-Machado, and Nicolas Robinson-Garcia. Bibliometric denialism. Scientometrics, 128(9):5357–5359, September 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04787-2. See letter [Rus23].

Torres-Salinas:2021:EWI

[TSAMT21] Daniel Torres-Salinas, Wenceslao Arroyo-Machado, and Mike Thelwall. Exploring WorldCat identities as an altmetric information source: a library catalog analysis experiment in the field of Scientometrics. Scientometrics, 126(2):1725–1743, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03814-w.

Torres-Salinas:2024:MPS

[TSDAMRG24] Daniel Torres-Salinas, Domingo Docampo, Wenceslao Arroyo-Machado, and Nicolas Robinson-Garcia. The many publics of science: using altmetrics to identify common communication channels by scientific field. Scientometrics, 129(7):3705–3723, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05077-1.

Thelwall:2023:WRF

[TSVV23] Mike Thelwall, Subreena Simrick, Ian Viney, and Peter Van den Besselaar. What is research funding, how does it influence research, and how is it recorded? Key dimensions of variation. Scientometrics, 128(11):6085–6106, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04836-w.

Turki:2020:FCW

[TTB20] Houcemeddine Turki, Mohamed Ali Hadj Taieb, and Mohamed Ben Aouicha. Facts to consider when analyzing the references of Nobel Prize scientific background. Scientometrics, 124(1):787–790, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03456-y. See [Bjø20b] and response

[Bjø20c, Bjø20a].

Turki:2022:ASB

[TTB22] Houcemeddine Turki, Mohamed Ali Hadj Taieb, and Mohamed Ben Aouicha. Awakening sleeping beauties during the COVID-19 pandemic influences the citation impact of their references. Scientometrics, 127(10):6047–6050, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04501-8.

Turki:2020:NSW

[TTBA20] Houcemeddine Turki, Mohamed Ali Hadj Taieb, Mohamed Ben Aouicha, and Ajith Abraham. Nature or Science: what Google Trends says. Scientometrics, 124(2):1367– 1385, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03511-8.

Tuesta:2025:DDS

[TTDH25] Geziel Fernandez Tuesta, Esteban Fernandez Tuesta, Luciano Antônio Digiampietri, and Dominik Hartmann. The diversity and (dis-)similarity of Brazilian universities’ research portfolios. Scientometrics, 130(1):237–266, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05191-0.

Taskin:2022:FAT

[TTDK22] Zehra Taşkın, Abdülkadir Taşkın, Güleda Doğan, and Emanuel Kulczycki. Factors affecting time to publication in information science. Scientometrics, 127(12):7499– 7515, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04296-8.

Tutuncu:2023:API

[Tut23] Lokman Tutuncu. All-pervading insider bias alters review time in Turkish university journals. Scientometrics, 128 (6):3743–3791, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04724-3.

Tutuncu:2024:GGI

[Tut24] Lokman Tutuncu. Gatekeepers or gatecrashers? The inside connection in editorial board publications of Turkish national journals. Scientometrics, 129(2):957–984, February 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04905-0.

Tian:2024:DGC

[TWH^{+}24] Wencan Tian, Yongzhen Wang, Zhigang Hu, Ruonan Cai, Guangyao Zhang, and Xianwen Wang. Does Granger causality exist between article usage and publication counts? A topic-level time-series evidence from IEEE Xplore. Scientometrics, 129(6):3285–3302, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05038-8.

Tian:2021:ANC

[TXL21] Shanwu Tian, Xiurui Xu, and Ping Li. Acknowledgement network and citation count: the moderating role of collaboration network. Scientometrics, 126(9):7837– 7857, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04090-y.

Tutuncu:2022:AFW

[TYS22] Lokman Tutuncu, Recep Yucedogru, and Idris Sarisoy. Academic favoritism at work: insider bias in Turkish national journals. Scientometrics, 127(5):2547–2576, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04355-0.

Teng:2024:MGD

[TZ24] Ziyou Teng and Xuezhong Zhu. Measuring the global and domestic technological impact of Chinese scientific output: a patent-to-paper citation analysis of science-technology linkage. Scientometrics, 129(9):5181–5210, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05110-3.

Taros:2023:RCA

[TZY^{+}23] Trenton Taros, Christopher Zoppo, Nathan Yee, Jack Hanna, and Christine MacGinnis. Retracted Covid-19 articles: significantly more cited than other articles within their journal of origin. Scientometrics, 128(5):2935–2943, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04707-4.

Tan:2023:SSS

[TZZZ23] Shicheng Tan, Tao Zhang, Shu Zhao, and Yanping Zhang. Self-supervised scientific document recommendation based on contrastive learning. Scientometrics, 128(9):5027– 5049, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04782-7.

Urbano:2020:CDC

[UA20] Cristóbal Urbano and Jordi Ardanuy. Cross-disciplinary collaboration versus coexistence in LIS serials: analysis of authorship affiliations in four European countries. Scientometrics, 124(1):575–602, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03471-z;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03471-z.pdf.

Ullah:2021:RRG

[UA21] Ahsan Ullah and Kanwal Ameen. Relating research growth, authorship patterns and publishing outlets: a bibliometric study of LIS articles produced by Pakistani authors. Scientometrics, 126(9):8029–8047, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04081-z.

Uribe-Bohorquez:2023:GDA

[UBROGS23] Maria-Victoria Uribe-Bohorquez, Juan-Camilo Rivera-Ordóñez, and Isabel-María García-Sánchez. Gender disparities in accounting academia: analysis from the lens of publications. Scientometrics, 128(7):3827–3865, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04718-1.

Bajwa:2020:EBU

[uHBKK20] Nida ul Habib Bajwa, Cornelius J. König, and Thiemo Kunze. Evidence-based understanding of introductions of research articles. Scientometrics, 124(1):195–217, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-020-03475-9. See correction

[uHBKK21].

Bajwa:2021:CEB

[uHBKK21] Nida ul Habib Bajwa, Cornelius J. König, and Thiemo Kunze. Correction to: Evidence-based understanding of introductions of research articles. Scientometrics, 126(10): 8773–8774, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04122-7.

See [uHBKK20].

Uddin:2021:RIS

[UIM21] Shahadat Uddin, Tasadduq Imam, and Mohammad Mozumdar. Research interdisciplinarity: STEM versus non-STEM. Scientometrics, 126(1):603–618, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03750-9;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03750-9.pdf.

Ujum:2021:NJP

[UKP21] Ephrance Abu Ujum, Sameer Kumar, and Gangan Prathap. A new journal power-weakness ratio to measure journal impact. Scientometrics, 126(11):9051–9068, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04132-5.

Usman:2021:RAA

[UMA21] Muhammad Usman, Ghulam Mustafa, and Muhammad Tanvir Afzal. Ranking of author assessment parameters using Logistic Regression. Scientometrics, 126(1): 335–353, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03769-y.

Uyar:2023:EQI

[UNK23] Ali Uyar, Khalil Nimer, and Cemil Kuzey. Education quality, internet access in schools, and research performance in management and accounting domains: a cross-country investigation. Scientometrics, 128(10):5441–5475, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04813-3.

Urban:2025:UFA

[Urb25] Timothy L. Urban. On the use of factor analysis for bibliometric indicators. Scientometrics, 130(7):4027–4060, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05359-2.

Vital:2022:CAL

[VA22] Adilson Vital and Diego R. Amancio. A comparative analysis of local similarity metrics and machine learning approaches: application to link prediction in author citation networks. Scientometrics, 127(10):6011–6028, October 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04484-6.

Valderrama:2025:CAM

[VAMBTS25] Pilar Valderrama, Wenceslao Arroyo-Machado, Adela Baca, and Daniel Torres-Salinas. Can altmetric mentions reflect the quality of evidence? A study in biomedical and life sciences. Scientometrics, 130(4):2345–2356, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05292-4.

Vermaut:2024:UIS

[VBF24] Benjamin Vermaut, Corentin Burnay, and Stéphane Faulkner. Unveiling the intellectual structure of soccer performance through keywords co-occurrence: a nested bibliometric approach. Scientometrics, 129(12):7501–7534, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05171-4.

Viggiani:2020:DFD

[VC20] Eloisa Viggiani and Luciana Calabró. Does faculty disciplinary background play a role in the publication pattern of an interdisciplinary research area? The case of science education in Brazil. Scientometrics, 125(2):893–908, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-020-03593-4.

Vieira:2022:IAC

[VC22] Elizabeth S. Vieira and Jorge Cerdeira. The integration of African countries in international research networks. Scientometrics, 127(4):1995–2021, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04297-7.

Vasconcelos:2023:BAS

[VCD^{+}23] Rodrigo N. Vasconcelos, Diego Pereira Costa, Soltan Galano Duverger, Jocimara S. B. Lobão, Elaine C. B. Cambuí, Carlos A. D. Lentini, André T. Cunha Lima, Juliano Schirmbeck, Deorgia Tayane Mendes, Washington J. S. Franca Rocha, and Milton J. Porsani. Bibliometric analysis of surface water detection and mapping using remote sensing in South America. Scientometrics, 128(3): 1667–1688, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04570-9.

Vakkari:2022:LCL

[VCJ22] Pertti Vakkari, Yu-Wei Chang, and Kalervo Järvelin. Largest contribution to LIS by external disciplines as measured by the characteristics of research articles. Scientometrics, 127(8):4499–4522, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04452-0.

Vervenne:2022:WED

[VCV22] Jan-Bart Vervenne, Julie Callaert, and Bart Van Looy. To what extent do SMEs contribute to Europe’s patent stock? A methodological outline for creating a Europewide SME technology indicator. Scientometrics, 127(6): 3049–3082, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04360-3.

vanDalen:2021:HPP

[vD21] Hendrik P. van Dalen. How the publish-or-perish principle divides a science: the case of economists. Scientometrics, 126(2):1675–1694, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03786-x;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03786-x.pdf.

vandenBesselaar:2025:CAI

[vdB25] Peter van den Besselaar. Circling around interdisciplinarity. Scientometrics, 130(6):3279–3291, June 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05346-7.

Viglioni:2020:IRL

[VdBC20] Marco Túlio Dinali Viglioni, Mozar José de Brito, and Cristina Lelis Leal Calegario. Innovation and R&D in Latin America and the Caribbean countries: a systematic literature review. Scientometrics, 125(3):2131–2167, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03714-z.

vandenBesselaar:2024:TGB

[vdBM24] Peter van den Besselaar and Charlie Mom. Is there gender bias in awarding cum laude for the PhD thesis? Scientometrics, 129(10):6349–6371, October 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04985-6.

Verreynne:2020:WMF

[VdOS^{+}20] Martie-Louise Verreynne, Rui Torres de Oliveira, John Steen, Marta Indulska, and Jerad A. Ford. What motivates ‘free’ revealing? Measuring outbound non-pecuniary openness, innovation types and expectations of future profit growth. Scientometrics, 124(1):271–301, July 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03434-4;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03434-4.pdf.

vanderPol:2024:IFI

[vdP24] Johannes van der Pol. The impact of funding on the 5G innovation ecosystem. Scientometrics, 129(4):2037–2056, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04954-z.

vanderPol:2025:FNR

[vdPF25] Janpieter van der Pol and Koen Frenken. Fragmentation of national research systems: the case of The Netherlands. Scientometrics, 130(6):3195–3211, June 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05310-5.

Vernon:2021:DMI

[VDY21] Marlo M. Vernon, C. Makenzie Danley, and Frances M. Yang. Developing a measure of innovation from research in higher education data. Scientometrics, 126(5):3919–3928, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03916-z.

Velez-Estevez:2022:WDP

[VEGSMMC22] A. Velez-Estevez, P. García-Sánchez, J. A. Moral-Munoz, and M. J. Cobo. Why do papers from international collaborations get more citations? A bibliometric analysis of Library and Information Science papers. Scientometrics, 127(12):7517–7555, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04486-4.

Ventura:2024:USM

[Ven24] Rafael Ventura. The use of scientific methods and models in the philosophy of science. Scientometrics, 129(3): 1255–1276, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04931-6.

Vahdati:2021:CQA

[VFL^{+}21] Sahar Vahdati, Said Fathalla, Christoph Lange, Andreas Behrend, Aysegul Say, Zeynep Say, and Sören Auer. A comprehensive quality assessment framework for scientific events. Scientometrics, 126(1):641–682, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03758-1;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03758-1.pdf.

Verma:2024:SDD

[VHM24] Shilpa Verma, Sandeep Harit, and Kundan Munjal. SRRS: Design and development of a Scholarly Reciprocal Recommendation System. Scientometrics, 129(11):6839–6866, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05143-8.

Vieira:2022:IRC

[Vie22] Elizabeth S. Vieira. International research collaboration in Africa: a bibliometric and thematic analysis. Scientometrics, 127(5):2747–2772, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04349-y.

Vieira:2023:IRC

[Vie23] Elizabeth S. Vieira. The influence of research collaboration on citation impact: the countries in the European Innovation Scoreboard. Scientometrics, 128(6):3555–3579, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04715-4.

Vieira:2024:RCK

[Vie24] Elizabeth S. Vieira. Research collaboration and knowledge development in Africa: the case of neonatal conditions. Scientometrics, 129(4):2407–2431, April 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04971-y.

Vinkler:2021:EPP

[Vin21] Péter Vinkler. Evaluation of publications by the partset method. Scientometrics, 126(4):2737–2757, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03841-7;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03841-7.pdf.

Vinkler:2023:INR

[Vin23] Péter Vinkler. Impact of the number and rank of coauthors on h-index and π-index. The part-impact method. Scientometrics, 128(4):2349–2369, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04643-3.

Valderrama:2022:IBI

[VJCV22] Pilar Valderrama, Evaristo Jiménez-Contreras, and Mariano J. Valderrama. Introducing a bibliometric index based on factor analysis. Scientometrics, 127(1):509–522, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04195-4.

Verma:2023:SAF

[VKY23] Manoj Kumar Verma, Daud Khan, and Mayank Yuvaraj. Scientometric assessment of funded scientometrics and bibliometrics research (2011–2021). Scientometrics, 128(8): 4305–4320, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04767-6.

Vara:2022:AMC

[VMF22] Narjes Vara, Mahdieh Mirzabeigi, and Seyed Mostafa Fakhrahmad. Application of k-means clustering algorithm to improve effectiveness of the results recommended by journal recommender system. Scientometrics, 127(6): 3237–3252, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04397-4.

Viiu:2021:CIA

[VP21a] Gabriel-Alexandru Vîiu and Mihai Paunescu. The citation impact of articles from which authors gained monetary rewards based on journal metrics. Scientometrics, 126(6): 4941–4974, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03944-9.

Viiu:2021:LMB

[VP21b] Gabriel-Alexandru Vîiu and Mihai Paunescu. The lack of meaningful boundary differences between journal impact factor quartiles undermines their independent use in research evaluation. Scientometrics, 126(2):1495– 1525, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03801-1.

Vercelli:2023:SCB

[VPC^{+}23] Stefano Vercelli, Leonardo Pellicciari, Andrea Croci, Cesare Maria Cornaggia, Francesca Cecchi, and Daniele Piscitelli. Self-citation behavior within the health allied professions’ scientific sector in Italy: a bibliometric analysis. Scientometrics, 128(2):1205–1217, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04599-w.

Vazquez:2022:VST

[VPDAG22] Manuel A. Vázquez, Jorge Pereira-Delgado, and Jerónimo Arenas-García. Validation of scientific topic models using graph analysis and corpus metadata. Scientometrics, 127(9):5441–5458, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04318-5.

Viegas:2022:SAP

[VPR22] Felipe Viegas, Antônio Pereira, and Leonardo Rocha. Semantic Academic Profiler (SAP): a framework for researcher assessment based on semantic topic modeling. Scientometrics, 127(8):5005–5026, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04449-9.

Vasiliadis:2023:IBD

[VPSF23] Giorgos Vasiliadis, Costas Panagiotakis, Iliana Stenaki, and John Fanourgiakis. The impact of brain-drain in country ranking: the case of computer science. Scientometrics, 128(2):1441–1450, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04613-1.

vanRaan:2021:LOP

[vR21a] Anthony F. J. van Raan. Laudation on the occasion of the presentation of the Derek de Solla Price Award 2021 to Prof. Ludo Waltman at the ISSI conference, Leuven, 2021. Scientometrics, 126(10):8235–8238, October 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04127-2.

vanRaan:2021:SBG

[vR21b] Anthony F. J. van Raan. Sleeping beauties gain impact in overdrive mode. Scientometrics, 126(5):4311–4332, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03910-5.

Vilchez-Roman:2021:UCC

[VRVH21] Carlos Vílchez-Román and Arístides Vara-Horna. Usage, content and citation in open access publication: any interaction effects? Scientometrics, 126(12):9457–9476, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04178-5.

vanRaan:2025:USP

[vRW25] Anthony F. J. van Raan and Jos J. Winnink. Urban scaling of patents and relation with socioeconomic strength for German cities and their urban areas. Scientometrics, 130(8):4309–4337, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05388-x.

vanSchalkwyk:2020:CSI

[vSDC20] François van Schalkwyk, Jonathan Dudek, and Rodrigo Costas. Communities of shared interests and cognitive bridges: the case of the anti-vaccination movement on Twitter. Scientometrics, 125(2):1499–1516, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03551-0.

Vanderstraeten:2021:IGW

[VV21] Raf Vanderstraeten and Frédéric Vandermoere. Inequalities in the growth of Web of Science. Scientometrics, 126(10):8635–8651, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04143-2.

Velicu:2025:QRP

[VWR25] Antonia Velicu, Fabian Winter, and Heiko Rauhut. Are questionable research practices considered a successful career strategy? A novel implementation of the implicit association test. Scientometrics, 130(7):3349–3366, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05357-4.

Veretennik:2023:IQS

[VY23] Elena Veretennik and Maria Yudkevich. Inconsistent quality signals: evidence from the regional journals. Scientometrics, 128(6):3675–3701, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04723-4.

Wada:2020:WDU

[Wad20] Tetsuo Wada. When do the USPTO examiners cite as the EPO examiners? An analysis of examination spillovers through rejection citations at the international family-to-family level. Scientometrics, 125(2):1591–1615, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03674-4;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03674-4.pdf.

Wada:2024:EEP

[Wad24] Tetsuo Wada. Experience effects of patent examiners: an empirical study of the career length and citation patterns on triadic patents. Scientometrics, 129(10):6333–6348, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05037-9.

Wagner:2025:ISC

[Wag25] Caroline S. Wagner. International science collaboration as a complex adaptive system in the work of Loet Leydesdorff. Scientometrics, 130(6):3155–3181, June 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05309-y.

Walters:2022:CDP

[Wal22] William H. Walters. Can differences in publisher size account for the relatively low prices of the journals available to master’s universities through commercial publishers’ databases? The importance of price discrimination and substitution effects. Scientometrics, 127(2):1065– 1097, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04205-5.

Wu:2025:NIE

[WAM^{+}25] Dengsheng Wu, Wenting Ao, Yuguo Min, Siting Li, and Jing Li. Normalising the H-index: the equal quantity H-index for journals. Scientometrics, 130(3):1443–1468, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05252-y.

Wang:2024:QCH

[Wan24a] Jiawei Wang. Quantifying cohesion in high citation research article titles: a cross-disciplinary and diachronic investigation. Scientometrics, 129(9):5075–5102, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05123-y.

Wang:2024:CCI

[Wan24b] Yongzhen Wang. Comparison of citation impact between pre-and post-publication peer-selected best papers: the case of Best Paper Awards recipients at computer science conferences. Scientometrics, 129(1):641–662, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04881-5.

Wang:2025:ETL

[Wan25] Juite Wang. Exploring technological landscape to uncover technological opportunities for immersive technologies in the Metaverse using patent data. Scientometrics, 130(7): 3483–3503, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05342-x.

Waqas:2023:HSE

[WAT23] Muhammad Waqas, Nadeem Anjum, and Afzal Muhammad Tanvir. A hybrid strategy to extract metadata from scholarly articles by utilizing support vector machine and heuristics. Scientometrics, 128(8):4349–4382, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04774-7.

Wohlrabe:2021:WBP

[WB21] Klaus Wohlrabe and Constantin Bürgi. What is the benefit from publishing a working paper in a journal in terms of citations? Evidence from economics. Scientometrics, 126 (6):4701–4714, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03942-x.

Wohlrabe:2022:ACA

[WB22] Klaus Wohlrabe and Lutz Bornmann. Alphabetized coauthorship in economics reconsidered. Scientometrics, 127 (5):2173–2193, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04322-9.

Wei:2020:ERB

[WC20] Mingkun Wei and Abdolreza Noroozi Chakoli. Evaluating the relationship between the academic and social impact of open access books based on citation behaviors and social media attention. Scientometrics, 125(3):2401– 2420, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03678-0.

Wang:2024:IIT

[WCCC24] Zhongyi Wang, Jing Chen, Jiangping Chen, and Haihua Chen. Identifying interdisciplinary topics and their evolution based on BERTopic. Scientometrics, 129(11):7359– 7384, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04776-5.

Woolcott:2020:CPE

[WCH^{+}20] Geoff Woolcott, Dan Chamberlain, Zachary Hawes, Michelle Drefs, Catherine D. Bruce, Brent Davis, Krista Francis, David Hallowell, Lynn McGarvey, Joan Moss, Joanne Mulligan, Yukari Okamoto, Nathalie Sinclair, and Walter Whiteley. The central position of education in knowledge mobilization: insights from network analyses of spatial reasoning research across disciplines. Scientometrics, 125(3):2323–2347, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03692-2.

Wang:2025:SIT

[WCL25] Haining Wang, Jason Clark, and Xiaozhong Liu. Science out of its Ivory Tower: improving accessibility with reinforcement learning. Scientometrics, 130(8):4495–4517, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05386-z.

Wagner:2020:CSS

[WCM20] Caroline S. Wagner, Xiaojing Cai, and Satyam Mukherjee. China’s scholarship shows atypical referencing patterns. Scientometrics, 124(3):2457–2468, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03579-2.

Wang:2022:ESS

[WCY22] Hei-Chia Wang, Jen-Wei Cheng, and Che-Tsung Yang. SentCite: a sentence-level citation recommender based on the salient similarity among multiple segments. Scientometrics, 127(5):2521–2546, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04339-0.

Weber:2021:DDA

[WD21] Rosina O. Weber and Kedma B. Duarte. Data-driven artificial intelligence to automate researcher assessment. Scientometrics, 126(4):3265–3281, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03859-x.

Wei:2020:RIE

[Wei20] Mingkun Wei. Research on impact evaluation of open access journals. Scientometrics, 122(2):1027–1049, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03306-6.

Wei:2024:ENE

[WFSZ24] Tingting Wei, Danyu Feng, Shiling Song, and Cai Zhang. An extraction and novelty evaluation framework for technology knowledge elements of patents. Scientometrics, 129(11):7417–7442, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04990-9.

Wang:2024:VTU

[WFW24] Yuefen Wang, Lipeng Fan, and Lei Wu. A validation test of the Uzzi et al. novelty measure of innovation and applications to collaboration patterns between institutions. Scientometrics, 129(7):4379–4394, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05071-7.

Wieland:2020:RBB

[WG20a] Martin Wieland and Juan Gorraiz. The rivalry between Bernini and Borromini from a scientometric perspective. Scientometrics, 125(2):1643–1663, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03514-5;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03514-5.pdf.

Wohlrabe:2020:UAA

[WG20b] Klaus Wohlrabe and Sabine Gralka. Using archetypoid analysis to classify institutions and faculties of economics. Scientometrics, 123(1):159–179, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03366-z;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03366-z.pdf.

Wren:2022:DAR

[WG22] Jonathan D. Wren and Constantin Georgescu. Detecting anomalous referencing patterns in PubMed papers suggestive of author-centric reference list manipulation. Scientometrics, 127(10):5753–5771, October 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04503-6.

Wei:2024:EUC

[WG24] Jiaqi Wei and Ying Guo. The effect of urban capacity in knowledge recombination on digital economy development. Scientometrics, 129(9):5373–5402, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05113-0.

Wendt:2022:IMR

[WGA22] Kaja Wendt, Hebe Gunnes, and Dag W. Aksnes. International migration of researchers and gender imbalance in academia — the case of Norway. Scientometrics, 127(12):7575–7591, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04365-y.

Wang:2020:IPL

[WGC20] Zhiqi Wang, Wolfgang Glänzel, and Yue Chen. The impact of preprints in Library and Information Science: an analysis of citations, usage and social attention indicators. Scientometrics, 125(2):1403–1423, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03612-4.

Wang:2023:PPM

[WGLN23] Jingbei Wang, Min Guo, Hui Liu, and Yafei Nie. Partners’ partners matter: the effect of partners’ centrality diversity on the focal organization’s innovation outputs. Scientometrics, 128(3):1547–1565, March 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04637-1.

Wang:2021:IAA

[WGW21] Chuanyi Wang, Fei Guo, and Qing Wu. The influence of academic advisors on academic network of physics doctoral students: empirical evidence based on scientometrics analysis. Scientometrics, 126(6):4899–4925, June 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03974-3.

Wu:2022:SDA

[WGZ22] Yue Wu, Xin Gu, and Zhaobohan Zhang. System dynamic analysis on industry-university-research institute synergetic innovation process based on knowledge flow. Scientometrics, 127(3):1317–1338, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04244-y.

Wang:2023:EDE

[WH23] Juite Wang and Tzu-Yen Hsu. Early discovery of emerging multi-technology convergence for analyzing technology opportunities from patent data: the case of smart health. Scientometrics, 128(8):4167–4196, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04760-z.

Wang:2024:ECR

[WH24] Qian Wang and Guangwei Hu. Expressions of confusion in research articles: a diachronic cross-disciplinary investigation. Scientometrics, 129(1):445–471, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04914-z.

Wang:2025:PCT

[WH25] Yunyun Wang and Guangwei Hu. Positive covariation or trade-off? A cross-disciplinary investigation of shell nouns and their congruent expressions in research articles. Scientometrics, 130(7):3597–3627, July 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05373-4.

Wieczorek:2024:MOS

[WHÅ24] Oliver Wieczorek, Olof Hallonsten, and Fredrik Åström. Is Management and Organizational Studies divided into (micro-)tribes? Scientometrics, 129(7):3871–3995, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05013-3.

Wijewickrema:2021:APA

[Wij21] Manjula Wijewickrema. Authors’ perception on abstracting and indexing databases in different subject domains. Scientometrics, 126(4):3063–3089, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-021-03896-0.

Wang:2022:IBU

[WJ22] Qi Wang and Tobias Jeppsson. Identifying benchmark units for research management and evaluation. Scientometrics, 127(12):7557–7574, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04413-7.

Wang:2024:EIC

[WJZ24] Congying Wang, Brent Jesiek, and Wei Zhang. Elevating international collaboration and academic outcomes through strategic research funding: a bibliometric analysis of China Scholarship Council funded publications. Scientometrics, 129(7):4329–4351, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05054-8.

Watson:2022:RBC

[WKC22] Don Watson, Manfred Krug, and Claus-Christian Carbon. The relationship between citations and the linguistic traits of specific academic discourse communities identified by using social network analysis. Scientometrics, 127(4): 1755–1781, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04287-9.

Whittington:2024:SSS

[WKC24] Kjersten Bunker Whittington, Molly M. King, and Isabella Cingolani. Structure, status, and span: gender differences in co-authorship networks across 16 region-subject pairs (2009–2013). Scientometrics, 129(1):147–179, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04885-1.

Wang:2021:ERB

[WL21a] Hui Wang and ZiChun Le. Expert recommendations based on link prediction during the COVID-19 outbreak. Scientometrics, 126(6):4639–4658, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03893-3.

Wang:2021:CLC

[WL21b] Wenjing Wang and Yiwei Liu. Community-level characteristics and member firms’ invention: evidence from university-industry innovation community in China. Scientometrics, 126(11):8913–8934, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04157-w.

Wen:2022:AAL

[WL22] Ju Wen and Lei Lei. Adjectives and adverbs in life sciences across 50 years: implications for emotions and readability in academic texts. Scientometrics, 127(8):4731–4749, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04453-z.

Wang:2024:HDO

[WL24] Yuzhuo Wang and Kai Li. How do official software citation formats evolve over time? A longitudinal analysis of R programming language packages. Scientometrics, 129(7): 3997–4019, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05064-6.

Wendzel:2020:AAE

[WLBC20] Steffen Wendzel, Cédric Lévy-Bencheton, and Luca Caviglione. Not all areas are equal: analysis of citations in information security research. Scientometrics, 122(1):

267–286, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03279-6.

Wang:2024:ORS

[WLHY24] Tiange Wang, Zixuan Li, Shan Huang, and Bo Yang. Is ORCID a reliable source for CV analysis? Exploring the data availability of ORCID academic profiles. Scientometrics, 129(3):1637–1662, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04944-1.

Wang:2025:CAE

[WLL25] Qianqian Wang, Hao Li, and Zhenhua Li. A contextaware enhanced local citation recommendation model integrating SciBERT and self-adaptive attention. Scientometrics, 130(8):4465–4493, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05382-3.

Wu:2020:DID

[WLLL20] Dengsheng Wu, Xiaoli Lu, Jianping Li, and Jing Li. Does the institutional diversity of editorial boards increase journal quality? The case economics field. Scientometrics, 124(2):1579–1597, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03505-6.

Wang:2025:DRP

[WLLW25a] Shuwen Wang, Minglu Li, Jianping Li, and Dengsheng Wu. Do researchers from prestigious universities deserve advantages in research funding? Evidence from the National Natural Science Foundation of China. Scientometrics, 130(3):1587–1615, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05278-2.

Wang:2025:MKF

[WLLW25b] Xi Wang, Dongqiao Li, Xiwen Liu, and Zhiqiang Wang. Measuring knowledge flow in the interdisciplinary field of biosecurity: full counting method or fractional counting method? Scientometrics, 130(2):1101–1128, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05249-7.

Wu:2024:ICA

[WLMZ24] Jinhong Wu, Tianye Liu, Keliang Mu, and Lei Zhou. Identification and causal analysis of predatory open access journals based on interpretable machine learning. Scientometrics, 129(4):2131–2158, April 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-04969-6.

Wei:2022:CPT

[WLS22] Wenjie Wei, Hongxu Liu, and Zhuanlan Sun. Cover papers of top journals are reliable source for emerging topics detection: a machine learning based prediction framework. Scientometrics, 127(8):4315–4333, August 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04462-y.

Wang:2022:RDL

[WLZ22] Shan Wang, Xiaojun Liu, and Jie Zhou. Readability is decreasing in language and linguistics. Scientometrics, 127(8):4697–4729, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04427-1.

Wang:2023:ESH

[WLZ^{+}23] Yue Wang, Ning Li, Bin Zhang, Qian Huang, Jian Wu, and Yang Wang. The effect of structural holes on producing novel and disruptive research in physics. Scientometrics, 128(3):1801–1823, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04635-3.

Wang:2022:IKC

[WMCL22] Shiyun Wang, Jin Mao, Yujie Cao, and Gang Li. Integrated knowledge content in an interdisciplinary field: identification, classification, and application. Scientometrics, 127(11):6581–6614, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04282-0.

Wijnen:2021:GBA

[WMK21] Marjolijn N. Wijnen, Jorg J. M. Massen, and Mariska E. Kret. Gender bias in the allocation of student grants. Scientometrics, 126(7):5477–5488, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03985-0.

Wouters:2025:CC

[Wou25] Paul Wouters. Complexity to the cube. Scientometrics, 130(6):3229–3255, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-025-
05326-x.

Wiechetek:2022:ASN

[WP22] Lukasz Wiechetek and Zbigniew Pastuszak. Academic social networks metrics: an effective indicator for university performance? Scientometrics, 127(3):1381–1401, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04258-6.

Wray:2024:URP

[WPBH24] K. Brad Wray, Søren R. Paludan, Lutz Bornmann, and Robin Haunschild. Using Reference Publication Year Spectroscopy (RPYS) to analyze the research and publication culture in immunology. Scientometrics, 129(6):3271–3283, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05001-7.

Wheeler:2022:IFA

[WPS22] Jonathan Wheeler, Ngoc-Minh Pham, and Justin D. Shanks. Impact factions: assessing the citation impact of different types of open access repositories. Scientometrics, 127(8):4977–5003, August 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04467-7.

Wang:2025:TBS

[WPSK25] Yuhang Wang, Lei Pei, Jianjun Sun, and Lele Kang. Trace on both sides: a two-step text mining method to identify academic inventors’ patent–paper pairs. Scientometrics, 130(2):833–860, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05207-9.

Waqas:2021:MHB

[WQ21] Humaira Waqas and Muhammad Abdul Qadir. Multilayer heuristics based clustering framework (MHCF) for author name disambiguation. Scientometrics, 126(9):7637– 7678, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04087-7.

Waqas:2022:CFA

[WQ22] Humaira Waqas and Abdul Qadir. Completing features for author name disambiguation (AND): an empirical analysis. Scientometrics, 127(2):1039–1063, February 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04229-x.

Wang:2021:CYS

[WR21] Zhiqi Wang and Ronald Rousseau. COVID-19, the Yule– Simpson paradox and research evaluation. Scientometrics, 126(4):3501–3511, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03830-w;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03830-w.pdf.

Wray:2020:HRS

[Wra20a] K. Brad Wray. How is a revolutionary scientific paper cited?: the case of Hess’ “History of Ocean Basins”. Scientometrics, 124(2):1677–1683, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03524-3.

Wray:2020:PSF

[Wra20b] K. Brad Wray. Paradigms in structure: finally, a count. Scientometrics, 125(1):823–828, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03669-1.

Widziewicz-Rzonca:2020:FSR

[WRT20] Kamila Widziewicz-Rzońca and Malwina Tytła. First systematic review on PM-bound water: exploring the existing knowledge domain using the CiteSpace software. Scientometrics, 124(3):1945–2008, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03547-w;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03547-w.pdf.

Wang:2024:EII

[WSC^{+}24] Jiajie Wang, Jing Shi, Yitong Chen, Lele Kang, and Jianjun Sun. Exploring the impact of inter-organizational knowledge potential difference: an empirical investigation of inventor mobility. Scientometrics, 129(10):5979–6006, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05021-3.

Wong:2023:AQS

[WSL23] Chan-Yuan Wong, Jeffrey Sheu, and Keun Lee. Assessing the quest of SMEs in pivoting for new technological ventures: comparing the patenting indexes of seven developed cities. Scientometrics, 128(7):4029–4064, July 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04729-y.

Williams:2023:LEL

[WSRM23] Jennifer S. Williams, Jenna C. Stone, Stacey A. Ritz, and Maureen J. MacDonald. Letter to the editor: Laxdal (2023) “The sex gap in sports and exercise medicine research: who does research on females?”.

Scientometrics, 128(7):4155–4160, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04741-2. See [Lax23b] and response

[Lax23a].

Wang:2021:ISP

[WSY21] Jean J. Wang, Sarah X. Shao, and Fred Y. Ye. Identifying ‘seed’ papers in sciences. Scientometrics, 126(7): 6001–6011, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03980-5.

Wang:2021:PAA

[WSZ21] Kehan Wang, Wenxuan Shi, and Liying Zhang. Prediction and application of article potential citations based on nonlinear citation-forecasting combined model. Scientometrics, 126(8):6533–6550, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04026-6.

Wieczorek:2021:MFP

[WUH21] Oliver Wieczorek, Saïd Unger, and Raphael Heiberger. Mapping the field of psychology: Trends in research topics 1995–2015. Scientometrics, 126(12):9699–9731, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04069-9.

Wheeler:2021:MCL

[WVH21] Melissa A. Wheeler, Ekaterina Vylomova, and Nick Haslam. More confident, less formal: stylistic changes in academic psychology writing from 1970 to 2016. Scientometrics, 126(12):9603–9612, December 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04166-9.

Wagner:2025:ETL

[WW25] Caroline Wagner and Paul Wouters. Editorial: A tribute to Loet Leydesdorff by his coauthors and friends. Scientometrics, 130(6):3257–3266, June 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05316-z.

Wang:2022:MIM

[WWC22] Zhongyi Wang, Keying Wang, and Haihua Chen. Measuring the innovation of method knowledge elements in scientific literature. Scientometrics, 127(5):2803–2827, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04350-5.

Wang:2021:EEC

[WWH21] Xiaoguang Wang, Hongyu Wang, and Han Huang. Evolutionary exploration and comparative analysis of the research topic networks in information disciplines. Scientometrics, 126(6):4991–5017, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03963-6.

Wang:2024:EIW

[WWL24a] Yining Wang, Qiang Wu, and Liangyu Li. Examining the influence of women scientists on scientific impact and novelty: insights from top business journals. Scientometrics, 129(6):3517–3542, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05014-2.

Wu:2024:CAP

[WWL24b] Dengsheng Wu, Huidong Wu, and Jianping Li. Citation advantage of positive words: predictability, temporal evolution, and universality in varied quality journals. Scientometrics, 129(7):4275–4293, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05074-4.

Wolfram:2020:OPR

[WWP20] Dietmar Wolfram, Peiling Wang, and Hyoungjoo Park. Open peer review: promoting transparency in open science. Scientometrics, 125(2):1033–1051, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-020-03488-4.

Wang:2023:ANA

[WWS23a] Bin Wang, Feng Wu, and Lukui Shi. AGSTA-NET: adaptive graph spatiotemporal attention network for citation count prediction. Scientometrics, 128(1):511–541, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04541-0.

Wang:2023:LCS

[WWS^{+}23b] Gui Wang, Hui Wang, Xinyi Sun, Nan Wang, and Li Wang. Linguistic complexity in scientific writing: a large-scale diachronic study from 1821 to 1920. Scientometrics, 128(1): 441–460, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04550-z.

Wang:2020:RCM

[WWY^{+}20] Weibin Wang, Zheng Wang, Tian Yu, CholMyong Pak, and Guang Yu. Research on citation mention times and contributions using a neural network. Scientometrics, 125(3):2383–2400, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03711-2.

Wang:2020:FRA

[WWZW20] Peiling Wang, Joshua Williams, Nan Zhang, and Qiang Wu. F1000Prime recommended articles and their citations: an exploratory study of four journals. Scientometrics, 122(2):933–955, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03302-w;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03302-w.pdf.

Wu:2024:DSD

[WXD24] Keye Wu, Ziyue Xie, and Jia Tina Du. Does science disrupt technology? Examining science intensity, novelty, and recency through patent-paper citations in the pharmaceutical field. Scientometrics, 129(9):5469–5491, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05126-9.

Wang:2022:DCP

[WXZ22] Yezhu Wang, Yundong Xie, and Rongting Zhou. Do cover papers get better citations and usage counts? An analysis of 42 journals in cell biology. Scientometrics, 127(7): 3793–3813, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04444-0.

Wu:2024:CSR

[WXZ24] Wenqing Wu, Haixu Xi, and Chengzhi Zhang. Are the confidence scores of reviewers consistent with the review content? Evidence from top conference proceedings in AI. Scientometrics, 129(7):4109–4135, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05070-8.

Wen:2023:CLS

[WY23] Ju Wen and Lan Yi. Comparing lay summaries to scientific abstracts for readability and jargon use: a case report. Scientometrics, 128(10):5791–5800, October 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04807-1.

Wu:2021:IAD

[WYL21] Hong Wu, Huifang Yi, and Chang Li. An integrated approach for detecting and quantifying the topic evolutions of patent technology: a case study on graphene field. Scientometrics, 126(8):6301–6321, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04000-2.

Wang:2021:DLA

[WYT21] Xue Wang, Xuemei Yang, and Xiaoli Tang. A deep learning approach for identifying biomedical breakthrough discoveries using context analysis. Scientometrics, 126(7):

5531–5549, July 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04003-z.

Wang:2021:SBP

[WZ21] Yajie Wang and Alesia Zuccala. Scholarly book publishers as publicity agents for SSH titles on Twitter. Scientometrics, 126(6):4817–4840, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03947-6.

Wu:2025:WDB

[WZ25] Bingyi Wu and Wenhao Zhou. Where do breakthroughs originate? Utilizing patent knowledge network to identify breakthrough technological innovations. Scientometrics, 130(5):2523–2550, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-025-
05276-4.

Weinberger:2021:DSM

[WZG21] Maor Weinberger and Maayan Zhitomirsky-Geffet. Diversity of success: measuring the scholarly performance diversity of tenured professors in the Israeli academia. Scientometrics, 126(4):2931–2970, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03823-9.

Wang:2025:WTA

[WZHW25] Yuanyuan Wang, Yang Zhang, Jianhua Hou, and Dongyi Wang. Who tweets about quantum physics research on Twitter: the impact of user types, tweet content and interaction patterns. Scientometrics, 130(3):1871–1899, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05274-6.

Wang:2020:ICI

[WZJ^{+}20] Mingyang Wang, Jiaqi Zhang, Shijia Jiao, Xiangrong Zhang, Na Zhu, and Guangsheng Chen. Important citation identification by exploiting the syntactic and contextual information of citations. Scientometrics, 125(3):2109– 2129, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03677-1.

Wang:2024:SLP

[WZJW24] Qi Wang, Bentao Zou, Jialin Jin, and Yuefen Wang. Studying the linkage patterns and incremental evolution of domain knowledge structure: a perspective of structure deconstruction. Scientometrics, 129(7):4249–4274, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05059-3.

Wagner:2022:DMT

[WZL22a] Caroline S. Wagner, Lin Zhang, and Loet Leydesdorff. A discussion of measuring the top-1% most-highly cited publications: quality and impact of Chinese papers. Scientometrics, 127(4):1825–1839, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04291-z.

Wang:2022:IDV

[WZL22b] Xuefeng Wang, Shuo Zhang, and Yuqin Liu. ITGInsight — discovering and visualizing research fronts in the scientific literature. Scientometrics, 127(11):6509–6531, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04190-9.

Wang:2022:RME

[WZL22c] Yuzhuo Wang, Chengzhi Zhang, and Kai Li. A review on method entities in the academic literature: extraction, evaluation, and application. Scientometrics, 127(5): 2479–2520, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04332-7.

Wu:2024:CKD

[WZM^{+}24] Mengjia Wu, Yi Zhang, Mark Markley, Caitlin Cassidy, Nils Newman, and Alan Porter. COVID-19 knowledge deconstruction and retrieval: an intelligent bibliometric solution. Scientometrics, 129(11):7229–7259, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04747-w.

Wei:2023:WDO

[WZNL23] Chunli Wei, Jingyi Zhao, Jue Ni, and Jiang Li. What does open peer review bring to scientific articles? Evidence from PLoS journals. Scientometrics, 128(5):2763–2776, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04683-9.

Wang:2021:RPD

[WZY21] Xuefeng Wang, Shuo Zhang, and Xuemei Yang. Revealing potential drug–disease–gene association patterns for precision medicine. Scientometrics, 126(5):3723–3748, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-03892-4.

Wang:2021:PAP

[WZZ^{+}21] Wenyan Wang, Jun Zhang, Fang Zhou, Peng Chen, and Bing Wang. Paper acceptance prediction at the institutional level based on the combination of individual and network features. Scientometrics, 126(2):1581–1597, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03813-x.

Wang:2023:ESC

[WZZ23] Ruijie Wang, Yuhao Zhou, and An Zeng. Evaluating scientists by citation and disruption of their representative works. Scientometrics, 128(3):1689–1710, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04631-7.

Wang:2024:CPH

[WZZ24] Shanshan Wang, Hongyan Zhou, and Tianyi Zhao. Configuration paths to high-value patents: evidence from patents winning the China Patent Awards. Scientometrics, 129 (5):2633–2658, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04976-7.

Wang:2025:LPA

[WZZ25] Xiaohong Wang, Jiyang Zhao, and Ben Zhang. The “leaky pipeline” in the academic growth: evidence from excellent young scientists of the NSFC. Scientometrics, 130(2): 1129–1158, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05229-3.

Wu:2020:EAH

[WZZZ20] Jie Wu, Ganggang Zhang, Qingyuan Zhu, and Zhixiang Zhou. An efficiency analysis of higher education institutions in China from a regional perspective considering the external environmental impact. Scientometrics, 122 (1):57–70, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03296-5.

Xu:2021:DSP

[XAA21] Shuo Xu, Mengjia An, and Xin An. Do scientific publications by editorial board members have shorter publication delays and then higher influence? Scientometrics, 126(8): 6697–6713, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04067-x.

Xu:2025:WPM

[XCML25] Yuan Xu, Xi Chen, Jin Mao, and Gang Li. Will patents with more interdisciplinary scientific knowledge have higher technological impact? Empirical evidence from USPTO patents. Scientometrics, 130(4):2037–2068, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05280-8.

Xia:2020:BRI

[XCT20] Qinghua Xia, Qinwei Cao, and Manqing Tan. Basic research intensity and diversified performance: the moderating role of government support intensity. Scientometrics, 125(1):577–605, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03635-x.

Xu:2022:DNC

[XDL22] Linhong Xu, Kun Ding, and Yuan Lin. Do negative citations reduce the impact of cited papers? Scientometrics, 127(2):1161–1186, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04214-4.

Xu:2023:DCP

[XDLZ23] Linhong Xu, Kun Ding, Yuan Lin, and Chunbo Zhang. Does citation polarity help evaluate the quality of academic papers? Scientometrics, 128(7):4065–4087, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04734-1.

Xu:2022:HFK

[XDW22] Lei Xu, Ronggui Ding, and Lei Wang. How to facilitate knowledge diffusion in collaborative innovation projects by adjusting network density and project roles. Scientometrics, 127(3):1353–1379, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04255-9.

Xu:2020:RER

[XHA^{+}20] Shuo Xu, Liyuan Hao, Xin An, Hongshen Pang, and Ting Li. Review on emerging research topics with key-route main path analysis. Scientometrics, 122(1):607–624, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03288-5.

Xu:2020:MIE

[XHQZ20] Guannan Xu, Weijie Hu, Yuanyuan Qiao, and Yuan Zhou. Mapping an innovation ecosystem using network clustering and community identification: a multi-layered framework. Scientometrics, 124(3):2057–2081, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03543-0;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03543-0.pdf.

Xie:2020:MAM

[Xie20] Shaoliang Xie. Multidimensional analysis of Master thesis abstracts: a diachronic perspective. Scientometrics, 123 (2):861–881, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03408-6.

Xie:2021:DHM

[Xie21] Zheng Xie. A distributed hypergraph model for simulating the evolution of large coauthorship networks. Scientometrics, 126(6):4609–4638, June 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03991-2.

Xie:2024:RSC

[XJZW24] Weixi Xie, Pengfei Jia, Guangyao Zhang, and Xianwen Wang. Are reviewer scores consistent with citations? Scientometrics, 129(8):4721–4740, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05103-2.

Xiang:2020:DCS

[XL20] Xuechun Xiang and Jing Li. A diachronic comparative study of research article titles in linguistics and literature journals. Scientometrics, 122(2):847–866, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03329-z.

Xu:2023:DAI

[XLA23] Shuo Xu, Ling Li, and Xin An. Do academic inventors have diverse interests? Scientometrics, 128(2):1023– 1053, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04587-0.

Xie:2020:APQ

[XLL20] Zheng Xie, Yanwu Li, and Zhemin Li. Assessing and predicting the quality of research master’s theses: an application of scientometrics. Scientometrics, 124(2):953–972, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03489-3.

Xia:2023:RSI

[XLL23] Wanjun Xia, Tianrui Li, and Chongshou Li. A review of scientific impact prediction: tasks, features and methods. Scientometrics, 128(1):543–585, January 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04547-8.

Xu:2025:RPN

[XLL25] Linming Xu, Baicun Li, and Meijuan Li. Research productivity and novelty under different funding models: evidence from NIH-funded research projects. Scientometrics, 130 (7):3725–3741, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05370-7.

Xu:2021:ADC

[XLY21] Shuo Xu, Ling Li, and Guancan Yang. An approach for detecting the commonality and specialty between scientific publications and patents. Scientometrics, 126(9):7445– 7475, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04085-9.

Xiong:2022:PFS

[XlYqZ22] Xi Xiong, Guo liang Yang, and De qun Zhou. A proposed fixed-sum carryovers reallocation DEA approach for social scientific resources of Chinese public universities. Scientometrics, 127(7):4097–4121, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04411-9.

Xu:2025:DPC

[XS25] Haotian Xu and Wenqin Shen. The double penalty of class and gender: the research productivity of married female doctoral students. Scientometrics, 130(7):4093–4115, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05375-2.

Xie:2025:CCB

[XW25] Qianqian Xie and Ludo Waltman. A comparison of citation-based clustering and topic modeling for science mapping. Scientometrics, 130(5):??, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05324-z.

Xi:2022:ACR

[XWGD22] Xiaowen Xi, Jiaqi Wei, Ying Guo, and Weiyu Duan. Academic collaborations: a recommender framework spanning research interests and network topology. Scientometrics, 127(11):6787–6808, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04555-8.

Xu:2025:MCC

[XXL25] Xiujuan Xu, Yueyue Xie, and Yu Liu. Mf-cite: citation intent classification in scientific papers based on multifeature fusion. Scientometrics, 130(8):4423–4464, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05365-4.

Xie:2025:QAP

[XYY25] Qianqian Xie and Alfredo Yegros-Yegros. A quantitative assessment of potential benefits and challenges of international researcher mobility for home and host countries: evidence from the Chinese Scholarship Council programmes. Scientometrics, 130(8):4519–4543, August 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05391-2.

Xie:2023:ECB

[XZ23] Qing Xie and Xinyuan Zhang. Exploring the correlation between acknowledgees’ contributions and their academic performance. Scientometrics, 128(11):6003–6027, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04828-w.

Xing:2019:SNE

[XZD19] Yanmeng Xing, An Zeng, and Zengru Di. The strong nonlinear effect in academic dropout. Scientometrics, 120(2): 793–805, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-019-03135-7.

Xiong:2022:CMS

[XZZ22a] Ting Xiong, Liang Zhou, and Xiaojuan Zhang. Correction to: Mining semantic information of co-word network to improve link prediction performance. Scientometrics, 127(6):3005, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04422-6.

See [XZZ22b].

Xiong:2022:MSI

[XZZ22b] Ting Xiong, Liang Zhou, and Xiaojuan Zhang. Mining semantic information of co-word network to improve link prediction performance. Scientometrics, 127(6):2981–3004, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04247-9. See correc-

tion [XZZ22a].

Ynalvez:2021:MPD

[YA21] Marcus Antonius Ynalvez and Jorge Luis Aviles. Mentoring practices, developmental networks, and doctoral science training experiences. Scientometrics, 126(8):6323– 6347, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04031-9.

Yamashita:2020:AIT

[Yam20] Yasuhiro Yamashita. An attempt to identify technologically relevant papers based on their references. Scientometrics, 125(2):1783–1800, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03673-5;   http://link.springer.
com/content/pdf/10.1007/s11192-020-03673-5.pdf.

Yao:2025:ACN

[YCL25] Le-Ye Yao, Kai-Yi Chen, and Peng-Hui Lyu. Academic collaboration networks study on library and information science community. Scientometrics, 130(8):4339–4367, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05334-x.

Yu:2020:HAP

[YCXY20] Houqiang Yu, Xueting Cao, Tingting Xiao, and Zhenyi Yang. How accurate are policy document mentions? A first look at the role of altmetrics database. Scientometrics, 125(2):1517–1540, November 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03558-7.

Yang:2024:USK

[YCYZ24] Jinqing Yang, Xiufeng Cheng, Guanghui Ye, and Yuchen Zhang. Understanding scientific knowledge evolution patterns based on egocentric network perspective. Scientometrics, 129(11):6719–6750, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05156-3.

Yalcin:2021:MRI

[YD21] Haydar Yalcin and Tugrul Daim. Mining research and invention activity for innovation trends: case of blockchain technology. Scientometrics, 126(5):3775–3806, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-03876-4.

Yenilmez:2025:UCA

[Yen25] Taylan Yenilmez. Understanding complexity in the author–journal space. Scientometrics, 130(7):3383–3403, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05344-9.

Yan:2024:IIE

[YF24] Zhaoping Yan and Kaiyu Fan. An integrated indicator for evaluating scientific papers: considering academic impact and novelty. Scientometrics, 129(11):6909–6929, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05150-9.

Yair:2020:AMP

[YG20] Gad Yair and Keith Goldstein. The Annus Mirabilis paper: years of peak productivity in scientific careers. Scientometrics, 124(2):887–902, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03544-z.

Yair:2022:TCA

[YGO22] Gad Yair, Keith Goldstein, and Anthony J. Olejniczak. The three cultures in American science: publication productivity in physics, history and economics. Scientometrics, 127(6):2967–2980, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04396-5.

Yang:2024:RDI

[YGW^{+}24] Alex J. Yang, Hongcun Gong, Yuhao Wang, Chao Zhang, and Sanhong Deng. Rescaling the disruption index reveals the universality of disruption distributions in science. Scientometrics, 129(1):561–580, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04889-x.

Yang:2024:TOP

[YH24] Chao Yang and Cui Huang. Target-oriented policy diffusion analysis: a case study of China’s information technology policy. Scientometrics, 129(3):1347–1376, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04895-z.

Ylikruuvi:2025:CRD

[YJJ25] Kaisa Ylikruuvi, Kalervo Järvelin, and Martti Juhola. Comparing representations of a discipline derived through LDA vs. intellectual content analysis: the case of information science. Scientometrics, 130(8):4283–4307, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05376-1.

Yoon:2021:MIT

[YJS21] So Yoon Yoon, Su Jung Jee, and So Young Sohn. Mapping and identifying technological coopetition: a multi-level approach. Scientometrics, 126(7):5797–5817, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04005-x.

Yang:2023:KIR

[YK23] Seolmin Yang and So Young Kim. Knowledge-integrated research is more disruptive when supported by homogeneous funding sources: a case of US federally funded research in biomedical and life sciences. Scientometrics, 128 (6):3257–3282, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04706-5.

Yoon:2022:DBL

[YKS22] Byungun Yoon, Songhee Kim, and Hyeonju Seol. Doc2vecbased link prediction approach using SAO structures: application to patent network. Scientometrics, 127(9):5385– 5414, September 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04187-4.

Yu:2020:EGG

[YKTM20] Mengyu Yu, Mazie Krehbiel, Samantha Thompson, and Tatjana Miljkovic. An exploration of gender gap using advanced data science tools: actuarial research community. Scientometrics, 123(2):767–789, May 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03412-w.

Yanhui:2024:IBP

[YL24] Song Yanhui and Lei Lixin. Inventor bibliographic-patent-coupling analysis and inventor-patent-classification-coupling analysis: a comparative analysis based on NPE. Scientometrics, 129(2):745–765, February 2024. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-023-04713-6.

Yang:2020:HRA

[YLC^{+}20] Chen Yang, Tingting Liu, Xiaohong Chen, Yiyang Bian, and Yuewen Liu. HNRWalker: recommending academic collaborators with dynamic transition probabilities in heterogeneous networks. Scientometrics, 123(1):429–449, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03374-z.

Yan:2020:SNF

[YLCY20] Weiwei Yan, Qian Liu, Ruoyu Chen, and Shengwei Yi. Social networks formed by follower–followee relationships on academic social networking sites: an examination of corporation users. Scientometrics, 124(3):2083–2101, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03553-y.

Yang:2024:UKA

[YLCY24] Jinqing Yang, Zhifeng Liu, Xiufeng Cheng, and Guanghui Ye. Understanding the keyword adoption behavior patterns of researchers from a functional structure perspective. Scientometrics, 129(6):3359–3384, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05031-1.

Yang:2025:HDS

[YLf25] Siluo Yang, Longfei Li, and Qian feng. How does social media mention academic papers? Evidence from WeChat in China. Scientometrics, 130(8):4591–4619, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05399-8.

Yang:2024:IFS

[YLH24] Jinqing Yang, Zhifeng Liu, and Yong Huang. From informal to formal: scientific knowledge role transition prediction. Scientometrics, 129(8):4909–4935, August 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05093-1.

Yanhui:2021:CSF

[YLJ21] Song Yanhui, Wu Lijuan, and Qiu Junping. A comparative study of first and all-author bibliographic coupling analysis based on Scientometrics. Scientometrics, 126(2): 1125–1147, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03798-7.

Yang:2024:DII

[YLM24a] Xi Yang, Xinyue Lin, and Xin Ma. Does interdisciplinarity improve research performance? An empirical study of faculty members at Chinese universities. Scientometrics, 129(12):7645–7666, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05182-1.

Yu:2024:IHD

[YLM^{+}24b] Qian Yu, Xiuwen Li, Ding Ma, Li Zhang, Kexin Chen, Qiang Xue, and Qingjie Zhang. Interdisciplinary hierarchical diversity driving disruption. Scientometrics, 129(12): 7833–7849, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-024-
05184-z.

Yi:2020:RPI

[YLW20] Yingting Yi, Jiangshui Luo, and Michael Wübbenhorst. Research on political instability, uncertainty and risk during 1953–2019: a scientometric review. Scientometrics, 123 (2):1051–1076, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03416-6.

Yu:2025:USS

[YLX25] Houqiang Yu, Yian Liang, and Yinghua Xie. Understanding the sustainability of supply–demand in peer review system: an analysis based on scholars’ research and review activities. Scientometrics, 130(3):1547–1569, March 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05264-8.

Yin:2020:LSN

[YMD20] Deyun Yin, Kazuyuki Motohashi, and Jianwei Dang. Large-scale name disambiguation of Chinese patent inventors (1985–2016). Scientometrics, 122(2):765–790, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03310-w.

Yu:2023:HCP

[YMLL23] Houqiang Yu, Biegzat Murat, Jiatong Li, and Longfei Li. How can policy document mentions to scholarly papers be interpreted? An analysis of the underlying mentioning process. Scientometrics, 128(11):6247–6266, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04826-y.

Yu:2020:CVE

[YMW^{+}20] Zhenglu Yu, Zheng Ma, Haiyan Wang, Jia Jia, and Lu Wang. Communication value of English-language S&T academic journals in non-native English language countries. Scientometrics, 125(2):1389–1402, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03594-3.

Yu:2021:HAT

[YMX21] Houqiang Yu, Biegzat Murat, and Tingting Xiao. How accurate are Twitter and Facebook altmetrics data? A comparative content analysis. Scientometrics, 126(5): 4437–4463, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03954-7.

Yao:2023:PPG

[YN23] Li Yao and He Ni. Prediction of patent grant and interpreting the key determinants: an application of interpretable machine learning approach. Scientometrics, 128(9):4933– 4969, September 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04736-z.

Ye:2025:ESC

[YNL25] Yingxin Estella Ye, Jin-Cheon Na, and Meky Liu. Examining scholarly communication on X (Twitter): insights from participants tweeting COVID-19 and ChatGPT publications. Scientometrics, 130(2):1045–1076, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05246-w.

Ye:2022:AAD

[YNO22] Yingxin Estella Ye, Jin-Cheon Na, and Poong Oh. Are automated accounts driving scholarly communication on Twitter? A case study of dissemination of COVID-19 publications. Scientometrics, 127(5):2151–2172, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04343-4.

Yoon:2020:PTS

[YP20] Jungwon Yoon and Han Woo Park. Pattern and trend of scientific knowledge production in North Korea by a semantic network analysis of papers in journal titled technological innovation. Scientometrics, 124(2):1421–1438, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03497-3.

Yang:2024:DCS

[YP24] Yiying Yang and Fan Pan. Diachronic changes in syntactic complexity of science research articles: a comparative study of medicine and mechanical engineering. Scientometrics, 129(3):1663–1686, March 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04891-3.

Yan:2021:FLS

[YRP21] Zheng Yan, Wenqian Robertson, and Sung Yong Park. Finding leading scholars in mobile phone behavior: a mixed-method analysis of an emerging interdisciplinary field. Scientometrics, 126(12):9499–9517, December 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04184-7.

Yu:2020:KDP

[YS20] Dejian Yu and Libo Sheng. Knowledge diffusion paths of blockchain domain: the main path analysis. Scientometrics, 125(1):471–497, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03650-y.

Yu:2023:MEC

[YS23] Bo Yu and Fei Shu. The Matthew Effect in China’s social sciences and humanities research: a comparative analysis of CSSCI and SSCI. Scientometrics, 128(11):6177– 6193, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04818-y.

Yang:2024:EJR

[YS24] Yan-Hong Yang and Ying-Hui Shao. Education journal rankings: a diversity-based Author Affiliation Index assessment methodology. Scientometrics, 129(5):2677–2700, May 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04979-4.

Yuan:2020:SBA

[YSW^{+}20] Sha Yuan, Zhou Shao, Xingxing Wei, Jie Tang, Wendy Hall, Yongli Wang, Ying Wang, and Ye Wang. Science behind AI: the evolution of trend, mobility, and collaboration. Scientometrics, 124(2):993–1013, August 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03423-7.

Yang:2024:OAI

[YSWC24] Puyu Yang, Ahad Shoaib, Robert West, and Giovanni Colavizza. Open access improves the dissemination of science: insights from Wikipedia. Scientometrics, 129(11): 7083–7106, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-024-
05163-4.

Yi:2025:REC

[YT25] Tao Yi and Chaoying Tang. Research exploration, collaborative partnerships, and scientific breakthrough: evidence from China’s State Key Laboratories. Scientometrics, 130(8):4397–4422, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05389-w.

Yu:2023:EEI

[YTJ23] Qian Yu, Rui Tao, and Shan Jiang. Exploring the evolution of interdisciplinary citation network by the colored network motifs: the case of Perovskite Materials. Scientometrics, 128(8):4421–4446, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04777-4.

Yue:2021:CIM

[YTM21] Mingliang Yue, Hongbo Tang, and Tingcan Ma. Consistency index: measuring the performances of scholar journal reviewers. Scientometrics, 126(8):7183–7195, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-021-04013-x.

Yan:2020:IPN

[YTZ20] Yan Yan, Shanwu Tian, and Jingjing Zhang. The impact of a paper’s new combinations and new components on its citation. Scientometrics, 122(2):895–913, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03314-6.

Yuret:2020:CWN

[Yur20] Tolga Yuret. Co-worker network: How closely are researchers who published in the top five economics journals related? Scientometrics, 124(3):2301–2317, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03589-0.

Yuret:2022:NAE

[Yur22] Tolga Yuret. Network analysis of econometric society fellows. Scientometrics, 127(12):7615–7631, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-022-04271-3.

Yuret:2023:CPR

[Yur23] Tolga Yuret. The citation performance of the references in the standard graduate-level microeconomics textbook: Mas–Collel et al. (1995). Scientometrics, 128(3):1473– 1484, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04650-4.

Yuret:2024:CPI

[Yur24] Tolga Yuret. Career paths of the International Mathematics Olympiad (IMO) medalists. Scientometrics, 129(6): 3469–3491, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05042-y.

Yu:2021:AKE

[YWB21] Qi Yu, Qi Wang, and Yi Bu. Analyzing knowledge entities about COVID-19 using entitymetrics. Scientometrics, 126 (5):4491–4509, May 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03933-y.

Yu:2022:RIA

[YWB22] Qi Yu, Qi Wang, and Yi Bu. Reply to issues about entitymetrics and paper-entity citation network. Scientometrics, 127(4):2127–2129, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04311-y. See [SLY22].

Yu:2023:TBU

[YWHS23] Houqiang Yu, Yue Wang, Shah Hussain, and Haoyang Song. Towards a better understanding of Facebook Altmetrics in LIS field: assessing the characteristics of involved paper, user and post. Scientometrics, 128(5): 3147–3170, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04678-6.

Yu:2025:MPI

[YWL25] Qian Yu, Zhongjun Wei, and Nian Liu. Mapping the path to interdisciplinary innovation: a study of key roles in disciplinary convergence. Scientometrics, 130(2):909– 935, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05225-7.

Yang:2025:IRE

[YWN25] Chen Yang, Hailong Wang, and Ben Niu. Interdisciplinary research enhances scientists’ career resilience and long-term benefits: evidence from a large-scale bibliographic analysis. Scientometrics, 130(7):3903–3928, July 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05349-4.

Yang:2022:RNI

[YWW22] Jiale Yang, Qing Wu, and Chuanyi Wang. Research networks and the initial placement of PhD holders in academia: evidence from social science fields. Scientometrics, 127(6):3253–3278, June 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04394-7.

Yao:2023:PRR

[YWW23] Mingxin Yao, Ying Wei, and Huiyu Wang. Promoting research by reducing uncertainty in academic writing: a large-scale diachronic case study on hedging in Science research articles across 25 years. Scientometrics, 128(8): 4541–4558, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04759-6.

Yang:2023:SAI

[YWX23] Ruilu Yang, Qiang Wu, and Yundong Xie. Are scientific articles involving corporations associated with higher citations and views? An analysis of the top journals in business research. Scientometrics, 128(10):5659–5685, October 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04808-0.

Yin:2020:COB

[YWY^{+}20] Xicheng Yin, Hongwei Wang, Pei Yin, Hengmin Zhu, and Zhenyu Zhang. A co-occurrence based approach of automatic keyword expansion using mass diffusion. Scientometrics, 124(3):1885–1905, September 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03601-7.

Yan:2023:DQA

[YWZ^{+}23] Weiwei Yan, Xin Wen, Yin Zhang, Sonali Kudva, and Qian Liu. The dynamics of Q&A in academic social networking sites: insights from participants, interaction network, response time, and discipline differences. Scientometrics, 128(3):1895–1922, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04624-y.

Yu:2024:ARD

[YWZ^{+}24] Chao Yu, Chuhan Wang, Tongyang Zhang, Yi Bu, and Jian Xu. Analyzing research diversity of scholars based on multi-dimensional calculation of knowledge entities. Scientometrics, 129(11):7329–7358, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04821-3.

Ye:2021:DFC

[YX21] Qin Ye and Xiaolei Xu. Determining factors of cities’ centrality in the interregional innovation networks of China’s biomedical industry. Scientometrics, 126(4):2801–2819, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03853-3.

Yang:2024:UDT

[YXDL24] Alex J. Yang, Huimin Xu, Ying Ding, and Meijun Liu. Unveiling the dynamics of team age structure and its impact on scientific innovation. Scientometrics, 129(10): 6127–6148, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-04987-4.

Yang:2021:CAB

[YXQG21] Siluo Yang, Xin Xing, Fan Qi, and Maria Cláudia Cabrini Grácio. Comparison of academic book impact from a disciplinary perspective: an analysis of citations and altmetric indicators. Scientometrics, 126(2):1101–1123, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03808-8.

Yue:2022:MKD

[YXYQ22] Zenghui Yue, Haiyun Xu, Guoting Yuan, and Yan Qi. Modeling knowledge diffusion in the disciplinary citation network based on differential dynamics. Scientometrics, 127(12):7593–7613, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04491-7.

Yu:2021:KDS

[YY21] Dejian Yu and Zhaoping Yan. Knowledge diffusion of supply chain bullwhip effect: main path analysis and science mapping analysis. Scientometrics, 126(10):8491–8515, October 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04105-8.

Yamamoto:2022:PRS

[YY22a] Koh Yamamoto and Takuo Yasunaga. A percentile rank score of group productivity: an evaluation of publication productivity for researchers from various fields. Scientometrics, 127(4):1737–1754, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04278-w.

Yu:2022:CML

[YY22b] Dejian Yu and Zhaoping Yan. Combining machine learning and main path analysis to identify research front: from the perspective of science-technology linkage. Scientometrics, 127(7):4251–4274, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04443-1.

Yu:2022:HAN

[YYC22] Houqiang Yu, Xinyun Yu, and Xueting Cao. How accurate are news mentions of scholarly output? A content analysis. Scientometrics, 127(7):4075–4096, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04382-x.

Yiru:2025:CBA

[YYZ25] Liu Yiru, Liu Yi, and Yuan Zihan. A comprehensive bibliometric analysis of retracted chapters based on OpenAlex database. Scientometrics, 130(4):2425–2444, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05289-z.

Yang:2023:SBB

[YZH23] Ning Yang, Zhiqiang Zhang, and Feihu Huang. A study of BERT-based methods for formal citation identification of scientific data. Scientometrics, 128(11):5865– 5881, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04833-z.

Yang:2024:DLB

[YZWH24] Zhenyu Yang, Wenyu Zhang, Zhimin Wang, and Xiaoling Huang. A deep learning-based method for predicting the emerging degree of research topics using emerging index. Scientometrics, 129(7):4021–4042, July 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05068-2.

Yin:2024:PSP

[YZZT24] Dehu Yin, Xi Zhang, Hongke Zhao, and Li Tang. Predicting scholar potential: a deep learning model on social capital features. Scientometrics, 129(12):7851–7879, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05196-9.

Zeng:2020:MCW

[ZA20] Tong Zeng and Daniel E. Acuna. Modeling citation worthiness by using attention-based bidirectional long short-term memory networks and interpretable models. Scientometrics, 124(1):399–428, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03421-9.

Zhang:2020:CIR

[ZAF20] Qin Zhang, Juneman Abraham, and Hui-Zhen Fu. Collaboration and its influence on retraction based on retracted publications during 1978–2017. Scientometrics, 125(1): 213–232, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03636-w.

Zanardello:2023:MFI

[Zan23] Chiara Zanardello. Market forces in Italian academia today (and yesterday). Scientometrics, 128(1):651–698, January 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04579-0.

Zhao:2022:RMG

[ZAZS22] Xinyi Zhao, Samin Aref, Emilio Zagheni, and Guy Stecklov. Return migration of German-affiliated researchers: analyzing departure and return by gender, cohort, and discipline using Scopus bibliometric data 1996–2020. Scientometrics, 127(12):7707–7729, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04351-4.

Zhang:2024:HEE

[ZBG^{+}24] Na Zhang, Yanran Bi, Wenyue Gao, Jia Deng, and Erzi Tang. Heterogeneous effects of environmental regulations on green technology innovation: measurement by using patents and evidence from China’s prefecture-level cities. Scientometrics, 129(12):7667–7698, December 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05200-2.

Zhang:2024:SCE

[ZBKW24] Yujie Zhang, Rujiang Bai, Ling Kong, and Xiaoyue Wang. 2SCE-4SL: a 2-stage causality extraction framework for scientific literature. Scientometrics, 129(11):7175–7195, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04817-z.

Zbonikowski:2025:WIN

[Zbo25] Rafał Zbonikowski. What influences the number of citations of scientific articles? Study on colloid and interface science. Scientometrics, 130(5):2551–2576, May 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05311-4.

Zanotto:2021:AAF

[ZC21] Edgar D. Zanotto and Vinicius Carvalho. Article age- and field-normalized tools to evaluate scientific impact and momentum. Scientometrics, 126(4):2865–2883, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-021-03877-3.

Zhao:2023:FMB

[ZC23] Yueyang Zhao and Lei Cui. Fusion matrix–based text similarity measures for clustering of retrieval results. Scientometrics, 128(2):1163–1186, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04596-z.

Zhang:2024:ARI

[ZC24] Biao Zhang and Yunwei Chen. Automated recognition of innovative sentences in academic articles: semiautomatic annotation for cost reduction and SAO reconstruction for enhanced data. Scientometrics, 129(9):5403– 5432, September 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05114-z.

Zhou:2020:DAG

[ZCL20] Ping Zhou, Xiaojing Cai, and Xiaozan Lyu. An in-depth analysis of government funding and international collaboration in scientific research. Scientometrics, 125(2):1331– 1347, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03595-2.

Zhou:2020:DTI

[ZCLW20] Yanbo Zhou, Hongbing Cheng, Qu Li, and Weihong Wang. Diversity of temporal influence in popularity prediction of scientific publications. Scientometrics, 123(1):383–392, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03354-3.

Zhang:2023:RII

[ZCS^{+}23] Na Zhang, Lu Cheng, Chao Sun, Julie Callaert, and Bart van Looy. The role of inter- and intra-organisational networks in innovation: towards requisite variety. Scientometrics, 128(7):4117–4136, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04748-9.

Zhang:2024:MIO

[ZCS^{+}24] Lin Zhang, Zhe Cao, Yuanyuan Shang, Gunnar Sivertsen, and Ying Huang. Missing institutions in OpenAlex: possible reasons, implications, and solutions. Scientometrics, 129(10):5869–5891, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04923-y.

Zhang:2024:IGR

[ZCSK24] Lin Zhang, Zhe Cao, Gunnar Sivertsen, and Dmitry Kochetkov. The influence of geopolitics on research activity and international collaboration in science: the case of Russia. Scientometrics, 129(10):6007–6021, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-04984-7.

Zhang:2021:TED

[ZCW21] Yi Zhang, Xiaojing Cai, and Caroline S. Wagner. Topic evolution, disruption and resilience in early COVID-19 research. Scientometrics, 126(5):4225–4253, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03946-7.

Zhou:2020:FET

[ZDL^{+}20] Yuan Zhou, Fang Dong, Yufei Liu, Zhaofu Li, JunFei Du, and Li Zhang. Forecasting emerging technologies using data augmentation and deep learning. Scientometrics, 123(1):1–29, April 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03351-6;   http://link.springer.com/content/
pdf/10.1007/s11192-020-03351-6.pdf.

Zhou:2021:DLF

[ZDLR21] Yuan Zhou, Fang Dong, Yufei Liu, and Liang Ran. A deep learning framework to early identify emerging technologies in large-scale outlier patents: an empirical study of CNC machine tool. Scientometrics, 126(2):969–994, February 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03797-8.

Zhang:2022:PPC

[ZF22] Qin Zhang and Hui-Zhen Fu. Productivity patterns, collaboration and scientific careers of authors with retracted publications in clinical medicine. Scientometrics, 127(4): 1883–1901, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04252-y.

Zhitomirsky-Geffet:2020:TWP

[ZGBH20] Maayan Zhitomirsky-Geffet, Ofer Bergman, and Shir Hilel. Towards a wider perspective in the social sciences using a network of variables based on thousands of results. Scientometrics, 123(3):1385–1406, June 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03446-0.

Zhang:2023:EAC

[ZGM^{+}23] Yuanrong Zhang, Wei Guo, Jian Ma, Zhonglin Fu, Zhixing Chang, and Lei Wang. Evolution analysis of cross-domain collaborative research topic: a case study of cognitive-based product conceptual design. Scientometrics, 128(12):6695–6718, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04865-5.

Zhang:2021:ECI

[Zha21a] Guiyang Zhang. Employee co-invention network dynamics and firm exploratory innovation: the moderation of employee co-invention network centralization and knowledge-employee network equilibrium. Scientometrics, 126(9):7811–7836, September 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04089-5.

Zhang:2021:WIP

[Zha21b] Tenghao Zhang. Will the increase in publication volumes ”dilute” prestigious journals’ impact factors? A trend analysis of the FT50 journals. Scientometrics, 126 (1):863–869, January 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/s11192-020-
03736-7;   http://link.springer.com/content/pdf/
10.1007/s11192-020-03736-7.pdf.

Zhang:2023:ERL

[Zha23a] Lin Zhang. Editorial response letter to Abramo et al. Scientometrics, 2022. Scientometrics, 128(2):1463–1464, February 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04608-y.

See [MS21a, MS22a, AAA^{+}23].

Zhang:2023:ICD

[Zha23b] Yinxian Zhang. Impactful COVID-19 discoveries from China are neglected in the media. Scientometrics, 128(8): 4523–4539, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04779-2.

Zong:2023:COA

[ZHH23a] Qianjin Zong, Zhihong Huang, and Jiaru Huang. Can open access increase LIS research’s policy impact? Using regression analysis and causal inference. Scientometrics, 128(8): 4825–4854, August 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04750-1.

Zong:2023:DOS

[ZHH23b] Qianjin Zong, Zhihong Huang, and Jiaru Huang. Do open science badges work? Estimating the effects of open science badges on an article’s social media attention and research impacts. Scientometrics, 128(6):3627–3648, June 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-023-04720-7.

Zhou:2021:SCC

[Zho21] Yangping Zhou. Self-citation and citation of top journal publishers and their interpretation in the journal-discipline context. Scientometrics, 126(7):6013–6040, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03995-y.

Zhou:2024:EBI

[Zho24] Qingqing Zhou. Evaluating book impacts via integrating multi-source reviews. Scientometrics, 129(11):6931– 6946, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05174-1.

Zhu:2025:SIT

[Zhu25] Yunu Zhu. A study of the index time of early access articles. Scientometrics, 130(1):159–186, January 2025. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05188-9.

Zielinski:2022:IMS

[Zie22] Andrea Zielinski. Impact of model settings on the textbased Rao diversity index. Scientometrics, 127(12):7751– 7768, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04312-x.

Zayet:2021:ITR

[ZIS21] Tasnim M. A. Zayet, Maizatul Akmar Ismail, and Sheena Kaur Jaswant Singh. Investigating transportation research based on social media analysis: a systematic mapping review. Scientometrics, 126(8):6383–6421, August 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04046-2.

Zheng:2022:CDM

[ZJ22] Lukun Zheng and Yuhang Jiang. Combining dissimilarity measures for quantifying changes in research fields. Scientometrics, 127(7):3751–3765, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04415-5.

Zhou:2025:ISA

[ZJ25] Yuanyuan Zhou and Jiaojiao Ji. The impact of scientific articles on Chinese social media: examining its correlation with novelty and citations. Scientometrics, 130(8): 4573–4590, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05383-2.

ZhengWei:2022:RMA

[ZJY22] Huang ZhengWei, Min JinTao, and Tian Ye. Recommendation method for academic journal submission based on doc2vec and XGBoost. Scientometrics, 127(5):2381–2394, May 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04354-1.

Zhu:2020:CLL

[ZL20a] Junwen Zhu and Weishu Liu. Comparing like with like: China ranks first in SCI-indexed research articles since 2018. Scientometrics, 124(2):1691–1700, August 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03525-2;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03525-2.pdf.

Zhu:2020:TTD

[ZL20b] Junwen Zhu and Weishu Liu. A tale of two databases: the use of Web of Science and Scopus in academic papers. Scientometrics, 123(1):321–335, April 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03387-8.

Zang:2023:CEC

[ZL23a] Dongyu Zang and Chunli Liu. Correction: Exploring the clinical translation intensity of papers published by the world’s top scientists in basic medicine. Scientometrics, 128(4):2417–2418, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04688-4. See [ZL23b].

Zang:2023:ECT

[ZL23b] Dongyu Zang and Chunli Liu. Exploring the clinical translation intensity of papers published by the world’s top scientists in basic medicine. Scientometrics, 128(4): 2371–2416, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04634-4.

See correction [ZL23a].

Zhou:2023:ICK

[ZL23c] Hao Zhou and Jie Lin. Impacts of codified knowledge index on the allocation of overseas inventors by emerging countries: evidence from PCT patent activities in China. Scientometrics, 128(2):877–899, February 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04563-8.

Zeng:2021:EDD

[ZLL21] Bei Zeng, Haihua Lyu, and Jiang Li. Exploring the direction and diversity of interdisciplinary knowledge diffusion: a case study of professor Zeyuan Liu’s scientific publications. Scientometrics, 126(7):6253–6272, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03886-2.

Zhu:2022:ESR

[ZLL22] Weidong Zhu, Shaorong Li, and Zhimin Li. Evaluation of scientific research projects on the basis of evidential reasoning approach under the perspective of expert reliability. Scientometrics, 127(1):275–298, January 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04201-9.

Zhang:2023:CDI

[ZLM^{+}23] Yujie Zhang, Hongzhen Li, Jingyi Mao, Guoxiu He, Yunhan Yang, Zhuoren Jiang, and Yufeng Duan. COVID-19: a disruptive impact on the knowledge support of references. Scientometrics, 128(8):4791–4823, August 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04764-9.

Zhang:2021:DCB

[ZLY21] Li Zhang, Ming Liu, and Peng Yang. Discovering communities based on mention distance. Scientometrics, 126(3): 1945–1967, March 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-03863-9.

Zhang:2023:DKE

[ZLZ^{+}23] Tingting Zhang, Baozhen Lee, Qinghua Zhu, Xi Han, and Ke Chen. Document keyword extraction based on semantic hierarchical graph model. Scientometrics, 128(5): 2623–2647, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04677-7.

Zhang:2021:DOD

[ZM21] Liwei Zhang and Liang Ma. Does open data boost journal impact: evidence from Chinese economics. Scientometrics, 126(4):3393–3419, April 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
021-03897-z;   http://link.springer.com/content/
pdf/10.1007/s11192-021-03897-z.pdf.

Zhang:2023:OSD

[ZM23a] Liwei Zhang and Liang Ma. Is open science a doubleedged sword?: data sharing and the changing citation pattern of Chinese economics articles. Scientometrics, 128(5): 2803–2818, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04684-8.

Zhu:2023:GRS

[ZM23b] Chen Zhu and Kazuyuki Motohashi. Government R&D spending as a driving force of technology convergence: a case study of the Advanced Sequencing Technology Program. Scientometrics, 128(5):3035–3065, May 2023. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04682-w.

Zhang:2024:DOA

[ZM24] Liwei Zhang and Liang Ma. Different open access routes, varying societal impacts: evidence from the Royal Society biological journals. Scientometrics, 129(6):3407–3431, June 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05032-0.

Zafar:2023:IFS

[ZMA23] Lubna Zafar, Nayyer Masood, and Samreen Ayaz. Impact of field of study (FoS) on authors’ citation trend. Scientometrics, 128(4):2557–2576, April 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04660-2.

Zhu:2022:TCA

[ZMK22] Wenjing Zhu, Bohong Ma, and Lele Kang. Technology convergence among various technical fields: improvement of entropy estimation in patent analysis. Scientometrics, 127(12):7731–7750, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04557-6.

Zhang:2024:ENE

[ZMLZ24] Chengzhi Zhang, Philipp Mayr, Wei Lu, and Yi Zhang. An editorial note on extraction and evaluation of knowledge entities from scientific documents. Scientometrics, 129(11):7169–7174, November 2024. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL https://link.springer.com/article/10.1007/
s11192-024-05166-1.

Zhang:2024:DDE

[ZMP24] Chengzhi Zhang, Yi Mao, and Shuyu Peng. Data-driven evolution of library and information science research methods (1990–2022): a perspective based on fine-grained method entities. Scientometrics, 129(12):7889–7912, December 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05202-0.

Zerva:2020:CTS

[ZNNA20] Chrysoula Zerva, Minh-Quoc Nghiem, Nhung T. H. Nguyen, and Sophia Ananiadou. Cited text span identification for scientific summarisation using pre-trained encoders. Scientometrics, 125(3):3109–3137, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03455-z;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03455-z.pdf.

Zwanenburg:2022:TGC

[ZNW22] Sander Zwanenburg, Maryam Nakhoda, and Peter Whigham. Toward greater consistency and validity in measuring interdisciplinarity: a systematic and conceptual evaluation. Scientometrics, 127(12):7769–7788, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-022-04310-z.

Zhong:2020:GAB

[ZO20] Hao Zhong and Defang Ouyang. The gap analysis between Chinese pharmaceutical academia and industry from 2000 to 2018. Scientometrics, 122(2):1113–1128, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03313-7.

Zivkovic:2020:DSE

[ZP20] Zivan Zivković and Marija Panić. Development of science and education in the Western Balkan countries: competitiveness with the EU. Scientometrics, 124(3):2319– 2339, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03554-x.

Zhao:2021:CAR

[ZPH21] Zhenyue Zhao, Xuelian Pan, and Weina Hua. Comparative analysis of the research productivity, publication quality, and collaboration patterns of top ranked library and information science schools in China and the United States. Scientometrics, 126(2):931–950, February 2021.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03796-9.

Zheng:2021:SIR

[ZPZ21] Han Zheng, L. G. Pee, and Dan Zhang. Societal impact of research: a text mining study of impact types. Scientometrics, 126(9):7397–7417, September 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04096-6.

Zhang:2023:CCS

[ZQML23] Liyin Zhang, Yuchen Qian, Chao Ma, and Jiang Li. Continued collaboration shortens the transition period of scientists who move to another institution. Scientometrics, 128(3):1765–1784, March 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04617-x.

Zhu:2022:EAO

[ZQS22] Hengmin Zhu, Li Qian, and Chao Shen. Evolution analysis of online topics based on ‘word-topic’ coupling network. Scientometrics, 127(7):3767–3792, July 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04439-x.

Zhao:2020:DNI

[ZS20a] Dangzhi Zhao and Andreas Strotmann. Deep and narrow impact: introducing location filtered citation counting. Scientometrics, 122(1):503–517, January 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03280-z.

Zhao:2020:TPV

[ZS20b] Dangzhi Zhao and Andreas Strotmann. Telescopic and panoramic views of library and information science research 2011–2018: a comparison of four weighting schemes for author co-citation analysis. Scientometrics, 124(1):

255–270, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03462-0.

Zeng:2021:TDJ

[ZS21] Ziqiang Zeng and Lantian Shi. A two-dimensional journal classification method based on output and input factors: perspectives from citation and authorship related indicators. Scientometrics, 126(5):3929–3964, May 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-03924-z.

Zhang:2024:AJC

[ZS24] Jiandong Zhang and Zhesi Shen. Analyzing journal category assignment using a paper-level classification system: multidisciplinary sciences journals. Scientometrics, 129(10):5963–5978, October 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-023-04913-0.

Zhang:2021:GDA

[ZSG21] Lin Zhang, Gunnar Sivertsen, and Wolfgang Glänzel. Gender differences in the aims and impacts of research. Scientometrics, 126(11):8861–8886, November 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04171-y.

Zhang:2024:MES

[ZSL24] Junsheng Zhang, Xiaoping Sun, and Zhihui Liu. Measuring the evolving stage of temporal distribution of research topic keyword in scientific literature by research heat curve. Scientometrics, 129(11):7287–7328, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-04937-0.

Zhang:2022:GDA

[ZSS22] Lin Zhang, Yuanyuan Shang, and Gunnar Sivertsen. Gender differences among active reviewers: an investigation based on Publons. Scientometrics, 127(1):145–179, January 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04209-1.

Zhang:2022:CPL

[ZSSH22] Lin Zhang, Beibei Sun, Fei Shu, and Ying Huang. Comparing paper level classifications across different methods and systems: an investigation of Nature publications. Scientometrics, 127(12):7633–7651, December 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04352-3.

Zhang:2024:PGC

[ZSX24] Xiaojuan Zhang, Shuqi Song, and Yuping Xiong. Personalized global citation recommendation with diversification awareness. Scientometrics, 129(7):3625–3657, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-024-05057-5.

Zhang:2023:NSG

[ZT23] Chengzhi Zhang and Liang Tian. Non-synchronism in global usage of research methods in library and information science from 1990 to 2019. Scientometrics, 128(7): 3981–4006, July 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04740-3.

Zhang:2024:SPR

[ZTY^{+}24] Lin Zhang, Ziyi Tu, Yifei Yu, Yuanyuan Shang, and Ying Huang. Spotting potential reviewers for interdisciplinary research: insights on active reviewers from Publons. Scientometrics, 129(9):5533–5556, September 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05129-6.

Zhang:2021:MAE

[ZW21a] Fang Zhang and Shengli Wu. Measuring academic entities’ impact by content-based citation analysis in a heterogeneous academic network. Scientometrics, 126(8):

7197–7222, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04063-1.

Zhang:2021:WAP

[ZW21b] Liwei Zhang and Jue Wang. What affects publications’ popularity on Twitter? Scientometrics, 126(11):9185– 9198, November 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04152-1.

Zhang:2021:EKH

[ZW21c] Shaopeng Zhang and Xiaohong Wang. Effect of knowledge hiding on knowledge innovative behavior of innovative team members. Scientometrics, 126(8):6423–6442, August 2021. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-021-04047-1.

Zhang:2024:PCI

[ZW24a] Fang Zhang and Shengli Wu. Predicting citation impact of academic papers across research areas using multiple models and early citations. Scientometrics, 129(7):4137–4166, July 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-024-05086-0.

Zhou:2024:MHT

[ZW24b] Jianlin Zhou and Jinshan Wu. Measuring hotness transfer of individual papers based on citation relationship. Scientometrics, 129(11):6659–6674, November 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05140-x.

Zagovora:2025:SCF

[ZW25] Olga Zagovora and Katrin Weller. Science communicators, flat-earthers, or fitness coaches: who is citing scientific publications in YouTube video descriptions? Scientometrics, 130(1):205–235, January 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05204-y.

Zhang:2022:AAP

[ZWC22] Yu Zhang, Min Wang, and Elizabeth Chang. Analysing academic paper ranking algorithms using test data and benchmarks: an investigation. Scientometrics, 127(7): 4045–4074, July 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04429-z.

Zhang:2020:OIE

[ZWD20] Gupeng Zhang, Xiao Wang, and Hongbo Duan. Obscure but important: examining the indirect effects of alliance networks in exploratory and exploitative innovation paradigms. Scientometrics, 124(3):1745–1764, September 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03586-3.

Zhang:2022:SOA

[ZWHS22] Lin Zhang, Yahui Wei, Ying Huang, and Gunnar Sivertsen. Should open access lead to closed research? The trends towards paying to perform research. Scientometrics, 127(12):7653–7679, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04407-5.

Zhao:2023:RPC

[ZWL23] Jingyi Zhao, Chunli Wei, and Jiang Li. Is the research performance of Chinese returnees better than that of their local counterparts? Scientometrics, 128(5):3091–3105, May 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-023-04693-7.

Zhang:2025:ITC

[ZWL25a] Mingze Zhang, Lili Wang, and Zexia Li. The impact of team compositions on disruptive and novel research in large-scale research infrastructures. Scientometrics, 130 (5):2963–2986, May 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05319-w.

Zhao:2025:EJP

[ZWL^{+}25b] Xinhang Zhao, Xuefeng Wang, Yuqin Liu, Hongshu Chen, and Rui Guo. Evolution of journal preference based on topic focus: a case study in the technology innovation management field. Scientometrics, 130(4):2137–2166, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL https://link.springer.com/
article/10.1007/s11192-025-05279-1.

Zhang:2020:KFT

[ZWSC20] Yu Zhang, Min Wang, Morteza Saberi, and Elizabeth Chang. Knowledge fusion through academic articles: a survey of definitions, techniques, applications and challenges. Scientometrics, 125(3):2637–2666, December 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03683-3.

Zhang:2021:OAU

[ZWW21] Guangyao Zhang, Yuqi Wang, and Xianwen Wang. The open access usage advantage: a temporal and spatial analysis. Scientometrics, 126(7):6187–6199, July 2021. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-020-03836-4.

Zhang:2022:CEA

[ZWX22] Xi Zhang, Xianhai Wang, and Hui Xiong. Correction to: An effectiveness analysis of altmetrics indices for different levels of artificial intelligence publications. Scientometrics, 127(3):1657, March 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04146-z. See [ZWZ^{+}19].

Zhang:2019:EAA

[ZWZ^{+}19] Xi Zhang, Xianhai Wang, Hongke Zhao, Patricia Ordóñez de Pablos, Yongqiang Sun, and Hui Xiong. An effectiveness analysis of altmetrics indices for different levels of artificial intelligence publications. Scientometrics, 119(3): 1311–1344, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03088-x.

Zhou:2022:PIP

[ZWZ22] Yuhao Zhou, Ruijie Wang, and An Zeng. Predicting the impact and publication date of individual scientists’ future papers. Scientometrics, 127(4):1867–1882, April 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04286-w.

Zhang:2025:AAI

[ZX25] Xinyuan Zhang and Qing Xie. Acknowledgment analysis: insight into biopharmaceutical company-funded papers and patents. Scientometrics, 130(2):999–1044, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05238-w.

Zong:2020:DOP

[ZXL20] Qianjin Zong, Yafen Xie, and Jiechun Liang. Does open peer review improve citation count? Evidence from a propensity score matching analysis of PeerJ. Scientometrics, 125(1):607–623, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03545-y.

Zhou:2022:IDE

[ZXL22] Yanbo Zhou, Xin-Li Xu, and Qu Li. The influence of disruption on evaluating the scientific significance of papers. Scientometrics, 127(10):5931–5945, October 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04505-4.

Zhang:2024:HPD

[ZXL^{+}24] Yang Zhang, Yinghua Xie, Longfei Li, Yian Liang, and Houqiang Yu. How is public discussion as reflected in WeChat articles different from scholarly research in China? An empirical study of metaverse. Scientometrics, 129(1): 473–495, January 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04892-2.

Zhang:2022:MEP

[ZXS22] Xinyuan Zhang, Qing Xie, and Min Song. Mining the evolutionary process of knowledge through multiple relationships between keywords. Scientometrics, 127(4): 2023–2053, April 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04272-2.

Zhang:2020:IAS

[ZY20a] Helena H. Zhang and Fred Y. Ye. Identifying ‘associatedsleeping-beauties’ in ‘swan-groups’ based on small qualified datasets of physics and economics. Scientometrics, 122(3): 1525–1537, March 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-020-03359-y.

Zhang:2020:EDT

[ZY20b] Jinzhu Zhang and Wenqian Yu. Early detection of technology opportunity based on analogy design and phrase semantic representation. Scientometrics, 125(1):551–576, October 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03641-z.

Zhang:2020:WBA

[ZY20c] Yajie Zhang and Qiang Yu. What is the best article publishing strategy for early career scientists? Scientometrics, 122(1):397–408, January 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03297-4.

Zhang:2022:EAB

[ZY22] Ruhao Zhang and Junpeng Yuan. Enhanced author bibliographic coupling analysis using semantic and syntactic citation information. Scientometrics, 127(12):7681– 7706, December 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-022-04333-6.

Zhuang:2020:IGL

[ZYL20] Liang Zhuang, Chao Ye, and Scott N. Lieske. Intertwining globality and locality: bibliometric analysis based on the top geography annual conferences in America and China. Scientometrics, 122(2):1075–1096, February 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03325-3.

Zhou:2025:IPC

[ZYL25] Lewei Zhou, Mingliang Yue, and Chundong Li. The impact of patent citation on the citation performance of academic papers. Scientometrics, 130(8):4187–4219, August 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.springer.

com/article/10.1007/s11192-025-05400-4.

Zhu:2020:NMB

[ZYPC20] Yongjun Zhu, Erjia Yan, Silvio Peroni, and Chao Che. Nine million book items and eleven million citations: a study of book-based scholarly communication using OpenCitations. Scientometrics, 122(2):1097–1112, February 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03311-9.

Zhang:2025:TEF

[ZYWA25] Zhengang Zhang, Chuanming Yu, Jingnan Wan, and Lu An. A temporal evolution and fine-grained information aggregation model for citation count prediction. Scientometrics, 130(4):2069–2091, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-025-05294-2.

Zhao:2023:IDF

[ZYZW23] Yiming Zhao, Jiaying Yin, Jin Zhang, and Linrong Wu. Identifying the driving factors of word co-occurrence: a perspective of semantic relations. Scientometrics, 128(12): 6471–6494, December 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https:

//link.springer.com/article/10.1007/s11192-023-
04851-x.

Zhang:2025:EKE

[ZYZZ25] Chengzhi Zhang, Xinyi Yan, Lei Zhao, and Yingyi Zhang. Enhancing keyphrase extraction from academic articles using section structure information. Scientometrics, 130(4): 2311–2343, April 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05286-2.

Zhou:2020:EWI

[ZZ20a] Qingqing Zhou and Chengzhi Zhang. Evaluating wider impacts of books via fine-grained mining on citation literatures. Scientometrics, 125(3):1923–1948, December 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-020-03676-2.

Zhu:2020:CWA

[ZZ20b] Xiang Zhu and Yunqiu Zhang. Co-word analysis method based on meta-path of subject knowledge network. Scientometrics, 123(2):753–766, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
020-03400-0.

Zhang:2023:MST

[ZZ23] Lingche Zhang and Qiuju Zhang. Mapping the scientific and technological landscape: an analysis of Nobel Prizeproducing institutions. Scientometrics, 128(11):6129– 6145, November 2023. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-023-04831-1.

Zhang:2024:EPM

[ZZ24] Yingyi Zhang and Chengzhi Zhang. Extracting problem and method sentence from scientific papers: a contextenhanced transformer using formulaic expression desensitization. Scientometrics, 129(6):3433–3468, June 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-024-05048-6.

Zhao:2025:RNM

[ZZ25] Yi Zhao and Chengzhi Zhang. A review on the novelty measurements of academic papers. Scientometrics, 130(2): 727–753, February 2025. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-025-05234-0.

Zhu:2022:DEE

[ZZC22] Lin Zhu, Junjie Zhang, and Scott W. Cunningham. Domain expertise extraction for finding rising stars. Scientometrics, 127(9):5475–5495, September 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04492-6.

Zhang:2020:DNF

[ZZL^{+}20] Lin Zhang, Wenjing Zhao, Jianhua Liu, Gunnar Sivertsen, and Ying Huang. Do national funding organizations properly address the diseases with the highest burden?: Observations from China and the UK. Scientometrics, 125(2):1733–1761, November 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03572-9.

Zhang:2022:EAI

[ZZMS22] Yi Zhang, Chengzhi Zhang, Philipp Mayr, and Arho Suominen. An editorial of “AI + informetrics”: multidisciplinary interactions in the era of big data. Scientometrics, 127(11):6503–6507, November 2022. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04561-w.

Zhang:2020:HSR

[ZZS^{+}20] Lin Zhang, Wenjing Zhao, Beibei Sun, Ying Huang, and Wolfgang Glänzel. How scientific research reacts to international public health emergencies: a global analysis of response patterns. Scientometrics, 124(1):747–773, July 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-020-03531-4;   http:
//link.springer.com/content/pdf/10.1007/s11192-
020-03531-4.pdf.

Zhang:2022:CTE

[ZZW^{+}22a] Yang Zhang, Rongying Zhao, Yufei Wang, Haihua Chen, Adnan Mahmood, Munazza Zaib, Wei Emma Zhang, and Quan Z. Sheng. Correction to: Towards employing native information in citation function classification.

Scientometrics, 127(11):6579, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-022-04451-1. See [ZZW^{+}22a].

Zhang:2022:TEN

[ZZW^{+}22b] Yang Zhang, Rongying Zhao, Yufei Wang, Haihua Chen, Adnan Mahmood, Munazza Zaib, Wei Emma Zhang, and Quan Z. Sheng. Towards employing native information in citation function classification. Scientometrics, 127(11):6557–6577, November 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11192-021-04242-0. See [?]Zhang:2022:CTE.

Zhang:2024:MDI

[ZZW24] Ziyan Zhang, Junyan Zhang, and Pushi Wang. Measurement of disruptive innovation and its validity based on improved disruption index. Scientometrics, 129(11):6477– 6531, November 2024. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-024-05134-9.

Zhang:2022:EKE

[ZZZ22] Chengzhi Zhang, Lei Zhao, and Yingyi Zhang. Enhancing keyphrase extraction from academic articles with their reference information. Scientometrics, 127(2):703– 731, February 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL https://link.

springer.com/article/10.1007/s11192-021-04230-4.

Zheng:2020:EVJ

[ZZZZ20] Minxian Zheng, Kuangji Zhao, Shikui Zhao, and Yantong Zhang. Effecting variables of journal’s ranking in forestry field. Scientometrics, 125(1):135–151, October 2020. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-020-03629-9.